A Complete Bibliography of Publications in Advances in Applied Probability

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

13 October 2023 Version 1.23

Title word cross-reference

−1 [Fre82]. 0 [JS14]. 1 [Hol11, JD89, Qui72]. [0, 1]^{d} [AR97a, AR97b]. ^{X} [EW14]. ^{Y} [BC89b]. _{2} [Hok80]. _{k} [Tru75a, Tru75b]. _{m} [Coh72, De 83a]. _{n} [Coh72]. α [AP98, HR16, JKR17, Num76]. AR(1) [OC10, PS10]. B [JVZ08]. β [GKT22]. m [Jan23]. U [Jan23]. · [KW90b, TW87b]. χ^{2} [AA85, AA86, AA88b, Cao99, Sha78a, Sha78b, Wor94, Wor01, Lin80]. cμ [BVW85]. d [Gne99, Gne00, JK94, Sch16, Tho07, VMM19]. D_{q} [Ran80]. dv_{T} = V_{T} − du_{T} + dl_{T} [Beh11]. E_{n} [Sto98]. ϵ [Ker78]. F [Cao99, CP90, Wor94, BD73, San98, TT94]. f(X_{n}) [BD73]. F_{ST} [WH03]. G [GC02]. H = λG [Sig91, Whi92a]. I [TWN12]. ∞ [EW14, Fal84b, Gly82, Gol16, GS95, Hal85b, KS93, Kuc83, PZ16, Ram84, RW91, Sat85, SK08, Sur08, Whi82a]. J [BV97]. K [Wei78, BBSW05, BBSW09, CP90, DFZ19, FRW12, Pap87, Pap88, PP15, ZM16]. L [BGS89, ENM89]. L^{2} [JY04, PD22]. Λ [Gri14, KW18, HMS03]. m [Kwe75, Pap88, Tam11]. H^{d} [Cza23]. Z [Ath18, GNR17, Hir17, Val96]. Z^{d} [Soo10]. G(\, {√_{⟩|}}) [LR12]. O(K^{∋}) [AM22]. μ [Ano93g, NP93, Pol88]. μc

[HK93, BDM85, HK92]. N [BFL04, BLT03, Fed78, CP90, DFZ19, GV76, HS77, Mat99, Pap87, Pap88, Pic74b]. n^{−1/α}S_{n} [Mij78]. p [Mat99, Van78c]. π^{∗} [Met80]. Q [Pol91, BOS23, Hal86]. R [BFL04, CK07]. R^{1} [KT01]. R^{2} [KN16]. R^{d} [CQ01, DMR88, FM95, GS77b, Møl89, Per97]. R^{n} [Kem73, JK94]. ρ [Qui72]. S [BR83, DLS00]. σ [MN14]. T [KP86, Cao99, Wor94]. X_{n} = ρX_{n−1} + B_{n} [Wol80b]. Y_{n+1} = A_{n}Y_{n} + B_{n} [Bra86]. Z [BMZ14, DM06].

-algebra [MN14]. -Classification [HMS03]. -coalescents [KW18]. -Content [Mat99]. -Convex [DLS00]. -cycle [Thé22]. -Delaunay [GKT22]. -Dependent [Hal86, Jan23]. -Dimensional [Gne99, Gne00, JK94, Tho07, Sch16]. -Ergodic [TT94, PD22]. -Fractions [BR83]. -Function [BV97]. -Functions [Van78c]. -homogeneous [JS14]. -Invariance [Pol88]. -Invariant [Ano93g, NP93]. -Mixing [AP98]. -Nearest-Neighbour [BBSW05, BBSW09]. -Nearest-Neighbours [FRW12]. -Network [GC02]. -Optimal [Ker78]. -out [PP15]. -out-of- [CP90, DFZ19, Pap87, Pap88, BFL04]. -out-of-2 [JD89]. -Parallelotope [Mat99]. -Person [Fed78]. -Process [Lin80]. -Processes [Met80, Pol91]. -Recurrent [Num76]. -root [JKR17]. -Rule [BDM85, HK92]. -scale [BOS23]. -Scan [CK07]. -Segments [TWN12]. -Space [Mat99]. -Stable [Hol11, HR16]. -statistics [Jan23]. -Superadditive [BGS89, ENM89]. -Tower [BLT03]. -trees [ZM16]. -Unit [GV76]. -Valued [Ran80]. -within- [Pap88].

/1 [BC89b, Coh72]. /G/ [EW14]. /G/1 [Tru75a, Tru75b]. /GI/1/N [KW90b]. /K [Coh72]. /m [Hok80]. /M/1 [TW87b]. /s [De 83a].

0 [Naw91, O’B80].

1 [AW88, Ali90, AFGR06, Asm82, BW90a, Ber78c, Box78, BK01, BC89b, Coh72, Coh74, Coh77a, CH80, DN78, DT80, Dos78, Fer93, Fri86, Fri87, GHT96, GHT97, Gal79b, Gly94, Gri94, HP96, Hau78, HK89a, Igl71, JS83, KS83b, KD08, Köl81, Kyp71, LPRS93, LZ95, LZ05a, LZ05b, MA95a, Min80, Neu76a, Neu77b, Neu86c, Num79, O’B80, Ott77a, Ott77b, Ott87a, PR06, Par87, Ram80, Ros73, Ros76a, Sch84, Sch90, Sen89a, Sen89b, SG92, SP80, ST90, Sys88, TH90, TV10, TV78, Tru75a, Tru75b, TW87b, Van78b, WW03, YC96]. 1-Type [MZ04]. 1-Valued [O’B80].

2 [Hok79, JD89]. 2D [SK96]. 2M [Imh92]. 2M-X [Imh92].

Abbreviation [Rog89]. Abdel [Ano95e]. Abdel-Hameed [Ano95e]. Abel [LP99]. Abel-Gontcharoff [LP99]. Abelian [LP96]. Abelian-Type [LP96].

Abnormal [WB71]. above [Imh86]. Absence [DAG95, Cur81]. Absolute [Cai07, Haj76, ZYY11]. absorbed [CV18]. Absorbing [GVZ13, Pak95, PR90, FS16b, GG17]. Absorption [BC81b, Ebe80, Hüb78, KZ20, Kes78, VY94, KW18]. Accelerating [TR19]. Acceptance [Rav77]. Access [FFH86, MY80, Szp94]. Accessible [Cow89b]. According [MRV02]. Accumulation [Hor77]. Achievable [BM98]. Acknowledgement [Ano88a, Ken79]. Acoustic [Sut79]. Acquisition [ÅH99]. action [HG15, ZCJW19, ZCLJW23]. Actions [Hüb78]. Active [Bro99]. Actual [Coh72]. Actuarial [SH89, LY15a]. Acyclic [BW89]. Ad [DGP10]. Ad-Hoc [DGP10]. Adaptable [BBR09]. Adaptation [YZ78]. adapted [MMP22]. Adaptive [Aas82, EF80, Led79]. Addendum [Gne00, MS04, WW91]. Addition [Mas79]. Additional [KZ94]. Additive [AS73, AK00, Bro77, BRT82, DGR02, Dur79, HL06, JLW05a, JLW05b, Jac05, Lev78, PZ86, AB15, BFP23, CKLP20, Fox16, Fro13]. Additive-Increase [DGR02]. Addresses [BÇNP84]. Adjustment [PGE96]. Admission [BPS15, HS89, ÖB05, JvL15]. Adsorption [PS09, PS11, Van06, Mey18]. Advanced [ØSZ11, MT16]. Advances [Gan69, Gan08, Ano14a, Ano14b, Ano15a, Ano15b, Ano13a, Ano13b]. Advancing [Dan77]. Advantageous [CW75]. Advertising [AW79]. Affected [Ste80a]. affects [CL14]. Affine [GKT21, KP10, BDLP14, Buc16, FG81, FJKR20, JKR17]. affines [FG81]. after [BSS17, Cow96, Cow97, Ste80a]. Age [Bar77, Big76, CW75, Coh77b, Coh77c, Eth90a, Fei76a, Fil72, Fil74, GH80, Kim83, Ney74a, O’C95, PT81, Ric11, RH89, Sav75b, Sim81, SS88b, Tav78, Tav79, Wau74, CH14]. Age- [Kim83]. Age-Dependent [Big76, GH80, Ney74a, Ric11, RH89, Sav75b, SS88b, CH14]. Age-Ordered [Eth90a]. Age-Structured [CW75]. Ageing [BS99, Baz74, BOR02, HM20]. Agents [Elv71, Gar71, Hir17]. Ages [Ano86f, DT86, Eth90b]. Aggregate [AEM94, LTW10]. Aggregated [BS88, Bal90, BMY91, BMTY97, GL79, JH92b, Jir13]. Aggregation [FS09a, HL02, MR09, PS10, RW85, YZB03, DS23, PSS20]. Aggregative [FZ96]. AIDS [BO93]. Air [Gra82, Ste84]. Air-Pollutant [Gra82]. Aléatoires [Mat72, Gal84]. Algebra [Hol78, Mai97, Ros79, Str76, MN14]. Algebraic [Ahm80a, BR76, LT73, Mas84b, Odo80, PL99, Wil71]. Algebras [Hol80, WB81]. Algorithm [BFP98, Bro95a, BL09b, CLW95, Dom07, GR96, IM02, KS78, MS01b, MS04, Oso10, Pfe85a, Pfe85b, RT07, Yin00, AM22, BMS18, BGZ16, CKL14, Cer17, RT23, TR19]. Algorithmic [Alf03, Neu75, Mas23]. Algorithms [Ano98e, AK06, BP94, Cer98, DGR02, EF80, FT78, FP96, HMV06, JY04, MSS01, Moo79, Neu76a, Neu77a, RR95, WCK96, KM14a, PHL16, SWL14]. Aligned [KO87]. alignment [GLM22]. Alignments [KK80]. all-time [HR16]. Allee [Lan13a]. Allele [Gri79b, Joy91, Kel77]. Alleles [Ano86f, BKS12b, DT86, EK81, Eth90a, Eth90b, Gri79a, Gri82, GP88, Hil76b, Wat74, Wat79b, Wat92]. Allelic [FM09, Hin94a, Sat78, Shi82, ALS18a].

Allocation [BGRSC10, BSC11, Bat80, Ben91, CL87, Chi80, ENS93, Epl86, FGRT17, Git69, MP08, NG77, SS92, Whi96, CR19, PHL16]. Allocations [BENP92, SS20]. Allowing [BP79, JH92b, Pak79, Pak89a, Pak10]. Almost [DMR88, DM03, Gat00, GM96, KT01, Lin74a, Nav03, Ott87a, Tjø76b, TL20, Wei95b, MS17]. Almost-Phase-Type [Ott87a]. Almost-Sure [Gat00, GM96, MS17]. Along [VGS10]. alphabets [Cas16]. alternate [Tam16b]. Alternating [Di 01, LOS03, MY01b, Sti72, Yeo78, AS14, PZ16]. Alternating-Priority [Sti72]. Alternative [CN86, CT08, Fac09, GF87, LT00, PP02]. Always [Pak95]. ambit [BNBV14]. American [EP15, Won08]. Ammunition [SSY91]. among [Ano86e, BL01, Fra85, Hil76b, KO87, KD92]. Amount [Ger77, Hal87, BWZ19]. Amplification [LJJ04, SW97]. Amplitude [Lin72, Ryc87]. Analogies [GSS01]. Analogue [Wol80b]. analogues [FT18]. Analogy [PL89]. Analyses [CFD85]. Analysis [ÅH99, Alf03, AANQ04, Ano12, AFGR06, AR92, AT81, BB93a, BKZ95, BFL04, Bar74b, Bar74c, BGZ95, BM03, BGZ16, BT08, BC89b, BC04, Cao90, Cao95, CF78, CNP91, Cow06, Cra97, CL00, DMC00, DGR02, EM11, FM88, FFH86, GHT88, GHT96, GM08, Gan75, GV76, Haj82, HK90, HM81, Hin94b, Hua97, HZ06, KM10a, KD08, Kog89, KZ92, KZ94, KKSS08, Lat88, LPRS93, LG05, LL98, MK01, MA95a, MS87, Mat98, Mit87, MH97, MZ98, Møl95b, MKS^{+}78, Mon95, Nar87, Nea06, Now88, Par77, Pfe85a, PL90, PC96, Pru78, Ram80, RS88, Sch71, Sch74a, SW93, SD98, Ste76a, SE91, Tak91b, TW87a, Tho82, TS99, VL98, VL10, WG83, Wei98, WDH98, Yas85, BMS18, BL22, CMT21, DFZ19, FR14, Fly18, HHBV20, HLLP16, HL14, dlI22]. Analytic [DSSW11, DHP12, GT08, Hwa99, Tak75a, Tak75b, Wen89]. Analytical [De 77, Rav83]. Analyticity [BH00, Zaz92]. Analyze [BM96]. Analyzing [Wei86]. Anatomy [Hin94a, WDH98]. Ancestor [Möh04, O’C95]. Ancestors [Cha99a, Cha99b, DWH^{+}99]. Ancestral [GB03, Möh00, LY15b]. Angiosperm [Heu80]. Animal [BC76, McN90]. Anisotropic [Ben95, HS11, Mec95, Mec98b, Sch94]. Annealing [HS91, MSS01, MRSV86, Van94]. Annihilating [Hol83, NS93, Sud99]. Annihilation [Jac74a]. Anomalies [Wil79]. Anticipative [EX11, PK96, PR94]. Antithetic [Geo76]. Any [RS07, SC83, Tam11]. Apollonian [KKV16, ZM16]. Apparent [JH92a]. Appearance [Sch90]. Application [AANQ04, AK95, Ano07e, Bal79, BD87, BB07b, BR83, BW97, Cao87, Cli07, Cli11, Dan89, Dom07, EL05, Gau02, HJ00, HS92, HZ06, KM00, Kha80, KS04a, Le01, MD79, Mih80, NM79, Num79, Off74, PTW10, Pap77, Phi78, RL77, Sen89b, SD98, Sze76, Tan96, TS99, Twe82, VK78, Van94, VJ08, Wau81, Wor95, Wor01, AD21, Bau14, BSHJ17, CDJR22, HL14, HNGS15, LTT20, Oça23, Yca84]. Applications [AK05, AE08, AH97, Ano03e, Ano04f, Ano13b, Ano14b, Ano15b, Asm82, AK00, AS96, BTZ00, BB01, Bal78, BMY94, BH89a, BGZ95, BG07, BY09, BT08, BFK92, BV90, BV02, CFW09, Cha92a, Coh88, Cov09, CDH87, DFT77b, Die06, Eco05, ENPS78, EGV79, EN80, EN88, FVK08, GPD00, GT08, GS77a, Haj82, HV83, JM86, Kas80, KK90, KK92, Ken77c, KS75, Kur78, Lam93, LP93, LS89, LS13, Lin80, Liu98, Mar77, Mat73, Mel79b, MSS01, MW96a, Nec78, Neu78, Now88, PT85a, Pic74a, Pic74b, Pri77, Pru78, Ris91, Rob76b, Sav75a, Sch77, Sen79a, Sen79b, SS88a, Shi93, SB09, Sti72, SK85, Tak91a, VP88, Van88a, VV01, Van74, Was78, Wil90, Wit79, YZ01, YZ02, ZCWS96, CKL14, Cha14, CYG14, Gal84, HA20, Jan23, KP22, LY15a]. applications [LSH16, Liu97, SMAL17]. Applied [Ano13a, Ano13b, Ano14a, Ano14b, Ano15a, Ano15b, DNP10, Gal79a, Gan78, Gan08, Igl73a, Jag18, JVZ08, Le 74, LSV18, Ano89e]. Applying [Lan82]. Appointment [Had76]. Appreciation [Jen03]. Approach [AWZ93, AFGR06, AG81, AB94, Ath79, Bal86, BGRSC10, Bäu02, BDM99, Bra78, Can74b, Can75b, CG80, Cap76, Cek97, CM00, CPR01, CJ11, De 76, De 80, DNPR01, DLS94, DS76a, DHP12, EF81, FV95, GM96, GLS01, Hak82, HCD94, Hui03a, Hwa99, Joy91, KKSS90, Kim83, Kit71, KSC09, Leó76, LT73, Mac75a, McG82, Miy04, NG77, Nic82, Odo80, OT02, PP02, Par87, PL96, QG11, RS86, RS88, Sch84, Sco73, Shi80, Tau75, TV75, Ven82, Wil85, Yas85, ACH21, HSZ18, LLX18, Miy17]. Approaches [Coh85, CSH^{+}80, DMC00, DK80, KS80]. Approximate [BKS12b, BS03b, Boy78, FJ09, LS89, RS12, Wah76]. Approximating [CP88, Hey74a, SY81]. Approximation [Ahm81, ARM99, Arr03, BE83, BU98, BM00, BNX02, BX06, BS03c, Cha09b, CPR01, Coh82b, CDH87, Dan74a, DN02, Duf04, Ell87, Epl86, GL04b, Har78, HK22, Hua90, HS14, Hwa99, KA81, KD85, Köl81, Küf94, LR80, Mij80, Nar91a, Nar91b, PP02, RR90a, RW03, Rei81, Rol87, RS07, Sel96, SÖ11, SZ11, Sto78, Szc90, VK78, WCK96, Wol80a, Yun86, Aaz13, BE21, BGZ16, CMT21, Chi14, CGR18, DHL17, GRR17, GLL20, MLZ21, MS15, MS16, SWL14, XY15]. Approximations [Als88, AG05, BL10, Bor09, BP82, CW03, Cek03, Coh82a, CDL99, CY99, DR92, DTV84, DLS00, DW96, EJMT85, EN80, EN88, Fad77, FQVJ71, Fuh97, GNR86, GNR87, God91, GS93, HC08, Hog86, Hua02, IL73, IL74, Jen88, KW90a, Kus94, LYA07, Loa91, NR21, PP07, PV92, Rao06, Roo93, Roo78, Ryc87, ST96, SC83, Sha97, Sie79, Smi83, TVF81, Van78d, Wal81b, Wan97, Wes76, WW03, Yuh82, Zha97, APZ21, BS15, BF19, CCHH18, GG17, LL18, NLM18]. Approximative [FR11]. APR [Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano17h, Ano17g, Ano18b, Ano18a, Ano18d, Ano18c, Ano18f, Ano18e, Ano18h, Ano18g, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano20h, Ano20g, Ano21b, Ano21a, Ano21d, Ano21c, Ano21f, Ano21e, Ano21h, Ano21g, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e]. Arbitrage [Pla02, PT23]. Arbitrarily [Gre89b, Swi75]. Arbitrary [BS12b, Can74a, GC20]. ARCH [Rud98, Sur08]. ARCH/GARCH [Rud98]. Archimedean [LN11]. Area [Bal86, Gri75b, HS98, JP00a, PLRC08, Rat98, Sch99a, PW18]. Areas [HQ02]. Arises [Kne00]. Arising [BP83, Can74b, Can75b, Cow10, FFH86, JH92b, Möh98b, RS95, Kar21]. Arm [BF80a]. ARMA [Ano86e, BR93, DR89, DP84, Fra85, Ger77, Hua90, KJH18, McK88]. Armed [Agr95, BF80a, BF80b, Gla83, ZCJW19, ZCLJW23]. Arms [BF80b]. Arnold [BS83]. Arrêt [PS83]. Arranged [Cow89b]. Arrays [SÖ11]. Arrival [FS89a, Fal88, Gre78, Had76, HK93, LMHN90, TV78, YZ78]. Arrivals [ARW11, CCP91, JS80, KS06, Lat88, MT97, ÖB05, Sha97, Van88b, BBRW14, LP21]. Arriving [CR04, LC08]. Arterial [SK96]. Artificial [Hil75, Whi91]. Ascertainment [Sim81, Ste79, Ste81]. Asian [JV18, PZ17, Sch06b]. Aspects [BW93]. assessment [Gof19]. Asset [DE12, MS11a]. Assignment [Rob92]. Associate [Asm82]. Associated [HM96, HM99, Kai78, KP86, KLRM05, KM95, Li98, MS87, RS09, SB09, Sze76, Woo80, Xu11, FGZ17, Gal84, HS14]. Association [LM90]. Associative [Wil79]. associées [Gal84]. Assortative [SV75, Vet78]. assortativity [vdHvdHLS20]. Assumptions [Cli11, Sch73, Ste81]. Astrophysics [Wor95]. Asymmetric [CCH00, Lin10, Qia90, Tay77, GML18]. Asymptotic [AMNW11, And99, ACF09, AK78, BN10, BN12, Bar75, BR13, BPP21, Bas79, BGZ95, BD96, BDK^{+}18, BS12b, Big78, BD74b, BD74a, BD75, Bob08, BV06, Cek97, Cer98, CTR00, Coh73, Coh76, Coh77d, Cra74a, Dav73b, DT80, Dei03, DM00, DEP91, DS00, Ebe80, FT12, FX13, Fin80, FE06, FL01, Fuh07, Gar99, Ger84a, GC20, GNR90, GLS11, GZ93, GR96, GB17, Her73, Her75, Hun74b, Iye88, Jan23, JL99, KS01b, Kne00, KD08, Koh78, KP11, LS98b, Lee99a, LY12, LW92, Mac75b, Mas84a, MV19, MM83, Mit87, MR09, MS94, MM85, Nag71, Nar87, Nar91a, Now88, Pak89b, Pak89a, PS11, PT73, Pho79, Pic77, Qia90, RC94, Reh80, Rem00, RS12, ST96, Saw77, ST85, Sch03b, SB77, Sch78c, Sie76, SRTM97, Sta98, SY97, SK20, Swi75]. Asymptotic [Szc93, Tak81, Tie81, Tie82, Tuc71, TW84b, VL10, Wal74, WD71, Wes74, Wil90, ZHCC14, BP17, BP22, CYG14, DM22, Fra89, GMS23, HG15, Bez20, Gal84]. Asymptotically [BM98, DM05, Gla83, Kün83, Oça23, ZCJW19, ZCLJW23]. Asymptotics [Ahl15, BM11, BFW22, CWW09, DS07, DP12, EG17, Gat85a, GRZ04, Han07, Hor02, Hwa99, Jan18, LS11, LZ05a, LZ05b, LM00, MY99, MZ04, OC12, PR06, PZ17, RZ13, RSS99, Sch02, TC13, AD21, DL20, GIMM14, KP22, KM14b, Kur18, Miy17, PB19, PW18]. asymptotique [Gal84]. asymptotiques [Fra89]. ATM [BDM99, BM98]. Attachment [Jor10, AH16, GS19, JSS19, JW15, PRR17, PP15, Ros13, WR22]. Attained [Coh91, SW90]. Attaining [Gue93]. Attraction [GR88, Hen87, MPS98, KXZ23]. Attrition [BC78, BC81a, Whi96, EG17]. augmented [LL18, Mas15]. Author [Ano89e]. Autocorrelation [Kas80]. Autocovariances [Ano86e, Fra85]. automata [BMM13, Cas16]. Automatic [DTM07, Ger77]. autonome [Fri86, Fri87]. Autonomous [GY78]. Autoregression [KS74a, LS03]. Autoregressions [BW97]. Autoregressive [Bha86, BLS88, BFL09, GL80, JL77, KS06, LL81, McK86, Rud98, Hue16]. Autoregressive-Moving [JL77]. aux [Gal84]. Availabilities [FN85, Gås12, Nat86]. Availability [DPY08, DG01, JD89, Maz89, Nat80, Pol75]. Available [Lee99a]. Avalanches [RR21]. Average [BS81, BR86, BLS88, Fas05, FT78, GY10, Hal86, HV83, JL77, McK86, MMC91, MS11b, Pfe85a, Sen97, SW80, Wal74, CCHH18, HSZ18, MVY19, PZ17]. Average-Case [Pfe85a]. Average-Cost [MMC91]. Averages [BKM92, MD94]. Averaging [GT12]. aware [MMP22]. Axiomatic [Ant79].

b [Kem75]. Båth [VJ08]. Back [Ano69a, Ano69b, Ano70a, Ano70b, Ano71a, Ano71b, Ano72a, Ano72b, Ano72c, Ano73a, Ano73b, Ano73c, Ano74a, Ano74b, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano75d, Ano76a, Ano76b, Ano76c, Ano76d, Ano77a, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d, Ano80a, Ano80b, Ano80c, Ano80d, Ano81a, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano82d, Ano83a, Ano83b, Ano83c, Ano83d, Ano84a, Ano84b, Ano84c, Ano84d, Ano85a, Ano85b, Ano85c, Ano85d, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano88b, Ano88c, Ano88d, Ano88e, Ano89a, Ano89b, Ano89c, Ano89d, Ano90a, Ano90b, Ano90c, Ano90d, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano92d]. Back [Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09a, Ano09b, Ano09c, Ano09d, Ano10a, Ano10b, Ano10c, Ano10d, Ano17a, Ano17c, Ano17e, Ano17g, Ano18a, Ano18c, Ano18e, Ano18g, Ano19a, Ano19c, Ano19e, Ano19g, Ano20a, Ano20c, Ano20e]. Back [Ano20g, Ano21a, Ano21c, Ano21e, Ano21g, Ano22a, Ano22c, Ano22e, Ano22g, Ano23a, Ano23c, Ano23e, Lin10, Möh04]. Background [Le89a, MZ04]. Backward [CEP10, Kal80, LMX05, MT09, ØSZ11, Aaz13, GLL20]. Backwards [AK79]. bacteria [GM14]. Bacterial [HMV06]. Baddeley [BV97, RS98]. Bailey [Dav70, WD71]. Balance [BB96, FE98, Pol87, Tör78, Whi86, JJL21]. Balanced [Ben75b, HPPT92, KM12, Jan20]. Balancing [AR06, BF19]. Ball [CS06, Gne00, Küf94]. Ballot [CS91a, Lef07]. Ballot-Type [Lef07]. Balls [AD93, BE06, Gne99, Sar98]. Banach [BR76, Kue75]. Bandit [Agr95, DPY08, DG01, Gla83, GRHK06, LY88, WW91]. Bandits [BF80a, BF80b, BK91, GNMA02, GKRH06, GHK13, NM01, AM22, HG15, ZCJW19, ZCLJW23]. Bandlimitedness [Wis80]. Bang [Bré76, SW92]. Barrier [CJPY97, Cre74, IM08, Ryc92, Fro13]. Barriers [Alp09, Lin15]. Bartlett [Bin79, Roy07]. Bartoszyński [PT85a]. base [MS14]. Based [Adl89, Agr95, ACS93, BL89, BE83, BO07, BL09b, Bul74, Cam84, DG01, Gås12, GP79, GAH03, HMM06, HPJ08, HJ02, HW09b, MGW96, Rai07, RT07, Ser75, Ste76a, Wei78, Wen89, ADP18, CP15, DFZ19, DJ17, HHHT19, JKR17, Mey18, MJM21, OS18, Pri15, SK20, Sva14, DE08]. Basic [DR89, Gri80a, Hun74c, Kön77, MNR02, MT11, PL96, Sch06a]. Basically [Had74]. Basis [And79]. Batch [BV90, DS73, Deb76, Fal88, MW96b, MT97, ÖB05, Van88b, YM98, Bar18]. Batches [IW70b]. Bayes [Han96]. Bayesian [BS99, BM03, BW96, CL00, GS13, Gla93c, Joy91, MZ20, Rie75, RS98, SK04]. Be [Hal87, JH92a, BP21]. beauty [GVW15]. Becoming [HP91]. before [BP23, Hol76a]. Behavior [AW87a, AW87b, AW88, BC89a, Bob08, FR12, FE06, FGM12, Fre82, Fri77, HM12, KD85, Kem75, KS06, Kne91, LN11, MS12, MB00, Mas84b, MRSV86, MS11b, Nar91a, Neu77a, Neu77c, PMG08, PT85b, Sie76, Ste02, Sur08, Tol88, VY94, WD71, BPP21, CMMS13, FX13, Gal84, Sto23]. Behaviour [Alb83, AA85, AA88b, BST09, BKO11, Bar75, BV06, Coh76, Coh77d, Coh88, DT80, DP84, Eth93, Gar80, GNR90, GM96, GZ93, Her73, Her75, HM10, Kos76, LT98, O’N95, Pak89b, Pak93, PR90, SB77, Sch78c, Wil90, BR13, MS17, NP17, Nea14, YS23]. Behavioural [Pru78]. Being [Don81]. Bellman [MY89, Sch82a]. Bernoulli [BK91, DLS94, Tak91a, Tam11, Ver73]. Bernstein [APP99, Odo80]. Bernstein-Type [APP99]. Berry [GZ11, Lan87]. Bessel [Bet98, Pit74, Pit75]. Best [CS81, Pet84, Pet82, Sam04b]. Best-Choice [Pet84, Pet82, Sam04b]. Bet [Pon05]. Bet-Dependent [Pon05]. beta [DY15, GIMM14]. beta-coalescents [DY15]. Better [AA88a, Che00, MS83, Sha84]. Betti [OT20]. Between [BS03a, TV10, AAL08, Ais79, Ano93g, BV02, BHB98, Bul74, DF18, DL82, DMPPG16, DMC00, Gar71, Gec77, GSS01, Gri92, Hil74, KS01b, Kol14, Lat77, Mån07, MY03, NP93, NS77, Ols07, Par77, PP15, Pfe85b, Pre77, SX08, She85, Smi80, SV76, Tam16b, Ver78]. beyond [AVLR23]. Bi [GKRH06]. Bi-Directional [GKRH06]. Bias [MD94]. Biased [Che00, DS03, KPO09, NS93, Pit96, Sud99]. Bidirectional [GHK13]. Bienaymé [SS09, SS15]. bifurcating [LY15b, PD22]. Bifurcation [Har71]. bilateral [dlI22]. Bilharziasis [Lew75]. Bilinear [DLO06, Liu90, TM91]. Binary [DM03, DH05, Ebr90, Gla96a, Maz89, Was78]. Binomial [AAL08, And99, JP98, Lit69a, McK89, McK86, ST85, IMS21, Ros13]. Biodiversity [AP05]. Biological [AS96, Mee82, Sch82b]. Biology [BFM^{+}84]. Biomathematics [Kit71]. Bipartite [CKW09, BGM13]. Bird [Hey78]. Birth [Aks75, AW79, Bar75, BL10, BK83, BC89a, Big76, BR83, BV90, BGR82, Bro85, CK73, Cav78, CZLR04, CY00, CC77, Con77, Dan77, Dav70, Ebe80, Fad90, FMP92, FG00, GM71, Gar95, GJL84, HMS03, Hel75, HH85, Hun83b, Hut80, Kei74a, Kes76, KNPV97, KK03, MT06b, Nat75, Qi08, QS00, Ser81, Sie76, SR82, Van80, Van85, Van91, Wu93, BMM16, BDK^{+}18, Bez20, CKO19, JS14, dlI22]. Birth-and-Death [Cav78, Ebe80, FG00, GJL84, HMS03, Kei74a, Nat75, BMM16, JS14]. birth-death [dlI22]. Birth-Growth [CY00]. Birthday [DKMJ10]. Births [Ste80a]. Bisexual [MA95b]. Bispectra [TM91]. Bivariate [Ahm81, AS95, AS96, BLS88, Che98, FZ96, GM79, GW13, GG90, Mil74, RS92, SY91, SK85, Wis72, ADL15, BCLF16]. Black [Pon05]. Blackwell [BCO95, Lin77]. Blaschke [Jen96a, Kid98, Wak85]. Blaschke-Petkantschin [Jen96a]. Block [MT06b, SB09, ZLB98, CGR18, GLM17, Mas15, Mas23, NRB22]. block-Hessenberg [Mas23]. block-monotone [Mas15]. Block-Structured [ZLB98]. blockchain [Gof19]. Blocking [CLW95, Cho93, GLM93, GC02, Kel86, Kog89, Mit87, SRTM97, Yun86, Zac91]. Blocks [BL10]. Blood [HW95]. BMAP [Alf03, Fer93]. BMAP/D/k [Alf03]. BMAP/GI/1 [Fer93]. Boarding [BBSS07]. Bodies [AB07, BFV10, BS95, Hof07b, Sto98, Vit95, Wei82, CD14]. Body [Rat96a]. Bold [JK85]. Bomber [SY90]. Bond [GW00, Wie81]. Bonding [Whi80b, Whi80c, Whi80d]. bonds [PHL16, US19]. Bonferroni [HS90, RS87]. Bonferroni-Type [HS90, RS87]. Bookie [BL06]. bookings [MT16]. Bookstein [Ken96, Ken98, KL00]. Boolean [Ano03e, BB01, Bal07, BS01, BRS99, CMP09, CSY98, DRS11, GKW96, HLW02, HW19, Kid96, Lan96, MP03, MS94, Mol95a, RT02, Sar97, Sar98, Sch03b, Sva14, VV01, Wei98, Wei01]. Bootstrap [AEM94, Bro95a, LY12]. Bootstrapping [FR97]. Borel [Bru80]. Borovkov [AH97]. Borrowing [Bor09]. Bose [LO04]. Both [Box79a, Box79b, CR11, GR88, HP83, Pau07, FZ13]. bottom [LT15]. bottom-up [LT15]. Bound [GZ11, Lan87, Maz89, SO85, WN12, FT18]. Boundaries [DM00, DG03, DG04c, GNR90, HCD94, Per85a, PL96, TF08, Ano09e, Oça23, STZ14]. Boundary [ACF09, Bar81, Beg12, Bet98, CRC04, DS93, Dan74a, LYA07, Ott76, PMG08, Sal88, Sys73, Van73b, Wor95, BHKK15, FS16a, PB19]. Boundary-Value [DS93]. Bounded [Kes95, Lem14, MZ77, Nak96, Pro76b, RL10, Sch03b, Sma76, Gal16, JS15]. Bounding [GGV89, MT11, Tru97]. Bounds [Agr74a, Agr74b, BMS89, Ber78c, BRS99, BB10, BS83, Buk01, Che98, DK97, Dia13, DS00, Fro04, FN85, Gås12, Haj82, HMM06, HV84, Hua02, Hüb80, Hun77, JVZ08, JY04, KS74b, Kwe75, LT01, Lep76a, Mer06, Nat80, Nat86, PGE96, Rai07, Ram84, Sat85, SÖ11, SS14, Tie82, TW84a, Van85, WW03, BCJW14, CKL14, HHHT19, Kub19, LSV18, LL18, LS15, Mas15, SS20, Ste15]. boxes [AS19]. Branch [BF05]. Branchement [Pia01, Liu97]. branches [DY15]. Branching [Agr74a, Agr74b, AP05, Ano79e, AK78, Ath85, AK98a, AK98b, Bac00, Ban09, BP79, BCM18, BA81, Big76, Big78, BB93b, BK04, BS05, BHJ11, BD74a, BD75, Bor98, CPP21, CLM08, CORR91, CLCZ12, DH97, DH91, Dur75, Eba78, Edl78, Eth93, Fil72, Fil74, Fuj79, Fuj80, Gat00, GB03, GLM90, Gre89a, GZ93, GP88, Gri92, Här81, Her73, Her75, Hey74b, Hol76b, Hop85, Hue16, Hui03a, IPV05, Iva80, Iva83, Iva89, JP98, Jag77, JN84, Jir74, JM86, Kap73, Kei78, Kes78, Kim83, KM10a, Kle82, Kle84, KM80, KP11, LJ04, LJJ04, Lam03, LRS10, Lev79, Li92, Li98, Lin74a, Lin74b, Lip75, Liu98, LMW96, Mac75b, Mat77, MT06a, MVY84, MY89, MA95b, NS93, Neu70, Neu76b, Ney74a, Ney74b, O’C95, Pak71a, Pak71b, Pak79, PT85a, Pak86]. Branching [Pak89b, Pak89a, Pak92, Pak93, Pak98, Pak10, Pia06, Pol74, RSY23, Ric11, Sav75a, Sav75b, Saw76, SB77, Sch78c, Sch79, Sch81, Sen74, Smi74, Spo84, Ste80b, SVW83, Sud99, SW97, Tan81, AD21, Ale13, AF18, BPP21, Bau14, BZ14, BMPR19, CN18, CH14, Che15, COC16, Clo17, Cra17, Duc21, EKKZ18, FHJK20, FMY22, GLL22, HLN13, HLZ18, Kal18, Kur18, Liu97, MYH18, RY14, SS15, Pia01]. Branching-Process [Spo84]. Branching-stable [BCM18]. BRAVO [DVN15]. Breakage [Jir74]. Breaking [GL04a]. Bridge [LS89, Ver78, BCSY15]. Brief [JH92a]. Bright [Sau78]. Brittle [GL04a]. Broad [Bri76, Sch74a]. Broadcast [FFH86]. Broken [FS89b]. Brown [OS18]. Brownian [AW87a, AW87b, Adl79, AP97, Att17, Bar81, Bat70, Bat76a, BCSY15, Beg12, BK05, Cau74, Cau76, CJPY97, CCH00, CH80, CK85, DMY93, Duf01, FS09a, FS08, GKL21, Gem80, GRW10, Imh86, KM15, KZ16, LW11, LS89, MPS98, MVY19, MY03, MZ20, Ott76, PV19, PZ17, Pit74, Pit75, Pyk75, RY14, RSY23, Sal88, Sch03c, Sch06b, SW13, Taq75, VJV10, Ver78, Yor92, Yuh82, ZÇ96]. BSDEs [Cha09b]. Bucy [CDJR22]. budget [CR19]. Buffer [Cao89, Kos76, Lee99a, Maj07, MS90, Mit88, RSS99, Sch76b, Zwa00]. Buffered [Szp94, DS15]. Buffers [AT90, Kel82, LM00, DVN15]. Buffon [DM90]. Bulk [BC89b, CR04, Dag75a, Dag75b, Dag78, LC08]. Bulk-Arriving [CR04, LC08]. Bulk-Service [BC89b]. Bundle [GL04a, PT73, Pho79, Tie81]. Bundles [Dan89, Tay79, Tie82]. Burke [Dad84]. Burn [CP15, Mi94]. Burn-in [CP15, Mi94]. Burst [GT75]. Burstiness [DR06]. Bursts [BD97, Bal97]. Bush [Fin80]. Bush-Mosteller [Fin80]. Busy [AB94, Box78, Coh72, DT80, Hau78, PR06, Ram84, Ros73, Ros76a, Sat85, Tru75a]. Busy-Period [DT80].

c [CK15, SW91, Tak81, TVF81]. Cache [Oso10]. calcul [Yca84]. Calculating [CHM91]. Calculation [Cao89, KS01a, Yca84]. Calculations [Smi76, DP23]. Calculus [AW96, PR94]. Calls [Fal84a]. Campbell [BB93a]. Can [Ano06e, ACDHT05, BP21, DEKS11, FS09a, Hal87, FM18]. Cancellations [LY77]. Cancer [BP83, Bar85b, HRY93, HJ00, Kel12]. cancerous [EKKZ18]. Candidate [CG08, QG11]. Candidates [HMV06]. Cannings [GRR17]. Cannot [JH92a]. Canonical [Hol78, Pre77]. Cantelli [Bru80, EPS79]. Cantor [Sha09]. Capacities [Cao89]. Capacity [AB15, DDSV87, FX00, Sch76b, Tak91b, TV78, WN12, YZ78, KN16]. Capital [DE08]. Caracterisation [Cap86, Dio78, Dio76]. Carcinogenesis [HRTY97]. Card [EL05]. Carlo [BCJW14, BT08, CR08, Dai11, GR08, Gla93b, LTT20, MT09, PH99, Pes75, PS94a, TKF15]. carry [WZ19]. Cascades [Bar01, Hui03a, Way96, Kir23]. Cascading [GL09]. Case [BP79, BFL04, BL06, CNP91, Fie76b, Fre82, Gat00, KR81a, KR81b, KT01, MY01b, NSHH86, Pak79, Pfe85a, Sch07, Bar18, CM19a, Jan18, KM14c, Man21]. cat [FPS22]. cat-and-mouse [FPS22]. Catastrophe [BGR82, Bro85]. Cauchy [RSS94, Tay77]. Caudal [Neu86a]. Causality [Liu90]. Cell [AS08, ABB^{+}80, Ban09, Cal02, Cal03b, Chi94, HRY93, HSS05, HW95, Jag77, MNW08, Ols07, Yao10, AG13, HH14, NB15]. Cells [BO07, BCM13, Cal03c, Cow10, HRS04, Mec96, Mee82, RSZ12, Sch94, Tau74, WB71, GC20, GKT22]. Cellular [Rob92, BMM13, Cas16, MS14]. Censored [Tho81b]. Central [AMS09, AK95, AK78, Bar74a, Bro95b, CQ01, FP12, GLS05, HM96, HM99, HS13, Hey79, JY20, Kha80, Lee99b, Nak79, PSS06, PD22, PR99, Qi08, Rog92, Rog94, Rud98, SU79, Sch02, ST10, Sco73, SGW96, Voi90, AS14, AM17, GW20, KWRK22, LPST14, Par98, TL20]. centraux [Par98]. Centre [PR90]. Certain [AS07, Als88, Bas79, Can74b, Can75b, Dav73a, Din94, Ell87, FKM82, Fin80, GM71, Hog86, KK73, Lin74b, Mil75, Mil95, NRT83, Ott75, PL89, PT85b, Sie79, Van84]. certaines [Die82b, Yca84]. chaînes [Let82, Yca84]. Chain [And99, BC04, Cha89, Coh79, DS93, FMP92, GP02, GH75, HV84, HS92, JLW05a, JLW05b, KD85, KR71a, Led04, Man88, PH99, PC96, RN70, Res72, RZ13, RS07, ST85, Sys77, Tol88, TS99, YS79, Als18b, Cas16, Eck19, FPS22, MLZ21, SWL14]. Chain-Binomial [And99, ST85]. chaines [Die82b]. Chains [AE95, AANQ04, AHO96, Ano93g, Ano04f, Ano04g, AN78a, AN78b, AK98a, AK98b, AL99, BMY94, BMTY97, Bat73a, Bat73b, Bat73c, Bat76b, BC87, Big87, Bob04, BFK92, BH04b, BC81b, Can74b, Can75b, Cao98, CPZC05, Coh76, Coh77d, Coh81, DL05, Eco05, EF80, EF81, EN80, EN88, GHT96, GHT97, Gal11, GTY70, Ger94, Gha08, GL79, GT08, HS94, HH03, HIS76, Hor02, Kai78, KA81, Kes76, Kes95, LT01, LL97, LZ05a, LZ05b, Man74, MY01a, MT92, MT93a, Möh98b, NP93, Neu78, Oka79, Per97, Pes75, Pit77b, Pol94, Pol87, PR90, PS94b, Ris91, Rob76b, Roo88, Rud73, Rud74, Sen80, Sen97, Sha97, SS77, Sys75, Sys79, Tav76, TV10, Tjø90, TT94, Twe76, Twe77, Twe82, VY94, Van90, YS94, YZ02, YZB03, ZL09, BMS18, Buc16, CCHH18]. chains [EP15, FS16b, HLLP16, KT20, Let82, LL18, Mas15, Mas23, MV19, NR21, PSTB17, PD22, XHW23, Yca84]. Champ [Pia01]. Chance [CW75]. Change [BK04, Jac80b, Neu77a, Neu77b, O’N95, Pic85a, YZ78]. Change-over [Neu77a, Neu77b]. Change-Points [Pic85a]. Changed [FL09, Gzy80]. Changes [BV90, Deh98, HJ00, Mor76, Neu71]. Changing [GJL84]. Channel [BMY94, BD97, Bal97, FFH86, IW70a, IW70b, Rob92, SY97, Whi70, WN12]. Channels [AW71, SSS04, Van71, Van73a]. Chaos [MS01b, MS04]. Chaotic [TC92, PJCB18]. Character [Bul74, LL97]. Characterisation [Bin79]. Characteristic [Ell87, Hey78, KPUU01, Luk72, Neu86a, SB98, Wor94, Wor95, JL15, TO21]. Characteristics [BRS99, Bro89, Cal03c, GS95, KM95, MT11, Sch87]. Characterization [BH04b, Cha14, CT08, DJ91, Fac09, Fei78, Gro78, KR85, KM12, KT20, Kot80, Let85, Mec77, NW05, Rao83, RS86, Rie80, RS92, SY91, TF08, Wes90, Ano09e, Dio76, Dio78, Fou18, Cap86]. Characterizations [Gup80, Mel79a, NZ77, Tak88]. Characterize [KD94]. Characterizing [JT13]. Characters [Smi79]. Charged [Lep76b]. Chart [Bat76a]. Cheeger [ACPP12, FS16b]. Chemical [AT80, Arn80, DSSW11, Git69, Kot87, PV92, JJL21]. chemostat [CMMS13]. Chen [BH89a]. Chi [SR76, AC22]. Chi-Square [SR76]. chi-squared [AC22]. Chiasma [KL83]. Children [Ste80a]. Choice [AGSC04, BW96, Pet84, Pet82, RR90b, Sam04b]. Choosing [CS81, Pes75]. Chord [AS08, DEKS11, HV01, HR04, HO09, Hei98, Muc10]. Chosen [Man94]. Chromosome [Ben75b, DAG95]. Chronic [Mes71]. Circle [CS91b, GACO00, Hui03b, Hui04, JM13]. Circuit [Kel86, Kel88, KF90, Kog89, Mit87, Mor94, RT90]. Circuit-Switched [Kel86, Kel88, Kog89, Mit87, Mor94, RT90]. Circuit-Switching [KF90]. Circular [Che89a]. Clade [BF05]. claim [BWZ19]. Claims [AEM94, AS07, BH98, GW13, HW09a, LTW10]. Clamped [Eli77]. Clark [Ano98e, FP96]. Class [AA88a, BW93, BE88, Bas79, BL08b, Bri76, CLM08, CHM91, CLW95, Coh78, Cro07, Fin80, FE06, GV98, Gla93c, GNMA02, GM79, HM96, HM99, Hös11, Jac05, Kar83, KL79, KKSS90, Kem75, KM12, Kog89, KP12, LG05, LMHN90, Mil98, Mit87, Nyr98, POGP95, Per97, Pie78, RL10, Sch81, SS85, TV78, VJ05, Voi90, Wis80, APZ21, AF18, BCJ^{+}17, CG16, Cha22, CMT21, CKO19, EC21, GB17, HH14, JLY22, Sak15]. Classes [AST10, Arj80, BP83, BMV03, CNP91, Cow10, Dav73a, EPS79, KR81a, KR81b, MM83, MM85, Mül97a, OM95, SM87, CH14]. Classic [MM96b, Mil98]. Classical [KW87, KM10b]. Classification [FH02, Ger77, HR78, HMS03, Pai01, XHW23]. Classifying [Twe76, Twe77]. Clines [Dow74, DS76a]. Clinical [WDH98]. Clique [Pen02, TYA16]. Closed [AS95, BS12b, BS03c, BGPS87, Cao87, Cao89, Dan74b, FM95, HCD94, KP83b, KOS03, Mat72, MM83, MM85, Nås96, Sch02, Sve93, Zha97]. Closed-Form [BS12b]. Cloud [GG12b, BF19]. Clouds [Stu88]. CLT [Rai07]. CLT-Related [Rai07]. Clump [BRS99]. Clumping [LMW96]. Cluster [BVM96, BK02, FS09a, Jan19, RPS01, SC83, Ser80, Wes74, BWZ19, Ber20, GW20, Hou18, Kir23, LZ17, LP21, US19]. Clustering [BD97, Bal97, BM00, BV90, QT97, SKF80, BY14, CL14, ORS18, vdHvdHLS20]. Clusters [Sar98, Seg03, Cza23]. Clutch [Hey78]. Co [Sar97, Sha80]. Co-Existence [Sar97]. Co-Optional [Sha80]. Coagulation [ILP18, Spo84]. Coalescence [BC04, Gor01, Hös11, Lam03, MP08, WH03]. Coalescent [Ano86f, BS12b, BF05, CK99, DG04a, DG04b, DT86, Gri06, Möh98a, Möh98b, Möh00]. Coalescents [Fou11, FM09, DY15, GIMM14, KW18]. Coarsening [Gil98]. Coefficient [LS03, PGE96, PS10, RH81, Rud98, Whi80a, PSS20]. Coefficients [Ahm80b, AK05, Bra86, Fie76a, Sen79a, Sen79b, Sen88, TW84b, ORS18]. Coexistence [DH04]. Cohen [Ano78e]. Coherent [Ano83f, BP21, DK97, Iye88, Maz89, Nat82, HA20, HM20, NSB13, SN16]. Coin [Che89a, Che00, KPO09, Soo10]. Coin-Tossing [Che89a]. Coincidence [Mac75a]. collaborating [AAM17]. Collapse [BPSZ06, Rob07, Pri22, Sim15]. Collapsibility [DR03]. Collection [Mah10, MS12]. Collective [PL90]. Collector [DP12, Sta90a]. collectors [FRLM20]. collegial [Ner18]. Collinear [Man90b]. Collinearity [Man90a]. Collision [CLCZ12]. collisions [Nak14]. Colon [Kel12]. Colonising [Ros79]. Color [AMS07, AMS09]. Combinations [GJ80, Tho81a]. Combinatorial [Pol75, Tak75a, Tak75b]. Combinatorics [FG00]. coming [BMM16]. commande [Lef88]. Comment [Hog86, Kon89b, Nat89, Sat84]. Comments [PT81, RSS94, RG84, Sys88]. Commodities [KP83a, KP84, Per85b]. Commodity [BSC11, BV13a]. Common [AT81, Cha99a, Cha99b, DWH^{+}99, HMM06, Möh04, O’C95, HI79]. Communicating [Bat73b]. Communication [FVK08, Mee82, Van71, Van73a, WR76b, vdHK04]. Communications [BB01, BY09, Hoo78]. communities [CL15, vdHKV21]. community [GLM17, Jag18]. Compact [FM95, Hög76, Mas79, MOK95, Mrk01]. Compactness [Sch73, Wij78]. compacts [Mat74]. companies [AAM17, Nyr16]. Comparative [MM98, Møl95b]. Compared [Had76]. Comparing [HI79, Son79a, Son79b, Whi81a, SN16]. Comparison [Coh85, CEP10, DLS04, DMC00, Hol76b, Ker86, LY94, OB89, RSW16, SS95b, Tru97, BY14, SK16]. Comparisons [BE85, BS99, BK99, JML92, LS11, Ste81, FN13, LSH16]. Compartmental [BD87, CF78, Fad77]. Compensation [AWZ93, Sys74]. Competing [AG81, BDM85, BL04]. Competition [Bul74, CM93, GW78, Sch82b, Wes79]. Complementary [Cow06, Mil71b]. Complete [Ano91e, CR19, Don81, Kai78, KM95, PS90b, Pla02, PT77, RK89, Tru97, Fox16]. Completion [Chi95, CY00, KNT87]. Complex [All82, CTS76, CTS78, Hak82, Miy83]. complexes [TYA16]. Complicated [GTY70]. Complications [Sin98]. Complimentary [Geo76]. Component [Ano79e, Bue88, CB99, CB00, HZ06, Lev79, MW98, Pak71a, RS21, Wal73, CYG14, HA20, Seg20]. Components [Cao99, DGN92, DMPG09, Ebr90, ENS93, HQ02, Hui03b, Ito79, Ito80, LM90, Lin78, Lin79, LC92, Nat80, DFZ19, HM20, HL17, MMZ22, NSB13, Rob18, SN16]. Comportement [Gal84]. Composition [Ben74, Ben75a, Gat85b, JN84, ZL09]. Compound [BU98, BM00, BNX02, CY05, Cek97, CW03, CGR18, DTV84, EJMT85, PP07, RR90a, Roo93, RS07, SV76, Teu79, BM16]. Compound-Geometric [SV76]. Compound-Poisson [SV76]. Comprehensive [Cow06]. Computable [BMS89, LS15, CKL14]. Computation [Ken88a, Sen97, Gal16]. Computational [BC89b, DNPR01, KS75, Mit87]. Computations [Duf04]. Compute [CLW95, MS01b, MS04, AM22]. Computer [BRB^{+}71, Hoo78, Kei76, KR97, VK78]. Computers [DHK^{+}84]. Computing [AP98, FZ85, GG17, GG12b, PS94b, Sen80, BGZ16]. Concave [CM98]. Concavity [AT90, MS90, RA98]. concentration [FRLM20, PJCB18]. Concentrations [Gra82]. concernant [Liu97]. concerning [CC99b, Liu97, Mil95, Smi88, Van83, Yu08, ZCWS96]. Concomitant [LT98]. Concurrent [YM98]. condensation [Yua22]. Condition [BL89, DS76b, Hel75, LK93, WV80, Che15, XCWC18].

Conditional [Ale13, Ano03e, AFH09, BD93, CDDP11, FT12, Gla93b, HL85, Jan84, KMW96, KNPV97, KS90, KR02, Lan96, Pap77, Rud74, Rüs91, Sal84, Sha03, VV01, WS11, BvdMS20, CV18, Hue16, JRR14, KT20]. Conditionally [BCL10, NR06]. Conditioned [Arj80, Asm82, Igl75, Ken74, ST94, Tav76, VS80]. Conditioning [JLW05a, JLW05b, EMOS14]. Conditions [AW87b, AM22, BH04b, CFPL12, DLS04, DW06, Fal84a, FT78, GH80, Gri75b, GY10, HJK05, Käm87, KVB99, LMX05, Nak96, Szp94, Van85, WCK96, BCHK18, Mas15, MT13, PD22]. Conductance [JY04]. Conductivities [BNK98]. Cone [Nos94, FJKR20]. Conference [ADW^{+}75, BBB^{+}76, MDB^{+}75, MKS^{+}78]. Confidence [PGE96, Rav77]. Configuration [Ano83e, CR11, LB10, Rub81, KWRK22]. configurations [DFRV14]. configured [MS14]. Conflict [BC76, Whi96]. Conflicts [Imh08]. Confluent [Gem80]. Congested [ST90]. Congestion [MAF07]. Congestion-Dependent [MAF07]. Conjecture [CCP91, Gre87, Mil95, Rol81, Tra74, Van83, Wak85]. Conjectured [Van84]. Conjugate [WHL17]. Connect [Ano06e, ACDHT05]. connectable [HJKP16]. Connected [Cao99, CB99, Hui03b, Ill98, LT00, MW98, CT18a, CT18b]. connectedness [JKM23]. Connecting [DL11, HPPT92, Pol86]. Connection [ENM89, Kai78, Mån07, Teo82, DF18, GLM21, Gri14, JKM23, LZ17]. Connections [PT77, BGV22, Yin23]. Connectivity [BBSW05, DGP10, GX07, Gar99, HNS13, IY12, MS01a]. Consanguineous [Haj76]. Consecutive [CP90, Hsi10, Pap87, Pap88]. Consecutive- [CP90, Pap87]. consensus [Hir17]. Consequences [Hio79, Ton82]. Conservation [LS96, NM01, Num79, Sig91]. Conservative [AS07, Rol78]. Consideration [SR76]. Consistency [ACO96b, ACO96a, Le98, Moo79, NB15, Ran77]. Consistent [DJ91, ZG88, CL22]. Constant [ACPP12, BBSW09, Bar78c, Cox80, Der83, Kes80, Maj07, Wes90, CDJR22, Ste15]. Constants [Sch82a, BCJW14]. Constrained [BR86, HS89, Ker86, LH07, CL22, Jan23, JMFPY19]. Constraint [BK99, CFW87, FOF06]. Constraints [Ala99, BMS89, CS12, DPY08, DTV84, DHP12, Gla91, Gla96b, Sek06, DP13, PS83, RW14, Sto17]. Constructed [CQZ03]. constructibility [Mas23]. Construction [DS23, Gat94, HS10, Jac80b, MNR02, MNW08, KTR19]. Constructions [CT08, Kur18, NB15]. consumer [Lan13b]. Consumers [Mit88]. consumption [FLR22, HS18, YY22]. Contact [BKO11, Gri80a, HBG99, Hei98, HLW02, LS98a, Mol78, Muc10, Wei82]. Contagion [DZ11a]. Contagions [CL15]. Containing [Cal02, Lin78, Lin79]. Contamination [Ban09]. contemporaneous [PSS20]. Content [Bro77, Mat99, RSS99, HS13]. Content-Dependent [Bro77]. contest [GVW15]. Context [JP00b]. Contingent [BH98]. Continual [MW78]. Continuation [Lev78]. Continued [FG00, Pea89, Woe85]. Continuity [Cli11, Esh72, Esh73, KA81, Sch73, Whi74]. Continuous [Aas82, Alp09, Ano93g, ACPP12, BMY94, BMTY97, BNJS98, BGRSC10, Bat76b, Bin75, BSS23, BW97, Bro99, Buc16, CPZC05, DL11, Epl86, Fad74, FZ85, Fei76b, GHL07, GY10, HK90, HMS03, HS74, HIS76, JK11, Kap73, KM80, LY94, Leh02, LM04, MW96a, MT93a, MT93b, MVY84, NP93, PT85a, PP96, PL89, Pes82b, Pfe85a, Pla02, Pol87, PR90, Sam04a, Sch07, SR00, Ste80b, SBT96, Van90, Wol80b, Wu93, BPP21, Bar18, BDK^{+}18, Bez20, DP15, EY14, EP15, FMY22, GVZ13, GHH15, GKPZ21, HMM19, Kur18, XHW23]. Continuous-Range [LM04]. continuous-space [BDK^{+}18, Bez20]. Continuous-State [PT85a, BPP21, FMY22, Kur18]. Continuous-Time [Aas82, Alp09, Ano93g, BMY94, BMTY97, BW97, Bro99, CPZC05, Epl86, Fei76b, GHL07, GY10, HMS03, HS74, JK11, KM80, MT93a, MT93b, MVY84, NP93, Pes82b, Pol87, PR90, Sam04a, Sch07, SR00, Ste80b, SBT96, Van90, BSS23, BPP21, BDK^{+}18, Bez20, DP15, EP15, GVZ13, GHH15, GKPZ21, Kur18, XHW23]. Continuously [Ryc92, Stu88]. Continuum [BBD02, CR11, Dei03, DH04, LBCEN21, Pen91, QT97, HT17, Hou18]. Contour [SL80]. contracting [Fly18]. Contradictory [BP21]. contrainte [PS83]. Contributed [BBM^{+}88, EFG^{+}92]. Control [AT75, Ahm80b, ADZ04, AS79, AH97, Bar79, Bat76a, Bäu02, Ben74, Ben75a, BK80, BDF^{+}84, CR04, CK85, DS73, Deb76, Dos78, EF80, EF81, Fad74, Far76, GRW10, HRY93, Has76, HS89, JS83, JS80, Kar83, Kli80, LW11, Man74, Mor73, MW74, Nec78, ØSZ11, ÖB05, Ott06, Rob76b, Rud72, Ser81, Sti78, SW92, TE96, TV78, Wan01, WS87, Wee02, WR76b, Whi81b, Yas85, BPS15, CS17, DF18, DP15, DM22, GKPZ21, HSZ18, JvL15, KMF22, Lef88, LF13, Lem14, MVY19]. Controllable [Gal79b]. Controlled [DH82, Dsh85, Duf77, Hor03, Rat75, Shi93, CCHH18]. Controlling [FR80]. Controls [Bré76, Gro78, XZ96, WHL17]. Convergence [Ais79, AP98, Ald80, Ano98e, BO94, BH89a, BW90b, BFP98, BS06, BS05, BS78, Bor79, BK99, BP94, Bro89, CB95, CMP09, Cer98, Coh82a, DE08, DJ91, Dou75, DW96, FP96, Fox14, Gol77, GH80, GLS01, GVW15, Han78, HS85, Hey77, Hor91, Igl73a, Igl73b, JY04, JM86, JRR14, Ken72a, Ken72b, KW88, Led79, Lin74a, LY12, Man90a, MY01a, MP03, MS76, MRSV86, Möh98b, Möh00, MS04, Oka79, Ott06, PM08, Pol79, Pri99, RW03, RT23, Res80, Res86, Roo83, RS88, Sat78, SU79, SF79, Smi82, Taq75, Teu79, TT94, Van91, Wil88, Yin00, BMS18, CV18, CM17, DSS17, DSS18, FN13, Fox16, GLL20, HSZ18, HL14, MY20, SWL14, Yua22]. Convergent [Coh79, VV78]. converges [CR19]. Convex [AD93, AB07, BFV10, Bat73c, BS95, Bia02, BY09, BV02, BHB98, Bro89, Cow07, CM09, CS12, DLS00, Der09, DHP12, DW96, EG81, EPS79, FH04, Gec77, Gro12, Hof07b, Hue94, Hue99, Ken85, KL87, LH07, MR02, Mas79, Mas00, Mil75, Mol93, MOK95, ON96, RT02, Sto98, VL07, Wei82, Buc13, CD14, FKV14, Lac20, TW18]. Convex-Ordered [BV02]. Convex-Polygonally [KL87]. Convexity [SS88a, SS90]. convolution [XCWC18]. Convolutions [GPD00, Ken77c]. Cooperation [Sch82b, CAR20, LR18]. Cooperative [PS09, PS11, CPP21, WW21]. Coordinates [Whi95]. Copulas [LN11]. corde [DM90]. core [KL13, SKF80]. Corollaries [KS01a]. Corrected [Fuh97, LYA07, Sie79, GLM17, Hog86]. Correction [Ano77e, Ano79e, Ano80e, Ano83e, Ano83f, Ano84e, Ano85e, Ano86f, Ano86e, Ano89e, Ano91e, Ano92e, Ano93e, Ano93f, Ano93g, Ano94e, Ano98e, Ano98f, Ano03e, Ano04e, Ano04f, Ano04g, Ano06e, Ano06f, Ano07e, Ano09e, Ano12, DSS18]. Corrections [Ano76e, Ano78e, Ano79f, Ano95e, DDH78, Dio78, Wor95, Yuh82]. Correlated [DGN92, Had80, HK90, Lat88, POGP95, SW13, SS85, GLM22]. Correlation [BD87, Bal07, BD93, CI00, DS16, DH05, FS08, Gne99, Gne00, Hil76a, Ols07, Smi80, SP80, Whi80a]. Correlations [AS96, FK89, Tho81a]. Correspondences [Fér77, GSS01, Tam16b]. Corrigendum [Bez20, ZCLJW23, BP22]. Cosmic [Gre78]. Cost [AK08, AD94, CD79, DHP12, FT78, MMC91, Pic83, RX91, BS15, DP13, GVZ13, MVY19]. Costs [BDM85, CFW09, Deb76, Hun89, IS06, KK92, Pau07, Rob76b, Tam08, CIL17, DK16, Mar16, MMP22]. Cotes [SK20]. Count [Hua02]. Countable [Ano04f, BS15, Coh77d, Ken00, Ken78a, YZ02, CM19a]. Countable-State-Space [Ano04f, YZ02]. Countably [MY01a, MZ04, MT06b, HLN13]. Counterexample [HK92, Smi88, Van83]. Counterexamples [BCO95]. Counterpart [Bru80]. Counters [BC87, Box79a, Box79b]. Counting [EL05, LR97, MS87, Neu87, PL96, SS88b, Whi81a, CG16]. Country [BRB^{+}71]. Counts [AK95, Ano06f, FT12, KO87, McK88, Rem00, Rød06, CGR18, OY22]. couple [Let82]. Coupled [GT12, Mit88, Wri92, GPHK21, JY20]. Coupling [Cao95, COC16, Lin92, Tho83, Whi86]. Couplings [LMS00]. Coupon [DP12, Mah10, MS12, FRLM20]. Coupons [MS12]. Covariance [BS03c, CB95, CB03, CFD85, FVK08, Gil02, GSS01, KOS03, Ma05, Mas83, Ott76, Ott77a, Ott77b, Rey75, Din19]. Covariances [KS12, WHZ10]. covariate [CP15]. covariates [ŠSKB13]. Covariogram [AB07, GLR15]. Covariograms [Bia02]. Cover [Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano17h, Ano17g, Ano18b, Ano18a, Ano18d, Ano18c, Ano18f, Ano18e, Ano18h, Ano18g, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano20h, Ano20g, Ano21b, Ano21a, Ano21d, Ano21c, Ano21f, Ano21e, Ano21h, Ano21g, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e, CCH00]. Coverage [ARS04, AFM91, BB01, KT01, MS03b, Ros89, Sar16]. Covering [HSS08, Hui03b, FRLM20]. Cowan [Mec89]. Cox [AB22, Bri99, HPJ08, JT22, MS15, MS16, Møl03, MT05]. Crack [NW05, TS98]. Creation [Jac74a]. Credit [VV06, SMAL17]. Criteria [Bha86, CPZC05, GS78, KK73, Kes76, MT92, MT93b, STC94, Twe76, Twe77, GVZ13, HP23, HLG18]. Criterion [Ada16, BW96, CD79, DHP12, Mer79, MS04, CCHH18, DP13, HSZ18]. Critical [AP05, BBSW09, BD75, ENM89, GH80, Gri81, Gri75b, Her73, Hol76b, Iva83, Kir23, LJ04, Maj07, Ney74b, Pak98, Pak10, Rah07, RT02, Sze76, VJV10, Whi82b, Wil74, Ada16, Bau14, GLL22, ILP18, OT20, Rob18, SS15, Sim15, TO21]. Criticality [GI10]. Critically [AMNW11, HK89b]. Crofton [Cal02, HRS04, Jen95]. Crop [CS81]. Cross [BV13a, Cur74]. Cross-commodity [BV13a]. Crosses [Cur74]. Crossing [BH04a, Beg12, Bet98, CP08, Col98, DGN92, DM00, Gri80b, Lef07, LN92, PL96, WA82, Wil88, PB19]. Crossings [AA86, BL08b, Hsi87, Kon89b, Lin78, Lin79, Lin80, Sha78a, Sha78b]. Crossover [KL79]. Crump [BD75, Ber20, Edl78, Gri92, Hol76b, Jan18]. Crump-Mode [BD75, Jan18]. Crystal [KS04a]. Crystalline [IM02]. Crystals [Møl95b]. Cubes [Adl79]. Cumulative [Ano94e, Gut90, Pel93, POGP95, Rog94, SS85, Yas85, DM22]. Current [CS81, Gan79, TS98]. Curvature [IM02, Rat96b]. Curve [Ano06e, ACDHT05, HJ00, Lie76, Neu86a, AZ18]. Curved [Bar81, Dan74a, DM00, DG03, DG04c, TF08, Ano09e]. Curves [CY99, Gol77]. Customer [BM77, RK89, SG92, TX10]. Customers [BBH84, BB93a, BGPS87, CKW09, De 83a, FX00, HR89, HP96, He96, HP91, Neu77a, Neu77b, Ros73, Web92]. Cusum [Sta92]. Cut [ACPP12, Lit69b]. Cutoff [MY01a]. cuts [TSvB17]. Cuttings [Hol11]. Cycle [Box78, Coh72, DS07, Har84, Thé22]. Cyclic [Box83, BY89, KMW96, LOS03]. Cyclic-Type [BY89]. cylinder [HS13]. Cylinders [Hof07b].

D [Alf03]. D. [Mil95, Van83]. Daley [Ano95e]. Dam [Kha86, Min80, PL89, Tur74]. Damage [Ano94e, GL09, Pel93, PZ86, Sob87]. Damped [Ors84]. Dams [Bro77, BPS80, Fad74, Kha80]. dans [Par98]. Darling [Sho79]. DARMA [JL79]. Darwinian [Sch82b]. Data [Ais79, Bad80, BT71, Cur81, DR06, DR08, DK97, FL09, Gla96a, Gor01, Hil81, Kos78, LA02, Maj07, MD89, MW94, Mas83, MKS^{+}78, Ste76a, Ste79, Wah76, ZCWS96, BN23, He16, KS16]. d’attente [Bré75]. Davidson [Pap77]. Day [Cha99a, Cha99b, DWH^{+}99]. de-preferential [BK18]. Dead [BGR06, GG12a, Gil02, Mån07]. Deadline [Bro99]. Death [Aks75, AW79, BD87, Bar75, BL10, BC89a, BR83, BV90, Bro85, BD93, CK73, Cav78, CZLR04, CC77, Con77, Dav70, Ebe80, FMP92, FG00, GM71, Gar95, GJL84, HMS03, Hel75, Hun83b, Hut80, Kei74a, Kes76, KNPV97, KK03, LP96, MW78, Mer79, MT06b, Nat75, PL99, Qi08, Ser81, Sie76, SR82, Van80, Van85, Van91, Wu93, BMM16, CKO19, JS14, dlI22]. debate [LT15]. Debit [Cai07, ZYY11]. Decay [DEKS11, GI10, LC08, MY01a, MT06b, Ros74b, TT79a, Van85, ZG04, GHL23]. decay-surge [GHL23]. Decentralization [Kel88]. Decentralized [Gha08]. Decision [Bal78, Bat73a, Bat73b, Bat73c, Bat78, BR86, DFT77a, DFT77b, DHP12, FT78, GY10, HV83, Hüb78, Hüb80, Ken78a, Rob76b, SF79, Sen97, SW80, Sza93, VW77, Van78d, Whi69, BS15, DP13, DP15, GVZ13, GHH15, GKPZ21, HLG18, PHL16]. Decision-Makers [Sza93]. Decomposability [BNPS01, FE98, KM95, Miy93, Tru97, Urb77]. Decomposable [Ito79, Ito80]. Decomposition [BM77, BS03b, CA94, CQZ03, ES04, Gås12, GP80b, Jac74b, Kal80, Koh80a, Koh80b]. Decompositions [Gre75, Maz89]. decoupled [Aaz13]. Decrease [DGR02]. Decreasing [Aas82, BD73, MVY84]. Decrements [Pak89a]. Default [CFW09]. Defaultable [BWY10, PHL16]. Deffuant [Hir17]. Define [Ano83f, Nat82]. Defined [GH75, MPP07, RN70, Res72]. Deflated [GW00]. Deformations [SCA00]. Deformed [BM03, Fou18]. Degeneracy [AS08, Yao10]. Degeneration [AS95]. Degraded [Hal87]. Degree [AR97a, AR97b, BJML07, DGP10, DGM09, Hen87, Iye23, Jor10, Mey18, Min05, TM91, GLM17, PRR17, ZM16]. degree-corrected [GLM17]. Degree-dependent [Mey18]. Degrees [DM06, HJ08, KKV16, LR12, Van78c]. Delaunay [BTZ00, CD18, ENR17, GKT22]. Delay [Le 74, ØSZ11, SW90, VL98, WHL17]. Delayed [FD83, HP83, LP21]. Delays [Gri80b]. Deletions [CK07]. Delivery [HP83]. Delta [RS98]. Demand [BV93, DTV84]. Demands [FZ85]. Dense [Ken00, NRB22]. densités [Lef88]. Densities [Ant78, BN10, BN12, Bin80, BNR87, GNRS89, GNR90, Gup80, Hof07a, KL87, Le89b, Lin05, Mae79, Sch06a, Sch06b, WW84, Wei01, FT18, HJKP17, HW19, Lef88, STZ14]. Density [Bro72, Cra74b, DN02, GM91, Gri79b, Hah98, Hea78, Leh02, Lom78, RL10, Ric75, Sar98, Sch03b, SM87, Cor14, GC20, LX13]. density-dependent [Cor14]. Denumerable [AANQ04, Bat76b, FT78, Fed78, Sen97]. Deny [RS86]. Departure [BW83, Cha92b, FS89a]. Departures [AMNW11, Ali90, MW96b]. depend [FM18, FZ13]. Dependence [BN12, DR08, DS03, DZ11b, HL02, Hüs89, Lip75, Pel83, RPS01, FHJK20, LY15b, Vaš16]. Dependencies [DS95]. Dependent [AK93, Ano79e, BB07a, BW89, Big76, BL08a, Bor98, BV91, BB10, Bro77, Cao90, CYZ99, CR04, Con77, Con83, DRV00, DSSW11, Ebr90, Edl78, Fil72, Fil74, FJ09, GW92, GH80, HJVN03, Hal86, HNT94, HK93, Jac80a, JP00b, JD89, KD92, KR71a, Kim83, Kle84, Kne00, LJ04, LJJ04, Lev79, LTW10, Maj07, MAF07, McK88, MP10, Mer79, MVY84, MY89, Ney74a, Pak71a, PP02, PMG08, Pon05, RG89a, Ric11, Rob07, Ros76a, RH89, Sav75b, Sch71, Slu84, SS88b, Wal73, Wau74, Web92, CT18b, CP15, CM19b, CH14, CKO19, Cor14, GLM21, HM20, He16, Jan23, JS14, Kom18, KS16, LP21, Mey18, STZ14, SJ18, SK16, BF80b]. Depending [HH81, Pet82]. Depletion [ENPS78, KS83b]. Deposition [Hui04]. Derivate [Tod77]. Derivation [Ant79, Bha86, CTS76, GM91, Miy83, Vit95, Yun86, CHY20]. Derivations [Mil75]. Derivative [Ano91e, PS90b, Tod76, Che15]. Derivatives [Ber80, Gla93a]. Derived [Fie76a, Lie76]. Dermal [Ros76b]. Derrida [Hui03a]. Descending [DL05]. Descent [SB09, Kar21]. Description [GKT22, PR90]. Descriptions [Nyr98]. Design [BK83, Che00, HJ02, Le 78, ADP18, Sva14]. Design-Based [HJ02, Sva14]. Detail [Hal87]. Detailed [JJL21, Ram80]. Detect [Car78]. Detectable [Don81]. Detected [JH92a, Pro76a]. Detection [GS13, HRTY97, Pro76b, SC77, Ste80a, GLM17]. Deteriorating [LY94, RG89b]. Deterioration [AH87, SM78]. Determination [Ben80, HW19, KL87, Kid02, Rat98]. Determined [May78, Hir17]. Determining [Bia02, BC81b]. Deterministic [Ano12, AT80, Arr03, Bal85, BL08b, BDD12, CT85, Daw75, DK80, FLZ12, GT12, KD85, Kin79, Lin79, Mat98, MW74, Nea06, LTT20, Fad79]. Development [Bro80, Rej82]. Developments [Bar78b, BSD^{+}84, Pyk79]. Deviation [CS06, CP08, DM05, Kub19, KT92, Loa91, Maj07, Rai07, SH89, Whi95, HJKP16, PB19]. Deviations [BB93b, Bor79, BMS09, CYZ99, Cho93, FVK08, GKT21, GP02, HMM06, PP96, RH81, Sel96, SW11, Shi93, SW93, BR15, CT18a, CT18b, CF14, DJ17, GW20, MZ16]. Deviations-Based [HMM06, DJ17]. Devices [AH87]. Diagnostics [HJ00]. Diagrams [MH08]. Diallel [Cur74]. Diameter [ANR02, GX07, MS19]. diary [MT16]. Dickman [IMS21, PW04]. Dickman-Type [PW04]. Dietrich [DKMJ10]. Difference [CT85, Hor03, Kar90, Rob74b, ST94, Ver77, Ver79, Wol80b]. Differences [Rob74b]. Different [De 83a, MY03, RK89, RT02, RS21, Sam04b, Son79b, Ste81, Str75, LSH16]. Differentiability [AE08]. Differentiable [Ryc92]. Differential [AT75, Boy78, CEP10, EX11, LMX05, Nar91a, NZ77, ØSZ11, Shi79, Aaz13, BGL20, FPP19, LLX18, Lin15, Luo16, NP17]. Differentials [Hil76a, Kas80]. Diffusion [ARMAG00, AS77, AML77, Ant79, ACV80, AG05, AT80, Ber82, Bro85, BF19, CC12, Cha96, CX04, CR90, CMT21, DLS94, Duf77, EN80, EK81, EN88, Eth90a, Eth90b, EG90, Fuh97, GM71, GS13, GNR86, GNR87, GNR90, Gri79b, Har78, Hog86, HT10, Iva83, KW90a, Ken77a, KW87, Ken88a, Ken96, Ken98, Kes78, Kot87, KW03b, McG82, Oka79, Pan06, PS90a, Pau07, PV92, Roe74, Sat78, Saw76, SS84a, Shi82, Sie79, Sto15, VK78, Wal81b, Wat79b, WWY12, Wu93, YZB03, DL20, FPZ13, Gri14, Kal18, LZ13, Luo16, STZ14, ZY16]. Diffusion-scale [Sto15]. Diffusions [CR08, HS10, Jac74b, Ken77c, Ken78b, Lin05, MA95b, PMG05, PMG08, BvdMS20, BZ20, CV18, CV23, Fra89, GPHK21, JLY22, Lem14, MLZ21, Zha15]. Diffusive [Mer79]. Digital [Ano92e, Bro99, Pit86, Sva14]. Digitization [KJ03]. digraphs [PP15]. Dilation [KR18, Rat98]. Dilemma [CC09]. Dimension [Cro07, HTO^{+}15]. Dimensional [AWZ93, AE95, Alm97, AK00, AFH09, BX01, BFP98, BB95, BP94, BP99, Cal03c, CC09, Cov09, Cre74, CJ11, Deh98, DGN92, Duf77, FKM82, Gar95, GNR90, Gne99, Gne00, Hal85b, HKT91, HS10, Hut80, Ill01, IM02, Iso00, JK94, KK80, KW87, Ken95, Ken77c, Kes76, MSV01, NR85, Pea89, Per85a, Pit74, Pit75, Sch86b, Tho07, TC92, Yao10, Ahl15, BCT13, BCJ^{+}17, BPS19, CV18, CDJR22, FT18, FPZ13, JS14, Kar21, KM14b, KZ16, MLZ21, Sch16, XHW23, Zha15]. Dimensions [AS08, CK85, GT12, Gri81, Hun74a, Hun74c, Hun74b, Hun77, Muc10, Pol94, Sar97, VJ99, Zäh96, BCJW14, MV19, Sar16]. Diminishing [KP12]. Diplo [Heu79]. Diplo-Diplo [Heu79]. Diploid [Esh73, Hin94a, Hin94b, Cor14]. Dirac [Cau74]. Direct [Mil75, Fra89, JKM23]. direct-connectedness [JKM23]. Directed [BLP06, BR04, Gat97, LC92, PW04, PW06, Was78, AZ18, WR22]. Direction [Rat96a]. Directional [GKRH06, Kid01, KJ03, NW98, RS12, WN99]. Directionally [AM01, BY09]. Directions [CM03, Gan79, HS11, LOS03, LO04]. directs [Fra89]. Dirichlet [BVH10, GRR17, Xu11]. Disaster [CZLR04]. Disasters [Hai80, PS01]. disc [AT18, FKV14]. Discarding [LA02]. Discipline [BL89, DN78, HK89a, RW91, Sch90]. Disciplines [MH97, Neu76a, Rit94, Tam74]. Discontinuities [Sak04, RT23]. Discontinuous [RS88]. Discount [GY10, CCHH18]. Discounted [Ass78, Ath85, CFW09, CD79, DPY08, GKL21, GNMA02, GHL07, LTW10, PP07, BS15]. Discounting [AEM94, KK92]. Discovered [Con77]. Discovery [FS93]. Discrete [Aaz13, Abr79, Alf03, ACPP12, Bal79, BNJS98, BSC11, Blu84, BC04, CW03, Cek03, Cha09b, De 74, DL11, Elv71, Epl86, GVM^{+}03, GLS05, HC08, HT92, HIS76, Hor91, HM11, Hwa99, KS06, KM10b, Lim97, LS13, MMP22, Mel79b, MT92, PP96, Pic85b, Pit96, Sha97, Sho74, Swi75, TT79a, Van90, WS11, YC96, YZB03, Als18b, EW14, ENR17, HL14, Mar16, Mas15, PZ17]. Discrete-Parameter [Blu84]. Discrete-State [Mel79b, Van90]. Discrete-Time [Cha09b, Elv71, Epl86, HC08, HT92, LS13, MT92, YC96, YZB03, Aaz13, MMP22, Mar16, Mas15, PZ17]. Discretely [Sch07]. Discretization [GW00, LL15]. Discriminating [NS77]. Discrimination [Ger80]. Discs [MH08, FKV14]. Discussants [Cha99b]. Discussion [BÇNP84, DWH^{+}99]. Disease [CR78, Lew75, SC77]. Diseases [Mes71]. Disequilibrium [AH78, Hil74, Hil76b]. Disk [HSS08]. Disks [Cal02]. Disorder [Hui04]. Disordered [Din94]. Disorders [Wil79]. Disparity [ORS18]. Dispersal [BLS^{+}80, Lan13a]. dispersed [Hir17]. Dispersion [ZÇ96]. Dispersive [AABK86, Bar85a, DK83, LMP83, Sau84]. Displacement [Bul74, US19]. Displacements [Bac00, CK07, AB15]. Disproof [Van84]. Dissociated [BE85, Sil76]. Distance [AAL08, AS77, BV02, BHB98, Gec77, HLW03, NR85, PP15, PvS21, Ros79, She85, BP17, BP22, DMPPG16, Kol14, Sch16]. Distances [Col74, DMM12, Möh00, RR90a, KKV16]. Distinguished [Fou11]. Distributed [Ais79, AAR08, BCL10, CQZ03, DM06, Di 01, GO73, MRV02, Pen98, Rie80, Szp94, Tak70, TE96, HM20]. Distribution [AA88a, Agr74b, AMS07, AS08, And99, Ano93e, Ano06f, Bal86, BB92, BW90a, BNX02, BN94, BNK98, BP94, Box78, BN92, Cal03b, Can74a, CS99, CP90, CM98, Cow89a, Cse92, DT80, De 83b, DEKS11, Dwa74, EGV79, Eth90a, FT12, Fil72, Fil74, Fre82, Fuj79, Fuj80, Gal79b, Gec77, GO73, GL04a, GR88, Gor01, Gre78, GR96, HV01, Har77, HR78, HO09, Hei98, Hey74b, Hol11, JH92a, JVZ08, JS15, Jen84, Kat90, Kei76, Ken77c, KD94, Kid01, LS98a, LB10, LP90, Let85, LBZ98, Lin10, Lit69a, Lit69b, LY12, Loa91, LR80, Man90b, MPB96, McK89, MT97, Muc10, Nås96, Neu76a, Nie98, Nie99, Nor82, ON96, OC10, OB89, Ott87a, PT81, PH99, Phi78, PT73, Pho79, PL90, PLPA96, Rat96a, Reh80, Rød06, Ruk02]. Distribution [ST85, Sch76a, Sen80, Sen88, Sen89b, SS84b, Sma76, SK08, Sta73, Sta80, Swi75, Sys76, Tak70, Tak81, Tak88, WC90, Wau74, Whi80a, Whi95, WN12, Wol78, Xu11, Cas16, Din19, DM22, Gol16, GR16, IMS21, KT20, KM14b, Kol14, NSB13, Ros13, Sch16, ZY16]. Distribution-Free [Dwa74]. Distributional [BL07, Beh11, BP82, FN13, TWN12]. Distributions [AW96, AAL08, Agr74a, Ais79, AMS07, Alm97, Ano91e, Ano92e, Ano93g, Arj80, Bar76b, Bar80, BP00, BFL04, Bar85a, BL09a, BV93, BVH10, BS12b, Bob04, Bon82, BJML07, Bro77, BB96, Cal02, Cal03c, CDDP11, Cav78, CAK00, CS91a, CC99a, Coh74, Cov09, DRV00, DS03, DR85a, DL82, DKS84, DM00, DJ91, DM91, Ebe80, Eco05, Fac09, FS89a, Fei76a, FFG07, FMP92, GM71, Gat94, Ghu87, Gla93c, GR88, GLS11, Gri79a, Gri06, Had76, HBG99, HP82, HZ07, HS77, Hue99, HLW02, HL02, Hüs89, Hwa99, JT13, KR81a, KR81b, KD92, Kem73, Ken85, KNPV97, KM10b, KLS02, Kot80, KS80, KP12, Leh02, Lep76a, MSM92, Mar75, Mar77, MD89, MW94, MGW96, MPS98, Mas95, McK86, Mej77, Mer06, MM96b, Mil98, Mol95a, Myl02, NP93, Neu76a]. Distributions [Neu87, NPOTC00, Nor75, Now88, O’C91, ON98, OMS06, Ott87a, PW04, PMG05, PMG08, PS90b, Pit96, Pit86, PS94b, RR90a, RS86, Rei81, RS12, RSS99, Roo78, Rud74, Rüs91, Ryc87, SW90, SL80, Sau84, Sch78b, SX08, SB09, SV76, Ste80b, Tav78, Tav79, TW84b, Twe82, VPCG10, Van91, YC96, Ale13, BE21, BGZ16, BFW22, But22, CV18, Chi14, EKKZ18, GRR17, GJRS17, HHHT19, Iye23, KTR19, Nea14, NLM18, PRR17, Sch16, Sto15, SMAL17, Tam16b]. Disturbance [LW92]. Divergent [Sie76]. Diverse [GS77a, Hun95, ZK89]. Diversification [GPD00]. Diversity [DMB11, ALS18a]. Divided [CC99a]. Dividend [Pau07, WWY12, AAM17, CKLP20, EL14, Fro13, JMFPY19, ZY16]. dividends [BFP23, CL22]. Divisibility [Ano91e, Ott74, Ott75, PS90b, Ste73, Ste75b, Van78c]. Divisible [AAMRM05, Bon82, Cek03, Chi12, HS77, HL06, Mas79, Mil74, Roy07, ST10, Ver79, Wol78, Chi14]. Division [Cow89a, CC96a, CC96b, Cow96, Cow97, Cow10, Mec89, NB15]. Divisions [MNW08]. DNA [Bob04, DR85a, HMV06, JP00b, Lit69b]. Do [Sam04b]. documents [JS15]. Doeblin [Coh81]. Domain [Gec77, GR88, Koh78, MPS98]. Domains [CQZ03, Sch03b, BFW22, HS13]. Dominance [Vet81]. Dominating [KM00]. Domination [LMS00, Sha03, FFI22]. Dots [Ano06e, ACDHT05]. Double [Sza93, TF08, Ano09e, BCSY15]. Doubly [CCP91, Cha09b, MS80b, Pro76b, Rud72]. Dove [MMC91]. down [BMM16, WZ19, WZ20]. downgrading [FGRT17]. downtime [Bré75]. Drainage [PW10]. drains [Dio76, Dio78]. Draw [WZ20, WZ19]. Draw-down [WZ20, WZ19]. Drawdown [Sek06, RW14]. drawdowns [Zha15]. Drawings [Sta90a]. drawups [Zha15]. Drift [AS77, ACLM77, AML77, Ant78, Ant79, ACV80, Can74a, Can74c, Can75a, Can76, CS99, Fuh07, Haj82, HV83, Imh86, Jan86, MPS98, Ott76, VJV10, Wac09, Mas15, MZ20, DMYT23]. Drifts [MY03]. Driven [AAMRM05, BT08, BK05, Fas05, GRW10, KKSS08, Mal90, MS11b, Zaz92, Aaz13, CDIM13, MMZ22, NP17, DMYT23, QDZ19]. Driving [Pes82a, Als18b]. droites [Par98]. Drug [QG11]. Drugs [CG08]. Dual [Gat94, Gri77, TV10, BFP23, EL14]. Duality [CS12, Fox16, HM11, LH07, Lan05, Sch99b, SR00, Tho07, Lac20, WHL17]. Duals [Roe73, SJ18]. Dufresne [MY03]. d’un [Let82]. d’une [Bré75, Fri86]. Duplications [CK07]. duration [Tam13, Tam16a]. during [Coh72, Deh98]. Dynamic [AR06, All82, Ass78, BV93, Bra98, BY89, CG80, CL87, CM93, CM98, DZ11a, GV98, Gla88, Hak82, Ker78, Kos76, Law92, Mar16, Nic82, ÖB05, RR90b, Rie75, Ros74a, Sch73, TE96, VV78, XZ96, ADL15, TYA16]. Dynamical [Le 78, Mal93, BFW22, GM14, KXZ23, QD15, Qia18]. Dynamics [Cau74, Imh08, Led04, Ren86, Sch05, Sve93, RT21, XHW23].

Early [But94a, SC77]. EARMA [JL77, Jac80a]. Ecological [LC76]. Ecology [Kit71]. Economic [Dyn78, Nec78, MVY19]. economical [LR18]. Ecosystem [Rej82]. Eddy [Roe74]. Edge [BK02, DMR13, HP82, Kho01, TS98]. Edges [BLP06, Muc05, VGS10, CGR18, MP18, vL18]. EDITOR [CM09, CC05]. Editorship [Gan08]. EEG [Ger77]. Effect [Ave76, AH78, Ave79, Hil76a, Lan13a]. Effective [Hil81]. Effects [BK02, Car78, HP82, HKH76, Mat98, Whi82b, CAR20]. Efficiency [AR92, IS83, LJJ04, Pfe85b]. Efficient [DM05, OS18, PS94b, QDZ19, LL15, SMAL17]. Ehrenfest [Bin91, Blo89, Che00, FR12, Pal93, Pal94, VS87]. Eigenproblems [JH92b]. eigenvalue [YS23]. Eigenvalues [FVK08, Ger77]. Einstein [LO04]. Einstein-Type [LO04]. Elastic [GL04a]. election [KM14a]. Electric [Mal90, Pal93, Pal94]. Electrical [BN94]. Electricity [Gat85b, BNBV14]. Electromagnetic [HR04]. Elementary [MOK95, CHY20]. Elements [Ran80]. ellipsoid [Sch16]. elliptic [BvdMS20, LLX18]. Elliptical [HL02]. Embedded [EC75, Hun84, KW88, Neu86c, Ser75, Wij78]. embedding [ACH21]. Embeddings [Die06]. Emergence [Had79, Whi82b]. Emerging [BKO11]. Emigration [Pak86, Pak89a]. Emphasis [Gup80, Hui03a]. Empirical [Ais79, Bal97, CY99, FS93, Gol77, KS09, Pyk79, Eck19, MZ16]. Empty [Han96, LS11, SG92]. Encoding [BE88]. encounter [GY17]. encounter-mating [GY17]. ended [Cza23]. Endemic [Nås96, Nea14]. Energy [BM05, DH82, Hui03a, Lep76b, Sys75, Sys79, Sys80, BV13a, BV13b]. Engelen [Rog89]. Engineering [RT90, CKL14]. Engset [FR12]. Enhancement [TS98]. enlargement [BSHJ17]. ensemble [CDJR22]. Ensembles [Mat72]. Entrance [Pak93]. Entries [MS87]. Entropies [AFN74]. Entropy [Ano86e, Fra85, Hua90, KR81a, KR81b, Pap77, ARS17]. Entry [Kos76, Wij78, YY22]. Enumeration [CAK00, BP17, BP22]. Envelope [HP78]. Environment [BMY94, Ban09, Bar77, Bar79, Cha76, DH97, DMB11, FV99, Iva83, KM80, LM90, Tan81, Bau14, CV23, DS16, GM14, GLL22, RSY23]. Environmental [GG90, KX94]. Environments [Ave76, Bal79, BM04, BV06, BC81b, D’S94, Eco05, Fuj79, Fuj80, GJL84, HM81, Hor03, Kap73, Kle82, Lin74a, Smi74, Zir01, Hue16, KDO16, SK16]. ephemerally [DP22]. Epidemic [AD94, And99, BT71, Bal86, BC93, BO93, BB07a, BST09, BS12a, But94a, But94b, Cap76, Dan74b, Elv71, Fad77, Fad79, Gan79, HKM^{+}84, LP90, LP93, LP99, Mol72, MW74, O’N95, PL90, PL99, Sve93, VML80, Wat79a, WD71, Wic77, Wil71, Woo74, BN23, LS21]. Epidemics [Aba74, Ano12, Bal85, BO94, BL01, BKO11, BRB^{+}71, BJML07, Dan71, GL09, KR71b, MT11, Nag71, Nea06, VJV10, Wei71, BLT19, CL14, GC16, LG16, Nea14, Sim15]. Epidemiology [ABA^{+}80, Gan71, Kit71, Tau71]. Epistatic [Car78]. Epoch [YZ78]. Epochs [Ver73, Wac09, Als18b]. Equal [AMS07, BS22]. equal-input [BS22]. Equality [SW90]. Equation [ARMAG00, BU98, Bra86, BNR87, Cau74, CDL99, DS76a, FFH86, FT78, Gly94, Kar90, Leh02, Nar91b, Neu86b, Pes82b, RSS94, Sac90, Shi79, Spo84, Ver77, Ver79, Wol80b, BS15, FS16a, FPZ13, HRW18, LZ17, Liu97, MV19, NP17, Liu97]. Equations [Ahm80b, AAMRM05, Ath79, Beg12, Ber80, Boy78, CLM08, CT85, CR90, CEP10, De 76, DN02, GR08, GNRS89, Hor03, KLRM05, Le 74, LMX05, MC09, Met80, MT09, Nar91a, ØSZ11, Ors90, Rob74b, Sen69, Tör78, Wah76, WCL01, Aaz13, BGV22, BL22, BGL20, DJ17, GS14, LLX18, MT13, Qia18]. Equilibria [Bro05, Hin94b]. Equilibrium [Bar80, BM77, Col74, CT94, DS76b, Dyn78, Heu79, RH81, Saw77, Sha87, SR76, AB22, Att17, GRR17, Lin15, MR20, Vid20]. Equivalence [BZ03, CH14, LSH16]. Equivalent [HJK05, Sim80, WCK96]. Equivariant [Soo10]. Erdos [Rob18, Sho79]. Ergodic [Alp09, Dav79, FGM12, Gri75a, KR71a, Kem75, Kli80, KZ92, KZ94, Lin78, Max77, Ott76, Ran80, Sch77, Sys77, TT94, NR21, Oça23, PD22]. Ergodicity [BH04b, CK73, CT85, Cha89, Coh77b, Coh77c, Fal84a, FJKR20, HS92, KW90b, Las04, Sen79a, Sen79b, Sen88, TT79a, TT79b, Van85, APZ21, BDLP14, CKO19, JKR17]. Erlang [DP19, DL82, Di 01, GLM93, JVZ08, LBZ98, MR20]. Erlang-A [DP19]. Erlang-Distributed [Di 01]. Erlangian [Con83]. Error [BCJW14, Dia13, Hua02, LL18, Mas15, Nak96, Roo78, Tru97, AB15, CKL14]. Escape [BHKK15]. Escaping [SS09]. ESS [Hin94a, Hin94b]. Esseen [GZ11, Lan87]. Essential [Whi69]. Estimate [McK89, Nak79]. Estimated [Ric75]. Estimates [AAL08, CP08, DR85b, Eng84, FGM12, Mae79, Nyr05, Ran77, RS88, SW13, SRTM97, Wal74, ZG88]. Estimating [Hal80, Hey74b, IL97, Joy91, KM96, KM99, KOS03, KL00, LW92, Mas84b, Pic85a, San98, TS81, TSvB17, Wat79b, Wit79, ZG04, Bré75]. Estimation [ARS17, ACF09, AK78, BT71, BC01, Ber78c, BA81, Bré75, Bue88, CB03, CSY98, CRC04, DE08, Edm76, EF81, ES94, EC75, Fei76b, FL09, GL95, God85, Hah98, HS98, HKT91, HBG99, HV01, Hey75, Hil81, HJ02, IPV05, JP98, JK94, KPUU01, Kas80, Kei74a, Kid01, KJ03, Kid08, KS75, LJ04, LJJ04, Le 78, LM04, LL98, Man74, Mas95, Mas83, MT06a, Mol95a, Mrk01, ON96, ON98, PLRC08, PS09, Pro76b, Rat96a, Rat04, RH81, Ros81, Sch96, Sch00, Smi80, Sri85, Tho81b, Van06, Wal73, Wei95a, Wei78, WHZ10, ADP18, CM17, CDJR22, Gol16, HP23]. Estimator [Lie76, LY12, PS11, RS97, RS12]. Estimators [DR81, FR97, Han96, Hea78, Koh78, MS94, Pic00, RC94, Sch99a, Sva14]. état [Bré75]. Étude [Fri86]. Euclidean [DMPPG16, Kol14, Lee99b, Mat99, Max77, MP03, Mil69, Mil71b, OD07, Pit77a, Sto98, vL18]. Euler [AO95, Rat04, TO21, Wor94, Wor95]. Evaluation [BNR87, GNRS89, Gla91, KS83a]. Evaluations [ST96]. Evasion [Lim97]. Even [FS09a]. Event [AK06, BKM92, Cao95, CDL12, PT77, LL15]. Events [CK07, Coh79, God91, JH92a, Kei74c, Kwe75, SW93, EMOS14, PSTB17]. Everywhere [Dav73a]. Evidence [Wil74]. Evolution [Ahm80b, AAR08, Bob04, Bob08, Deh98, Esh72, Esh73, Hai75, JP00b, KR71b, Pes82a, Sch82b, Ton82, Vet81, BE21, TYA16]. Evolutionarily [Ben80, CG80, CM93]. Evolutionary [CC09, Imh08]. Evolutions [Blu84, LS13]. evolving [KKV16]. EVT [DE08]. EVT-Based [DE08]. Exact [AD21, AAL08, Ahm81, Ano93e, Ano03e, Ano06f, BPS19, BZ20, Buc13, BN92, CR08, CMB19, Chi12, Dai11, DN02, DEP91, GPHK21, Gri79a, Kel77, Ken85, Kne00, Kog89, LP99, Mar75, PR06, Rød06, TM91, VV01, MS16, PB19]. Example [Fad77]. Examples [Bat73a, DMY93, HM11, Pak95, Sys73, AD15, dlI22]. Exceedance [Alp09, FT12]. Exceedances [Roo88]. Excess [KLS02]. Excesses [RW03]. Excessive [VW77]. Exchange [Won08]. Exchange-Option [Won08]. Exchangeability [MS11a]. Exchangeability-Type [MS11a]. Exchangeable [BE83, Fou11, FM09, HJ08, Möh00, Möh04, Wat86a]. excitable [RR21]. Excitation [BM02, CGMT20]. Excited [Ors84, Bau14, Pin16]. Exciting [CS97, DP22]. exclusion [US19]. Excursion [AST10, Cao99, CH80, Tak91a, Ver78, Wor94, Wor95, Fou18, KN16, PW18]. Excursions [CJPY97, Gau02, Ill01]. Exhaustive [LZ95, VBvdM15]. Existence [BBD99, BW97, Der09, DMPG09, FMP92, Kar90, May78, Sar97, Ber20]. Exit [AT81, BDD12, DM00, Eme73, GM11, Jan86, Kei74b, KJH18, LS89, Rog90, Sal88, SW13, Sys76]. Exits [Ill98, KK92]. exogenously [Lem14]. expanding [HS13]. Expansion [DDSV87, Gri79b, Mei77, Van88b]. Expansions [Ano04g, BC89a, BG07, Cek97, FL01, Fuh07, HH03, Jan19, JVZ08, LP96, MM83, MM85, But22, HR16, JL15]. Expansive [BH00]. Expectation [CFW09, BMS18, Li23]. expectation-maximization [BMS18]. Expectations [AP98, Ker86, FLR22]. Expected [Ali90, ASC96, CFGW87, DGM09, DHP12, ENR17, Fro04, Gha08, HMM06, Jan84, KSC09, MP08, PR82, PP07, SC83, Sta92, VL98, Wor94, Wor95, CR19, CS17, DP13]. Experimentally [Sch82b]. Expert [BS03a]. Explanation [Hey78]. Explicit [Cal03b, DL20, KL87, Neu77c, SY97, Wad07, LR18, Yca84, dlI22]. explicite [Yca84]. Exploiting [Igl80]. Exploration [SK85]. Explosive [Pak93]. Explosiveness [Gre89a]. Exponent [BBR09, Cek97]. Exponential [APZ21, Bal90, BLL12, BLS88, Box83, Bri76, BP99, But22, CK73, CNP91, De 73a, DL82, DL11, DS00, Fac09, FS08, GI10, HZ07, HW09a, Hol76b, JL77, JKR17, KD94, LL81, LY15b, MY03, Sen89b, SSS04, Szc90, TT79a, TT79b, Van85, Van88b, Yor92, AZ18, BP23, GKPZ21, HHBV20, JV18, Lac20, Sch06b]. Exponentiality [Tak81]. exponents [AB15]. Expressed [AP88]. Expression [Cal03b]. Expressions [Whi95]. Extended [Als88, BZ22, Eck19, KS90, Pea89, Pes82b, CT18a, Cha22, KM14c]. Extension [Bal78, Gyi75, MS76, Pit78, RS92, Lac20, Liu97]. Extensions [DDH78, Dio78, McD76, RS86, Sho79, Wat79a]. External [BF05, DY15]. Extinction [Agr74a, Agr74b, Bro85, CLCZ12, ENPS78, Fuj79, Fuj80, GZ93, HLN13, Hel75, JKS07, Jir76, KK73, Li09, MP10, Nor82, Pak89b, Pak89a, Pol74, SS09, Sze76, Tho79, Cor14]. extra [MP18]. extraction [FY18]. Extraneous [Neu71]. Extrema [AA85, AA86, AA88b, CMB19]. Extremal [BS95, CHM91, HK89a, HT89, LN11, RR73a, RR73b, Res74a, Res74b, Res75, RR90b, Shi93, Sho74, Smi88, Ste02, CM19a, MS17]. Extreme [Ber87, BVH10, Coh82a, DR85b, DZ11b, Gal79a, GR88, Lin80, Mej77, PSTB17, Per97, Pic83, PC96, Rei81, Roo93, Smi82, Tak88, ZG04, EMOS14]. Extreme-Value [Coh82a, DR85b, GR88, Smi82]. Extremes [BNX02, Ber82, BMPR19, BFL09, CH16, Coh82b, DL18, Fas05, FS09b, Hal80, HL02, LT98, Pen98, Roo83, Seg03, ST98, BN20, NW22, Seg20].

Faced [McN90]. Faces [CC96a, CC96b, HS11, Sch09]. Facilities [Coh78, IL73, IL74]. factor [BDLP14, CD18, JKR17]. Factorisation [Dag75a, Dag75b, Gre75, Ric75]. Factorization [Die06, Pra73, XZ20]. factorizations [Thé22]. Factors [Cao90]. Failure [AA88a, AHO20, Arj80, Ave83, Bar85a, Ber78b, De 83b, FE06, HP82, LL97, LM03, LM04, Pho79, RG89a, SS84b, SW98, Tay79, Tie81, Tie82, GB17, HA20]. Failures [CAK00, GL09]. Fair [Che89a, PR22, KM14a]. Fairness [AIL04, GMS23]. Falling [CT94]. Falling-Leaves [CT94]. Familial [Wil79]. Families [BLL12, GRHK06, LS93, Wat92]. Family [Bob08, Büh74, Gla93c, Hil76a, LP90, Möh06, RS07, Ste76a, Ste79, Ste80a]. Fast [BDM99, DEKS11, Hös11, APZ21]. Fatigue [Sob87, Tie82]. Fatou [Pes82b]. FCFS [Bro05, CKW09, VBvdM15]. feasible [PV19]. Features [Ger77]. Featuring [BKO11, O’N95]. Fed [DS00, LM00]. Feed [SW11]. Feed-Forward [SW11]. Feedback [BM77, DKS84, FD83, Hun83a, Hun84, Hun85, Rei88]. Feedbacks [Hau78]. Feller [Bob08, Sac90]. Fermés [Mat72]. Ferromagnetic [Geo84]. few [MP18]. Feynman [GR08]. Fiber [PT73, Pho79, Tay79]. Fibers [GL04a, Tie81, Tie82]. Fibre [AS95, Hah98, Kid01, WN94]. Fibres [Dan89]. Fibrous [De 83b, HP82, KPUU01]. Field [Cha09a, CORR91, Iva80, Nar91b, Nos94, RSZM02, Swi75, Wil85, Wor01, AS13, Pia01, Sto23, VMM19, YS23].

Fields [Adl76, Adl77, AST10, Ber87, BY09, BMR05, Cao99, Dob77, Häg00, Kar86, KLRM05, KMW96, MW98, OB89, Pic80, Pic85b, Pre73, Rav83, Roz78, Sch05, Ser80, TP92, WS11, Whi92b, WA82, Wil88, Wor94, Wor95, BNBV14, DHL17, Fou18, Gal16, KN16, LL15, LX13]. FIFO [HK89a]. Fighter [BSC11]. File [LRS10, Bré75, Fri86, Fri87]. File-Sharing [LRS10]. Files [MY80]. filter [BCJ^{+}17, CDJR22, NLM18]. Filtering [CC12, CPR01, Hun83a, Hun83b, Hun84, Hun85, Kar85a, Kri79, MS80b, MT09, BJ77, CM17, JLY22, QD15, Qia18]. Filters [Cam88, HC08, LM03, LM04, CHJ16, JY20]. filtrage [BJ77]. filtrations [BSHJ17]. Final [And99, BC93, BST09, But94a, But94b, CM98, LP90, LP99, PL90, ST85]. Final-Size [ST85]. Finance [Gau02]. Financial [ACDP12, Duf04, CKL14, LY15a]. financing [ZY16]. Finds [SG92]. Finetti [Pit78]. finies [Yca84]. Finite [AE95, AT90, AK79, Ave76, Ave79, BB92, Bat73a, Bat73b, Bat73c, BR88, BKS12b, Cao89, Coh77c, Coh76, CM98, Cse92, Dsh85, EF80, EF81, Fad74, GJL84, God85, GW13, GHH15, HV84, HW09a, Hil74, Hil76b, Hol81, Hüb80, LT01, Mar70, MS90, Mil69, Min80, MRSV86, OD07, Ott74, Pak79, Roe73, Ros73, Sar98, Sch86b, Sha97, SS20, Sma76, Son79a, Son79b, Sta80, Ste75b, SW92, Tak91b, TT04, TW87a, TV78, Tru75a, Tru75b, Zwa00, DVN15, GLL22, HMM19, HHHT19, HLLP16, KR18, Mar16, YS23, Yca84, ZCJW19, ZCLJW23]. Finite- [GW13, Sch86b, TT04]. Finite-Alleles [BKS12b]. Finite-Capacity [Tak91b]. Finite-Dimensional [AE95]. Finite-Horizon [CM98, GHH15, Mar16, ZCJW19, ZCLJW23]. Finite-Stage [Hüb80]. Finite-State [AK79, Coh77c, YS23]. finite-state-space [GLL22]. Finite-Time [HW09a, MRSV86]. fire [GK19]. Firing [Fie75, Kry76]. First [AV19, BH04a, BR93, BOR02, BVH10, Big76, Bor09, BNR87, Cox80, Der83, DGN92, DNPR01, DW06, FL01, Fuh07, GL80, GNRS89, GNR90, Gut74, HBG99, HNGS15, Hol81, Jan86, Kes80, KK15, KW03b, KZ16, Lal84, LL97, LL89, Lef07, Leh02, Mat98, McD83, PL96, Pim07, ST96, Sal88, Sha84, SW77, SW78, SZ11, TF08, US19, Van83, Whi95, Wie77, Wis72, Yuh82, Ahl15, Ano09e, But22, DL20, DS15, Kub19, Lef88, Man21, MP18, STZ14]. First- [Big76, Bor09]. First-Crossing [Lef07]. First-Crossing-Time [DGN92]. first-in-first-match [Man21]. First-Order [BR93, GL80, Mat98]. First-Passage [BH04a, BOR02, Cox80, Der83, Jan86, Kes80, Lal84, McD83, Sha84, SW77, SW78, Van83, Whi95, Wie77, Yuh82, KZ16, Ahl15, But22, Lef88, MP18]. First-Passage-Time [BNR87, DNPR01, GNRS89, GNR90, ST96]. Fischer [Vet80]. Fisher [Bob08, BC04, CV23, GPHK21, Gri14, Ken77c, LC76, Vet81]. Fit [BK99, AC22, EY14]. fitnesses [Iye23]. Five [Ken95]. Five-Dimensional [Ken95]. Fixation [Hol76a, Hol81]. Fixed [AT75, Ber82, Cal03a, DN02, DMPG09, FN85, IJ02, IS06, JK94, Liu98, Nat80, Nat86, Pau07, Rob92, Tam08, AF18, CIL17, HRW18]. Fixed-Point [DN02, AF18]. Fixed-Size [DMPG09]. Flag [Tau75]. Flat [BL09a, Kid02, Kid08, Sch96, Sch00, Spo01]. Flats [HLW03, Kri77, Mil69, Mil71b, Sch99b]. Fleming [Gri14, Fou11, Hir00]. flexibility [ARW13]. Flexible [ADZ04, AR06, ARW11, HTO^{+}15, PWZ17, Sto15]. Flights [Tak84]. Flips [Soo10]. Flooding [Gan69]. Flow [Fei76a, HO98, Hun85, IM02, LS96, RSS99, TE96, Tur74, HJKP17]. Flows [AS07, Beu83, BZ03, KKSS90, TX10, ZÇ96]. Flowtime [Käm87]. Fluctuating [Cha76, DMB11, Pol78]. Fluctuation [Adl89, And74, Bin73a, Bin75, Bin91, CKLP20, GS77a, GP80a, GP80b, IPV05, Kei74c, Pal93, Sta78a, Tak76a, Tak76b]. Fluctuations [CORR91, GS77b, Kot87, Lep76b, Bow21, Jan18]. Fluid [Bäu02, DW06, DS00, FS09a, GY17, KD08, KKSS08, KM95, LW11, LA02, Mit88, Oso10, PTW10, RSS99, Zwa00, BOP21, DL18, GG17, Kin21, RFZ14]. Fluids [AB94]. FM [Hua97]. fMRI [Wor01]. Focusing [Pen02]. Fokker [Sac90]. Following [DR85a]. fonctionnelle [Liu97]. fonctions [Let82]. Force [CY05]. Forces [Pes82a]. Forecasting [Hal86, Kas78]. Forecasts [Coh85]. Forest [LT94, Myl02, VT94]. Forests [Dai08, HNS13]. Fork [BMS89, Dai11, KS04b, LF13]. Fork-Join [BMS89, Dai11, KS04b, LF13]. Form [BB92, BS12b, Coh82b, DTM07, Eco05, FS89a, Miy93, MT97, PMG05, PMG08, RT90, YM98, YS94, BZ22, But22, Mas23]. Formal [FG00]. Formation [KL83, YY22]. Forms [BV91, GO73]. Formula [BB93a, JVZ08, Jen95, Jen96a, Ott06, Rat96b, Vit95, Xu11]. Formulae [BKS12b, Cal03c, Chi94, LR97, Mil75, Buc13]. Formulas [And79, Ben95, CTR00, Neu76b, Neu77c, Wei98, DS16]. fortement [Cap86]. Fortes [Dio78, Dio76]. Forward [BMS18, Fly18, Rus10, SW11, VK78, Aaz13, FPZ13, GLL20, WZ19]. forward-backward [GLL20]. Forward-reverse [BMS18]. Forwards [AK79, QG11, BV13b]. Foschini [WN12]. Foster [MT93b]. Four [CC99b, LOS03]. Fourth [Bad77a]. Foward [SW80]. Fractal [Cro07, VJ99, KW20]. Fractals [HJ03, HR00]. Fraction [DSS99, Hen87, Sch99a]. Fractional [ARMAG00, AAMRM05, Beg12, BK05, FS09a, GRW10, KLRM05, Ken77c, KRS07, KMP06, LW11, MC09, Taq75, BM16, BD22, BPS15, KM15, MZ20]. Fractional-Dimensional [Ken77c]. Fractions [BR83, FG00, Mån07, Pea89, Woe85]. Fragment [DR85a]. Fragmentation [BM05]. Fragmentations [Duc21]. Frailty [GAH03, GB17]. Framework [CTR00, BZ22]. Fraud [Gof19]. Fréchet [KL00, KL03, LK00, Le01, MY20]. Free [BM96, Dad82, Dwa74, GHT97, LYA07, SW80, Van73b, Wei76, AV19, GC16, vdHvdHLS20]. French [Bré75, BJ77, Cap86, Dio76, Dio78, DM90, FG81, Fra89, Fri86, Fri87, Gal84, Lef88, Let82, Liu97, Mat72, Mat74, Par98, Pia01, PS83, Tau75, Yca84]. Frequencies [Eth90a, Gar80, TS81, ZG88]. Frequency [FL09, Hin94a, Joy91, KOS03, Rob76a, Ruk02, Fas14]. Frequently [KS78]. frog [GNR17, Kub19]. Front [Ano69c, Ano69d, Ano70c, Ano70d, Ano71c, Ano71d, Ano72d, Ano72e, Ano72f, Ano73d, Ano73e, Ano73f, Ano74e, Ano74f, Ano74g, Ano74h, Ano75e, Ano75f, Ano75g, Ano75h, Ano76f, Ano76g, Ano76h, Ano76i, Ano77f, Ano77g, Ano77h, Ano77i, Ano78f, Ano78g, Ano78h, Ano78i, Ano79g, Ano79h, Ano79i, Ano79j, Ano80f, Ano80g, Ano80h, Ano80i, Ano81e, Ano81f, Ano81g, Ano81h, Ano82e, Ano82f, Ano82g, Ano82h, Ano83g, Ano83h, Ano83i, Ano83j, Ano84f, Ano84g, Ano84h, Ano84i, Ano85f, Ano85g, Ano85h, Ano85i, Ano86g, Ano86h, Ano86i, Ano86j, Ano87e, Ano87f, Ano87g, Ano87h, Ano88f, Ano88g, Ano88h, Ano88i, Ano89f, Ano89g, Ano89h, Ano89i, Ano90e, Ano90f, Ano90g, Ano90h, Ano91f, Ano91g, Ano91h, Ano91i, Ano92f, Ano92g, Ano92h, Ano92i]. Front [Ano93h, Ano93i, Ano93j, Ano93k, Ano94f, Ano94g, Ano94h, Ano94i, Ano95f, Ano95g, Ano95h, Ano95i, Ano96e, Ano96f, Ano96g, Ano96h, Ano97e, Ano97f, Ano97g, Ano97h, Ano98g, Ano98h, Ano98i, Ano98j, Ano99e, Ano99f, Ano99g, Ano99h, Ano00e, Ano00f, Ano00g, Ano00h, Ano01e, Ano01f, Ano01g, Ano01h, Ano02e, Ano02f, Ano02g, Ano02h, Ano03f, Ano03g, Ano03h, Ano03i, Ano04h, Ano04i, Ano04j, Ano04k, Ano05e, Ano05f, Ano05g, Ano05h, Ano06g, Ano06h, Ano06i, Ano06j, Ano07f, Ano07g, Ano07h, Ano07i, Ano08e, Ano08f, Ano08g, Ano08h, Ano09f, Ano09g, Ano09h, Ano09i, Ano10e, Ano10f, Ano10g, Ano10h, Ano17b, Ano17d, Ano17f, Ano17h, Ano18b, Ano18d, Ano18f, Ano18h, Ano19b, Ano19d, Ano19f, Ano19h, Ano20b, Ano20d, Ano20f]. Front [Ano20h, Ano21b, Ano21d, Ano21f, Ano21h, Ano22b, Ano22d, Ano22f, Ano22h, Ano23b, Ano23d, Ano23f, Lin10]. Frost [LP90, PL90]. Froths [CK99]. Fuel [SW92]. Fulfilling [CFPL12]. Full [BCM13, Pet82, XHW23, Tam16a]. Full- [BCM13]. Full-Information [Pet82, Tam16a]. Fully [GT12, LT00]. Function [AR92, Bal07, BV97, CB03, CT85, DP84, Ell87, FV95, Gec77, Han96, Han07, HO09, HS10, Hun77, Jen84, Kem75, Ker78, LG05, Off74, Ott76, Ott77a, Ott77b, PP07, Ros75, San98, TS99, DM22, JKM23, Kol14]. Functional [AK95, AD94, Ano98e, Bar74a, CM98, DM00, FFH86, FP96, GW20, Igl71, JLW05a, JLW05b, Mec77, Qi08, Rah07, RSS94, Sen69, Ser75, TO21, VS80, KWRK22, KN16, Liu97, SS14]. Functionals [Ahm80a, AP98, AO95, ACF09, BKZ95, DMY93, DEP91, IL97, Jac74a, Lev77, Lev78, Lin74b, MY03, MZ98, PSS06, PR99, Sch06a, Sch06b, Sch07, Seg03, Ser89, Yor92, Buc13, DHL17, JV18, KW20, LPST14]. Functions [AM90, AB07, BGS89, CB95, CTS78, Cao98, DK97, Ebr90, ENM89, EGV79, Gne99, Gne00, GSS01, GGV89, HK22, Kei76, Koh80a, Koh80b, LS11, LT01, Led04, Lin80, LW92, Luk72, Mat73, Muc10, Mül97a, NR85, OT02, PT77, RX91, RS88, Sak04, Sav75a, Sav75b, Sen74, SP80, Stu78, SM87, Van84, Van78c, VW77, Wei98, Wol78, Ath18, AV19, BOS23, CRZ18, ENR17, JL15, Let82, LY15a, LY19, LX13, Yin23]. Fundamental [Arj72a, Arj72b, AL99, EL05, LC76, Vet80, Wis72]. Further [CSH^{+}80, Heu80, JK85, Kij92, MM96b, MW72, Pra73, Sys88, Tjø83, Can75b]. future [FZ13]. futures [BNBV14]. Fuzzy [Fér77].

G [Ali90, Asm82, BW90a, Ber78c, Box78, BK01, Cha92b, Coh74, Coh77a, CH80, CK15, DT80, Dos78, EW14, FS89a, Fal84b, Fri86, Fri87, GHT96, GHT97, Gal79b, Gly82, Gly94, Gol16, Gri94, Hal85b, HP96, Hau78, Hok79, Hok85, Igl71, JS83, KS83b, Köl81, Kuc83, Kyp71, LZ95, LZ05a, LZ05b, MA95a, Mil95, Min80, MZ04, Naw91, Neu76a, Neu77b, Neu86c, Num79, Ott77a, Ott77b, Ott87a, PR06, PZ16, Ram84, Ram80, RW91, Ros73, Ros76a, Sat85, Sch84, Sch90, Sen89a, SW91, ST90, TH90, TV10, TV78, TVF81, Tru75a, Tru75b, Van78b, Whi82a, WW03, YC96, BW90a, BC89b, GHT96, GHT97, HK89a, JD89, Ott87b, PZ16, RW91, Sen89a, SW91, SP80]. G. [Tra74]. G/G/ [PZ16]. G/G/1 [BW90a, Sen89a]. G/G/c [SW91]. G/G/m/ [RW91]. G/IFR/1 [HK89a]. G/M/1 [GHT96, GHT97, SP80]. GA [Kuc83]. GA/G/ [Kuc83]. Gabriel [BBD02, DGM09]. Gain [Aas82]. gains [BFP23]. Galam [LT15]. Galaxies [MPB96]. Galton [SS09, SS15, BD74b, BD74a, BM08, D’S94, Dou75, FQVJ71, FV99, ILP18, Jir76, KM10a, KRS07, KXZ23, Lin74a, Lin74b, Min05, Qui72, SJ18, Sen69, Sri85, Sze76, Ven82, Wau81]. gambler [PR22, Sti06]. Gambling [HS74, Pes82b, Ros74a]. Game [Dro89, EL05, EX11, Hai75, Imh08, KPO09, Lim97, MMC91, MS01b, MS04, AH16, DK16, GY16, GML18, HMM19, Luo16]. Games [And75, CC09, Fed78, FW81, GHL07, Phi78, AB22, Att17, DF18, FPP19, Lin15, Mon17]. Gametic [Esh73]. Gametophytically [May78]. Gametophytically-Determined [May78]. Gamma [AAR08, BL09a, Bri99, CQZ03, GL80, Let85, MZ96, Wen89, Wes90, Zuy99, But22, MMZ22]. Gamma-Distributed [AAR08, CQZ03]. Gamma-Type [MZ96, Zuy99]. Gani [Fei76a]. Gani-Type [Fei76a]. gap [LS15]. Gaps [Hol89]. GARCH [Ano07e, Cli07, Cli11, Rud98]. Gating [BD97, Bal97]. Gauss [DNPR01, IM02, MW72]. Gaussian [Abr79, AAMRM05, Ano98f, BN94, BNK98, BV13b, Bow21, CP08, CS99, Dan89, Dav73b, Dav73a, DHL17, DL18, DR16, GLM22, Gly82, HTO^{+}15, HR05b, Ill01, JVZ08, KZ20, KM12, Koh78, Kom96, KN16, LL15, Lin72, Lin80, LX13, Mah10, Mer06, OS18, PR99, PLPA96, Sak15, Sel96, Sha79, TP92, Vit01, Whi81b, WA82]. Gegenbauer [Gal84]. Gelfand [Let82]. Gene [Rob76a, Ros74b, Tho75, Tho78, Tho81a, Tho82]. Genealogical [Kur18, Llo73, Tho76, AD21]. Genealogies [BSS17]. Genealogy [Edw79, O’C95, STW84, Kal18]. General [Ano98f, Ano13a, Ano14a, Ano15a, AN78a, AN78b, Ave83, BT71, BO93, BW93, BFM^{+}84, Bat73c, Ber78b, BV93, BKS12b, BGR06, BL08b, BG93, Box83, BM02, BRT82, Cap76, CPZC05, CTR00, Coh78, CEP10, Dag75a, Dag75b, Dag78, DFT77a, DFT77b, De 73b, DM91, ES94, EX11, Fad77, FE06, GV98, GLS11, GW13, HS91, Häg00, HJK05, Han07, Här81, HR78, Her73, Her75, Hös11, JMJ01, KP10, Kas80, Ker78, Kom96, LMX05, LG05, Liu90, McD81, McD83, MH97, Nak96, Nat75, Num79, Nyr98, Oso10, PT73, QS00, RX91, Rit94, Ros81, RS88, Ser05, SS84b, Sha97, SY81, Teo82, TT79a, TT79b, TT79c, Twe76, Twe77, VL07, Wil71, Cas16, FHJK20, GB17, HLG18, Nea14, NLM18, Sto17, WZ19, ZCJW19, ZCLJW23, ZY16, BLT19, Coh77b]. General-Purpose [Ros81]. Generalisation [Kas80]. generalisations [FPS22]. Generalised [BC93, Bat76a, BC78, CC77, Con77, Con83, JH92b, MT05, Ros73, IMS21]. Generalising [Adl76, Adl77]. Generalization [BB90, EDH90, Fei76a, Hol76b, McL75, Nai86, WD71]. Generalizations [Baz74, Dav70, Dvo77, KSC09, Neu86b, Sch16]. Generalized [BS81, Bar01, BMV03, Bri99, Cao95, Cra17, DL82, Eba78, Eco05, Fou11, HLW02, Jen96a, KM15, Lem74, LZ95, Li09, Liu98, Mah10, MS12, MC09, Mei77, MSS95, Pre73, RW03, Ric77, Ris91, Sch76a, Sch78b, Won08, Cha14, Cha22, CDIM13, Hir17, Hue16, KXZ23]. Generalizing [Lor79]. Generated [AS07, Har71, Här81, JL79, KL87, KT92, Mül97b, Rud73, Sha78a, Sha78b, SS88b]. Generating [DM06, Mec77, Mül97a, PS94a, Tjø76a, SS14]. Generation [AO95, Edl78, Fil74, Møl95b, vdHK04]. Generation-Dependent [Edl78]. Generations [Cha74, Hil75]. Generators [Gzy80, MT06b]. Genes [Hil74, Hil81]. Genetic [ACLM77, ACV80, AH78, Can74a, Can74c, Cer98, Dom07, Don81, HMV06, HKM^{+}84, HKH76, Hol78, Hol80, RT07, Ros79, Ros81, Sen78, Smi76, Wat78, WB81, Cer17]. Genetical [Edw74]. Genetics [ABA^{+}80, Can74b, Can75b, EN80, EN88, HT10, Kin79, Mar80, Möh98b, NM79, Oka79, RL77, Ros76b, Sat78, Saw76, Sen86, SS84a, Stu78, Tav76, Tav78]. Genic [Wat79b]. Genome [SW97]. Genotype [HH81]. Geodesics [Ken95]. Geographically [Saw77]. GEOM [EW14]. Geometric [ALS08, Ano91e, ACO98, BDM85, BS95, BG07, BLS88, Cal03c, Cha89, DL82, DMPG09, Duf01, FRW12, GKL21, Gat85a, GI10, HT92, HNS13, HS82, IY12, Jor10, KW20, Le95, MY99, McK86, MT97, Neu78, Pen02, Phi78, PS90b, Sch03c, SF79, SK04, SV76, Twe82, Urb77, YM98, CYG14, DMPPG16, GP15, Gan15, HNGS15, JW15, LM17, Mas15, OY22, PZ17, XY15]. Geometrical [Bad77a, Bad77b, BM11, GM91, Le89a, Lit74, Mor69]. Geometrically [GT78, NR21]. Geometry [Amb77a, Amb77b, Ano13b, Ano14b, Ano15b, AS77, ACLM77, AML77, Ant78, Ant79, ACV80, DRS11, Rip77, Sch03b, GKT22, TL20]. Georges [Jeu00]. geostatistical [AT18]. Gerber [LG05, PP07]. Germ [DSS99, HM96, HM99, LS98a, LS11, PSS06, KL13]. Germ-Grain [HM96, HM99, LS98a, PSS06]. Germination [CQ01]. Germination-Growth [CQ01]. GI [Ano04e, Fer93, KW90b, Ali90, Asm82, Ber78c, Cha92b, Coh74, DN78, De 83a, Fri86, Fri87, Gly82, Gri94, Hau78, Hok85, Igl71, Köl81, Kyp71, LZ05a, LZ05b, Min80, MZ04, Naw91, Num79, Ott87b, Ott87a, PR06, PR00, Sen89b, ST90, Whi82a, WW03, YC96, BB10, TV10]. GI/G [Ott87b]. GI/G/ [Gly82, Whi82a]. GI/G/1 [Ali90, Asm82, Ber78c, Coh74, Fri86, Fri87, Gri94, Hau78, Igl71, Köl81, Kyp71, LZ05a, LZ05b, Min80, MZ04, Num79, Ott87a, PR06, ST90, WW03, YC96]. GI/G/1-Type [MZ04]. GI/G/1/0 [Naw91]. GI/G/1/N [Cha92b].

GI/G/s [Hok85]. GI/H [De 83a]. GI/M/1 [DN78, TV10]. GI/M/s [De 83a]. GI/PH/1 [Sen89b]. GI/PH/N [PR00]. GI/PHA/V [Ano04e]. Giant [LC92, Ber20]. Gibbs [BBD99, Jan19, KT92, Mas84a, MM98, MJM21, NZ77, Pre73, Pre77, SS14]. Gibbsian [Dob77, XY15]. Gilbert [AT18, BCM13, BC16]. GINAR [Lat97]. Ginibre [MS14]. Gittins [WG91]. Given [HS11, Kar86, Min05, PvS21, SC83, BN23, CM19a, KTR19]. giving [Tam16b]. Glivenko [EPS79]. Glivenko-Cantelli [EPS79]. Global [Hor02, Ken88b, Yin00, BMZ14, JJL21]. Goal [BF80a, BF80b]. Goals [Bro99]. Goldie [CY99]. Goldstein [Hol76b, SO85]. Gontcharoff [LP99]. Good [Hüb80]. goodness [AC22]. goodness-of-fit [AC22]. Goofspiel [Dro89]. Gordon [Dad82, Dad84]. gossip [BR13]. Governed [GTY70, NPOTC00, Ors90, Ste84]. Grade [Bar77, Bar79]. Gradient [EF81, GL95, Hah98, HS98, HP23, Kar21]. gradual [GKPZ21]. gradual-impulse [GKPZ21]. Grain [Deh98, HM96, HM99, LS98a, LS11, PSS06, KL13]. Grained [Sch03b]. Grains [DSS99, Der09, MR02]. Grand [Pre77]. Graph [AR97a, AR97b, ACPP12, BTZ00, BBHM11, BBD02, CGH^{+}99, CM03, CC96a, CC96b, Cro10, FRW12, Gri75a, HJ04, LR12, Naw91, NR06, AZ18, DMPPG16, Kin21, Rob18, TSvB17, XZ20]. Graphs [ACS93, BBSW05, BJML07, CT08, DM06, DGM09, DMPG09, GI10, Häg00, HNS13, IY12, Jor10, LC92, McD81, Now88, Sha09, Tak84, VJV10, Wad07, Was78, Whi92b, AH16, Ahl15, CYG14, CKM19, DMPPG16, Gan15, HNGS15, JW15, KWRK22, KK15, LSV18, LM17, MS19, OY22, PRR17, Ros13, vL18, vdHKV21, vdHvdHM23]. Grassmann [Mar77]. greedy [GMS23, HG15]. Grids [Gat85b, Tho07, LM17]. ground [BCLF16]. Group [FS89a, Ger84b, Lom78, Pak86, PS94a, Sch74b, SM78, Sta90a]. Group-Arrival [FS89a]. Group-Departure [FS89a]. Group-Valued [PS94a]. Grouping [Mor76, DFZ19]. Groups [Ott75]. Growth [BPSZ06, CY00, CQ01, Coh77c, D’S94, Deh98, Dei03, DH04, DMB11, FS09a, GM14, Gat97, GG12b, GM98, HM10, JN84, KK73, KR71a, Kes70a, Kes70b, Kes72a, Kes72b, KS04a, KK03, LMX05, Mel87, Mol74, QS00, Rob07, SK96, Sch94, Sma76, Tam08, Wil74, Ada16, Fox16, LY15b, MV19, Nyr16, Pri22]. Growth-Collapse [BPSZ06, Rob07, Pri22]. GSMP [Miy93]. Guidelines [Pes75]. Gumbel [MY20].

H [De 83a]. habit [YY22]. hail [BCLF16]. Half [BCM13, Ken74, BC16]. Half-Gilbert [BCM13, BC16]. Half-Line [Ken74]. Halfin [Ano04e, AJV13, GM08, PR00]. Halflines [Hoo89]. Hameed [Ano95e]. Hamiltonian [NG77]. Hammersley [Van83]. Handed [Sti06]. Handling [Kos78, SE91]. Haploid [Can75a, Can75b, STW84, Can74b]. Hard [KL13, SKF80, HT17, MJM21]. Hard-core [KL13, SKF80]. hard-sphere [HT17, MJM21]. Hardy [Cha74, MS74, RH81, SR76]. Harmonic [Pia06, Sav75a, Sav75b, Wal73]. Harris [MY89, Sch82a]. Hartley [AFN74].

Hastings [JRR14]. Hausdorff [BHB98, Cro07, HTO^{+}15]. Having [Phi78, SM87]. Hawk [MMC91]. Hawkes [Bou16, BM02, CMT21, CGMT20, CHY20, HL17, MR05]. Hazard [AK93, BLL12, LS11, SS87]. Hazards [Tie83]. ’Head [DN78]. Headstarts [Yas85]. Heat [ARMAG00, KLRM05]. heaviest [BMPR19]. Heavy [AR92, AK06, BGZ95, BZ03, CDL12, Dag78, DS07, EKKZ18, FQVJ71, Fal89, FR97, GPD00, GW91, Gri94, GW13, HR89, Har78, HW09a, IW70a, IW70b, Jac80a, JT13, Ken72a, Ken72b, Köl81, Kus94, Kyp71, LZ05a, MM85, OC10, Ott77b, PY13, Rat75, Rei88, Sch87, Sta78a, Ste02, Szc90, Szc99, VL98, VBvdM15, Whi70, WH80, Whi82a, BPS15, CRZ18, He16, KS16, LF13, MZ16]. Heavy-Tailed [BZ03, CDL12, DS07, GPD00, GW13, HW09a, JT13, LZ05a, Ste02, EKKZ18, CRZ18, He16, KS16, MZ16]. Heavy-Traffic [GW91, WH80, Whi82a, PY13, VBvdM15]. Heavy-Usage [MM85]. Hedging [BH98, CS12, KP10]. Height [Fuh07, Gau02, Sch06a, Sch06b]. Height- [Gau02]. hemisphere [CM09]. Hendricks [Die82a]. Heredity [Ser86]. Hermite [TM91]. Hessenberg [Mas23]. Heston [AE08]. Heterogeneity [Sta80]. Heterogeneous [BMV03, But94a, But94b, LM00, SS95a, Sad95, Web92, Miy17, SN16, Sto17]. Heterozygosity [HA79, Str74]. Heterozygotes [Hol76a, Smi79]. Heuristic [Mil95, Sto78, ZCJW19, ZCLJW23]. heuristics [HG15]. Heyde [HLZ18]. Hidden [AE95, DMR13, FS82, HR05a, HK22, ZG88]. Hierarchical [CL00, HJ04, Kel12, OT02, LT15]. High [AS08, BB95, DE08, FL09, JLW05b, Muc10, NR85, Sar98, Sch03b, BB17, BCJW14, BCJ^{+}17, CDJR22, Fas14, MV19]. High-Density [Sch03b]. High-Dimensional [BB95, NR85, BCJ^{+}17, CDJR22]. High-Frequency [FL09]. Higher [Gar95, KP86, KP94]. Higher-Dimensional [Gar95]. Highly [Nak96]. Hilbert [Ahm80b, Van74, Ver75]. Hill [RS97]. Hindsight [DL11]. Hindy [FLR22]. Histories [Arj80, DG04a, DG04b]. history [Pin16]. Hitchhiking [Leo09]. Hitting [BB07b, FPZ13, Gha08, Haj82, LL89, MY01a, Sal88, TF08, Zuy99, Ano09e, STZ14, JLW05b]. Hitting-Time [Haj82]. HJMM [Rus10]. HNBUE [POGP95]. Hoare [GR96]. Hoc [DGP10]. Hölder [DMYT23]. Holes [MRV02]. Holme [ORS18]. Homecomings [Kin73]. Homogeneous [Ano98f, Coh76, Coh77d, Coh81, GNR87, Gre75, HRY93, HQ02, Hor88, Jir76, Kom96, Kot87, MM96a, Mat77, MNW08, Mil71b, Swi75, Van74, Woo74, Zir01, DHL17, JS14]. Homogenization [SCA00]. homology [DFRV14]. Homothetic [BN12]. Honest [Bar78a]. Honeycomb [Wie81]. Honour [Mon95]. Hopf [Dag75a, Dag75b, Gre75, Pra73]. Hopfield [Pai01]. Horizon [CM98, Nos94, FPP19, GHH15, HMM19, Mar16, Tam13, Tam16a, ZCJW19, ZCLJW23]. Horizona [Tam11]. Horse [BL06]. host [AG13]. host-parasite [AG13]. hostile [GM14]. Hotel [LY77]. Hotelling [AB22]. Household [Ano12, BST09, BS12a, Nea06]. Households [BL01]. Huang [FLR22]. Hull [AD93, BHB98, Bro89, EG81, Hue94, Hue99, Mas00]. Hulls [Cow07, FH04, Gro12, Mol93, Buc13, CM09]. Human [HKH76, Hio79, Ste76a, Stu78, DAG95]. Hunting [Sud99]. Hybrid [KKSS90, PTW10, Pic98]. Hydrology [DGB^{+}84]. Hyperbolic [Iso00, OY22, TC13]. hypercube [Nes17]. hypergraphs [RM21]. Hypergroups [Voi90]. Hyperplane [HS11, Mec96, Mec98a, MP03, Sch09, HH14, HS14]. Hyperplanes [Sch03a, Mat74]. Hyperplans [Mat74]. Hyperplasia [WB71]. Hypersurfaces [Hof07a]. hypo [BvdMS20]. hypo-elliptic [BvdMS20]. Hypothesis [Dav77, Shi93].

I.D. [Hüs89]. i.i.d [ASC96, BR91, Ito79, Ito80, KK92, Mij80, Shi93]. I.I.D. [FJ09, Yu08]. Identical [Box79a, Box79b, RX91]. Identically [Ais79, BCL10, DM06, GO73, Rie80, Tak70, HM20]. Identifiability [AAR08, BL04, CH14]. Identification [Fou18, Moo79]. Identifying [HCD94]. Identities [Cow07, Gre75, GP80a, GP80b, Tho75, CKLP20, Pri15]. Identity [Arj72a, Arj72b, Kry76, Saw77, SB09, Tho78]. Idiosyncratic [PS10]. Idle [CR04, WW03]. IFR [HK89a, RX91, WC90]. Iglehart [Ber85]. Ignatov [Rog89]. II [LP96, AW87b, ACLM77, AK98b, BP79, Bat73b, BF80b, BD75, BC81a, Box79b, CT18b, Can75b, Coh77b, Dag75b, DG04b, DFT77b, DG04c, EN88, Her75, Hin94b, Hun83b, IW70b, IL74, JLW05b, KR81b, KR71a, Ken72b, Kes70b, Kes72b, Koh80b, Le89b, Lee99b, Man79, MT93a, Mil71b, Pic77, Pro76b, Sad95, Sch06b, Son79b, Szc86b, Tjø83, Whi80c, Wie77]. III [LP99, AML77, Bat73c, Coh77c, HW09b, Hun84, MT93b, SW78, Whi70, Whi80d]. Illustration [Coh77c]. Image [ÅH99, BE88, BS78, HK90, Mon95, SD98]. Images [BE06, Gla96a, HCD94]. imbalance [LF13]. Imbedding [BPS80]. Imbeddings [AN78b]. Imitation [CC09]. Immediate [HP83]. Immigration [Aks75, Ano79e, Ban09, BP79, BA81, BGR82, Fou11, GT84, Her73, Her75, IPV05, Lev79, Li98, MVY84, MY89, Pak71a, Pak79, Pak86, Rah07, Ric11, SVW83, Ven82, BPP21, Bau14, FMY22, MYH18]. Immigration-Branching [Her73, Her75]. Immunisation [Aba74, Wic77]. Impact [EX11]. Impatience [Sta90b]. Impatient [BBH84, MA95a, Thö99]. Imperfect [CD79, RG89b]. Implications [Kit71]. Implicit [JOC12]. Implied [Haj82, Hun89, JL15]. Importance [AHO96, BL08a, CDL12, CRZ18, DG04a, DG04b, GKT21, GJS08, LN92, LY12, DJ17, MJM21]. importance-sampling-based [MJM21]. imposing [DK16]. Impossible [Tho78, Tho81a]. Improved [AK06, GS93, Lep76a, Nat80, Nat86, RS07, Sat85, XZ20]. improvement [CR19]. Improving [vdHK04]. Impulse [Ano86e, CS17, Fra85, GKPZ21]. Impulsive [DP15]. In-Degrees [HJ08]. In-Forest [LT94]. Inbeddability [Ott75]. Inbreeding [Fie76a, Fie76b, Fie80, Hol80, MS74, RH81]. Incidence [DK97]. Inclined [Ros89]. including [Kem75]. Incompatibility [Heu79, Heu80, May78]. Incomplete [BH98, Pri99, Rav83]. incompressible [MT13]. Incorporating [BS88, BMY91, Bal97]. Incorrect [FS79]. Increase [Ben75b, DGR02, Pit77a, HHHT19]. Increasing [BLL12, BOR02, Bro05, LMX05]. Increments [AK05, Ber78a, Dav73a, Pic74a, Pic74b, Sco73, Sel96]. Independence [BN10, Che89b, Hüs89, MR09, Rem00, Tuc71, ŠSKB13]. Independent [AG81, Bac00, DM06, Gås12, GW13, Jac74a, Ker86, LR80, McD78, Ott87b, RW85, Sch74a, Tak70, Tam11, Tuc71, ACH21, MY20]. Index [Agr95, Ano89e, Ano13a, Ano13b, Ano14a, Ano14b, Ano15a, Ano15b, BK98, CHM91, DPY08, DG01, Fre82, GNMA02, Llo73, Smi88, WW91, AM22, HG15]. Index-Based [DG01]. Indexability [GHK13, NM01, AM22]. Indexable [GRHK06]. Indexed [Jof75, SW98]. Indicators [HS90, PP02]. Indices [Epl86, WG91]. Indifference [FS08]. Individual [Lin06]. Individuals [BS12a, Cha99a, Cha99b, DWH^{+}99, GB03, GP79, SR82]. Induced [JT13, LB10, Nie98, Nie99, Sch03a, SW93]. Induction [Jen96a, QG11]. Inequalities [HR04, HS90, KR81a, KR81b, Lep76a, Mec95, Mec96, Mec98a, RS87, SS77, WCL01, FS16b, KT20]. Inequality [CN86, Sen86, Tur74, Wat86a, Yu08]. Infection [Gan15, Fox14, LG16]. infectious [Nea14]. Infective [Gar71]. Infectives [Bal86, But94a, But94b]. Infector [BB07a, BLT19]. Infector-Dependent [BB07a]. Inference [Ano03e, Bas79, Bro73, Dav73b, Dav79, FM95, FS79, Gla96a, Gor01, Hal85b, Hey85, HW09b, Kar85b, Kar86, Kei78, Kün83, KP11, Leh85, Par77, Pic77, Pic85b, Qia90, TJM98, Tho76, VV01, Woe07, EW14, EMOS14, PC21]. Infinite [Ano91e, AT81, BP79, BL10, CKW09, DSSW11, Dur79, Dur80, GT12, GW91, GT08, Gri79a, Gri82, GP88, GW13, Jac74a, KS74a, Kha86, KS12, MY01a, MW78, MRV02, Mij80, MY01b, Ott74, Ott75, Pak10, PS90b, PS10, Sch86b, Sch87, SB77, Sch78c, Sch79, Sch76b, Sen80, Shi82, Sta78b, Ste73, Ste75b, TT04, Van78c, Wol80a, ZLB98, BCT13, BD15, HHHT19, HL17, MT16, Mas23, PSS20, SS20]. Infinite-Alleles [Gri82, GP88]. Infinite-Allelic [Shi82]. Infinite-Capacity [Sch76b]. Infinite-Dimensional [Sch86b]. Infinite-Server [GW91, Sch87, MT16]. Infinite-State [GT08]. Infinite-Time [GW13, TT04]. Infinitely [AAMRM05, Bon82, Cek03, Chi12, EK81, Eth90b, HS77, HL06, Mas79, Mil74, Roy07, ST10, Ver79, Wol78, Chi14, HLN13]. Infinitely-Many-Neutral-Alleles [EK81, Eth90b]. Infinitesimal [Gzy80]. infinity [BMM16]. Influence [CGH^{+}99, DR08, GW78, Wes79]. Influenced [HA79]. Influenza [BRB^{+}71]. Information [Ano71f, Ano72g, Ano73h, Ano74k, Ano75j, Ano76l, Ano77k, Ano78k, Ano79k, Ano80j, Ano81i, Ano82i, Ano83k, Ano84k, Ano85j, Ano86k, Ano87i, Ano88j, Ano89j, Ano90i, Ano91j, Ano92j, Ano93l, Ano94j, Ano95j, Ano96i, Ano97i, Ano98k, Ano99i, Ano00i, Ano01i, Ano02i, Ano03j, Ano04l, Ano05i, Ano06k, Ano06l, Ano07j, Ano07k, Ano08i, Ano08j, BK91, Bha86, Cur74, Don81, Eba78, EX11, Ger77, JOC10, Kit71, Pet82, RK89, SE91, ACH21, ADL15, BSHJ17, HS18, Tam13, Tam16a]. Information-Handling [SE91]. Informative [Neu86c]. Ingersoll [MS15, MS16]. Inheritance [CR78, Odo80, Ste76b, Sti80]. inhibition [CGMT20]. Inhomogeneous [BM08, HJVN03, HLW02, LR12, Pia06, RC94, VV06, HLLP16, KK15, MYH18, Oça23, RSW16, vdHvdHM23]. Initial [AW87b, But94a, But94b, CR11, BCHK18]. Inner [HSS05, Wei82]. Innovations [PS10]. Input [BZ03, Bro77, BPS80, CYZ99, DH82, Dsh85, Fal84b, FS09a, GRW10, JS80, KP84, LW11, MW78, MSS95, Nat75, Ott87b, Rol81, Sch74a, Tak75c, Zwa00, BS22, BPS19, BPS15, XY15, XZ20]. Input-Output [JS80]. input-queued [XZ20]. Inputs [BRT82, Cam88, Le 78, MB00, Miy83, Tur74, Wil98, Wri92]. inradius [CH16]. Inscribed [BFV10]. Insensitive [BS81, BB10, Jan84, MSS95, Sch86a]. Insensitivity [Ano80e, Bur81, Hen83, Miy93, RH89, Sch78a, Sch78b, Sto78, Tay89, VMM19, Whi86]. insider [ACH21, BSHJ17]. Insight [JK85]. Inspection [AH87, LY94]. Inspired [HMV06]. Instantaneous [CM00, CZLR04]. Insurance [Bor09, ES94, PM08, AAM17, BZ22, BEJ16, Buc16, LS21, Nyr16]. Integer [DLO06, McD78, SVW83]. Integer-Valued [DLO06, McD78, SVW83]. Integers [Mel87]. Integral [Amb77a, Amb77b, ADP18, Ber80, BNR87, De 76, DS76a, Duf01, GNRS89, GKW96, Le 74, Leh02, Mil75, Mül97a, Neu86b, NZ77, Rat96b, Rob74b, Sch03c, TF08, Wei98, Ano09e, BD22]. Integrals [Bad77b, BS78, Ell87, GM71, Gat85a, Har80, Mül97b, OB89, Rao83, Rie80, Roo78, Zak77, LX13]. Integrated [Alp09, BGZ95, CA94, CS99, LL89, RSS94]. Intensities [Åbe07, BG08, Bor98, BD93, Pap77, RT02, YS79]. Intensity [Hsi87, Kon89b, Sch96, Sch00, CD18, Gal16, LP21]. Inter [GH75]. Inter-Record [GH75]. Interacting [BMTY97, BP83, BG08, CLCZ12, CB00, CLM23, CDL99, FLS^{+}80, Hor02, KK73, LRS10, PS04, Sud93, Tau74, Lan13a, YS23]. Interaction [Adl89, CORR91, CC09, Der09, Gar71, JP00a, KVB99, MH08, Pie78, Sto23]. Interactions [Dur79, Gri75a, Sch81, CT18b, CPP21]. Interactive [Nar87]. Interarrival [BV93, Jac80a, Kom18]. Intercepts [EC75]. Interchangeability [TW87b, Web92]. interchangeable [WW21]. Interconnected [PT94]. Interdependence [NW98]. Interdependences [WN99]. Interdependent [Nat80]. interdisciplinary [Yin23]. Interest [CY05, Cai07, HJK05, Pol78, ZYY11]. Interesting [CC99b, Gan79]. interface [Sak15]. Interference [KL79, KL83, KS04a]. Interior [BHB98, Man94]. Intermixture [Hio79]. Internally [HPPT92]. Internet [Ott06]. Interparticle [BD87, BD93]. interpoint [Sch16]. Interpolating [SV76]. Interpolation [ÅH99, AS13]. Interpretation [DL82, Gar71]. Interrelations [Hen87]. interruptions [PZ16]. Intersection [Adl89, Hof07a, Kwe75, vdHKV21]. Intersections [Spo01]. Intertwining [HM11]. Interval [BS88, Bal90, BMY91, Bal97, Ber82, FN85, HK88, JH92b, KI92, LS89, LL98, Man76, Man79, Maz89, Nat80, Nat86, Naw91]. Intervals [CFW87, JML92, Rav77]. Intervention [DW02, Wan01, Lem14]. Intrapolymer [Whi80b, Whi80c, Whi80d]. Intrinsic [BFV10, Mat73, Vit01, Sva14]. Introduction [Ano71e, Ano73g, Ano74i, Ano74j, Ano75i, Ano76j, Ano77j, Ano78j, Ano84j, CO98a, Jen96b, Ken77b, McN78, PG85, Tau80, Tav92, Van78a].

Introductory [Tau71]. Invariance [Bal75, Ben74, Ben75a, Bor79, McD85, Miy83, Pol88, Sen89a]. Invariant [Ano93g, Ass78, BW90a, BPS80, CC99a, Coh79, KXZ23, Man90b, NP93, Neu78, O’C91, Pit96, Pol91, Wol80a, BMM13, Cas16, Sto15]. Invariants [FS93, MGW96]. Inventory [DTV84, ENPS78, FZ85, HP83, KP83a, KP84, LY77, Per85b, PS01, TV78, BEJ16, HSZ18, PL89]. Inverse [ARM99, BN94, BNK98, Edm76, FV95, RS12, SZ11, ŠSKB13]. Inverses [CY99, Gol77, Res74b, Ris91]. Inversion [Eng84, Sak04, CKL14]. Inversions [Kat90]. Invertebrates [Fri82]. Invertibility [Hal86, Liu90]. Investigation [AD78, Tau71]. Investigations [Bar84, Heu80]. Investment [ABS00, Bro98, EX11, GW13, HW09a, FS16a, HS18]. Investments [PG97, TT04]. Investor [BH98]. Invited [AAB^{+}92, BGH^{+}88]. Involved [Bob08]. Involving [Bar84, BS95, Neu77a, Pet84, Pie78]. Ion [BMY94, SY97]. Irreducibility [Cli11]. Irreducible [Mai97, Zha97]. Irregular [Sch79]. Irregularity [Dav73a]. Irregularly [Mas83]. Irregularly-Spaced [Mas83]. Irreversible [FS16a]. Ising [KW03a, Kün83, Pic77, Qia90]. Island [Gor01]. Isomorphism [Par77]. ’Isotropic [Gne00, Bro95a, Bro95b, Gne99, JK94, Kid02, Mat98, Mil71a, Nie98, Nie99, Rem00, Cow13, HH14]. Isotropy [Hof07b, Fou18]. Issue [Xu92, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano17h, Ano17g, Ano18b, Ano18a, Ano18d, Ano18c, Ano18f, Ano18e, Ano18h, Ano18g, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano20h, Ano20g, Ano21b, Ano21a, Ano21d, Ano21c, Ano21f, Ano21e, Ano21h, Ano21g, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e]. Itô [BZ20, CF14, Lem14, Wee02]. Items [PS01]. Iterated [FV95, Hey79, Kue75, MS03a, Ver75, CRZ18]. Iterates [BH00]. Iteration [MNW08, NW05, ST10, SF79]. Iterative [BGRSC10, BS06, Cow10, PS94a]. Ito [AT75]. IV [Hun85]. IV. [Ant78].

J [Van83]. J. [Van83]. Jackson [AK95, Cao87, FK89, HS92, KDO16, LS15, Mel79a, TX10, WW21]. Jacksonian [WV80]. Jagers [Ber20, Edl78, Gri92, Hol76b, Jan18, Ner18, Sen18a]. Jante [KV18]. jigsaw [CKM19]. Jirina [BD75, Li09]. Jitter [Kei74b, Kei74c]. Job [BV91, CM93, KRW09, KNT87, MM83, Xu92]. Job-Search [CM93]. Jobs [CNP91, GF87, RX91, Rit94, Wei95b]. Jockeying [XZ96]. John [GN23]. Johnson [Chi95, Hor88, Møl92, Møl95b]. Johnson-Mehl [Møl92, Møl95b]. Join [BMS89, Dai11, HR89, KS04b, LF13]. Joint [CAK00, Cse92, DK97, GM71, GR16, Kat90, Lim97, LR12, MS12, PRR17, PSS20, STZ14, Sch87, LZ13, NSB13]. Joseph [Jen03]. Journal [Gan08]. Jump [AK95, BNBJ00, CC12, CP88, Dro86, FLZ12, GHL07, Hir00, HT10, KW03b, Mal90, PMG08, Ste84, Aaz13, BS15, Chi14]. Jump-Driven [Mal90].

Jump-Type [Hir00]. Jumping [Sud99]. Jumps [ASS99, Dia13, Har77, IM08, Jac05, Kes95, PMG05, BGL20, JS14, KM14b].

Kac [GR08]. Kakutani [KI92]. Kalman [CDJR22]. Karhunen [Mei77]. Karlin [Yu08]. Keilson [Sys74]. Kelly [BW96, WW21]. Kendall [Mil95]. Kernels [HM11, RZ13, HL14]. Kesten [FT18, Liu97]. Keynesian [GVW15]. Keynote [BÇNP84]. Khinchin [DRV00, Neu86b]. Kiefer [RT23]. Killed [CR08, CKLP20]. Kim [ORS18]. Kimura [Edw74]. Kin [Wau81]. Kinds [Bal85, WC90]. Kinetic [AAMRM05, DK12, MC09]. Kinetics [DSSW11]. Kingman [Get74, Yua22]. Knightian [EC21]. Known [BF80a]. Kohonen [BFP98, BP94, BP99]. Kolmogorov [FPZ13]. KPKVB [MS19]. Kramers [Nar91a, Nar91b]. Krein [And79]. Kreps [FLR22]. kth [SG92]. Kushner [Ano98e, CDL99, FP96]. Kushner-Stratonovitch [CDL99].

L [Mon95, San98]. L-Function [San98]. Ladder [Als18b, Fuh07, Sch06a, Sch06b, Wac09]. Lagrange [Lin06, Lin10]. Lagrangian [Åbe07, Zir01]. Laguerre [LZ08, Sch06a, Sch06b, SK85]. Landmark [MD89, MW94]. Landscape [GP02]. Lane [Kon89a]. Language [Nic82]. Laplace [AW88, AW96, CKL14, DLS94, Eck19, KMP06, LZ13, Pri15, RL77, Sak04, SM87, Xu11]. Laplacian [BGV22]. Large [BR15, BB93b, Bor79, BMS09, BH98, But94a, But94b, CT18a, CT18b, CS06, CP08, CYZ99, CF91, Cho93, DM05, FVK08, GLM93, GKT21, Gre89b, GP02, HMM06, HJKP16, HRS04, Ill01, Kel86, KT92, Kus94, LM00, Loa91, Maj07, MZ16, MP10, Mil95, Mil98, Möh06, PP96, Rai07, SS95a, Sad95, Sel96, SW11, Shi93, SW93, SRTM97, SH89, Sto17, Sto23, Wad07, Wei86, Wei78, Whi95, Yao10, YS23, Ber20, BSS17, CD18, CF14, DS15, He16, HJKP17, KS16, LP21, MR20, Miy17, NP17, PB19, Rob18, Sen18b, Sto15]. Large-Deviation [Loa91, SH89, Whi95]. Large-Dimensional [Yao10]. large-intensity [LP21]. Large-scale [Sto17, Sto23, DS15, Sto15]. Large-time [YS23, NP17]. Largest [GI10, Wei78, BB17, CYG14, Sch16]. Laslett [BS01]. Last [Big76, Hsi10, Jan86, KK92, Sal88, Sys76, Tam11, BP23]. Last-Birth [Big76]. Last-Exit [Jan86, Sys76]. Latency [GG12b]. Latent [Can74b, Can74c, Can75b]. Lattice [Cox80, FG00, Gat00, HJ04, Hol90, Mas84a, Pic77, Ros89, Sch81, SW77, SW78, Ste80b, TS98, Tjø83, Wie77, WB71, Jan18, Lan13b]. Lattices [Ano98f, AFH09, Kom96, Wie81]. Law [Cha74, GG12b, Hol76b, Kue75, LS96, MRV02, MS74, Sig91, VJ08, Ver75, Wes90, He16, KV18, KS16, KL13, Sen18b]. Laws [ANR02, BD73, JOC10, MY03, NM01, Pak93, Pak95, RS95, RN70, Ton82, Wad07, Cap86, FPZ13, Fra89, Ros13]. LCFS [RW91]. LCFS-P [RW91]. Lead [Pro76a, Pro76b]. Leader [RR90b, KM14a]. Learning [Baz74, CX04, FM88, Fin80, DR16]. ’Least [KS78, Fin69, LW92, Moo79, Wal74]. Least-Squares [LW92]. Leaves [BGR06, BMS09, CT94, GG12a, Gil02, Mån07, Myl02, BP17, BP22]. Leaving [HK88]. Lecar [LTT20]. Lectures [Hey85]. ledger [Kin21]. Lemma [Bru80, Hol76b, HT17, BR91]. Length [AK08, AS08, BF05, DEKS11, DKS84, Gal11, Gau02, HV01, HO09, HMM06, Hei98, Kes80, Lin72, Mar70, Muc10, PLRC08, PW06, Ros73, Sch87, SGW96, ST94, Tru75b, AD21, DY15, HL17]. Length-Related [Gau02]. Lengths [DM03, HR04, SS95a, Sad95, VGS10]. LEPT [HK93]. Letter [CC05, KO87, CM09]. Level [AA86, BH04a, BL08b, Coh77a, DSSW11, DTV84, Eth93, HP82, Imh86, JLW05b, LN92, Lin78, Lin79, Rej82, SÖ11, WA82, CKO19, OT20]. Level-Crossing [BH04a, LN92]. Level-Dependent [DSSW11, CKO19]. Levels [Hau78, TV78]. leverage [BV13a]. Lévy [Aaz13, BP23, Beh11, BOS23, BBR09, BD22, CM19b, Cov09, DL11, Dia13, Die06, DM22, EK12, EY14, Fas05, FS16a, FL09, GP15, GAH03, GP80a, GM11, GR16, Har77, HPJ08, HW09a, KS09, Kar76, KM15, KR02, Lac04, MS11b, PM08, PvS21, QD15, Qia18, QDZ19, Rog90, Sch07, Sur08, WZ19, WZ20, Woe07, YZ01, BB07b]. Lévy-Based [HPJ08]. Lévy-Driven [Fas05, QDZ19]. Lévy-Type [Woe07]. Libraries [Die82a]. Lie [PS94a]. Liénard [Nar91b]. Lieshout [BV97]. Life [AA88a, BZ22, Buc16]. Lifecycle [Kas78]. Lifelengths [LM90]. Lifetime [Arj80, Bue88, DR03, Fil72, Fil74, Ghu87, Per85b]. Lifetimes [Ano83e, BS99, ENS93, Rub81, ADL15, NSB13, YBC20]. Light [BB93a, BN12, BZ03, DR92, Igl71, LZ05b, Wan97, Kol18]. Light-Tailed [BN12, BZ03, LZ05b]. Lightbulb [GZ11]. Like [FS09a, Har84, KF90]. Likelihood [Ano03e, BA81, DP84, Fei76b, Gla93b, GL95, HW09b, Kei74a, PS09, PS11, Ran77, Ros81, VV01, Van06, Ath18]. Likelihood-Based [HW09b]. Likelihoods [CTS76]. Lilypond [CV04, DL05]. Limit [AMS07, AMS09, AK95, AD94, ANR02, AT80, AK78, Asm82, AN78a, AK98a, AK98b, Bac00, Bad80, BD73, Bar74a, BP79, Ber77, Ber87, BD17, Bin73a, Blu84, Bro95b, Chi95, CQ01, CP90, Clo17, CH80, Coh88, Cre79, DMR88, Din94, DM00, Ell86, FHJK20, Fas14, FGZ17, FDJ20, FPS22, FMY22, FP12, Gly82, GW91, GLS05, GLL22, Gut74, HR89, Här81, HM96, HM99, HSS05, Hey74a, Hey79, Hol76b, HR05b, Igl71, Igl75, IPV05, Iva89, JK11, Jir13, KD92, KS83b, KW87, KK90, KK92, Kes72a, Kes72b, Kes95, Kha80, KR02, KS12, KT92, Lal84, Lee99b, Lem74, Lin74b, Mae79, McD78, Mej77, Mij78, MY01b, Mol93, Myl02, Nak79, Nor75, Num76, Oso10, OT20, PTW10, Pak71b, Pak79, Pak92, Pak98, Pak10, PSS06, Pap87, Pap88, PV19, PW10]. Limit [PR99, PT85b, Qi08, RS95, Rah07, RN70, Ric11, Rog92, Rog94, Roo78, Rud98, SU79, Sch02, ST10, Sco73, Ser75, Shi80, Sil76, ST98, Sma76, Smi83, Sta78a, Sta78b, SGW96, Sys80, ST94, TV75, Tol88, TW18, VJ08, VS80, Voi90, VML80, WH80, Whi82a, Woo80, YZB03, ZK89, vdHvdHLS20, AS14, BN20, BOP21, Che15, GW20, GY17, HS13, JY20, KWRK22, Kin21, KM15, LPST14, NW22, Par98, PD22, RFZ14, Sch16, TO21, TL20, VMM19]. Limited [HK88, Min80, Pic98, Pol75, HHBV20]. limites [Par98]. Limiting [Alb83, Ano92e, BC89a, De 83b, Eth93, KNPV97, KP12, Pak93, Pit86, Sch76a, Sta78b]. Limits [Bal75, Cro10, DW06, FLZ12, JP00a, MP10, NW03, SS95a, Sad95, SKF80, Sch05, TH90, VGS10, AS19, BB17, FM18, ILP18, Kal18, PY13, VBvdM15, vdHvdHM23]. Lindelöf [Ros78]. Lindley [BB90, KS83a]. Line [AK08, BGPS87, CB95, Cal03c, Cow13, DN78, FH74, HR89, Ken74, Kid01, Kus94, CH16, FT18, LBCEN21]. line-of-sight [LBCEN21]. Lineage [Gri06]. linéaires [FG81]. Linear [Ahm80b, BR87, BK18, BDM85, BH04a, BK80, BD93, CY00, CS99, DS03, DDH78, DH76, FR97, Fin69, GO73, Hal85a, Han79, HK84, Hor03, Hua97, Hut80, Kas80, Kes76, KRS07, Kli80, Lal84, Le 78, Led04, LP96, LRS10, LL98, MW93, PL96, QS00, Ros75, Sel96, Sie76, ST04, Tjø90, Wah76, ZG88, FG81, Jir13, KT20, Luo16, Whi81b]. Linear/Quadratic/Gaussian [Whi81b]. lines [Par98]. Linguistic [Gan75]. Link [GI10]. Linkage [AH78, Don81, Hil74, Hil76b, HA79, Stu78, TS99]. Linked [Ave79, HH81, Hil81, Hol78, Wat82, BSS17]. Linking [Cow07, NW22]. Linnik [Gyi75]. Lists [MY80]. Littlewood [McK89]. Litvin [Jof77]. Living [DMR13, Ton82]. LLN [Mas00]. Load [AR06, Bro05, HP82, Mal90, RS21, Smi83, Tie82, BF19, DS23]. Load-Increasing [Bro05]. Load-Sharing [RS21, Smi83, DS23]. Loaded [AMNW11, HK89b]. Loading [Maj07, Mal00]. Loads [GJ80, Tay79]. Local [Adl89, BMZ14, BJML07, CC09, DLS04, DS07, Fri77, Hah98, Han07, HJ08, HP82, KKSS08, Mae79, McD78, McG82, MP03, PW18, Sva14, Tie82, Wor94, vdHvdHM23, AC22, BZ14, HT17, JJL21, Nes17, TK13]. Local-Time [KKSS08]. Localization [AK05, AK05]. Locally [FLS^{+}80, Hor02, KM00, Mas84a, Tau74, BSS23]. Locally-Interacting [Hor02]. Locating [Le01, Sau78]. Location [HJVN03, HS82, MY99, Wor01]. Location-Dependent [HJVN03]. Locations [CQ01, CL00, Lim97]. Loci [Ave79, BS12b, FFT75, Fie80, HH81, Hol78, BSS17]. Locus [Bob04, CR78, EG90, Fie76b, GJS08, HH81, Hin94a, May78, Str75, DAG95]. Loève [Mei77]. Log [CDJR22, Duf04, Agr95]. Log-Normal [Duf04]. Log-normalization [CDJR22]. Logarithm [Hey79, Kue75, Ver75]. Logarithmic [DRS11, Sad95, MV19]. Logconcave [Bac00, LY19]. Logistic [Nor82]. lois [Cap86, Fra89]. Long [CYZ99, DRV00, FGM12, GG12b, GKT21, HM10, JLW05a, KO87, KS12, LS03, LM00, Maj07, MB00, McN90, MSV01, Mil95, PR90, RPS01, Sek06, Sur08, Tam08, WHZ10, CMMS13, GS14, He16, HSZ18, KS16, Vaš16]. Long-Memory [KS12, WHZ10]. Long-Range [CYZ99, DRV00, Maj07, MSV01, He16, KS16, Vaš16]. Long-Standing [Mil95]. Long-Tailed [LM00, MB00]. Long-Term [McN90, PR90, Sek06, Tam08, HSZ18]. Long-Time [FGM12, GKT21, HM10, CMMS13]. Longest [DW06, HMM06, DS15]. Longest-Queue-First [DW06, DS15]. Look [Ald80, FS09a, Ken95, Pal94]. Lookback [DL11]. Looking [Ken95]. Loop [Han07]. Lorden [CN86]. Lorenz [CY99, Gol77]. Loss [AJV13, BGZ95, CLW95, Cho93, Cou90, FS89a, GLM93, HK89b, Hun89, Hun95, KS01b, LM00, MRV02, Mor94, ÖB05, PS90a, RR90a, RS98, SRTM97, Wil98, Yun86, Zac91, ZK89, ALS18a, BD15, VMM19, WZ19]. loss-carry-forward [WZ19]. Losses [Lep76b]. Lovász [HT17]. Low [AK08, Bro72, DGP10]. Low-Cost [AK08]. Lower [Che98, Kat90, KS74b, PL96, WN12]. Lumpability [Pai01]. Lundberg [Pfe83, WCL01]. Lyapounov [Ano98e, FP96]. Lyapunov [CT85, MT93b]. Lying [KPO09].

M [AW88, Ali90, AFGR06, Cha92b, Cho93, DN78, De 83a, GHT96, GHT97, GS95, KS93, KD08, KS04b, LPRS93, Par87, SG92, SK08, SP80, Sys88, TV10, TW87b, Van83, RW91, AW88, Ali90, AFGR06, Box78, BK01, BB10, BC89b, Cha92b, Cho93, Coh77a, CH80, CK15, DT80, Dos78, FS89a, Fal84b, GHT96, GHT97, Gal79b, Gly94, Gol16, GS95, Hal85b, HP96, Hok79, Hok80, JS83, KS83b, KS93, KD08, KS04b, LPRS93, LZ95, MA95a, Neu76a, Neu77b, Ott77a, Ott77b, Ott87b, Par87, Ram84, Ros73, Ros76a, Sat85, Sch84, Sch90, SG92, SK08, Sys88, TH90, TV10, TV78, TVF81, Van78b]. M/G [BC89b, Ott87b]. M/G/ [Gol16, Hal85b, Ram84, Sat85]. M/G/1 [Box78, BK01, Coh77a, CH80, DT80, Dos78, GHT96, GHT97, Gal79b, Gly94, HP96, JS83, KS83b, LZ95, MA95a, Neu76a, Neu77b, Ott77a, Ott77b, Ros73, Ros76a, Sch84, Sch90, TH90, TV10, TV78, Van78b]. M/G/1/ [Fal84b]. M/G/2 [Hok79]. M/G/c [CK15, TVF81]. M/G/k [FS89a]. M/Gi [BB10]. M/K [Hok80]. M/M/ [GS95, KS93, SK08]. M/M/1 [AW88, Ali90, AFGR06, KD08, LPRS93, Par87, SG92, Sys88]. M/M/1/N [Cha92b]. M/M/N [Cho93]. M/M/s [KS04b]. MA [GPD00]. Mabinogion [Cha96, Ste18]. Machine [DG01, GF87, GW92, LBZ98, MGW96, Bré75]. Machines [Wei95b]. Maclaurin [Cao98]. Macromolecules [Sch82b]. Macroscopic [Hal85a]. Made [Ott76]. Maintained [Nat80]. Maintaining [Bar77]. Maintenance [AH87, CD79, Ger84b, GV76, Mi94, RG89b, HA20]. Majority [BS03a]. Majorization [Cha92a, LS93, Sha87, STC94]. Makers [Sza93]. Makespan [PR82]. Makespans [CFGW87]. Malaria [Die71, Pic98]. Malliavin [AE08]. Mammalian [Ben75b]. Management [Ala99, Bro99, DE12, FX00, AM17, LY15a]. Manifold [Bad77b, PR90]. Manifolds [HZ06, Mar77]. Many [Ano06e, ACDHT05, DR92, De 73a, De 73c, De 73b, De 74, De 76, De 77, EK81, Eth90b, ES04, MB00, MZ04, MT06b, Pol94, WH80, DVN15, HLN13, JvL15, PY13]. Many-Server [DR92, De 76, WH80, DVN15, JvL15, PY13]. Many-Sources [ES04]. Map [ZYY11]. Mapping [Stu78]. Mappings [DEP91, HJ08]. Maps [BH00, BP99, LL98, TC92]. marches [Gal84]. Marcinkiewicz [He16, KS16]. Marginal [Ben98, McK86, NLM18]. marginals [CM19a, KTR19]. Marked [AG81, BKS12a, He96, KS09, Kön77, MM98, ARS17, BWZ19, ŠSKB13]. Market [BWY10, BH98, FX00]. Markets [Har80, Pla02, Pri99, TW87a, BV13a, BV13b, DK16]. marking [AT18]. Markov [AWZ93, AE95, Ahm80a, Als18b, AANQ04, AK95, AK79, AHO96, Ano93g, Ano98f, Ano04f, Ano04g, APZ21, Arj72a, Arj72b, AS73, AB94, AK00, AN78a, AN78b, Ath79, AK98a, AK98b, AL99, BS22, BTZ00, BVM96, BS88, Bal90, BMY91, BMY94, BD97, Bal97, BMTY97, Bar74a, Bar76b, Bar80, BS81, BHKK15, BNBJ00, BE88, Bar18, Bat73a, Bat73b, Bat73c, Bat76b, Bat78, BMS18, BOR02, Ber87, BR86, BL88, BC87, Big87, BS15, Bob04, BFK92, BH04b, BL08b, BG93, BC81b, BDD12, Buc16, BC04, But22, Can74b, Can75b, Cao95, Cao98, Cas16, CCHH18, Cha89, CF91, CPZC05, CB99, CB00, Çin69, Çin74, Çin75, Çin78, Coh76, Coh77d, Coh79, Coh81, Cra17, CP88, Cra74a, Cra74b, CM17, Cse92, CKLP20, DS93, DP23, Daw75, DFT77a, DFT77b, DSS17, DSS18]. Markov [DNPR01, Die82b, DHP12, DP13, DP15, Eck19, Eco05, EF80, EF81, EP15, EN80, EN88, FKM82, FT78, FPS22, FS79, FLZ12, FS82, Fro13, FS16b, FJ09, Fuh97, FL01, Fuh07, FP12, GHT96, GHT97, GT12, GTY70, Ger94, Ger95, Gha08, GL79, Gly82, GT08, Gre89a, GGV89, GP02, GHL07, GY10, GVZ13, GHH15, GKPZ21, GH75, Gzy76, Häg00, HS94, HV84, HH03, HLLP16, Hen83, HPPT92, HL14, Hop85, HIS76, HV83, HS92, Hor02, HK22, HJM^{+}16, HLG18, Hüb78, HM11, Hun69, Hun83a, Hun83b, Hun84, Hun85, JLW05a, JLW05b, Jac05, Jac80b, JH92a, JH92b, JN98, Kai78, KA81, Kap73, KD85, KD92, Kei74b, KS93, Ken78a, KMW96, Kes76, Kes95, KT20, Kin73, KM80, Kom96, KM95, Kur18, Lam93, LT01, Led04, LL97, Leh02, LTT20, Let82]. Markov [Lev77, LZ05a, LZ05b, LS13, LL18, Man74, Man88, MZ16, MY01a, MS87, Mas15, Mas23, Mat72, McD85, McG82, MLZ21, MSS95, Möh98b, MW98, Mon17, Mor73, NM79, NP93, NR21, Neu70, Neu76b, Neu78, Neu86c, NS77, Num76, Oça23, Oka79, Ott74, Ott75, PT81, Pak89a, Pak93, Pak95, Pak10, PSTB17, PH99, Pea89, PD22, PHL16, Per97, Pes75, Pic80, Pic85b, Pit77b, Pol94, Pol86, Pol87, PR90, PS94b, PC96, PZ86, Pre73, RSZM02, RL10, Rav83, RN70, Res72, RZ13, Ris91, Rob76b, Rol78, Roo88, RS07, Roz78, Roz80, Rud73, Rud74, RSW16, Sch71, Sch76a, Sch78b, Sch05, SF79, Seg20, Sen80, Sen89b, Sen97, SP73, Ser89, Ser05, Sha80, SD98, Sim80, SS77, SE91, Sys75, Sys77, Sys79, Sze76, Tak75c, Tav76, Tav78, TV10, Teu73]. Markov [Tho81b, Tjø90, Tod76, Tol88, TS99, TT79a, TT79b, TT79c, TT94, Twe76, Twe77, Twe82, VY94, VP88, Van88a, Van90, VW77, Van78d, VV06, Wes76, Whi83, Wij78, Wil85, WD90, XHW23, YS79, YA00, Yca84, YS94, YZ02, YZB03, ZL09, vL18]. Markov-Dependent [FJ09, KD92]. Markov-Modulated [AB94, KM95, SE91, Bar18, HJM^{+}16]. Markov-Renewal [Sim80]. Markovian [AAR08, ADP18, BN23, BPSZ06, BPS80, BP82, CR04, CPR01, Coh77b, Coh77c, CC99b, DS95, DS16, DGR02, FE98, GLL22, Hüb80, JKS07, LT01, LC08, LMS00, Mal90, MP18, Mel79b, MT92, MT93a, MT93b, Mol78, Nak96, Neu77c, PC21, PS01, PL99, Pri22, Pru78, Sma76, SW80, Tur74]. Markoviens [Mat72]. Marks [RPS01]. Marriages [Haj76]. Marshall [SMAL17]. Marshall-Olkin [SMAL17]. Martin [Sys73]. Martingale [AK78, AK00, GLS01, HJK05, Hey75, Hey77, KS74b, McG82, MZ77, Miy04, Pri99, Shi80, Sys80, VL07, CM19a, Che15]. Martingale-Type [VL07]. Martingales [Dvo77, Far76, Har80, Hey74a, McL75, Sco73, TW87a, Bré75, BJ77]. Mass [DR85a, Kem73]. Match [CJ11, Man21]. Matchbox [Sta98]. Matches [KO87]. Matching [CKW09, DL05, KKT06, BGM13, Jan23, RM21]. Matchings [Soo10, Man21]. Material [KPUU01]. Materials [De 83b, HP82, Sob87]. Matérn [HW09b]. Maternal [Car78]. Mathematical [Arn80, CSH^{+}80, Gar71, Gra82, KR71b, Mes71, Nag71]. Matheron [Jeu00]. Mating [HA79, SV75, Sta80, Vet78, Wil79, GY17]. Matrices [FVK08, Ken78a, Kho01, Mai97, Ott74, Ott75, SB09, ZLB98, BS22, Din19, FJKR20, GLM22]. Matrix [AL99, Dag75a, Dag75b, DP23, DSSW11, Edm76, Fac09, FM88, GT08, HZ07, KS93, Neu78, Pes75, Sen89b, Smi76, Wol80a, YS94, Mas23]. Matrix-Analytic [DSSW11, GT08]. Matrix-Exponential [Sen89b]. Matrix-Geometric [Neu78]. matrix-infinite-product-form [Mas23]. Matter [Ano69a, Ano69b, Ano69c, Ano69d, Ano70a, Ano70b, Ano70c, Ano70d, Ano71a, Ano71b, Ano71c, Ano71d, Ano72a, Ano72b, Ano72c, Ano72d, Ano72e, Ano72f, Ano73a, Ano73b, Ano73c, Ano73d, Ano73e, Ano73f, Ano74a, Ano74b, Ano74c, Ano74d, Ano74e, Ano74f, Ano74g, Ano74h, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano75f, Ano75g, Ano75h, Ano76a, Ano76b, Ano76c, Ano76d, Ano76f, Ano76g, Ano76h, Ano76i, Ano77a, Ano77b, Ano77c, Ano77d, Ano77f, Ano77g, Ano77h, Ano77i, Ano78a, Ano78b, Ano78c, Ano78d, Ano78f, Ano78g, Ano78h, Ano78i, Ano79a, Ano79b, Ano79c, Ano79d, Ano79g, Ano79h, Ano79i, Ano79j, Ano80a, Ano80b, Ano80c, Ano80d, Ano80f, Ano80g, Ano80h, Ano80i, Ano81a, Ano81b, Ano81c, Ano81d, Ano81e, Ano81f, Ano81g, Ano81h]. Matter [Ano82a, Ano82b, Ano82c, Ano82d, Ano82e, Ano82f, Ano82g, Ano82h, Ano83a, Ano83b, Ano83c, Ano83d, Ano83g, Ano83h, Ano83i, Ano83j, Ano84a, Ano84b, Ano84c, Ano84d, Ano84f, Ano84g, Ano84h, Ano84i, Ano85a, Ano85b, Ano85c, Ano85d, Ano85f, Ano85g, Ano85h, Ano85i, Ano86a, Ano86b, Ano86c, Ano86d, Ano86g, Ano86h, Ano86i, Ano86j, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano87f, Ano87g, Ano87h, Ano88b, Ano88c, Ano88d, Ano88e, Ano88f, Ano88g, Ano88h, Ano88i, Ano89a, Ano89b, Ano89c, Ano89d, Ano89f, Ano89g, Ano89h, Ano89i, Ano90a, Ano90b, Ano90c, Ano90d, Ano90e, Ano90f, Ano90g, Ano90h, Ano91a, Ano91b, Ano91c, Ano91d, Ano91f, Ano91g, Ano91h, Ano91i, Ano92a, Ano92b, Ano92c, Ano92d, Ano92f, Ano92g, Ano92h, Ano92i]. Matter [Ano93a, Ano93b, Ano93c, Ano93d, Ano93h, Ano93i, Ano93j, Ano93k, Ano94a, Ano94b, Ano94c, Ano94d, Ano94f, Ano94g, Ano94h, Ano94i, Ano95a, Ano95b, Ano95c, Ano95d, Ano95f, Ano95g, Ano95h, Ano95i, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano96f, Ano96g, Ano96h, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano97h, Ano98a, Ano98b, Ano98c, Ano98d, Ano98g, Ano98h, Ano98i, Ano98j, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano99f, Ano99g, Ano99h, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano03a, Ano03b, Ano03c, Ano03d, Ano03f, Ano03g, Ano03h, Ano03i].

Matter [Ano04a, Ano04b, Ano04c, Ano04d, Ano04h, Ano04i, Ano04j, Ano04k, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano06a, Ano06b, Ano06c, Ano06d, Ano06g, Ano06h, Ano06i, Ano06j, Ano07a, Ano07b, Ano07c, Ano07d, Ano07f, Ano07g, Ano07h, Ano07i, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano09a, Ano09b, Ano09c, Ano09d, Ano09f, Ano09g, Ano09h, Ano09i, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Lep76b, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano17h, Ano17g, Ano18b, Ano18a, Ano18d, Ano18c, Ano18f, Ano18e, Ano18h, Ano18g, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f]. matter [Ano20e, Ano20h, Ano20g, Ano21b, Ano21a, Ano21d, Ano21c, Ano21f, Ano21e, Ano21h, Ano21g, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e]. Max [Hüs89, WS10, WS11, EMOS14, Mai97, DR89]. Max-ARMA [DR89]. Max-Stable [WS10, WS11, EMOS14]. Maxima [Bal77, BX01, Bin73b, BK05, Cha09a, CNT05, GKL21, Gre75, Ker86, Mer06, Pak98, Pak00, RN70, Res72, Roo88, Sam04a, TV75, TW84b, Wor94, MY20]. Maximal [Ano86e, Fra85, Han07, Hua90, KD92, MZ77]. Maximally [DSS99]. Maximises [McN90]. maximising [SS20]. Maximization [HMV06, JD89, Tam08, BMS18]. maximizes [PR22]. Maximizing [GY16, Ros75, Tam11]. Maximum [Ano85e, Ano03e, AR97a, AG05, BT98, BA81, Dan74b, DS85, Dan89, DS07, DGM09, Edw74, Fei76b, GP80a, GP80b, GM98, Kei74a, Nav03, Naw91, PS09, PS11, Ran77, Ros81, Rüs82, Sho79, Tak70, Teu79, TS99, VV01, Van06, HR16]. may [FZ13, Gre89b]. MCARMA [Fas14]. MCMC [AVLR23, FDJ20, SWL14]. Mean [Agr95, Ano85e, AB94, BP79, BP00, Ber78c, Blo83, Blo89, BGL20, Bro89, CB03, CORR91, Chi94, CS12, DS85, Dan89, GLL20, Hoo78, Jan20, Kid02, Kid08, LH07, LJ04, Le95, Le96, Le98, LK00, Mil75, MY01b, Nar91b, Pak79, Pro76b, RW91, Rit94, Rus10, Sch01, SB77, Sch78c, Sch79, Wat92, Wei95a, Wei97, Wit79, Pia01, Sto23, VMM19, YS23]. Mean-Field [CORR91, Nar91b, Sto23, YS23]. Mean-Value [Chi94]. Mean-Variance [CS12, LH07]. Meander [CH80]. Means [GKW96, IPV05, Kid98, Kid02, KL00, KL03, Le96, LK00, Le01, Mol95a, Rat98]. Measurable [Mec77, GLR15]. Measure [AO95, AFM91, BK04, Cov09, Eth93, EG90, Ger77, GLM90, HJK05, Kid02, KJ03, Kid08, KM96, KM99, KM10a, Li92, MS80b, Pri99, Sta78b, VJ05, Wol80a, Clo17, MZ16, Qia18]. Measure-Branching [Eth93]. Measure-Valued [EG90, Clo17, Qia18]. Measured [Sen88]. Measurements [HLW03]. Measures [Ano77e, Ano93g, ACF09, BKZ95, BS07, BR76, BHJ11, BS78, BY09, Bri99, EK12, FV95, Gra77, Gri77, HM96, HM99, KL79, Kar76, MY80, MS76, NP93, Pel83, Pit77b, Pol91, Rat96b, Rat98, Roy07, Sch01, Urb77, YZB03, Zäh96, BCM18, BMM13, CM19a, Eck19, Vaš16]. Measuring [AIL04, Coh85, KL83, WR22]. Mechanical [Orn77]. Mechanics [Cor78, KR97]. Mechanisms [Adl89, Jir74, BF19]. Mecke [Jen03]. Mehl [Møl92, Møl95b, Chi95, Hor88]. Melamed [Ano79f]. Membrane [MRV02]. Memory [Gal11, KS12, LS03, Sur08, WHZ10, AH16, FPZ13, GS14, HL17, RR21]. Mendelian [Odo80]. Merge [Cra97]. Merge-Sort’ [Cra97]. Message [LA02]. Metapopulation [Arr03]. Metapopulations [MP10]. Metastatic [Bar85b]. Method [Ano98e, AR92, Bar76a, BH89a, Baz74, BS06, BM96, Bra98, CSY98, CP88, DLU12, DFT77a, DFT77b, FP96, Lan82, LW92, Rai07, Sys74, Tru97, Wei71, Bré75, GLM17, LL18, Ros13]. méthode [Bré75]. Methodology [Bri76]. Methods [Ben74, Ben75a, BT08, BDD12, Dai08, De 77, Ger84a, GT08, Gre75, Hey75, HS90, HIS76, IS83, JS83, Ken88b, MM98, Mes71, MT09, Mit87, Nag71, Ott87a, Pes75, Rav83, Sch06a, Sch07, Tak75a, Tak75b, Tan96, WDH98, FDJ20, LTT20, MT13]. Metric [Deh98, Pol79, SX08, KXZ23]. Metrics [HR00, Mül97a, RR95]. Metropolis [AVLR23, JY04, JRR14, MV19]. MGF [PM08]. Microorganisms [KR71b]. microscopic [RT21]. Microtubule [HM10]. Migrating [JP98]. Migration [Aks75, Dav70, Dur75, Gor01, Hel75, HKH76, Hös11, Ser05]. Mine [CL00]. Minima [AG81, Gre75]. Minimal [Ave83, Bac00, BLP06, BF05, Fin92, GH80, HNS13, Lee99b, Nij74, Pen98, PW04, PW06, Pri99, DK16, Thé22]. Minimax [BL06]. minimization [DK16]. Minimize [RX91]. Minimizing [ASC96, CFGW87, KSC09, PR82]. Minimum [AR97b, Jan86]. Minkowski [Mas79]. Mises [Ken77c]. Missing [DR81]. Mixed [Alb83, Fas05, JL77, Mec96, Mec98a, MS11b, Rat96b, SÖ11, SK73, Wei01, HW19, KMF22]. Mixed-Level [SÖ11]. Mixing [AP98, LR11, McL75, Pel83, Roy07, VPCG10, ABL15]. Mixture [BZ03, FE06, NSB13]. Mixtures [BLL12, JL79, SS77, Wil90, BE21]. Mobile [BTZ00, FVK08, MY11]. Mode [Ber20, BD75, Edl78, Gri92, Hol76b, Jan18]. Model [Aas82, ACH21, AW79, AP05, AAR08, ACDP12, ACV80, AT80, Arr03, ABS00, AD78, Ave83, BB01, Bal78, BBSW09, BC93, BB07a, BS12a, BB90, BS01, BND81, BB07b, BR93, Bat76a, Ber78b, BRS99, Blo89, Bob08, BGR06, Bor98, Box79a, Box79b, BPSZ06, Bro85, Bul74, CMP09, Cam79, Car78, CR11, CY00, CV04, CR78, DSS99, DL05, DRS11, Dav70, DH91, DFT77a, Dei03, DH04, DK12, Dom07, DLS94, DLO06, Dyn78, ENPS78, ES94, EK81, Eth90a, Eth90b, EG90, EX11, FH02, FRW12, FS79, GG12a, Gal79b, Gar80, Ger77, Gil02, Gor01, GKP84, Gri79a, Gri79b, Gri80b, Gri82, GJS08, Gue93, GG90, Had79, HL85, HJK05, HRTY97, HMV06, HW95, HH81, HJ02, HV83, HM10, Hui03a, JMJ01, Kap73, Kel77, Kel12, Ken96, Ken98, Ker78]. Model [Kes70a, Kes70b, KD08, KI92, Lan05, Lan96, LL81, Lew75, LTW10, LR12, Lip75, Liu90, LA02, Mån07, MPP07, Mal90, MGW96, Mar80, MT11, MP03, Mit88, MS94, Mol95a, Nås96, Nat75, Neu77a, Nor82, Pai01, PT85a, Pak92, PS90a, Pen91, PL89, Pfe85b, Pfe91, Pic98, PP07, Pru78, PT85b, Qia90, QS00, QT97, Rob74a, Ros74b, SK96, Sat78, SKF80, Sch82b, Sch76b, SP73, SS84b, Shi79, Sma76, Ste84, Ste76b, Str75, Sud99, Sur08, Sut79, Tol88, VS87, Van78b, VV06, Wal81a, Wat79a, Wat79b, WWY12, Wen89, Whi91, Wil88, Woo80, Yeh90, ZYY11, ABL15, AT18, AG13, AZ18, AS19, BDLP14, BV13b, BEJ16, BFP23, BGM13, CN18, EL14, Fox14, Fri87, GNR17, GY17, GKT22, Hir17, HT17, Hou18, IMS21, JKM23, JKR17, KWRK22, Kin21]. model [Kub19, Lan13b, LT15, LZ17, LTT20, MS14, NRB22, ORS18, RM21, RFZ14, Ste18, Yua22, ZY16]. Model-Based [HJ02, MGW96]. Model-independent [ACH21]. modèle [Fri87]. Modeling [BP83, BGR06, Ebr93, KK03, BV13a, BJ77, Cor14]. modélisation [BJ77]. Modelling [All82, Ano84e, BO93, BMY94, BNBV14, BRB^{+}71, Bri76, Hin94a, Hin94b, HJ00, LJJ04, Pes82a, RS95, Tjø78, Tjø83]. Models [AE95, ABB^{+}80, ABA^{+}80, AANQ04, AK79, Ano93e, Ano93f, Ano94e, Ano03e, ADW^{+}75, ACS93, Arn80, AR92, Bal86, BD87, BD97, Bal97, Bal07, BP00, Bar84, BFM^{+}84, Bäu02, BBD99, Bha86, BR76, BKS12b, BC76, BM03, BLS88, Bro80, BN92, Can74b, Can74c, Can75a, Can75b, Can76, Cha96, CS97, CPR01, CQ01, CLW95, Cli07, Cli11, CL00, DR06, DR08, DG04b, DDH78, DP84, DL11, Die71, Dow74, DS76a, Dro86, DR03, DH76, Elv71, Fad77, Fad79, FKM82, FZ85, Fei76a, Fie75, FL09, Fin80, Gan75, Gar71, GW78, Gat85b, GL09, GJL84, GAH03, Gla93c, GKW96, GLS05, Gra82, GKT21, Gri75b, Gut90, HK90, Hal86, HK84, HM96, HM99, HKM^{+}84, HK22, Hua97, HLW02, JML92, JP00b, KX94, KP10, KS93, KW03a, Kes72a, Kes72b, Kid96]. Models [Kin79, Koc90, KM95, Kry76, KP12, KS12, Kün83, LS98a, Las04, LS11, LR97, LP99, Lin06, Ma05, MSM92, Mar70, Mas84a, McK88, Möh98a, Möh00, Möh04, Møl95b, Mor76, MS11b, NPOTC00, O’N95, Odo80, PTW10, PSS06, Pel93, POGP95, PL89, Pim07, PC96, PZ86, RSZM02, Ren86, RR90b, RS71, Ros74a, RT02, Rud98, Rus10, RS21, San98, Sar97, Sar98, Sch86a, Sch03b, SS84a, SC77, Shi82, SD98, Sin98, SJ89, SW98, Sob87, SZ90, SY97, Ste81, SE91, SS85, Tau71, Twe82, VK78, VV01, Was78, Wei98, Wei01, Wes79, Woe07, Woo74, Ada16, BBRW14, BK18, BWZ19, BOP21, BCJ^{+}17, BSS23, BPS19, CM19b, CMMS13, CV23, CGR18, CM17, CDJR22, DS23, EY14, EKKZ18, Fas14, Fox16, GRR17, GS19, GHL23, GLM21, GLM17, GB17, HHBV20, HLPW14]. models [HW19, KL13, MJM21, NLM18, NRB22, ORT16, RT21, Sak15, Sva14, VBvdM15, vdHvdHLS20, Ano07e]. Moderate [CYZ99, DJ17, SH89, GW20]. Moderate- [SH89]. moderately [GLM17]. Modification [BO93, Cha76]. Modifications [Pic00]. Modified [KS83a, Phi78]. Modular [Maz89, GC16]. Modulated [AB94, BS03b, KS93, KD08, KM95, Sha97, SE91, Bar18, HJM^{+}16]. Modulated-Poisson [BS03b]. Modules [Gås12]. Molecular [AAR08, DR85a]. Molecules [Reh80]. Moment [BK99, CL22, HHHT19, Lep76a, Neu76b, PM08, SS20]. Moment-constrained [CL22]. Moments [BB10, Coh77c, CWW09, Fuh07, HR04, Hun69, Ill98, Jan86, Kar90, KP86, Kem87, KS74b, KM96, KM99, KP94, LPST14, Mec98b, MW96a, MM83, Pia06, Pic00, Pri22, CHY20, DP23]. Monotone [Ano91e, Bal75, Bar85a, Ben91, Bue88, FN85, Gås12, Gal79b, GHK13, KM14c, LS93, Nat86, Nat89, PS90b, SR00, BS22, BZ14, LY15a, Mas15]. Monotonicity [CCP91, Coh88, Gla93b, Kij92, LS98b, LMX05, Van80, Fox14]. Monte [BCJW14, BT08, CR08, Dai11, GR08, Gla93b, LTT20, MT09, PH99, Pes75, PS94a, TKF15]. Moran [AD78, BS83, CV23, FX13, HL85, Kel77]. Morris [LTT20]. Morse [ENR17]. Mosaic [CT94, Hal85b, Hal85a]. Mosaics [HS11, ENR17]. Most [Möh04, O’C95]. Mosteller [Fin80]. Motility [ABB^{+}80]. Motion [AW87a, AW87b, Bar81, Bat70, Bat76a, BK05, Cau74, Cau76, CK85, DMY93, Duf01, FS09a, GRW10, Imh86, KMP06, LW11, LOS03, MPS98, Mer79, Ott76, Pit74, Pit75, Sal88, Sch03c, Sch06b, ST04, Taq75, VJV10, Yor92, Yuh82, Att17, BD22, KM15, KZ16, MZ20, PZ17, RY14, RSY23]. Motions [Adl79, CCH00, De 10, Di 01, GKL21, Gem80, Jac74a, Lac04, LO04, MY03, Max77, Ors90, OD07, Pie78, SW13]. Motivated [BO93, DR85b]. mouse [FPS22]. Movement [HW95]. Movements [CM00, YM98]. Moving [Bad77b, BLS88, Cha09a, Fas05, Hal86, JL77, McK86, MS11b, Sal88, Sto80, Wal74, Oça23]. Moving-Average [Hal86, McK86]. Moyen [Pia01]. Multi [Agr95, ARM99, AK00, CLW95, FFH86, Gla83, Gri79b, Kes76, Kos76, Sat78, SRTM97, SK73, Son79a, Son79b, Van73a, AF18, BPP21, BPS19, BLT19, CMT21, FHJK20, FT18, Fox16, HG15, JS14, Seg20, YBC20, ZCJW19, ZCLJW23]. Multi-Access [FFH86]. multi-action [HG15, ZCJW19, ZCLJW23]. Multi-Allele [Gri79b]. Multi-Allelic [Sat78]. Multi-Armed [Agr95, Gla83, ZCJW19, ZCLJW23]. Multi-Class [CLW95, AF18, CMT21]. multi-component [Seg20]. Multi-Dimensional [AK00, Kes76, BPS19, FT18, JS14]. Multi-Entry [Kos76]. Multi-Rate [CLW95, SRTM97]. Multi-Resolution [ARM99]. Multi-Resource [CLW95]. Multi-Server [SK73, Son79a, Son79b]. multi-state [YBC20]. Multi-Terminal [Van73a]. multi-type [BPP21, BLT19, FHJK20, Fox16]. Multicasting [RSZM02]. Multichain [SF79, Mon17]. Multichannel [BGPS87, Fal84a, Szc93]. Multiclass [Ano04e, Fal88, PR00, Rei88]. multicolor [ZHCC14]. Multicompartment [Tol88]. Multicomponent [SM78, Szá76]. Multidimensional [DMR13, FL01, GS77a, HS85, KS75, Lin80, Sev71, WC90, Ada16, BPS15, EC21, GVW15, KTR19, RT21]. Multidirectional [Roz80]. Multifractal [BHJ11, KM10a, MNR02]. Multifractality [ALS08]. Multifractional [Lac04, ST04]. Multigene [Wat92]. multigraph [GK19]. Multilevel [DH91, Wu93, LTT20, LLX18]. multilinear [AS13]. Multilocus [Hol81, KL79]. Multimode [KNT87]. Multinomial [SR76, ZL09]. Multinormal [FFG07]. Multiparameter [Kur78, Smy76]. Multiphase [Mac75b]. multiplayer [GY16]. Multiple [ARW11, BE85, BS06, BV90, BV91, CH76, De 80, Fie80, IW70a, IW70b, Imh08, MM83, MM85, Tho75, Whi70, Whi96, CGR18, HLG18, KMF22, Li23]. Multiple-Server [ARW11, De 80]. Multiple-Trial [Imh08, Whi96]. Multiplexing [JL99]. Multiplication [Hol78]. Multiplicative [Bar01, DGR02, Lev78, PT23]. Multiplicative-Decrease [DGR02]. Multipurpose [Has76]. Multirate [GLM93]. Multiresolution [KW03a]. Multiscaling [ACDP12, Way96]. Multiserver [Dsh85, GHT88, GM08, KRW09, PT94, SS95a, Sad95, Miy17]. Multistage [KF90, Sim15]. Multistate [Ano83f, Bue88, FJ09, FN85, Gås12, Iye88, Nat82, Nat86, NSHH86, Nat89]. Multitrees [Buk01]. Multitype [AD94, And99, AK78, BC93, BL01, Big76, BK04, BS05, Can76, GB03, GLM90, Iva89, JM86, KM80, Li92, LMW96, MYH18, Pim07, Pol74, Qui72, RY14, SS09, Tan81, Ale13, BMPR19, CN18]. Multitype-Branching [LMW96]. Multivariable [Pri77]. Multivariate [BNPS01, BS99, BG08, BM16, Ber78d, Ber78e, Beu83, BS05, EK12, Ger77, Ghu87, HK84, Hua02, HL02, KR81b, Kei74b, Lat97, MD79, MS11b, OMS06, Pri77, RR90b, Ros76b, SL80, SS87, Slu84, Tak88, BN20, BV13b, BZ20, CG16, Cha22, KTR19, LPST14, NW22, PB19, PV19, TKF15]. Mutant [CW75, Hil76b, Wat82]. Mutants [Pak92]. Mutation [BKS12b, BB96, FM09, RT07, BB17, CV23, GRR17, Yua22]. Mutation-Selection [BB96]. Mutations [Ben75b, CN18]. mutator [BL22]. Mutual [DL05, Sch86b, Sch77, ADL15]. Mutual-Nearest-Neighbour [DL05]. Mutually [Hol83, Tak70].

n [Agr95, Thé22, Cha92b, Cho93, KW90b, PR00, Ram80]. N/G/1 [Ram80]. Nash [AB22, Att17, Bro05, Lin15]. Natural [AFN74, BLL12, Edw74, KL79, LC76, Rob76a, Vet80]. Nature [GVM^{+}03]. Navier [MT13]. Navigation [BM11]. NBU [MS80a, RA98]. NEAR [LL81, BPS05, LJ04, Pak00, Tru97, Kot87, Qui72, Rob18, SS15, WA82]. Near-Critical [LJ04, Rob18, SS15]. Near-Maxima [Pak00]. Near-Record [BPS05]. Nearest [BBSW05, BBSW09, BBD99, CM03, Col74, DL05, FRW12, GI10, Hen87, NRT83, NR85, Sha09, Van94, Wad07, vL18]. Nearest-Neighbor [GI10, Sha09]. Nearest-Neighbour [BBD99, Hen87, Van94, vL18]. Nearest-Neighbour-Type [Wad07]. Nearly [KM95, OC10]. Necessary [Käm87, WCK96, Che15, XCWC18]. Negative [Bal90, CS99, Har77, HP96, Imh86, Ken78a, McK86, MW96a, PvS21, SM87, Tuc71, Ver79, Vet78, Wac09, YZ01, BP23, BOS23, CAR20, CKLP20, EY14, GR16, IMS21, Mar16, Ros13, WZ20]. Negative-Binomial [McK86]. Neighbor [GI10, Sch86b, Sha09]. Neighbors [NRT83, NR85]. Neighbour [BBSW05, BBSW09, BBD99, Col74, DL05, Hen87, Van94, Wad07, vL18]. Neighbourhood [Chi94]. Neighbours [CX04, CM03, FRW12]. Nerve [AS95]. Nested [ABH80, LS98c, STW84, CM17, GC16]. Network [BST09, BS12a, DR06, DR08, ES04, FK89, GC02, Hoo78, HS92, Hun95, KF90, LW11, MT97, Mor94, SW11, SRTM97, TJM98, TE96, TW83, VK78, BGV22, DS16, FR14, GC16, MS14, ORS18, PY13, RR21, WR22]. Networks [AK08, AK95, Ano80e, BTZ00, BBW09, BBSW10, BW89, Bar76a, BN94, Bäu02, BGZ95, BM77, BVH10, BM98, BV90, BV91, Bra78, BP82, Cao87, Cao89, Cao90, CM00, Dad82, Dad84, DS95, Dai11, DGP10, DMM12, FE98, Fri82, GX07, Har84, HT92, HNT94, HK89b, Hun89, Hun95, IW70b, Igl80, IS83, Jan84, KKSS90, Kel76, KP83b, Kel86, Kel88, KS04b, Kog89, KW90b, LY88, Law92, LT00, LRS10, LT94, Lom72, Mal93, MK01, Mas84b, MS01a, Mel79a, Mit87, MW96b, Pal93, Pal94, PW10, PBM00, RSZM02, RT90, Sch78a, Ser89, Sto78, TX10, Tör78, TW84a, VT94, WV80, WS87, YM98, Zac91, Zha97, ZK89, AM17, BR13, CT18a, CT18b, CL14, CL15, EW14, HJKP16, HJKP17, JJL21, KKV16, KDO16, LS15, MR20, PWZ17, PC21]. networks [SS20, WW21, ZM16, vdHvdHLS20]. Neural [Fri82, Wal81a, CT18a, CT18b]. Neuro [WDH98]. Neuro-Anatomy [WDH98]. Neuron [Fie75, Kry76, PTW10, Whi91]. Neutral [BF05, EK81, Eth90b, EG90, Gri79a, Hil74, Hil76b, Kel77, Ric11, Wat74, BSS17]. Neveu [Hui03a]. Nevzorov [Pfe91]. Newell [Dad82, Dad84]. Newman [ABL15]. Newton [SK20]. Nikodym [Tod76, Tod77]. Niu [Gre87]. No [Ano89e, Har77, IL73, PT23, Tam13]. no-information [Tam13]. Node [BGPS87, FK89, HK92]. Nodes [PBM00, SS92]. Noise [AAMRM05, Beh11, BY09, Bri99, HK90, HKT91, HS85, Hsi87, HT89, IJ02, KD85, Kon89b, Lan87, Lin72, Mar75, Møl03, MT05, Ors84, RPS01, Ric77, WCK96, Wei76, BCHK18, BFW22, CF14, LP21, Qia18]. Noises [BM02, KS09, LP21, DMYT23, QD15]. Noisy [Wah76]. Non [AW87b, AH97, ASS99, BH00, BD73, Bow21, BD93, CQ01, Coh76, Coh77d, Coh81, DSS99, Dav79, DMC00, DM22, FE98, FMP92, Fin69, GHT88, GHT97, Ger80, GNR87, GNRS89, Gri80b, Gri77, Had74, Hal86, HRY93, Hor88, Hua97, Hüb78, Jir76, Ken78a, Ker78, Kid01, Kli80, Lal84, Lea76, Lee79, LP90, LP96, Lin78, LMHN90, Mae79, MW93, Mas83, MW96a, MT97, Neu87, OD07, Par77, PL96, Pol79, Pru78, RX91, Rob74b, Rol81, Ros75, SR76, Sma76, Tan96, Tjø76a, Tjø90, Tur74, Van74, Ver79, Wis80, Woo74, BV13b, BZ22, CLM23, DR16, Lac20, MMP22, MP18, Nes17, VBvdM15, BB07b]. non-adapted [MMP22]. Non-asymptotic [DM22]. Non-Bandlimitedness [Wis80]. Non-Decreasing [BD73]. Non-Ergodic [Dav79, Lin78]. Non-Euclidean [OD07]. Non-Expansive [BH00]. non-exponential [Lac20]. non-FCFS [VBvdM15]. Non-Gaussian [Bow21, BV13b, DR16]. Non-Homogeneous [Coh76, Coh77d, Coh81, HRY93, Jir76, Van74, Woo74]. Non-Identical [RX91]. Non-Integral [Rob74b]. Non-Lévy [BB07b]. non-life [BZ22]. Non-Linear [BD93, Hua97, Kli80, Lal84, LP96, MW93, PL96, Ros75, Tjø90]. non-local [Nes17]. Non-Markovian [FE98, Pru78, Sma76, Tur74, MP18]. Non-Multinomial [SR76]. Non-Negative [Ken78a, MW96a, Ver79]. Non-Optimal [Hüb78]. Non-Overlapping [DSS99]. Non-Parametric [DMC00, Kid01, Mas83, Par77]. Non-Poisson [CQ01, Had74]. Non-Preemptive [GHT88]. Non-Random [Gri77]. Non-Renewal [LMHN90, Neu87]. Non-Separable [Pol79]. Non-Signalized [Gri80b]. Non-Singular [GNRS89]. Non-Skip-Free [GHT97]. Non-Spherical [Tan96]. Non-Standard [ASS99, LP90, MT97]. Non-Stationary [AH97, Ger80, Hal86, Ker78, Lee79, Rol81, Tjø76a]. Non-Stochastic [Fin69]. non-synchronization [CLM23]. Non-Topological [Lea76]. Non-Trivial [FMP92]. Non-Uniform [Mae79]. Non-Zero [AW87b]. nonautonomous [KDO16]. Nonbounded [Der09]. Noncrossing [KS01a]. Nondependence [CR11]. Nonexplosion [CLM08]. nonhomogeneous [BMZ14]. Nonlattice [CW03, Cek03]. Nonlinear [Ano07e, CC12, Cli07, MT09, QD15, Rao06, YA00, CLM23, FLR22, GS14, Li23, Qia18]. Nonlocally [Der09]. Nonnegative [Chi12, JLW05a, JLW05b, Sen18b]. Nonnormal [Chi14]. Nonoverlapping [MR02]. Nonparametric [ACF09, EW14, FL09, Gol16, HV01]. Nonstationary [Hof07a, Lef07, Rao06, Sch03a, DL18, ZCJW19, ZCLJW23]. nontrivial [Che15]. Nonuniform [MP08]. nonzero [Att17, DF18]. nonzero-sum [Att17, DF18]. Normal [Ahm81, BX06, BE21, Coh82b, Duf04, GO73, Hal80, HS85, HK22, Hua02, Hue94, Kid02, Kid08, Lin79, Roo83, Sch01, Ton82, XY15, TKF15]. normalising [BCJW14]. Normality [GLS11, Mas84a, PS11, Jan23]. normalization [CDJR22]. Normalized [ACPP12, FP12, Sho79]. Normally [Pen98]. Normed [BH04b]. Norming [Bal77]. Norms [Wes76]. Note [Alb83, Bad77a, BN94, Bar78c, Beu83, Big87, Bue88, Cha89, CV04, Dan71, Fri87, Gre89a, Har84, HJ02, Hop85, Hor88, Hüs89, KP86, Ken88b, Li92, Lit74, Liu90, Mec89, MSV01, MS76, Mor69, Pel83, Pfe83, Qia90, Sek06, Sha03, Sig91, TH90, TC92, TW83, TW84a, WG91, Yeh88, Fri87]. Notes [Fin92]. Notion [Adl76, Adl77]. Notions [BS99]. Nucleation [Mel87, Møl95b]. Nucleation-Growth [Mel87]. Nucleotide [Wat82]. null [vdHvdHLS20]. Number [Ali90, BPS05, BX01, BR88, BS12b, BL88, Bin80, But94a, But94b, Cal03b, CS06, ENM89, FM09, Gar99, GP79, GLS01, GLS05, GLS11, Gri82, Hol76a, Hol11, IM08, Ill98, KP86, Kat90, Kem87, KP94, Mas00, MS87, MS01a, Mil69, Min05, Möh06, Moy72, Myl02, Pak00, Pen02, Pet84, Pit77a, Rat04, RS07, Ros73, Ruk02, Son79a, Wat92, Wau81, Wei95a, FN13, HL17, IMS21, JS15, Nak14]. Number-Theoretic [Bin80]. Numbers [Haj76, Pak92, Roo93, Son79b, Wad07, He16, KS16, OT20, Sen18b, HI79]. Numerical [AD78, BDD12, Bra78, De 77, Mer06, Sak04, SK85, BMS18, CMT21].

O [Agr95]. Obituary [Jeu00]. Object [HCD94, RS98, Van94]. Objects [Cow89b, FLS^{+}80, Ken95, Tan96]. Observable [Ger84b, LM03, LM04]. Observation [Pet82]. Observations [BPS05, CC12, CY99, DR81, JP98, LM04, MT06a, Pet84, Pet82, Wei78, Zir01, GLS15]. Observed [Kar80, KP11, Rud74, JLY22, PJCB18]. Obtaining [Eng84]. Occasion [DKMJ10]. Occlusion [BGR06]. Occupancy [JH92a, Lat77, RS09, Nak14]. Occupation [Haj82, Hor77, Pit77b, YZB03, Zha15, FPZ13, GR16, LZ13]. Occupation-Time [Haj82]. Occupied [Sar97]. Occurrence [Alb83, Che89a, Pfe83]. Occurrences [RS07]. Ocean [Sut79]. Odds [BL09b]. Off [DS00, JL99, Lat77, Nyr16, Ste02]. Offspring [BP79, Can74a, Hey74b, LJ04, Pak79, Ros75]. Often [BDM85]. Oldest [Joy91]. Olkin [SMAL17]. Omega [CKLP20]. Omega-killed [CKLP20]. Omission [BS88, Bal90, BMY91, Bal97, JH92b]. On-off [JL99]. ON-Periods [Ste02]. On/Off [DS00, Ste02]. One [AFH09, BFP98, BP94, Box83, BGPS87, BP99, CC09, Cov09, Duf77, GNR90, Hal85b, Hel75, HS10, HH81, Hin94a, Kar21, KW87, LY19, Mer06, Mer79, Per85a, Pit74, Pit75, Sch74a, Seg20, Sta92, TC92, Ahl15, CV18, CV23, FPZ13, MLZ21, XHW23, Zha15, BF80a, Whi80b]. One- [Mer06, Seg20]. One-Dimensional [AFH09, BFP98, BP94, BP99, CC09, Cov09, Duf77, GNR90, Hal85b, HS10, KW87, Per85a, Pit74, Pit75, TC92, Kar21, Ahl15, CV18, FPZ13, MLZ21, XHW23, Zha15]. One-Locus [Hin94a]. One-Node [BGPS87]. One-Sided [Sta92, LY19, CV23]. online [AS14]. Only [SM78, BN23]. Onset [Sim81]. Open [HS92, LY88, Mas84b, MM83, PR82, SY97]. Openings [BD97, Bal97]. Operating [CFW09, NPOTC00]. Operational [AW96]. Operator [Ahm80b, BW90a, Twe82, Wah76, BD22]. Operator-Geometric [Twe82]. Operator-Valued [Ahm80b]. Operators [APP99, Cli11, Moy72, Tjø76a]. Ophthalmic [SD98]. opinions [Hir17]. Optimal [Aba74, AVLR23, AT75, ADZ99, ADZ04, AAM17, AM01, AS14, ABS00, AGSC04, Ave83, BBHM11, BBSS07, BDM85, BB07b, BGRSC10, BSC11, Bat70, Bat73a, Bat73b, Bat73c, Bat76b, Bäu02, BCSY15, Ben74, BK80, Ber78b, BR86, BWY10, BG93, Cam84, CL87, CG08, Che98, Chi80, CIL17, CFW87, Coh77a, DDSV87, DE12, DS73, Deb76, DR16, Dos78, DW02, Dur75, ENS93, EX11, Fad74, FR11, Far76, FZ85, FS89b, FLR22, FW81, Gal79b, GV98, GKL21, Gat85b, Ger84b, GRW10, GP79, Git69, Gla83, God85, Gro78, HRY93, HS10, HV83, HK92, Hüb78, IS06, JMJ01, JS83, KP10, KK90, KK92, Ker78, Ker86, LY88, LYA07, Li23, MS01b, MS04, Mor73, MW74, NG77, ØSZ11, PT94, Pau07, RG89b, Sch99a, Sek06, Sen97, Ser81, SS92, SZ90, Sti78, Sto80]. Optimal [Tam13, Tam16a, Tho82, Tie83, Van73b, VT94, WZ19, WS87, WWY12, Wei95b, Wic77, YY22, Yeh88, ZY16, Att17, CR19, CL22, DF18, EY14, EL14, FS16a, FY18, GP15, HMM19, HS18, Lef88, LY19, Mar16, MMP22, PS83, RW14, Tam16b, ZCJW19, ZCLJW23, PS83]. optimale [Lef88]. Optimality [FOF06, FT78, FS91, GY10, HK93, JMFPY19, Käm87, KS78, Lee99a, Rob92, SSS04, Str78, BS15, GHH15, HG15]. Optimally [PvS21]. Optimisation [CM98]. Optimization [Bal78, Gat85a, Kel88, Ken88b, LH07, Lan82, PK96, Yin00]. Optimum [Rob74a]. Option [AE08, BB07b, Sch06b, Won08, EP15, GR16]. Optional [Sha80]. Options [BB07b, Bro99, CJPY97, DL11, Pri99, BV13b, DK16, JV18, PZ17]. Oracle [KPO09]. Order [Ahm81, Als88, BR93, Ben95, BENP92, Bor09, BW97, BR91, DW06, Ell87, Fei78, GL80, GVM^{+}03, Gla91, GM96, Köl81, LP93, LO04, Mat98, Mej77, Ors90, Pic74b, Pic83, Rei81, RW77, SB98, ST10, SY91, WN94, Yuh82, VBvdM15]. Ordered [BV02, Cro10, DJ91, Eth90a, KM00, MY80, SS90, WR76a, YBC20]. Ordering [AABK86, AK93, BMS89, Bar85a, BY09, Cha92a, DK83, ENS93, KM10b, LMP83, Rüs91, Sat84, Sau84, SS87, TW87b, Wat86a]. Orderings [Ano93f, Ano94e, DLU12, Koc90, Pel93, SJ89, VL07]. Orders [APP99, BB95, KLS02, Mül97b, RS92]. Ordinary [Boy78]. Organization [All82, Fri82, HMV06, Ton82]. Orientation [KPUU01]. Oriented [Rat96a]. Origin [AW87a]. Origins [TS81]. Ornstein [ALS08, BR15, BPS15, HJM^{+}16, LZ17, Sal84]. Orthant [PMG05, PMG08]. Orthogonal [SÖ11, Zäh96]. Orthotropic [Eli77]. Oscillators [MW93, Nar87]. Oscillatory [HP78]. Other [Duf04, Hor77, Lin80, MZ96, RS21]. Our [O’C95]. Out-Migration [HKH76]. Out-of-equilibrium [Vid20]. outages [Fri86, Fri87]. Outcome [BST09, LP99]. Outcrossing [Str74]. outer [He16, KS16]. Output [Dal76, Fad74, Had74, JS80, KP84, Nat75, Rub74]. Overall [FE98]. Overbooking [LY77]. Overdominant [Mar80]. Overflow [KS01b, Kos78, MB00, Sch76b]. Overlap [JM13]. Overlapping [Cha74, DSS99, Hil75, Wei95a, CL15, HLPW14]. overloaded [RFZ14]. Overly [Hir17]. Overshoot [CWW09, GM98]. Overshoots [DG03, DG04c, PM08]. Overtake [BM96, Dad82]. Overtake-Free [Dad82]. Overtaking [WV80]. Overview [Kar85a].

p [Ano89e, JL79, Lat97, RW91]. Packet [MK01]. Packets [Hoo78]. Packing [KI92, Man76, Man79, Sto17]. Page [KS78]. Pair [Bal07, Bob04, Let82]. Paired [MW94]. Pairs [BH89b, AD15]. Pairwise [JP00a]. Palm [BS03c, DRV00, Las06, Miy04, RSS99, Tho07, Whi92a]. Palm-Duality [Tho07]. Palm-Martingale [Miy04]. Palm-Stationary [RSS99]. Panmixia [Ste76a]. panne [Bré75]. Paper [Coh81, Get74, Jof77, Kon89b]. Papers [AAB^{+}92, BBM^{+}88, BGH^{+}88, EFG^{+}92, RSS94]. Para [Pak98]. Para-Critical [Pak98]. Parabolic [VJV10]. paradigm [AV19]. Paradox [Tho76]. paradoxes [DS23]. Parallel [ADZ99, AT81, BW93, BENP92, CNP91, GV98, HJ04, HK93, Hun79, KKT06, Lom72, SS92, Vit95, Wei95b, Wri92, Xu92, AHO20, TR19]. Parallel-Processing [GV98]. Parallelogram [Hol90]. Parallelotope [Mat99]. Parameter [Blu84, Duf77, Ger84b, Le 78, Leh85, LM04, LC92, Mer06, Ros81, Van85, Xu11, CM17, MS80b]. Parameters [BT71, Ber78c, FH02, JP98, MS94, RC94, San98, Sch71, Ste84, Tho81b, Wei78]. Parametric [CSY98, DMC00, DR81, Hea78, Kid01, LL98, Mas83, Par77]. Parametrisation [Kos78]. Parasite [BP00, AG13]. Parasitic [Lew75]. Parent [Man94, Ros75]. Parent-Offspring [Ros75]. Paretian [RW85]. Pareto [EK12, RW03]. Parisian [CJPY97, WZ20]. Parity [SR82]. Parker [Gar80]. Parking [Yin23]. Parsimony [SGW96]. Parsing [Edw79]. Part [Her73, Her75, Bat73a, Bat73b, Bat73c, Dag75a, Dag75b, KR71a, Mil69, Mil71b]. Partial [ARW13, Ano93f, Ano94e, Bal77, Bin73b, CJ11, ENS93, Fad79, GM98, KD92, Koc90, KS12, Mij80, NM01, Pel93, Pol87, SJ89, TV75, Tuc71, Whi86, Yao10, Yin00, HS18, LLX18]. Partially [Cre74, Gue93, Had80, Kar80, KP11, LM03, LM04, Per85a, Rud74, JLY22, PJCB18]. Particle [CB03, CLM08, CDL99, DLS04, Dur79, Dur80, EC75, FH74, GKP84, Jac74a, Kid98, Li98, Mas95, MW78, Moy72, ON98, Sch01, Sud93, Wei95a, Wei97, Wu93, ZÇ96, BCJ^{+}17, CHJ16, CM17, FX13, JY20, Sto23, US19, YS23]. Particles [Adl89, Hol83, Lep76b, Mer79, PS90a, HTO^{+}15]. Partition [Ala99]. Partitioned [BL01]. Partitions [Las06]. Parts [Tuc71]. Pass [Ano06e, ACDHT05]. Passage [BH04a, BOR02, BVH10, BNR87, Cox80, Dad82, Dad84, Der83, DNPR01, FL01, GNRS89, GNR90, Gut74, Jan86, Kes80, KW03b, Lal84, LL97, Leh02, McD83, MW96a, Pim07, ST96, Sha84, SW77, SW78, SZ11, Van83, Whi95, Wie77, Yuh82, Ahl15, AV19, But22, DL20, HNGS15, KK15, KZ16, Kub19, Lef88, MP18]. Passage-Time [MW96a]. Passengers [BBSS07]. Passing [Lep76b]. past [FZ13]. PASTA [KS90]. Patch [BBZ10, MP10]. Patch-Dependent [MP10]. patches [Lan13a]. Path [Ano85e, AA88b, AB94, DS85, Dan89, DM03, FG00, GP80b, Kal80, Kes80, Sig91, STC94, VGS10, HNGS15]. Paths [Ano92e, BTZ00, BBHM11, Dad82, JKS07, Pit86, BCT13]. Pathwise [SSS04, SJ18]. Pattern [Bar74b, Bar74c, Jan23]. Patterns [AK08, BT98, BBB^{+}76, CAK00, Che89a, DMC00, FJ09, God91, GS93, HBG99, MR02, Tan96, VJ99, WDH98]. Payments [Pau07]. Payoffs [GKL21, GHL07, Lin15]. PCR [LJJ04, Pia01]. PDE [Kne00]. Peak [BZ03]. Pedestrian [Gri80b, AD15]. Pedigree [Fie76a, Tho79, Tho82]. Pedigrees [CTS76, CTS78, Ste75a]. Peeling [DNP10]. Penalties [VT94]. Penalty [Gla83, Han07, PP07]. Pentultimate [Coh82a]. Penultimate [Coh82b]. percolated [LM17]. Percolation [AFH09, BBD02, BVH10, Cox80, DGP10, Der83, GLM21, Gri81, Häg00, Hou18, IY12, Kes80, McD81, McD83, MT11, MSV01, Mol74, Pen91, PP96, Pim07, QT97, SW77, SW78, Van83, Van84, Wie77, Wie81, Ahl15, Ber20, CKM19, Cza23, LBCEN21, HNGS15, JT22, KW20, KK15, LSV18, MP18]. Percolative [Geo84]. Perfect [BK91, BD15, CK15, Dai08, Gal11, KM00, MR05, Thö99, BMM13, MJM21]. Perfectly [FFG07]. Performance [BM98, BL09b, Cao95, Cao98, DG01, Hil76a, Iye88, MY80, Nat89, vdHK04]. Performing [Mer79]. perimeter [KR18, TSvB17]. Period [DT80, Hau78, PR06, Ram84, Ros73, Sat85, Tru75a, WW03, Nea14]. Periodic [Fal89, KM80, Pro76a, Pro76b, Rol87, Wil98, Woe85, MT13, Oça23, RSY23]. Periodicity [Çin74]. Periods [AB94, GG12b, Ros76a, Ste02]. Perishable [KP83a, KP84, Per85b, PS01]. Permanental [MM06]. Permutation [LY15a, Pit96]. permutations [Die82b, Jan23]. Perpetual [BB07b, Won08]. Perpetuities [GG96, KM14a, PH99, Kol18]. Persistence [BCHK18, Hai80, Sak15]. Persistent [Gre89b]. Person [And75, Fed78, Fei76a, Phi78, Pon05, Mon17]. Person-Flow [Fei76a]. Personalized [VL10]. Perspective [Wil88, Ner18]. perspectives [DP19]. Perturbation [AANQ04, AFGR06, BKZ95, Dan71, HV84, HLLP16, KKSS90, Lat88, MH97, RS88, Sen88, VP88, Van88a, Wei71, LL18].

Perturbed [AG05, AL99, CY05, YZB03]. Peter [Ner18, Sen18a]. Petersburg [GML18]. Petkantschin [Jen96a]. PH [PR00, Sen89b, Tak81, HHHT19, Neu86c, Tak81]. PH/G/1 [Neu86c]. PH/PH/c [Tak81]. PHA [Ano04e]. Pharmaceutical [WG91]. Phase [Abr79, BFL04, JS83, JW15, Lan05, Lat88, MK01, MSM92, Neu71, Neu76a, NPOTC00, O’C91, Ott87a, Per85a, Rem00, Sar97, AZ18, GLM21]. Phase-Type [BFL04, Lat88, MSM92, O’C91, Rem00]. Phases [KM12, Lan82, Mah10, MT06b]. Phenomena [Baz74, BFK92, FR97, Rub78, Shi82, Teo82, Wac09, LSV18]. Phenomenon [MK01]. Phenotypic [DMB11, BE21]. Phragmén [Ros78]. Phylogenetic [FS93]. Physical [And79]. Physics [Ano83e, Rub81]. Pick [BGPS87]. Pick-up [BGPS87]. Piecewise [AS13, BL08b, BDD12, FLZ12, GT12, LL98, Rol78, Sel96, LTT20]. Piecewise-Deterministic [BL08b, BDD12, FLZ12, GT12]. Piecewise-multilinear [AS13]. Piles [Die82a]. Pinned [CP08]. Pitman [FGZ17, Imh92]. PL [Ros78]. Place [LL89]. plan [Par98]. Planar [BC16, CC96a, CC96b, Gat94, Gat97, KJ03, LOS03, LO04, NW98, NW03, Rat96a, Rem00, WN94, WN99, CD18, Cow13, JV18]. Planck [Sac90]. Plane [BBB^{+}76, BM11, Cal02, CI00, DRS11, DM91, EC75, HQ00, HQ02, HS82, Sch05, ST10, She85, Wak85, Wei95a, WB71, Zak77, Par98, TC13]. planning [CL22]. Plans [Ker78]. Plantations [GW78, Wes79]. Plants [Heu79, Heu80]. Plates [Eli77, GKRH06]. player [FPP19, HMM19]. Playing [HK88, PR22]. Plenary [Fie76a]. Poets [Had79]. Point [Ano77e, Ano93e, ANR02, ACS93, AG81, AS96, BB93a, BT98, BV97, Ber78d, Ber78e, Bor98, Bré76, BM02, BK02, BN92, CDDP11, CFD85, CS97, CB99, CM03, Çin78, CQZ03, Cow06, DO74, DS93, DRV00, Dav77, DNP10, DN02, DMC00, Dob77, Ell86, Ell87, Fei78, GL80, GM96, Gra77, GL04b, Gri77, HJVN03, Här81, HPJ08, HW09b, JL77, JL73, JP00a, Jan19, JK94, JN98, Kar85b, KM00, Kid96, Kim83, Kön77, KW88, Kri78, KT92, Le95, Lea76, Led79, Lie76, LR80, Mac75a, MD79, MC09, Mec77, MSS95, MS10, MB12, NRT83, NR85, Pap77, Pie78, RC94, Rem00, Res80, Res86, Ric75, Rip77, Rud73, SC83, San98, Sau78, SB98, SX08, SCA00, SS95b, Shi80, Str78, Thö99, Tod77, Van94, VJ99, WG83, WW84]. Point [WDH98, Wis72, ARS17, AB15, AF18, BCM18, BY14, DP22, HRW18, HT17, JT22, NP17, Pri15, QDZ19, SS14, vL18, BJ77]. Point-Based [ACS93]. Point-Measures [Gri77]. Point-Process [Ell87, GM96, Kim83, MSS95]. Points [Ano06e, AR97a, AR97b, ACDHT05, BLP06, BM00, BM11, Cal03a, Gec77, HJ03, HSS08, Hen87, Hol90, IJ02, KK80, Liu98, Mij78, Pen98, Pic85a, Pit77a, She85, SK20]. Pointwise [Ais79]. Poisson [AAL08, Alb83, AK08, AS08, Alm97, ARS17, BTZ00, BB01, BCLF16, BB92, BE83, BH89a, BU98, BM00, BNX02, BWZ19, BL07, BL09a, BM16, Ber87, Beu83, BVH10, BE06, BS03b, BT08, CY05, Cal02, Cal03b, Cal03c, Cek97, CW03, Cha09a, CCP91, CK07, CH16, CD18, CVS96, CQ01, CGR18, Cow89a, CQZ03, Cow06, DSS99, DRS11, Dav77, DNP10, DTV84, Dwa74, ENR17, EJMT85, FS09a, FG81, FZ96, FS82, LBCEN21, GW20, Gar95, Gly94, God91, GS93, GL04b, GM79, GY17, Had74, HQ00, HQ02, HS13, Hof07a, HH14, HRS04, HS11, HS14, HLPW14, Hwa99, IJ02, Iso00, JP00a, JVZ08, Kar80, Kar86, KP84, Kem73, KS01a, Lan87, LPST14, MSM92, MN18, MS80b, McK88, Mec95, Mec96, Mec98a, Mec98b, Mel79a, Mel79b, MRV02, MP03, MV19]. Poisson [Mil69, Mil71b, MM96b, Mil98, MW72, Mil74, MYH18, MZ98, MZ96, MS10, MB12, Mrk01, Muc05, Muc10, NP17, Now88, Ols07, PP02, PP07, PBM00, Pri15, RR90a, RC94, Roo93, RS07, Rud72, SKF80, Sch74a, Sch99b, Sch03a, Sch09, Ser89, SJ89, SV76, Str78, Teu79, VPCG10, Wes74, Wil90, Xu11, Yao10, Zaz92]. Poisson-Delaunay [CD18]. Poisson-Driven [BT08, Zaz92]. Poissonian [Mat72, Mat74, NR06, Par98]. Poissoniens [Mat72, Mat74, Par98]. Pol [Nar87]. Poles [SM87]. Policies [Aba74, AH87, Agr95, Ala99, BBSS07, Bat76b, CG08, DPY08, DG01, FZ85, FS91, Gal79b, GNMA02, GKRH06, GHK13, HV83, Hüb80, IS06, JK85, Käm87, Mi94, RG89b, Rob92, Sen97, SS95b, VBvdM15]. Policy [JMJ01, Kha86, Lee99a, McN90, WW91, Bar18, HHBV20]. Polish [GHL07]. Pollaczek [De 76, Neu86b]. Polling [Tak91b, VL98, HHBV20, RFZ14, VBvdM15]. Pollutant [Gra82]. Pollutants [Ste84]. Pollution [KX94]. Pólya [AC22, Bin79, Cha14, Cha22, EJMT85, Jan20, KM12, Pfe83, RW77, WG83, ZHCC14]. Polydisperse [Spo84]. Polyèdres [Mat72]. Polygamy [Can74a]. Polygenic [Vet78]. Polygon [Gro12, HQ02, MM96b, Mil98]. Polygonal [ACS93, Sch05]. Polygonally [KL87]. Polygons [Bia02, HO09, Mec98b, Mil95, Wak85, FKV14]. Polyhedra [Mat72]. Polyhedrons [OM95, ON96]. Polymerisation [Whi80b, Whi80c, Whi80d]. polymorphism [DAG95]. polynômes [Gal84]. Polynomial [Voi90, XHW23]. Polynomials [HS90, LP90, TW84b, Gal84]. Polytope [Mec95]. Polytopes [BFV10, CS06, HR05b, Küf94, O’C91, HS14, KZ20, MN18, TW18]. ponctuels [BJ77]. Pooling [Law92]. Population [Ave79, BT71, BL01, BS12a, Bar74a, Bar76b, Bar80, Bar84, BBZ10, Bor98, Bou16, But94a, But94b, CW75, Coh85, Dal85, DG04b, DMB11, EN80, EN88, GT84, Ger95, GJS08, Hil76b, Hil81, HT10, KK73, Kes70a, Kes70b, Kin79, Kle84, Lip75, Lom78, Mar80, Möh98a, Möh98b, Möh00, Möh04, NM79, Nor82, Oka79, ORT16, Ren86, Sat78, STW84, Saw76, Saw77, SS84a, Sma76, Sve93, Wat78, AD21, AG13, BHKK15, BSS17, FHJK20, GM14, GHL23, SWL14]. Population-Size-Dependent [Kle84]. Populations [Ave76, Ben75b, Büh74, Coh77b, Coh77c, Esh73, Hai80, HRY93, Heu79, Hil74, Hil75, Hil76a, Hin94a, Hin94b, HKH76, Hio79, Hol81, Hös11, JN84, Möh06, Moy72, Pak00, Rob76a, Sta80, Wau74, WH03, Woo74, CN18, Cor14, Kal18]. Portfolio [BWY10, BW96, Bro98, Bro99, GPD00, IS06, LH07, PK96, Tam08, CIL17]. Portfolios [Sek06]. Position [Bac00]. Positions [DM00]. Positive [AK98a, AK98b, Cli11, Eme73, FK89, Jan86, Ken74, KR02, PMG05, Rog90, Sad95, Sch07, Tuc71, FJKR20]. Positivity [BGRSC10, LL97]. possessing [Cas16]. Possibility [HP91]. Possible [Hin94b, Mol72]. Post [Gor01]. Post-Data [Gor01]. Potential [Sys77, JJL21, KP22]. Potentially [AHO96, CAR20]. pour [DM90, Par98]. Powder [Pes82a]. Powder-Product [Pes82a]. Power [Alp09, BM02, BMR05, Hol89, JOC10, JOC12, MH08, Nyr05, Ros13, KL13]. power-law [KL13]. powers [Ath18, JT22]. Practical [BT98]. Precedence [BL89, Gla96b]. Precedence-Based [BL89]. Precise [Cal03c]. Precision [Kiê98]. Precludes [Mih80]. precommitment [CL22]. Predation [McN90]. Predictable [Gri77]. Predicting [BP23]. Prediction [And79, ACO98, DR89, Fin69, GACO02, Lin78, Lin79]. Predictions [BP21]. Preempt [KS83a]. Preempt/Resume [KS83a]. Preemptive [GHT88, Nai86, Wei95b, HHBV20]. preferences [FLR22]. Preferential [JSS19, Jor10, AH16, BK18, GS19, JW15, PRR17, PP15, Ros13, WR22]. premier [Lef88]. Prescribed [Ano86e, DM06, Fra85, Ruk02]. Presence [BR87, Cre74, HK90, TT04, EC21, RT23, Cur81]. Presence/Absence [Cur81]. Present [Cha99a, Cha99b, DWH^{+}99]. Present-Day [Cha99a, Cha99b, DWH^{+}99]. Preservation [APP99, Ano93f, Koc90, MSM92, SJ89]. Preserve [Cam88]. Preserving [Pol87]. prevent [Hir17]. Preventive [Ger84b, GV76]. Price [EX11, BV13a, RT21]. Prices [GW00, Kel88, MS11a]. Pricing [AW79, AE08, BV13b, MAF07, PR94, ACH21, GR16]. Principal [Cal03c, HZ06]. Principle [Ant79, Bor79, Edw74, Gla88, KT92, McD85, Rol78, Eck19, EY14]. principles [HJKP16]. Prior [Gla93c]. Priority [BY89, DN78, FX00, GHT88, Hau78, Käm87, KS83a, Ken79, Nai86, PT94, Sch74a, Sti72, Wei95b]. Prisms [Cow96, Cow97]. Prisoner [CC09]. Probabilistic [AFGR06, BL22, BMM13, Cas16, CJ11, De 76, De 80, FM88, GV76, JP00b, Lin77, MT13, MY11, RS95, Stu88, Vit95, HMM19]. Probabilities [Ano80e, AP88, ASS99, BH04a, BS95, Beg12, BC81b, CS06, Cao89, Che98, CH76, CLW95, Cho93, DGN92, DM05, FJ09, Fuh97, GLM93, GW13, HL85, Hal80, Har71, HV84, HT92, Kel86, KS01a, Las06, LN92, LS89, Mal00, MP10, MS01b, MS04, Mor94, Nyr05, Pol74, Pon05, ST96, Sch78a, Sch86b, Sch76b, SRTM97, SH89, Smi76, Sta98, Sze76, TT04, Tho79, TVF81, Wei82, Yuh82, BB17, HLN13, KZ20, Nak14, PB19, TKF15, Yua22]. Probability [AW96, AS07, Ano13a, Ano14a, Ano15a, Bad77a, Bad77b, Bet98, BR76, Buk01, BNR87, CTS78, Cao87, CP08, CLCZ12, Cor78, Cou90, CL00, DH97, DMPG09, Don81, FFI22, Gal79a, Gan78, Gan08, Gat85a, GNRS89, GZ93, Gue93, HP82, Hel75, HW09a, HR00, Igl73a, KR81a, KR81b, Kel12, KS01b, Kol14, Kot80, KS80, Kwe75, LZ05a, LZ05b, Lit74, Lit69a, MY03, Mec77, Mer79, MS76, Mor69, Mül97a, Neu78, Pro76b, RR95, RL10, RS86, Ros78, Rud74, SS95a, Sad95, Saw77, Sch74b, SW13, SG92, Sta73, SM87, Tam11, Urb77, Van88b, Wol80a, Yun86, BFP23, Jag18, Rob18, Sak15, SS14, XY15, Ano14a, Ano14b, Ano15a, Ano15b, Ano89e, Ano13a, Ano13b]. Probability-Generating [Mec77]. Probes [ACO98]. Problem [Abr79, Agr95, AS79, ARM99, AS73, ASC96, BSC11, Bat76a, BL04, BR88, BWY10, BLT03, Cha96, CS91a, CCH00, CM93, CO98b, Dag75a, Dag75b, DS93, Dan74a, DTV84, DP12, EDH90, Eme73, FOF06, FV95, Gla83, Gla96b, GGV89, Hal86, HS82, IL73, IL74, Kin77, Kne91, KS04a, KSC09, LY77, Lor79, MMC91, NM79, Odo80, Pap77, Rog90, Ros79, Ryc92, Sen78, SSY91, SY90, SY81, SZ11, Sta90a, Sta98, SW92, Tau75, Van74, Wau81, Won08, Yeh88, AF18, CKLP20, DM90, EL14, FY18, HS18, JMFPY19, KMF22, Li23, MVY19, Nak14, PB19, Tam13, Tam16a, TK13, YS23]. problème [DM90]. Problems [Ass78, Bal78, Bat70, BS95, BK83, Ben74, Ben75a, Ben98, Ben91, Big76, BLT03, BC04, CAK00, CC99b, DPY08, DE12, DNPR01, DG01, FR11, FT78, Gan78, GS13, GKL21, Gla91, GRHK06, HS74, Hüb78, Hüb80, KR85, Käm87, Kar83, KR71a, LYA07, Leh85, LT73, MY99, Nag71, Neu75, Pet84, Pet82, Sam04b, Sch74b, SF79, Sie79, Tak76a, Tak76b, TJM98, Van73b, Wan01, AV19, EC21, EY14, GLR15, JM13, LS21, LY19, Mar16, PHL16, RW14, Tam16b, Hog86]. Procedure [DNP10, Fad74, PS94b, CP15]. Procedures [Bat73a, Bat73b, Bat73c, KS75]. Process [Agr74b, Aks75, AP98, Alb83, AP05, Alm97, Ano93e, AG81, Ath79, BB01, BP79, BK83, BK80, Ber82, Bet98, Big76, BB93b, Bin79, BD74b, BD74a, Bor98, BDD12, Bro85, BN92, CY05, Cal03c, Che89a, CK07, CM03, Cra74a, Cra74b, D’S94, DH97, DH82, Dal85, Dan77, Dan89, DZ11a, Dav77, DNP10, Dur75, Dwa74, Ell87, Eme73, FQVJ71, Fal84b, FS09a, Fil72, Fil74, FGM12, FZ96, FS82, Fuj79, Fuj80, GT84, Gat97, GT12, Geo76, GT75, GM96, GZ11, GLM90, GZ93, GP88, HS91, Här81, Har77, HR78, HP78, HQ02, Hel75, Hey74b, HW09a, Hsi10, Hsi87, HP91, Hui04, Ill01, IL97, JP98, JL77, JL79, Jac05, Jac80b, JH92a, KL83, Kar80, Kei74a, KR71a, KS83a, Kim83, KM80, KW03b, KT92, LJ04, LJJ04, Lam03, Lat97, LL89, Li92]. Process [LG05, Lie76, Lin80, MS87, MM06, MP08, Mec89, MNW08, Mel87, MP03, MSS95, MZ98, Nar91a, NS93, Neu70, Neu76b, Neu86c, Neu87, NZ77, NR06, OC10, Ott77a, Pak71a, Pak71b, Pak79, Pak86, Pak89b, Pak92, Pak98, Pak10, Pfe83, PvS21, Pit74, Pit75, PS10, Qui72, RS95, Ric75, RW77, Ros80, Rud72, Sac90, San98, Sau78, SB98, Sch01, ST10, Sch07, Sch82a, Sen69, Sha78a, Sha78b, Sha79, Sie76, Spo84, Sri85, Sud99, SS88b, Tak75c, Taq75, Tay77, Teu79, Tho81b, Tod76, Tod77, Van85, Ven82, Wau81, WG83, Wen89, ZG88, AB22, BGV22, BP23, BR15, BD17, BPS15, Cha14, CP15, CDIM13, DP22, EKKZ18, FG81, FGZ17, Fro13, Gri14, GR16, HRW18, HT17, IMS21, JKR17, KV18, KM14c, MMZ22, MS15, MS16, NP17]. process [OT20, PHL16, STZ14, SS14, TO21, US19, Kon89b]. process-level [OT20]. Processer [Coh78]. Processes [AWZ93, AP97, Agr74a, AK05, AM01, Alp09, AK95, AK79, And99, ALS08, Ano77e, Ano79e, Ano86e, Ano98f, ABH80, Arj72a, AS73, AA85, AA86, AK78, Asm82, AP88, ASS99, AK00, AN78a, AS96, BB93a, BVM96, BD87, BS88, Bal90, BMY91, Ban09, Bar74a, Bar75, Bar76b, Bar80, BS81, BNJS98, BNBJ00, BE88, BBB^{+}76, BP83, BH04a, Bas79, BLS^{+}80, Bat78, BG08, BL09a, BL10, BV97, BBR09, BOR02, Ber77, Ber78d, Ber78e, BW83, BC89a, Beu83, BR86, BA81, BC87, BHJ11, Bin73b, BD74a, BD75, BR83, BO07, BL08b, BG93, BV90, BFL09, Bré76, BM02, BMR05, Bri99, BK02, BGR82, BRT82, BM08, Bro73, BD93, BK05, CK73, Cao95, CP08, CDDP11, Cav78, CSY98, CI00, CFD85, Cha92b, CORR91, CS97, CZLR04]. Processes [CLCZ12, CB99, CB00, CS99, Çin75, Çin78, Col74, CC77, CQZ03, Cow06, CI86, CP88, Cse92, DO74, Dal76, DRV00, Dav73a, DR89, DSSW11, De 10, DZ11b, Der09, DNPR01, Dia13, Die06, Dou75, Duf77, DHP12, Ebe80, Edl78, Ell86, EJMT85, Esh72, Fad90, Fas05, Fei76b, Fei78, FR12, FG00, FV99, Fou11, FS79, Fra85, FLZ12, GS78, GM71, GS13, Gar95, GL80, GB03, GTY70, Ger95, GNR90, GAH03, God85, GH80, Gra77, GL04b, Gre75, GS77a, GP80a, Gre89a, Gri80a, GM11, GM79, Gri92, GGV89, GHL07, GY10, GW13, Gzy76, Gzy80, Hah98, HS98, HJVN03, Hal86, Här81, HMS03, HPJ08, Hen83, HPPT92, Her73, Her75, Hey74a, Hey79, Hey85, HL05, Hir00, Hof07a, Hof07b, HH85, Hol76b, Hop85, HV83, HK93, Hor77, Hor91, HT89, Hua90, HW09b]. Processes [HLW03, Hui03a, HL06, Hun69, Hun84, Hun85, Hut80, HT10, IPV05, Ito79, Ito80, JL73, Jag77, JH92b, JP00a, Jan19, JL99, JN98, Jir74, Jir76, JM86, KS09, KK73, KL79, Kar85b, Kas78, Kas80, Kei78, Kei74b, KW87, KVB99, KM00, Ken78a, Kes76, Kha80, KS01a, Kid96, Kid98, Kid01, Kij92, KNPV97, Kim83, Kin73, Kle84, KRS07, Kom96, Kön77, KW88, KR02, KKSS08, Kot80, KK03, Kri78, Kri79, KS75, KP11, LY88, LL91, Lam93, LS96, Lea76, LR97, Lee79, LP90, LP96, Leh02, Leh85, LRS10, Lev77, Lev79, Li98, Li09, Lin78, Lin79, Lin80, Lin74a, Lin74b, Lin92, LMHN90, Mac75a, Mac75b, MD79, MNR02, MS80a, MS83, MM98, Mat98, Mat77, McD85, McK86, MDB^{+}75, Mec77, MT06a, Mel79b, MW96a, Mer06]. Processes [Met80, MT92, MT93a, MT93b, Mih80, MW72, Mil74, MVY84, MY89, MY01b, Möh00, MZ96, Møl03, MT05, MR05, MH08, MS10, MB12, Mol74, Mol78, Mor73, MS11b, MT06b, Mrk01, NM79, Nat75, NS77, NRT83, NR85, Ney74a, Ney74b, Nij74, Num76, Nyr98, O’C95, Off74, Oka79, Orn77, Oso10, PT81, PT85a, Pak89a, Pak93, Pak95, Pap77, PM08, Pau07, Pea89, Pes82a, Pia06, PL90, PL96, PL99, Pic74a, Pic74b, Pie78, PR99, Pol74, Pol86, Pol91, PBM00, Pru78, Pyk79, Qi08, Rah07, RPS01, RL10, Rao83, Rao06, RC94, Rem00, RR73a, RR73b, Res74a, Res74b, Res75, Res80, Res86, RR90b, Rey75, Ric11, Rie80, Rip77, Rob07, RS09, Rog90, Rog92, Rog94, Rol78, Roz80, Rud72, Rud73, Ryc92, SS09, Sal84, Sam04a, SC83, Sav75a]. Processes [Sav75b, Saw76, ST85, Sch71, Sch76a, Sch78b, Sch96, Sch00, Sch07, SB77, Sch78c, Sch79, SX08, Sch81, Sch77, Sco73, Sel96, Sen74, Sen89b, SCA00, Ser75, Ser81, Ser86, Ser89, Ser05, Sev71, SS95b, Sha84, Shi93, Shi80, Sho74, Sim80, Slu84, Smi74, Smy76, Spo01, SR82, Ste02, Ste84, Ste80b, SVW83, Sti72, SW80, SBT96, Str78, SK85, SW97, Sze76, Tan81, Tav78, Tay89, TM91, Thö99, Tho83, Tjø76b, Tjø76a, Tjø83, TF08, TT79a, TT79b, TT79c, VP88, Van88a, Van80, Van91, Van94, Van78d, Vit01, VML80, Wal74, Wan97, WS10, Wee02, WW84, Wei97, WN94, Wes74, Wes78, Whi81a, Whi69, Whi80b, Whi80c, Whi80d, Whi83, Wij78, Wis80, Wis72, WHZ10, YZ01, YA00, Yeo78, Aaz13, Ale13, ARS17, AB15, Ano09e]. processes [APZ21, AF18, BMM16, BR13, BHKK15, BPP21, Bau14, BB17, BM16, BOS23, BCM18, BDK^{+}18, BY14, BS15, Bou16, BJ77, BLT19, Buc16, But22, CPP21, CMB19, CG16, CM19b, Cha22, CH14, CMT21, CF14, Clo17, CKO19, CGMT20, CLM23, CHY20, CKLP20, DS16, DL20, DP23, DSS17, DSS18, DP13, DP15, DM22, EMOS14, FHJK20, FFI22, FMY22, FJKR20, GP15, GW20, GLL22, GVZ13, GHH15, GKPZ21, HLN13, HS13, HL17, HJM^{+}16, HLG18, Hue16, HS14, ILP18, JT22, Jan18, JV18, Jir13, JS14, KWRK22, KM15, Kol14, Kom18, KJH18, Kur18, LZ17, LTT20, LP21, Liu97, LY15b, Luo16, MS17, MZ16, MN18, MYH18, NB15, Oça23, OS18, Par98, PV19, Pia01, PSS20, Pri15, Pri22, QDZ19, RSW16, SS15, SJ18, Sak15, WZ19, WZ20, dlI22, vL18]. processes [Bez20]. Processing [BW93, GV98, GRW10, GW92, HKT91, Sch90, PWZ17]. Processor [BMV03, BB10, CFGW87, Gri92, Gri94, GRZ04, MM83, MM85, Sch84, TW83, Xu92]. Processor-Sharing [Gri92, GRZ04, MM83, MM85, Sch84, TW83]. Processors [CNP91, FR80, HK93, RX91, Wri92]. Processus [BJ77, Pia01, FG81, Liu97, Par98]. Procrustean [Le96, Le98]. producer [Lan13b]. producer-consumer [Lan13b]. Producers [Mit88]. Product [BV91, DTM07, DTV84, Eco05, FS89a, Miy93, MT97, Pes82a, PMG05, PMG08, RT90, YM98, YS94, Mas23, XCWC18]. Product-Form [Eco05, FS89a, Miy93, MT97, PMG05, PMG08, RT90, YS94]. Production [DTV84, Bar18, PL89]. Production-Inventory [DTV84]. Production/Inventory [PL89]. Products [ALS08, BK98, Mai97, MNR02, Max77, Fra89, He16, KS16]. Produits [Fra89]. Professor [Mon95]. profile [ZM16]. Profiles [DH05, Stu88]. Progeny [Cur81, Pak71b]. Program [Ros81]. Programming [Ass78, CG80, FR97, Gla88, Ker78, Rie75, Ros74a, Sch73, VV78, Wes76, Mar16]. Progression [Bar85b]. Prohorov [Pol80]. Projected [Kid96]. Projection [LB10, Mil80]. Projections [Gar95, Gri77, Mil75, Nie98, Nie99, Rat04, Zäh96, BPP21]. Projective [GS14, MGW96]. Proliferation [BLS^{+}80, Jag77]. Promotion [Bar79]. Proof [Ber85, CN86, Dro89, Imh92, Jen96a, Lin77, Mil95, Pol80, Rog89, Van90, Vet80, Wil71]. Proofreading [Tie83]. Propagation [HR04, Nij74, Sut79, BDK^{+}18, Bez20]. Properties [APP99, AR92, BVM96, Bal75, BR93, Beh11, BOR02, BD74b, BD74a, BD75, BH89b, BB96, Con77, CT94, CFPL12, DR89, DW06, EC75, Fin80, Gar99, Gha08, GL79, GT75, Hal85a, Hal86, HJ04, HJ08, HS92, HT89, KD94, Kij92, Koh78, LR11, Lan05, LS98b, Las04, LW92, Maj07, MNR02, McG82, Miy04, MS94, MS11a, MZ96, PT85a, Pak86, Pic74a, Pic74b, Pol75, RC94, RS12, Roy07, SK96, Saw77, ST10, SS84b, Sha78b, Sha80, ST04, Str78, Tam74, TE96, Urb77, ZLB98, Zuy99, AHO20, BY14, FPS22, GC20, HS14, KXZ23, PJCB18, ZHCC14]. Property [AK05, Che89a, Eba78, Fei78, Gly82, HK89a, Hua90, KT01, KS90, LC08, Num79, POGP95, Wal74, HHHT19]. Proportion [AMS07, HQ00, Ken74, Pro76a]. Proportional [Bro98, IS06, Pau07, Sha87, Tam08, BFP23]. Prorok [Ano76e]. Protocol [FFH86, Ott06]. Prove [Ran77]. Proving [BH89a]. Proximate [Pak95]. proxy [MT16]. pruning [KXZ23]. Pseudo [Maz89]. Pseudo-Modular [Maz89]. Pseudolikelihood [BT98]. Pseudopolynomials [LP99, PL96]. Pseudosystematic [GACO02]. public [LT15]. pulsating [RSY23]. Pure [BK83, Dro89, Dro86, Aaz13]. Purpose [Ros81].

q [JL79]. QBD [DSSW11]. QED [DVN15]. Quad [KW03a]. Quad-Trees [KW03a]. quadrant [FFI22]. Quadratic [DMY93, Kes70a, Kes70b, Sur08, Whi81b, Luo16]. quadratures [SK20]. Quadtrees [CJ11]. QuanTA [TR19]. Quantile [LY12, SL80, TR19]. Quantiles [DE08, Wei78]. Quantitative [AH78, FGM12, Hil76a, HA79, Kit71, Smi79]. Quantities [Jan86, NW98, PT85b, WN99]. Quantum [Cor78, KR97, Moy72]. Quarante [EL05]. Quasi [Ano93g, Bar76b, BL10, BC89a, Bob04, Cav78, CM00, Daw75, Fal84b, FMP92, HPPT92, MPS98, MT06b, Nås96, NP93, Pak95, PS94b, TW83, Van91, Wij78, BGZ16, BFW22, CKO19, GJRS17, Mas23, Oça23]. quasi-algorithmic [Mas23]. Quasi-Birth-and-Death [BL10, BC89a, MT06b, CKO19]. Quasi-Compactness [Wij78]. Quasi-Deterministic [Daw75]. Quasi-Input [Fal84b]. Quasi-Reversibility [CM00]. Quasi-Reversible [HPPT92, TW83]. Quasi-Stationarity [Van91, Oça23]. Quasi-Stationary [Ano93g, Bar76b, Bob04, Cav78, FMP92, MPS98, Nås96, NP93, Pak95, PS94b, Van91, BGZ16, BFW22, GJRS17]. quasispecies [Cer17]. Quelques [Par98]. Queries [CJ11]. Quermass [Der09, KVB99]. Quermass-Interaction [Der09, KVB99]. Queue [AW88, ADZ99, AH05, Ano04e, AFGR06, Asm82, AT81, BMS89, Box83, BK01, Coh74, CH80, Dag75a, Dag75b, Dag78, De 73a, De 74, De 76, De 80, De 83a, DW06, DKS84, Dsh85, DS00, Fal84b, Fal88, Fer93, GHT88, GM08, GY78, Gly94, Gri94, GGV89, Had74, Hal85b, HP96, Hau78, Hok79, Hok80, Hok85, HK92, Igl71, KS83a, KS06, KD08, Köl81, Kuc83, Kyp71, Lat77, Lat88, LPRS93, LK93, LMHN90, MA95a, Min80, MZ04, Neu71, Neu76a, Neu77b, Neu77c, Neu86c, Ott77a, Ott77b, Ott87b, Ott87a, Par87, PR00, Ram84, Ram80, RG84, Rei88, RW91, Ros73, SS95a, Sad95, Sat85, Sch84, Sch90, Sch87, Sen89a, Sen89b, SG92, SK73, SK08, ST90, Sti72, Sys88, Szc90, ST94, Tak81, Tam74, TVF81, Tru75a, Tru75b, TW83, Van88b]. Queue [WW03, YC96, Zwa00, Bré75, DP19, DS15, Fri86, Fri87, Gol16, Miy17, XZ20]. Queue-Length [DKS84, Sch87]. queue-size [XZ20]. queued [XZ20]. Queueing [AL94, AANQ04, Ano04f, AR92, ASS99, BL89, BW89, BW90a, BM04, Bäu02, BW83, BM96, BM77, Bor06, BV91, Box78, Box79a, Box79b, Bra78, Bro05, BP82, Bur81, BSD^{+}84, Cao87, Cao89, Cao90, CM00, Coh72, Coh73, Coh77a, CC77, Con77, CDH87, Dal76, Dos78, Eco05, FE98, Fal84a, FS91, FK89, Gal79b, GV98, GNR86, GNR87, HR89, Har84, HT92, HNT94, HV83, HS89, HP91, Hun79, Igl73b, Jan84, JML92, JS83, KR85, KP83b, Ken72b, KM95, LY88, Law92, LT00, MS90, Mel79a, Mel79b, Miy83, MT97, Nat75, Neu75, Neu77a, Neu78, PL89, Pol75, PBM00, RA98, Ser89, SW91, Sto78, TX10, TE96, TV78, Tör78, TW84a, Twe82, Van78b, VP88, Van88a, Van74, Wan97, XZ96, YZ78, YM98, YZ02, Zha97, BPS19, BEJ16, EW14]. queueing [PY13, PC21]. queueing/inventory [BEJ16]. Queues [AMNW11, Ali90, AT90, Ano80e, Arj72b, AB94, AT81, AIL04, BBH84, BBRW14, Bal79, BW89, BDM85, Bar76a, BDM99, BMV03, BZ03, BC89b, Bri76, BY89, CCP91, CYZ99, Cha92b, Che89b, CR04, Con83, DR92, DTM07, DT80, De 73c, De 73b, De 77, DS73, Deb76, DKS84, Fal89, FD83, GW91, Gre87, Gri92, GRZ04, Had80, Har78, He96, HK89a, Hun83a, Hun83b, Hun84, Hun85, IW70a, IW70b, Igl80, IS83, Jac80a, KW90a, Kel76, Ken72a, Ken72b, Kne00, KW90b, KZ92, KZ94, KKSS08, Lem74, LZ95, LC08, MT16, MB00, Mas84b, MW96b, MH97, Miy04, Neu86a, Neu86b, Num79, Num81, PR06, PT94, Rat75, Rey75, Rit94, Rob76b, Roe73, Rol81, Rol87, RSS99, Ros76a, Rub74, Sch71, Sch74a, Sch78a, Sch87, Ser81, Sha97, Son79a, Son79b]. Queues [Sta90b, Sti72, Szc86a, Szc86b, Szc90, Szc93, Tak75a, Tak75b, Tak76a, Tak76b, Tam74, TW87b, WS87, Web92, Whi70, Whi74, WH80, Whi81a, Whi82a, WD90, AF18, CK15, DL18, EW14, MR20, PZ16]. Queuing [Lan82]. Quickest [GS13]. QuickSelect [FN13]. Quicksort [HRW18]. Quite [Sam04b].

radial [BCT13]. Radio [Rob92]. Radioactive [Lit69b]. Radiotherapy [HRY93]. Radix [EM11]. Radon [Pan06, Tod76, Tod77]. Railways [Gar99]. Random [Adl76, Adl77, AST10, AD93, Ahm80b, Ano77e, Ano85e, Ano92e, Ano06e, AS77, ACLM77, AML77, Ant78, Ant79, ACV80, ANR02, ACS93, AG05, ACDHT05, Asm82, ASC96, Ath85, ARS04, Ave76, AFM91, AS95, BKZ95, BH00, BBHM11, Bac00, BLP06, BBSW05, BBW09, BMY94, BST09, BS12a, BKS12a, BM04, Ban09, BFV10, BE85, BM00, BX06, BNK98, BS07, Bar78c, BK98, Beg12, Ber78a, Ber87, BD96, BVH10, BR04, BE06, Big78, BS05, BHJ11, Bin73b, Bin80, BC81a, BG07, BL08a, BY09, Blu84, BS03c, BB95, BM11, BV06, BK99, BC81b, BV02, Boy78, BHB98, BM02, BMR05, Bro80, BGPS87, Bro95a, Bro95b, BR91, BMS09, CS06, Cam88, CGH^{+}99, Cha09a, CDL12, CS91a, CNT05, Chi12, Chi94, CCH00, Col98, CD14, CC96a, CC96b, Cow96, Cow97]. Random [Cow10, Cre74, Cre79, Cro07, Cro10, CWW09, DH97, Dai08, DO74, DS85, Dav73a, DMR88, DG18, De 10, DZ11b, DM06, DS07, DMM12, DN02, DGN92, Di 01, DMPG09, Din94, Dom07, DM00, DEP91, DH05, DW96, DW02, Eco05, Eli77, FH02, FRW12, FFG07, Fér77, FM95, FH04, FV99, FLS^{+}80, FS09b, Fuh97, FL01, Fuh07, FP12, GLR15, Gal16, GM71, Gar99, Gat00, GJ80, GPD00, Geo84, GP79, Gil98, Gra77, Gre75, GS77b, GT78, GP80b, Gre89b, Gri77, GP02, GG90, GI10, GH75, Had76, Häg00, Hai80, HJ03, HJ04, HSS08, HJ08, Har71, HI79, HM96, HM99, HSS05, Hen87, HM81, HNS13, HS82, Hog86, Hög76, Hol89, Hol90, Hol11, Hor03, Hui03a, Hui03b, HL06, HO98, HR00, Hwa99, Igl75, IM08, IM02, Iva80]. Random [Iva89, IY12, Jan86, JK11, JML92, JK94, Jof75, Kai78, KA81, Kap73, KO87, Kar76, Kar86, KP86, Kat90, KLRM05, Kem87, KK80, Ken85, Ken00, Ken74, KK90, KK92, KMW96, Ker86, Kha86, Kho01, Kid98, KJ03, KM96, KM99, KP94, Kle82, KOS03, KI92, Kri77, Kry76, Kue75, Küf94, LYA07, LZ08, Le89a, Le89b, Lee79, Leh85, LS03, LO04, LS98c, LS93, LS13, LM03, LM04, LR12, Lin74a, Liu98, LC92, LR80, MR02, Mai97, Man76, Man79, Man90a, Man90b, Man94, Mas79, Mas00, Mat99, Mat73, Max77, McD81, MW78, McD78, MY99, MS01a, MNW08, MY80, Mij80, Mil69, Mil71a, Mil80, Mil95, MM96b, Mil98, Mol93, MOK95, Møl89, Møl92, Møl95b, MB12, Muk75, Myl02, NW05, Naw91, Nie99, Nos94, ON96, OC12, OMS06]. Random [Orn77, OB89, Ors90, OD07, PW04, PW06, PW10, Per85a, Per85b, Pet84, Pfe85a, Phi78, PS94a, Pic00, Pie78, Pit77a, Pit96, Pit86, Pre73, PT77, PS10, RS95, RSZM02, Ran80, Rao06, Rav83, Reh80, RFZ14, RN70, Res72, RS71, RW84, RW85, Roy07, Rub74, Rud98, Rüs82, Sch02, Sch05, ST10, Ser80, Ser81, SS90, Sha87, She85, Sie79, Sil76, ST98, Smi74, Sta80, Sto98, Swi75, Szp94, Tak70, Tak84, Tam11, Tan96, TJM98, Tan81, Tau74, Tay79, Thé22, Tho07, Tie81, Tjø76b, Tuc71, TW84b, Val96, VJV10, Van06, VJ05, Ver75, Ver79, Wac09, Wad07, Wan01, WS11, Wat86a, Wat86b, Way96, WW84, WR76a, Wei77, Wes78, Whi70, Whi92b, Wis80, Woe85, Wor95, Wor01, Yu08, Zir01, vdHKV21, AS13, Als18b]. random [AB15, AZ18, AD15, Ath18, AS19, AV19, BCHK18, Bau14, BMZ14, BZ14, BMPR19, BP17, BP22, Bow21, CYG14, Che15, CRZ18, CKM19, CV23, CL14, CL15, CDIM13, DS16, DFRV14, DMPPG16, FPP19, FKV14, Fou18, Gal84, Gan15, GLL20, GC20, GK19, GLS15, GLM21, HHBV20, HTO^{+}15, HLZ18, HL14, HNGS15, Hou18, HS14, HR16, JT22, JKM23, Jan23, JW15, Kin21, KM14b, KK15, KKV16, KN16, KDO16, LZ17, LSV18, LY15a, LL15, LM17, LLX18, LX13, Mey18, MV19, MY20, MS16, MS19, Nes17, OY22, PRR17, PW18, PP15, PSS20, Pin16, RW14, Sak15, SK16, ŠSKB13, Tam13, Tam16a, TW18, US19, Vaš16, Vid20, Yua22, vdHvdHM23, Mat72, Pak89a]. Random-Access [MY80]. Random-Coefficient [PS10, PSS20]. Random-Parameter [Leh85]. Randomised [Bat80]. Randomization [Lim97, YS79]. Randomized [DS15, FOF06, GVW15, SK20]. Randomly [AMS07, AMS09, CC99a, Cow89b, GJL84, HM12, LB10, Lit69b, Bow21, CT18a, CT18b, LP21, ZHCC14]. Randomly-Divided [CC99a]. Randomness [Lie76, Ruk02]. Range [CYZ99, DRV00, LM04, Maj07, MSV01, RPS01, Val96, GNR17, He16, KS16, Vaš16]. Ranging [BB01]. Rank [ASC96, HR05a, KSC09, SJ18]. Rank-dependent [SJ18]. Ranked [SK08]. Ranking [JOC10, KKT06]. Ranks [Cam84, GP79]. Rao [BCO95]. Rapid [BW90b]. Rapidly [MR09]. Rare [AK06, CDL12, God91, Kei74c, LL15, Pak92, RS07, SW93]. Rare-Event [CDL12, LL15]. Rate [AA88a, AP98, AK93, BLL12, Bar85a, Bor79, Bro77, BPS80, Bro05, BP99, CLW95, Coh77c, Cra74a, Dos78, FS82, GG12b, GM98, GI10, Hey77, JY04, JS83, KR71a, LL97, LY12, Maj07, Mer79, Naw91, Ott06, RW03, Rol78, Roo83, Rus10, Sig91, SRTM97, SE91, Tam08, Teu79, TV78, YZ78, AHO20, GLL20, HL14]. Rate-Conservative [Rol78]. Rates [AAR08, BW90b, BGPS87, Cao90, Coh77b, Coh77c, Coh82a, D’S94, DW96, FE06, HJK05, Han78, HS85, Hol81, JP00b, Ken72a, Ken72b, MY01a, Möh00, MT06b, Pol78, SS87, Sha87, Smi82, TT94, WS87, BS15, DSS17, DSS18, GHH15, GB17, SWL14, XHW23]. Ratio [BP00, GL95, KD85, Kes95, KS01b]. Rationing [SSY91]. Ratios [Gla93b]. Ray [BE06, Gre78]. Reaching [Bro99]. Reaction [BLS^{+}80, CR90, PV92, Pia01, JJL21, Pia01]. Reaction-Diffusion [CR90]. Reaction-Dispersal [BLS^{+}80]. Reactions [AT80, Arn80, Kot87]. Real [Ais79, BB07b, FH74, BLT19, FT18, Fra89, Nyr16]. real-time [BLT19]. realizability [GLR15]. Realizable [TM91]. Realization [Cao87, Cao89, Cao90, GS78]. Reallocatable [Miy93]. Reallocation [MSS95]. Recall [Pet84, Pet82]. receivers [HJKP16]. Recently [Con77]. récents [BJ77]. Reciprocal [BNK98]. reciprocity [WR22]. Recognition [RS98, Van94]. Recombination [BS12b, Cha76, DR85a, HH81, KL79, Str75, Str76, AML77]. Reconstruction [RSZ12]. Record [Ano93e, BPS05, BN92, GM96, GH75, Hsi10, Pfe85b, Pfe91, Res74a, Wes78]. Records [BR87, GLS01, GLS05, GLS11, GLS15, Hol11, Kat90, Tho81b, Wes78]. Recovered [Hal87]. Recovery [Fro04, Lit69b, Wic77]. Rectangle [Ros89]. Rectangles [CC99a]. Rectangular [BCM13, Cow97, MM96a]. Recurrence [Ano79e, Bau14, Ber78a, HR78, HS92, Hut80, Kes76, Lev79, Pea89, Pro76b, Sad95, TT79c, Wis72]. Recurrency [FT78]. Recurrent [AN78b, AK98a, AK98b, Gly94, Num76]. Recursion [RS95, SW80]. Recursions [Ano07e, Cli07, GL95, SR00]. Recursive [AH97, BV06, CTS76, DH05, EF80, EF81, HJ03, LM03, RR95, CM17, DJ17, Iye23]. Red [Pon05]. Red-and-Black [Pon05]. Reduced [BZ03, FV99, KM96, KM99, BZ22]. reduced-form [BZ22]. Reduced-Peak [BZ03]. reducible [GJRS17]. Reducing [LF13, Pic00]. Reduction [AHO96, Gla88, Nie98, OS18]. Reductions [Nie99]. Redundancy [BENP92]. Reed [LP90, PL90]. reelles [Fra89]. Reference [Aas82]. ’Refined [Maz89, CMP09]. Reflected [Cha09b, LMX05, Lin05, PMG05, BGL20]. Reflecting [Cre74, Ott76, Per85a, FFI22, KM14b, Lin15]. Reflections [PMG08]. refraction [JMFPY19]. Regeneration [AN78b, Gal11, Num81]. Regenerative [Bal79, Ber78d, Ber78e, Gla93a, HI79, IS83, Kal80, Mai80, MN18, Mih80, Rub78, Sti72, Tho83, Wan97, FZ13]. Regime [AJV13, ES04, FS09a, GM08, LS03, PR00, WWY12, APZ21, CPP21, Cer17, FY18, LP21, TO21, ZY16, Ano04e]. Regime-Switching [WWY12]. regimes [KMF22, OT20]. Region [DRS11, Geo84, Ill98, Ill01]. Regions [NRT83, SL80, She85]. Regression [DR03, Han78, KS74a, Ros75, Ryc87, Wes90, ŠSKB13]. Regret [Agr95]. Regular [Ano06e, ACDHT05, AF18, DMR13, EK12, HO09, HR05a, HL06, Pit77a, Pol91, Res80, Res86, Sta73, Tak84, Teu73, KZ20, MW21, Seg20, Zha15].

Regularity [HBG99]. Regularly [BL08a, Fas05, Fre82, Sen74]. Regulated [AW87a, AW87b]. Regulating [BBZ10]. Regulation [Lom78, Ste80a]. Regulator [Led79]. Reinforced [AMS07, AMS09, AM01, PS04, CLM23, ZHCC14]. Reinforced-Urn [PS04]. Reinforcement [AMS07, AC22]. Rejection [MJM21]. Rejection- [MJM21]. Related [AMS09, ACF09, BMS89, Bar75, Bob04, BMR05, Bro85, BLT03, CM03, CI86, FZ96, Gau02, HNS13, Jan86, KL79, KS83b, LOS03, MZ96, Oka79, Rey75, Sch74b, Sta98, Sto98, VL07, Sak15, Rai07]. Relatedness [Llo73]. Relating [Asm82]. Relation [Ais79, LP93, MY03, Pre77, Ver78, DMPPG16]. Relations [Coh73, DL82, Hok85, HM11, Miy83, Pea89, SY91, Wis72]. Relationship [Ano93g, Gri92, KS01b, NP93, Pfe85b, Sen89a, Sig91, TV10]. Relationships [Don81, Muc05, WC11]. Relative [Cam84, GP79, HMM06, Hil76a, Nak96, Nav03, HM20]. Relatives [Smi80]. Relativistic [Cau76]. Relaxation [Beg12, Sen78]. Relaxed [Whi83]. Release [Bro77, BRT82, HR78, HT92, Las04]. Reliability [Bar78b, CF91, CS97, CTR00, DFZ19, Dro86, ENPS78, Ger84a, GL04b, Gup80, LR97, Lin80, Maz89, NSHH86, Pap87, Pap88, RS71, RS21, Smi83, Szá76, Wen89]. Reliable [Nak96]. Remainder [BW90b, Eng84, Nak79, Rog92]. Remaining [Sch90]. Remark [BV97, Der83, Jof77]. Remarks [Coh74, Get74, Gri80a, Jen84, Pfe91, Sch86a, WB81]. reminiscences [Sen18a]. Removal [Ste84, BN23]. Rendezvous [Lim97]. Rendezvous-Evasion [Lim97]. Renewal [Ahm81, Als88, ABH80, AN78a, Ath79, BS07, Ber77, BW83, BC87, BG07, CN86, Çin69, Çin74, Col74, CGMT20, Fre82, FL01, Gat00, GTY70, Gup80, Hun69, Hun74a, Hun74c, Hun74b, Hun77, Hun83a, Hun83b, Hun84, Hun85, Ito79, Ito80, JOC12, Jen84, KL83, KP84, Kij92, Kom18, Lal84, LL91, Lam93, LG05, LTW10, Lin77, Lin92, LMHN90, McD76, MY01b, MOK95, Neu70, Neu76b, Neu86c, Neu87, OT02, Rog92, Sch71, Sim80, Slu84, SS88b, Szá76, Teu73, VY94, WCL01, Yeo78, BPS19, MZ16, PZ16]. Renewal-Type [VY94]. Renovation [AH97]. Rényi [KI92, Rob18]. Repair [Ave83, FS89b, Fin92, GV76, IL73, IL74, JMJ01, NPOTC00, SZ90, Yeh90, JMJ01]. Repair/Replacement [JMJ01]. Repairable [BFL04, BL04, Ebr93, JD89, LS98b, Las04, LBZ98, NPOTC00]. Repairman [IL73, IL74, Kne91]. repeat [HHBV20]. Repeated [BC87, Big87, Cow96, Cow97, Fal84a, FW81]. Repetition [Blo83]. Replacement [Ave83, Ber78b, BO07, Geo76, Ger94, Ger95, HV83, KS78, RG89b, RG89a, SS95b, SM78, SZ90, Yeh88, Yeh90, JMJ01]. Replenished [Stu88]. Replicated [DMC00, WDH98]. Replication [KRW09]. replicator [BL22]. replicator-mutator [BL22]. Reply [Cha99b]. Representation [AM90, BS03c, CA94, Lac04, Sco73, Szc86a, Szc86b, Ver79, Wei76]. Representations [CLM08, WS10, BD22, NSB13, NW22, Yca84]. Represented [OB89]. Reproduction [KRS07]. Repulsive [HW09b]. Request [Oso10]. Requirements [BDM85]. Resampling [CDL12]. rescaled [AC22]. Research [Bad77a, Git69, Lit74, Mor69, WG91]. Resequencing [Xu92]. Reservation [FOF06]. reservations [MT16]. Reserve [Asm82]. Reserved [FX00]. Reservoir [Has76]. Residence [BS81]. Residual [Arj80, BS99, DZ11b, RL77]. Residuals [Wal73]. Resistance [KR71b]. Resistors [TS98]. Resnick [OS18]. Resolution [ARM99]. Resource [Ala99, BGRSC10, BGZ95, Chi80, CLW95, Git69, Law92, NG77]. Resources [SS92, Whi96, FGRT17]. Respect [AK05, Mas79, NW05, RR90a, Wes76, KXZ23]. Response [KS04b, Wal81a, LF13]. Responses [Ano86e, Fra85]. Restarting [MSS01, Yin00]. Restless [GNMA02, GRHK06, GKRH06, GHK13, NM01, WW91, AM22, HG15, ZCJW19, ZCLJW23]. Restoration [CSY98]. Restoration-Estimation [CSY98]. Restricted [AB07, MT06a, Lem14]. Restrictions [PLRC08]. Result [Ber85, DM03, Jac74b]. résultats [BJ77]. Results [AMS09, ACF09, Ave79, BL07, BD96, BS05, CCP91, CNP91, CC77, Con83, Dag78, De 73c, De 73b, DEP91, Gzy76, Hey79, HS77, Hun74c, Hun74b, Jac80a, JL99, Kel77, KZ94, Lef07, LMP83, Mac75b, MW72, Möh98a, MZ96, MY11, Pak89a, Pra73, RS86, Sch86b, Sch06a, Sch79, SY90, SY97, Ste73, TWN12, Tjø83, Wes74, Zuy99, BJ77, Hou18, MMZ22, PV19]. Resume [Nai86, KS83a]. Resurrection [CZLR04]. retirement [Sen18a]. Retransmissions [JT13, JS15]. Retrial [Fal88, KS93, LK93]. Retrials [AJV13, JvL15]. Retrieving [AB07]. Return [Bro98, GW13, HW09a, PG97]. Returns [KP86, Kem87, KP94, TT04]. Revenue [FX00]. Reversal [HT10, NM79, SW93, TC92]. reverse [BMS18]. Reversed [Hen83]. Reversibility [CM00, Pol88, BD22]. Reversible [HPPT92, Ken78b, Pol86, Ser86, Ser05, TW83, NR21]. Reversion [Rus10, Tav79]. Review [BKM92, Ger84a]. Revisited [BVW85, DP12, FH04, PLPA96, LT15]. Reward [CM98, Gla83, Ker78, Lat77, SB09, SW80, VP88, Van88a, LY19]. Reward-Penalty [Gla83]. Rewards [BC81a, DR16]. Richardson [CR11]. Ricker [FH02]. Ridges [Ros76b]. Riemannian [HZ06]. Right [KS75, Sev71, SW80]. Right-Shift [KS75, Sev71]. Rinott [Yu08]. Risk [Asm82, ASS99, ABS00, BG08, CY05, DE08, DK16, ES94, HLG18, Jac05, LG05, LTW10, PM08, PP07, SH89, VV06, Whi81b, YZ01, ZYY11, AM17, BBRW14, BN20, BEJ16, BFP23, CCHH18, Fro13, Gof19, LY15a, MMP22, SMAL17, WZ19]. risk-aware [MMP22]. Risk-Sensitive [Whi81b, HLG18, CCHH18]. Risks [AG81, BL04, DMR13, McN90, MR09]. Roads [Kon89a]. Robust [PWZ17, Wit79]. Rods [CO98b]. Role [Kei74c, Lan13a, LR18]. Rolling [CFPL12]. Rolling-Type [CFPL12]. Room [Dsh85, Ros73, Tru75a, Tru75b]. Rooms [Son79a, Son79b]. Root [Ott06, JKR17]. Rooted [BLP06, Har71, Kho01]. Roots [Can74b, Can74c, Can75b]. Rose [Spo01]. Rosenblatt [Taq75]. Rosenlund [Ano76e]. Ross [CCP91, MS15, MS16, Pfe85b, Rol81]. Rotated [LB10]. Rotational [Jen95]. Rotations [RW84]. Rough [Nyr98, RT21]. roulette [Cer17]. round [Eli77]. Routeing [BTZ00, LT00, ARW13, VMM19]. Routes [AK08]. Routing [BV93, Hun95, Kel88, Law92, RK89, RH89, TE96, XZ96, ZK89]. Ruin [AS07, AP88, CY05, Cai07, CCH00, ES94, Fro04, Fuh97, GW13, HW09a, Jac05, KR71a, LS21, Mal00, Nyr98, Nyr05, PG97, Pau07, Sti06, TT04, ZYY11, BFP23, EG17, GY16, Nyr16, PR22, WZ20]. Rule [BDM85, BRT82, BVW85, HR78, HK92, SSS04, Tam13, Tam16a]. Rules [AM01, BS03a, BY89, Gla83, HS10, HK93, Wei95b, Wic77]. Run [MPP07, Nyr16]. run-off [Nyr16]. Runs [FJ09, God91].

s [FZ85, De 83a, Hok85, KS04b]. S. [Jof77]. Saddle [LR80]. Saddlepoint [Jen88]. Safety [Mal00]. salesman [Ste15]. Same [DK83, Sam04b, Son79a]. Sample [Agr95, AA88b, AB94, EG81, FVK08, Gri82, Hue94, KS12, RS88, Sig91, ST98, STC94, Wit79, Din19]. Sample-Path [Sig91]. Sampled [MT93a, Sch07, Swi75, Fas14]. samplers [BCJW14, JRR14]. Samples [BN20, DMR88, FM09, God85, Gro12, HS82, Hue99, Kar86, MY99, Wat78]. Sampling [AHO96, Ano03e, ACO96b, ACO96a, ACO98, BCL10, BKS12b, BS12b, BL08a, CDL12, Chi12, Dai08, DG04a, DG04b, DJ91, ENPS78, FFG07, GR08, GKT21, Gri79a, GJS08, GACO00, GACO02, Kiê98, Lee79, LP93, LN92, LY12, Pes75, Pic83, Ste75a, Tho80, Tho82, VV01, WS11, Wat74, Xu11, AD15, BD15, BPS19, BMM13, CRZ18, DJ17, GMS23, LLX18, MJM21, SK20, Tam16b]. Sand [BND81]. Sandwiched [DMYT23]. Sanov [Lac20]. Santaló [Mon95]. Scalar [CR90]. Scale [ÅH99, Mat98, Wor01, AV19, BOS23, DS15, GC16, Sto15, Sto17, Sto23, vdHvdHLS20]. scale-free [GC16, vdHvdHLS20]. Scale-Space [ÅH99]. Scales [Ano04f, EN80, EN88, SW98, YZ02]. Scaling [ACDP12, AS19, Cro10, HJVN03, O’B80, VGS10, AVLR23, FR14, HLZ18, ILP18, XZ20]. Scan [Alm97, KS09, Loa91, Pen02, CK07]. Scanning [AP97]. Scattering [Hor91, KW87]. Scenario [MY11]. Sceneries [FS09b]. Schedules [WD90]. Scheduling [ADZ99, BW93, BM04, CNP91, DW06, GV98, GF87, Gla91, GW92, Käm87, LY88, LBZ98, LT94, PT94, RX91, SSS04, VT94, Wei95b, Xu92, ARW13, DS15, PWZ17]. Scheme [Gal11, Had76, PS94a, Sta98, CM17]. Schemes [Ave76, HH85, KM12, PV92, RS09, Tho80, Yas85, FGRT17]. Schweitzer [Ano79f]. Science [Kit71, Jag18]. Score [AR92, KD92]. Scores [Han07]. Screening [Pro76a, Pro76b, SC77]. SDEs [GLL20, DMYT23]. Sea [Lin06]. Search [CM93, DM03, DH05, Gha08, Gla93c, Ken88b, MSS01, Pfe85a, Sto80, VL10, BP17, BP22]. Searching [FH74, Smi79]. Seas [Åbe07]. Secants [Col98]. Second [Als88, Bal79, Ben95, Bor09, DW06, Ell87, GVM^{+}03, HK92, Köl81, Mor69, ST10, WN94, Yuh82, YS23]. Second-Order [Als88, Ben95, Bor09, DW06, Ell87, GVM^{+}03, Köl81, ST10, WN94, Yuh82]. secondary [EKKZ18]. Secrecy [Sar16]. Secretary [ASC96, KSC09, Lor79]. Section [EC75, GKW96, Kid98]. Sectional [CVS96]. Sections [Kid96, Kid02, Kid08]. Securities [Har80]. Security [TW87a]. Sedimented [Lit69b]. Seed [Mih80]. Seeking [DMR13]. Segment [CSY98]. Segmental [KD92]. segmented [CHJ16]. Segments [FX00, TWN12, Cow13]. Segregating [Gor01, Gri82, Möh06]. Segregating-Site [Gor01]. Seismology [VJ08]. Selecting [Bha86, Cur74, Hsi10]. Selection [ACV80, ABS00, Ave76, Ave79, BBSW10, BK99, BB96, Cam84, CG08, Chi80, CFW87, Edw74, EM11, FFT75, GR08, GP79, GR96, Hil75, Hil76a, LC76, Leó76, Lom78, Pet84, Pet82, QG11, RT07, Rob76a, SS84a, Vet80, Wat79b, Whi70, AS14, Cer17, CIL17, CV23, GY16, NRB22]. Selective [HKH76, Leo09, BSS17]. Selectively [Wat74]. Self [All82, BNPS01, BS07, BBZ10, BE06, BHJ11, CS97, Cro07, FP12, GVM^{+}03, Hue99, Lom78, May78, VJ05, CGMT20, DP22, Duc21]. Self-Decomposability [BNPS01]. self-excitation [CGMT20]. Self-Exciting [CS97, DP22]. Self-Incompatibility [May78]. Self-Normalized [FP12]. Self-Organization [All82]. Self-Regulating [BBZ10]. Self-Regulation [Lom78]. Self-Similar [BS07, BHJ11, Cro07, Hue99, VJ05, Duc21]. Self-Similarity [BE06, GVM^{+}03]. Selfing [Möh98b]. Semi [Arj72a, Arj72b, AN78a, BCT13, BMY91, BD97, Bal97, BS81, BR86, BG93, Cao95, Çin75, FT78, Ger94, Hun84, LS13, MS87, Mat72, McD85, MSS95, Neu77c, Num76, Ott74, Ott75, Pic98, PZ86, Sch76a, Sch78b, SP73, Tak75c, Tho81b, VV06, Bré75, But22, Fra89, HLG18, Mon17, Fra89]. Semi-direct [Fra89]. semi-directs [Fra89]. Semi-Hybrid [Pic98]. Semi-infinite [BCT13]. Semi-Markov [Arj72a, Arj72b, AN78a, BMY91, BD97, Bal97, BS81, BR86, BG93, Cao95, Çin75, FT78, Ger94, Hun84, LS13, MS87, Mat72, McD85, MSS95, Num76, Ott74, Ott75, PZ86, Sch76a, Sch78b, SP73, Tak75c, Tho81b, VV06, But22, HLG18, Mon17]. Semi-Markovian [Neu77c]. Semi-Markoviens [Mat72]. semi-martingales [Bré75]. semidefinite [FJKR20]. Semigroup [Cap76]. Semigroups [Hög76, Muk75]. Semilinear [CLM08]. Semimartingales [JM86, PT23]. Semiperiodic [AH05]. semistationary [PV19]. Senescence [Leó76]. Seneta [HLZ18, Sch82a]. Sense [Tjø76b]. Sensitive [Whi81b, CCHH18, HLG18]. Sensitivities [Cao89, RS21]. Sensitivity [AR92, BB93a, BT08, Cao90, Cao95, KZ92, KZ94, Fly18]. Sentry [BBSW10]. Separable [Pol79, SE91]. Sequence [Cam84, CG08, Che89a, Fre82, GH75, JL77, JP00b, Kat90, Man90a, Man90b, Nav03, RN70, Ver73, Wat86b, Wei77]. Sequences [AH97, Ano06f, Bal77, BCL10, BC87, Big87, CC99a, CHM91, CC99b, GL80, Gla93a, IW70b, JM86, Jor10, KA81, KO87, Ken72b, Lef07, McK88, McL75, NW03, O’B80, Pel83, Rød06, Roo83, WCK96, ZG04, FZ13]. Sequential [Bat80, BL09b, CDL12, Chi80, HH85, LY94, PS09, PS11, Sta98, Tho80, Van06, Whi69, ALS18a, BCJW14, FDJ20, Mey18, MZ20]. Serial [Whi80a]. Series [Abb75, AT90, ACDP12, Ano84e, Ano86e, Ano04g, Bal79, BENP92, Cao98, CF91, CA94, Cli11, Cou90, Dav73b, Dav79, DDH78, DLO06, DH76, DR81, Fin69, Fra85, GPD00, Ger80, HJ04, Han78, HK84, HH03, Hol89, Hua97, HR16, JL79, Kel82, Koh80a, Koh80b, Lac04, LL81, Liu90, Lom72, Neu70, Par77, Pic85a, PC96, Sch06a, Sch06b, SS92, Tjø78, Tjø83, Tjø90, Ven82, Wal73, Whi80a, Wit79, BSS23]. Series-Parallel [HJ04, Lom72]. Served [Ros73]. Server [ADZ04, ARW11, BBH84, Box83, BC89b, DR92, De 73a, De 73c, De 73b, De 74, De 76, De 77, De 80, DKS84, FS89b, FS91, Fer93, GY78, GW91, GV76, Jac80a, KW90a, KD94, Lem74, LK93, LMHN90, MH97, Miy04, Ott87b, Sch74a, Sch87, SK73, Son79a, Son79b, Van88b, Whi70, WH80, Bar18, BD15, DVN15, Fri86, Fri87, JvL15, MT16, PY13, Lat77]. Servers [ADZ99, CKW09, HP91, Neu70, RK89, SK08, Son79a, Son79b, WW21]. serveur [Fri86, Fri87]. Service [AT81, BDM85, BV90, Box78, Box79a, Box79b, BK01, BC89b, BGPS87, Bro05, Cao90, CS91b, Coh78, Cou90, DTV84, De 73a, DS73, Deb76, Dos78, Gal79b, GY78, GG12b, Had76, Jac80a, JS83, KF90, KD94, LZ95, MAF07, MH97, MY11, Neu76a, Neu77c, PBM00, RA98, RW91, Rit94, Rub74, Sti78, TV78, WS87, Web92, YZ78, CAR20, Fri86, Fri87, Gol16, SS20, Sto15, Sto17, VBvdM15]. Service-Level [DTV84]. Service-Rate [JS83]. Service-Time [Neu76a]. ses [Liu97]. Set [ANR02, AFM91, BC01, BM11, BHB98, CB95, CB03, Cre79, Fer93, Gil02, HS94, Her73, Her75, Led79, Leo09, Mil75, RSZ12, SS90, Sha09, NW22]. Set-Chains [HS94]. Set-up [Fer93]. Sets [AST10, ARS04, AS95, BKS12a, BS07, BS03c, Cao99, CFPL12, DMR88, DW96, EPS79, FM95, Kal80, KK80, Ken85, Ken00, Ken78a, Kid98, KJ03, KOS03, Mai80, Mas79, Mat72, Mol93, MOK95, Mrk01, Pic00, Sch02, WW84, Wor94, Wor95, Zuy99, Fou18, GLR15, KR18, KN16, Pri15, ŠSKB13]. setting [Sva14]. Settings [FS08]. Settling [ST90]. Seventieth [DKMJ10]. Several [Bal85, BS12a, BGPS87, CK85, Neu77a, Neu77b, Whi80d]. Severity [BC93, BB07a, PL90]. Sex [Ben80, Dal85, GT84, HM81, Möh98a]. Sex-Determination [Ben80]. Sexes [Str75]. Shadow [Kel88]. Shannon [AFN74]. Shape [BCLF16, Big78, CC99b, Dei03, DM91, Gat94, GM91, HJ00, HZ06, Ken77a, KL87, Ken95, Ken96, Ken98, KL00, KL03, Le89b, Le96, LK00, Le01, LB10, Lin06, MD89, MD94, MW94, Pan06, Pim07, Wei95a, Wei97, BN20, BDK^{+}18, Bez20]. Shape-Densities [KL87]. Shaped [BC01]. Shapes [CC99a, Cow96, Cow97, Har71, Ken85, Ken88a, Le95, Le96, Le98, Man88, Man90a, Man90b, Mil95, RT02, Wat86b]. Shared [Kei76, Sch86b, Kin21, NSB13]. Sharing [BMV03, BB10, Coh78, Gri92, Gri94, GRZ04, HP82, LRS10, Maj07, MM83, MM85, RS21, Sch84, Smi83, Tie82, TW83, DS23]. Sharp [AAL08, KS74b, CKM19]. Sharply [Bar81]. Sharpness [FRW12]. Shattering [GKP84]. Shearer [HT17]. Shedler [Ber85]. Sheep [Cha96]. Sheets [Pyk75]. Sheffield [Gan71]. Shell [MP03]. Shift [AK05, KS75, Sev71, BR15]. Shiu [LG05, PP07]. Shock [Ano93f, Ano94e, Dro86, Gut90, Koc90, MSM92, Pel93, POGP95, PZ86, SS84b, SJ89, SS85]. Shocks [RG89a]. Shops [PR82]. Short [AK08, LS89, RR21]. Short-Length [AK08]. short-memory [RR21]. Shortest [GGV89, HR89, Sch90, VGS10, HNGS15]. shortest-path [HNGS15]. Shot [BY09, BM02, Bri99, GT75, HS85, Hsi87, HT89, IJ02, Kon89b, Lan87, LP21, Møl03, MT05, RPS01, Ric77, Wei76]. Shot-Noise [Bri99, Lan87]. Showers [Gre78]. Shuffled [Ano06f, Rød06]. SI [KWRK22]. Sibships [Tho80]. Sided [Jac05, Rog90, Sta92, Yas85, CKL14, CV23, LY19]. Sides [Cal03b]. sight [LBCEN21]. Signal [Hal87, HKT91]. Signalized [Gri80b]. Signalling [HNT94]. Signals [Hor02, Wor01]. signatures [YBC20]. Signed [Adl89]. Similar [BS07, BHJ11, BH89b, Cro07, Hue99, VJ05, Duc21]. Similarity [BE06, GVM^{+}03]. Simple [ACDP12, BP79, Ber85, BJML07, Cro10, DK12, DLO06, Dro89, ENPS78, GP88, HMV06, HH81, Hey74b, Imh92, KP86, Kem87, KP94, Kog89, KSC09, LR97, Lin92, Mit87, Mol72, Mor94, Pak79, Pak89b, Pak98, Par87, PV92, RG84, SC83, Sen74, Shi80, Ste75b, Val96, Van90, Ven82, WD71, Yun86, Cer17, US19]. Simplex [DS03, KL03]. Simplices [Hin94a, Mil71a, KZ20]. Simplicial [DFRV14]. Simply [Eli77, Ill98]. Simulated [HS91, MSS01, MRSV86, SK96]. Simulating [HIS76, LN92, Tan96, Wat78]. Simulation [AK06, BvdMS20, Bon82, BK02, CR08, CDL12, CR90, Dai11, DN02, DM05, DHK^{+}84, Elv71, Gal11, Geo76, GJRS17, HI79, Igl80, IS83, KM00, Lac04, Lan96, MM98, Mih80, MR05, PH99, Thö99, AD21, BZ20, CMB19, CK15, GPHK21, LL15, MS16, MJM21, OS18, QDZ19, SMAL17]. Simulations [Nak96, TKF15]. Simultaneous [Gri75b, KM10a, WR76b, BBRW14, Cra97]. Single [BBH84, BD97, Bal97, BGZ95, BC89b, DTV84, DKS84, FOF06, FS91, Fie75, GF87, GW92, Jac80a, Kon89a, Lem74, LK93, LMHN90, MH97, Miy04, Ott87b, Sch74a, Tho82, Van88b, AF18, SWL14]. single- [AF18]. single-chain [SWL14]. Single-Lane [Kon89a]. Single-Machine [GW92]. Single-Product [DTV84]. Single-Server [BBH84, BC89b, DKS84, FS91, Jac80a, LK93, LMHN90, MH97, Miy04, Ott87b, Van88b]. Singpurwalla [BB90]. Singular [Din19, GNRS89, Kar83, DF18, KMF22]. singular/switching [KMF22]. Singularity [BNBJ00, BS07]. Singularly [AL99, YZB03]. SINR [BBHM11, JT22]. SIR [BL01, BS12a, BN23, GC16, GL09, LP99, LG16]. SIS [Ano12, Nås96, Nea06, Nea14]. Site [Gor01]. Sites [Gri82, Möh06, Wat82]. Situation [SR76]. Size [And99, AT81, Ave76, Ave79, Bal86, BC93, BRS99, BF05, Cao99, CGH^{+}99, DS03, Dan74b, DMPG09, Fil74, HRTY97, Hey78, Hil81, Hui03b, KS06, Kle84, LJ04, LJJ04, Le95, LP90, Leo09, Lip75, MW94, Mas95, ON96, ON98, PL90, Pit96, SC83, ST85, ST98, Ste80a, Szc90, Wau74, YC96, Yin00, BN23, BSS17, CYG14, XZ20]. Size-and-Shape [MW94]. Size-and-Shapes [Le95]. Size-Biased [DS03, Pit96]. Size-Dependent [LJ04, LJJ04, Wau74]. Sizes [BK99, Möh06, Sar98, ENR17, KL13, LSH16]. Skeletons [SS15, TT79a]. Skewness [Lep76a]. Skip [GHT97, SW80, AV19]. Skip-Free [SW80, AV19]. Skorokhod [ACH21, Sco73]. Slabs [Rat04]. Slacks [Tie81]. Slepian [Lin06]. Slice [JK94]. Sliced [ŠSKB13]. Slope [Lin10]. Slopes [Åbe07]. Slow [AJV13, FS09a, US19]. Small [But94a, CT08, Dia13, Fuh07, GX07, Gha08, Hil76a, KD85, KKSS90, MS17, Wac09, ABL15, BCHK18, BR13, BFW22, Chi14, CF14]. Small-Noise [KD85]. Small-time [MS17]. Small-World [CT08, GX07, Gha08, ABL15, BR13]. Smallest [BK99, Cal02, Lee99a]. Smallest-Fit [BK99]. Smith [GZ93]. Smoluchowski [Nar91a, Nar91b]. Smooth [BFV10, Sch03b, EY14, FKV14, HP23]. Smooth-Grained [Sch03b]. Smoothed [MH97]. Smoothing [ÅH99, Cal03a, Liu98, RS97, NLM18]. Smoothness [Leh02, PLRC08]. Software [CS97]. Sojourn [BB10, Che89b, Cse92, Hoo78, Jan84, Kei74b, Kei74c, KP83b, Rit94, Tav76, WV80]. Sojourns [Ber82]. Solution [Boy78, Bra78, FKM82, Hok79, Hok80, Par87, Sac90, SY81, Spo84, Sys88, Wah76, Mas23, NP17]. Solutions [Beh11, DSSW11, DN02, FR11, GR08, GT08, Kne00, Neu75, Sam04b, YS94, HRW18, LY19, Sto23]. Solvable [Gat85b, EC21]. Solving [BU98, Ben74, Ben75a, BC04, CR90, FV95]. Some [Agr74a, Ano93e, Ano93f, AK98a, AK98b, AS96, BCO95, Bar75, BE83, BH89a, BNJS98, BL07, Ben75a, BD96, BS03b, BY09, BH89b, Bra78, BP82, Bro89, BMS09, BN92, Can74c, Cha96, CTR00, Coh74, Dav70, De 73b, DGN92, DMY93, EGV79, Fin92, GS78, Gan75, Gan78, Gan79, GS13, Get74, Ghu87, GNR86, GNR87, Gla83, GRHK06, GL04b, Gri80a, Gup80, Gut74, Gzy76, Hey78, Hol89, HS77, Hüb78, HT10, JK11, Jen84, Koc90, Kot80, KS80, KS75, LS98b, LL97, LP93, Lef07, Li09, Lin74b, LT73, Mac75b, MNR02, Mar75, MMZ22, McK88, Mer06, MS76, MH97, Nec78, Neu77c, OM95, Pak71b, PT85a, Pak86, PH99, Pfe91, Pol74, PV92, RSS94, Rao06, RG84, Rob07, Ros80, SK96, ST85, Sch79, SY91, Sha78b, Shi80]. Some [SY90, SJ89, Smi74, SZ90, SB09, Ste73, Szp94, TJM98, Thö99, Tjø83, VML80, Wan01, Wat79a, Wau74, Wil79, Yor92, BPS19, CYG14, Clo17, KTR19, LPST14, Sen18a, Yca84, dlI22, Par98]. Sort’ [Cra97]. Sorting [BND81]. Sources [DS00, ES04, KS93, LM00]. sous [FG81]. sous-variétés [FG81]. Space [ÅH99, Ahm80b, AL94, Ano04f, ADW^{+}75, ARS04, BBHM11, BC89a, Cow89a, Cow89b, Fed78, Gzy76, HS91, Han96, Ken95, KOS03, Kue75, KL00, LS11, MY01a, MS87, Mat99, Mec89, MS90, MP03, MS03b, OD07, Pit77a, Sav75a, Sav75b, SCA00, Sto98, Ver75, YZ02, BCJ^{+}17, BDK^{+}18, Bez20, CM17, GLL22, MT13, NLM18, Sim15]. Space-Time [BBHM11, Sav75a, Sav75b]. Spaced [Mas83]. Spaces [AN78a, AN78b, BR76, GHL07, HZ06, Iso00, JY04, KM00, KL03, Le01, Mec77, Mil69, Mil71b, Pol79, Ser05, SK08, DG18, GJRS17]. Spacings [Roo93]. Spanning [BLP06, BR04, HNS13, Lee99b, Pen98, PW04, PW06, BCT13]. Spans [WR76a]. Spares [IL73, IL74]. Sparse [Kho01, MW21, PS94b, GLM17, OT20]. Spatial [ABA^{+}80, Ano84e, Bad80, BT98, Bar74b, Bar74c, BR93, BBD99, Bor06, BO07, BV90, CDDP11, CORR91, Dan77, DMC00, Ell86, HJVN03, HBG99, Iva89, Kiê98, Kri79, MPB96, MDB^{+}75, MT11, MS10, MB12, MKS^{+}78, PW10, Qi08, RC94, Ser05, TWN12, Tjø78, Tjø83, WC11, WDH98, GC16, Lan13b, Man21, vdHvdHM23]. Spatially [Mat77, LR18, CT18a]. Spatio [Ma05]. Spatio-Temporal [Ma05]. Special [MY89, PT23]. Species [BCL10, Bul74, ENPS78, Hey78, IMS21]. Specific [Had76, MS01a]. Specifically [RT07]. Specified [BJML07, Hen87].

Spectra [BHJ11, BM02, BMR05, LW92]. Spectral [CK73, Cra74b, GHT96, GLM22, HL14, Kho01, KOS03, Koh80a, Koh80b, LSV18, LL98, Ric75, Tjø76a, dlI22, GLM17, LS15]. Spectrally [Eme73, KR02, PvS21, Rog90, WS11, YZ01, BP23, BOS23, CKLP20, EY14, GR16, WZ20]. Spectrum [Roy07]. Speed [BMM16, Bar78c, FRLM20, Li09, SU79, BDK^{+}18, Bez20, Pin16]. Speeding [MP18]. Speeds [BK01, Sch78b, Tay89, Duc21]. Spent [Imh86, Ken74]. Sphere [Bro95a, Bro95b, CGH^{+}99, GACO00, GACO02, Ols07, ZCWS96, HT17, MJM21]. Spheres [Adl79, Gha08, Ken77c, HLPW14]. Spherical [DSS99, LS98a, Le89a, Le89b, Rat98, Tan96, KZ20, TKF15, Let82]. Spherically [EG81, Stu88]. sphériques [Let82]. Spike [BMR05]. Spin [Geo84]. Spinning [GKRH06]. Split [Hol11]. Splitting [Bar78a, GP80a, KI92]. Sponge [Gri81]. spontaneous [Fox14]. spot [BV13a]. Spots [Sau78]. Spread [BLS^{+}80, WB71, Gan15, GK19]. spreading [LSV18]. Spreads [VV06]. SQ [VMM19]. Squad [GKRH06]. ’Square [Ott06, Cox80, Ros89, SR76, SW77, SW78, Wie77, GLL20]. squared [AC22]. Squares [Fin69, LW92, Moo79, Sur08, Wal74]. St. [GML18]. Stability [BL89, BW89, BW93, BNPS01, BK98, BS06, Bor06, BGM13, Cam88, Cli07, Dag78, DW06, DM00, DS76b, FH02, GM11, Hin94b, Hor03, JS14, KR85, KVB99, LK93, Luk77, May78, MT92, MT93a, MT93b, Nav03, Ott74, Rej82, Res72, Rob07, SS20, Szp94, YA00, Ano07e]. Stable [AST10, AS07, Alp09, Ben80, Bin73b, BP99, CG80, CM93, Der09, Fei76a, Fri77, Hol11, KM00, NW05, Ott77b, Rie80, Sam04a, ST10, Ste80b, ST04, Sur08, WS10, WS11, BCM18, BCJ^{+}17, CMB19, EMOS14, HLZ18, HR16, JSS19, Thé22]. stable-type [Thé22]. Stably [CK99]. Stack [Han07]. Staffing [JvL15]. Stage [ADZ99, Hüb80]. Stages [Bar76a, Baz74, Cou90, LG16]. standalone [Fri86, Fri87]. Standard [AD94, ASS99, LP90, MT97, Wat79a]. Standardized [Pic00]. Standing [Mil95]. Standpoint [Amb77a, Amb77b]. Star [BC01, Hun95, KF90, Sat84]. Star-like [KF90]. Star-Ordering [Sat84]. Star-Shaped [BC01]. Starshapeness [SW91]. Starting [AW87a, FM18]. Starvation [McN90]. State [AK79, Ano79e, Ano04f, AN78a, AN78b, BW89, BC89a, BL08a, BV91, BB10, But94b, Cao90, CR04, Coh77c, CM98, Con77, Con83, FT78, Fed78, GM08, Ger84b, GT08, Gzy76, HS91, HNT94, Hok79, Hok80, JH92a, JY04, Led04, Lev79, MY01a, MS87, Mel79b, Mer79, MVY84, MY89, Nat89, Neu77a, Neu77c, Pak71a, PT85a, Pak95, PMG08, Rob07, Sch71, Sch78b, Sch90, Sen89b, Sen97, Sim15, SK73, Teo82, TVF81, Van90, Wil98, YC96, YZ02, APZ21, BPP21, BCJ^{+}17, BS15, Bré75, CM19b, CM17, FM18, FMY22, GLL22, GJRS17, JS14, Kur18, NLM18, YS23, YBC20, CT18b]. State-Dependent [Ano79e, BW89, BL08a, BV91, BB10, Cao90, CR04, Con77, Con83, HNT94, Lev79, Mer79, MVY84, MY89, PMG08, Rob07, Sch71, CM19b, JS14, CT18b]. state-space [BCJ^{+}17, CM17, NLM18]. States [Ano83e, But94a, Coh77b, Daw75, Gat97, Kot87, MZ04, Pre73, Pre77, Rub81, RS21, Tho78, MR20].

Station [FE98, XZ96]. Stationarity [BDLP14, DLS94, Hof07b, LW11, Szc93, Van91, Oça23]. Stationary [AK05, AS07, Alp09, AH97, Ano80e, Ano93g, Ano04g, Ant78, Bal07, Bar76b, BNJS98, Bat76b, BL09a, BK80, Ber78a, Bob04, Bra86, Bro77, BB96, Cav78, CA94, CHM91, CS99, Coh74, CK85, CY99, Dav73b, Dav73a, DM06, DKS84, DR81, Eco05, FMP92, Gat97, Ger80, Hal86, HR78, HV84, HH03, Hei98, HSS05, Hey74a, Hüb80, HT10, Ill01, Kar90, Kar85b, Kar86, Kem73, Ken00, Ker78, Kid96, Kid98, Kid01, Koh80a, Koh80b, Kön77, KZ92, KZ94, LS98a, Las06, Lee79, LZ05a, LZ05b, Lie76, MS03a, MPS98, Mat98, Mec96, Mec98a, Miy83, MZ04, Miy04, Mor73, Nås96, NW03, NW05, NP93, Nav03, Nor75, Ott87a, Pak95, PSS06, Pic74a, Pic74b, PMG05, PMG08, PS94b, RL10, Rao06, Ric75, Rol78, Rol81, RSS99, Roo83, Roo88, RS07]. Stationary [Sam04a, Sch78a, Sch96, Sch00, Sco73, Sel96, Sen80, Sen88, Ser86, Sha78a, Sha78b, Spo01, Szc86a, Szc86b, Szc99, Tjø76b, Tjø76a, Twe82, Van91, VGS10, Wal73, Wee02, WW84, Whi80a, Wis72, ZG04, AD21, AB15, BSS23, BGZ16, BFW22, FG81, GLS15, GJRS17, HS13, KT20, KM14b, LZ17]. stationnaires [FG81]. stations [MS14]. Statistic [Fei78, Pen02]. Statistical [Ano76k, Ano83e, Ano84e, Ano13b, Ano14b, Ano15b, Bal78, Bar74b, Bar74c, Bro73, CFPL12, Edm76, EMOS14, FM95, FS79, HCD94, IS83, Kas80, Kei78, Lep76b, MPB96, Now88, Par77, Pic85b, Pru78, Rub81, Str78, Tjø78, Tjø83, FG81]. Statistics [Alm97, Bad80, BE83, Bat78, BR91, CK07, Dwa74, GM96, Hua02, HLPW14, KS09, Kho01, Kri77, Kri78, LO04, Loa91, MPP07, Mej77, Mol95a, Pic83, Rei81, RW77, Jan23, XY15]. Statistique [FG81]. Stay [JLW05a, JLW05b]. Steady [GM08, Hok79, Hok80, Nat89, Neu77a, Neu77c, Sch78b, Sch90, Sen89b, SK73, TVF81, Wil98, YC96, APZ21]. Steady-State [GM08, Hok79, Hok80, Nat89, Neu77a, Neu77c, Sch78b, Sch90, TVF81, YC96, APZ21]. steepest [HHHT19]. Stein [BH89a, BU98, DLU12, Rai07, Ros13]. Steiner [Mat74, Vit95]. Steinhaus [HS82]. Step [Yin00, GR16]. Stepping [Mar70, Ren86, Shi79]. Stepping-Stone [Ren86, Shi79]. Steps [BR88]. Stereological [HS98, JK94, Mas95, ON96, ON98]. Stereology [ACO96b, BCO95, Ben95, Ben98, HJ02, Mon95, OM95, WN94, TK13]. sticky [MLZ21]. Stiefel [Mar77]. Stieltjes [BR83]. Stienen [Mån07, Ols07]. Stigum [Liu97]. Stimuli [BP94]. STIT [Cow13, LR11, MN14, NW05, TWN12]. Stochastic [Aas82, ÅH99, APP99, AEM94, AT75, Ahm80b, ADZ99, AH97, Amb77a, Amb77b, AT90, ARM99, Ano98e, Ano07e, Ano12, Ano13b, Ano14b, Ano15b, ADW^{+}75, AT80, Arr03, BMS89, BT71, Bal78, Bal85, BC93, BO93, BO94, BL01, BND81, Bar77, Bar79, Bar84, BP83, Bas79, BS99, Ben75a, BDF^{+}84, Ber80, BV93, Beu83, BENP92, BT08, Bor98, Bra86, Bro73, BK05, Cap76, CF78, CT85, CCP91, Cha92a, CR90, CK99, Cli07, CFW87, CEP10, Coh88, CMMS13, Cor14, Cra97, DLU12, DSSW11, DF18, De 10, DK12, DLS04, DGB^{+}84, DK80, DG01, Dyn78, Elv71, EX11, Fad77, Far76, Fed78, Fei76b, Fie75, Fin69, FP96, FZ13, FS08, GS78, Gan75, GRW10, Gla93b, GF87, Gla91, GW92, GL95, God85, GT08, GKT21, GKP84, Gri80b, Gri75b, Gue93].

Stochastic [GW13, Had79, HRTY97, Har80, HI79, HW95, Hey85, HJ00, Hor03, Hua97, Igl80, Imh08, IL97, JML92, JS80, KX94, KP10, Käm87, Kar83, Kar85a, Kar90, Kas78, KLRM05, Ken88b, Kes72a, Kes72b, Kha80, Kin79, KM10b, Kli80, KLS02, Kön77, KW88, KKSS08, Kot80, KS12, Lan13b, Le 74, LM90, LP93, Leh85, LMX05, LS93, LBZ98, LY77, Lin06, Lin10, LT94, LMS00, Mac75a, MNR02, MAF07, MW93, MS80b, MW96a, Met80, Mih80, Mit88, MS01b, MS04, Mor76, MS11b, Mül97b, Nar91a, Nar91b, NG77, Nea06, Nec78, Nij74, Nor82, Off74, ØSZ11, OB89, Pai01, PTW10, PT94, PG97, Pic74a, Pic74b, PR82, PR94, Pri77, RS95, Rao83, Rie80, RS92, Rip77, Rob07, Rob74b, Roo78, Ros74b, Roz78, Rud72, Rüs91, SK96, Sch86a, Sch06a]. Stochastic [Sch06b, Sch77, Ser75, Ser86, Sha03, SS87, SS88a, Sha87, SY91, Shi79, SR00, Sin98, Sob87, Sti78, ST04, SW92, Sut79, TT04, Tau75, Tru97, TW84a, TW87b, Van80, Van94, VT94, VL07, Ver77, Ver79, Wal81a, WCK96, Wee02, Wei95b, Whi91, Wil71, Woe07, Wol80a, Wol80b, Xu92, Yin00, Zak77, Zaz92, Aaz13, Ada16, BGV22, BOP21, BV13a, BV13b, BGZ16, BGL20, BFW22, CGR18, CLM23, DJ17, FPP19, FR14, Fly18, GS14, GLM17, HRW18, JJL21, Kar21, Lef88, LTT20, Lin15, Luo16, NP17, NRB22, ORT16, QD15, Qia18, RM21, SWL14, TL20, US19, WHL17, AS77, Fad79]. Stochastic/Deterministic [Fad79]. Stochastically [Gås12, LS93, SS90, BSS17, PR22]. stochastique [Lef88]. Stocks [KK03]. Stokes [MT13]. Stone [Mar70, Ren86, Shi79]. Stop [RR90a]. Stop-Loss [RR90a]. Stopped [Als88, CDH87, GT78, KK90, KK92, LC08, WZ19]. Stopping [AP98, AM01, And75, AS73, AGSC04, Bat70, Bat76a, BS06, Ben91, BG93, CL87, DE12, DW02, FR11, GKL21, Gau02, HS10, KK90, KK92, Kur78, LYA07, PvS21, Sta92, Sza93, Tam11, Van73b, ZCWS96, Zuy99, Att17, BCSY15, DF18, EC21, EY14, GP15, HP23, Li23, LY19, PS83, Pri15, Tam13, Tam16a, Tam16b]. Stopping-Allocation [Ben91]. Stops [LOS03]. Storage [AP88, BRT82, DH82, Gan69, GS77b, GS77a, HR78, Hun79, MS90, PT85b, SP73, SY81, Tak75c, Tol88, Woo80, BF19]. Store [VK78]. Store-and-Forward [VK78]. Stoyan [DKMJ10]. Strain [GL04a]. Strassen [Ver75]. Strategically [Hor03]. Strategies [ABS00, BJML07, Bro98, FW81, LY94, SZ90, AAM17, HA20, JMFPY19]. Strategy [Ave83, CG80, CM93, Dro89, Gar80, Gla91, SW92, WWY12]. Strategy-Frequencies [Gar80]. Stratified [Whi80c]. Stratonovitch [CDL99]. Stream [Rol81, Wei76]. Streams [BM77, Bro72, Maj07, Mel79a, Ott87b, Sch74a]. Strength [Dan89, HP82, PT73]. Stress [GG90, Las04]. Stretch [CD18]. Strict [Edl78, Tjø76b]. String [Kin77, Wil85, DM90]. strings [Jan23]. Strong [BO94, BZ14, GG12b, GLS01, Kla85, Moo79, Ran77, Zha97, Dio76, Dio78]. Strongly [VW77, VV78, Cap86]. Structural [CNP91, ENPS78, Gha08, Tau71, FPS22]. Structure [Ahm80a, BST09, BS12a, Bar77, Bar79, BGS89, Büh74, BB96, CK73, CI00, CF91, CPR01, Çin75, Coh77b, Coh77c, Ebr90, ENM89, GJS08, Gue93, Hil76a, HA79, Igl80, JK85, Kei76, KL87, Mas84b, MSS95, PL99, PS94b, Rej82, RR73a, RR73b, Rey75, Sch79, Sen79a, Sen79b, Tho79, Tod76, Tod77, TT79c, VV06, WS10, Whi82b, WH03, WA82, YS94, AB22, BC16, FZ13, Mon17]. Structured [CW75, Hös11, Saw77, ZLB98, FRLM20]. Structures [CPZC05, HSS05, Kiê98, VL07]. Stuart [Ano78e]. Student [HL05]. Studied [Neu70]. Studies [MPB96, RPS01]. Study [Cap76, KR71b, MM98, Mes71, Mül97b, NSHH86, Pri77, Ros76b, Sch06a, Sch06b, Sch07, Ven82, Fri86]. Sub [DS00, Kes95, OY22]. Sub-Exponential [DS00]. Sub-tree [OY22]. Subadditive [Hey79, Smy76]. Subcritical [FV99, PT85b, Ven82, CPP21]. Subdivided [BBZ10, DG04b, GJS08, WH03]. Subexponential [ASS99, CNT05, EGV79, GR88, JL99, KP22, LTW10, Miy04, OMS06, Zwa00, FT18]. Subexponentiality [Van84, XCWC18]. Subgeometric [DSS17, DSS18, TT94]. subgraph [CGR18]. Subgraphs [GS19]. Subject [AH87, Ano89e, BL04, GV76, Hai80, LM03, LM04, Neu71, Pau07, RG89a, SM78, GLS15, Lan13a]. Subjected [LM90, Tay79, YZ78]. submanifolds [FG81]. Submartingales [Sud99]. Suboptimal [KSC09]. Subordinated [Sta73]. Subordinates [Rub78]. Subordination [BNPS01, DSS17, DSS18]. Subordinators [Cov09, Fri77, Mai80, SMAL17]. subpopulations [BB17]. Subsamples [STW84]. subsequence [AS14]. Subsequences [HMM06]. Subset [MS87]. Subsets [FM95, Gzy76, Tho07]. Subspace [SK04]. Substitution [JP00b]. Substitutions [Ant78, Wat82]. Substochastic [GT08]. Success [MPP07, Ver73]. Successes [BL88, CAK00, Tam11]. Succession [Rej82]. Successional [Lan05]. Successive [BSHJ17, TV75, Van78d, Wes76]. Sufficiency [DS76b, Nij74]. Sufficient [Bat78, BH04b, DW06, Fal84a, Mer79, WCK96, Che15, XCWC18]. Suggestions [Ano83f, Nat82]. Sum [BPS05, BK99, CFW87, GM98, GHL07, LR80, Pak00, PP02, Yas85, Att17, DF18, FPP19, HMM19, Mon17]. summands [Sen18b]. Sums [Als88, BE85, BX06, Bin73b, BG07, BR91, CW03, Cha09a, CNT05, CDH87, HM12, JML92, KD92, KS12, McD78, Mij80, OC12, OMS06, Phi78, RR90a, Rüs82, Sho79, Tak70, TV75, Tuc71, Yu08, BM16]. Super [BD75, BP99, Wil74]. Super-Critical [BD75, Wil74]. Super-Stable [BP99]. Superadditive [AABK86, BGS89, ENM89]. Supercritical [BB93b, BD74b, BD74a, Edl78, GB03, Her75, LJ04, Ric11, Sch82a, BPP21]. Superinfection [Pic98]. Superposition [BMTY97, Ber77, LL91, Lam93, MB12, Ste02, Yeo78]. SupOU [MS11b]. supplies [Bar18]. Support [Wei98, CM19a, HHHT19]. Supported [Eli77, Sch16]. Suprema [Bin73b, Coh72, KK92, CS17]. Supremum [Box78, CS99, Har77, OB89, PvS21, Wes76, LL15]. Sure [DMR88, DM03, Gat00, GM96, KT01, Lin74a, Nav03, MS17, TL20]. Surface [Bal07, Hah98, HS98, LS96, Mil75, Mil80, PLRC08, HS13]. Surfaces [AA88b, TW18]. surge [GHL23]. Surplus [Nyr98]. Surprising [Sch99a]. Survey [And74, Fie75, Jag77, Luk72, Pyk75, Ren86, Rey75]. Survival [BM08, CR11, CW75, DH97, DK97, McN90, Pol74, Sud93, BMZ14, BZ14, JLW05a]. susceptibility [BLT19]. Susceptibles [Bal85]. Swaps [EM11]. Sweep [Leo09, BSS17]. Switch [TV78]. Switch-over [TV78]. Switched [Kel86, Kel88, Kog89, Mit87, Mor94, RT90]. switches [XZ20]. Switching [Coh77a, Deb76, GKL21, KKSS90, KF90, LS03, VT94, WWY12, YA00, YZB03, APZ21, FY18, HMM19, KMF22, Mar16, MMP22, ZY16]. switchover [PY13]. symbol [FN13]. Symbolic [Ken88a]. symétrique [Let82]. Symmetric [AM90, Cal03a, EG81, Gor01, Hun95, MW96b, Stu88, WH03, FJKR20, Let82]. symmetrizations [CD14]. Symmetry [ACLM77]. Symposium [Ano76k]. Synchronization [AW71, BMS89, CLM23, PBM00]. Synchronous [CS91b]. Synergetics [Hak82]. Synge [DM90]. System [ARW11, Alf03, Ano83f, Ben80, Box78, Bra98, BGPS87, Bro05, Bue88, CF91, CP90, Coh72, Coh77a, DK97, Dos78, Dur79, FS89a, Ger84b, GRW10, GV76, GS95, HR89, HS89, HP83, Iye88, Jac74a, JS80, JD89, KS83a, KS93, Las04, Le 78, LY77, LM03, LM04, MAF07, May78, MW78, Nat82, Naw91, Pap87, Pap88, Per85b, PS01, Pol75, PT77, RG89b, SS84b, SM78, SP80, Ste84, Sto98, Wu93, XZ96, YZ78, Yun86, Bar18, FX13, Kar21, LT15, LSH16, Sto15, Sto23, YBC20]. Systematic [ACO96b, ACO96a, ACO98, GR08, GACO00, Kiê98]. Systems [Aas82, Adl89, AT75, All82, Ano04f, Ath79, AJV13, BMS89, BW93, BFL04, Bat73b, Bat73c, BGZ95, BL04, BW83, BM96, BENP92, Bor06, BT08, BB10, Bur81, Cho93, CFGW87, Con77, CDL99, Daw75, DLS04, Din94, Dur80, Ebr93, Fad77, Fal84a, FS91, FVK08, FLS^{+}80, FV95, FN85, Gås12, GV98, GLM93, Geo84, GNR86, GNR87, GKRH06, GS77b, GS77a, Hak82, Han79, HI79, Heu80, Hun79, Jan84, JS83, Kai78, KP83a, KP84, Kei76, KS83b, Ken72b, KRW09, Kli80, KNT87, Kus94, LY94, Lan82, LS98b, Li98, Mal93, MD79, Maz89, MS90, Mel79b, MM83, Miy83, MM85, Nak96, Nat80, Nat86, Nat89, ON96, ÖB05, Orn77, PS04, PTW10, Pri77, RA98, RG89a, Rob92, SS92, SW91, Smi83, SE91, Sti78, Sud93, Szá76, Szp94, Tak91b, TH90]. Systems [Tau74, Teo82, TV78, Ton82, VL98, VP88, Wei95a, Wei86, Wil98, YZ02, Zaz92, ZÇ96, vdHK04, AHO20, BGV22, BMZ14, BD15, BF19, BFW22, CT18a, CAR20, DVN15, DG18, DS15, DFZ19, Fly18, HA20, HM20, JvL15, KT20, Lan13b, LR18, LF13, LSH16, NSB13, PJCB18, QD15, Qia18, SN16, Sto17, VMM19, YS23].

tagged [US19]. Tail [ASS99, DR85b, DK83, Fre82, GRZ04, GW13, HM12, KS06, KM14b, LT98, Lit69a, MZ04, MS11b, RZ13, Sat84, Sta73, Tak81, Van88b, Wil90, ZG04, BMPR19, But22, DHL17, MN14]. Tail-Ordering [DK83, Sat84]. Tailed [BN12, BZ03, CDL12, DS07, FR97, GPD00, GW13, HW09a, JT13, LZ05a, LZ05b, LM00, MB00, Pak00, Ste02, CRZ18, EKKZ18, He16, KS16, MZ16]. Tails [AK06, GG96, Gre89b, JOC12, LTW10, OC10, BPS15, Kol18]. Taken [FM09]. tan [Kem75]. Tandem [ADZ99, Box79a, Box79b, Che89b, FR80, GC02, Gre87, Har78, HK92, LW11, MS90, Num79, Num81, Sch87, TW87b, Web92]. Tangential [Mil80].

Targets [Sto80]. Tax [Gla96b]. taxation [WZ19]. Taylor [Ano04g, HH03]. Teams [KKT06]. Technique [Dan74a, Eng84, VK78, Wei86]. Techniques [Ano76k, CSH^{+}80, Dan71, DFT77a, Pri77]. Technology [CX04]. Telecommunication [Kos78]. Telecommunications [Le 74]. Telegraph [FGM12, CDIM13, Kol14, MMZ22]. Teletraffic [MD79, RT90]. Teller [BDM99]. tempering [TR19]. Temporal [Iva89, Ma05, PSS20]. Temporally [SS84a]. temps [Bré75]. Tenable [KM12]. Tends [Mal00]. Tension [MRV02]. Term [CPR01, Eng84, McN90, Nak79, PR90, Rog92, Sek06, Tam08, VV06, HSZ18]. Terminal [AT75, CM98, Lom72, Van71, Van73a, RW14]. Terms [AP88, Far76, Neu70]. Tessellation [BB01, BM03, BGR06, Cal03c, Chi94, CVS96, Deh98, HQ00, Hei98, Hor88, HO98, KT01, KT92, Muc05, Muc10, TJM98, Yao10, CH16, Cow13, GKT22]. Tessellations [AS08, BL07, BS03b, BCM13, Chi95, Cow10, HSS05, Iso00, LR11, LZ08, MM96a, MS03a, Mec96, Mec98a, MNW08, MM96b, Mil98, Møl89, Møl92, NW98, NW03, NW05, Sch03a, Sch09, ST10, TWN12, VGS10, WN99, WC11, BC16, LBCEN21, HH14, MN14]. Test [KLS02, Rob76a, Sha79, SR76, AC22, MZ20]. Testable [Sch82b]. Testing [Cur81, Dav77, Pic85a, Shi93, Wat79b, Wor01]. Tests [Hüb78, Lie76, Rav77, Ruk02, Str78, Wit79]. Tetrahedra [LS98c]. Tetrahedron [Man94]. th [Pic74b]. th-Order [Pic74b]. Their [Ahm80a, AH97, Bia02, CNT05, CY99, DRV00, Epl86, GM71, Gol77, KR85, LS13, Mar77, MD94, Mas84b, Mat73, MW96b, Mül97a, Pap77, Rao06, Roe73, Sch03a, Tan96, TT79a, Was78, Wei98, ZLB98, ENR17, Gal16, Kal18, SJ18]. Them [Sch74a]. Theorem [Abb75, AMS09, AK95, AD94, Ano98e, Bad80, BCO95, Bar74a, Ber77, Blu84, Bro95b, CQ01, CEP10, Cow06, Dad84, Din94, DM00, Ell86, EL05, FP96, FP12, Gat00, GW91, Gyi75, HM96, HM99, Hey77, Imh92, KS83b, Kes95, Kha80, KT92, Lee99b, LC76, Lin77, Mae79, Max77, McD76, Mil71b, Möh98b, MOK95, Nak79, Pak98, Pap87, PR99, Pit78, Pol80, Qi08, Ran80, RS86, Rog89, Rog92, Rog94, Rud98, SU79, Sho79, SGW96, VJ08, VS80, Vet80, WH80, Whi82a, Wil71, AS14, HMM19, KWRK22, Lac20, Liu97, Oça23, PD22]. théorème [Liu97]. théorèmes [Par98]. Theorems [AK78, Asm82, AN78a, AK98a, AK98b, Bac00, BP79, Ber87, Bin73a, BG93, Chi95, CH80, Cre79, EPS79, Gol77, GLS05, Gri75a, Gut74, HR89, Här81, HSS05, HR05b, Igl71, Igl75, Iva89, JL73, JK11, KK90, KK92, Kes72a, Kes72b, KR02, KS12, Lal84, Lee79, Lem74, Lin74b, Luk77, McD78, MY01b, Mol93, Nor75, Num76, PTW10, Pak71b, Pak79, Pak92, Pak10, PSS06, PW10, Pim07, Rah07, Ric11, Ros80, Ros78, Sch02, Sco73, Ser75, Shi80, Sil76, ST98, Smi83, Sta78a, Sta78b, ST94, Voi90, VML80, Woo80, ZK89, BCLF16, BD17, Clo17, FHJK20, FGZ17, FDJ20, FPS22, FMY22, GW20, GLL22, HS13, JY20, Jir13, Kom18, LPST14, OT20, Par98, PV19, TO21, TL20, TW18, vdHvdHLS20]. Theoretic [Bin80]. Theoretical [CMT21, Jag18]. Theory [Adl89, Ahm81, Als88, AH97, And74, Ano93e, Arj72a, Arj72b, AK78, AN78a, Bal78, Bar78b, Bas79, BVH10, Bin73a, Bin75, Bin91, BG07, BR83, BN92, CG80, CHJ16, Cha92a, Çin69, Çin74, Coh73, Coh82a, Cow89b, CDH87, DR85b, Eba78, Edw74, Fad79, FG00, FL01, Gal79a, Gan69, Gan79, Gat85a, Ger84a, GS77a, Gri81, GR96, Hai75, Har80, HS94, Hey74a, Hey75, Hös11, Hun74a, Hun74c, Hun74b, Hun77, Igl73b, JOC12, Kar85a, KKSS90, Kha80, Kli80, KOS03, KR71b, Lal84, LP93, Luk72, Mar75, Mel79b, Mit88, Moy72, Nag71, NSHH86, Nec78, Neu78, Neu86b, Off74, Pal93, PG97, Per97, PC96, Pro76a, Pro76b, RS71, Ros78, Sch74b, ST10, Sch77, Sen74, SH89, Smi82, Sta78a, Sti72, Szá76, Tak75a, Tak75b, Tak76a, Tak76b, Teu73, Van84, VP88, Van88a]. Theory [VW77, Van74, Vet81, Wat74, Wei71, Wen89, Wil88, YZ01, Fas14, Nyr16]. Thermal [DH82]. Thermodynamic [Sch05]. Thermodynamics [Ton82]. These [Sam04b]. Thick [HJ03, Kid96, MP03]. Thin [BBSS07, Cza23, GG96, HJ03, Pak00, Rat04]. Thin-ended [Cza23]. Thin-Tailed [Pak00]. Thinning [BO07, CFD85, JL73, LTT20, MS10, Sha03]. Thinning-Replacement [BO07]. Thinnings [Ser80, KL13]. Third [Lit74, Ors90, SB98, vdHK04]. Third-Generation [vdHK04]. Third-Order [Ors90]. Three [Abb75, AST10, Ala99, FFT75, Fie76b, FK89, Iso00, JL73, LO04, LC92, May78, Pit74, Pit75, Sar97, Sti06, Van71]. Three-Dimensional [Iso00, Pit74, Pit75]. Three-Handed [Sti06]. Three-Locus [Fie76b, May78]. Three-Node [FK89]. Three-Parameter [LC92]. Three-Terminal [Van71]. Threshold [BST09, BKO11, BFL09, BW97, Cli11, FS91, KK90, KK92, SBT96, VML80, Wil71, Bar18, CKM19, Mey18]. threshold-based [Mey18]. Threshold-Stopped [KK90, KK92]. Throughput [AT90, Kel82, MS90, Naw91, TE96]. Throughputs [Cao89]. Thrown [Kin77, Wil85]. Ties [Cam84]. Tight [SS95a]. Tightness [Szc99, Sto15]. Time [Aas82, Agr74a, Agr74b, AT75, Alf03, Alp09, Ano86e, Ano93g, Ano04f, Ant78, ADW^{+}75, BBHM11, BS88, Bal90, BMY91, BMY94, Bal97, BMTY97, BNJS98, BB07b, Bat76b, Ber78c, BV93, BR86, Bin75, Blo83, Bob04, BDD12, BKM92, Bro85, BW97, Bro99, BNR87, Cai07, Cha09b, CA94, CR04, CPZC05, Chi95, CY00, Cli11, Coh91, Cox80, CI86, DN78, Dav73b, Dav79, De 10, De 74, De 83b, DDH78, Der83, DGN92, DNPR01, DKS84, DLO06, DH76, DR81, EDH90, Elv71, Epl86, EN80, EN88, FS89b, Fei76b, FL09, Fin69, FGM12, Fra85, Fre82, Fro04, Fuj79, Fuj80, FN85, Gal79b, GPD00, Ger80, GNR87, GNRS89, GNR90, GKT21, GM11, GHL07, GY10, GW13, Gzy80, Haj82, Hal86, Han78, HK84, HMS03, HS74, HK88, HC08, HT92, HW09a, Hoo78]. Time [Hor88, HIS76, HM10, Hua97, HT10, Imh86, JLW05a, JL79, Jac05, Jac80b, JH92b, JK11, Kap73, Kei76, KS83b, KS83a, Ken74, Ken78b, Kes80, KD94, KS06, Kim83, KM80, Kne00, KLS02, Koh78, Koh80a, Koh80b, KKSS08, KNT87, Kyp71, LL81, Leh02, LS89, LTW10, Lie76, LS13, Lin10, Liu90, LT73, McG82, MP08, MW96a, MT92, MT93a, MT93b, MVY84, MM83, MRSV86, Möh04, MM85, NM79, NP93, Nat80, Nat86, Neu76a, Nor82, ØSZ11, Ott77a, Ott87a, Pak89b, Pak89a, Par77, Pes82b, Pho79, Pic85a, Pol87, PR90, PC96, Pro76a, Pro76b, RW91, Rit94, Ros80, Ros76a, ST96, SW90, Sal88, Sam04a, SC83, Sav75a, Sav75b, Sch90, Sch07, SS84b, Sha97, Sho74, SW93, SR00, SZ11, Sta92, ST90, Ste80b, SBT96, Sys76, Szc90, ST94]. Time [Szc99, Tak81, TH90, TT04, Tav79, Tay79, Tie81, Tie82, Tjø90, TC92, TF08, Van90, Van88b, Ven82, Wal73, Wit79, YC96, YZ02, YZB03, ZCWS96, Aaz13, ABL15, Ano09e, BPP21, BP23, BDK^{+}18, Bez20, BSS23, BD22, BLT19, CMMS13, CM17, DP15, EW14, EP15, EG17, FRLM20, Gol16, GY16, GR16, GVZ13, GHH15, GKPZ21, HHBV20, HMM19, HR16, KW18, Kub19, Kur18, LF13, MS17, Mar16, MMP22, Mas15, NP17, Oça23, PR22, PZ17, RSW16, STZ14, US19, WZ19, XHW23, YS23]. Time-Advanced [ØSZ11]. Time-Average [BR86]. Time-Changed [FL09, Gzy80]. Time-Dependent [Kim83, Kne00, LTW10, Ros76a, STZ14]. time-inhomogeneous [Oça23, RSW16]. time-limited [HHBV20]. Time-Non-Homogeneous [GNR87, Hor88]. time-reversibility [BD22]. Time-Reversible [Ken78b]. Time-Series [Dav73b]. Time-Shared [Kei76]. Time-Varying [Hal86]. Timed [Wal81b]. Times [Adl89, Alb83, BS81, Bar78a, BH04a, BOR02, Blo89, Box78, Box79a, Box79b, BB10, Che89b, CCH00, Coh72, Cse92, Dad82, Dad84, De 73a, Di 01, DM00, DW02, Fer93, FL01, Gau02, Gha08, GW92, Gor01, Gre78, Gut74, GH75, Had76, Hai80, Har84, Hor77, Jac80a, JH92a, Jan84, Jan86, Kal80, Kei74b, Kei74c, KP83b, KS04b, KW03b, Kur78, Lal84, Lam03, LL97, MY01a, Min80, Neu77a, Neu77b, Neu77c, NPOTC00, Pfe83, Pol74, Pro76b, RA98, Res74a, RL77, SS95a, Sad95, Sha84, Sha80, SY97, Tam74, Tav76, VY94, Van78b, WV80, Wan01, Web92, WH03, Wis72, Ale13, DL20, FPZ13, GG17, Kom18, KJH18, KZ16, Lem14, LZ13, MS16, ORT16, RW14, STZ14, Zha15]. Timid [JK85]. to/from [Sal88]. Tommassen [Rog89]. Tomography [KR97]. Tools [Ste75b]. Topics [Mai80]. Topological [Deh98, Lea76, Ott75, WC11]. topology [TYA16]. Torus [Qia90]. Tossing [Che89a]. Total [AAL08, AD94, Bal86, BGRSC10, CFW09, Cou90, DHP12, Hol76a, Hüs89, KS83a, Ker78, KLS02, LL97, Möh00, Pak71b, PW06, Tur74, AD21, BWZ19, DY15, DP13, FFI22, GVZ13]. Tournament [KKT06]. Tower [BLT03]. Trace [SS77]. Tracing [BKO11, Tho82]. Tracking [SK04]. Trade [Lat77]. Trade-off [Lat77]. Trading [CS12, RW14]. Traffic [AH05, AR92, BB93a, BGZ95, BMV03, Bro72, Dag78, DR92, FQVJ71, Fal89, GW91, Gri94, HR89, Har78, HJKP17, IW70a, IW70b, Igl71, Jac80a, Ken72a, Ken72b, Köl81, KM95, Kos78, Kus94, Kyp71, Le 74, Maj07, Mar75, Mel79a, Mel79b, Ott77b, Rat75, Rei88, Sch87, Sta78a, Szc90, Szc99, VL98, Wan97, Wei76, Wei86, Whi70, WH80, Whi82a, Yeo78, LF13, PY13, VBvdM15]. Trains [Fie75]. Trait [AH78, Sim81, Vet78]. Traits [Hil76a, HA79, BE21]. Trajectorial [GKT21]. Trajectory [Bal86]. Transaction [IS06, Tam08, CIL17, DK16]. transactions [Gof19]. Transceiver [BBW09]. Transects [CB95]. Transfer [LW92]. Transfers [MT97]. Transform [BS01, HR05a, KLS02, Ott87a, SK85, Swi75, Xu11, Pri15]. Transformation [Cal03a, Liu98, Whi92a]. Transformations [Bal75, BK98, JN98, Kem75, Kes70a, Kes70b, KRS07, O’B80, Buc16].

Transformed [CM19b, CK07]. Transforming [MB12]. Transforms [AW88, AW96, GKW96, IJ02, RL77, Sak04, SM87, LZ13]. Transience [Ano79e, Hut80, Lev79, Bau14, Die82b]. Transient [AW87a, AW87b, AW88, BD96, BFK92, Cho93, Con83, DN78, FR12, GS95, HHBV20, Kei76, Kne91, LPRS93, Mas84b, Par87, Yun86, FR14, FFI22, GNR17]. Transients [Cam79]. Transit [KKSS90]. Transition [Abr79, AM90, Blo89, CPZC05, CH76, Gri79b, Lin05, MK01, Per85a, Pes75, PS94b, SB09, Wac09, ZLB98, AZ18, GLM21, GHH15, XHW23]. Transitions [BV91, Lan05, Rud73, JW15]. Translation [Soo10]. Translation-Equivariant [Soo10]. Translative [Rat96b]. Translocation [Ste76b]. Transmission [Ott06, Fox14]. Transparent [GG12a]. Transport [DK12, HJKP17]. trapped [Bow21]. Traps [Bra98]. Traveling [Sto23, RY14, RSY23, Ste15]. Treatment [BS83]. Treatments [Bat80]. Tree [Ano92e, BLP06, BR04, Har71, Har84, Jof75, KM10a, Min05, Pen98, PW06, Pit86, RSZM02, SK96, YS94, AD21, BCT13, BP17, BP22, OY22]. Tree-like [Har84]. Tree-Shapes [Har71]. Trees [BNK98, BV06, BMS09, Cro07, Cro10, DM03, DH05, Hol11, JOC10, JOC12, KW03a, Kho01, Lee99b, LMW96, PW04, SGW96, Ber20, COC16, Cra17, HNGS15, Iye23, KXZ23, Mey18, Seg20, ZM16]. treewidth [LM17]. Trend [BR87, Nos94, GLS15]. Trente [EL05]. Triads [Ken88a]. Trial [Imh08, Whi96]. Trials [BE83, Bat80, BL88, FJ09, Tam11, Ver73, CDIM13]. Triangle [Ken96, Ken98, Man88, Man90a, Man90b]. Triangles [Gat94, HQ00, Ken85, Le89a, Le89b, Wat86b]. Triangular [HZ06, Wie81]. triangulation [CD18]. Tribute [DKMJ10]. Tridiagonal [MT06b, SB09]. Trigonometric [TW84b]. trimming [GML18]. Trisomics [TS81]. trivariate [FPZ13]. Trivial [FMP92, MN14]. Truncated [Cov09, FFG07, MSV01]. Truncating [VPCG10]. truncation [LL18]. truncations [Mas15]. Trunk [FOF06, Kus94]. Tsetlin [Die82a]. Tumor [HRTY97]. tumors [EKKZ18]. Tumour [Mol74, Wil74]. Tunnel [BND81]. Two [AWZ93, ADZ99, AMS07, AMS09, Alm97, And75, Ano83f, Ano04f, AGSC04, AT81, Ave79, Bal79, BDM85, BX01, BL06, BP83, BF80a, BF80b, Bor98, BK01, BW97, Cal03c, CR11, CNP91, Coh85, Cre74, CJ11, CR78, Dal85, DS93, Deh98, DH04, DGN92, Elv71, Eth93, EN80, EN88, EG90, FH02, FKM82, FX00, FPP19, Fin92, FS08, GT84, Gat85a, Gec77, GV76, Gor01, GJS08, HKT91, HP82, Hau78, Hol83, Hoo89, Hun74a, Hun74c, Hun74b, Hun77, Hut80, IL73, IL74, Ill01, IM02, Jac05, JH92a, KK80, Lom72, MA95a, MS80b, MSV01, Mer06, Möh98a, Nat82, Neu70, NS77, PT94, Pea89, Pel83, Phi78, Pon05, Roe73, Rog90, Ryc92, Sch74a, Sch86a, SW13, She85, SW98, Str75, Sza93, TV78, Wat79b, Wri92, XZ96, Xu11, Yas85]. Two [YZ02, AAM17, AG13, BCT13, BDLP14, BSS17, CKL14, HMM19, JKR17, KM14b, Kol14, KZ16, Lin15, Mon17, NSB13, NLM18, PP15, STZ14, Sar16, Tam16b]. Two-Alleles [Wat79b]. Two-Armed [BF80a, BF80b]. Two-Barrier [Ryc92]. Two-Choice [AGSC04]. Two-Color [AMS07, AMS09]. Two-Dimensional [AWZ93, Alm97, BX01, Cal03c, Cre74, CJ11, DGN92, FKM82, HKT91, Hut80, Ill01, IM02, KK80, MSV01, Pea89, BCT13, KM14b, KZ16]. two-factor [BDLP14, JKR17]. two-filter [NLM18]. Two-Horse [BL06]. Two-Island [Gor01]. Two-Level [Eth93, HP82]. Two-Locus [CR78, EG90, GJS08]. Two-Parameter [Mer06, Xu11, MS80b]. Two-Person [And75, Phi78, Pon05, Mon17]. Two-player [FPP19, HMM19]. Two-Point [DS93]. Two-Server [GV76]. Two-Sex [Dal85, GT84, Möh98a]. Two-Sided [Jac05, Rog90, Yas85, CKL14]. Two-Stage [ADZ99]. Two-State [JH92a]. Two-Station [XZ96]. Two-Terminal [Lom72]. Two-Type [CR11, DH04, AG13]. Type [APP99, ALS08, BFL04, BY89, CR11, CFPL12, Dag75a, Dag75b, DH04, Fei76a, GHT96, GHT97, Gal79b, GY78, Han96, Hir00, HS90, Kem73, Lat88, LP96, Lef07, LO04, LZ05a, LZ05b, MSM92, MY89, MS11a, MZ96, Nar87, Neu76a, NPOTC00, O’C91, Ott87a, PW04, RS87, Rem00, TV10, VY94, VL07, Wad07, Whi80b, Woe07, ZL09, Zuy99, AG13, BPP21, BLT19, FHJK20, Fox16, Thé22, HW09b, MZ04]. Type-III [HW09b]. Types [BS12a, Bor98, BGPS87, CR11, De 83a, Esh72, Esh73, Fin92, FM09, GB03, Her73, Her75, Hol83, KK73, MS12, MA95a, Neu77a, Neu77b, PL99, Whi80d, HLN13]. Typical [AS08, Cal02, Cal03b, Cal03c, Chi94, DMM12, GB03, Ols07, Yao10, GKT22].

Uhlenbeck [ALS08, BR15, BPS15, HJM^{+}16, Sal84]. Uhlenbeck-Type [ALS08]. Ultimate [Coh82a, PvS21]. ultrametric [DG18]. Ultrasmall [DMM12]. Unavailability [Nat80]. Unbiased [HS98, JLY22, JK94, Kün83, LLX18]. Unbounded [Gal11, JY04, Rob76b, Van88a, BS15, BFW22, GHH15, KM14b, DMYT23]. Uncertainty [AS79, Chi80, Coh85, DDSV87, Pet84, Pet82, EC21, FS16a]. Underlying [RT02, Tol88, GC20]. Undershoot [CWW09]. Undershoots [PM08]. Undiscounted [Gla96b, Mon17]. unequal [Nak14]. Unification [CS97]. Unified [Bal86, BDM99, BM96, CM00, KI92, Mül97b, PL90, PL96, RS86, Yas85, Miy17]. Uniform [Ais79, AR97a, AR97b, BG07, BP94, CV18, CFGW87, CM17, DHL17, Gro12, HC08, Jen84, Jen88, Mae79, Mil69, MS15, MS16, Rei81, Smi82]. Uniformization [Van90]. Uniformly [Min80]. Unilateral [BR93, CI00, Pic80, Tjø83, TP92]. Unimodal [BN10, Lep76a, Cap86]. unimodales [Cap86]. Unimodality [Kla85, Wol78]. Union [Buk01, Kwe75]. Unions [MH08, Sch02]. Uniqueness [CPZC05, Qia18, RSY23]. Unit [GV76, HSS08, Whi80b, Whi80d]. Units [Gat85b, MA95a]. Univariate [KR81a, MS80a, Cha22]. universal [ILP18]. Unknown [Duf77, Wor01, YZ78]. Unlimited [AHO96, Ada16]. Unobservables [Whi69]. Unreliable [Gat85b, KD94, LS15]. Unsigned [Adl89]. Unsolved [Kin77]. Unstable [OC10]. until [Pau07]. Unusual [LP93]. unusually [Rob18]. Upcrossings [Adl76, Adl77, HP78, Sha79]. Updating [BL09b]. Upper [Buk01, Cov09, Fro04, GW13, HMM06, Kat90, KS74b, Nat86, Mas23]. Uptime [LBZ98]. upward [KM14b]. upwards [AV19]. Urn [AMS07, AMS09, Bin91, Che00, ENPS78, KM12, KP12, MPP07, PS04, Pal93, Pal94, Sta98, Tam16b, AC22, BK18, Blo89, HH85, RW77, Ste18]. urns [Jan20, ZHCC14]. Usage [MM85]. Use [Arj72b, CT85, Gri75b, Sha79, Wei98]. Used [AA88a, KS78, MS83, Sha84, FN13]. User [Thö99]. Using [Bar79, BGRSC10, CB03, Cur74, Cur81, DMR13, DW06, FV95, GR08, Gha08, HI79, JS83, KM00, Koh80a, Koh80b, MB12, Pes75, PR90, San98, Sch76b, Sco73, TW87a, CDJR22, HSZ18, TSvB17, Lef88]. USSR [BRB^{+}71]. Usual [Bal78]. Utilisation [Lef88]. utilities [HLG18]. Utility [FS08, GKPZ21, SS20]. utility-maximising [SS20].

V [Ano04e, Ant79]. vacances [Fri86, Fri87]. Vacant [Sar97]. Vacation [TH90, Van78b, WD90, Bar18]. Vacations [FS91, Fer93, KW90a, LZ95, LMHN90]. Vaccination [BJML07]. Validity [Rob74a]. Valuation [FS08, EP15]. Value [Ber87, BVH10, Cai07, Chi94, Coh82a, DS93, DR85b, Gal79a, GR88, HS10, Ker86, Per97, PC96, Res74a, SF79, Smi82, Sta92, Tak88]. Value-Iteration [SF79]. Valued [Ahm80b, DLO06, EG90, Kue75, McD78, O’B80, PS94a, Ran80, SVW83, Ver75, Clo17, Qia18]. Values [AAL08, GM96, HH81, Lin80, Per97, Pfe85b, Pri99, Sch06b, Str75, WH03, ZG04]. Vanishing [CCHH18, CIL17]. Variability [BD87]. Variable [AFGR06, Fil72, Fil74, Gal11, KP86, Kem87, Led79, Sim81, Sto98, TV78, HL17]. Variables [AK93, ASC96, BE85, Bin73b, BV02, BR91, CW03, Cek03, CS91a, CNT05, Chi12, DN02, GM71, GPD00, GH75, HS77, Jof75, KK90, KK92, Ker86, Kue75, LL81, LS93, LR80, McD78, Mij80, Phi78, PS94a, PT77, RN70, Res72, RW85, Rüs82, SS90, Sha87, Sil76, Tak70, Tuc71, Ver75, Ver79, Wat86a, Yu08]. Variance [AMNW11, AHO96, ACO98, AH78, BP00, CS12, DP12, FH04, GP15, GACO02, HKT91, Hey74b, HJ02, KP10, KS74a, KS12, LH07, Lie76, LY12, Mij80, Pak10, Pic00, Pit77a, PS10, Ram84, Roe74, Sat85, Sch02, SO85, Sma76, Wit79, GY16, Jan20, MV19, OS18, PSS20, SK20]. Variance-Optimal [KP10]. Variance-Time [Lie76]. Variance-to-Mean [BP00]. Variances [HR05b, Mrk01, Yao10, Buc13]. Variant [FGM12, VS87]. Variates [Geo76]. Variation [AAL08, BW90b, DMR13, EK12, Fad90, Gra82, HR05a, HA79, HL06, MZ77, Möh00, Res80, Res86, Rob76a, Sta73, Teu73, AF18, Gal16, Lem14, MW21, Seg20]. Variational [Ant79, Leó76, MZ98, PC21]. Variations [Abb75]. variétés [FG81]. Variograms [GSS01, Ma05]. Various [Neu76a, NR85, Sud93, Tak91a, Tam74]. Varying [BL08a, D’S94, Fas05, Fre82, Fuj79, Fuj80, GNR90, Hal86, Iva83, Kle82, Lin74a, MR09, Sar98, Sen74, SS84a, BSS17]. Vector [Bar01, DDH78, DH76, LS98a, Led04, Neu78, Din19]. Vectors [DZ11b, ENM89, HL06, LZ05a, LZ05b, OMS06, LY15a]. Vehicle [BV93]. Velocities [Di 01, Mol72]. Velocity [De 10, Nar91a, OD07, CDIM13]. verification [HMM19]. Version [GP88]. Versions [Jen95, KS90, LC76]. versus [Che00, DK80, HJ02, Seg20, SWL14]. Vertex [AR97a, AR97b, LR12, ST10]. Vertices [CS06, Man94, Mas00, Min05, BP17, BP22, TW18]. Vervaat [Rog89]. Very [Gan78]. VI [ACV80]. Via [BHJ11, DGP10, AW88, AW96, AV19, BB93a, Bäu02, CLM08, Cho93, GG17, GNRS89, Hey74a, JML92, KS83a, KT20, KOS03, KK03, Lac20, LF13, LL18, MV19, MS16, NW22, Pal93, Pal94, PR94, SS95b, SW93, XZ20, ZG04, AK08]. Vibrations [Eli77, Ors84]. Vidanges [Dio78, Dio76]. View [Dob77, GW91, Le95, Lea76, MC09, Sen78, GR16]. viewpoint [Bou16]. Viot [Fou11, Gri14, Hir00]. Virtual [BB93a, Coh72, CI86, Kyp71, Min80, Ott77a, Ros80, SW90]. Virtually [Mih80]. Virus [PT85a]. Visibility [KM10a, TC13]. Visible [Gal11]. Vision [MGW96]. Visit [Hoo89]. Vital [Coh77b, Coh77c]. Volatility [AE08, GKT21, KP10, KS12, MS11b, Sin98, Woe07, BV13a, BV13b, JL15, RT21]. Volume [AS08, Ano71f, Ano72g, Ano73h, Ano74k, Ano75j, Ano76l, Ano77k, Ano78k, Ano79k, Ano80j, Ano81i, Ano82i, Ano83k, Ano84k, Ano85j, Ano86k, Ano87i, Ano88j, Ano89j, Ano90i, Ano91j, Ano92j, Ano93l, Ano94j, Ano95j, Ano96i, Ano97i, Ano98k, Ano99i, Ano00i, Ano01i, Ano02i, Ano03j, Ano04l, Ano05i, Ano06k, Ano06l, Ano07j, Ano07k, Ano08i, Ano08j, Ano13a, Ano13b, Ano14a, Ano14b, Ano15a, Ano15b, CB03, DSS99, JK94, Mån07, Man94, Mat98, Nie98, Nie99, Vit95, Yao10, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano17h, Ano17g, Ano18b, Ano18a, Ano18d, Ano18c, Ano18f, Ano18e, Ano18h, Ano18g, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano20h]. volume [Ano20g, Ano21b, Ano21a, Ano21d, Ano21c, Ano21f, Ano21e, Ano21h, Ano21g, Ano22b, Ano22a, Ano22d, Ano22c, Ano22f, Ano22e, Ano22h, Ano22g, Ano23b, Ano23a, Ano23d, Ano23c, Ano23f, Ano23e, HS13, HH14, Pri15]. Volumes [BFV10, Cow07, NR85, Ols07, Rat98, Sch02, Vit01, Wei01, Yao10, HW19, KR18, Sva14]. Voronoi [AS08, BB01, BL07, BS03b, BO07, Cal02, Cal03b, Cal03c, CVS96, FZ96, LBCEN21, GC20, HQ00, HQ02, Hei98, HRS04, Iso00, KT01, KT92, Muc05, Muc10, NRT83, NR85, Ols07, RSZ12, Sch94, Yao10]. vortex [Cor14]. Voter [Sud99]. voting [GMS23].

Waiting [Ber78c, Blo83, BGPS87, Coh72, Coh91, CI86, DN78, DKS84, Dsh85, EDH90, Fre82, KS06, Kyp71, LT73, Min80, MM83, MM85, Neu76a, Ott77a, Ott87a, RL77, RW91, Ros80, Ros73, SS95a, Sad95, SW90, SC83, Son79a, Son79b, Szc90, ST94, Szc99, Tak81, TH90, Tam74, Tru75a, Tru75b, Van88b, Ale13]. Waiting-Time [CI86, DKS84, Fre82, Neu76a, Ott77a, Ott87a, Tak81, Van88b]. Wald [Kry76, BR91]. Walk [Ano85e, AG05, Asm82, Bac00, Big78, Bro80, Cre74, CWW09, DO74, DS85, DS07, DGN92, Dom07, Fuh97, GS77b, GT78, GP80b, Hog86, Iva89, KP86, Kem87, KP94, Kle82, Kry76, LYA07, Liu98, Pak89a, Per85a, Sie79, Val96, WR76a, Als18b, BMZ14, Che15, GLL20, GN23, HLZ18, KM14b, MV19, Nes17, PW18, Pin16]. Walking [GY78]. Walks [Ath85, Bar78c, Ber78a, BD96, BS05, Bin80, BL08a, CDL12, CCH00, Cro10, DM00, FS09b, FL01, Fuh07, FP12, Gat00, Gre89b, Hög76, Igl75, IM08, Jan86, JK11, Ken74, Muk75, Ser81, Wac09, Woe85, Ath18, AV19, Bau14, BZ14, BMPR19, Bow21, Gal84, HL14, HR16, MY20, SK16, Vid20]. Walsh [Koh80a, Koh80b]. Wandering [Shi82]. War [BC78, BC81a, Whi96, EG17]. Wasted [SK08]. Watson [BD74b, BD74a, BM08, D’S94, Dou75, FQVJ71, FV99, ILP18, Jir76, KM10a, KRS07, KXZ23, Lin74a, Lin74b, Min05, Qui72, SS09, SJ18, Sen69, Sri85, Sze76, Ven82, Wau81, SS15]. Watterson [Tra74]. Watts [ABL15]. Wave [Åbe07, Dan77, HR04, Lin72, PR99, Sut79, Sto23]. Wave-Functionals [PR99]. Wave-Length [Lin72]. Wavelength [Ryc87]. Wavelet [AK05]. Waves [Lin06, Lin10, RY14, RSY23]. Weak [Ald80, Bal75, BB17, BK99, CY99, Dou75, Hof07b, Hor91, Igl73a, Igl73b, JM86, KW88, MS76, Pol79, Res80, Res86, SCA00, SWL14, STC94, Taq75, Whi86, Wil88, HSZ18, Sen18b]. Weakly [Hol11, Lat88, Lie76]. Wealth [KLS02, PHL16, RW14]. Wear [RG89a, RS71, CM19b]. Wear-Dependent [RG89a]. Weather [Gri75b]. Web [VL10]. Weibull [Pap88]. weight [GLM21, ZM16]. weight-dependent [GLM21]. Weighted [CNT05, Dom07, HM12, Käm87, Lee99a, OC12, OT02, Sch09, Wes76, Yu08, COC16]. Weinberg [Cha74, MS74, RH81, SR76]. Welfare [EX11]. Well [Wal81b]. Well-Timed [Wal81b]. Welsh [Van83]. WFK [Ant79]. wheel [Cer17]. Which [Cam88, Che00, Gla83, HR89, JH92a, Mas79, McN90, Neu78, SM78, SW80, YZ78]. White [HW95, Ors84]. Whitt [Ano04e, AJV13, GM08, PR00]. Whittle [AM22]. Who [BLT19]. Whole [BRB^{+}71, MS03b, SW97]. Whose [Ano85e, DS85, Dan89, FS82, Man94, Sen89b]. Wicksell [TK13]. Width [DRS11, DH05]. Wiener [BK80, Dag75a, Dag75b, Gre75, LL89, Pra73]. Wilkinson [GZ93]. Williams [Cha96]. Win [Pon05]. Wind [BND81]. Wind-Tunnel [BND81]. Winding [BR88]. Windings [JV18]. wins [BMPR19]. Wireless [BB01, BBSW10, BY09, SSS04, vdHK04, HJKP16]. Withdrawal [Kha86]. within [Gof19, Pap88]. without [AW71, BK02, Le 74, PP15]. Wolfowitz [RT23]. Word [Ano06f, GS93, Hua02, KO87, Rød06, RS07]. Words [Ruk02]. Work [Gan71, Rey75, Smi74]. Workers [AR06]. Working [RS21]. Workload [Hok85, Las04, BPS15]. World [CT08, GX07, Gha08, ABL15, BR13]. Worst [CNP91]. Worst-Case [CNP91]. Worth [BK91]. Wrapped [BB92]. Wright [Bob08, BC04, CV23, GPHK21, Gri14].

x [Kem75, BE06, Imh92]. X-Ray [BE06]. Xerographic [Bro80]. XX [Ben80]. XX/XY [Ben80]. XY [Ben80].

Yaglom [BOP21, FM18]. Yao [DGM09]. years [Jag18]. Yor [FGZ17]. Yule [BB17, RS09]. zag [BD17]. Zernike [LZ17]. Zero [AW87b, DTM07, GHL07, KM12, Mal00, Tay89, VPCG10, BP23, FPP19, HMM19, HH14, MN18, Mon17]. Zero-Automatic [DTM07]. Zero-Balanced [KM12]. Zero-Sum [GHL07, FPP19, HMM19, Mon17]. Zero-Truncating [VPCG10]. Zeros [SM87]. ZFY [DAG95]. zig [BD17]. zig-zag [BD17]. Zinger [Gyi75]. Zone [GW78, Wes79]. Zonoids [DS03].

References

Aronowich:1985:BPE

[AA85] M. Aronowich and R. J. Adler. Behaviour of χ^{2} processes at extrema. Advances in Applied Probability, 17(2):280–297, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427141.

Aronowich:1986:ELC

[AA86] Michael Aronowich and Robert J. Adler. Extrema and level crossings of χ^{2} processes. Advances in Applied Probability, 18 (4):901–920, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427255.

Abouammoh:1988:NBT

[AA88a] A. M. Abouammoh and A. N. Ahmed. The new better than used failure rate class of life distribution. Advances in Applied Probability, 20(1):237–240, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427280.

Aronowich:1988:SPB

[AA88b] Michael Aronowich and Robert J. Adler. Sample path behaviour of χ^{2} surfaces at extrema. Advances in Applied Probability, 20 (4):719–738, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427357.

Aldous:1992:IP

[AAB^{+}92] D. J. Aldous, R. A. Arratia, A. D. Barbour, B. Bollobás, P. Diaconis, P. J. Donnelly, P. Erdős, L. H. Harper, B. Harris, V. F. Kolchin, A. M. Odlyzko, J. W. Pitman, B. G. Pittel, L. A. Shepp, J. Spencer, L. Takács, V. A. Vatutin, A. M. Vershik, O. V. Viskov, and H. S. Wilf. Invited papers. Advances in Applied Probability, 24(4):763–773, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427710.

Ahmed:1986:DSO

[AABK86] A. N. Ahmed, A. Alzaid, J. Bartoszewicz, and S. C. Kochar. Dispersive and superadditive ordering. Advances in Applied Probability, 18(4):1019–1022, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427262.

Adell:2008:EVS

[AAL08] José A. Adell, José M. Anoz, and Alberto Lekuona. Exact values and sharp estimates for the total variation distance between binomial and Poisson distributions. Advances in Applied Probability, 40(4):1033–1047, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443623.

Albrecher:2017:ODS

[AAM17] Hansjörg Albrecher, Pablo Azcue, and Nora Muler. Optimal dividend strategies for two collaborating insurance companies. Advances in Applied Probability, 49(2):515–548, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/optimal-dividend-
strategies-for-two-collaborating-insurance-companies/
726C969E06470DC887D50CE4D89722F4.

Angulo:2005:FKE

[AAMRM05] J. M. Angulo, V. V. Anh, R. McVinish, and M. D. Ruiz-Medina. Fractional kinetic equations driven by Gaussian or infinitely divisible noise. Advances in Applied Probability, 37 (2):366–392, June 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/30037332.

Altman:2004:PAD

[AANQ04] Eitan Altman, Konstantin E. Avrachenkov, and Rudesindo Núñez-Queija. Perturbation analysis for denumerable Markov chains with application to queueing models. Advances in Applied Probability, 36(3):839–853, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/4140411.

Allman:2008:IMM

[AAR08] Elizabeth S. Allman, Cécile Ané, and John A. Rhodes. Identifiability of a Markovian model of molecular evolution with gamma-distributed rates. Advances in Applied Probability, 40 (1):229–249, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443578.

Aase:1982:SCT

[Aas82] Knut Kristian Aase. Stochastic continuous-time model reference adaptive systems with decreasing gain. Advances in Applied Probability, 14(4):763–788, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427023.

Aazizi:2013:DTA

[Aaz13] Soufiane Aazizi. Discrete-time approximation of decoupled forward–backward stochastic differential equations driven by pure jump Lévy processes. Advances in Applied Probability, 45(3):791–821, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1377868539.

Asmussen:1994:SPA

[AB94] Søren Asmussen and Mogens Bladt. A sample path approach to mean busy periods for Markov-modulated queues and fluids. Advances in Applied Probability, 26(4):1117–1121, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427907.

Averkov:2007:RCB

[AB07] Gennadiy Averkov and Gabriele Bianchi. Retrieving convex bodies from restricted covariogram functions. Advances in Applied Probability, 39(3):613–629, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443532.

Anantharam:2015:CEE

[AB15] Venkat Anantharam and François Baccelli. Capacity and error exponents of stationary point processes under random additive displacements. Advances in Applied Probability, 47(1):1–26, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1427814578.

Anantharam:2022:NES

[AB22] Venkat Anantharam and François Baccelli. Nash equilibrium structure of Cox process Hotelling games. Advances in Applied Probability, 54(2):570–598, June 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/nash-equilibrium-structure-of-cox-
process-hotelling-games/161705657FFBC42DBF93469501BD6B38.

Abakuks:1974:OIP

[Aba74] Andris Abakuks. Optimal immunisation policies for epidemics. Advances in Applied Probability, 6(3):494–511, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426230.

Altenburg:1980:SME

[ABA^{+}80] Hans-Peter Altenburg, Wolfgang J. Bühler, Donald G. Aronson, Norman T. J. Bailey, Konrad Bögel, Odo Diekmann, P. C. Fife, L. A. Peletier, Denis Mollison, F. H. Rijsdijk, J. C. Zadoks, F. Rothe, Stanley Sawyer, K. Schumacher, Horst R. Thieme, and Hans F. Weinberger. Spatial models in epidemiology and genetics. Advances in Applied Probability, 12(3):559– 566, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426417.

Abbott:1975:VTS

[Abb75] J. H. Abbott. Variations on the three series theorem. Advances in Applied Probability, 7(2):252, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426066.

Alt:1980:MCM

[ABB^{+}80] Wolfgang Alt, George I. Bell, Howard C. Berg, Evelyn Fox Keller, I. Richard Lapidus, Michael Levandowsky, Douglas Lauffenburger, and Ralph Nossal. Models for cell motility. Advances in Applied Probability, 12(3):566–569, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426418.

Aaberg:2007:WIS

[Åbe07] Sofia Åberg. Wave intensities and slopes in Lagrangian seas. Advances in Applied Probability, 39(4):1020–1035, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/20443556.

Ansell:1980:NRP

[ABH80] J. Ansell, A. Bendell, and S. Humble. Nested renewal processes. Advances in Applied Probability, 12(4):880–892, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426746.

Addario-Berry:2015:MTN

[ABL15] Louigi Addario-Berry and Tao Lei. The mixing time of the Newman–Watts small-world model. Advances in Applied Probability, 47(1):37–56, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1427814580.

Abraham:1979:PTD

[Abr79] D. B. Abraham. The phase transition in the discrete Gaussian problem. Advances in Applied Probability, 11(2):303–304, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426834.

Assaf:2000:OSR

[ABS00] David Assaf, Yuliy Baryshnikov, and Wolfgang Stadje. Optimal strategies in a risk selection investment model. Advances in Applied Probability, 32(2):518–539, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428202.

Aletti:2022:RPU

[AC22] Giacomo Aletti and Irene Crimaldi. The rescaled Pólya urn: local reinforcement and chi-squared goodness-of-fit test. Advances in Applied Probability, 54(3):849–879, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/rescaled-polya-urn-local-
reinforcement-and-chisquared-goodnessoffit-test/AD8BD0C92F5AAD335183BB4EEDEC501C.

Arias-Castro:2005:CDH

[ACDHT05] Ery Arias-Castro, David L. Donoho, Xiaoming Huo, and Craig A. Tovey. Connect the dots: How many random points can a regular curve pass through? Advances in Applied Probability, 37(3):571–603, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/30037345. See correction [HH03].

Andreoli:2012:SMF

[ACDP12] Alessandro Andreoli, Francesco Caravenna, Paolo Dai Pra, and Gustavo Posta. Scaling and multiscaling in financial series: a simple model. Advances in Applied Probability, 44(4):1018–1051, December 2012. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
41805989.

Armendariz:2009:NEB

[ACF09] Inés Armendáriz, Antonio Cuevas, and Ricardo Fraiman. Nonparametric estimation of boundary measures and related functionals: Asymptotic results. Advances in Applied Probability, 41 (2):311–322, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/27793879.

Acciaio:2021:MIP

[ACH21] Beatrice Acciaio, Alexander M. G. Cox, and Martin Huesmann. Model-independent pricing with insider information: a Skorokhod embedding approach. Advances in Applied Probability, 53(1):30–56, March 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/modelindependent-pricing-with-insider-
information-a-skorokhod-embedding-approach/EDD32B604CAF3272599F8F1CAF5C4059.

Antonelli:1977:GRDb

[ACLM77] Peter L. Antonelli, Jared Chapin, G. Mark Lathrop, and Kenneth Morgan. The geometry of random drift II. The symmetry of random genetic drift. Advances in Applied Probability, 9 (2):250–259, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426384.

Arnau:1996:CSSb

[ACO96a] X. Gual Arnau and L. M. Cruz-Orive. Consistency in systematic sampling. Advances in Applied Probability, 28(4):982–992, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428160.

Arnau:1996:CSSa

[ACO96b] X. Gual Arnau and L. M. Cruz-Orive. Consistency in systematic sampling for stereology. Advances in Applied Probability, 28 (2):329–330, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428037.

Arnau:1998:VPU

[ACO98] Ximo Gual Arnau and Luis M. Cruz-Orive. Variance prediction under systematic sampling with geometric probes. Advances in Applied Probability, 30(4):889–903, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428020.

Arias-Castro:2012:NGC

[ACPP12] Ery Arias-Castro, Bruno Pelletier, and Pierre Pudlo. The normalized graph cut and Cheeger constant: from discrete to continuous. Advances in Applied Probability, 44(4):907–937, December 2012. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
41805984.

Arak:1993:PBP

[ACS93] T. Arak, P. Clifford, and D. Surgailis. Point-based polygonal models for random graphs. Advances in Applied Probability, 25 (2):348–372, June 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427657.

Antonelli:1980:GRG

[ACV80] Peter L. Antonelli, Jared Chapin, and B. H. Voorhees. The geometry of random genetic drift VI. A random selection diffusion model. Advances in Applied Probability, 12(1):50–58, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426493.

Attia:1978:NIM

[AD78] F. A. Attia and D. Y. Downham. A numerical investigation of a model of Moran. Advances in Applied Probability, 10(1): 1–3, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426706.

Affentranger:1993:CHR

[AD93] Fernando Affentranger and Rex A. Dwyer. The convex Hull of random balls. Advances in Applied Probability, 25(2):373– 394, June 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427658.

Andersson:1994:FLT

[AD94] Håkan Andersson and Boualem Djehiche. A functional limit theorem for the total cost of a multitype standard epidemic. Advances in Applied Probability, 26(3):690–697, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427815.

Arratia:2015:RSP

[AD15] Richard Arratia and Stephen DeSalvo. On the random sampling of pairs, with pedestrian examples. Advances in Applied Probability, 47(1):292–305, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1427814592.

Abraham:2021:ESG

[AD21] Romain Abraham and Jean-François Delmas. Exact simulation of the genealogical tree for a stationary branching population and application to the asymptotics of its total length. Advances in Applied Probability, 53(2):537–574, June 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/exact-simulation-of-the-
genealogical-tree-for-a-stationary-branching-population-
and-application-to-the-asymptotics-of-its-total-length/
29CF7634AC1D54270B6D65CFE68074F5.

Adam:2016:CUG

[Ada16] Etienne Adam. Criterion for unlimited growth of critical multidimensional stochastic models. Advances in Applied Probability, 48(4):972–988, December 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1482548425.

Adler:1976:GNU

[Adl76] Robert J. Adler. On generalising the notion of upcrossings to random fields. Advances in Applied Probability, 8(4):789–805, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425934.

Adler:1977:GNU

[Adl77] Robert J. Adler. On generalising the notion of upcrossings to random fields. Advances in Applied Probability, 9(2):226– 227, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426370.

Adler:1979:SCB

[Adl79] Robert J. Adler. Spheres, cubes, and Brownian motions. Advances in Applied Probability, 11(2):289, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426819.

Adler:1989:FTS

[Adl89] Robert J. Adler. Fluctuation theory for systems of signed and unsigned particles with interaction mechanisms based on intersection local times. Advances in Applied Probability, 21(2):334– 356, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427163.

Ahmadi:2015:DMI

[ADL15] Jafar Ahmadi, Antonio Di Crescenzo, and Maria Longobardi. On dynamic mutual information for bivariate lifetimes. Advances in Applied Probability, 47(4):1157–1174, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1449859804.

Azais:2018:IEB

[ADP18] Romain Azaïs, Bernard Delyon, and François Portier. Integral estimation based on Markovian design. Advances in Applied Probability, 50(3):833–857, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/integral-estimation-based-
on-markovian-design/0BE2E561A0639ED694717FA9F16B29FA.

Appleton:1975:CSM

[ADW^{+}75] D. R. Appleton, H. E. Daniels, Tony Woods, Kevin Gough, John Steele, Rex Howell, Peter J. Diggle, Peter Clifford, Byron Morgan, Chris Preston, and Joel E. Cohen. Conference on stochastic models in space and time. Advances in Applied Probability, 7(3): 461–467, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426122.

Ahn:1999:OSS

[ADZ99] Hyun-Soo Ahn, Izak Duenyas, and Rachel Q. Zhang. Optimal stochastic scheduling of a two-stage tandem queue with parallel servers. Advances in Applied Probability, 31(4):1095–1117, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428346.

Ahn:2004:OCF

[ADZ04] Hyun-Soo Ahn, Izak Duenyas, and Rachel Q. Zhang. Optimal control of a flexible server. Advances in Applied Probability, 36 (1):139–170, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428357.

Aggoun:1995:FDM

[AE95] Lakhdar Aggoun and Robert J. Elliott. Finite-dimensional models for hidden Markov chains. Advances in Applied Probability, 27(1):146–160, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428101.

Alos:2008:MDH

[AE08] Elisa Alòs and Christian-Oliver Ewald. Malliavin differentiability of the Heston volatility and applications to option pricing. Advances in Applied Probability, 40(1):144–162, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/20443574.

Aebi:1994:SDA

[AEM94] M. Aebi, P. Embrechts, and T. Mikosch. Stochastic discounting, aggregate claims, and the bootstrap. Advances in Applied Probability, 26(1):183–206, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427586.

Asmussen:2018:RVF

[AF18] Søren Asmussen and Sergey Foss. Regular variation in a fixed-point problem for single- and multi-class branching processes and queues. Advances in Applied Probability, 50(A):47–61, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/regular-variation-in-a-fixedpoint-
problem-for-single-and-multiclass-branching-processes-
and-queues/175DB7BE85611B94E042CA32963EC59C.

Antunes:2006:PAV

[AFGR06] Nelson Antunes, Christine Fricker, Fabrice Guillemin, and Philippe Robert. Perturbation analysis of a variable M/M/1 queue: A probabilistic approach. Advances in Applied Probability, 38(1):263–283, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443438.

Axelson-Fisk:2009:CPO

[AFH09] Marina Axelson-Fisk and Olle Häggström. Conditional percolation on one-dimensional lattices. Advances in Applied Probability, 41(4):1102–1122, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/27793921.

Ayala:1991:RSC

[AFM91] Guillermo Ayala, Juan Ferrandiz, and Francisco Montes. Random set and coverage measure. Advances in Applied Probability, 23(4):972–974, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427687.

Aczel:1974:WSH

[AFN74] J. Aczél, B. Forte, and C. T. Ng. Why the Shannon and Hartley entropies are ‘natural’. Advances in Applied Probability, 6(1):

131–146, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426210.

Arjas:1981:CRI

[AG81] Elja Arjas and Priscilla Greenwood. Competing risks and independent minima: A marked point process approach. Advances in Applied Probability, 13(4):669–680, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426966.

Araman:2005:DAM

[AG05] Victor F. Araman and Peter W. Glynn. Diffusion approximations for the maximum of a perturbed random walk. Advances in Applied Probability, 37(3):663–680, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037349.

Alsmeyer:2013:HPM

[AG13] Gerold Alsmeyer and Sören Gröttrup. A host-parasite model for a two-type cell population. Advances in Applied Probability, 45(3):719–741, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1377868536.

Agresti:1974:BETa

[Agr74a] A. Agresti. Bounds on the extinction time distributions of some branching processes. Advances in Applied Probability, 6(2):228– 229, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426266.

Agresti:1974:BETb

[Agr74b] Alan Agresti. Bounds on the extinction time distribution of a branching process. Advances in Applied Probability, 6(2):322– 335, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426296.

Agrawal:1995:SMB

[Agr95] Rajeev Agrawal. Sample mean based index policies with o(log n) regret for the multi-armed bandit problem. Advances in Applied Probability, 27(4):1054–1078, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427934.

Assaf:2004:TCO

[AGSC04] David Assaf, Larry Goldstein, and Ester Samuel-Cahn. Twochoice optimal stopping. Advances in Applied Probability, 36 (4):1116–1147, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/4140391.

Avery:1978:ELD

[AH78] P. J. Avery and W. G. Hill. The effect of linkage disequilibrium on the genetic variance of a quantitative trait. Advances in Applied Probability, 10(1):4–6, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426707.

Abdel-Hameed:1987:IMP

[AH87] Mohamed Abdel-Hameed. Inspection and maintenance policies of devices subject to deterioration. Advances in Applied Probability, 19(4):917–931, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427108. See corrections [Ano95e].

Altman:1997:ABR

[AH97] Eitan Altman and Arie Hordijk. Applications of Borovkov’s renovation theory to non-stationary stochastic recursive sequences and their control. Advances in Applied Probability, 29(2):388– 413, June 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428009.

Aastrom:1999:SAI

[ÅH99] Kalle Åström and Anders Heyden. Stochastic analysis of image acquisition, interpolation and scale-space smoothing. Advances in Applied Probability, 31(4):855–894, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428332.

Alvarez:2005:QST

[AH05] Juan Alvarez and Bruce Hajek. A queue with semiperiodic traffic. Advances in Applied Probability, 37(1):160–184, March 2005.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/30037320.

Acan:2016:MGP

[AH16] Hüseyin Acan and Pawel Hitczenko. On a memory game and preferential attachment graphs. Advances in Applied Probability, 48(2):585–609, June 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1465490764.

Ahlberg:2015:AFP

[Ahl15] Daniel Ahlberg. Asymptotics of first-passage percolation on one-dimensional graphs. Advances in Applied Probability, 47 (1):182–209, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1427814587.

Ahmad:1980:MFT

[Ahm80a] R. Ahmad. Markov functionals and their algebraic structure. Advances in Applied Probability, 12(2):290–291, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426563.

Ahmed:1980:SCH

[Ahm80b] N. U. Ahmed. Stochastic control on Hilbert space for linear evolution equations with random operator-valued coefficients. Advances in Applied Probability, 12(2):288, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426560.

Ahmad:1981:EON

[Ahm81] Ibrahim A. Ahmad. The exact order of normal approximation in bivariate renewal theory. Advances in Applied Probability, 13 (1):113–128, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426470.

Andradottir:1996:PUV

[AHO96] Sigrún Andradóttir, Daniel P. Heyman, and Teunis J. Ott. Potentially unlimited variance reduction in importance sampling of Markov chains. Advances in Applied Probability, 28(1):166– 188, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427916.

Arab:2020:FRP

[AHO20] Idir Arab, Milto Hadjikyriakou, and Paulo Eduardo Oliveira. Failure rate properties of parallel systems. Advances in Applied Probability, 52(2):563–587, June 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/failure-rate-properties-
of-parallel-systems/CD8AC5C74F3780E8EDBB8B62275225C9.

Avi-Itzhak:2004:MFQ

[AIL04] Benjamin Avi-Itzhak and Hanoch Levy. On measuring fairness in queues. Advances in Applied Probability, 36(3):919–936, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
4140415.

Aisbett:1979:RBP

[Ais79] C. W. Aisbett. The relation between pointwise and uniform convergence of the empirical distributions of identically distributed real data. Advances in Applied Probability, 11(2):280–281, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426811.

Avram:2013:LSS

[AJV13] F. Avram, A. J. E. M. Janssen, and J. S. H. Van Leeuwaarden. Loss systems with slow retrials in the Halfin–Whitt regime. Advances in Applied Probability, 45(1):274–294, March 2013. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/41806008.

Asmussen:1978:MCL

[AK78] Søren Asmussen and Niels Keiding. Martingale central limit theorems and asymptotic estimation theory for multitype branching processes. Advances in Applied Probability, 10(1):109–129, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426721.

Anderson:1979:FBM

[AK79] B. D. O. Anderson and T. Kailath. Forwards and backwards models for finite-state Markov processes. Advances in Applied Probability, 11(1):118–133, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426771.

Aly:1993:HRO

[AK93] Emad-Eldin A. A. Aly and Subhash C. Kochar. On hazard rate ordering of dependent variables. Advances in Applied Probability, 25(2):477–482, June 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427664.

Anantharam:1995:FCL

[AK95] Venkat Anantharam and Takis Konstantopoulos. A functional central limit theorem for the jump counts of Markov processes with an application to Jackson networks. Advances in Applied Probability, 27(2):476–509, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427836.

Athreya:1998:SLTa

[AK98a] Krishna B. Athreya and Hye-Jeong Kang. Some limit theorems for positive recurrent branching Markov chains: I. Advances in Applied Probability, 30(3):693–710, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428146.

Athreya:1998:SLTb

[AK98b] Krishna B. Athreya and Hye-Jeong Kang. Some limit theorems for positive recurrent branching Markov chains: II. Advances in Applied Probability, 30(3):711–722, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428147.

Asmussen:2000:MDM

[AK00] Søren Asmussen and Offer Kella. A multi-dimensional martingale for Markov additive processes and its applications. Advances in Applied Probability, 32(2):376–393, June 2000. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428194.

Albeverio:2005:LPW

[AK05] Sergio Albeverio and Shuji Kawasaki. On a localization property of wavelet coefficients for processes with stationary increments, and applications. I. Localization with respect to shift. Advances in Applied Probability, 37(4):938–962, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037368.

Asmussen:2006:IAR

[AK06] Søren Asmussen and Dirk P. Kroese. Improved algorithms for rare event simulation with heavy tails. Advances in Applied Probability, 38(2):545–558, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443455.

Aldous:2008:SLR

[AK08] David J. Aldous and Wilfrid S. Kendall. Short-length routes in low-cost networks VIA Poisson line patterns. Advances in Applied Probability, 40(1):1–21, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443567.

Aksland:1975:BDM

[Aks75] M. Aksland. A birth, death and migration process with immigration. Advances in Applied Probability, 7(1):44–60, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425852.

Altman:1994:QS

[AL94] Eitan Altman and Hanoch Levy. Queueing in space. Advances in Applied Probability, 26(4):1095–1116, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427906.

Avrachenkov:1999:FMS

[AL99] Konstantin E. Avrachenkov and Jean B. Lasserre. The fundamental matrix of singularly perturbed Markov chains. Advances in Applied Probability, 31(3):679–697, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428391.

Alanyali:1999:TMP

[Ala99] Murat Alanyali. Three management policies for a resource with partition constraints. Advances in Applied Probability, 31(3): 795–818, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428397.

Albrecht:1983:NLB

[Alb83] Peter Albrecht. A note on a limiting behaviour of the occurrence times of a mixed Poisson process. Advances in Applied Probability, 15(2):460, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426445.

Aldous:1980:NLW

[Ald80] D. J. Aldous. A new look at weak convergence. Advances in Applied Probability, 12(2):266–267, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426532.

Alexander:2013:CDW

[Ale13] H. K. Alexander. Conditional distributions and waiting times in multitype branching processes. Advances in Applied Probability, 45(3):692–718, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1377868535.

Alfa:2003:AAB

[Alf03] Attahiru Sule Alfa. Algorithmic analysis of the BMAP/D/k system in discrete time. Advances in Applied Probability, 35(4): 1131–1152, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428503.

Ali:1990:END

[Ali90] Hydar Ali. Expected number of departures in M/M/1 and GI/G/1 queues. Advances in Applied Probability, 22(3):770–772, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427474.

Allen:1982:MDS

[All82] P. M. Allen. Modelling dynamic self-organization in complex systems. Advances in Applied Probability, 14(2):195–196, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426506.

Alm:1997:DSS

[Alm97] Sven Erick Alm. On the distributions of scan statistics of a twodimensional Poisson process. Advances in Applied Probability, 29(1):1–18, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427857.

Alparslan:2009:EPB

[Alp09] Ugur Tuncay Alparslan. Exceedance of power barriers for integrated continuous-time stationary ergodic stable processes. Advances in Applied Probability, 41(3):874–892, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/27793906.

Alsmeyer:1988:SOA

[Als88] Gerold Alsmeyer. Second-order approximations for certain stopped sums in extended renewal theory. Advances in Applied Probability, 20(2):391–410, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427396.

Anh:2008:MPG

[ALS08] V. V. Anh, Nikolai N. Leonenko, and Narn-Rueih Shieh. Multifractality of products of geometric Ornstein–Uhlenbeck-type processes. Advances in Applied Probability, 40(4):1129–1156, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443627.

Achaz:2018:SLA

[ALS18a] Guillaume Achaz, Amaury Lambert, and Emmanuel Schertzer. The sequential loss of allelic diversity. Advances in Applied Probability, 50(A):13–29, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/sequential-loss-of-allelic-diversity/
AA3C0194A048E62FF9EEBAB8571C3D21.

Alsmeyer:2018:LEL

[Als18b] Gerold Alsmeyer. Ladder epochs and ladder chain of a Markov random walk with discrete driving chain. Advances in Applied Probability, 50(A):31–46, December 2018. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/ladder-epochs-and-ladder-chain-of-
a-markov-random-walk-with-discrete-driving-chain/3954040C1CBCD963391FDAAE68E162E0.

Anderson:1990:RST

[AM90] W. J. Anderson and P. M. McDunnough. On the representation of symmetric transition functions. Advances in Applied Probability, 22(3):548–563, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427457.

Allaart:2001:OSR

[AM01] Pieter Allaart and Michael Monticino. Optimal stopping rules for directionally reinforced processes. Advances in Applied Probability, 33(2):483–504, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428264.

Avram:2017:CMR

[AM17] Florin Avram and Andreea Minca. On the central management of risk networks. Advances in Applied Probability, 49(1):221–237, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-the-central-management-of-risk-
networks/7CB3497E38D71CA374570BE951E478A1.

Akbarzadeh:2022:CIR

[AM22] Nima Akbarzadeh and Aditya Mahajan. Conditions for indexability of restless bandits and an O(K^{∋}) algorithm to compute Whittle index. Advances in Applied Probability, 54 (4):1164–1192, December 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/conditions-for-indexability-of-restless-
bandits-and-an-mathcaloleftk3right-algorithm-to-compute-
whittle-index/05749A3BD36DAC46166F96F9BEF8B9D7.

Ambartzumian:1977:SGSa

[Amb77a] R. V. Ambartzumian. Stochastic geometry from the standpoint of integral geometry. Advances in Applied Probability, 9(3):425– 426, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426088.

Ambartzumian:1977:SGSb

[Amb77b] R. V. Ambartzumian. Stochastic geometry from the standpoint of integral geometry. Advances in Applied Probability, 9(4): 792–823, December 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426701.

Antonelli:1977:GRDc

[AML77] Peter L. Antonelli, Kenneth Morgan, and G. Mark Lathrop. The geometry of random drift III. Recombination and diffusion. Advances in Applied Probability, 9(2):260–267, June 1977. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426385.

AlHanbali:2011:AVD

[AMNW11] A. Al Hanbali, M. Mandjes, Y. Nazarathy, and W. Whitt. The asymptotic variance of departures in critically loaded queues. Advances in Applied Probability, 43(1):243–263, March 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/23024534.

Aletti:2007:DLP

[AMS07] Giacomo Aletti, Caterina May, and Piercesare Secchi. On the distribution of the limit proportion for a two-color, randomly reinforced urn with equal reinforcement distributions. Advances in Applied Probability, 39(3):690–707, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443536.

Aletti:2009:CLT

[AMS09] Giacomo Aletti, Caterina May, and Piercesare Secchi. A central limit theorem, and related results, for a two-color randomly reinforced urn. Advances in Applied Probability, 41(3):829–844, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
27793904.

Athreya:1978:LTS

[AN78a] K. B. Athreya and P. Ney. Limit theorems for semi-Markov processes and renewal theory for Markov chains on general state spaces. Advances in Applied Probability, 10(2):292–293, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426896.

Athreya:1978:RIR

[AN78b] K. B. Athreya and P. Ney. Regeneration and imbeddings for recurrent Markov chains on general state spaces. Advances in Applied Probability, 10(2):292, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426895.

Andersen:1974:SFT

[And74] E. Sparre Andersen. Survey of fluctuation theory. Advances in Applied Probability, 6(2):213–214, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426252.

Andre:1975:TPS

[And75] C. André. Two-person stopping games. Advances in Applied Probability, 7(2):244–246, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426057.

Anderson:1979:PBK

[And79] B. D. O. Anderson. A physical basis for Krein’s prediction formulas. Advances in Applied Probability, 11(2):297–298, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426828.

Andersson:1999:AFS

[And99] Mikael Andersson. The asymptotic final size distribution of multitype chain-binomial epidemic processes. Advances in Applied Probability, 31(1):220–234, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427999.

Anonymous:1969:BMa

[Ano69a] Anonymous. Back matter. Advances in Applied Probability, 1(1):??, Spring 1969. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426412.

Anonymous:1969:BMb

[Ano69b] Anonymous. Back matter. Advances in Applied Probability, 1(2):??, Autumn 1969. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426221.

Anonymous:1969:FMa

[Ano69c] Anonymous. Front matter. Advances in Applied Probability, 1(1):??, Spring 1969. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426406.

Anonymous:1969:FMb

[Ano69d] Anonymous. Front matter. Advances in Applied Probability, 1(2):??, Autumn 1969. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426215.

Anonymous:1970:BMa

[Ano70a] Anonymous. Back matter. Advances in Applied Probability, 2(1):??, Spring 1970. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/3518348.

Anonymous:1970:BMb

[Ano70b] Anonymous. Back matter. Advances in Applied Probability, 2(2):??, Autumn 1970. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426326.

Anonymous:1970:FMa

[Ano70c] Anonymous. Front matter. Advances in Applied Probability, 2(1):??, Spring 1970. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/3518343.

Anonymous:1970:FMb

[Ano70d] Anonymous. Front matter. Advances in Applied Probability, 2(2):??, Autumn 1970. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426317.

Anonymous:1971:BMa

[Ano71a] Anonymous. Back matter. Advances in Applied Probability, 3(1):??, Spring 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426334.

Anonymous:1971:BMb

[Ano71b] Anonymous. Back matter. Advances in Applied Probability, 3(2):??, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426179.

Anonymous:1971:FMa

[Ano71c] Anonymous. Front matter. Advances in Applied Probability, 3(1):??, Spring 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426327.

Anonymous:1971:FMb

[Ano71d] Anonymous. Front matter. Advances in Applied Probability, 3(2):??, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426150.

Anonymous:1971:I

[Ano71e] Anonymous. [Introduction]. Advances in Applied Probability, 3(2):193, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426151.

Anonymous:1971:VI

[Ano71f] Anonymous. Volume information. Advances in Applied Probability, 3(2):??, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426149.

Anonymous:1972:BMa

[Ano72a] Anonymous. Back matter. Advances in Applied Probability, 4 (1):??, April 1972. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425811.

Anonymous:1972:BMb

[Ano72b] Anonymous. Back matter. Advances in Applied Probability, 4(2):??, August 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426006.

Anonymous:1972:BMc

[Ano72c] Anonymous. Back matter. Advances in Applied Probability, 4 (3):??, December 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425994.

Anonymous:1972:FMa

[Ano72d] Anonymous. Front matter. Advances in Applied Probability, 4 (1):??, April 1972. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425804.

Anonymous:1972:FMb

[Ano72e] Anonymous. Front matter. Advances in Applied Probability, 4(2):??, August 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425995.

Anonymous:1972:FMc

[Ano72f] Anonymous. Front matter. Advances in Applied Probability, 4 (3):??, December 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425987.

Anonymous:1972:VI

[Ano72g] Anonymous. Volume information. Advances in Applied Probability, 4(3):??, December 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1425986.

Anonymous:1973:BMa

[Ano73a] Anonymous. Back matter. Advances in Applied Probability, 5 (1):??, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425971.

Anonymous:1973:BMb

[Ano73b] Anonymous. Back matter. Advances in Applied Probability, 5(2):??, August 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426043.

Anonymous:1973:BMc

[Ano73c] Anonymous. Back matter. Advances in Applied Probability, 5 (3):??, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425838.

Anonymous:1973:FMa

[Ano73d] Anonymous. Front matter. Advances in Applied Probability, 5 (1):??, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425939.

Anonymous:1973:FMb

[Ano73e] Anonymous. Front matter. Advances in Applied Probability, 5(2):??, August 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426031.

Anonymous:1973:FMc

[Ano73f] Anonymous. Front matter. Advances in Applied Probability, 5 (3):??, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425826.

Anonymous:1973:I

[Ano73g] Anonymous. [Introduction]. Advances in Applied Probability, 5 (1):1, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425940.

Anonymous:1973:VI

[Ano73h] Anonymous. Volume information. Advances in Applied Probability, 5(3):632–633, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425825.

Anonymous:1974:BMa

[Ano74a] Anonymous. Back matter. Advances in Applied Probability, 6 (1):??, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426214.

Anonymous:1974:BMb

[Ano74b] Anonymous. Back matter. Advances in Applied Probability, 6 (2):??, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426301.

Anonymous:1974:BMc

[Ano74c] Anonymous. Back matter. Advances in Applied Probability, 6 (3):??, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426237.

Anonymous:1974:BMd

[Ano74d] Anonymous. Back matter. Advances in Applied Probability, 6 (4):??, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426193.

Anonymous:1974:FMa

[Ano74e] Anonymous. Front matter. Advances in Applied Probability, 6 (1):??, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426194.

Anonymous:1974:FMb

[Ano74f] Anonymous. Front matter. Advances in Applied Probability, 6 (2):185–186, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426238.

Anonymous:1974:FMc

[Ano74g] Anonymous. Front matter. Advances in Applied Probability, 6 (3):??, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426222.

Anonymous:1974:FMd

[Ano74h] Anonymous. Front matter. Advances in Applied Probability, 6 (4):??, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426181.

Anonymous:1974:Ia

[Ano74i] Anonymous. Introduction. Advances in Applied Probability, 6 (2):187, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426239.

Anonymous:1974:Ib

[Ano74j] Anonymous. [Introduction]. Advances in Applied Probability, 6 (3):407, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426223.

Anonymous:1974:VI

[Ano74k] Anonymous. Volume information. Advances in Applied Probability, 6(4):784–786, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426180.

Anonymous:1975:BMa

[Ano75a] Anonymous. Back matter. Advances in Applied Probability, 7 (1):??, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425862.

Anonymous:1975:BMb

[Ano75b] Anonymous. Back matter. Advances in Applied Probability, 7 (2):??, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426085.

Anonymous:1975:BMc

[Ano75c] Anonymous. Back matter. Advances in Applied Probability, 7 (3):??, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426135.

Anonymous:1975:BMd

[Ano75d] Anonymous. Back matter. Advances in Applied Probability, 7 (4):??, December 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426405.

Anonymous:1975:FMa

[Ano75e] Anonymous. Front matter. Advances in Applied Probability, 7 (1):??, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425839.

Anonymous:1975:FMb

[Ano75f] Anonymous. Front matter. Advances in Applied Probability, 7 (2):227–228, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426044.

Anonymous:1975:FMc

[Ano75g] Anonymous. Front matter. Advances in Applied Probability, 7 (3):??, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426120.

Anonymous:1975:FMd

[Ano75h] Anonymous. Front matter. Advances in Applied Probability, 7 (4):??, December 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426395.

Anonymous:1975:I

[Ano75i] Anonymous. Introduction. Advances in Applied Probability, 7 (2):229, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426045.

Anonymous:1975:VI

[Ano75j] Anonymous. Volume information. Advances in Applied Probability, 7(4):878–880, December 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426394.

Anonymous:1976:BMa

[Ano76a] Anonymous. Back matter. Advances in Applied Probability, 8 (1):??, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426030.

Anonymous:1976:BMb

[Ano76b] Anonymous. Back matter. Advances in Applied Probability, 8 (2):??, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425913.

Anonymous:1976:BMc

[Ano76c] Anonymous. Back matter. Advances in Applied Probability, 8 (3):??, September 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426148.

Anonymous:1976:BMd

[Ano76d] Anonymous. Back matter. Advances in Applied Probability, 8 (4):??, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425938.

Anonymous:1976:CPR

[Ano76e] Anonymous. Corrections to Prorok and Rosenlund. Advances in Applied Probability, 8(3):610–611, September 1976. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426147. See

[Pro76a, Ros76a].

Anonymous:1976:FMa

[Ano76f] Anonymous. Front matter. Advances in Applied Probability, 8 (1):??, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426007.

Anonymous:1976:FMb

[Ano76g] Anonymous. Front matter. Advances in Applied Probability, 8 (2):209–210, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425863.

Anonymous:1976:FMc

[Ano76h] Anonymous. Front matter. Advances in Applied Probability, 8 (3):??, September 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426136.

Anonymous:1976:FMd

[Ano76i] Anonymous. Front matter. Advances in Applied Probability, 8 (4):??, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425915.

Anonymous:1976:I

[Ano76j] Anonymous. Introduction. Advances in Applied Probability, 8 (2):211–212, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425864.

Anonymous:1976:SST

[Ano76k] Anonymous. Symposium on statistical techniques. Advances in Applied Probability, 8(2):254–256, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425903.

Anonymous:1976:VI

[Ano76l] Anonymous. Volume information. Advances in Applied Probability, 8(4):847–850, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425914.

Anonymous:1977:BMa

[Ano77a] Anonymous. Back matter. Advances in Applied Probability, 9 (1):??, March 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425824.

Anonymous:1977:BMb

[Ano77b] Anonymous. Back matter. Advances in Applied Probability, 9 (2):??, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426393.

Anonymous:1977:BMc

[Ano77c] Anonymous. Back matter. Advances in Applied Probability, 9 (3):??, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426119.

Anonymous:1977:BMd

[Ano77d] Anonymous. Back matter. Advances in Applied Probability, 9 (4):??, December 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426704.

Anonymous:1977:CPP

[Ano77e] Anonymous. Correction: Point processes and random measures. Advances in Applied Probability, 9(4):861, December 1977. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426703. See [Gra77].

Anonymous:1977:FMa

[Ano77f] Anonymous. Front matter. Advances in Applied Probability, 9 (1):??, March 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425812.

Anonymous:1977:FMb

[Ano77g] Anonymous. Front matter. Advances in Applied Probability, 9 (2):187–188, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426335.

Anonymous:1977:FMc

[Ano77h] Anonymous. Front matter. Advances in Applied Probability, 9 (3):423, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426086.

Anonymous:1977:FMd

[Ano77i] Anonymous. Front matter. Advances in Applied Probability, 9 (4):??, December 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426695.

Anonymous:1977:I

[Ano77j] Anonymous. Introduction. Advances in Applied Probability, 9 (2):189, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426336.

Anonymous:1977:VI

[Ano77k] Anonymous. Volume information. Advances in Applied Probability, 9(4):862–866, December 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426694.

Anonymous:1978:BMa

[Ano78a] Anonymous. Back matter. Advances in Applied Probability, 10 (1):??, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426728.

Anonymous:1978:BMb

[Ano78b] Anonymous. Back matter. Advances in Applied Probability, 10 (2):??, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426947.

Anonymous:1978:BMc

[Ano78c] Anonymous. Back matter. Advances in Applied Probability, 10 (3):??, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426643.

Anonymous:1978:BMd

[Ano78d] Anonymous. Back matter. Advances in Applied Probability, 10 (4):??, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426667.

Anonymous:1978:CCS

[Ano78e] Anonymous. Corrections to Cohen and Stuart. Advances in Applied Probability, 10(3):702, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426642. See [Coh77c,

Stu78].

Anonymous:1978:FMa

[Ano78f] Anonymous. Front matter. Advances in Applied Probability, 10 (1):??, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426705.

Anonymous:1978:FMb

[Ano78g] Anonymous. Front matter. Advances in Applied Probability, 10 (2):255–258, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426868.

Anonymous:1978:FMc

[Ano78h] Anonymous. Front matter. Advances in Applied Probability, 10 (3):??, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426630.

Anonymous:1978:FMd

[Ano78i] Anonymous. Front matter. Advances in Applied Probability, 10 (4):??, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426645.

Anonymous:1978:I

[Ano78j] Anonymous. Introduction. Advances in Applied Probability, 10 (2):259, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426869.

Anonymous:1978:VI

[Ano78k] Anonymous. Volume information. Advances in Applied Probability, 10(4):913–918, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426644.

Anonymous:1979:BMa

[Ano79a] Anonymous. Back matter. Advances in Applied Probability, 11 (1):??, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426777.

Anonymous:1979:BMb

[Ano79b] Anonymous. Back matter. Advances in Applied Probability, 11 (2):??, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426851.

Anonymous:1979:BMc

[Ano79c] Anonymous. Back matter. Advances in Applied Probability, 11 (3):??, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426961.

Anonymous:1979:BMd

[Ano79d] Anonymous. Back matter. Advances in Applied Probability, 11 (4):??, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426867.

Anonymous:1979:CTR

[Ano79e] Anonymous. Correction: Transience and recurrence of statedependent branching processes with an immigration component. Advances in Applied Probability, 11(4):889, December 1979. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426866. See [Lev79].

Anonymous:1979:CMS

[Ano79f] Anonymous. Corrections to Melamed and Schweitzer. Advances in Applied Probability, 11(2):456, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426850. See [Mel79b, SF79].

Anonymous:1979:FMa

[Ano79g] Anonymous. Front matter. Advances in Applied Probability, 11 (1):??, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426755.

Anonymous:1979:FMb

[Ano79h] Anonymous. Front matter. Advances in Applied Probability, 11 (2):257–259, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426790.

Anonymous:1979:FMc

[Ano79i] Anonymous. Front matter. Advances in Applied Probability, 11 (3):??, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426948.

Anonymous:1979:FMd

[Ano79j] Anonymous. Front matter. Advances in Applied Probability, 11 (4):??, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426853.

Anonymous:1979:VI

[Ano79k] Anonymous. Volume information. Advances in Applied Probability, 11(4):890–893, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426852.

Anonymous:1980:BMa

[Ano80a] Anonymous. Back matter. Advances in Applied Probability, 12 (1):??, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426504.

Anonymous:1980:BMb

[Ano80b] Anonymous. Back matter. Advances in Applied Probability, 12 (2):??, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426613.

Anonymous:1980:BMc

[Ano80c] Anonymous. Back matter. Advances in Applied Probability, 12 (3):??, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426433.

Anonymous:1980:BMd

[Ano80d] Anonymous. Back matter. Advances in Applied Probability, 12 (4):??, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426754.

Anonymous:1980:CIS

[Ano80e] Anonymous. Correction: The insensitivity of stationary probabilities in networks of queues. Advances in Applied Probability, 12(2):541, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426612. See [Sch78a].

Anonymous:1980:FMa

[Ano80f] Anonymous. Front matter. Advances in Applied Probability, 12 (1):??, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426478.

Anonymous:1980:FMb

[Ano80g] Anonymous. Front matter. Advances in Applied Probability, 12 (2):263–265, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426531.

Anonymous:1980:FMc

[Ano80h] Anonymous. Front matter. Advances in Applied Probability, 12 (3):543–544, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426413.

Anonymous:1980:FMd

[Ano80i] Anonymous. Front matter. Advances in Applied Probability, 12 (4):??, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426742.

Anonymous:1980:VI

[Ano80j] Anonymous. Volume information. Advances in Applied Probability, 12(4):1019–1026, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426741.

Anonymous:1981:BMa

[Ano81a] Anonymous. Back matter. Advances in Applied Probability, 13 (1):??, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426477.

Anonymous:1981:BMb

[Ano81b] Anonymous. Back matter. Advances in Applied Probability, 13 (2):??, June 1981. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426693.

Anonymous:1981:BMc

[Ano81c] Anonymous. Back matter. Advances in Applied Probability, 13 (3):??, September 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426789.

Anonymous:1981:BMd

[Ano81d] Anonymous. Back matter. Advances in Applied Probability, 13 (4):??, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426977.

Anonymous:1981:FMa

[Ano81e] Anonymous. Front matter. Advances in Applied Probability, 13 (1):??, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426450.

Anonymous:1981:FMb

[Ano81f] Anonymous. Front matter. Advances in Applied Probability, 13 (2):??, June 1981. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426682.

Anonymous:1981:FMc

[Ano81g] Anonymous. Front matter. Advances in Applied Probability, 13 (3):??, September 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426778.

Anonymous:1981:FMd

[Ano81h] Anonymous. Front matter. Advances in Applied Probability, 13 (4):??, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426963.

Anonymous:1981:VI

[Ano81i] Anonymous. Volume information. Advances in Applied Probability, 13(4):860–862, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426962.

Anonymous:1982:BMa

[Ano82a] Anonymous. Back matter. Advances in Applied Probability, 14 (1):??, March 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426740.

Anonymous:1982:BMb

[Ano82b] Anonymous. Back matter. Advances in Applied Probability, 14 (2):??, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426530.

Anonymous:1982:BMc

[Ano82c] Anonymous. Back matter. Advances in Applied Probability, 14 (3):??, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426681.

Anonymous:1982:BMd

[Ano82d] Anonymous. Back matter. Advances in Applied Probability, 14 (4):??, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427031.

Anonymous:1982:FMa

[Ano82e] Anonymous. Front matter. Advances in Applied Probability, 14 (1):??, March 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426729.

Anonymous:1982:FMb

[Ano82f] Anonymous. Front matter. Advances in Applied Probability, 14 (2):??, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426505.

Anonymous:1982:FMc

[Ano82g] Anonymous. Front matter. Advances in Applied Probability, 14 (3):??, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426668.

Anonymous:1982:FMd

[Ano82h] Anonymous. Front matter. Advances in Applied Probability, 14 (4):??, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427018.

Anonymous:1982:VI

[Ano82i] Anonymous. Volume information. Advances in Applied Probability, 14(4):912–914, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427017.

Anonymous:1983:BMa

[Ano83a] Anonymous. Back matter. Advances in Applied Probability, 15 (1):??, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426998.

Anonymous:1983:BMb

[Ano83b] Anonymous. Back matter. Advances in Applied Probability, 15 (2):??, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426449.

Anonymous:1983:BMc

[Ano83c] Anonymous. Back matter. Advances in Applied Probability, 15 (3):??, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426629.

Anonymous:1983:BMd

[Ano83d] Anonymous. Back matter. Advances in Applied Probability, 15 (4):??, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427334.

Anonymous:1983:CLC

[Ano83e] Anonymous. Correction: The lifetimes of configuration states in statistical physics. Advances in Applied Probability, 15(1): 223, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1426997. See [Rub81].

Anonymous:1983:CTS

[Ano83f] Anonymous. Correction: Two suggestions of how to define a multistate coherent system. Advances in Applied Probability, 15(1):223, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426996. See [Nat82].

Anonymous:1983:FMa

[Ano83g] Anonymous. Front matter. Advances in Applied Probability, 15 (1):??, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426978.

Anonymous:1983:FMb

[Ano83h] Anonymous. Front matter. Advances in Applied Probability, 15 (2):??, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426434.

Anonymous:1983:FMc

[Ano83i] Anonymous. Front matter. Advances in Applied Probability, 15 (3):??, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426614.

Anonymous:1983:FMd

[Ano83j] Anonymous. Front matter. Advances in Applied Probability, 15 (4):??, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427318.

Anonymous:1983:VI

[Ano83k] Anonymous. Volume information. Advances in Applied Probability, 15(4):894–896, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427317.

Anonymous:1984:BMa

[Ano84a] Anonymous. Back matter. Advances in Applied Probability, 16 (1):??, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427235.

Anonymous:1984:BMb

[Ano84b] Anonymous. Back matter. Advances in Applied Probability, 16 (2):??, June 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427080.

Anonymous:1984:BMc

[Ano84c] Anonymous. Back matter. Advances in Applied Probability, 16 (3):??, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427299.

Anonymous:1984:BMd

[Ano84d] Anonymous. Back matter. Advances in Applied Probability, 16 (4):??, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427352.

Anonymous:1984:CSS

[Ano84e] Anonymous. Correction: Statistical spatial series modelling. Advances in Applied Probability, 16(1):220, March 1984. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427234. See [Tjø78].

Anonymous:1984:FMa

[Ano84f] Anonymous. Front matter. Advances in Applied Probability, 16(1):1–2, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427214.

Anonymous:1984:FMb

[Ano84g] Anonymous. Front matter. Advances in Applied Probability, 16 (2):??, June 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427067.

Anonymous:1984:FMc

[Ano84h] Anonymous. Front matter. Advances in Applied Probability, 16 (3):??, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427283.

Anonymous:1984:FMd

[Ano84i] Anonymous. Front matter. Advances in Applied Probability, 16 (4):??, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427336.

Anonymous:1984:I

[Ano84j] Anonymous. Introduction. Advances in Applied Probability, 16(1):3–4, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427215.

Anonymous:1984:VI

[Ano84k] Anonymous. Volume information. Advances in Applied Probability, 16(4):936–940, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427335.

Anonymous:1985:BMa

[Ano85a] Anonymous. Back matter. Advances in Applied Probability, 17 (1):??, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427066.

Anonymous:1985:BMb

[Ano85b] Anonymous. Back matter. Advances in Applied Probability, 17 (2):??, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427157.

Anonymous:1985:BMc

[Ano85c] Anonymous. Back matter. Advances in Applied Probability, 17 (3):??, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427127.

Anonymous:1985:BMd

[Ano85d] Anonymous. Back matter. Advances in Applied Probability, 17 (4):??, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427097.

Anonymous:1985:CMR

[Ano85e] Anonymous. Correction: The maximum of a random walk whose mean path has a maximum. Advances in Applied Probability, 17(2):475, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427156. See [DS85].

Anonymous:1985:FMa

[Ano85f] Anonymous. Front matter. Advances in Applied Probability, 17 (1):??, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427048.

Anonymous:1985:FMb

[Ano85g] Anonymous. Front matter. Advances in Applied Probability, 17 (2):??, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427128.

Anonymous:1985:FMc

[Ano85h] Anonymous. Front matter. Advances in Applied Probability, 17 (3):??, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427115.

Anonymous:1985:FMd

[Ano85i] Anonymous. Front matter. Advances in Applied Probability, 17 (4):??, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427082.

Anonymous:1985:VI

[Ano85j] Anonymous. Volume information. Advances in Applied Probability, 17(4):915–917, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427081.

Anonymous:1986:BMa

[Ano86a] Anonymous. Back matter. Advances in Applied Probability, 18(1):287, March 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427250.

Anonymous:1986:BMb

[Ano86b] Anonymous. Back matter. Advances in Applied Probability, 18 (2):??, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427316.

Anonymous:1986:BMc

[Ano86c] Anonymous. Back matter. Advances in Applied Probability, 18 (3):??, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427193.

Anonymous:1986:BMd

[Ano86d] Anonymous. Back matter. Advances in Applied Probability, 18 (4):??, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427264.

Anonymous:1986:CAP

[Ano86e] Anonymous. Correction: ARMA processes have maximal entropy among time series with prescribed autocovariances and impulse responses. Advances in Applied Probability, 18(2): 580, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1427315. See [Fra85].

Anonymous:1986:CAA

[Ano86f] Anonymous. Correction: The ages of alleles and a coalescent. Advances in Applied Probability, 18(4):1023, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427263. See

[DT86].

Anonymous:1986:FMa

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427236.

Anonymous:1986:FMb

[Ano86h] Anonymous. Front matter. Advances in Applied Probability, 18 (2):??, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427300.

Anonymous:1986:FMc

[Ano86i] Anonymous. Front matter. Advances in Applied Probability, 18 (3):??, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427178.

Anonymous:1986:FMd

[Ano86j] Anonymous. Front matter. Advances in Applied Probability, 18 (4):??, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427252.

Anonymous:1986:VI

[Ano86k] Anonymous. Volume information. Advances in Applied Probability, 18(4):1024–1025, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427251.

Anonymous:1987:BMa

[Ano87a] Anonymous. Back matter. Advances in Applied Probability, 19 (1):??, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427386.

Anonymous:1987:BMb

[Ano87b] Anonymous. Back matter. Advances in Applied Probability, 19 (2):??, June 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427434.

Anonymous:1987:BMc

[Ano87c] Anonymous. Back matter. Advances in Applied Probability, 19 (3):??, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427420.

Anonymous:1987:BMd

[Ano87d] Anonymous. Back matter. Advances in Applied Probability, 19 (4):1002, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427114.

Anonymous:1987:FMa

[Ano87e] Anonymous. Front matter. Advances in Applied Probability, 19 (1):??, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427370.

Anonymous:1987:FMb

[Ano87f] Anonymous. Front matter. Advances in Applied Probability, 19 (2):??, June 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427421.

Anonymous:1987:FMc

[Ano87g] Anonymous. Front matter. Advances in Applied Probability, 19 (3):??, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427405.

Anonymous:1987:FMd

[Ano87h] Anonymous. Front matter. Advances in Applied Probability, 19 (4):??, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427099.

Anonymous:1987:VI

[Ano87i] Anonymous. Volume information. Advances in Applied Probability, 19(4):999–1001, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427098.

Anonymous:1988:A

[Ano88a] Anonymous. Acknowledgement. Advances in Applied Probability, 20(4):917, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427368.

Anonymous:1988:BMa

[Ano88b] Anonymous. Back matter. Advances in Applied Probability, 20(1):244, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427282.

Anonymous:1988:BMb

[Ano88c] Anonymous. Back matter. Advances in Applied Probability, 20 (2):488, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427404.

Anonymous:1988:BMc

[Ano88d] Anonymous. Back matter. Advances in Applied Probability, 20 (3):??, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427047.

Anonymous:1988:BMd

[Ano88e] Anonymous. Back matter. Advances in Applied Probability, 20 (4):??, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427369.

Anonymous:1988:FMa

[Ano88f] Anonymous. Front matter. Advances in Applied Probability, 20 (1):1, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427265.

Anonymous:1988:FMb

[Ano88g] Anonymous. Front matter. Advances in Applied Probability, 20 (2):??, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427387.

Anonymous:1988:FMc

[Ano88h] Anonymous. Front matter. Advances in Applied Probability, 20 (3):??, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427032.

Anonymous:1988:FMd

[Ano88i] Anonymous. Front matter. Advances in Applied Probability, 20 (4):??, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427354.

Anonymous:1988:VI

[Ano88j] Anonymous. Volume information. Advances in Applied Probability, 20(4):918–920, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427353.

Anonymous:1989:BMa

[Ano89a] Anonymous. Back matter. Advances in Applied Probability, 21 (1):??, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427213.

Anonymous:1989:BMb

[Ano89b] Anonymous. Back matter. Advances in Applied Probability, 21 (2):??, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427177.

Anonymous:1989:BMc

[Ano89c] Anonymous. Back matter. Advances in Applied Probability, 21 (3):??, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427650.

Anonymous:1989:BMd

[Ano89d] Anonymous. Back matter. Advances in Applied Probability, 21 (4):??, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427785.

Anonymous:1989:CBP

[Ano89e] Anonymous. Correction: Applied Probability author and subject index no. 7, p. 50. Advances in Applied Probability, 21(4):958, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427784.

Anonymous:1989:FMa

[Ano89f] Anonymous. Front matter. Advances in Applied Probability, 21 (1):??, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427194.

Anonymous:1989:FMb

[Ano89g] Anonymous. Front matter. Advances in Applied Probability, 21 (2):??, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427158.

Anonymous:1989:FMc

[Ano89h] Anonymous. Front matter. Advances in Applied Probability, 21 (3):??, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427631.

Anonymous:1989:FMd

[Ano89i] Anonymous. Front matter. Advances in Applied Probability, 21 (4):??, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427762.

Anonymous:1989:VI

[Ano89j] Anonymous. Volume information. Advances in Applied Probability, 21(4):959–961, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427761.

Anonymous:1990:BMa

[Ano90a] Anonymous. Back matter. Advances in Applied Probability, 22 (1):??, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427614.

Anonymous:1990:BMb

[Ano90b] Anonymous. Back matter. Advances in Applied Probability, 22 (2):??, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427558.

Anonymous:1990:BMc

[Ano90c] Anonymous. Back matter. Advances in Applied Probability, 22 (3):??, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427475.

Anonymous:1990:BMd

[Ano90d] Anonymous. Back matter. Advances in Applied Probability, 22 (4):??, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427573.

Anonymous:1990:FMa

[Ano90e] Anonymous. Front matter. Advances in Applied Probability, 22 (1):??, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427593.

Anonymous:1990:FMb

[Ano90f] Anonymous. Front matter. Advances in Applied Probability, 22 (2):??, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427535.

Anonymous:1990:FMc

[Ano90g] Anonymous. Front matter. Advances in Applied Probability, 22 (3):??, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427454.

Anonymous:1990:FMd

[Ano90h] Anonymous. Front matter. Advances in Applied Probability, 22 (4):??, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427560.

Anonymous:1990:VI

[Ano90i] Anonymous. Volume information. Advances in Applied Probability, 22(4):969–971, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427559.

Anonymous:1991:BMa

[Ano91a] Anonymous. Back matter. Advances in Applied Probability, 23 (1):??, March 1991. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427519.

Anonymous:1991:BMb

[Ano91b] Anonymous. Back matter. Advances in Applied Probability, 23 (2):??, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427760.

Anonymous:1991:BMc

[Ano91c] Anonymous. Back matter. Advances in Applied Probability, 23 (3):??, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427630.

Anonymous:1991:BMd

[Ano91d] Anonymous. Back matter. Advances in Applied Probability, 23 (4):??, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427689.

Anonymous:1991:CDC

[Ano91e] Anonymous. Correction: Distributions with complete monotone derivative and geometric infinite divisibility. Advances in Applied Probability, 23(2):436, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427759. See [PS90b].

Anonymous:1991:FMa

[Ano91f] Anonymous. Front matter. Advances in Applied Probability, 23 (1):??, March 1991. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427508.

Anonymous:1991:FMb

[Ano91g] Anonymous. Front matter. Advances in Applied Probability, 23 (2):??, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427745.

Anonymous:1991:FMc

[Ano91h] Anonymous. Front matter. Advances in Applied Probability, 23 (3):??, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427615.

Anonymous:1991:FMd

[Ano91i] Anonymous. Front matter. Advances in Applied Probability, 23 (4):??, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427668.

Anonymous:1991:VI

[Ano91j] Anonymous. Volume information. Advances in Applied Probability, 23(4):978–980, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427667.

Anonymous:1992:BMa

[Ano92a] Anonymous. Back matter. Advances in Applied Probability, 24(1):240, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427744.

Anonymous:1992:BMb

[Ano92b] Anonymous. Back matter. Advances in Applied Probability, 24 (2):??, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427706.

Anonymous:1992:BMc

[Ano92c] Anonymous. Back matter. Advances in Applied Probability, 24 (3):??, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427491.

Anonymous:1992:BMd

[Ano92d] Anonymous. Back matter. Advances in Applied Probability, 24 (4):??, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427725.

Anonymous:1992:CPR

[Ano92e] Anonymous. Correction: Paths in a random digital tree: Limiting distributions. Advances in Applied Probability, 24(3):759, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1427490. See [Pit86].

Anonymous:1992:FMa

[Ano92f] Anonymous. Front matter. Advances in Applied Probability, 24 (1):??, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427726.

Anonymous:1992:FMb

[Ano92g] Anonymous. Front matter. Advances in Applied Probability, 24 (2):??, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427690.

Anonymous:1992:FMc

[Ano92h] Anonymous. Front matter. Advances in Applied Probability, 24 (3):??, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427476.

Anonymous:1992:FMd

[Ano92i] Anonymous. Front matter. Advances in Applied Probability, 24 (4):761, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427708.

Anonymous:1992:VI

[Ano92j] Anonymous. Volume information. Advances in Applied Probability, 24(4):1012–1015, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427707.

Anonymous:1993:BMa

[Ano93a] Anonymous. Back matter. Advances in Applied Probability, 25 (1):??, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427507.

Anonymous:1993:BMb

[Ano93b] Anonymous. Back matter. Advances in Applied Probability, 25 (2):??, June 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427666.

Anonymous:1993:BMc

[Ano93c] Anonymous. Back matter. Advances in Applied Probability, 25 (3):??, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427534.

Anonymous:1993:BMd

[Ano93d] Anonymous. Back matter. Advances in Applied Probability, 25 (4):??, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427806.

Anonymous:1993:CED

[Ano93e] Anonymous. Correction: Exact distribution theory for some point process record models. Advances in Applied Probability, 25(1):261, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427506. See [BN92].

Anonymous:1993:CPS

[Ano93f] Anonymous. Correction: On preservation of some partial orderings under shock models. Advances in Applied Probability, 25 (4):1013, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427805. See [Koc90].

Anonymous:1993:CRB

[Ano93g] Anonymous. Correction: On the relationship between μinvariant measures and quasi-stationary distributions for continuous-time Markov chains. Advances in Applied Probability, 25(3):717–719, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427533. See [NP93].

Anonymous:1993:FMa

[Ano93h] Anonymous. Front matter. Advances in Applied Probability, 25 (1):??, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427492.

Anonymous:1993:FMb

[Ano93i] Anonymous. Front matter. Advances in Applied Probability, 25 (2):??, June 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427651.

Anonymous:1993:FMc

[Ano93j] Anonymous. Front matter. Advances in Applied Probability, 25 (3):??, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427520.

Anonymous:1993:FMd

[Ano93k] Anonymous. Front matter. Advances in Applied Probability, 25 (4):??, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427787.

Anonymous:1993:VI

[Ano93l] Anonymous. Volume information. Advances in Applied Probability, 25(4):1015–1017, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427786.

Anonymous:1994:BMa

[Ano94a] Anonymous. Back matter. Advances in Applied Probability, 26 (1):??, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427592.

Anonymous:1994:BMb

[Ano94b] Anonymous. Back matter. Advances in Applied Probability, 26 (2):??, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427453.

Anonymous:1994:BMc

[Ano94c] Anonymous. Back matter. Advances in Applied Probability, 26 (3):??, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427824.

Anonymous:1994:BMd

[Ano94d] Anonymous. Back matter. Advances in Applied Probability, 26 (4):??, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427908.

Anonymous:1994:CPO

[Ano94e] Anonymous. Correction: Partial orderings under cumulative damage shock models. Advances in Applied Probability, 26(1): 280, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1427591. See [Pel93].

Anonymous:1994:FMa

[Ano94f] Anonymous. Front matter. Advances in Applied Probability, 26 (1):??, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427574.

Anonymous:1994:FMb

[Ano94g] Anonymous. Front matter. Advances in Applied Probability, 26 (2):??, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427435.

Anonymous:1994:FMc

[Ano94h] Anonymous. Front matter. Advances in Applied Probability, 26 (3):??, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427807.

Anonymous:1994:FMd

[Ano94i] Anonymous. Front matter. Advances in Applied Probability, 26 (4):??, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427892.

Anonymous:1994:VI

[Ano94j] Anonymous. Volume information. Advances in Applied Probability, 26(4):1122–1124, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427891.

Anonymous:1995:BMa

[Ano95a] Anonymous. Back matter. Advances in Applied Probability, 27 (1):??, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428108.

Anonymous:1995:BMb

[Ano95b] Anonymous. Back matter. Advances in Applied Probability, 27 (2):??, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427841.

Anonymous:1995:BMc

[Ano95c] Anonymous. Back matter. Advances in Applied Probability, 27 (3):??, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428138.

Anonymous:1995:BMd

[Ano95d] Anonymous. Back matter. Advances in Applied Probability, 27 (4):??, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427939.

Anonymous:1995:CAH

[Ano95e] Anonymous. Corrections to Abdel-Hameed and Daley. Advances in Applied Probability, 27(2):584, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427840. See [AH87, Dal85].

Anonymous:1995:FMa

[Ano95f] Anonymous. Front matter. Advances in Applied Probability, 27 (1):??, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428089.

Anonymous:1995:FMb

[Ano95g] Anonymous. Front matter. Advances in Applied Probability, 27 (2):??, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427825.

Anonymous:1995:FMc

[Ano95h] Anonymous. Front matter. Advances in Applied Probability, 27 (3):??, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428124.

Anonymous:1995:FMd

[Ano95i] Anonymous. Front matter. Advances in Applied Probability, 27 (4):??, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427926.

Anonymous:1995:VI

[Ano95j] Anonymous. Volume information. Advances in Applied Probability, 27(4):1189–1191, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427925.

Anonymous:1996:BMa

[Ano96a] Anonymous. Back matter. Advances in Applied Probability, 28 (1):??, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427924.

Anonymous:1996:BMb

[Ano96b] Anonymous. Back matter. Advances in Applied Probability, 28 (2):??, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428074.

Anonymous:1996:BMc

[Ano96c] Anonymous. Back matter. Advances in Applied Probability, 28 (3):??, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428188.

Anonymous:1996:BMd

[Ano96d] Anonymous. Back matter. Advances in Applied Probability, 28 (4):??, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428171.

Anonymous:1996:FMa

[Ano96e] Anonymous. Front matter. Advances in Applied Probability, 28 (1):??, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427909.

Anonymous:1996:FMb

[Ano96f] Anonymous. Front matter. Advances in Applied Probability, 28 (2):327, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428035.

Anonymous:1996:FMc

[Ano96g] Anonymous. Front matter. Advances in Applied Probability, 28 (3):??, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428172.

Anonymous:1996:FMd

[Ano96h] Anonymous. Front matter. Advances in Applied Probability, 28 (4):??, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428158.

Anonymous:1996:VI

[Ano96i] Anonymous. Volume information. Advances in Applied Probability, 28(4):1227–1229, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428157.

Anonymous:1997:BMa

[Ano97a] Anonymous. Back matter. Advances in Applied Probability, 29 (1):??, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427871.

Anonymous:1997:BMb

[Ano97b] Anonymous. Back matter. Advances in Applied Probability, 29 (2):??, June 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428017.

Anonymous:1997:BMc

[Ano97c] Anonymous. Back matter. Advances in Applied Probability, 29 (3):??, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428088.

Anonymous:1997:BMd

[Ano97d] Anonymous. Back matter. Advances in Applied Probability, 29 (4):??, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427855.

Anonymous:1997:FMa

[Ano97e] Anonymous. Front matter. Advances in Applied Probability, 29 (1):??, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427856.

Anonymous:1997:FMb

[Ano97f] Anonymous. Front matter. Advances in Applied Probability, 29 (2):??, June 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428003.

Anonymous:1997:FMc

[Ano97g] Anonymous. Front matter. Advances in Applied Probability, 29 (3):??, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428075.

Anonymous:1997:FMd

[Ano97h] Anonymous. Front matter. Advances in Applied Probability, 29 (4):??, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427843.

Anonymous:1997:VI

[Ano97i] Anonymous. Volume information. Advances in Applied Probability, 29(4):1081–1083, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427842.

Anonymous:1998:BMa

[Ano98a] Anonymous. Back matter. Advances in Applied Probability, 30 (1):??, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427890.

Anonymous:1998:BMb

[Ano98b] Anonymous. Back matter. Advances in Applied Probability, 30 (2):??, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427985.

Anonymous:1998:BMc

[Ano98c] Anonymous. Back matter. Advances in Applied Probability, 30 (3):??, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428156.

Anonymous:1998:BMd

[Ano98d] Anonymous. Back matter. Advances in Applied Probability, 30 (4):??, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428034.

Anonymous:1998:CCS

[Ano98e] Anonymous. Correction: Convergence of stochastic algorithms: From the Kushner–Clark theorem to the Lyapounov functional method. Advances in Applied Probability, 30(1):270, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427889.

See [FP96].

Anonymous:1998:CHG

[Ano98f] Anonymous. Correction: Homogeneous Gaussian Markov processes on general lattices. Advances in Applied Probability, 30 (1):269, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427888. See [Kom96].

Anonymous:1998:FMa

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427872.

Anonymous:1998:FMb

[Ano98h] Anonymous. Front matter. Advances in Applied Probability, 30 (2):271, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427940.

Anonymous:1998:FMc

[Ano98i] Anonymous. Front matter. Advances in Applied Probability, 30 (3):??, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428139.

Anonymous:1998:FMd

[Ano98j] Anonymous. Front matter. Advances in Applied Probability, 30 (4):??, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428019.

Anonymous:1998:VI

[Ano98k] Anonymous. Volume information. Advances in Applied Probability, 30(4):1157–1160, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428018.

Anonymous:1999:BMa

[Ano99a] Anonymous. Back matter. Advances in Applied Probability, 31 (1):??, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428002.

Anonymous:1999:BMb

[Ano99b] Anonymous. Back matter. Advances in Applied Probability, 31 (2):??, June 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428123.

Anonymous:1999:BMc

[Ano99c] Anonymous. Back matter. Advances in Applied Probability, 31 (3):??, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428400.

Anonymous:1999:BMd

[Ano99d] Anonymous. Back matter. Advances in Applied Probability, 31 (4):??, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428348.

Anonymous:1999:FMa

[Ano99e] Anonymous. Front matter. Advances in Applied Probability, 31 (1):??, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427986.

Anonymous:1999:FMb

[Ano99f] Anonymous. Front matter. Advances in Applied Probability, 31 (2):??, June 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428109.

Anonymous:1999:FMc

[Ano99g] Anonymous. Front matter. Advances in Applied Probability, 31 (3):??, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428383.

Anonymous:1999:FMd

[Ano99h] Anonymous. Front matter. Advances in Applied Probability, 31 (4):??, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428331.

Anonymous:1999:VI

[Ano99i] Anonymous. Volume information. Advances in Applied Probability, 31(4):1151–1154, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428330.

Anonymous:2000:BMa

[Ano00a] Anonymous. Back matter. Advances in Applied Probability, 32 (1):??, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428239.

Anonymous:2000:BMb

[Ano00b] Anonymous. Back matter. Advances in Applied Probability, 32 (2):??, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428205.

Anonymous:2000:BMc

[Ano00c] Anonymous. Back matter. Advances in Applied Probability, 32 (3):924, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428419.

Anonymous:2000:BMd

[Ano00d] Anonymous. Back matter. Advances in Applied Probability, 32 (4):??, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428524.

Anonymous:2000:FMa

[Ano00e] Anonymous. Front matter. Advances in Applied Probability, 32 (1):??, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428222.

Anonymous:2000:FMb

[Ano00f] Anonymous. Front matter. Advances in Applied Probability, 32 (2):??, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428189.

Anonymous:2000:FMc

[Ano00g] Anonymous. Front matter. Advances in Applied Probability, 32 (3):??, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428401.

Anonymous:2000:FMd

[Ano00h] Anonymous. Front matter. Advances in Applied Probability, 32 (4):??, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428507.

Anonymous:2000:VI

[Ano00i] Anonymous. Volume information. Advances in Applied Probability, 32(4):1193–1197, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428506.

Anonymous:2001:BMa

[Ano01a] Anonymous. Back matter. Advances in Applied Probability, 33 (1):??, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428453.

Anonymous:2001:BMb

[Ano01b] Anonymous. Back matter. Advances in Applied Probability, 33 (2):??, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428267.

Anonymous:2001:BMc

[Ano01c] Anonymous. Back matter. Advances in Applied Probability, 33 (3):??, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428484.

Anonymous:2001:BMd

[Ano01d] Anonymous. Back matter. Advances in Applied Probability, 33 (4):??, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428314.

Anonymous:2001:FMa

[Ano01e] Anonymous. Front matter. Advances in Applied Probability, 33 (1):??, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428436.

Anonymous:2001:FMb

[Ano01f] Anonymous. Front matter. Advances in Applied Probability, 33 (2):??, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428253.

Anonymous:2001:FMc

[Ano01g] Anonymous. Front matter. Advances in Applied Probability, 33 (3):??, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428472.

Anonymous:2001:FMd

[Ano01h] Anonymous. Front matter. Advances in Applied Probability, 33 (4):??, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428299.

Anonymous:2001:VI

[Ano01i] Anonymous. Volume information. Advances in Applied Probability, 33(4):944–946, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428298.

Anonymous:2002:BMa

[Ano02a] Anonymous. Back matter. Advances in Applied Probability, 34 (1):??, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428382.

Anonymous:2002:BMb

[Ano02b] Anonymous. Back matter. Advances in Applied Probability, 34 (2):??, June 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428297.

Anonymous:2002:BMc

[Ano02c] Anonymous. Back matter. Advances in Applied Probability, 34 (3):??, September 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428252.

Anonymous:2002:BMd

[Ano02d] Anonymous. Back matter. Advances in Applied Probability, 34 (4):??, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428329.

Anonymous:2002:FMa

[Ano02e] Anonymous. Front matter. Advances in Applied Probability, 34 (1):??, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428367.

Anonymous:2002:FMb

[Ano02f] Anonymous. Front matter. Advances in Applied Probability, 34 (2):??, June 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428285.

Anonymous:2002:FMc

[Ano02g] Anonymous. Front matter. Advances in Applied Probability, 34 (3):??, September 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428240.

Anonymous:2002:FMd

[Ano02h] Anonymous. Front matter. Advances in Applied Probability, 34 (4):??, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428316.

Anonymous:2002:VI

[Ano02i] Anonymous. Volume information. Advances in Applied Probability, 34(4):923–925, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428315.

Anonymous:2003:BMa

[Ano03a] Anonymous. Back matter. Advances in Applied Probability, 35 (1):??, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428284.

Anonymous:2003:BMb

[Ano03b] Anonymous. Back matter. Advances in Applied Probability, 35 (2):??, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428435.

Anonymous:2003:BMc

[Ano03c] Anonymous. Back matter. Advances in Applied Probability, 35 (3):??, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428221.

Anonymous:2003:BMd

[Ano03d] Anonymous. Back matter. Advances in Applied Probability, 35 (4):??, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428505.

Anonymous:2003:CES

[Ano03e] Anonymous. Correction: Exact sampling from conditional Boolean models with applications to maximum likelihood inference. Advances in Applied Probability, 35(2):362, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428426. See

[VV01].

Anonymous:2003:FMa

[Ano03f] Anonymous. Front matter. Advances in Applied Probability, 35 (1):??, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428268.

Anonymous:2003:FMb

[Ano03g] Anonymous. Front matter. Advances in Applied Probability, 35 (2):??, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428420.

Anonymous:2003:FMc

[Ano03h] Anonymous. Front matter. Advances in Applied Probability, 35 (3):??, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428206.

Anonymous:2003:FMd

[Ano03i] Anonymous. Front matter. Advances in Applied Probability, 35 (4):??, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428486.

Anonymous:2003:VI

[Ano03j] Anonymous. Volume information. Advances in Applied Probability, 35(4):1169–1172, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428485.

Anonymous:2004:BMa

[Ano04a] Anonymous. Back matter. Advances in Applied Probability, 36 (1):??, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428366.

Anonymous:2004:BMb

[Ano04b] Anonymous. Back matter. Advances in Applied Probability, 36 (2):??, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428471.

Anonymous:2004:BMc

[Ano04c] Anonymous. Back matter. Advances in Applied Probability, 36 (3):??, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/4140418.

Anonymous:2004:BMd

[Ano04d] Anonymous. Back matter. Advances in Applied Probability, 36 (4):??, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/4140400.

Anonymous:2004:CBM

[Ano04e] Anonymous. Correction: The Multiclass GI/PHA/V Queue in the Halfin–Whitt Regime. Advances in Applied Probability, 36 (3):971, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/4140417. See [PR00].

Anonymous:2004:CCS

[Ano04f] Anonymous. Correction: Countable-state-space Markov chains with two time scales and applications to queueing systems. Advances in Applied Probability, 36(1):324, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428365. See [YZ02].

Anonymous:2004:CTS

[Ano04g] Anonymous. Correction: Taylor series expansions for stationary Markov chains. Advances in Applied Probability, 36(4):1300, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

4140399. See [HH03].

Anonymous:2004:FMa

[Ano04h] Anonymous. Front matter. Advances in Applied Probability, 36 (1):??, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428349.

Anonymous:2004:FMb

[Ano04i] Anonymous. Front matter. Advances in Applied Probability, 36 (2):??, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428454.

Anonymous:2004:FMc

[Ano04j] Anonymous. Front matter. Advances in Applied Probability, 36 (3):??, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/4140401.

Anonymous:2004:FMd

[Ano04k] Anonymous. Front matter. Advances in Applied Probability, 36 (4):??, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/4140383.

Anonymous:2004:VI

[Ano04l] Anonymous. Volume information. Advances in Applied Probability, 36(4):1301–1304, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/4140382.

Anonymous:2005:BMa

[Ano05a] Anonymous. Back matter. Advances in Applied Probability, 37 (1):??, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
30037325.

Anonymous:2005:BMb

[Ano05b] Anonymous. Back matter. Advances in Applied Probability, 37 (2):??, June 2005. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
30037342.

Anonymous:2005:BMc

[Ano05c] Anonymous. Back matter. Advances in Applied Probability, 37 (3):??, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/30037360.

Anonymous:2005:BMd

[Ano05d] Anonymous. Back matter. Advances in Applied Probability, 37 (4):??, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/30037377.

Anonymous:2005:FMa

[Ano05e] Anonymous. Front matter. Advances in Applied Probability, 37 (1):??, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
30037312.

Anonymous:2005:FMb

[Ano05f] Anonymous. Front matter. Advances in Applied Probability, 37 (2):??, June 2005. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
30037327.

Anonymous:2005:FMc

[Ano05g] Anonymous. Front matter. Advances in Applied Probability, 37 (3):??, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/30037344.

Anonymous:2005:FMd

[Ano05h] Anonymous. Front matter. Advances in Applied Probability, 37 (4):??, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/30037363.

Anonymous:2005:VI

[Ano05i] Anonymous. Volume information. Advances in Applied Probability, 37(4):??, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/30037362.

Anonymous:2006:BMa

[Ano06a] Anonymous. Back matter. Advances in Applied Probability, 38 (1):??, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443440.

Anonymous:2006:BMb

[Ano06b] Anonymous. Back matter. Advances in Applied Probability, 38 (2):??, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443459.

Anonymous:2006:BMc

[Ano06c] Anonymous. Back matter. Advances in Applied Probability, 38 (3):??, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443474.

Anonymous:2006:BMd

[Ano06d] Anonymous. Back matter. Advances in Applied Probability, 38 (4):??, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443493.

Anonymous:2006:CCD

[Ano06e] Anonymous. Correction: Connect the dots: How many random points can a regular curve pass through? Advances in Applied Probability, 38(2):579, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443457. See [ACDHT05].

Anonymous:2006:CED

[Ano06f] Anonymous. Correction: Exact distribution of word counts in shuffled sequences. Advances in Applied Probability, 38(2): 580, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

20443458. See [Rød06].

Anonymous:2006:FMa

[Ano06g] Anonymous. Front matter. Advances in Applied Probability, 38 (1):??, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443424.

Anonymous:2006:FMb

[Ano06h] Anonymous. Front matter. Advances in Applied Probability, 38 (2):??, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443442.

Anonymous:2006:FMc

[Ano06i] Anonymous. Front matter. Advances in Applied Probability, 38 (3):??, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443461.

Anonymous:2006:FMd

[Ano06j] Anonymous. Front matter. Advances in Applied Probability, 38 (4):??, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443477.

Anonymous:2006:VIa

[Ano06k] Anonymous. Volume information. Advances in Applied Probability, 38(1):??, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443423.

Anonymous:2006:VIb

[Ano06l] Anonymous. Volume information. Advances in Applied Probability, 38(4):??, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443476.

Anonymous:2007:BMa

[Ano07a] Anonymous. Back matter. Advances in Applied Probability, 39 (1):??, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443510.

Anonymous:2007:BMb

[Ano07b] Anonymous. Back matter. Advances in Applied Probability, 39 (2):??, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443529.

Anonymous:2007:BMc

[Ano07c] Anonymous. Back matter. Advances in Applied Probability, 39 (3):??, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443543.

Anonymous:2007:BMd

[Ano07d] Anonymous. Back matter. Advances in Applied Probability, 39 (4):??, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443562.

Anonymous:2007:CBN

[Ano07e] Anonymous. Correction: [“Stability of Nonlinear Stochastic Recursions with Application to Nonlinear AR–GARCH Models]. Advances in Applied Probability, 39(4):1115–1116, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/20443561.

See [Cli07].

Anonymous:2007:FMa

[Ano07f] Anonymous. Front matter. Advances in Applied Probability, 39 (1):??, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443496.

Anonymous:2007:FMb

[Ano07g] Anonymous. Front matter. Advances in Applied Probability, 39 (2):??, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443512.

Anonymous:2007:FMc

[Ano07h] Anonymous. Front matter. Advances in Applied Probability, 39 (3):??, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443531.

Anonymous:2007:FMd

[Ano07i] Anonymous. Front matter. Advances in Applied Probability, 39 (4):??, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443546.

Anonymous:2007:VIa

[Ano07j] Anonymous. Volume information. Advances in Applied Probability, 39(1):??, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443495.

Anonymous:2007:VIb

[Ano07k] Anonymous. Volume information. Advances in Applied Probability, 39(4):??, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443545.

Anonymous:2008:BMa

[Ano08a] Anonymous. Back matter. Advances in Applied Probability, 40 (1):??, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443581.

Anonymous:2008:BMb

[Ano08b] Anonymous. Back matter. Advances in Applied Probability, 40 (2):??, June 2008. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443597.

Anonymous:2008:BMc

[Ano08c] Anonymous. Back matter. Advances in Applied Probability, 40 (3):??, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443614.

Anonymous:2008:BMd

[Ano08d] Anonymous. Back matter. Advances in Applied Probability, 40 (4):??, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443632.

Anonymous:2008:FMa

[Ano08e] Anonymous. Front matter. Advances in Applied Probability, 40 (1):??, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443565.

Anonymous:2008:FMb

[Ano08f] Anonymous. Front matter. Advances in Applied Probability, 40 (2):??, June 2008. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443583.

Anonymous:2008:FMc

[Ano08g] Anonymous. Front matter. Advances in Applied Probability, 40 (3):??, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443599.

Anonymous:2008:FMd

[Ano08h] Anonymous. Front matter. Advances in Applied Probability, 40 (4):??, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443617.

Anonymous:2008:VIa

[Ano08i] Anonymous. Volume information. Advances in Applied Probability, 40(1):??, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443564.

Anonymous:2008:VIb

[Ano08j] Anonymous. Volume information. Advances in Applied Probability, 40(4):??, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443616.

Anonymous:2009:BMa

[Ano09a] Anonymous. Back matter. Advances in Applied Probability, 41 (1):??, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
27793876.

Anonymous:2009:BMb

[Ano09b] Anonymous. Back matter. Advances in Applied Probability, 41 (2):??, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
27793893.

Anonymous:2009:BMc

[Ano09c] Anonymous. Back matter. Advances in Applied Probability, 41 (3):??, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/27793908.

Anonymous:2009:BMd

[Ano09d] Anonymous. Back matter. Advances in Applied Probability, 41 (4):??, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/27793926.

Anonymous:2009:CCF

[Ano09e] Anonymous. Correction: A characterization of the first hitting time of double integral processes to curved boundaries. Advances in Applied Probability, 41(1):309, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/27793875. See [TF08].

Anonymous:2009:FMa

[Ano09f] Anonymous. Front matter. Advances in Applied Probability, 41 (1):??, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
27793861.

Anonymous:2009:FMb

[Ano09g] Anonymous. Front matter. Advances in Applied Probability, 41 (2):??, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
27793878.

Anonymous:2009:FMc

[Ano09h] Anonymous. Front matter. Advances in Applied Probability, 41 (3):??, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/27793895.

Anonymous:2009:FMd

[Ano09i] Anonymous. Front matter. Advances in Applied Probability, 41 (4):??, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/27793910.

Anonymous:2010:BMa

[Ano10a] Anonymous. Back matter. Advances in Applied Probability, 42 (1):??, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25683817.

Anonymous:2010:BMb

[Ano10b] Anonymous. Back matter. Advances in Applied Probability, 42 (2):??, June 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25683835.

Anonymous:2010:BMc

[Ano10c] Anonymous. Back matter. Advances in Applied Probability, 42 (3):??, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25799074.

Anonymous:2010:BMd

[Ano10d] Anonymous. Back matter. Advances in Applied Probability, 42 (4):??, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25799091.

Anonymous:2010:FMa

[Ano10e] Anonymous. Front matter. Advances in Applied Probability, 42 (1):??, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25683803.

Anonymous:2010:FMb

[Ano10f] Anonymous. Front matter. Advances in Applied Probability, 42 (2):??, June 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25683819.

Anonymous:2010:FMc

[Ano10g] Anonymous. Front matter. Advances in Applied Probability, 42 (3):??, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25799058.

Anonymous:2010:FMd

[Ano10h] Anonymous. Front matter. Advances in Applied Probability, 42 (4):??, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25799076.

Anonymous:2012:CSD

[Ano12] Anonymous. Correction: Stochastic and deterministic analysis of SIS household epidemics. Advances in Applied Probability, 44 (1):309–310, March 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/23207866. See [Nea06].

Anonymous:2013:BAPa

[Ano13a] Anonymous. Advances in Applied Probability volume 45 (2013): Index general applied probability. Advances in Applied Probability, 45(4):1198–1200, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1386857864.

Anonymous:2013:BAPb

[Ano13b] Anonymous. Advances in Applied Probability volume 45 (2013): Index stochastic geometry and statistical applications. Advances in Applied Probability, 45(4):1201, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1386857865.

Anonymous:2014:AAPa

[Ano14a] Anonymous. Advances in Applied Probability volume 46 (2014): Index general applied probability. Advances in Applied Probability, 46(4):??, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1418396249.

Anonymous:2014:AAPb

[Ano14b] Anonymous. Advances in Applied Probability volume 46 (2014): Index stochastic geometry and statistical applications. Advances in Applied Probability, 46(4):??, December 2014. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1418396250.

Anonymous:2015:AAPa

[Ano15a] Anonymous. Advances in Applied Probability volume 47 (2015): Index general applied probability. Advances in Applied Probability, 47(4):1235–1237, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1449859808.

Anonymous:2015:AAPb

[Ano15b] Anonymous. Advances in Applied Probability volume 47 (2015): Index stochastic geometry and statistical applications. Advances in Applied Probability, 47(4):1238, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1449859809.

Anonymous:2017:AVIb

[Ano17a] Anonymous. APR volume 49 issue 1 cover and back matter. Advances in Applied Probability, 49(1):b1–b2, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-49-
issue-1-cover-and-back-matter/9DD49569C65D89D78445F94A0295CE0E.

Anonymous:2017:AVIa

[Ano17b] Anonymous. APR volume 49 issue 1 cover and front matter. Advances in Applied Probability, 49(1):f1–f4, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-49-issue-1-cover-
and-front-matter/7C0639717054F7C586D7332B6AA525DF.

Anonymous:2017:AVId

[Ano17c] Anonymous. APR volume 49 issue 2 cover and back matter. Advances in Applied Probability, 49(2):b1–b2, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-49-issue-2-cover-
and-back-matter/540E3402FC6A6F0668E46F0D524E195A.

Anonymous:2017:AVIc

[Ano17d] Anonymous. APR volume 49 issue 2 cover and front matter. Advances in Applied Probability, 49(2):f1–f2, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-49-issue-2-cover-
and-front-matter/91A71D29786F3D4D30F5C9B33024938A.

Anonymous:2017:AVIf

[Ano17e] Anonymous. APR volume 49 issue 3 cover and back matter. Advances in Applied Probability, 49(3):b1–b2, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-49-
issue-3-cover-and-back-matter/024F4EF7AADAFA25DA8AA0C948F6E40F.

Anonymous:2017:AVIe

[Ano17f] Anonymous. APR volume 49 issue 3 cover and front matter. Advances in Applied Probability, 49(3):f1–f2, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-49-
issue-3-cover-and-front-matter/93A8B40E0CFA6E4A1393DC45E429C42F.

Anonymous:2017:AVIh

[Ano17g] Anonymous. APR volume 49 issue 4 cover and back matter. Advances in Applied Probability, 49(4):b1–b6, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-49-
issue-4-cover-and-back-matter/4969F3D5795715B659445F282F22BC68.

Anonymous:2017:AVIg

[Ano17h] Anonymous. APR volume 49 issue 4 cover and front matter. Advances in Applied Probability, 49(4):f1–f2, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-49-
issue-4-cover-and-front-matter/A14442209C99820583CDB9F33DFB78B2.

Anonymous:2018:AVIb

[Ano18a] Anonymous. APR volume 50 issue 1 cover and back matter. Advances in Applied Probability, 50(1):b1–b2, March 2018.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-50-
issue-1-cover-and-back-matter/844E51628252799F932124C9D3890E5B.

Anonymous:2018:AVIa

[Ano18b] Anonymous. APR volume 50 issue 1 cover and front matter. Advances in Applied Probability, 50(1):f1–f4, March 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-50-issue-1-cover-
and-front-matter/03FD860D4C00F63C7609C8F4DB32DCC8.

Anonymous:2018:AVId

[Ano18c] Anonymous. APR volume 50 issue 2 cover and back matter. Advances in Applied Probability, 50(2):b1–b2, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-50-issue-2-cover-
and-back-matter/F8D97241DE2E5AABC6C6C1CAD446ED00.

Anonymous:2018:AVIc

[Ano18d] Anonymous. APR volume 50 issue 2 cover and front matter. Advances in Applied Probability, 50(2):f1–f2, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-50-issue-2-cover-
and-front-matter/FB1A50529A27ED6631A47255D8BEEE75.

Anonymous:2018:AVIf

[Ano18e] Anonymous. APR volume 50 issue 3 cover and back matter. Advances in Applied Probability, 50(3):b1–b2, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-50-
issue-3-cover-and-back-matter/80401BAC65EAA13788C77234E2165FB7.

Anonymous:2018:AVIe

[Ano18f] Anonymous. APR volume 50 issue 3 cover and front matter. Advances in Applied Probability, 50(3):f1–f2, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-50-
issue-3-cover-and-front-matter/FCD45BF95A95AABEE73D7ABAB20F9C8A.

Anonymous:2018:AVIh

[Ano18g] Anonymous. APR volume 50 issue 4 cover and back matter. Advances in Applied Probability, 50(4):b1–b6, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-50-
issue-4-cover-and-back-matter/6D9534333D06C50FFF4F601DEAA00D0D.

Anonymous:2018:AVIg

[Ano18h] Anonymous. APR volume 50 issue 4 cover and front matter. Advances in Applied Probability, 50(4):f1–f2, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-50-
issue-4-cover-and-front-matter/EE7E3A443CD181A8C4BAF1C2866933F4.

Anonymous:2019:AVIb

[Ano19a] Anonymous. APR volume 51 issue 1 cover and back matter. Advances in Applied Probability, 51(1):b1–b2, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-51-
issue-1-cover-and-back-matter/84A8E34C136BD274D235367B379B4841.

Anonymous:2019:AVIa

[Ano19b] Anonymous. APR volume 51 issue 1 cover and front matter. Advances in Applied Probability, 51(1):f1–f4, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-51-issue-1-cover-
and-front-matter/34A16464B2F231E81C92F17BE4263E43.

Anonymous:2019:AVId

[Ano19c] Anonymous. APR volume 51 issue 2 cover and back matter. Advances in Applied Probability, 51(2):b1–b2, June 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-51-issue-2-cover-
and-back-matter/D5855998E0E6C855B36FE4370AF8D3BF.

Anonymous:2019:AVIc

[Ano19d] Anonymous. APR volume 51 issue 2 cover and front matter. Advances in Applied Probability, 51(2):f1–f2, June 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-51-issue-2-cover-
and-front-matter/9D2398BBE60499D5BBDCE7C25D877FB4.

Anonymous:2019:AVIf

[Ano19e] Anonymous. APR volume 51 issue 3 cover and back matter. Advances in Applied Probability, 51(3):b1–b2, September 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-51-
issue-3-cover-and-back-matter/55AC181F0E4FBEC280861AF88029C6E3.

Anonymous:2019:AVIe

[Ano19f] Anonymous. APR volume 51 issue 3 cover and front matter. Advances in Applied Probability, 51(3):f1–f2, September 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-51-
issue-3-cover-and-front-matter/E3C9B54305605116D30435334C46AB64.

Anonymous:2019:AVIh

[Ano19g] Anonymous. APR volume 51 issue 4 cover and back matter. Advances in Applied Probability, 51(4):b1–b2, December 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-51-
issue-4-cover-and-back-matter/517E20652021A58FF958EE433AD4A9D0.

Anonymous:2019:AVIg

[Ano19h] Anonymous. APR volume 51 issue 4 cover and front matter. Advances in Applied Probability, 51(4):f1–f2, December 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-51-
issue-4-cover-and-front-matter/8D2B0F0628FC709AE6FA57AE2AAAB28F.

Anonymous:2020:AVIb

[Ano20a] Anonymous. APR volume 52 issue 1 cover and back matter. Advances in Applied Probability, 52(1):b1–b2, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-52-
issue-1-cover-and-back-matter/0E4BF095F4E2AED51BC763E0D6C52F07.

Anonymous:2020:AVIa

[Ano20b] Anonymous. APR volume 52 issue 1 cover and front matter. Advances in Applied Probability, 52(1):f1–f2, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-52-issue-1-cover-
and-front-matter/07D05B93B2E593615C58BA2B886AC0DA.

Anonymous:2020:AVId

[Ano20c] Anonymous. APR volume 52 issue 2 cover and back matter. Advances in Applied Probability, 52(2):b1–b2, June 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-52-issue-2-cover-
and-back-matter/E946D03B39AF61858DED39D277749AD9.

Anonymous:2020:AVIc

[Ano20d] Anonymous. APR volume 52 issue 2 cover and front matter. Advances in Applied Probability, 52(2):f1–f2, June 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-52-issue-2-cover-
and-front-matter/ED351C014CCFCC8B9A6137DA4724EE21.

Anonymous:2020:AVIf

[Ano20e] Anonymous. APR volume 52 issue 3 cover and back matter. Advances in Applied Probability, 52(3):b1–b2, September 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-52-
issue-3-cover-and-back-matter/6E9B28B62E54BA5AB9C347C18FB310A1.

Anonymous:2020:AVIe

[Ano20f] Anonymous. APR volume 52 issue 3 cover and front matter. Advances in Applied Probability, 52(3):f1–f2, September 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-52-
issue-3-cover-and-front-matter/C53B0DC058EFFC59B08CB0D11B5FFB52.

Anonymous:2020:AVIh

[Ano20g] Anonymous. APR volume 52 issue 4 cover and back matter. Advances in Applied Probability, 52(4):b1–b2, December 2020.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-52-
issue-4-cover-and-back-matter/246C29B19A79F3FDA0DD9D2139FA6565.

Anonymous:2020:AVIg

[Ano20h] Anonymous. APR volume 52 issue 4 cover and front matter. Advances in Applied Probability, 52(4):f1–f2, December 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-52-
issue-4-cover-and-front-matter/FA3658250F003C44450A6FF1961441D3.

Anonymous:2021:AVIb

[Ano21a] Anonymous. APR volume 53 issue 1 cover and back matter. Advances in Applied Probability, 53(1):b1–b2, March 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-53-
issue-1-cover-and-back-matter/57AD28BF799F189D64301AB60CD10A88.

Anonymous:2021:AVIa

[Ano21b] Anonymous. APR volume 53 issue 1 cover and front matter. Advances in Applied Probability, 53(1):f1–f2, March 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-53-issue-1-cover-
and-front-matter/80A548CE365FC2815E3FF213E471AFCE.

Anonymous:2021:AVId

[Ano21c] Anonymous. APR volume 53 issue 2 cover and back matter. Advances in Applied Probability, 53(2):b1–b2, June 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-53-issue-2-cover-
and-back-matter/AE722FE2F736C76C66F646112EC3600F.

Anonymous:2021:AVIc

[Ano21d] Anonymous. APR volume 53 issue 2 cover and front matter. Advances in Applied Probability, 53(2):f1–f2, June 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-53-issue-2-cover-
and-front-matter/B6C3E3921C74BCDAC555991A0E3FB82E.

Anonymous:2021:AVIf

[Ano21e] Anonymous. APR volume 53 issue 3 cover and back matter. Advances in Applied Probability, 53(3):b1–b2, September 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-53-
issue-3-cover-and-back-matter/E4915C9BCF85B5C46CCDDDA48E6BF6FC.

Anonymous:2021:AVIe

[Ano21f] Anonymous. APR volume 53 issue 3 cover and front matter. Advances in Applied Probability, 53(3):f1–f2, September 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-53-
issue-3-cover-and-front-matter/D62B910F835E120FC3AD9F4EA1773748.

Anonymous:2021:AVIh

[Ano21g] Anonymous. APR volume 53 issue 4 cover and back matter. Advances in Applied Probability, 53(4):b1–b2, December 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-53-
issue-4-cover-and-back-matter/1CD412A4D81422A0EE5AB4B898C0F124.

Anonymous:2021:AVIg

[Ano21h] Anonymous. APR volume 53 issue 4 cover and front matter. Advances in Applied Probability, 53(4):f1–f2, December 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-53-
issue-4-cover-and-front-matter/0B756DE8077205D4AAD96D233DC31921.

Anonymous:2022:AVIb

[Ano22a] Anonymous. APR volume 54 issue 1 cover and back matter. Advances in Applied Probability, 54(1):b1–b2, March 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-54-
issue-1-cover-and-back-matter/E2DA6843C653EE52A8F5F602E018C037.

Anonymous:2022:AVIa

[Ano22b] Anonymous. APR volume 54 issue 1 cover and front matter. Advances in Applied Probability, 54(1):f1–f2, March 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-54-issue-1-cover-
and-front-matter/F985D511364D342E920F45CD0AFB84C8.

Anonymous:2022:AVId

[Ano22c] Anonymous. APR volume 54 issue 2 cover and back matter. Advances in Applied Probability, 54(2):b1–b2, June 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-54-issue-2-cover-
and-back-matter/F8E80AC7D0E618C7AD7984450E4470F0.

Anonymous:2022:AVIc

[Ano22d] Anonymous. APR volume 54 issue 2 cover and front matter. Advances in Applied Probability, 54(2):f1–f2, June 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-54-issue-2-cover-
and-front-matter/F6F2696830C426CCD9569BACE59B40A8.

Anonymous:2022:AVIf

[Ano22e] Anonymous. APR volume 54 issue 3 cover and back matter. Advances in Applied Probability, 54(3):b1–b2, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-54-
issue-3-cover-and-back-matter/E18789C85587E278B2605B114C8843C1.

Anonymous:2022:AVIe

[Ano22f] Anonymous. APR volume 54 issue 3 cover and front matter. Advances in Applied Probability, 54(3):f1–f2, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-54-
issue-3-cover-and-front-matter/F3417897DCA441D64DBD4BD466E4A876.

Anonymous:2022:AVIh

[Ano22g] Anonymous. APR volume 54 issue 4 cover and back matter. Advances in Applied Probability, 54(4):b1–b2, December 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-54-
issue-4-cover-and-back-matter/3123CF58F02B89872A4B4CEDEAD9B965.

Anonymous:2022:AVIg

[Ano22h] Anonymous. APR volume 54 issue 4 cover and front matter. Advances in Applied Probability, 54(4):f1–f2, December 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-54-
issue-4-cover-and-front-matter/EE748BF7B481B9EB7CE7A7051A6B7B40.

Anonymous:2023:AVIb

[Ano23a] Anonymous. APR volume 55 issue 1 cover and back matter. Advances in Applied Probability, 55(1):b1–b2, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-55-
issue-1-cover-and-back-matter/C66400D24FFF3F95F719937688AD8644.

Anonymous:2023:AVIa

[Ano23b] Anonymous. APR volume 55 issue 1 cover and front matter. Advances in Applied Probability, 55(1):f1–f2, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-55-issue-1-cover-
and-front-matter/6B21E67054DF7A1C5F8777566676593F.

Anonymous:2023:AVId

[Ano23c] Anonymous. APR volume 55 issue 2 cover and back matter. Advances in Applied Probability, 55(2):b1–b2, June 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-55-issue-2-cover-
and-back-matter/6B1270BBF9B3A8EB9DE3652E78D74784.

Anonymous:2023:AVIc

[Ano23d] Anonymous. APR volume 55 issue 2 cover and front matter. Advances in Applied Probability, 55(2):f1–f2, June 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/apr-volume-55-issue-2-cover-
and-front-matter/E8F34CFB39D749FE51707FB89A7F6DEF.

Anonymous:2023:AVIf

[Ano23e] Anonymous. APR volume 55 issue 3 cover and back matter. Advances in Applied Probability, 55(3):b1–b2, September 2023.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-55-
issue-3-cover-and-back-matter/4189F2DDB786E7A9B829650E8A4A0F14.

Anonymous:2023:AVIe

[Ano23f] Anonymous. APR volume 55 issue 3 cover and front matter. Advances in Applied Probability, 55(3):f1–f2, September 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/apr-volume-55-
issue-3-cover-and-front-matter/1E5DCFEFD290BACA0B104C1E3017059D.

Appel:2002:LLD

[ANR02] Martin J. B. Appel, Christopher A. Najim, and Ralph P. Russo. Limit laws for the diameter of a random point set. Advances in Applied Probability, 34(1):1–10, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428368.

Antonelli:1978:GRD

[Ant78] Peter L. Antonelli. The geometry of random drift IV. random time substitutions and stationary densities. Advances in Applied Probability, 10(3):563–569, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426634.

Antonelli:1979:GRD

[Ant79] Peter L. Antonelli. The geometry of random drift V. Axiomatic derivation of the WFK diffusion from a variational principle. Advances in Applied Probability, 11(3):502–509, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426951.

Ambartzumian:1995:MGE

[AO95] R. V. Ambartzumian and V. K. Oganian. Measure generation by Euler functionals. Advances in Applied Probability, 27(3): 606–626, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428126.

Asmussen:1988:RPE

[AP88] Søren Asmussen and Søren Schock Petersen. Ruin probabilities expressed in terms of storage processes. Advances in Applied Probability, 20(4):913–916, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427367.

Adler:1997:SBP

[AP97] Robert J. Adler and Ron Pyke. Scanning Brownian processes. Advances in Applied Probability, 29(2):295–326, June 1997. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428004.

Alaya:1998:RCC

[AP98] Mohamed Ben Alaya and Gilles Pagès. Rate of convergence for computing expectations of stopping functionals of an α-mixing process. Advances in Applied Probability, 30(2):425–448, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427976.

Aldous:2005:CBP

[AP05] David Aldous and Lea Popovic. A critical branching process model for biodiversity. Advances in Applied Probability, 37(4): 1094–1115, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/30037375.

Adell:1999:SOP

[APP99] Jose A. Adell and Ana Perez-Palomares. Stochastic orders in preservation properties by Bernstein-type operators. Advances in Applied Probability, 31(2):492–509, June 1999. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428119.

Arapostathis:2021:EES

[APZ21] Ari Arapostathis, Guodong Pang, and Yi Zheng. Exponential ergodicity and steady-state approximations for a class of Markov processes under fast regime switching. Advances in Applied Probability, 53(1):1–29, March 2021. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/exponential-
ergodicity-and-steadystate-approximations-for-a-class-
of-markov-processes-under-fast-regime-switching/71604D53D14A7A01E3EC070B61BFED3A.

Asmussen:1992:EHT

[AR92] Søren Asmussen and Reuven Y. Rubinstein. The efficiency and heavy traffic properties of the score function method in sensitivity analysis of queueing models. Advances in Applied Probability, 24(1):172–201, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427735.

Appel:1997:MVDa

[AR97a] Martin J. B. Appel and Ralph P. Russo. The maximum vertex degree of a graph on uniform points in [0, 1]^{d}. Advances in Applied Probability, 29(3):567–581, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428076.

Appel:1997:MVDb

[AR97b] Martin J. B. Appel and Ralph P. Russo. The minimum vertex degree of a graph on uniform points in [0, 1]^{d}. Advances in Applied Probability, 29(3):582–594, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428077.

Ahn:2006:DLB

[AR06] Hyun-Soo Ahn and Rhonda Righter. Dynamic load balancing with flexible workers. Advances in Applied Probability, 38(3): 621–642, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443464.

Arjas:1972:FIT

[Arj72a] E. Arjas. On a fundamental identity in the theory of semi-Markov processes. Advances in Applied Probability, 4(2):258– 270, August 1972. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425998.

Arjas:1972:UFI

[Arj72b] E. Arjas. On the use of a fundamental identity in the theory of semi-Markov queues. Advances in Applied Probability, 4(2): 271–284, August 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425999.

Arjas:1980:CRL

[Arj80] Elja Arjas. Classes of residual lifetime distributions conditioned on failure histories. Advances in Applied Probability, 12(2): 267, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426533.

Angulo:1999:MRA

[ARM99] J. M. Angulo and M. D. Ruiz-Medina. Multi-resolution approximation to the stochastic inverse problem. Advances in Applied Probability, 31(4):1039–1057, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428343.

Angulo:2000:FDF

[ARMAG00] J. M. Angulo, M. D. Ruiz-Medina, V. V. Anh, and W. Grecksch. Fractional diffusion and fractional heat equation. Advances in Applied Probability, 32(4):1077–1099, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428518.

Arnold:1980:MMC

[Arn80] L. Arnold. Mathematical models of chemical reactions. Advances in Applied Probability, 12(2):304, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426580.

Arrigoni:2003:DAS

[Arr03] Francesca Arrigoni. Deterministic approximation of a stochastic metapopulation model. Advances in Applied Probability, 35(3): 691–720, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428214.

Athreya:2004:CSR

[ARS04] Siva Athreya, Rahul Roy, and Anish Sarkar. On the coverage of space by random sets. Advances in Applied Probability, 36(1): 1–18, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428350.

Alonso-Ruiz:2017:EEP

[ARS17] P. Alonso-Ruiz and E. Spodarev. Estimation of entropy for Poisson marked point processes. Advances in Applied Probability, 49(1):258–278, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/estimation-of-entropy-for-poisson-
marked-point-processes/93FC4B0BA666B59FA2A02953A30127C8.

Akgun:2011:MSS

[ARW11] Osman T. Akgun, Rhonda Righter, and Ronald Wolff. Multipleserver system with flexible arrivals. Advances in Applied Probability, 43(4):985–1004, December 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/23208035.

Akgun:2013:PFR

[ARW13] Osman T. Akgun, Rhonda Righter, and Ronald Wolff. Partial flexibility in routeing and scheduling. Advances in Applied Probability, 45(3):673–691, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1377868534.

Arjas:1973:SPM

[AS73] E. Arjas and T. P. Speed. A stopping problem in Markov additive processes. Advances in Applied Probability, 5(1):2–3, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425941.

Antonelli:1977:GRDa

[AS77] Peter L. Antonelli and Curtis Strobeck. The geometry of random drift I. Stochastic distance and diffusion. Advances in Applied Probability, 9(2):238–249, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426383.

Ali:1979:CPU

[AS79] M. A. Ali and N. Singh. Control problem under uncertainty. Advances in Applied Probability, 11(2):308, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426840.

Ayala:1995:BRC

[AS95] Guillermo Ayala and Amelia Simó. Bivariate random closed sets and nerve fibre degeneration. Advances in Applied Probability, 27(2):293–305, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427826.

Ayala:1996:CBP

[AS96] Guillermo Ayala and Amelia Simó. Correlations in bivariate point processes: Some biological applications. Advances in Applied Probability, 28(2):330, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1428038.

Alparslan:2007:RPC

[AS07] Ugur Tuncay Alparslan and Gennady Samorodnitsky. Ruin probability with certain stationary stable claims generated by conservative flows. Advances in Applied Probability, 39(2):360– 384, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443518.

Alishahi:2008:VDT

[AS08] Kasra Alishahi and Mohsen Sharifitabar. Volume degeneracy of the typical cell and the chord length distribution for Poisson– Voronoi tessellations in high dimensions. Advances in Applied Probability, 40(4):919–938, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443618.

Abramowicz:2013:PMI

[AS13] Konrad Abramowicz and Oleg Seleznjev. Piecewise-multilinear interpolation of a random field. Advances in Applied Probability, 45(4):945–959, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1386857852.

Arlotto:2014:OOS

[AS14] Alessandro Arlotto and J. Michael Steele. Optimal online selection of an alternating subsequence: a central limit theorem. Advances in Applied Probability, 46(2):536–559, June 2014. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1401369706.

Aurzada:2019:SLR

[AS19] F. Aurzada and S. Schwinn. Scaling limits for a random boxes model. Advances in Applied Probability, 51 (3):773–801, September 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/scaling-limits-for-a-random-boxes-
model/C07B0C2C75C19A09B4B9B9E77DC30D61.

Assaf:1996:SPM

[ASC96] David Assaf and Ester Samuel-Cahn. The secretary problem: Minimizing the expected rank with i.i.d. random variables. Advances in Applied Probability, 28(3):828–852, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428183.

Asmussen:1982:CLT

[Asm82] Søren Asmussen. Conditioned limit theorems relating a random walk to its associate, with applications to risk reserve processes and the GI/G/1 queue. Advances in Applied Probability, 14 (1):143–170, March 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426737.

Assaf:1978:IPD

[Ass78] David Assaf. Invariant problems in discounted dynamic programming. Advances in Applied Probability, 10(2):472–490, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426946.

Asmussen:1999:TPN

[ASS99] Søren Asmussen, Hanspeter Schmidli, and Volker Schmidt. Tail probabilities for non-standard risk and queueing processes with subexponential jumps. Advances in Applied Probability, 31(2): 422–447, June 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428117.

Adler:2010:EST

[AST10] Robert J. Adler, Gennady Samorodnitsky, and Jonathan E. Taylor. Excursion sets of three classes of stable random fields.

Advances in Applied Probability, 42(2):293–318, June 2010. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/25683820.

Ahmed:1975:OCS

[AT75] N. U. Ahmed and K. L. Teo. Optimal control of stochastic Ito differential systems with fixed terminal time. Advances in Applied Probability, 7(1):154–178, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1425858.

Arnold:1980:DLS

[AT80] L. Arnold and M. Theodosopulu. Deterministic limit of the stochastic model of chemical reactions with diffusion. Advances in Applied Probability, 12(2):367–379, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426602.

Atkinson:1981:ATP

[AT81] C. Atkinson and M. E. Thompson. Analysis of two parallel queues with common exit service and infinite queue size. Advances in Applied Probability, 13(1):147–166, March 1981. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426472.

Anantharam:1990:SCT

[AT90] Venkat Anantharam and Pantelis Tsoucas. Stochastic concavity of throughput in series of queues with finite buffers. Advances in Applied Probability, 22(3):761–763, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427471.

Ahlberg:2018:GDM

[AT18] Daniel Ahlberg and Johan Tykesson. Gilbert’s disc model with geostatistical marking. Advances in Applied Probability, 50(4):1075–1094, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/gilberts-disc-model-with-geostatistical-
marking/C2FB56E7A363AD35FF64883C67D6EC8B.

Athreya:1979:MPA

[Ath79] K. B. Athreya. A Markov process approach to systems of renewal equations. Advances in Applied Probability, 11(2):275– 276, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426805.

Athreya:1985:DBR

[Ath85] K. B. Athreya. Discounted branching random walks. Advances in Applied Probability, 17(1):53–66, March 1985. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427052.

Athreya:2018:PLF

[Ath18] Krishna B. Athreya. On powers of likelihood functions of random walks on Z. Advances in Applied Probability, 50(A):63–66, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-powers-of-likelihood-functions-
of-random-walks-on/EFA04ED7CFB89EAF39840E5B20967A2D.

Attard:2017:NEN

[Att17] Natalie Attard. Nash equilibrium in nonzero-sum games of optimal stopping for Brownian motion. Advances in Applied Probability, 49(2):430–445, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/nash-equilibrium-in-nonzerosum-games-
of-optimal-stopping-for-brownian-motion/74847E03E48C42FEDA5CB0375FFA375C.

Avram:2019:FPP

[AV19] Florin Avram and Matija Vidmar. First passage problems for upwards skip-free random walks via the scale functions paradigm. Advances in Applied Probability, 51(2):408–424, June 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/first-
passage-problems-for-upwards-skipfree-random-walks-via-
the-scale-functions-paradigm/1A05FE0AFA01A28B2693C0F1ADEEBBC1.

Avery:1976:ERE

[Ave76] P. J. Avery. The effect of random environments on selection schemes in populations of a finite size. Advances in Applied Probability, 8(4):613–616, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1425916.

Avery:1979:EFP

[Ave79] P. J. Avery. The effect of finite population size on selection results for two linked loci. Advances in Applied Probability, 11(1): 1–2, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426756.

Aven:1983:ORU

[Ave83] Terje Aven. Optimal replacement under a minimal repair strategy: A general failure model. Advances in Applied Probability, 15(1):198–211, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426990.

Agrawal:2023:OSM

[AVLR23] Sanket Agrawal, Dootika Vats, Krzysztof Latuszyński, and Gareth O. Roberts. Optimal scaling of MCMC beyond Metropolis. Advances in Applied Probability, 55(2):492–509, June 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/optimal-scaling-
of-mcmc-beyond-metropolis/2E5767C32506E5F1860E59B3BECECF59.

Ahlswede:1971:CS

[AW71] R. Ahlswede and J. Wolfowitz. Channels without synchronization. Advances in Applied Probability, 3(2):383–403, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426177.

Albright:1979:BDM

[AW79] S. Christian Albright and Wayne Winston. A birth–death model of advertising and pricing. Advances in Applied Probability, 11 (1):134–152, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426772.

Abate:1987:TBRa

[AW87a] Joseph Abate and Ward Whitt. Transient behavior of regulated Brownian motion, i: Starting at the origin. Advances in Applied Probability, 19(3):560–598, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427408.

Abate:1987:TBRb

[AW87b] Joseph Abate and Ward Whitt. Transient behavior of regulated Brownian motion, II: Non-zero initial conditions. Advances in Applied Probability, 19(3):599–631, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427409.

Abate:1988:TBM

[AW88] Joseph Abate and Ward Whitt. Transient behavior of the M/M/1 queue via Laplace transforms. Advances in Applied Probability, 20(1):145–178, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427274.

Abate:1996:OCP

[AW96] Joseph Abate and Ward Whitt. An operational calculus for probability distributions via Laplace transforms. Advances in Applied Probability, 28(1):75–113, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427914.

Adan:1993:CAT

[AWZ93] I. J. B. F. Adan, J. Wessels, and W. H. M. Zijm. A compensation approach for two-dimensional Markov processes. Advances in Applied Probability, 25(4):783–817, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427792.

Aristoff:2018:PTC

[AZ18] David Aristoff and Lingjiong Zhu. On the phase transition curve in a directed exponential random graph model. Advances in Applied Probability, 50(1):272–301, March 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/on-the-phase-transition-
curve-in-a-directed-exponential-random-graph-model/7897E4AE59B96358DDF252D3829294C9.

Bhat:1981:MLE

[BA81] B. R. Bhat and S. R. Adke. Maximum likelihood estimation for branching processes with immigration. Advances in Applied Probability, 13(3):498–509, September 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426782.

Bachmann:2000:LTM

[Bac00] Markus Bachmann. Limit theorems for the minimal position in a branching random walk with independent logconcave displacements. Advances in Applied Probability, 32(1):159–176, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428232.

Baddeley:1977:FNR

[Bad77a] Adrian Baddeley. A fourth note on recent research in geometrical probability. Advances in Applied Probability, 9(4):824–860, December 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426702.

Baddeley:1977:IMM

[Bad77b] Adrian Baddeley. Integrals on a moving manifold and geometrical probability. Advances in Applied Probability, 9(3):588–603, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426116.

Baddeley:1980:LTS

[Bad80] Adrian Baddeley. A limit theorem for statistics of spatial data. Advances in Applied Probability, 12(2):447–461, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426605.

Balkema:1975:IPW

[Bal75] A. A. Balkema. Invariance properties for weak limits under monotone transformations. Advances in Applied Probability, 7 (2):252–253, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426067.

Balkema:1977:NSP

[Bal77] A. A. Balkema. Norming sequences for partial maxima. Advances in Applied Probability, 9(2):190, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426337.

Balder:1978:EUM

[Bal78] E. J. Balder. An extension of the usual model in statistical decision theory with applications to stochastic optimization problems. Advances in Applied Probability, 10(2):325–326, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426924.

Balagopal:1979:QDR

[Bal79] K. Balagopal. On queues in discrete regenerative environments, with application to the second of two queues in series. Advances in Applied Probability, 11(4):851–869, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426863.

Ball:1985:DSE

[Bal85] Frank Ball. Deterministic and stochastic epidemics with several kinds of susceptibles. Advances in Applied Probability, 17(1): 1–22, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427049.

Ball:1986:UAD

[Bal86] Frank Ball. A unified approach to the distribution of total size and total area under the trajectory of infectives in epidemic models. Advances in Applied Probability, 18(2):289–310, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427301.

Ball:1990:AMP

[Bal90] Frank Ball. Aggregated Markov processes with negative exponential time interval omission. Advances in Applied Probability, 22(4):802–830, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427563.

Ball:1997:ECB

[Bal97] Frank Ball. Empirical clustering of bursts of openings in Markov and semi-Markov models of single channel gating incorporating time interval omission. Advances in Applied Probability, 29(4): 909–946, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427847.

Ballani:2007:SPC

[Bal07] Felix Ballani. The surface pair correlation function for stationary Boolean models. Advances in Applied Probability, 39(1): 1–15, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443497.

Bansaye:2009:CCB

[Ban09] Vincent Bansaye. Cell contamination and branching processes in a random environment with immigration. Advances in Applied Probability, 41(4):1059–1081, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/27793919.

Barbour:1974:FCL

[Bar74a] Andrew D. Barbour. On a functional central limit theorem for Markov population processes. Advances in Applied Probability, 6(1):21–39, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426205.

Bartlett:1974:SASa

[Bar74b] M. S. Bartlett. The statistical analysis of spatial pattern. Advances in Applied Probability, 6(2):188–190, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426240.

Bartlett:1974:SASb

[Bar74c] M. S. Bartlett. The statistical analysis of spatial pattern. Advances in Applied Probability, 6(2):336–358, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426297.

Barbour:1975:ABB

[Bar75] Andrew D. Barbour. The asymptotic behaviour of birth and death and some related processes. Advances in Applied Probability, 7(1):28–43, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1425851.

Barbour:1976:NQM

[Bar76a] Andrew D. Barbour. Networks of queues and the method of stages. Advances in Applied Probability, 8(3):584–591, September 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426145.

Barbour:1976:QSD

[Bar76b] Andrew D. Barbour. Quasi-stationary distributions in Markov population processes. Advances in Applied Probability, 8(2):296– 314, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425906.

Bartholomew:1977:MGA

[Bar77] D. J. Bartholomew. Maintaining a grade or age structure in a stochastic environment. Advances in Applied Probability, 9(1): 1–17, March 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425813.

Barlow:1978:HTS

[Bar78a] Martin Barlow. Honest times and splitting times. Advances in Applied Probability, 10(2):280, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426884.

Barlow:1978:RDR

[Bar78b] Richard E. Barlow. Recent developments in reliability theory. Advances in Applied Probability, 10(2):261–262, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426870.

Bartlett:1978:NRW

[Bar78c] M. S. Bartlett. A note on random walks at constant speed. Advances in Applied Probability, 10(4):704–707, December 1978.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426647.

Bartholomew:1979:CGS

[Bar79] D. J. Bartholomew. The control of a grade structure in a stochastic environment using promotion control. Advances in Applied Probability, 11(3):603–615, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426957.

Barbour:1980:EDM

[Bar80] A. D. Barbour. Equilibrium distributions Markov population processes. Advances in Applied Probability, 12(3):591–614, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426422.

Barbour:1981:BMS

[Bar81] A. D. Barbour. Brownian motion and a sharply curved boundary. Advances in Applied Probability, 13(4):736–750, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426970.

Bartlett:1984:RII

[Bar84] M. S. Bartlett. Recent investigations involving stochastic population models. Advances in Applied Probability, 16(3):449–470, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427284.

Bartoszewicz:1985:DOM

[Bar85a] J. Bartoszewicz. Dispersive ordering and monotone failure rate distributions. Advances in Applied Probability, 17(2):472– 474, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427155.

Bartoszynski:1985:MPC

[Bar85b] Robert Bartoszyński. On metastatic progression of cancer. Advances in Applied Probability, 17(2):245–246, June 1985. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427131.

Barral:2001:GVM

[Bar01] Julien Barral. Generalized vector multiplicative cascades. Advances in Applied Probability, 33(4):874–895, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428310.

Barron:2018:TPM

[Bar18] Yonit Barron. A threshold policy in a Markov-modulated production system with server vacation: the case of continuous and batch supplies. Advances in Applied Probability, 50 (4):1246–1274, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/threshold-policy-in-a-markovmodulated-
production-system-with-server-vacation-the-case-of-continuous-
and-batch-supplies/96B84C6B29BA806FC80C2E63AF8AF8E4.

Basawa:1979:ATI

[Bas79] I. V. Basawa. Asymptotic theory of inference for a certain class of stochastic processes. Advances in Applied Probability, 11(2):300–301, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426831.

Bather:1970:OSP

[Bat70] John Bather. Optimal stopping problems for Brownian motion. Advances in Applied Probability, 2(2):259–286, Autumn 1970. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426320.

Bather:1973:ODPa

[Bat73a] John Bather. Optimal decision procedures for finite Markov chains. Part I: Examples. Advances in Applied Probability, 5 (2):328–339, August 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426039.

Bather:1973:ODPb

[Bat73b] John Bather. Optimal decision procedures for finite Markov chains. Part II: Communicating systems. Advances in Applied Probability, 5(3):521–540, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1425832.

Bather:1973:ODPc

[Bat73c] John Bather. Optimal decision procedures for finite Markov chains. Part III: General convex systems. Advances in Applied Probability, 5(3):541–553, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1425833.

Bather:1976:CCM

[Bat76a] J. A. Bather. A control chart model and a generalised stopping problem for Brownian motion. Advances in Applied Probability, 8(2):213–215, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425865.

Bather:1976:OSP

[Bat76b] John Bather. Optimal stationary policies for denumerable Markov chains in continuous time. Advances in Applied Probability, 8(1):144–158, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426026.

Bather:1978:SSM

[Bat78] J. A. Bather. Sufficient statistics for Markov decision processes. Advances in Applied Probability, 10(2):307–308, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426907.

Bather:1980:RAT

[Bat80] John Bather. Randomised allocation of treatments in sequential trials. Advances in Applied Probability, 12(1):174–182, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426500.

Bauerle:2002:OCQ

[Bäu02] Nicole Bäuerle. Optimal control of queueing networks: An approach via fluid models. Advances in Applied Probability, 34 (2):313–328, June 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428290.

Bauernschubert:2014:RTC

[Bau14] Elisabeth Bauernschubert. Recurrence and transience of critical branching processes in random environment with immigration and an application to excited random walks. Advances in Applied Probability, 46(3):687–703, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1409319555.

Bazovsky:1974:GMS

[Baz74] Igor Bazovsky, Jr. On generalizations of the method of stages to ageing and learning phenomena. Advances in Applied Probability, 6(2):215–216, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426253.

Bandyopadhyay:1990:GML

[BB90] Dipankar Bandyopadhyay and Asit P. Basu. On a generalization of a model by Lindley and singpurwalla. Advances in Applied Probability, 22(2):498–500, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427552.

Ball:1992:FFW

[BB92] Frank Ball and Paul Blackwell. A finite form for the wrapped Poisson distribution. Advances in Applied Probability, 24(1): 221–222, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427738.

Baccelli:1993:VCS

[BB93a] F. Baccelli and P. Brémaud. Virtual customers in sensitivity and light traffic analysis via Campbell’s formula for point processes. Advances in Applied Probability, 25(1):221–234, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427503.

Biggins:1993:LDS

[BB93b] J. D. Biggins and N. H. Bingham. Large deviations in the supercritical branching process. Advances in Applied Probability, 25(4):757–772, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427790.

Bollobas:1995:RHD

[BB95] Béla Bollobás and Graham Brightwell. Random highdimensional orders. Advances in Applied Probability, 27(1):

161–184, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428102.

Burger:1996:SDU

[BB96] Reinhard Bürger and Immanuel M. Bomze. Stationary distributions under mutation-selection balance: Structure and properties. Advances in Applied Probability, 28(1):227–251, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427919.

Baccelli:2001:CPR

[BB01] François Baccelli and Bartlomiej Blaszczyszyn. On a coverage process ranging from the Boolean model to the Poisson–Voronoi tessellation with applications to wireless communications. Advances in Applied Probability, 33(2):293–323, June 2001. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428254.

Ball:2007:EMI

[BB07a] Frank Ball and Tom Britton. An epidemic model with infectordependent severity. Advances in Applied Probability, 39(4):949– 972, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443553.

Barrieu:2007:OHT

[BB07b] P. Barrieu and N. Bellamy. Optimal hitting time and perpetual option in a Non-Lévy model: Application to real options. Advances in Applied Probability, 39(2):510–530, June 2007. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443524.

Brandt:2010:IBM

[BB10] Andreas Brandt and Manfred Brandt. Insensitive bounds for the moments of the sojourn times in M/Gi systems under statedependent processor sharing. Advances in Applied Probability, 42(1):246–267, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/25683815.

Baur:2017:WLL

[BB17] Erich Baur and Jean Bertoin. Weak limits for the largest subpopulations in Yule processes with high mutation probabilities. Advances in Applied Probability, 49(3):877–902, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/weak-
limits-for-the-largest-subpopulations-in-yule-processes-
with-high-mutation-probabilities/45730FA963C0400A9C73FFD7CD1BAC23.

Bartholomew:1976:CPP

[BBB^{+}76] D. J. Bartholomew, R. J. Bennett, Julian Besag, Don McNeil, Andrew D. Cliff, Keith Ord, Pierre Delfiner, Peter J. Diggle, R. F. Galbraith, D. Walley, S. Gregory, R. P. Haining, R. J. Johnston, R. Mead, Denis Mollison, S. Openshaw, Robin Sibson, R. W. Thomas, John Whitelegg, and Neil Wrigley. Conference on patterns and processes in the plane. Advances in Applied Probability, 8(4):651–658, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1425928.

Bertin:1999:ENN

[BBD99] Etienne Bertin, Jean-Michel Billiot, and Remy Drouilhet. Existence of ‘nearest-neighbour’ spatial Gibbs models. Advances in Applied Probability, 31(4):895–909, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428333.

Bertin:2002:CPG

[BBD02] Etienne Bertin, Jean-Michel Billiot, and Rémy Drouilhet. Continuum percolation in the Gabriel graph. Advances in Applied Probability, 34(4):689–701, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428317.

Baccelli:1984:SSQ

[BBH84] F. Baccelli, P. Boyer, and G. Hebuterne. Single-server queues with impatient customers. Advances in Applied Probability, 16 (4):887–905, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427345.

Baccelli:2011:OPS

[BBHM11] François Baccelli, Bartlomiej Blaszczyszyn, and Mir-Omid Haji-Mirsadeghi. Optimal paths on the space-time SINR random graph. Advances in Applied Probability, 43(1):131–150, March 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/23024529.

Balkema:1988:CP

[BBM^{+}88] A. A. Balkema, R. Ballerini, W. P. McCormick, J. Beirlant, H. A. Brozius, J. P. Cohen, P. Deheuvels, Z. Govindarajulu, P. Purdue, L. De Haan, I. Weissman, T. Hsing, A. Janssen, R.- D. Reiss, D. M. Mason, D. Pfeifer, and S. Yakowitz. Contributed papers. Advances in Applied Probability, 20(1):8–13, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427267.

Bekker:2009:LPA

[BBR09] R. Bekker, O. J. Boxma, and J. A. C. Resing. Lévy processes with adaptable exponent. Advances in Applied Probability, 41 (1):177–205, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/27793869.

Badila:2014:QRM

[BBRW14] E. S. Badila, O. J. Boxma, J. A. C. Resing, and E. M. M. Winands. Queues and risk models with simultaneous arrivals. Advances in Applied Probability, 46(3):812–831, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1409319561.

Bachmat:2007:OBP

[BBSS07] Eitan Bachmat, Daniel Berend, Luba Sapir, and Steven Skiena. Optimal boarding policies for thin passengers. Advances in Applied Probability, 39(4):1098–1114, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443560.

Balister:2005:CRN

[BBSW05] Paul Balister, Béla Bollobás, Amites Sarkar, and Mark Walters. Connectivity of random k-nearest-neighbour graphs. Advances in Applied Probability, 37(1):1–24, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037313.

Balister:2009:CCN

[BBSW09] Paul Balister, Béla Bollobás, Amites Sarkar, and Mark Walters. A critical constant for the k-nearest-neighbour model. Advances in Applied Probability, 41(1):1–12, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/27793862.

Balister:2010:SSW

[BBSW10] Paul Balister, Béla Bollobás, Amites Sarkar, and Mark Walters. Sentry selection in wireless networks. Advances in Applied Probability, 42(1):1–25, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/25683804.

Balister:2009:RTN

[BBW09] Paul Balister, Béla Bollobás, and Mark Walters. Random transceiver networks. Advances in Applied Probability, 41(2): 323–343, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/27793880.

Belhadji:2010:SRP

[BBZ10] Lamia Belhadji, Daniela Bertacchi, and Fabio Zucca. A selfregulating and patch subdivided population. Advances in Applied Probability, 42(3):899–912, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/25799073.

Bishop:1976:MAC

[BC76] D. T. Bishop and C. Cannings. Models of animal conflict. Advances in Applied Probability, 8(4):616–621, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425917.

Bishop:1978:GWA

[BC78] D. T. Bishop and C. Cannings. The generalised war of attrition. Advances in Applied Probability, 10(1):6–7, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426708.

Bishop:1981:WAR

[BC81a] D. T. Bishop and C. Cannings. The war of attrition with random rewards. II. Advances in Applied Probability, 13(1):1, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426451.

Bourgin:1981:DAP

[BC81b] Richard D. Bourgin and Robert Cogburn. On determining absorption probabilities for Markov chains in random environments. Advances in Applied Probability, 13(2):369–387, June 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426689.

Biggins:1987:MRP

[BC87] J. D. Biggins and C. Cannings. Markov renewal processes, counters and repeated sequences in Markov chains. Advances in Applied Probability, 19(3):521–545, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427406.

Beuerman:1989:SSE

[BC89a] Steven L. Beuerman and Edward J. Coyle. State space expansions and the limiting behavior of quasi-birth-and-death processes. Advances in Applied Probability, 21(2):284–314, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427161.

Briere:1989:CAS

[BC89b] G. Brière and M. L. Chaudhry. Computational analysis of single-server bulk-service queues, M/G^{Y} /1. Advances in Applied Probability, 21(1):207–225, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427204.

Ball:1993:FSS

[BC93] Frank Ball and Damian Clancy. The final size and severity of a generalised stochastic multitype epidemic model. Advances in Applied Probability, 25(4):721–736, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427788.

Baillo:2001:ESS

[BC01] Amparo Baíllo and Antonio Cuevas. On the estimation of a star-shaped set. Advances in Applied Probability, 33(4):717–726, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428300.

Buss:2004:SFW

[BC04] Samuel R. Buss and Peter Clote. Solving the Fisher–Wright and coalescence problems with a discrete Markov chain analysis. Advances in Applied Probability, 36(4):1175–1197, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/4140393.

Burridge:2016:PTH

[BC16] James Burridge and Richard Cowan. Planar tessellations that have the half-Gilbert structure. Advances in Applied Probability, 48(2):574–584, June 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1465490763.

Baker:2018:PSN

[BCHK18] J. Baker, P. Chigansky, K. Hamza, and F. C. Klebaner. Persistence of small noise and random initial conditions. Advances in Applied Probability, 50(A):67–81, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/persistence-
of-small-noise-and-random-initial-conditions/98550294B3C69342EFAF4FA75FF37B69.

Beskos:2017:SPF

[BCJ^{+}17] Alexandros Beskos, Dan Crisan, Ajay Jasra, Kengo Kamatani, and Yan Zhou. A stable particle filter for a class of highdimensional state-space models. Advances in Applied Probability, 49(1):24–48, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/stable-particle-filter-for-a-class-
of-highdimensional-statespace-models/904D64E95E2E5CCCB4AFE95E00EDD4FA.

Beskos:2014:EBN

[BCJW14] Alexandros Beskos, Dan O. Crisan, Ajay Jasra, and Nick Whiteley. Error bounds and normalising constants for sequential Monte Carlo samplers in high dimensions. Advances in Applied Probability, 46(1):279–306, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1396360114.

Bassetti:2010:CID

[BCL10] Federico Bassetti, Irene Crimaldi, and Fabrizio Leisen. Conditionally identically distributed species sampling sequences. Advances in Applied Probability, 42(2):433–459, June 2010. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/25683828.

Baccelli:2016:STP

[BCLF16] François Baccelli, Héctor A. Chang-Lara, and Sergey Foss. Shape theorems for Poisson hail on a bivariate ground. Advances in Applied Probability, 48(2):525–543, June 2016. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1465490761.

Burridge:2013:FHG

[BCM13] James Burridge, Richard Cowan, and Isaac Ma. Full- and half-Gilbert tessellations with rectangular cells. Advances in Applied Probability, 45(1):1–19, March 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/41805997.

Bertoin:2018:BSP

[BCM18] Jean Bertoin, Aser Cortines, and Bastien Mallein. Branchingstable point measures and processes. Advances in Applied Probability, 50(4):1294–1314, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/branchingstable-point-measures-and-
processes/976F4773C475ECBC2E52167B1D369863.

Bartlett:1984:KAD

[BÇNP84] M. S. Bartlett, Erhan Çinlar, M. F. Neuts, and P. Purdue. Keynote addresses and discussion. Advances in Applied Probability, 16(1):4–6, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427216.

Baddeley:1995:RBT

[BCO95] A. J. Baddeley and L. M. Cruz-Orive. The Rao–Blackwell theorem in stereology and some counterexamples. Advances in Applied Probability, 27(1):2–19, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428091.

Baurdoux:2015:ODS

[BCSY15] Erik J. Baurdoux, Nan Chen, Budhi A. Surya, and Kazutoshi Yamazaki. Optimal double stopping of a Brownian bridge. Advances in Applied Probability, 47(4):1212–1234, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1449859807.

Baccelli:2013:SIP

[BCT13] François Baccelli, David Coupier, and Viet Chi Tran. Semiinfinite paths of the two-dimensional radial spanning tree. Advances in Applied Probability, 45(4):895–916, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1386857849.

Balkema:1973:LLN

[BD73] A. A. Balkema and L. De Haan. Limit laws for f(X_{n}), f nondecreasing. Advances in Applied Probability, 5(1):4–5, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425942.

Bingham:1974:APSb

[BD74a] N. H. Bingham and R. A. Doney. Asymptotic properties of supercritical branching processes i: The Galton–Watson process. Advances in Applied Probability, 6(4):711–731, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426188.

Bingham:1974:APSa

[BD74b] N. H. Bingham and R. A. Doney. Asymptotic properties of the supercritical Galton–Watson process. Advances in Applied Probability, 6(2):229–230, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426267.

Bingham:1975:APS

[BD75] N. H. Bingham and R. A. Doney. Asymptotic properties of super-critical branching processes II: Crump-mode and Jirina processes. Advances in Applied Probability, 7(1):66–82, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425854.

Ball:1987:ICD

[BD87] Frank Ball and Peter Donnelly. Interparticle correlation in death processes with application to variability in compartmental models. Advances in Applied Probability, 19(4):755–766, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427100.

Brown:1993:CII

[BD93] Timothy C. Brown and Peter Donnelly. On conditional intensities and on interparticle correlation in non-linear death processes. Advances in Applied Probability, 25(1):255–260, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427505.

Bertoin:1996:SAR

[BD96] J. Bertoin and R. A. Doney. Some asymptotic results for transient random walks. Advances in Applied Probability, 28(1):207– 226, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427918.

Ball:1997:CBO

[BD97] Frank Ball and Sue Davies. Clustering of bursts of openings in Markov and semi-Markov models of single channel gating. Advances in Applied Probability, 29(1):92–113, March 1997. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427862.

Blanchet:2015:PSI

[BD15] Jose Blanchet and Jing Dong. Perfect sampling for infinite server and loss systems. Advances in Applied Probability, 47(3): 761–786, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1444308881.

Bierkens:2017:LTZ

[BD17] Joris Bierkens and Andrew Duncan. Limit theorems for the zig-zag process. Advances in Applied Probability, 49 (3):791–825, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/limit-theorems-for-the-zigzag-process/
5DF88453667E7E2A8F48183EC5EDCE92.

Boniece:2022:OFL

[BD22] B. Cooper Boniece and Gustavo Didier. On operator fractional Lévy motion: integral representations and time-reversibility. Advances in Applied Probability, 54(2):493–535, June 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/on-operator-
fractional-levy-motion-integral-representations-and-
timereversibility/258A1C6CA4D25D5B2C36CD1F435FA651.

Brandejsky:2012:NME

[BDD12] Adrien Brandejsky, Benoîte De Saporta, and François Dufour. Numerical methods for the exit time of a piecewise-deterministic Markov process. Advances in Applied Probability, 44(1):196– 225, March 2012. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
23207861.

Benes:1984:SC

[BDF^{+}84] V. E. Benes, M. H. A. Davis, Wendell H. Fleming, Ioannis Karatzas, Steven E. Shreve, C. A. Glasbey, M. B. McGechan, Gary Gottlieb, Benny Levikson, A. Hordijk, F. A. Van Der Duyn Schouten, and Domokos Vermes. Stochastic control. Advances in Applied Probability, 16(1):14–16, March 1984. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427219.

Bezborodov:2018:ASS

[BDK^{+}18] Viktor Bezborodov, Luca Di Persio, Tyll Krueger, Mykola Lebid, and Tomasz Ozański. Asymptotic shape and the speed of propagation of continuous-time continuous-space birth processes. Advances in Applied Probability, 50(1):74–101, March 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/

journals/advances-in-applied-probability/article/asymptotic-
shape-and-the-speed-of-propagation-of-continuoustime-
continuousspace-birth-processes/3232E000771D08A0CF2A9566970A137C.

See corrigendum [Bez20].

Barczy:2014:SEA

[BDLP14] Mátyás Barczy, Leif Döring, Zenghu Li, and Gyula Pap. Stationarity and ergodicity for an affine two-factor model. Advances in Applied Probability, 46(3):878–898, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1409319564.

Baras:1985:TCQ

[BDM85] J. S. Baras, A. J. Dorsey, and A. M. Makowski. Two competing queues with linear costs and geometric service requirements: The μc-rule is often optimal. Advances in Applied Probability, 17(1):186–209, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427059.

Beck:1999:UAF

[BDM99] B. Beck, A. R. Dabrowski, and D. R. McDonald. A unified approach to fast Teller queues and ATM. Advances in Applied Probability, 31(3):758–787, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428395.

Barbour:1983:PAS

[BE83] A. D. Barbour and G. K. Eagleson. Poisson approximation for some statistics based on exchangeable trials. Advances in Applied Probability, 15(3):585–600, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426620.

Barbour:1985:MCS

[BE85] A. D. Barbour and G. K. Eagleson. Multiple comparisons and sums of dissociated random variables. Advances in Applied Probability, 17(1):147–162, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427057.

Barnsley:1988:NCM

[BE88] Michael F. Barnsley and John H. Elton. A new class of Markov processes for image encoding. Advances in Applied Probability, 20(1):14–32, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427268.

Bierme:2006:PRB

[BE06] Hermine Biermé and Anne Estrade. Poisson random balls: Selfsimilarity and X-ray images. Advances in Applied Probability, 38(4):853–872, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/20443478.

Bartoszek:2021:NAM

[BE21] Krzysztof Bartoszek and Torkel Erhardsson. Normal approximation for mixtures of normal distributions and the evolution of phenotypic traits. Advances in Applied Probability, 53(1):162–188, March 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/normal-approximation-for-mixtures-
of-normal-distributions-and-the-evolution-of-phenotypic-
traits/7F91127E3F243605FA24051202B6CC08.

Beghin:2012:FRE

[Beg12] L. Beghin. Fractional relaxation equations and Brownian crossing probabilities of a random boundary. Advances in Applied Probability, 44(2):479–505, June 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/41714060.

Behme:2011:DPS

[Beh11] Anita Diana Behme. Distributional properties of solutions of dv_{T} = V_{T} − du_{T} + dl_{T} with Lévy noise. Advances in Applied Probability, 43(3):688–711, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/23024571.

Boxma:2016:QII

[BEJ16] Onno Boxma, Rim Essifi, and Augustus J. E. M. Janssen. A queueing/inventory and an insurance risk model. Advances in Applied Probability, 48(4):1139–1160, December 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1482548432.

Benes:1974:CIM

[Ben74] V. E. Benes. Composition and invariance methods for solving optimal control problems. Advances in Applied Probability, 6 (2):190–191, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426241.

Benes:1975:CIM

[Ben75a] V. E. Benes. Composition and invariance methods for solving some stochastic control problems. Advances in Applied Probability, 7(2):299–329, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426079.

Bengtsson:1975:IBC

[Ben75b] Bengt Olle Bengtsson. The increase of balanced chromosome mutations in mammalian populations. Advances in Applied Probability, 7(1):1–3, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425840.

Bengtsson:1980:HES

[Ben80] Bengt O. Bengtsson. How evolutionarily stable is the XX/XY sex-determination system? Advances in Applied Probability, 12(1):1–3, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426479.

Benhenni:1991:MSA

[Ben91] Robert A. Benhenni. Monotone stopping-allocation problems. Advances in Applied Probability, 23(1):24–45, March 1991. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427510.

Benes:1995:SOF

[Ben95] Viktor Benes. On second-order formulas in anisotropic stereology. Advances in Applied Probability, 27(2):326–343, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427828.

Benes:1998:MPS

[Ben98] Viktor Benes. Marginal problems in stereology. Advances in Applied Probability, 30(2):273–274, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427943.

Boland:1992:SOR

[BENP92] Philip J. Boland, Emad El-Neweihi, and Frank Proschan. Stochastic order for redundancy allocations in series and parallel systems. Advances in Applied Probability, 24(1):161–171, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427734.

Berbee:1977:LTS

[Ber77] H. C. P. Berbee. A limit theorem for the superposition of renewal processes. Advances in Applied Probability, 9(2):235–236, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426380.

Berbee:1978:RRW

[Ber78a] H. C. P. Berbee. Recurrence for random walks with stationary increments. Advances in Applied Probability, 10(2):293– 294, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426897.

Bergman:1978:ORU

[Ber78b] Bo Bergman. Optimal replacement under a general failure model. Advances in Applied Probability, 10(2):431–451, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426944.

Bergmann:1978:EMW

[Ber78c] R. Bergmann. On the estimation of the mean waiting time and bounds for waiting time parameters in GI/G/1. Advances in Applied Probability, 10(2):313–314, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426913.

Berman:1978:RMPa

[Ber78d] Mark Berman. Regenerative multivariate point processes. Advances in Applied Probability, 10(2):301, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426903.

Berman:1978:RMPb

[Ber78e] Mark Berman. Regenerative multivariate point processes. Advances in Applied Probability, 10(2):411–430, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426943.

Berger:1980:SDI

[Ber80] Marc A. Berger. Stochastic derivatives and integral equations. Advances in Applied Probability, 12(2):289–290, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426562.

Berman:1982:SED

[Ber82] Simeon M. Berman. Sojourns and extremes of a diffusion process on a fixed interval. Advances in Applied Probability, 14(4): 811–832, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427025.

Berman:1985:SPR

[Ber85] Mark Berman. A simple proof of a result of iglehart and shedler. Advances in Applied Probability, 17(1):239–241, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427065.

Berman:1987:PEV

[Ber87] Simeon M. Berman. Poisson and extreme value limit theorems for Markov random fields. Advances in Applied Probability, 19 (1):106–122, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427375.

Berzunza:2020:EGC

[Ber20] G. Berzunza. The existence of a giant cluster for percolation on large Crump–Mode–Jagers trees. Advances in Applied Probability, 52(1):266–290, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/existence-of-a-giant-cluster-for-
percolation-on-large-crumpmodejagers-trees/51F2F419B69E2A2B5BF64F7C7384BB97.

Betensky:1998:BCP

[Bet98] Rebecca A. Betensky. A boundary crossing probability for the Bessel process. Advances in Applied Probability, 30(3):807–830, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428152.

Beutler:1983:NMP

[Beu83] Frederick J. Beutler. A note on multivariate Poisson flows on stochastic processes. Advances in Applied Probability, 15(1): 219–220, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426994.

Bezborodov:2020:CAS

[Bez20] V. Bezborodov. Corrigendum to: Asymptotic shape and the speed of propagation of continuous-time continuous-space birth processes. Advances in Applied Probability, 52(4): 1325–1327, December 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/corrigendum-to-asymptotic-shape-
and-the-speed-of-propagation-of-continuoustime-continuousspace-
birth-processes/C802B13E5C9E8383DF94B609A6D3ABDE. See

[BDK^{+}18].

Berry:1980:TABa

[BF80a] Donald A. Berry and Bert Fristedt. Two-armed bandits with a goal, I. One arm known. Advances in Applied Probability, 12(3): 775–798, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426431.

Berry:1980:TABb

[BF80b] Donald A. Berry and Bert Fristedt. Two-armed bandits with a goal, II. Dependent arms. Advances in Applied Probability, 12 (4):958–971, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426751.

Blum:2005:MCS

[BF05] Michael G. B. Blum and Olivier François. Minimal clade size and external branch length under the neutral coalescent. Advances in Applied Probability, 37(3):647–662, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/30037348.

Budhiraja:2019:DAL

[BF19] Amarjit Budhiraja and Eric Friedlander. Diffusion approximations for load balancing mechanisms in cloud storage systems. Advances in Applied Probability, 51(1):41–86, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/diffusion-approximations-
for-load-balancing-mechanisms-in-cloud-storage-systems/
3687FDE0E99E191EE32337EF27B5D4AF.

Borovkov:1992:TPM

[BFK92] A. A. Borovkov, G. Fayolle, and D. A. Korshunov. Transient phenomena for Markov chains and applications. Advances in Applied Probability, 24(2):322–342, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427695.

Barron:2004:ARS

[BFL04] Yonit Barron, Esther Frostig, and Benny Levikson. Analysis of R-out-of-N repairable systems: The case of phase-type distributions. Advances in Applied Probability, 36(1):116–138, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428356.

Brachner:2009:EAT

[BFL09] Claudia Brachner, Vicky Fasen, and Alexander Lindner. Extremes of autoregressive threshold processes. Advances in Applied Probability, 41(2):428–451, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/27793885.

Bartoszynski:1984:GMB

[BFM^{+}84] Robert Bartoszyński, M. J. Faddy, P. D. M. MacDonald, S. F. L. Gallot, James H. King, Joshua Chover, Israel David, and Uri Yechiali. General models in biology. Advances in Applied Probability, 16(1):23–25, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427221.

Benaim:1998:COD

[BFP98] Michel Benaïm, Jean-Claude Fort, and Gilles Pagès. Convergence of the one-dimensional Kohonen algorithm. Advances in Applied Probability, 30(3):850–869, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428154.

Boxma:2023:DRM

[BFP23] Onno Boxma, Esther Frostig, and Zbigniew Palmowski. A dual risk model with additive and proportional gains: ruin probability and dividends. Advances in Applied Probability, 55(2):549–580, June 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/dual-risk-model-with-additive-and-
proportional-gains-ruin-probability-and-dividends/761120EC5B3E6F06E5841975B83C7156.

Barany:2010:IVI

[BFV10] I. Bárány, F. Fodor, and V. Vígh. Intrinsic volumes of inscribed random polytopes in smooth convex bodies. Advances in Applied Probability, 42(3):605–619, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/25799060.

Budhiraja:2022:AQS

[BFW22] Amarjit Budhiraja, Nicolas Fraiman, and Adam Waterbury. Asymptotics of quasi-stationary distributions of small noise stochastic dynamical systems in unbounded domains. Advances in Applied Probability, 54(1):64–110, March 2022. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/asymptotics-
of-quasistationary-distributions-of-small-noise-stochastic-
dynamical-systems-in-unbounded-domains/5A3D95567E30C7294CA029FE6CF346D0.

Boshuizen:1993:GOS

[BG93] Frans A. Boshuizen and José M. Gouweleeuw. General optimal stopping theorems for semi-Markov processes. Advances in Applied Probability, 25(4):825–846, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427794.

Blanchet:2007:URT

[BG07] J. Blanchet and P. Glynn. Uniform renewal theory with applications to expansions of random geometric sums. Advances in Applied Probability, 39(4):1070–1097, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443559.

Bauerle:2008:MRP

[BG08] Nicole Bäuerle and Rudolf Grübel. Multivariate risk processes with interacting intensities. Advances in Applied Probability, 40 (2):578–601, June 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443596.

Berman:1988:IP

[BGH^{+}88] S. M. Berman, J. Galambos, C. C. Heyde, M. R. Leadbetter, E. Parzen, S. I. Resnick, H. Taylor, J. Tiago De Oliveira, and I. Weissman. Invited papers. Advances in Applied Probability, 20(1):2–8, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427266.

Briand:2020:MRS

[BGL20] Phillippe Briand, Abir Ghannoum, and Céline Labart. Mean reflected stochastic differential equations with jumps. Advances in Applied Probability, 52(2):523–562, June 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/mean-reflected-stochastic-
differential-equations-with-jumps/302A9ACEF1E376490D8D23A6B4DA8637.

Busic:2013:SBM

[BGM13] Ana Bušić, Varun Gupta, and Jean Mairesse. Stability of the bipartite matching model. Advances in Applied Probability, 45 (2):351–378, June 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1370870122.

Bronshtein:1987:ONC

[BGPS87] O. Bronshtein, I. Gertsbakh, B. Pittel, and S. Shahaf. One-node closed multichannel service system: Several types of customers and service rates, and random pick-up from the waiting line.

Advances in Applied Probability, 19(2):487–504, June 1987. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427429.

Brockwell:1982:BIC

[BGR82] P. J. Brockwell, J. Gani, and S. I. Resnick. Birth, immigration and catastrophe processes. Advances in Applied Probability, 14 (4):709–731, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427020.

Bordenave:2006:DLM

[BGR06] Charles Bordenave, Yann Gousseau, and François Roueff. The dead leaves model: A general tessellation modeling occlusion. Advances in Applied Probability, 38(1):31–46, March 2006. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443426.

Bartroff:2010:OAC

[BGRSC10] Jay Bartroff, Larry Goldstein, Yosef Rinott, and Ester Samuel-Cahn. On optimal allocation of a continuous resource using an iterative approach and total positivity. Advances in Applied Probability, 42(3):795–815, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/25799068.

Block:1989:SSF

[BGS89] Henry W. Block, William S. Griffith, and Thomas H. Savits. Lsuperadditive structure functions. Advances in Applied Probability, 21(4):919–929, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427774.

Bagchi:2022:SPN

[BGV22] Amitabha Bagchi, Iqra Altaf Gillani, and Pooja Vyavahare. A stochastic process on a network with connections to Laplacian systems of equations. Advances in Applied Probability, 54(1):254–278, March 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/stochastic-process-on-a-network-
with-connections-to-laplacian-systems-of-equations/BFC44A295068B74529DBCF81CDEEDB55.

Bean:1995:AAS

[BGZ95] N. G. Bean, R. J. Gibbens, and S. Zachary. Asymptotic analysis of single resource loss systems in heavy traffic, with applications to integrated networks. Advances in Applied Probability, 27(1): 273–292, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428107.

Blanchet:2016:ASA

[BGZ16] J. Blanchet, P. Glynn, and S. Zheng. Analysis of a stochastic approximation algorithm for computing quasi-stationary distributions. Advances in Applied Probability, 48(3):792–811, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1474296315.

Barbour:1989:SAS

[BH89a] A. D. Barbour and Lars Holst. Some applications of the Stein– Chen method for proving Poisson convergence. Advances in Applied Probability, 21(1):74–90, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427198.

Blom:1989:SPS

[BH89b] Gunnar Blom and Lars Holst. Some properties of similar pairs. Advances in Applied Probability, 21(4):941–944, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427779.

Buckdahn:1998:HCC

[BH98] Rainer Buckdahn and Ying Hu. Hedging contingent claims for a large investor in an incomplete market. Advances in Applied Probability, 30(1):239–255, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427886.

Baccelli:2000:AIR

[BH00] François Baccelli and Dohy Hong. Analyticity of iterates of random non-expansive maps. Advances in Applied Probability, 32(1):193–220, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428234.

Basak:2004:LCP

[BH04a] Gopal K. Basak and Kwok-Wah Remus Ho. Level-crossing probabilities and first-passage times for linear processes. Advances in Applied Probability, 36(2):643–666, June 2004. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428470.

Borovkov:2004:CSC

[BH04b] A. A. Borovkov and A. Hordijk. Characterization and sufficient conditions for normed ergodicity of Markov chains. Advances in Applied Probability, 36(1):227–242, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428360.

Bhansali:1986:DIC

[Bha86] R. J. Bhansali. A derivation of the information criteria for selecting autoregressive models. Advances in Applied Probability, 18(2):360–387, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427304.

Braker:1998:HDB

[BHB98] H. Bräker, T. Hsing, and N. H. Bingham. On the Hausdorff distance between a convex set and an interior random convex Hull. Advances in Applied Probability, 30(2):295–316, June 1998. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427969.

Biggins:2011:MSR

[BHJ11] J. D. Biggins, B. M. Hambly, and O. D. Jones. Multifractal spectra for random self-similar measures via branching processes. Advances in Applied Probability, 43(1):1–39, March 2011. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/23024523.

Barbour:2015:EBM

[BHKK15] A. D. Barbour, K. Hamza, Haya Kaspi, and F. C. Klebaner. Escape from the boundary in Markov population processes. Advances in Applied Probability, 47(4):1190–1211, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1449859806.

Bianchi:2002:DCP

[Bia02] Gabriele Bianchi. Determining convex polygons from their covariograms. Advances in Applied Probability, 34(2):261–266, June 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428286.

Biggins:1976:FLB

[Big76] J. D. Biggins. The first- and last-birth problems for a multitype age-dependent branching process. Advances in Applied Probability, 8(3):446–459, September 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426138.

Biggins:1978:ASB

[Big78] J. D. Biggins. The asymptotic shape of the branching random walk. Advances in Applied Probability, 10(1):62–84, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426719.

Biggins:1987:NRS

[Big87] J. D. Biggins. A note on repeated sequences in Markov chains. Advances in Applied Probability, 19(3):739–742, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427415.

Bingham:1973:LTF

[Bin73a] N. H. Bingham. Limit theorems in fluctuation theory. Advances in Applied Probability, 5(3):554–569, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425834.

Bingham:1973:MPS

[Bin73b] N. H. Bingham. Maxima of partial sums of random variables and suprema of stable processes. Advances in Applied Probability, 5(1):5–6, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425943.

Bingham:1975:FTC

[Bin75] N. H. Bingham. Fluctuation theory in continuous time. Advances in Applied Probability, 7(4):705–766, December 1975.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426397.

Binet:1979:BCP

[Bin79] F. E. Binet. On Bartlett’s characterisation of the Pólya process. Advances in Applied Probability, 11(2):276–277, June 1979. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426806.

Bingham:1980:RWN

[Bin80] N. H. Bingham. Random walks and number-theoretic densities. Advances in Applied Probability, 12(2):280–281, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426551.

Bingham:1991:FTE

[Bin91] N. H. Bingham. Fluctuation theory for the Ehrenfest urn. Advances in Applied Probability, 23(3):598–611, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427624.

Bremaud:1977:PPM

[BJ77] P. Brémaud and J. Jacod. Processus ponctuels et martingales: résultats récents sur la modélisation et le filtrage. (French) [Point processes and martingales: Recent results on modeling and filtering]. Advances in Applied Probability, 9(2):362– 416, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426391.

Britton:2007:GSD

[BJML07] Tom Britton, Svante Janson, and Anders Martin-Löf. Graphs with specified degree distributions, simple epidemics, and local vaccination strategies. Advances in Applied Probability, 39(4): 922–948, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443552.

Benes:1980:OSL

[BK80] Václav E. Benes and Ioannis Karatzas. Optimal stationary linear control of the Wiener process. Advances in Applied Probability, 12(2):288–289, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426561.

Becker:1983:DPP

[BK83] Gerhard Becker and Götz Kersting. Design problems for the pure birth process. Advances in Applied Probability, 15(2):255– 273, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426436.

Berry:1991:WPI

[BK91] Donald A. Berry and Robert P. Kertz. Worth of perfect information in Bernoulli bandits. Advances in Applied Probability, 23(1):1–23, March 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427509.

Baxendale:1998:SIP

[BK98] P. H. Baxendale and R. Z. Khasminskii. Stability index for products of random transformations. Advances in Applied Probability, 30(4):968–988, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428025.

Boshuizen:1999:SFS

[BK99] Frans A. Boshuizen and Robert P. Kertz. Smallest-fit selection of random sizes under a sum constraint: Weak convergence and moment comparisons. Advances in Applied Probability, 31(1): 178–198, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427997.

Boxma:2001:MGQ

[BK01] O. J. Boxma and I. A. Kurkova. The M/G/1 queue with two service speeds. Advances in Applied Probability, 33(2):520– 540, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428266.

Brix:2002:SCP

[BK02] Anders Brix and Wilfrid S. Kendall. Simulation of cluster point processes without edge effects. Advances in Applied Probability, 34(2):267–280, June 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428287.

Biggins:2004:MCM

[BK04] J. D. Biggins and A. E. Kyprianou. Measure change in multitype branching. Advances in Applied Probability, 36(2):544–581, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428466.

Buchmann:2005:MSP

[BK05] Boris Buchmann and Claudia Klüppelberg. Maxima of stochastic processes driven by fractional Brownian motion. Advances in Applied Probability, 37(3):743–764, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037353.

Bandyopadhyay:2018:LPU

[BK18] Antar Bandyopadhyay and Gursharn Kaur. Linear depreferential urn models. Advances in Applied Probability, 50 (4):1176–1192, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/linear-depreferential-urn-models/
B98982111752C7E21B2DADC21EE8D70E.

Bremaud:1992:ETA

[BKM92] Pierre Brémaud, Raghavan Kannurpatti, and Ravi Mazumdar. Event and time averages: A review. Advances in Applied Probability, 24(2):377–411, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427697.

Ball:2011:TBE

[BKO11] Frank G. Ball, Edward S. Knock, and Philip D. O’Neill. Threshold behaviour of emerging epidemics featuring contact tracing. Advances in Applied Probability, 43(4):1048–1065, December 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/23208038.

Ballani:2012:RMS

[BKS12a] F. Ballani, Z. Kabluchko, and M. Schlather. Random marked sets. Advances in Applied Probability, 44(3):603–616, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/41714067.

Bhaskar:2012:ASF

[BKS12b] Anand Bhaskar, John A. Kamm, and Yun S. Song. Approximate sampling formulae for general finite-alleles models of mutation. Advances in Applied Probability, 44(2):408–428, June 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/41714057.

Baccelli:1995:PAF

[BKZ95] François Baccelli, Maurice Klein, and Sergei Zuyev. Perturbation analysis of functionals of random measures. Advances in Applied Probability, 27(2):306–325, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427827.

Bhat:1988:NSM

[BL88] U. Narayan Bhat and Ram Lal. Number of successes in Markov trials. Advances in Applied Probability, 20(3):677–680, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427041.

Baccelli:1989:SCP

[BL89] François Baccelli and Zhen Liu. On the stability condition of a precedence-based queueing discipline. Advances in Applied Probability, 21(4):883–898, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427772.

Ball:2001:SMS

[BL01] Frank Ball and Owen D. Lyne. Stochastic multitype SIR epidemics among a population partitioned into households. Advances in Applied Probability, 33(1):99–123, March 2001. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428443.

Bedford:2004:IPR

[BL04] Tim Bedford and Bo H. Lindqvist. The identifiability problem for repairable systems subject to competing risks. Advances in Applied Probability, 36(3):774–790, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/4140407.

Barry:2006:MBT

[BL06] Daniel Barry and Caroline Lynch. The minimax bookie: The two-horse case. Advances in Applied Probability, 38(4):899–915, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443481.

Baumstark:2007:SDR

[BL07] Volker Baumstark and Günter Last. Some distributional results for Poisson–Voronoi tessellations. Advances in Applied Probability, 39(1):16–40, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443498.

Blanchet:2008:SDI

[BL08a] Jose H. Blanchet and Jingchen Liu. State-dependent importance sampling for regularly varying random walks. Advances in Applied Probability, 40(4):1104–1128, December 2008. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443626.

Borovkov:2008:LCG

[BL08b] K. Borovkov and G. Last. On level crossings for a general class of piecewise-deterministic Markov processes. Advances in Applied Probability, 40(3):815–834, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443609.

Baumstark:2009:GDS

[BL09a] Volker Baumstark and Günter Last. Gamma distributions for stationary Poisson flat processes. Advances in Applied Probability, 41(4):911–939, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/27793911.

Bruss:2009:OAB

[BL09b] F. Thomas Bruss and Guy Louchard. The odds algorithm based on sequential updating and its performance. Advances in Applied Probability, 41(1):131–153, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/27793867.

Bean:2010:AQB

[BL10] Nigel Bean and Guy Latouche. Approximations to quasi-birth-and-death processes with infinite blocks. Advances in Applied Probability, 42(4):1102–1125, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/25799086.

Bo:2022:PAR

[BL22] Lijun Bo and Huafu Liao. Probabilistic analysis of replicatormutator equations. Advances in Applied Probability, 54 (1):167–201, March 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/probabilistic-analysis-of-replicatormutator-
equations/C6705AD60C844C5DC0F9F56601CDFC38.

Bar-Lev:2012:IHR

[BLL12] Shaul K. Bar-Lev and Gérard Letac. Increasing hazard rate of mixtures for natural exponential families. Advances in Applied Probability, 44(2):373–390, June 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/41714055.

Blom:1983:MWT

[Blo83] Gunnar Blom. The mean waiting time to a repetition. Advances in Applied Probability, 15(1):216–218, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426993.

Blom:1989:MTT

[Blo89] Gunnar Blom. Mean transition times for the Ehrenfest urn model. Advances in Applied Probability, 21(2):479–480, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427173.

Bai:2006:REM

[BLP06] Z. D. Bai, Sungchul Lee, and Mathew D. Penrose. Rooted edges of a minimal directed spanning tree on random points. Advances in Applied Probability, 38(1):1–30, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443425.

Baserga:1980:PSR

[BLS^{+}80] Renato Baserga, Brian I. Lord, R. Schofield, Christopher S. Potten, A. D. Barbour, Robert Bartoszyński, John D. Biggins, James M. Greenberg, F. C. Hoppensteadt, W. Jäger, Harry Kesten, Masao Nagasawa, G. Oster, Werner Rittgen, Henry C. Tuckwell, Michael Turelli, Arthur T. Winfree, and Carla Wofsy. Proliferation, spread and reaction-dispersal processes. Advances in Applied Probability, 12(3):546–555, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426415.

Block:1988:BEG

[BLS88] H. W. Block, N. A. Langberg, and D. S. Stoffer. Bivariate exponential and geometric autoregressive and autoregressive moving average models. Advances in Applied Probability, 20(4):798–821, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427361.

Bruss:2003:TPR

[BLT03] F. Thomas Bruss, Guy Louchard, and John W. Turner. On the N-tower problem and related problems. Advances in Applied Probability, 35(1):278–294, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428283.

Britton:2019:WIG

[BLT19] Tom Britton, Ka Yin Leung, and Pieter Trapman. Who is the infector? General multi-type epidemics and realtime susceptibility processes. Advances in Applied Probability, 51(2):606–631, June 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/who-is-the-infector-general-multitype-
epidemics-and-realtime-susceptibility-processes/40917C81DF35BA57A2763D776FF0FB4A.

Blum:1984:LTD

[Blu84] Gilles A. Blum. A limit theorem for discrete-parameter random evolutions. Advances in Applied Probability, 16(2):281– 292, June 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427070.

Beutler:1977:DCS

[BM77] Frederick J. Beutler and Benjamin Melamed. Decomposition and customer streams of feedback queueing networks in equilibrium. Advances in Applied Probability, 9(2):215–216, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426358.

Bertsimas:1996:UMA

[BM96] Dimitris Bertsimas and Georgia Mourtzinou. A unified method to analyze overtake free queueing systems. Advances in Applied Probability, 28(2):588–625, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428073.

Borst:1998:AAP

[BM98] Sem Borst and Debasis Mitra. Asymptotically achievable performance in ATM networks. Advances in Applied Probability, 30(2):568–585, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427983.

Barbour:2000:CPA

[BM00] A. D. Barbour and Marianne Månsson. Compound Poisson approximation and the clustering of random points. Advances in Applied Probability, 32(1):19–38, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428224.

Bremaud:2002:PSG

[BM02] P. Brémaud and L. Massoulié. Power spectra of general shot noises and Hawkes point processes with a random excitation. Advances in Applied Probability, 34(1):205–222, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428379.

Blackwell:2003:BAD

[BM03] Paul G. Blackwell and Jesper Møller. Bayesian analysis of deformed tessellation models. Advances in Applied Probability, 35(1):4–26, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428270.

Bambos:2004:QSR

[BM04] Nicholas Bambos and George Michailidis. Queueing and scheduling in random environments. Advances in Applied Probability, 36(1):293–317, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428363.

Bertoin:2005:FE

[BM05] Jean Bertoin and Servet Martínez. Fragmentation energy. Advances in Applied Probability, 37(2):553–570, June 2005. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/30037341.

Broman:2008:SIG

[BM08] Erik Broman and Ronald Meester. Survival of inhomogeneous Galton–Watson processes. Advances in Applied Probability, 40 (3):798–814, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/20443608.

Bonichon:2011:AGN

[BM11] Nicolas Bonichon and Jean-François Marckert. Asymptotics of geometrical navigation on a random set of points in the plane. Advances in Applied Probability, 43(4):899–942, December 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/23208032.

Beghin:2016:MFP

[BM16] Luisa Beghin and Claudio Macci. Multivariate fractional Poisson processes and compound sums. Advances in Applied Probability, 48(3):691–711, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1474296310.

Busic:2013:PCA

[BMM13] Ana Bušić, Jean Mairesse, and Irène Marcovici. Probabilistic cellular automata, invariant measures, and perfect sampling. Advances in Applied Probability, 45(4):960–980, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1386857853.

Bansaye:2016:SCI

[BMM16] Vincent Bansaye, Sylvie Méléard, and Richard Mathieu. Speed of coming down from infinity for birth-and-death processes. Advances in Applied Probability, 48(4):1183–1210, December 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1482548434.

Bhattacharya:2019:EMB

[BMPR19] Ayan Bhattacharya, Krishanu Maulik, Zbigniew Palmowski, and Parthanil Roy. Extremes of multitype branching random walks: heaviest tail wins. Advances in Applied Probability, 51(2):514–540, June 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/extremes-of-multitype-branching-
random-walks-heaviest-tail-wins/FEB97BD554AF91E8191DD353D0588F4F.

Bremaud:2005:PSR

[BMR05] Pierre Brémaud, Laurent Massoulié, and Andrea Ridolfi. Power spectra of random spike fields and related processes. Advances in Applied Probability, 37(4):1116–1146, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037376.

Baccelli:1989:FJQ

[BMS89] François Baccelli, Armand M. Makowski, and Adam Shwartz. The fork-join queue and related systems with synchronization constraints: Stochastic ordering and computable bounds. Advances in Applied Probability, 21(3):629–660, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427640.

Bryc:2009:LDL

[BMS09] Wlodek Bryc, David Minda, and Sunder Sethuraman. Large deviations for the leaves in some random trees. Advances in Applied Probability, 41(3):845–873, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/27793905.

Bayer:2018:FRE

[BMS18] Christian Bayer, Hilmar Mai, and John Schoenmakers. Forwardreverse expectation-maximization algorithm for Markov chains: convergence and numerical analysis. Advances in Applied Probability, 50(2):621–644, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/forwardreverse-expectationmaximization-
algorithm-for-markov-chains-convergence-and-numerical-
analysis/E3106A90B4C70E6BE89DE908D5051037.

Ball:1997:SIA

[BMTY97] Frank Ball, Robin K. Milne, Ian D. Tame, and Geoffrey F. Yeo. Superposition of interacting aggregated continuous-time Markov chains. Advances in Applied Probability, 29(1):56–91, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427861.

Borst:2003:GPS

[BMV03] Sem Borst, Michel Mandjes, and Miranda Van Uitert. Generalized processor sharing queues with heterogeneous traffic classes. Advances in Applied Probability, 35(3):806–845, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428220.

Ball:1991:ASM

[BMY91] Frank Ball, Robin K. Milne, and Geoffrey F. Yeo. Aggregated semi-Markov processes incorporating time interval omission. Advances in Applied Probability, 23(4):772–797, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427675.

Ball:1994:CTM

[BMY94] Frank Ball, Robin K. Milne, and Geoffrey F. Yeo. Continuoustime Markov chains in a random environment, with applications to ion channel modelling. Advances in Applied Probability, 26(4): 919–946, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427898.

Bertacchi:2014:LGS

[BMZ14] Daniela Bertacchi, Fábio Prates Machado, and Fabio Zucca. Local and global survival for nonhomogeneous random walk systems on Z. Advances in Applied Probability, 46(1):256–278, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1396360113.

Bunge:1992:EDT

[BN92] J. A. Bunge and H. N. Nagaraja. Exact distribution theory for some point process record models. Advances in Applied Probability, 24(1):20–44, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427728. See correction [Ano93e].

Barndorff-Nielsen:1994:NEN

[BN94] O. E. Barndorff-Nielsen. A note on electrical networks and the inverse Gaussian distribution. Advances in Applied Probability, 26(1):63–67, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427579.

Balkema:2010:AIU

[BN10] Guus Balkema and Natalia Nolde. Asymptotic independence for unimodal densities. Advances in Applied Probability, 42(2):411– 432, June 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25683827.

Balkema:2012:ADL

[BN12] Guus Balkema and Natalia Nolde. Asymptotic dependence for light-tailed homothetic densities. Advances in Applied Probability, 44(2):506–527, June 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/41714061.

Balkema:2020:SLS

[BN20] Guus Balkema and Natalia Nolde. Samples with a limit shape, multivariate extremes, and risk. Advances in Applied Probability, 52(2):491–522, June 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-

probability/article/samples-with-a-limit-shape-multivariate-
extremes-and-risk/0F0EBCA47B2607A69A3049106B0555FF.

Ball:2023:SMS

[BN23] Frank Ball and Peter Neal. The size of a Markovian SIR epidemic given only removal data. Advances in Applied Probability, 55(3):895–926, September 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/size-of-a-markovian-sir-
epidemic-given-only-removal-data/E4238B5A0578B1487B9A0C962EDB1B34.

Barndorff-Nielsen:2000:MJP

[BNBJ00] Ole E. Barndorff-Nielsen, Fred Espen Benth, and Jens Ledet Jensen. Markov jump processes with a singularity. Advances in Applied Probability, 32(3):779–799, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428413.

Barndorff-Nielsen:2014:MEF

[BNBV14] Ole E. Barndorff-Nielsen, Fred Espen Benth, and Almut E. D. Veraart. Modelling electricity futures by ambit fields. Advances in Applied Probability, 46(3):719–745, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1409319557.

Barndorff-Nielsen:1981:SMS

[BND81] O. Barndorff-Nielsen and J. N. Darroch. A stochastic model for sand sorting in a wind-tunnel. Advances in Applied Probability, 13(2):282–297, June 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426684.

Barndorff-Nielsen:1998:SSP

[BNJS98] O. E. Barndorff-Nielsen, J. L. Jensen, and M. Sørensen. Some stationary processes in discrete and continuous time. Advances in Applied Probability, 30(4):989–1007, December 1998. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428026.

Barndorff-Nielsen:1998:TRC

[BNK98] O. E. Barndorff-Nielsen and A. E. Koudou. Trees with random conductivities and the (reciprocal) inverse Gaussian distribution. Advances in Applied Probability, 30(2):409–424, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427975.

Barndorff-Nielsen:2001:MSS

[BNPS01] Ole E. Barndorff-Nielsen, Jan Pedersen, and Ken-Iti Sato. Multivariate subordination, self-decomposability and stability. Advances in Applied Probability, 33(1):160–187, March 2001. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428446.

Buonocore:1987:NIE

[BNR87] A. Buonocore, A. G. Nobile, and L. M. Ricciardi. A new integral equation for the evaluation of first-passage-time probability densities. Advances in Applied Probability, 19(4):784–800, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427102.

Barbour:2002:CPA

[BNX02] A. D. Barbour, S. Y. Novak, and A. Xia. Compound Poisson approximation for the distribution of extremes. Advances in Applied Probability, 34(1):223–240, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428380.

Ball:1993:MGS

[BO93] Frank Ball and Philip O’Neill. A modification of the general stochastic epidemic motivated by AIDS modelling. Advances in Applied Probability, 25(1):39–62, March 1993. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427495.

Ball:1994:SCS

[BO94] Frank Ball and Philip O’Neill. Strong convergence of stochastic epidemics. Advances in Applied Probability, 26(3):629–655, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427812.

Borovkov:2007:STR

[BO07] K. A. Borovkov and D. A. Odell. On spatial thinningreplacement processes based on Voronoi cells. Advances in Applied Probability, 39(2):293–306, June 2007. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443513.

Bobrowski:2004:QSD

[Bob04] A. Bobrowski. Quasi-stationary distributions of a pair of Markov chains related to time evolution of a DNA locus. Advances in Applied Probability, 36(1):57–77, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428353.

Bobrowski:2008:ABF

[Bob08] Adam Bobrowski. Asymptotic behavior of a Feller evolution family involved in the Fisher–Wright model. Advances in Applied Probability, 40(3):734–758, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443606.

Bondesson:1982:SID

[Bon82] Lennart Bondesson. On simulation from infinitely divisible distributions. Advances in Applied Probability, 14(4):855–869, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427027.

Bean:2021:YLS

[BOP21] Nigel G. Bean, Małgorzata M. O’Reilly, and Zbigniew Palmowski. Yaglom limit for stochastic fluid models. Advances in Applied Probability, 53(3):649–686, September 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/yaglom-limit-for-stochastic-
fluid-models/B7FB46E1C67DBAC84DCD349B3BE01533.

Borovkov:1979:RCL

[Bor79] A. A. Borovkov. On the rate of convergence and large deviations in the invariance principle. Advances in Applied Probability, 11 (2):260–261, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426791.

Born:1998:PPM

[Bor98] Eike Born. A point process model with stochastic intensities for a branching population of two dependent types. Advances in Applied Probability, 30(3):723–739, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428148.

Belzunce:2002:APF

[BOR02] Félix Belzunce, Eva-María Ortega, and José M. Ruiz. On ageing properties of first-passage times of increasing Markov processes. Advances in Applied Probability, 34(1):241–259, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428381.

Bordenave:2006:SSQ

[Bor06] C. Bordenave. Stability of spatial queueing systems. Advances in Applied Probability, 38(2):487–504, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443452.

Borovkov:2009:IBF

[Bor09] A. A. Borovkov. Insurance with borrowing: First- and secondorder approximations. Advances in Applied Probability, 41(4): 1141–1160, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/27793923.

Behme:2023:SFS

[BOS23] Anita Behme, David Oechsler, and René Schilling. On q-scale functions of spectrally negative Lévy processes. Advances in Applied Probability, 55(1):56–84, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-qscale-functions-of-spectrally-
negative-levy-processes/53039CB5EF5DC58BFD41CD6BD59850AC.

Boumezoued:2016:PVH

[Bou16] Alexandre Boumezoued. Population viewpoint on Hawkes processes. Advances in Applied Probability, 48(2):463–480, June 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1465490758.

Bowditch:2021:NGF

[Bow21] Adam Bowditch. Non-Gaussian fluctuations of randomly trapped random walks. Advances in Applied Probability, 53 (3):801–838, September 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/nongaussian-fluctuations-of-randomly-
trapped-random-walks/FA77844B00D53995F575BF80EF3A2AA6.

Boxma:1978:DSS

[Box78] O. J. Boxma. On the distribution of the supremum of the service times in a busy cycle of the M/G/1 queueing system. Advances in Applied Probability, 10(2):314–315, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426914.

Boxma:1979:TQMa

[Box79a] O. J. Boxma. On a tandem queueing model with identical service times at both counters, I. Advances in Applied Probability, 11(3):616–643, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426958.

Boxma:1979:TQMb

[Box79b] O. J. Boxma. On a tandem queueing model with identical service times at both counters, II. Advances in Applied Probability, 11(3):644–659, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426959.

Boxma:1983:CQO

[Box83] O. J. Boxma. The cyclic queue with one general and one exponential server. Advances in Applied Probability, 15(4):857–873, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427328.

Boyce:1978:ASR

[Boy78] William E. Boyce. Approximate solution of random ordinary differential equations. Advances in Applied Probability, 10(1): 172–184, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426724.

Barbour:1979:LTS

[BP79] A. D. Barbour and Anthony G. Pakes. Limit theorems for the simple branching process allowing immigration, II. The case of infinite offspring mean. Advances in Applied Probability, 11(1): 63–72, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426768.

Brown:1982:SDA

[BP82] T. C. Brown and P. K. Pollett. Some distributional approximations in Markovian queueing networks. Advances in Applied Probability, 14(3):654–671, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426679.

Bartoszynski:1983:TCI

[BP83] Robert Bartoszyński and Prem S. Puri. On two classes of interacting stochastic processes arising in cancer modeling. Advances in Applied Probability, 15(4):695–712, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427319.

Bouton:1994:CDO

[BP94] Catherine Bouton and Gilles Pagès. Convergence in distribution of the one-dimensional Kohonen algorithms when the stimuli are not uniform. Advances in Applied Probability, 26(1):80– 103, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427581.

Burton:1999:ODK

[BP99] Robert M. Burton and David C. Plaehn. One-dimensional Kohonen maps are super-stable with exponential rate. Advances in Applied Probability, 31(2):367–393, June 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428115.

Barbour:2000:VMR

[BP00] Andrew D. Barbour and Andrea Pugliese. On the varianceto-mean ratio in models of parasite distributions. Advances in Applied Probability, 32(3):701–719, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428409.

Bona:2017:RST

[BP17] Miklós Bóna and Boris Pittel. On a random search tree: asymptotic enumeration of vertices by distance from leaves. Advances in Applied Probability, 49(3):850–876, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/on-a-random-search-tree-
asymptotic-enumeration-of-vertices-by-distance-from-
leaves/EEE6903D24CC8938929C7ACECD533DF3. See corrigen-

dum [BP22].

Burdzy:2021:CCP

[BP21] Krzysztof Burdzy and Soumik Pal. Can coherent predictions be contradictory? Advances in Applied Probability, 53(1):133–161, March 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/can-coherent-predictions-be-contradictory/
8ECCC61D49E2E3A380A2E2BFCCDA506B.

Bona:2022:RST

[BP22] Miklós Bóna and Boris Pittel. On a random search tree: asymptotic enumeration of vertices by distance from leaves — Corrigendum. Advances in Applied Probability, 54(2):337–339, June 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/on-
a-random-search-tree-asymptotic-enumeration-of-vertices-
by-distance-from-leaves-corrigendum/D4A9683847D6E30FDFA0E19FDC71227E.

See [BP17].

Baurdoux:2023:PLZ

[BP23] Erik J. Baurdoux and José M. Pedraza. Predicting the last zero before an exponential time of a spectrally negative Lévy process. Advances in Applied Probability, 55(2):611–642, June 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/predicting-the-
last-zero-before-an-exponential-time-of-a-spectrally-
negative-levy-process/3141C1709602DBA22DD101A8E5A6D3F5.

Barczy:2021:ABP

[BPP21] Mátyás Barczy, Sandra Palau, and Gyula Pap. Asymptotic behavior of projections of supercritical multi-type continuous-state and continuous-time branching processes with immigration. Advances in Applied Probability, 53(4):

1023–1060, December 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/asymptotic-behavior-of-projections-
of-supercritical-multitype-continuousstate-and-continuoustime-
branching-processes-with-immigration/028B4E035DEBC26E8434FEA6D7224134.

Brockwell:1980:IID

[BPS80] P. J. Brockwell and N. Pacheco-Santiago. Invariant imbedding and dams with Markovian input rate. Advances in Applied Probability, 12(2):304–305, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426581.

Balakrishnan:2005:NSN

[BPS05] N. Balakrishnan, A. G. Pakes, and A. Stepanov. On the number and sum of near-record observations. Advances in Applied Probability, 37(3):765–780, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/30037354.

Budhiraja:2015:ACM

[BPS15] Amarjit Budhiraja, Vladas Pipiras, and Xiaoming Song. Admission control for multidimensional workload input with heavy tails and fractional Ornstein–Uhlenbeck process. Advances in Applied Probability, 47(2):476–505, June 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1435236984.

Blanchet:2019:ESS

[BPS19] Jose Blanchet, Yanan Pei, and Karl Sigman. Exact sampling for some multi-dimensional queueing models with renewal input. Advances in Applied Probability, 51(4): 1179–1208, December 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/exact-sampling-for-some-multidimensional-
queueing-models-with-renewal-input/BEDC4FD16B83E2E58C4A80D95C0CA59E.

Boxma:2006:MGC

[BPSZ06] Onno Boxma, David Perry, Wolfgang Stadje, and Shelemyahu Zacks. A Markovian growth-collapse model. Advances in Applied Probability, 38(1):221–243, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443436.

Bharucha-Reid:1976:AMP

[BR76] A. T. Bharucha-Reid. Algebraic models for probability measures on Banach spaces. Advances in Applied Probability, 8(2):215– 217, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425866.

Bordes:1983:AST

[BR83] G. Bordes and B. Roehner. Application of Stieltjes theory for S-fractions to birth and death processes. Advances in Applied Probability, 15(3):507–530, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426617.

Beutler:1986:TAO

[BR86] Frederick J. Beutler and Keith W. Ross. Time-average optimal constrained semi-Markov decision processes. Advances in Applied Probability, 18(2):341–359, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427303.

Ballerini:1987:RPL

[BR87] Rocco Ballerini and Sidney I. Resnick. Records in the presence of a linear trend. Advances in Applied Probability, 19(4):801–828, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427103.

Berger:1988:WNP

[BR88] M. A. Berger and P. H. Roberts. On the winding number problem with finite steps. Advances in Applied Probability, 20 (2):261–274, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427389.

Bruss:1991:WLS

[BR91] F. Thomas Bruss and James B. Robertson. ‘Wald’s Lemma’ for sums of order statistics of i.i.d. random variables. Advances in Applied Probability, 23(3):612–623, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427625.

Basu:1993:PSU

[BR93] Sabyasachi Basu and Gregory C. Reinsel. Properties of the spatial unilateral first-order ARMA model. Advances in Applied Probability, 25(3):631–648, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427527.

Bhatt:2004:RDS

[BR04] Abhay G. Bhatt and Rahul Roy. On a random directed spanning tree. Advances in Applied Probability, 36(1):19–42, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428351.

Barbour:2013:ABG

[BR13] A. D. Barbour and G. Reinert. Asymptotic behaviour of gossip processes and small-world networks. Advances in Applied Probability, 45(4):981–1010, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1386857854.

Bercu:2015:LDO

[BR15] Bernard Bercu and Adrien Richou. Large deviations for the Ornstein–Uhlenbeck process with shift. Advances in Applied Probability, 47(3):880–901, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1444308886.

Brandwajn:1978:ANS

[Bra78] A. Brandwajn. An approach to the numerical solution of some queueing networks. Advances in Applied Probability, 10(2): 333, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426932.

Brandt:1986:SES

[Bra86] Andreas Brandt. The stochastic equation Y_{n+1} = A_{n}Y_{n} + B_{n} with stationary coefficients. Advances in Applied Probability, 18 (1):211–220, March 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427243.

Brandiere:1998:DSM

[Bra98] Odile Brandière. The dynamic system method and the traps. Advances in Applied Probability, 30(1):137–151, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427880.

Baroyan:1971:CMI

[BRB^{+}71] O. V. Baroyan, L. A. Rvachev, U. V. Basilevsky, V. V. Ermakov, K. D. Frank, M. A. Rvachev, and V. A. Shashkov. Computer modelling of influenza epidemics for the whole country (USSR). Advances in Applied Probability, 3(2):224–226, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426167.

Bremaud:1975:EEF

[Bré75] P. Brémaud. Estimation de l’état d’une file d’attente et du temps de panne d’une machine par la méthode de semimartingales. (French) [Estimating the state of a queue, and the downtime of a machine by the method of semi-martingales]. Advances in Applied Probability, 7(4):845–863, December 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426403.

Bremaud:1976:BBC

[Bré76] P. Brémaud. Bang–Bang controls of point processes. Advances in Applied Probability, 8(2):385–394, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425910.

Brill:1976:NMM

[Bri76] Percy H. Brill. A new methodology for modelling a broad class of exponential queues. Advances in Applied Probability, 8(2): 242, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425890.

Brix:1999:GGM

[Bri99] Anders Brix. Generalized gamma measures and shot-noise Cox processes. Advances in Applied Probability, 31(4):929–953, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428335.

Brown:1972:LDT

[Bro72] Mark Brown. Low density traffic streams. Advances in Applied Probability, 4(1):177–192, April 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425810.

Brown:1973:SIS

[Bro73] Mark Brown. Statistical inference for stochastic processes. Advances in Applied Probability, 5(1):6–8, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425944.

Brockwell:1977:SDD

[Bro77] P. J. Brockwell. Stationary distributions for dams with additive input and content-dependent release rate. Advances in Applied Probability, 9(3):645–663, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426118.

Bromberg:1980:RWM

[Bro80] Neil Bromberg. Random walk models in xerographic development. Advances in Applied Probability, 12(2):305–306, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426582.

Brockwell:1985:ETB

[Bro85] P. J. Brockwell. The extinction time of a birth, death and catastrophe process and of a related diffusion model. Advances in Applied Probability, 17(1):42–52, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427051.

Brozius:1989:CMS

[Bro89] Henk Brozius. Convergence in mean of some characteristics of the convex Hull. Advances in Applied Probability, 21(3):526–542, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427634.

Brown:1995:BAI

[Bro95a] Jason J. Brown. A bootstrap algorithm for the isotropic random sphere. Advances in Applied Probability, 27(3):627–641, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428127.

Brown:1995:CLT

[Bro95b] Jason J. Brown. A central limit theorem for the isotropic random sphere. Advances in Applied Probability, 27(3):642–651, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428128.

Browne:1998:RIP

[Bro98] Sid Browne. The return on investment from proportional portfolio strategies. Advances in Applied Probability, 30(1):216– 238, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427885.

Browne:1999:RGD

[Bro99] Sid Browne. Reaching goals by a deadline: Digital options and continuous-time active portfolio management. Advances in Applied Probability, 31(2):551–577, June 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428122.

Brooms:2005:NEF

[Bro05] A. C. Brooms. On the Nash equilibria for the FCFS queueing system with load-increasing service rate. Advances in Applied Probability, 37(2):461–481, June 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/30037336.

Blaszczyszyn:1999:BCS

[BRS99] Bartlomiej Blaszczyszyn, Christian Rau, and Volker Schmidt. Bounds for clump size characteristics in the Boolean model. Advances in Applied Probability, 31(4):910–928, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428334.

Brockwell:1982:SPG

[BRT82] P. J. Brockwell, S. I. Resnick, and R. L. Tweedie. Storage processes with general release rule and additive inputs. Advances in Applied Probability, 14(2):392–433, June 1982. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426528.

Bruss:1980:CBC

[Bru80] F. Thomas Bruss. A counterpart of the Borel–Cantelli lemma. Advances in Applied Probability, 12(2):281–282, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426552.

Birnbaum:1978:CIM

[BS78] D. Birnbaum and R. Serfozo. Convergence of image measures and integrals. Advances in Applied Probability, 10(2):286– 288, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426889.

Barbour:1981:IAR

[BS81] A. D. Barbour and R. Schassberger. Insensitive average residence times in generalized semi-Markov processes. Advances in Applied Probability, 13(4):720–735, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426969.

Buckley:1983:ATM

[BS83] M. J. Buckley and E. Seneta. On Arnold’s treatment of Moran’s bounds. Advances in Applied Probability, 15(1):212–213, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426991.

Ball:1988:AMP

[BS88] Frank Ball and Mark Sansom. Aggregated Markov processes incorporating time interval omission. Advances in Applied Probability, 20(3):546–572, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1427035.

Bauer:1995:EPG

[BS95] Christina Bauer and Rolf Schneider. Extremal problems for geometric probabilities involving convex bodies. Advances in Applied Probability, 27(1):20–34, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428092.

Bassan:1999:SCR

[BS99] Bruno Bassan and Fabio Spizzichino. Stochastic comparisons for residual lifetimes and Bayesian notions of multivariate ageing. Advances in Applied Probability, 31(4):1078–1094, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428345.

Barbour:2001:LTB

[BS01] A. D. Barbour and V. Schmidt. On Laslett’s transform for the Boolean model. Advances in Applied Probability, 33(1):1– 5, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428437.

Berend:2003:BEM

[BS03a] Daniel Berend and Luba Sapir. Between the expert and majority rules. Advances in Applied Probability, 35(4):941–960, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428494.

Blaszczyszyn:2003:ADS

[BS03b] Bartlomiej Blaszczyszyn and René Schott. Approximate decomposition of some modulated-Poisson Voronoi tessellations. Advances in Applied Probability, 35(4):847–862, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428487.

Bohm:2003:PRA

[BS03c] Stephan Böhm and Volker Schmidt. Palm representation and approximation of the covariance of random closed sets. Advances in Applied Probability, 35(2):295–302, June 2003. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428421.

Biggins:2005:CRM

[BS05] J. D. Biggins and A. Rahimzadeh Sani. Convergence results on multitype, multivariate branching random walks. Advances in Applied Probability, 37(3):681–705, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037350.

Bender:2006:IMM

[BS06] Christian Bender and John Schoenmakers. An iterative method for multiple stopping: Convergence and stability. Advances in Applied Probability, 38(3):729–749, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443468.

Barral:2007:RSS

[BS07] Julien Barral and Stéphane Seuret. Renewal of singularity sets of random self-similar measures. Advances in Applied Probability, 39(1):162–188, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443505.

Ball:2012:SEM

[BS12a] Frank Ball and David Sirl. An SIR epidemic model on a population with random network and household structure, and several types of individuals. Advances in Applied Probability, 44(1):63– 86, March 2012. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
23207856.

Bhaskar:2012:CFA

[BS12b] Anand Bhaskar and Yun S. Song. Closed-form asymptotic sampling distributions under the coalescent with recombination for an arbitrary number of loci. Advances in Applied Probability, 44(2):391–407, June 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/41714056.

Blok:2015:CSM

[BS15] H. Blok and F. M. Spieksma. Countable state Markov decision processes with unbounded jump rates and discounted cost: optimality equation and approximations. Advances in Applied Probability, 47(4):1088–1107, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1449859801.

Baake:2022:EIM

[BS22] Michael Baake and Jeremy Sumner. On equal-input and monotone Markov matrices. Advances in Applied Probability, 54(2):460–492, June 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-equalinput-and-monotone-markov-
matrices/D8E3238A51F77706CE21452A37784ACA.

Bartroff:2011:FPO

[BSC11] Jay Bartroff and Ester Samuel-Cahn. The fighter problem: Optimal allocation of a discrete commodity. Advances in Applied Probability, 43(1):121–130, March 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/23024528.

Burman:1984:RDQ

[BSD^{+}84] David Y. Burman, Donald R. Smith, Ralph L. Disney, Jagadeesh Chandramohan, William A. Massey, Marcel F. Neuts, E. Arjas, P. Haara, R. L. Tweedie, J. S. Baras, A. J. Dorsey, A. M. Makowski, Jos H. A. De Smit, Kyung Y. Jo, Arnold Greenland, P. K. Reeser, David D. W. Yao, and J. A. Buzacott. Recent developments in queueing. Advances in Applied Probability, 16(1):6–9, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427217.

Blanchet-Scalliet:2017:SEF

[BSHJ17] Christophette Blanchet-Scalliet, Caroline Hillairet, and Ying Jiao. Successive enlargement of filtrations and application to insider information. Advances in Applied Probability, 49 (3):653–685, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/successive-enlargement-of-filtrations-
and-application-to-insider-information/5E1A00A98CB3421EA0556813185D9CCB.

Brink-Spalink:2017:GTL

[BSS17] Rebekka Brink-Spalink and Charline Smadi. Genealogies of two linked neutral loci after a selective sweep in a large population of stochastically varying size. Advances in Applied Probability, 49(1):279–326, March 2017. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/genealogies-
of-two-linked-neutral-loci-after-a-selective-sweep-in-
a-large-population-of-stochastically-varying-size/4A9506CBA5407D8605056768ECB6AD73.

Bitter:2023:CTL

[BSS23] Annemarie Bitter, Robert Stelzer, and Bennet Ströh. Continuoustime locally stationary time series models. Advances in Applied Probability, 55(3):965–998, September 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/continuoustime-locally-stationary-
time-series-models/883D58F742D18DDDD6A063903CF1D549.

Ball:2009:TBF

[BST09] Frank Ball, David Sirl, and Pieter Trapman. Threshold behaviour and final outcome of an epidemic on a random network with household structure. Advances in Applied Probability, 41 (3):765–796, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/27793902.

Bailey:1971:EPP

[BT71] Norman T. J. Bailey and Anthony S. Thomas. The estimation of parameters from population data on the general stochastic epidemic. Advances in Applied Probability, 3(2):211–214, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426161.

Baddeley:1998:PMP

[BT98] A. J. Baddeley and R. Turner. Practical maximum pseudolikelihood for spatial point patterns. Advances in Applied Probability, 30(2):273, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427942.

Bordenave:2008:MCM

[BT08] Charles Bordenave and Giovanni Luca Torrisi. Monte Carlo methods for sensitivity analysis of Poisson-driven stochastic systems, and applications. Advances in Applied Probability, 40 (2):293–320, June 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443584.

Baccelli:2000:MPP

[BTZ00] F. Baccelli, K. Tchoumatchenko, and S. Zuyev. Markov paths on the Poisson–Delaunay graph with applications to routeing in mobile networks. Advances in Applied Probability, 32(1):1– 18, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428223.

Barbour:1998:SSE

[BU98] A. D. Barbour and Sergey Utev. Solving the Stein equation in compound Poisson approximation. Advances in Applied Probability, 30(2):449–475, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427977.

Buchta:2013:EFV

[Buc13] Christian Buchta. Exact formulae for variances of functionals of convex hulls. Advances in Applied Probability, 45(4): 917–924, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1386857850.

Buchardt:2016:CAP

[Buc16] Kristian Buchardt. Continuous affine processes: transformations, Markov chains and life insurance. Advances in Applied Probability, 48(2):423–442, June 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1465490756.

Bueno:1988:NCL

[Bue88] Vanderlei Costa Bueno. A note on the component lifetime estimation of a multistate monotone system through the system lifetime. Advances in Applied Probability, 20(3):686–689, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427044.

Buhler:1974:FSP

[Büh74] W. J. Bühler. On the family structure of populations. Advances in Applied Probability, 6(2):192–193, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426242.

Bukszar:2001:UBP

[Buk01] József Bukszár. Upper bounds for the probability of a union by multitrees. Advances in Applied Probability, 33(2):437– 452, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428262.

Bulmer:1974:MCD

[Bul74] M. G. Bulmer. A model for character displacement based on competition between species. Advances in Applied Probability, 6(1):1–3, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426195.

Burman:1981:IQS

[Bur81] David Y. Burman. Insensitivity in queueing systems. Advances in Applied Probability, 13(4):846–859, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426976.

Butler:1994:EFS

[But94a] Steven M. Butler. The early and final states of an epidemic in a large heterogeneous population with a small initial number of infectives. Advances in Applied Probability, 26(3):671–689, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427814.

Butler:1994:FSE

[But94b] Steven M. Butler. The final state of an epidemic in a large heterogeneous population with a large initial number of infectives. Advances in Applied Probability, 26(3):656–670, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427813.

Butler:2022:EGF

[But22] Ronald W. Butler. Exponential and gamma form for tail expansions of first-passage distributions in semi-Markov processes. Advances in Applied Probability, 54(4):1291–1319, December 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/exponential-
and-gamma-form-for-tail-expansions-of-firstpassage-distributions-
in-semimarkov-processes/94F003AEF1FBF2DEF2A24E9EDC451E67.

Boucherie:1990:SBD

[BV90] Richard J. Boucherie and Nico M. Van Dijk. Spatial birth– death processes with multiple changes and applications to batch service networks and clustering processes. Advances in Applied Probability, 22(2):433–455, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427544.

Boucherie:1991:PFQ

[BV91] Richard J. Boucherie and Nico M. Van Dijk. Product forms for queueing networks with state-dependent multiple job transitions. Advances in Applied Probability, 23(1):152–187, March 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427516.

Bertsimas:1993:SDV

[BV93] Dimitris J. Bertsimas and Garrett Van Ryzin. Stochastic and dynamic vehicle routing with general demand and interarrival time distributions. Advances in Applied Probability, 25(4):947– 978, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427801.

Bedford:1997:RVL

[BV97] T. Bedford and J. Van Den Berg. A remark on the Van Lieshout and Baddeley J-function for point processes. Advances in Applied Probability, 29(1):19–25, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427858.

Boutsikas:2002:DBC

[BV02] Michael V. Boutsikas and Eutichia Vaggelatou. On the distance between convex-ordered random variables, with applications. Advances in Applied Probability, 34(2):349–374, June 2002. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428292.

Borovkov:2006:ABR

[BV06] K. A. Borovkov and V. A. Vatutin. On the asymptotic behaviour of random recursive trees in random environments. Advances in Applied Probability, 38(4):1047–1070, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443487.

Benth:2013:CCS

[BV13a] F. E. Benth and L. Vos. Cross-commodity spot price modeling with stochastic volatility and leverage for energy markets. Advances in Applied Probability, 45(2):545–571, June 2013. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1370870129.

Benth:2013:PFO

[BV13b] F. E. Benth and L. Vos. Pricing of forwards and options in a multivariate non-Gaussian stochastic volatility model for energy markets. Advances in Applied Probability, 45(2):572–594, June 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1370870130.

Bierkens:2020:SEH

[BvdMS20] Joris Bierkens, Frank van der Meulen, and Moritz Schauer. Simulation of elliptic and hypo-elliptic conditional diffusions. Advances in Applied Probability, 52(1):173–212, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/simulation-of-
elliptic-and-hypoelliptic-conditional-diffusions/3CBDD8A76C2FE03EF87ED987BAACFDB5.

Bhamidi:2010:EVT

[BVH10] Shankar Bhamidi, Remco Van Der Hofstad, and Gerard Hooghiemstra. Extreme value theory, Poisson–Dirichlet distributions, and first passage percolation on random networks. Advances in Applied Probability, 42(3):706–738, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/25799065.

Baddeley:1996:MPC

[BVM96] A. J. Baddeley, M. N. M. Van Lieshout, and J. Møller. Markov properties of cluster processes. Advances in Applied Probability, 28(2):346–355, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428060.

Buyukkoc:1985:RR

[BVW85] C. Buyukkoc, P. Varaiya, and J. Walrand. The cμ rule revisited. Advances in Applied Probability, 17(1):237–238, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427064.

Berman:1983:QSR

[BW83] Mark Berman and Mark Westcott. On queueing systems with renewal departure processes. Advances in Applied Probability, 15(3):657–673, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426624.

Bambos:1989:SSD

[BW89] Nicholas Bambos and Jean Walrand. On stability of statedependent queues and acyclic queueing networks. Advances in Applied Probability, 21(3):681–701, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427642.

Bambos:1990:IDG

[BW90a] Nicholas Bambos and Jean Walrand. An invariant distribution for the G/G/1 queueing operator. Advances in Applied Probability, 22(1):254–256, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427610.

Beirlant:1990:RVR

[BW90b] J. Beirlant and E. Willekens. Rapid variation with remainder and rates of convergence. Advances in Applied Probability, 22(4): 787–801, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427562.

Bambos:1993:SSA

[BW93] Nicholas Bambos and Jean Walrand. Scheduling and stability aspects of a general class of parallel processing systems. Advances in Applied Probability, 25(1):176–202, March 1993. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427501.

Browne:1996:PCB

[BW96] Sid Browne and Ward Whitt. Portfolio choice and the Bayesian Kelly criterion. Advances in Applied Probability, 28(4):1145– 1176, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428168.

Brockwell:1997:EAC

[BW97] P. J. Brockwell and R. J. Williams. On the existence and application of continuous-time threshold autoregressions of order two. Advances in Applied Probability, 29(1):205–227, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427867.

Bo:2010:OPP

[BWY10] Lijun Bo, Yongjin Wang, and Xuewei Yang. An optimal portfolio problem in a defaultable market. Advances in Applied Probability, 42(3):689–705, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/25799064.

Basrak:2019:TCA

[BWZ19] Bojan Basrak, Olivier Wintenberger, and Petra Zugec. On the total claim amount for marked Poisson cluster models. Advances in Applied Probability, 51(2):541–569, June 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/on-
the-total-claim-amount-for-marked-poisson-cluster-models/
FDEC12806B7F704EE4CAAF07FB599513.

Barbour:2001:NTD

[BX01] A. D. Barbour and A. Xia. The number of two-dimensional maxima. Advances in Applied Probability, 33(4):727–750, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428301.

Barbour:2006:NAR

[BX06] A. D. Barbour and Aihua Xia. Normal approximation for random sums. Advances in Applied Probability, 38(3):693–728, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443467.

Browne:1989:DPR

[BY89] Sid Browne and Uri Yechiali. Dynamic priority rules for cyclictype queues. Advances in Applied Probability, 21(2):432–450, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427168.

Blaszczyszyn:2009:DCO

[BY09] Bartlomiej Blaszczyszyn and D. Yogeshwaran. Directionally convex ordering of random measures, shot noise fields, and some applications to wireless communications. Advances in Applied Probability, 41(3):623–646, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/27793896.

Blaszczyszyn:2014:CCP

[BY14] Bartłomiej Błaszczyszyn and D. Yogeshwaran. On comparison of clustering properties of point processes. Advances in Applied Probability, 46(1):1–20, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1396360100.

Borst:2003:RPE

[BZ03] Sem Borst and Bert Zwart. A reduced-peak equivalence for queues with a mixture of light-tailed and heavy-tailed input flows. Advances in Applied Probability, 35(3):793–805, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428219.

Bertacchi:2014:SLS

[BZ14] Daniela Bertacchi and Fabio Zucca. Strong local survival of branching random walks is not monotone. Advances in Applied Probability, 46(2):400–421, June 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1401369700.

Blanchet:2020:ESM

[BZ20] Jose Blanchet and Fan Zhang. Exact simulation for multivariate Itô diffusions. Advances in Applied Probability, 52 (4):1003–1034, December 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/exact-simulation-for-multivariate-
ito-diffusions/580E7AF2DC0A4081487A88762918CF31.

Biagini:2022:ERF

[BZ22] Francesca Biagini and Yinglin Zhang. Extended reduced-form framework for non-life insurance. Advances in Applied Probability, 54(3):945–973, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/extended-reducedform-framework-for-
nonlife-insurance/ED18167A2E1A576C4B487012BD975E02.

Chen:1994:RDI

[CA94] Zhao-Guo Chen and Oliver D. Anderson. The representation and decomposition of integrated stationary time series. Advances in Applied Probability, 26(3):799–819, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427821.

Cai:2007:TVA

[Cai07] Jun Cai. On the time value of absolute ruin with debit interest. Advances in Applied Probability, 39(2):343–359, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/20443517.

Chadjiconstantinidis:2000:JDS

[CAK00] S. Chadjiconstantinidis, D. L. Antzoulakos, and M. V. Koutras. Joint distributions of successes, failures and patterns in enumeration problems. Advances in Applied Probability, 32(3):866–884, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428417.

Calka:2002:DSD

[Cal02] Pierre Calka. The distributions of the smallest disks containing the Poisson–Voronoi typical cell and the Crofton cell in the plane. Advances in Applied Probability, 34(4):702–717, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428318.

Caliebe:2003:SFP

[Cal03a] Amke Caliebe. Symmetric fixed points of a smoothing transformation. Advances in Applied Probability, 35(2):377–394, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428428.

Calka:2003:EED

[Cal03b] Pierre Calka. An explicit expression for the distribution of the number of sides of the typical Poisson–Voronoi cell. Advances in Applied Probability, 35(4):863–870, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428488.

Calka:2003:PFD

[Cal03c] Pierre Calka. Precise formulae for the distributions of the principal geometric characteristics of the typical cells of a twodimensional Poisson–Voronoi tessellation and a Poisson line process. Advances in Applied Probability, 35(3):551–562, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428207.

Cameron:1979:MT

[Cam79] M. A. Cameron. A model for transients. Advances in Applied Probability, 11(2):298–299, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426829.

Campbell:1984:OSB

[Cam84] Gregory Campbell. Optimal selection based on relative ranks of a sequence with ties. Advances in Applied Probability, 16 (1):131–146, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427228.

Cambanis:1988:RFW

[Cam88] Stamatis Cambanis. Random filters which preserve the stability of random inputs. Advances in Applied Probability, 20(2):275– 294, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427390.

Cannings:1974:GDP

[Can74a] C. Cannings. Genetic drift with polygamy and arbitrary offspring distribution. Advances in Applied Probability, 6(1):3– 4, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426196.

Cannings:1974:LRC

[Can74b] C. Cannings. The latent roots of certain Markov chains arising in genetics: A new approach, I. Haploid models. Advances in Applied Probability, 6(2):260–290, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426293.

Cannings:1974:LRS

[Can74c] C. Cannings. The latent roots of some genetic drift models. Advances in Applied Probability, 6(2):232–233, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426271.

Cannings:1975:DHM

[Can75a] C. Cannings. Drift in haploid models. Advances in Applied Probability, 7(1):4–5, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425841.

Cannings:1975:LRC

[Can75b] C. Cannings. The latent roots of certain Markov chains arising in genetics: A new approach, II. Further haploid models. Advances in Applied Probability, 7(2):264–282, June 1975. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426077.

Cannings:1976:MDM

[Can76] C. Cannings. Multitype drift models. Advances in Applied Probability, 8(1):1–2, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426008.

Cao:1987:RPC

[Cao87] X. R. Cao. Realization probability in closed Jackson queueing networks and its application. Advances in Applied Probability, 19(3):708–738, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427414.

Cao:1989:CST

[Cao89] Xi-Ren Cao. Calculation of sensitivities of throughputs and realization probabilities in closed queueing networks with finite buffer capacities. Advances in Applied Probability, 21(1):181– 206, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427203.

Cao:1990:RFS

[Cao90] Xi-Ren Cao. Realization factors and sensitivity analysis of queueing networks with state-dependent service rates. Advances in Applied Probability, 22(1):178–210, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427604.

Cao:1995:ECP

[Cao95] Xi-Ren Cao. Event coupling and performance sensitivity analysis of generalized semi-Markov processes. Advances in Applied Probability, 27(3):741–769, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428132.

Cao:1998:MSP

[Cao98] Xi-Ren Cao. The Maclaurin series for performance functions of Markov chains. Advances in Applied Probability, 30(3):676–692, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428145.

Cao:1999:SCC

[Cao99] J. Cao. The size of the connected components of excursion sets of χ^{2}, t and F fields. Advances in Applied Probability, 31(3): 579–595, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428384.

Capasso:1976:SAS

[Cap76] Vincenzo Capasso. A semigroup approach to the study of a general stochastic epidemic. Advances in Applied Probability, 8 (2):228–229, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425876.

Caperaa:1986:CLF

[Cap86] P. Capéraà. Caracterisation des lois fortement unimodales. (French) [Characterization of strongly unimodal laws]. Advances in Applied Probability, 18(1):277–278, March 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427247.

Carbonell:1978:MDM

[Car78] Emilio Carbonell. A model to detect maternal and epistatic effects. Advances in Applied Probability, 10(1):7–10, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426709.

Chung:2020:PNE

[CAR20] Hakjin Chung, Hyun-Soo Ahn, and Rhonda Righter. The potentially negative effects of cooperation in service systems. Advances in Applied Probability, 52(1):319–347, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/potentially-
negative-effects-of-cooperation-in-service-systems/27C21602CA8E8563B5EDFEC7E236F121.

Casse:2016:PCA

[Cas16] Jérôme Casse. Probabilistic cellular automata with general alphabets possessing a Markov chain as an invariant distribution. Advances in Applied Probability, 48(2):369–391, June 2016. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1465490753.

Caubet:1974:DED

[Cau74] J.-P. Caubet. Dirac equation and dynamics of Brownian motion. Advances in Applied Probability, 6(2):219, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426257.

Caubet:1976:RBM

[Cau76] Jean-Pierre Caubet. Relativistic Brownian motion. Advances in Applied Probability, 8(2):230–231, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425878.

Cavender:1978:QSD

[Cav78] James A. Cavender. Quasi-stationary distributions of birth-and-death processes. Advances in Applied Probability, 10(3):570–586, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426635.

Cabo:1995:LTC

[CB95] A. J. Cabo and A. J. Baddeley. Line transects, covariance functions and set convergence. Advances in Applied Probability, 27 (3):585–605, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428125.

Chin:1999:CCM

[CB99] Y. C. Chin and A. J. Baddeley. On connected component Markov point processes. Advances in Applied Probability, 31 (2):279–282, June 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428110.

Chin:2000:MIC

[CB00] Y. C. Chin and A. J. Baddeley. Markov interacting component processes. Advances in Applied Probability, 32(3):597–619, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428402.

Cabo:2003:EMP

[CB03] Annoesjka Cabo and Adrian Baddeley. Estimation of mean particle volume using the set covariance function. Advances in Applied Probability, 35(1):27–46, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428271.

Conolly:1977:GBD

[CC77] B. W. Conolly and J. Chan. Generalised birth and death queueing processes: Recent results. Advances in Applied Probability, 9(1):125–140, March 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425820.

Cowan:1996:RDFa

[CC96a] Richard Cowan and Simone Chen. The random division of faces in a planar graph. Advances in Applied Probability, 28(2): 331, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428040.

Cowan:1996:RDFb

[CC96b] Richard Cowan and Simone Chen. The random division of faces in a planar graph. Advances in Applied Probability, 28(2):377– 383, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428062.

Chen:1999:IDS

[CC99a] Francis K. C. Chen and Richard Cowan. Invariant distributions for shapes in sequences of randomly-divided rectangles. Advances in Applied Probability, 31(1):1–14, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427987.

Cowan:1999:FIP

[CC99b] Richard Cowan and Francis K. C. Chen. Four interesting problems concerning Markovian shape sequences. Advances in Applied Probability, 31(4):954–968, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428336.

Cowan:2005:LE

[CC05] Richard Cowan and S. N. Chiu. Letter to the Editor. Advances in Applied Probability, 37(3):857–858, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037359.

Chen:2009:EPD

[CC09] Hsiao-Chi Chen and Yunshyong Chow. Evolutionary Prisoner’s Dilemma games with one-dimensional local interaction and imitation. Advances in Applied Probability, 41(1):154–176, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/27793868.

Ceci:2012:NFJ

[CC12] Claudia Ceci and Katia Colaneri. Nonlinear filtering for jump diffusion observations. Advances in Applied Probability, 44(3): 678–701, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/41714071.

Chong:2000:RPC

[CCH00] K. S. Chong, Richard Cowan, and Lars Holst. The ruin problem and cover times of asymmetric random walks and Brownian motions. Advances in Applied Probability, 32(1):177–192, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428233.

Cavazos-Cadena:2018:VDA

[CCHH18] Rolando Cavazos-Cadena and Daniel Hernández-Hernández. Vanishing discount approximations in controlled Markov chains with risk-sensitive average criterion. Advances in Applied Probability, 50(1):204–230, March 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/vanishing-discount-approximations-
in-controlled-markov-chains-with-risksensitive-average-
criterion/4988FBB33A95D41332899B2CDD5D818A.

Chang:1991:MRQ

[CCP91] Cheng-Shang Chang, Xiu Li Chao, and Michael Pinedo. Monotonicity results for queues with doubly stochastic Poisson arrivals: Ross’s conjecture. Advances in Applied Probability, 23 (1):210–228, March 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427518.

Chan:1979:DCC

[CD79] P. K. W. Chan and T. Downs. Discounted cost criterion for imperfect maintenance. Advances in Applied Probability, 11(2): 304, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426835.

Coupier:2014:RSC

[CD14] D. Coupier and Yu. Davydov. Random symmetrizations of convex bodies. Advances in Applied Probability, 46(3):603– 621, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1409319551.

Chenavier:2018:SFP

[CD18] Nicolas Chenavier and Olivier Devillers. Stretch factor in a planar Poisson-Delaunay triangulation with a large intensity.

Advances in Applied Probability, 50(1):35–56, March 2018. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/stretch-factor-
in-a-planar-poissondelaunay-triangulation-with-a-large-
intensity/A1098E5721C7A5045B988511DB3B766D.

Caron:2011:CDS

[CDDP11] François Caron, Pierre Del Moral, Arnaud Doucet, and Michele Pace. On the conditional distributions of spatial point processes. Advances in Applied Probability, 43(2):301–307, June 2011. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/23024506.

Csorgo:1987:ASS

[CDH87] Miklós Csörgő, Paul Deheuvels, and Lajos Horváth. An approximation of stopped sums with applications in queueing theory. Advances in Applied Probability, 19(3):674–690, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427412.

Crimaldi:2013:GTP

[CDIM13] Irene Crimaldi, Antonio Di Crescenzo, Antonella Iuliano, and Barbara Martinucci. A generalized telegraph process with velocity driven by random trials. Advances in Applied Probability, 45(4):1111–1136, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1386857860.

Crisan:2022:LNC

[CDJR22] Dan Crisan, Pierre Del Moral, Ajay Jasra, and Hamza Ruzayqat. Log-normalization constant estimation using the ensemble Kalman–Bucy filter with application to highdimensional models. Advances in Applied Probability, 54 (4):1139–1163, December 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/lognormalization-constant-estimation-
using-the-ensemble-kalmanbucy-filter-with-application-
to-highdimensional-models/68B11B3FCBD87552E23EE9B6C0D8417F.

Crisan:1999:IPS

[CDL99] D. Crisan, P. Del Moral, and T. J. Lyons. Interacting particle systems approximations of the Kushner-stratonovitch equation. Advances in Applied Probability, 31(3):819–838, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428398.

Chan:2012:RES

[CDL12] Hock Peng Chan, Shaojie Deng, and Tze-Leung Lai. Rare-event simulation of heavy-tailed random walks by sequential importance sampling and resampling. Advances in Applied Probability, 44(4):1173–1196, December 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/41805994.

Cekanavicius:1997:AEE

[Cek97] V. Cekanavicius. Asymptotic expansions in the exponent: A compound Poisson approach. Advances in Applied Probability, 29(2):374–387, June 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428008.

Cekanavicius:2003:IDA

[Cek03] V. Cekanavicius. Infinitely divisible approximations for discrete nonlattice variables. Advances in Applied Probability, 35(4):982– 1006, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428496.

Cohen:2010:GCT

[CEP10] Samuel N. Cohen, Robert J. Elliott, and Charles E. M. Pearce. A general comparison theorem for backward stochastic differential equations. Advances in Applied Probability, 42(3):878–898, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25799072.

Cerf:1998:ACG

[Cer98] Raphaël Cerf. Asymptotic convergence of genetic algorithms. Advances in Applied Probability, 30(2):521–550, June 1998. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427981.

Cerf:2017:QRS

[Cer17] Raphaël Cerf. The quasispecies regime for the simple genetic algorithm with roulette wheel selection. Advances in Applied Probability, 49(3):903–926, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/quasispecies-regime-for-
the-simple-genetic-algorithm-with-roulette-wheel-selection/
4DDE4D060C6A34B860FDA1E81BF12172.

Capasso:1978:SCA

[CF78] V. Capasso and S. L. Paveri Fontana. Stochastic compartmental analysis. Advances in Applied Probability, 10(2):323–325, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426923.

Chao:1991:RLS

[CF91] M. T. Chao and James C. Fu. The reliability of a large series system under Markov structure. Advances in Applied Probability, 23(4):894–908, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427682.

Chiarini:2014:LDS

[CF14] Alberto Chiarini and Markus Fischer. On large deviations for small noise Itô processes. Advances in Applied Probability, 46(4):1126–1147, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1418396246.

Chandramohan:1985:TPP

[CFD85] Jagadeesh Chandramohan, Robert D. Foley, and Ralph L. Disney. Thinning of point processes: Covariance analyses. Advances in Applied Probability, 17(1):127–146, March 1985. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427056.

Coffman:1987:MEM

[CFGW87] E. G. Coffman, Jr., L. Flatto, M. R. Garey, and R. R. Weber. Minimizing expected makespans on uniform processor systems. Advances in Applied Probability, 19(1):177–201, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427379.

Cuevas:2012:SPS

[CFPL12] A. Cuevas, R. Fraiman, and B. Pateiro-López. On statistical properties of sets fulfilling rolling-type conditions. Advances in Applied Probability, 44(2):311–329, June 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/41714052a.

Coffman:1987:OSS

[CFW87] E. G. Coffman, Jr., L. Flatto, and R. R. Weber. Optimal selection of stochastic intervals under a sum constraint. Advances in Applied Probability, 19(2):454–473, June 1987. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427427.

Cai:2009:ETD

[CFW09] Jun Cai, Runhuan Feng, and Gordon E. Willmot. On the expectation of total discounted operating costs up to default and its applications. Advances in Applied Probability, 41(2):495– 522, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
27793888.

Cannings:1980:DPA

[CG80] C. Cannings and D. Gardiner. A dynamic programming approach to evolutionarily stable strategy theory. Advances in Applied Probability, 12(1):3–5, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426480.

Charalambous:2008:OSP

[CG08] C. Charalambous and J. C. Gittins. Optimal selection policies for a sequence of candidate drugs. Advances in Applied Probability, 40(2):359–376, June 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443587.

Cha:2016:CMC

[CG16] Ji Hwan Cha and Massimiliano Giorgio. On a class of multivariate counting processes. Advances in Applied Probability, 48 (2):443–462, June 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1465490757.

Chalker:1999:SRS

[CGH^{+}99] T. K. Chalker, A. P. Godbole, P. Hitczenko, J. Radcliff, and O. G. Ruehr. On the size of a random sphere of influence graph. Advances in Applied Probability, 31(3):596–609, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428385.

Costa:2020:RHP

[CGMT20] Manon Costa, Carl Graham, Laurence Marsalle, and Viet Chi Tran. Renewal in Hawkes processes with self-excitation and inhibition. Advances in Applied Probability, 52(3): 879–915, September 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/renewal-in-hawkes-processes-with-
selfexcitation-and-inhibition/29602BD9DA9F1EFE521C77E7D162F57C.

Coulson:2018:CPA

[CGR18] Matthew Coulson, Robert E. Gaunt, and Gesine Reinert. Compound Poisson approximation of subgraph counts in stochastic block models with multiple edges. Advances in Applied Probability, 50(3):759–782, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/compound-poisson-approximation-of-
subgraph-counts-in-stochastic-block-models-with-multiple-
edges/A97FB752EE26DF710DDC6404732D3FFC.

Chiang:1976:MTP

[CH76] C. L. Chiang and J. P. Hsu. On multiple transition probabilities. Advances in Applied Probability, 8(2):217–219, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425867.

Cohen:1980:BEB

[CH80] J. W. Cohen and G. Hooghiemstra. Brownian excursion, Brownian meander and limit theorems for the M/G/1 queue. Advances in Applied Probability, 12(2):297–298, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426573.

Chen:2014:CEI

[CH14] Rui Chen and Ollivier Hyrien. On classes of equivalence and identifiability of age-dependent branching processes. Advances in Applied Probability, 46(3):704–718, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1409319556.

Chenavier:2016:EIP

[CH16] Nicolas Chenavier and Ross Hemsley. Extremes for the inradius in the Poisson line tessellation. Advances in Applied Probability, 48(2):544–573, June 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1465490762.

Charlesworth:1974:HWL

[Cha74] Brian Charlesworth. The Hardy–Weinberg law with overlapping generations. Advances in Applied Probability, 6(1):4–6, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426197.

Charlesworth:1976:RMF

[Cha76] Brian Charlesworth. Recombination modification in a fluctuating environment. Advances in Applied Probability, 8(1):2– 4, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426009.

Chan:1989:NGE

[Cha89] K. S. Chan. A note on the geometric ergodicity of a Markov chain. Advances in Applied Probability, 21(3):702–704, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427643.

Chang:1992:NOS

[Cha92a] Cheng-Shang Chang. A new ordering for stochastic majorization: Theory and applications. Advances in Applied Probability, 24(3):604–634, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427482.

Chao:1992:DPM

[Cha92b] Xiuli Chao. On the departure processes of M/M/1/N and GI/G/1/N queues. Advances in Applied Probability, 24(3):751– 754, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427488.

Chan:1996:SDM

[Cha96] Terence Chan. Some diffusion models for the Mabinogion sheep problem of Williams. Advances in Applied Probability, 28(3): 763–783, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428180.

Chang:1999:RCA

[Cha99a] Joseph T. Chang. Recent common ancestors of all present-day individuals. Advances in Applied Probability, 31(4):1002–1026, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428340.

Chang:1999:RDR

[Cha99b] Joseph T. Chang. Reply to discussants: Recent common ancestors of all present-day individuals. Advances in Applied Probability, 31(4):1036–1038, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428342. See [DAG95].

Chan:2009:MMS

[Cha09a] Hock Peng Chan. Maxima of moving sums in a Poisson random field. Advances in Applied Probability, 41(3):647–663, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
27793897.

Chassagneux:2009:DTA

[Cha09b] Jean-François Chassagneux. A discrete-time approximation for doubly reflected BSDEs. Advances in Applied Probability, 41 (1):101–130, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/27793866.

Cha:2014:CGP

[Cha14] Ji Hwan Cha. Characterization of the generalized Pólya process and its applications. Advances in Applied Probability, 46(4):1148–1171, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1418396247.

Cha:2022:ECU

[Cha22] Ji Hwan Cha. An extended class of univariate and multivariate generalized Pólya processes. Advances in Applied Probability, 54(3):974–997, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/an-extended-class-of-univariate-
and-multivariate-generalized-polya-processes/390030E00037905C4E9C59D39BC2AFF5.

Chen:1989:CPO

[Che89a] Robert W. Chen. A circular property of the occurrence of sequence patterns in the fair coin-tossing process. Advances in Applied Probability, 21(4):938–940, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427778.

Chen:1989:IST

[Che89b] Thomas M. Chen. On the independence of sojourn times in tandem queues. Advances in Applied Probability, 21(2):488– 489, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427176.

Chen:1998:OLB

[Che98] Tuhao Chen. Optimal lower bounds for bivariate probabilities. Advances in Applied Probability, 30(2):476–492, June 1998. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427978.

Chen:2000:WDB

[Che00] Yung-Pin Chen. Which design is better? Ehrenfest urn versus biased coin. Advances in Applied Probability, 32(3):738–749, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428411.

Chen:2015:NSC

[Che15] Xinxin Chen. A necessary and sufficient condition for the nontrivial limit of the derivative martingale in a branching random walk. Advances in Applied Probability, 47(3):741–760, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1444308880.

Chikte:1980:OSS

[Chi80] Shirish D. Chikte. Optimal sequential selection and resource allocation under uncertainty. Advances in Applied Probability, 12(4):942–957, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426750.

Chiu:1994:MVF

[Chi94] S. N. Chiu. Mean-value formulae for the neighbourhood of the typical cell of a random tessellation. Advances in Applied Probability, 26(3):565–576, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1427808.

Chiu:1995:LTT

[Chi95] S. N. Chiu. Limit theorems for the time of completion of Johnson–Mehl tessellations. Advances in Applied Probability, 27 (4):889–910, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427927.

Chi:2012:ESN

[Chi12] Zhiyi Chi. On exact sampling of nonnegative infinitely divisible random variables. Advances in Applied Probability, 44(3): 842–873, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/41714078.

Chi:2014:NSJ

[Chi14] Zhiyi Chi. Nonnormal small jump approximation of infinitely divisible distributions. Advances in Applied Probability, 46(4): 963–984, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1418396239.

Chan:2016:TSP

[CHJ16] Hock Peng Chan, Chiang-Wee Heng, and Ajay Jasra. Theory of segmented particle filters. Advances in Applied Probability, 48(1):69–87, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1457466156.

Chernick:1991:CEI

[CHM91] Michael R. Chernick, Tailen Hsing, and William P. McCormick. Calculating the extremal index for a class of stationary sequences. Advances in Applied Probability, 23(4):835–850, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427679.

Choukri:1993:TBP

[Cho93] T. Choukri. The transient blocking probabilities in M/M/N loss systems via large deviations. Advances in Applied Probability, 25(2):483–486, June 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427665.

Cui:2020:EDM

[CHY20] Lirong Cui, Alan Hawkes, and He Yi. An elementary derivation of moments of Hawkes processes. Advances in Applied Probability, 52(1):102–137, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/an-elementary-derivation-of-moments-
of-hawkes-processes/79A7355542F08087C8AE828C664A3049.

Cox:1986:VWT

[CI86] D. R. Cox and Valerie Isham. The virtual waiting-time and related processes. Advances in Applied Probability, 18(2):558– 573, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427312.

Champagnat:2000:CSU

[CI00] F. Champagnat and J. Idier. On the correlation structure of unilateral AR processes on the plane. Advances in Applied Probability, 32(2):408–425, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428196.

Christensen:2017:OPS

[CIL17] Sören Christensen, Albrecht Irle, and Andreas Ludwig. Optimal portfolio selection under vanishing fixed transaction costs. Advances in Applied Probability, 49(4):1116–1143, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/optimal-portfolio-
selection-under-vanishing-fixed-transaction-costs/9972119C0B22B6C7869C4D19946728F6.

Cinlar:1969:MRT

[Çin69] Erhan Çinlar. Markov renewal theory. Advances in Applied Probability, 1(2):123–187, Autumn 1969. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426216.

Cinlar:1974:PMR

[Çin74] Erhan Çinlar. Periodicity in Markov renewal theory. Advances in Applied Probability, 6(1):61–78, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426207.

Cinlar:1975:SSM

[Çin75] Erhan Çinlar. On the structure of semi-Markov processes. Advances in Applied Probability, 7(2):230, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426046.

Cinlar:1978:MPP

[Çin78] Erhan Çinlar. Markov point processes. Advances in Applied Probability, 10(2):262–263, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1426871.

Curien:2011:PMQ

[CJ11] Nicolas Curien and Adrien Joseph. Partial match queries in two-dimensional quadtrees: a probabilistic approach. Advances in Applied Probability, 43(1):178–194, March 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/23024531.

Chesney:1997:BEP

[CJPY97] Marc Chesney, Monique Jeanblanc-Picqué, and Marc Yor. Brownian excursions and Parisian barrier options. Advances in Applied Probability, 29(1):165–184, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427865.

Callaert:1973:EES

[CK73] Herman Callaert and Julian Keilson. On exponential ergodicity and spectral structure for birth–death processes. Advances in Applied Probability, 5(1):9–11, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1425946.

Cox:1985:SCB

[CK85] R. Mitchell Cox and Ioannis Karatzas. Stationary control of Brownian motion in several dimensions. Advances in Applied Probability, 17(3):531–561, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427119.

Clark:1999:SCS

[CK99] J. M. C. Clark and V. Katsouros. Stably coalescent stochastic froths. Advances in Applied Probability, 31(1):199–219, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427998.

Chen:2007:SSP

[CK07] Chingfer Chen and Samuel Karlin. r-Scan statistics of a Poisson process with events transformed by duplications, deletions, and displacements. Advances in Applied Probability, 39(3):799–825, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443541.

Connor:2015:PSM

[CK15] Stephen B. Connor and Wilfrid S. Kendall. Perfect simulation of M/G/c queues. Advances in Applied Probability, 47(4):1039– 1063, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1449859799.

Cai:2014:TSL

[CKL14] Ning Cai, S. G. Kou, and Zongjian Liu. A two-sided Laplace inversion algorithm with computable error bounds and its applications in financial engineering. Advances in Applied Probability, 46(3):766–789, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1409319559.

Czarna:2020:FIO

[CKLP20] Irmina Czarna, Adam Kaszubowski, Shu Li, and Zbigniew Palmowski. Fluctuation identities for omega-killed spectrally negative Markov additive processes and dividend problem. Advances in Applied Probability, 52(2):404–432, June 2020. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/fluctuation-
identities-for-omegakilled-spectrally-negative-markov-
additive-processes-and-dividend-problem/950900D4533421BFD06CB03AF9B0DB11.

Cooley:2019:STJ

[CKM19] Oliver Cooley, Tobias Kapetanopoulos, and Tamás Makai. The sharp threshold for jigsaw percolation in random graphs. Advances in Applied Probability, 51(2):378–407, June 2019. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/sharp-threshold-
for-jigsaw-percolation-in-random-graphs/214C80305CB8141668ECDBDCA30C90A5.

Cordeiro:2019:ECL

[CKO19] James D. Cordeiro, Jeffrey P. Kharoufeh, and Mark E. Oxley. On the ergodicity of a class of level-dependent quasibirth-and-death processes. Advances in Applied Probability, 51(4):1109–1128, December 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-the-ergodicity-of-a-class-of-
leveldependent-quasibirthanddeath-processes/3171939A1584BB544DCA9175471649B6.

Caldentey:2009:FIB

[CKW09] René Caldentey, Edward H. Kaplan, and Gideon Weiss. FCFS infinite bipartite matching of servers and customers. Advances in Applied Probability, 41(3):695–730, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/27793900.

Chang:1987:OSD

[CL87] Fu Chang and Tze Leung Lai. Optimal stopping and dynamic allocation. Advances in Applied Probability, 19(4):829–853, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427104.

Cressie:2000:HPM

[CL00] Noel Cressie and Andrew B. Lawson. Hierarchical probability models and Bayesian analysis of mine locations. Advances in Applied Probability, 32(2):315–330, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428190.

Coupechoux:2014:HCA

[CL14] Emilie Coupechoux and Marc Lelarge. How clustering affects epidemics in random networks. Advances in Applied Probability, 46(4):985–1008, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1418396240.

Coupechoux:2015:CRN

[CL15] Emilie Coupechoux and Marc Lelarge. Contagions in random networks with overlapping communities. Advances in Applied Probability, 47(4):973–988, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1449859796.

Christensen:2022:MCO

[CL22] Sören Christensen and Kristoffer Lindensjö. Moment-constrained optimal dividends: precommitment and consistent planning. Advances in Applied Probability, 54(2):404–432, June 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/momentconstrained-
optimal-dividends-precommitment-and-consistent-planning/
0DB19D4BAE5443BD429E94D331A85F11.

Chen:2012:EPI

[CLCZ12] Anyue Chen, Junping Li, Yiqing Chen, and Dingxuan Zhou. Extinction probability of interacting branching collision processes. Advances in Applied Probability, 44(1):226–259, March 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/23207862.

Cline:2007:SNS

[Cli07] Daren B. H. Cline. Stability of nonlinear stochastic recursions with application to nonlinear AR–GARCH models. Advances in Applied Probability, 39(2):462–491, June 2007. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443522. See correc-

tion [Ano07e].

Cline:2011:ICA

[Cli11] Daren B. H. Cline. Irreducibility and continuity assumptions for positive operators with application to threshold GARCH time series models. Advances in Applied Probability, 43(1):49– 76, March 2011. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
23024525.

Chakraborty:2008:NCS

[CLM08] Santanu Chakraborty and Jose Alfredo López-Mimbela. Nonexplosion of a class of semilinear equations via branching particle representations. Advances in Applied Probability, 40(1):250– 272, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443579.

Crimaldi:2023:INR

[CLM23] Irene Crimaldi, Pierre-Yves Louis, and Ida G. Minelli. Interacting nonlinear reinforced stochastic processes: Synchronization or non-synchronization. Advances in Applied Probability, 55(1):275–320, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/interacting-nonlinear-reinforced-
stochastic-processes-synchronization-or-nonsynchronization/
0E09826E21309B9D15E4DA032F5B3069.

Cloez:2017:LTS

[Clo17] Bertrand Cloez. Limit theorems for some branching measurevalued processes. Advances in Applied Probability, 49(2):549– 580, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL https://www.cambridge.org/
core/journals/advances-in-applied-probability/article/
limit-theorems-for-some-branching-measurevalued-processes/
3B5FEF99874C7084A7792E61910BC11A.

Choudhury:1995:ACB

[CLW95] Gagan L. Choudhury, Kin K. Leung, and Ward Whitt. An algorithm to compute blocking probabilities in multi-rate multi-class multi-resource loss models. Advances in Applied Probability, 27 (4):1104–1143, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427936.

Collins:1993:JSP

[CM93] E. J. Collins and J. M. McNamara. The job-search problem with competition: An evolutionarily stable dynamic strategy. Advances in Applied Probability, 25(2):314–333, June 1993. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427655.

Collins:1998:FHD

[CM98] E. J. Collins and J. M. McNamara. Finite-horizon dynamic optimisation when the terminal reward is a concave functional of the distribution of the final state. Advances in Applied Probability, 30(1):122–136, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427879.

Chao:2000:QNI

[CM00] Xiuli Chao and Masakiyo Miyazawa. Queueing networks with instantaneous movements: A unified approach by quasireversibility. Advances in Applied Probability, 32(1):284–313, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428238.

Chiu:2003:NGR

[CM03] S. N. Chiu and I. S. Molchanov. A new graph related to the directions of nearest neighbours in a point process. Advances in Applied Probability, 35(1):47–55, March 2003. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428272.

Cowan:2009:LEC

[CM09] Richard Cowan and R. E. Miles. LETTER TO THE EDITOR: Convex hulls on a hemisphere. Advances in Applied Probability, 41(4):1002–1004, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/27793915.

Crisan:2017:UCT

[CM17] Dan Crisan and Joaquín Míguez. Uniform convergence over time of a nested particle filtering scheme for recursive parameter estimation in state-space Markov models. Advances in Applied Probability, 49(4):1170–1200, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/uniform-convergence-over-
time-of-a-nested-particle-filtering-scheme-for-recursive-
parameter-estimation-in-statespace-markov-models/8413116D5EC4CE32B370C06F11876B3F.

Campi:2019:SEM

[CM19a] Luciano Campi and Claude Martini. On the support of extremal martingale measures with given marginals: the countable case. Advances in Applied Probability, 51(2):570–605, June 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/on-
the-support-of-extremal-martingale-measures-with-given-
marginals-the-countable-case/9528B789F39057B4589BD70DDB4FAA8F.

Cha:2019:TLP

[CM19b] Ji Hwan Cha and Sophie Mercier. Transformed Lévy processes as state-dependent wear models. Advances in Applied Probability, 51(2):468–486, June 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/transformed-levy-processes-as-statedependent-
wear-models/8EBB0B07BA4270BB161F7E6AEFEA4D05.

Cazares:2019:ESE

[CMB19] Jorge I. González Cázares, Aleksandar Mijatović, and Gerónimo Uribe Bravo. Exact simulation of the extrema of stable processes. Advances in Applied Probability, 51(4):967–993, December 2019.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/exact-simulation-
of-the-extrema-of-stable-processes/E08EF62D67DFDF49C8311BB8F47DC1EA.

Collet:2013:SMC

[CMMS13] Pierre Collet, Servet Martínez, Sylvie Méléard, and Jaime San Martín. Stochastic models for a chemostat and long-time behavior. Advances in Applied Probability, 45(3):822–837, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1377868540.

Calka:2009:RCB

[CMP09] Pierre Calka, Julien Michel, and Katy Paroux. Refined convergence for the Boolean model. Advances in Applied Probability, 41(4):940–957, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/27793912.

Chevallier:2021:DAM

[CMT21] Julien Chevallier, Anna Melnykova, and Irene Tubikanec. Diffusion approximation of multi-class Hawkes processes: Theoretical and numerical analysis. Advances in Applied Probability, 53(3):716–756, September 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/diffusion-approximation-of-multiclass-
hawkes-processes-theoretical-and-numerical-analysis/
9329B1AE1775361620772A06C59E6E2D.

Carlsson:1986:APL

[CN86] Hasse Carlsson and Olle Nerman. An alternative proof of Lorden’s renewal inequality. Advances in Applied Probability, 18 (4):1015–1016, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427260.

Chaumont:2018:MBM

[CN18] Loïc Chaumont and Thi Ngoc Anh Nguyen. On mutations in the branching model for multitype populations. Advances in Applied Probability, 50(2):543–564, June 2018. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-mutations-in-the-branching-model-
for-multitype-populations/20C77BF913C1E13D9321463BF9E4783A.

Chang:1991:STC

[CNP91] Cheng-Shang Chang, Randolph Nelson, and Michael Pinedo. Scheduling two classes of exponential jobs on parallel processors: Structural results and worst-case analysis. Advances in Applied Probability, 23(4):925–944, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427684.

Chen:2005:WSS

[CNT05] Yiqing Chen, Kai W. Ng, and Qihe Tang. Weighted sums of subexponential random variables and their maxima. Advances in Applied Probability, 37(2):510–522, June 2005. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/30037338.

Cruz-Orive:1998:I

[CO98a] Luis M. Cruz-Orive. Introduction. Advances in Applied Probability, 30(2):272, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427941.

Cruz-Orive:1998:RP

[CO98b] Luis M. Cruz-Orive. The rods problem. Advances in Applied Probability, 30(2):276–277, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427946.

Chen:2016:CWB

[COC16] Ningyuan Chen and Mariana Olvera-Cravioto. Coupling on weighted branching trees. Advances in Applied Probability, 48 (2):499–524, June 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1465490760.

Cohen:1972:SAV

[Coh72] J. W. Cohen. The suprema of the actual and virtual waiting times during a busy cycle of the K_{m} /K_{n} /1 queueing system. Advances in Applied Probability, 4(2):339–356, August 1972.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426003.

Cohen:1973:ARQ

[Coh73] J. W. Cohen. Asymptotic relations in queueing theory. Advances in Applied Probability, 5(1):8–9, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1425945.

Cohen:1974:SRS

[Coh74] J. W. Cohen. Some remarks on stationary distributions for the GI/G/1 queue. Advances in Applied Probability, 6(2):252– 253, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426287.

Cohn:1976:FNH

[Coh76] Harry Cohn. Finite non-homogeneous Markov chains: Asymptotic behaviour. Advances in Applied Probability, 8(3):502–516, September 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426141.

Cohen:1977:OSL

[Coh77a] J. W. Cohen. On the optimal switching level for an M/G/1 queueing system. Advances in Applied Probability, 9(2):191, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426338.

Cohen:1977:EASa

[Coh77b] Joel E. Cohen. Ergodicity of age structure in populations with Markovian vital rates. II. General states. Advances in Applied Probability, 9(1):18–37, March 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425814.

Cohen:1977:EASb

[Coh77c] Joel E. Cohen. Ergodicity of age structure in populations with Markovian vital rates, III: Finite-state moments and growth rate; an illustration. Advances in Applied Probability, 9(3): 462–475, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/ stable/1426109. See correction [Ano78e].

Cohn:1977:CNH

[Coh77d] Harry Cohn. Countable non-homogeneous Markov chains: Asymptotic behaviour. Advances in Applied Probability, 9(3): 542–552, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426113.

Cohen:1978:GCP

[Coh78] J. W. Cohen. On a general class of processer sharing service facilities. Advances in Applied Probability, 10(2):334–335, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426934.

Cohn:1979:IEC

[Coh79] Harry Cohn. On the invariant events of a convergent Markov chain. Advances in Applied Probability, 11(2):277, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426807.

Cohn:1981:PDN

[Coh81] Harry Cohn. On a paper by Doeblin on non-homogeneous Markov chains. Advances in Applied Probability, 13(2):388– 401, June 1981. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426690.

Cohen:1982:CRU

[Coh82a] Jonathan P. Cohen. Convergence rates for the ultimate and pentultimate approximations in extreme-value theory. Advances in Applied Probability, 14(4):833–854, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427026.

Cohen:1982:PFA

[Coh82b] Jonathan P. Cohen. The penultimate form of approximation to normal extremes. Advances in Applied Probability, 14(2):324– 339, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426524.

Cohen:1985:MUP

[Coh85] Joel E. Cohen. Measuring the uncertainty of population forecasts: A comparison of two approaches. Advances in Applied Probability, 17(2):246–247, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427132.

Cohn:1988:LBS

[Coh88] Harry Cohn. Limit behaviour for stochastic monotonicity and applications. Advances in Applied Probability, 20(2):331–347, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427393.

Cohen:1991:AWT

[Coh91] J. W. Cohen. On the attained waiting time. Advances in Applied Probability, 23(3):660–661, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427628.

Coleman:1974:NND

[Col74] R. Coleman. Nearest neighbour distances for equilibrium renewal processes. Advances in Applied Probability, 6(2):220, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426258.

Coleman:1998:CRS

[Col98] Rodney Coleman. The crossing of random secants. Advances in Applied Probability, 30(2):275–276, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427945.

Conolly:1977:RDP

[Con77] B. W. Conolly. Recently discovered properties of generalised birth and death state-dependent queueing systems. Advances in Applied Probability, 9(2):216–217, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426359.

Conolly:1983:MTR

[Con83] B. W. Conolly. More transient results for generalised statedependent Erlangian queues. Advances in Applied Probability, 15(3):688–690, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426627.

Corbett:1978:PQM

[Cor78] J. V. Corbett. Probability in quantum mechanics. Advances in Applied Probability, 10(4):725–729, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426653.

Coron:2014:SMD

[Cor14] Camille Coron. Stochastic modeling of density-dependent diploid populations and the extinction vortex. Advances in Applied Probability, 46(2):446–477, June 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1401369702.

Chauvin:1991:FSB

[CORR91] B. Chauvin, P. Olivares-Rieumont, and A. Rouault. Fluctuations of spatial branching processes with mean-field interaction. Advances in Applied Probability, 23(4):716–732, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427672.

Courtois:1990:TLP

[Cou90] P.-J. Courtois. On the total loss probability of service stages in series. Advances in Applied Probability, 22(1):260–262, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427612.

Covo:2009:ODD

[Cov09] Shai Covo. One-dimensional distributions of subordinators with upper truncated Lévy measure, and applications. Advances in Applied Probability, 41(2):367–392, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/27793883.

Cowan:1989:DSP

[Cow89a] Richard Cowan. The division of space and the Poisson distribution. Advances in Applied Probability, 21(1):233–234, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427208.

See note [Mec89].

Cowan:1989:OAR

[Cow89b] Richard Cowan. Objects arranged randomly in space: An accessible theory. Advances in Applied Probability, 21(3):543–569, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427635.

Cowan:1996:SPA

[Cow96] Richard Cowan. Shapes of prisms after repeated random division. Advances in Applied Probability, 28(2):331, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428039.

Cowan:1997:SRP

[Cow97] Richard Cowan. Shapes of rectangular prisms after repeated random division. Advances in Applied Probability, 29(1):26– 37, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427859.

Cowan:2006:MCC

[Cow06] Richard Cowan. A more comprehensive complementary theorem for the analysis of Poisson point processes. Advances in Applied Probability, 38(3):581–601, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443462.

Cowan:2007:ILV

[Cow07] Richard Cowan. Identities linking volumes of convex hulls. Advances in Applied Probability, 39(3):630–644, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/20443533.

Cowan:2010:NCR

[Cow10] Richard Cowan. New classes of random tessellations arising from iterative division of cells. Advances in Applied Probability, 42(1):26–47, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25683805.

Cowan:2013:LSI

[Cow13] Richard Cowan. Line segments in the isotropic planar STIT tessellation. Advances in Applied Probability, 45(2):295–311, June 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1370870119.

Cox:1980:TCF

[Cox80] J. Theodore Cox. The time constant of first-passage percolation on the square lattice. Advances in Applied Probability, 12(4): 864–879, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426745.

Crank:1988:MAM

[CP88] Keith N. Crank and Prem S. Puri. A method of approximating Markov jump processes. Advances in Applied Probability, 20 (1):33–58, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427269.

Chryssaphinou:1990:LDC

[CP90] Ourania Chryssaphinou and Stavros G. Papastavridis. Limit distribution for a consecutive-k-out-of-n: F system. Advances in Applied Probability, 22(2):491–493, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427550.

Caramellino:2008:LDE

[CP08] Lucia Caramellino and Barbara Pacchiarotti. Large deviation estimates of the crossing probability for pinned Gaussian processes. Advances in Applied Probability, 40(2):424–453, June 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/20443590.

Cha:2015:BPB

[CP15] Ji Hwan Cha and Gianpaolo Pulcini. Burn-in procedure based on a dependent covariate process. Advances in Applied Probability, 47(2):506–529, June 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1435236985.

Casanova:2021:BPI

[CPP21] Adrián González Casanova, Juan Carlos Pardo, and José Luis Pérez. Branching processes with interactions: subcritical cooperative regime. Advances in Applied Probability, 53(1):251–278, March 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/branching-processes-with-interactions-
subcritical-cooperative-regime/03E7D952430A4283F29990C34B951330.

Chiarella:2001:FMT

[CPR01] Carl Chiarella, Sara Pasquali, and Wolfgang J. Runggaldier. On filtering in Markovian term structure models: An approximation approach. Advances in Applied Probability, 33(4):794–809, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428306.

Chen:2005:UCC

[CPZC05] Anyue Chen, Phil Pollett, Hanjun Zhang, and Ben Cairns. Uniqueness criteria for continuous-time Markov chains with general transition structures. Advances in Applied Probability, 37 (4):1056–1074, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/30037373.

Chiu:2001:CLT

[CQ01] S. N. Chiu and M. P. Quine. Central limit theorem for germination-growth models in R^{d} with non-Poisson locations. Advances in Applied Probability, 33(4):751–755, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428302.

Cowan:2003:DGD

[CQZ03] Richard Cowan, Malcolm Quine, and Sergei Zuyev. Decomposition of gamma-distributed domains constructed from Poisson point processes. Advances in Applied Probability, 35(1):56– 69, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428273.

Curnow:1978:TLM

[CR78] R. N. Curnow and J. H. Roger. A two-locus model for the inheritance of disease. Advances in Applied Probability, 10(1): 10, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426710.

Chauvin:1990:SSS

[CR90] B. Chauvin and A. Rouault. A stochastic simulation for solving scalar reaction-diffusion equations. Advances in Applied Probability, 22(1):88–100, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427598.

Chen:2004:MBA

[CR04] Anyue Chen and Eric Renshaw. Markovian bulk-arriving queues with state-dependent control at idle time. Advances in Applied Probability, 36(2):499–524, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428464.

Casella:2008:EMC

[CR08] Bruno Casella and Gareth O. Roberts. Exact Monte Carlo simulation of killed diffusions. Advances in Applied Probability, 40 (1):273–291, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443580.

Carstens:2011:TTC

[CR11] Sebastian Carstens and Thomas Richthammer. The two-type continuum Richardson model: Nondependence of the survival of both types on the initial configuration. Advances in Applied Probability, 43(3):597–615, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/23024566.

Chen:2019:CEI

[CR19] Ye Chen and Ilya O. Ryzhov. Complete expected improvement converges to an optimal budget allocation. Advances in Applied Probability, 51(1):209–235, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/complete-expected-improvement-converges-
to-an-optimal-budget-allocation/5F61B13648DDAED930B5AB0889B5D134.

Craven:1974:ARM

[Cra74a] B. D. Craven. Asymptotic rate of a Markov process. Advances in Applied Probability, 6(4):732–746, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426189.

Craven:1974:SDM

[Cra74b] B. D. Craven. The spectral density of a Markov process. Advances in Applied Probability, 6(2):216–217, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426254.

Cramer:1997:SAS

[Cra97] M. Cramer. Stochastic analysis of ‘Simultaneous Merge-Sort’. Advances in Applied Probability, 29(3):669–694, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428081.

Crane:2017:GMB

[Cra17] Harry Crane. Generalized Markov branching trees. Advances in Applied Probability, 49(1):108–133, March 2017. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/generalized-
markov-branching-trees/288CC7B8E5680DA3E0ADA50197745813.

Cuevas:2004:BE

[CRC04] Antonio Cuevas and Alberto Rodríguez-Casal. On boundary estimation. Advances in Applied Probability, 36(2):340–354, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428456.

Cressie:1974:TDR

[Cre74] N. A. Cressie. A two-dimensional random walk in the presence of a partially reflecting barrier. Advances in Applied Probability, 6(2):248–249, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426283.

Cressie:1979:RSL

[Cre79] Noel Cressie. Random set limit theorems. Advances in Applied Probability, 11(2):281–282, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426812.

Croydon:2007:HDC

[Cro07] D. A. Croydon. The Hausdorff dimension of a class of random self-similar fractal trees. Advances in Applied Probability, 39(3):

708–730, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443537.

Croydon:2010:SLS

[Cro10] D. A. Croydon. Scaling limits for simple random walks on random ordered graph trees. Advances in Applied Probability, 42 (2):528–558, June 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25683832.

Chen:2018:ISH

[CRZ18] Bohan Chen, Chang-Han Rhee, and Bert Zwart. Importance sampling of heavy-tailed iterated random functions. Advances in Applied Probability, 50(3):805–832, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/importance-sampling-of-heavytailed-
iterated-random-functions/A45C7CE80945A169903BE977715613EF.

Campbell:1981:CBC

[CS81] Gregory Campbell and Stephen M. Samuels. Choosing the best of the current crop. Advances in Applied Probability, 13(3): 510–532, September 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426783.

Chao:1991:DBP

[CS91a] Chern-Ching Chao and Norman C. Severo. Distributions of ballot problem random variables. Advances in Applied Probability, 23(3):586–597, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427623.

Chen:1991:SSC

[CS91b] Robert W. Chen and Larry A. Shepp. Synchronous service on a circle. Advances in Applied Probability, 23(1):140–151, March 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427515.

Chen:1997:USR

[CS97] Yiping Chen and Nozer D. Singpurwalla. Unification of software reliability models by self-exciting point processes. Advances in Applied Probability, 29(2):337–352, June 1997. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428006.

Choe:1999:SDI

[CS99] Jinwoo Choe and Ness B. Shroff. On the supremum distribution of integrated stationary Gaussian processes with negative linear drift. Advances in Applied Probability, 31(1):135–157, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427995.

Calka:2006:LDP

[CS06] Pierre Calka and Tomasz Schreiber. Large deviation probabilities for the number of vertices of random polytopes in the ball. Advances in Applied Probability, 38(1):47–58, March 2006. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443427.

Czichowsky:2012:CDM

[CS12] Christoph Czichowsky and Martin Schweizer. Convex duality in mean-variance hedging under convex trading constraints. Advances in Applied Probability, 44(4):1084–1112, December 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/41805991.

Christensen:2017:ICE

[CS17] Sören Christensen and Paavo Salminen. Impulse control and expected suprema. Advances in Applied Probability, 49(1):238–257, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/impulse-control-and-expected-suprema/
E0590BFAB2E8DCEFC32EF400353F512C.

Csenki:1992:JDS

[Cse92] Attila Csenki. The joint distribution of sojourn times in finite Markov processes. Advances in Applied Probability, 24(1):141– 160, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427733.

Conley:1980:FMA

[CSH^{+}80] C. Conley, Joel Smoller, J. M. Hammersley, David G. Kendall, Nguyen Xuan Xanh, George C. Papanicolaou, and Robert T.

Smythe. Further mathematical approaches and techniques. Advances in Applied Probability, 12(3):571–573, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426420.

Chadoeuf:1998:PEB

[CSY98] Joël Chadoeuf, Rachid Senoussi, and Jian-Feng Yao. Parametric estimation in Boolean segment processes by a restorationestimation method. Advances in Applied Probability, 30(2):274– 275, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427944.

Chan:1985:UDL

[CT85] K. S. Chan and H. Tong. On the use of the deterministic Lyapunov function for the ergodicity of stochastic difference equations. Advances in Applied Probability, 17(3):666–678, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427125.

Cowan:1994:FLM

[CT94] Richard Cowan and Albert K. L. Tsang. The falling-leaves mosaic and its equilibrium properties. Advances in Applied Probability, 26(1):54–62, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427578.

Cont:2008:SWG

[CT08] Rama Cont and Emily Tanimura. Small-world graphs: Characterization and alternative constructions. Advances in Applied Probability, 40(4):939–965, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443619.

Cabana:2018:LDRa

[CT18a] Tanguy Cabana and Jonathan D. Touboul. Large deviations for randomly connected neural networks: I. Spatially extended systems. Advances in Applied Probability, 50 (3):944–982, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/large-deviations-for-randomly-connected-
neural-networks-i-spatially-extended-systems/6BE054BD0958DAD2C3622B1527DADDC2.

Cabana:2018:LDRb

[CT18b] Tanguy Cabana and Jonathan D. Touboul. Large deviations for randomly connected neural networks: II. Statedependent interactions. Advances in Applied Probability, 50 (3):983–1004, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/large-deviations-for-randomly-connected-
neural-networks-ii-statedependent-interactions/DC0C97934074F7EAAC09B5CDB6A73A36.

Cocozza-Thivent:2000:GFS

[CTR00] Christiane Cocozza-Thivent and Michel Roussignol. A general framework for some asymptotic reliability formulas. Advances in Applied Probability, 32(2):446–467, June 2000. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428198.

Cannings:1976:RDL

[CTS76] C. Cannings, E. A. Thompson, and H. H. Skolnick. The recursive derivation of likelihoods on complex pedigrees. Advances in Applied Probability, 8(4):622–625, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425918.

Cannings:1978:PFC

[CTS78] C. Cannings, E. A. Thompson, and M. H. Skolnick. Probability functions on complex pedigrees. Advances in Applied Probability, 10(1):26–61, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426718.

Curnow:1974:SCU

[Cur74] R. N. Curnow. Selecting crosses using information from a diallel cross. Advances in Applied Probability, 6(1):6–7, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426198.

Curnow:1981:PTU

[Cur81] R. N. Curnow. Progeny testing using Presence/Absence data. Advances in Applied Probability, 13(1):1–2, March 1981. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426452.

Cotar:2004:NLM

[CV04] Codina Cotar and Stanislav Volkov. A note on the lilypond model. Advances in Applied Probability, 36(2):325–339, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428455.

Champagnat:2018:UCC

[CV18] Nicolas Champagnat and Denis Villemonais. Uniform convergence of conditional distributions for absorbed one-dimensional diffusions. Advances in Applied Probability, 50(1):178–203, March 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/uniform-
convergence-of-conditional-distributions-for-absorbed-
onedimensional-diffusions/68C293CD2D1957DC35BCCD32DCE8433F.

Cordero:2023:MMW

[CV23] Fernando Cordero and Grégoire Véchambre. Moran models and Wright–Fisher diffusions with selection and mutation in a one-sided random environment. Advances in Applied Probability, 55(3):701–767, September 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/moran-models-and-wrightfisher-diffusions-
with-selection-and-mutation-in-a-onesided-random-environment/
95583233B46BA361659258BAD24FC1D9.

Chiu:1996:SPV

[CVS96] S. N. Chiu, R. Van De Weygaert, and D. Stoyan. The sectional Poisson Voronoi tessellation is not a Voronoi tessellation. Advances in Applied Probability, 28(2):356–376, June 1996. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428061.

Charlesworth:1975:CSA

[CW75] Brian Charlesworth and John A. Williamson. The chance of survival of an advantageous mutant in an age-structured population. Advances in Applied Probability, 7(1):5–7, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425842.

Cekanavicius:2003:CPA

[CW03] V. Cekanavicius and Y. H. Wang. Compound Poisson approximations for sums of discrete nonlattice variables. Advances in Applied Probability, 35(1):228–250, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428281.

Cui:2009:AMO

[CWW09] Zhaolei Cui, Yuebao Wang, and Kaiyong Wang. Asymptotics for the moments of the overshoot and undershoot of a random walk. Advances in Applied Probability, 41(2):469–494, June 2009. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/27793887.

Chatterjee:2004:TDL

[CX04] Kalyan Chatterjee and Susan H. Xu. Technology diffusion by learning from neighbours. Advances in Applied Probability, 36 (2):355–376, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428457.

Csorgo:1999:WAE

[CY99] Miklós Csörgő and Hao Yu. Weak approximations for empirical Lorenz curves and their Goldie inverses of stationary observations. Advances in Applied Probability, 31(3):698–719, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428392.

Chiu:2000:TCL

[CY00] S. N. Chiu and C. C. Yin. The time of completion of a linear birth-growth model. Advances in Applied Probability, 32(3): 620–627, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428403.

Cai:2005:RPC

[CY05] Jun Cai and Hailiang Yang. Ruin in the perturbed compound Poisson risk process under interest force. Advances in Applied Probability, 37(3):819–835, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037357.

Chen:2014:ASL

[CYG14] Ge Chen, Changlong Yao, and Tiande Guo. The asymptotic size of the largest component in random geometric graphs with some applications. Advances in Applied Probability, 46(2):307–324, June 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1401369696.

Chang:1999:LDM

[CYZ99] Cheng-Shang Chang, David D. Yao, and Tim Zajic. Large deviations, moderate deviations, and queues with long-range dependent input. Advances in Applied Probability, 31(1):254– 278, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428001.

Czajkowski:2023:TEC

[Cza23] Jan Czajkowski. Thin-ended clusters in percolation in H^{d}. Advances in Applied Probability, 55(2):581–610, June 2023. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/thinended-clusters-
in-percolation-in-mathbbhd/BE09D77D45D85CCE379F99567966B3BA.

Chen:2004:BDP

[CZLR04] Anyue Chen, Hanjun Zhang, Kai Liu, and Keith Rennolls. Birth–death processes with disaster and instantaneous resurrection. Advances in Applied Probability, 36(1):267–292, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428362.

Daduna:1982:PTO

[Dad82] Hans Daduna. Passage times for overtake-free paths in Gordon– Newell networks. Advances in Applied Probability, 14(3):672– 686, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426680.

Daduna:1984:BTP

[Dad84] Hans Daduna. Burke’s theorem on passage times in Gordon– Newell networks. Advances in Applied Probability, 16(4):867– 886, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427344.

Dagsvik:1975:GBQa

[Dag75a] John Dagsvik. The general bulk queue as a matrix factorisation problem of the Wiener–Hopf type. Part I. Advances in Applied Probability, 7(3):636–646, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426132.

Dagsvik:1975:GBQb

[Dag75b] John Dagsvik. The general bulk queue as a matrix factorisation problem of the Wiener–Hopf type. Part II. Advances in Applied Probability, 7(3):647–655, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426133.

Dagsvik:1978:SHT

[Dag78] John Dagsvik. Stability and heavy traffic results for the general bulk queue. Advances in Applied Probability, 10(1):213– 231, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426726.

Dorit:1995:APZ

[DAG95] R. L. Dorit, H. Akashi, and W. Gilberg. Absence of polymorphism at the ZFY locus on the human Y chromosome. Science, 268(5214):1183–1185, May 26, 1995. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/content/268/5214/1183. See reply

[Cha99b].

Dai:2008:PSM

[Dai08] Hongsheng Dai. Perfect sampling methods for random forests. Advances in Applied Probability, 40(3):897–917, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/20443613.

Dai:2011:EMC

[Dai11] Hongsheng Dai. Exact Monte Carlo simulation for fork-join networks. Advances in Applied Probability, 43(2):484–503, June 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/23024516.

Daley:1976:QOP

[Dal76] D. J. Daley. Queueing output processes. Advances in Applied Probability, 8(2):395–415, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425911.

Daley:1985:WTS

[Dal85] D. J. Daley. What is a two-sex population process? Advances in Applied Probability, 17(1):230, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427061. See corrections [Ano95e].

Daniels:1971:NPT

[Dan71] H. E. Daniels. A note on perturbation techniques for epidemics. Advances in Applied Probability, 3(2):214–218, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426162.

Daniels:1974:ATC

[Dan74a] H. E. Daniels. An approximation technique for a curved boundary problem. Advances in Applied Probability, 6(2):194– 196, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426243.

Daniels:1974:MSC

[Dan74b] H. E. Daniels. The maximum size of a closed epidemic. Advances in Applied Probability, 6(4):607–621, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426182.

Daniels:1977:AWS

[Dan77] H. E. Daniels. The advancing wave in a spatial birth process. Advances in Applied Probability, 9(2):191–192, June 1977. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426339.

Daniels:1989:MGP

[Dan89] H. E. Daniels. The maximum of a Gaussian process whose mean path has a maximum, with an application to the strength of bundles of fibres. Advances in Applied Probability, 21(2):315– 333, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427162.

Davis:1970:SGB

[Dav70] A. W. Davis. Some generalizations of Bailey’s birth death and migration model. Advances in Applied Probability, 2(1):83– 109, Spring 1970. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
3518345.

Davies:1973:EIC

[Dav73a] P. L. Davies. Everywhere irregularity of certain classes of random processes with stationary Gaussian increments. Advances in Applied Probability, 5(1):103–118, April 1973. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425966.

Davies:1973:AIS

[Dav73b] Robert B. Davies. Asymptotic inference in stationary Gaussian time-series. Advances in Applied Probability, 5(3):469–497, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425830.

Davies:1977:THP

[Dav77] Robert B. Davies. Testing the hypothesis that a point process is Poisson. Advances in Applied Probability, 9(4):724–746, December 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426698.

Davies:1979:INE

[Dav79] R. B. Davies. Inference from non-ergodic time series. Advances in Applied Probability, 11(2):261–262, June 1979. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426792.

Dawson:1975:QDS

[Daw75] D. A. Dawson. Quasi-deterministic states for Markov systems. Advances in Applied Probability, 7(2):231–232, June 1975. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426047.

Deistler:1978:VLT

[DDH78] M. Deistler, W. Dunsmuir, and E. J. Hannan. Vector linear time series models: Corrections and extensions. Advances in Applied Probability, 10(2):360–372, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426940. See [DH76].

Davis:1987:OCE

[DDSV87] M. H. A. Davis, M. A. H. Dempster, S. P. Sethi, and D. Vermes. Optimal capacity expansion under uncertainty. Advances in Applied Probability, 19(1):156–176, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427378.

DeSmit:1973:MSQ

[De 73a] J. H. A. De Smit. On the many server queue with exponential service times. Advances in Applied Probability, 5(1):170– 182, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425970.

DeSmit:1973:SGR

[De 73b] J. H. A. De Smit. Some general results for many server queues. Advances in Applied Probability, 5(1):153–169, April 1973. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425969.

DeSmit:1973:RRM

[De 73c] Jos H. A. De Smit. Recent results for many server queues. Advances in Applied Probability, 5(1):11–12, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425947.

DeSmit:1974:MSQ

[De 74] J. H. A. De Smit. The many server queue in discrete time. Advances in Applied Probability, 6(2):253, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426288.

DeSmit:1976:PAP

[De 76] J. H. A. De Smit. A probabilistic approach to Pollaczek’s integral equations for the many-server queue. Advances in Applied Probability, 8(2):243, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425891.

DeSmit:1977:ANM

[De 77] J. H. A. De Smit. Analytical and numerical methods for many server queues. Advances in Applied Probability, 9(2):192– 193, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426340.

DeSmit:1980:PAM

[De 80] Jos H. A. De Smit. A probabilistic approach to the multipleserver queue. Advances in Applied Probability, 12(2):279, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426549.

DeSmit:1983:QGM

[De 83a] Jos H. A. De Smit. The queue GI/M/s with customers of different types or the queue GI/H_{m} /s. Advances in Applied Probability, 15(2):392–419, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426442.

DeSouzaBorges:1983:LDF

[De 83b] Wagner De Souza Borges. On the limiting distribution of the failure time of fibrous materials. Advances in Applied Probability, 15(2):331–348, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426439.

DeGregorio:2010:SVM

[De 10] Alessandro De Gregorio. Stochastic velocity motions and processes with random time. Advances in Applied Probability, 42 (4):1028–1056, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/25799083.

Degen:2008:EBE

[DE08] Matthias Degen and Paul Embrechts. EVT-Based estimation of risk capital and convergence of high quantiles. Advances in Applied Probability, 40(3):696–715, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443604.

Dayanik:2012:OSP

[DE12] Savas Dayanik and Masahiko Egami. Optimal stopping problems for asset management. Advances in Applied Probability, 44 (3):655–677, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/41714070.

Deb:1976:OCB

[Deb76] Rajat K. Deb. Optimal control of batch service queues with switching costs. Advances in Applied Probability, 8(1):177–194, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426028.

Dehoff:1998:TMC

[Deh98] Robert T. Dehoff. Topological and metric changes during tessellation evolution in two dimensional grain growth. Advances in Applied Probability, 30(2):277–278, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427947.

Deijfen:2003:ASC

[Dei03] Maria Deijfen. Asymptotic shape in a continuum growth model. Advances in Applied Probability, 35(2):303–318, June 2003. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428422.

Demichel:2011:HFC

[DEKS11] Yann Demichel, Anne Estrade, Marie Kratz, and Gennady Samorodnitsky. How fast can the chord length distribution decay? Advances in Applied Probability, 43(2):504–523, June 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/23024517.

Donnelly:1991:FRM

[DEP91] P. J. Donnelly, W. J. Ewens, and S. Padmadisastra. Functionals of random mappings: Exact and asymptotic results. Advances in Applied Probability, 23(3):437–455, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427616.

Derriennic:1983:RTC

[Der83] Yves Derriennic. A remark on the time constant in first-passage percolation. Advances in Applied Probability, 15(1):214–215, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426992.

Dereudre:2009:EQI

[Der09] David Dereudre. The existence of quermass-interaction processes for nonlocally stable interaction and nonbounded convex grains. Advances in Applied Probability, 41(3):664–681, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
27793898.

DeAngelis:2018:SNS

[DF18] Tiziano De Angelis and Giorgio Ferrari. Stochastic nonzerosum games: a new connection between singular control and optimal stopping. Advances in Applied Probability, 50(2):347–372, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/stochastic-nonzerosum-games-a-new-
connection-between-singular-control-and-optimal-stopping/
B9D231CDFFC18881871E3F83077B72AF.

Decreusefond:2014:SHR

[DFRV14] L. Decreusefond, E. Ferraz, H. Randriambololona, and A. Vergne. Simplicial homology of random configurations. Advances in Applied Probability, 46(2):325–347, June 2014. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1401369697.

DeLeve:1977:GMDa

[DFT77a] G. De Leve, A. Federgruen, and H. C. Tijms. A general Markov decision method I: Model and techniques. Advances in Applied Probability, 9(2):296–315, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426388.

DeLeve:1977:GMDb

[DFT77b] G. De Leve, A. Federgruen, and H. C. Tijms. A general Markov decision method II: Applications. Advances in Applied Probability, 9(2):316–335, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426389.

Ding:2019:RAK

[DFZ19] Weiyong Ding, Rui Fang, and Peng Zhao. Reliability analysis of k-out-of-n systems based on a grouping of components. Advances in Applied Probability, 51(2):339–357, June 2019. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/reliability-
analysis-of-koutofn-systems-based-on-a-grouping-of-components/
CE7A30DED0C4BC48B737D96D422AA3C2.

Dunn:2001:PIB

[DG01] R. T. Dunn and K. D. Glazebrook. The performance of indexbased policies for bandit problems with stochastic machine availability. Advances in Applied Probability, 33(2):365–390, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428258.

Doney:2003:OCB

[DG03] R. A. Doney and P. S. Griffin. Overshoots over curved boundaries. Advances in Applied Probability, 35(2):417–448, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428430.

DeIorio:2004:ISCa

[DG04a] Maria De Iorio and Robert C. Griffiths. Importance sampling on coalescent histories. I. Advances in Applied Probability, 36 (2):417–433, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428460.

DeIorio:2004:ISCb

[DG04b] Maria De Iorio and Robert C. Griffiths. Importance sampling on coalescent histories. II: Subdivided population models. Advances in Applied Probability, 36(2):434–454, June 2004. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428461.

Doney:2004:OCB

[DG04c] R. A. Doney and P. S. Griffin. Overshoots over curved boundaries. II. Advances in Applied Probability, 36(4):1148–1174, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
4140392.

Dawson:2018:RSU

[DG18] D. A. Dawson and L. G. Gorostiza. Random systems in ultrametric spaces. Advances in Applied Probability, 50 (A):83–97, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/random-systems-in-ultrametric-spaces/
CE9945603EB31F0C46492F596C3DF04D.

Delleur:1984:SH

[DGB^{+}84] J. W. Delleur, Vijay K. Gupta, R. N. Bhattacharya, V. Klemes, Richard L. Smith, P. Todorovic, Stuart J. Deutsch, José A. Ramos, Van-Thanh-Van Nguyen, David K. Pickard, Elmer M. Tory, James A. Smith, Alan F. Karr, and Samuel W. Woolford. Stochastic hydrology. Advances in Applied Probability, 16 (1):17–23, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427220.

Devroye:2009:EMD

[DGM09] Luc Devroye, Joachim Gudmundsson, and Pat Morin. On the expected maximum degree of Gabriel and Yao graphs. Advances in Applied Probability, 41(4):1123–1140, December 2009. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/27793922.

DiCrescenzo:1992:SFC

[DGN92] A. Di Crescenzo, V. Giorno, and A. G. Nobile. On some firstcrossing-time probabilities for a two-dimensional random walk with correlated components. Advances in Applied Probability, 24(2):441–454, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427699.

DeSantis:2010:LDC

[DGP10] Emilio De Santis, Fabrizio Grandoni, and Alessandro Panconesi. Low degree connectivity of ad-hoc networks via percolation. Advances in Applied Probability, 42(2):559–576, June 2010. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/25683833.

Dumas:2002:MAA

[DGR02] Vincent Dumas, Fabrice Guillemin, and Philippe Robert. A Markovian analysis of additive-increase multiplicative-decrease algorithms. Advances in Applied Probability, 34(1):85–111, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428373.

Dunsmuir:1976:VLT

[DH76] W. Dunsmuir and E. J. Hannan. Vector linear time series models. Advances in Applied Probability, 8(2):339–364, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425908. See

corrections [DDH78].

Daley:1982:TES

[DH82] D. J. Daley and J. Haslett. A thermal energy storage process with controlled input. Advances in Applied Probability, 14 (2):257–271, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426520.

Dawson:1991:MBM

[DH91] D. A. Dawson and K. J. Hochberg. A multilevel branching model. Advances in Applied Probability, 23(4):701–715, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427671.

DSouza:1997:SPB

[DH97] J. C. D’Souza and B. M. Hambly. On the survival probability of a branching process in a random environment. Advances in Applied Probability, 29(1):38–55, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427860.

Deijfen:2004:CTT

[DH04] Maria Deijfen and Olle Häggström. Coexistence in a two-type continuum growth model. Advances in Applied Probability, 36 (4):973–980, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/4140384.

Drmota:2005:PRT

[DH05] Michael Drmota and Hsien-Kuei Hwang. Profiles of random trees: Correlation and width of random recursive trees and binary search trees. Advances in Applied Probability, 37(2):321– 341, June 2005. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
30037330.

Dishi:1984:CS

[DHK^{+}84] B. T. Dishi, H. Heffes, Alan G. Konheim, Guy Latouche, Yukio Takahashi, Søren Asmussen, Yoshifusa Ito, G. Koch, P. J. C. Spreij, F. J. Massey III, and D. F. Miller. Computers and simulation. Advances in Applied Probability, 16(1):10–13, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427218.

Debicki:2017:UTA

[DHL17] Krzysztof Debicki, Enkelejd Hashorva, and Peng Liu. Uniform tail approximation of homogeneous functionals of Gaussian fields. Advances in Applied Probability, 49(4):1037–1066, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/uniform-
tail-approximation-of-homogenous-functionals-of-gaussian-
fields/AA99410F8868AB423AE785B7D3B89C1A.

Dufour:2012:ETC

[DHP12] François Dufour, M. Horiguchi, and A. B. Piunovskiy. The expected total cost criterion for Markov decision processes under constraints: a convex analytic approach. Advances in Applied Probability, 44(3):774–793, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/41714075.

DiCrescenzo:2001:RMV

[Di 01] Antonio Di Crescenzo. On random motions with velocities alternating at Erlang-distributed random times. Advances in Applied Probability, 33(3):690–701, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428482.

Dia:2013:EBS

[Dia13] E. H. A. Dia. Error bounds for small jumps of Lévy processes. Advances in Applied Probability, 45(1):86–105, March 2013. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/41806001.

Dietz:1971:MM

[Die71] Klaus Dietz. Malaria models. Advances in Applied Probability, 3 (2):208–210, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426159.

Dies:1982:HLT

[Die82a] Jacques-Edouard Dies. Hendricks libraries and Tsetlin piles. Advances in Applied Probability, 14(1):37–55, March 1982. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426732.

Dies:1982:TCC

[Die82b] Jacques-Edouard Dies. Sur la transience de certaines chaines de Markov sur les permutations. Advances in Applied Probability, 14(3):526–542, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426672.

Dieker:2006:AFE

[Die06] A. B. Dieker. Applications of factorization embeddings for Lévy processes. Advances in Applied Probability, 38(3):768–791, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443470.

Ding:1994:LTC

[Din94] Xinhong Ding. A limit theorem for certain disordered random systems. Advances in Applied Probability, 26(4):1022–1043, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427903.

Ding:2019:SVD

[Din19] Xiucai Ding. Singular vector distribution of sample covariance matrices. Advances in Applied Probability, 51 (1):236–267, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/singular-vector-distribution-of-
sample-covariance-matrices/67B20ABE4FE8FCA3DE140DD9086F8918.

Dion:1976:CVF

[Dio76] Jean-Guy Dion. Sur une caractérisation des vidanges fortes. (French) [On a characterization of strong drains]. Advances in Applied Probability, 8(1):105–126, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426024. See correc-

tions [Dio78].

Dion:1978:CVF

[Dio78] Jean-Guy Dion. Sur une caractérisation des vidanges fortes: Corrections and extensions. (French) [On a characterization of strong drains: corrections and extensions]. Journal of Applied Probability, 15(2):268–279, June 1978. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.

jstor.org/stable/3213400. See [Dio76].

Donnelly:1991:COS

[DJ91] Peter Donnelly and Paul Joyce. Consistent ordered sampling distributions: Characterization and convergence. Advances in Applied Probability, 23(2):229–258, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427746.

Dupuis:2017:MDB

[DJ17] Paul Dupuis and Dane Johnson. Moderate deviations-based importance sampling for stochastic recursive equations. Advances in Applied Probability, 49(4):981–1010, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/moderate-deviationsbased-
importance-sampling-for-stochastic-recursive-equations/
0BC81B8BDF999F5246A5F13ACE7D5848.

Doss:1980:SVD

[DK80] Halim Doss and Thomas G. Kurtz. Stochastic versus deterministic approaches. Advances in Applied Probability, 12(3):569–570, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426419.

Deshpande:1983:DOS

[DK83] J. V. Deshpande and S. C. Kochar. Dispersive ordering is the same as tail-ordering. Advances in Applied Probability, 15(3): 686–687, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426626.

Deshpande:1997:BJS

[DK97] J. V. Deshpande and S. R. Karia. Bounds for the joint survival and incidence functions through coherent system data. Advances in Applied Probability, 29(2):478–497, June 1997. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428013.

Dekking:2012:SSK

[DK12] Michel Dekking and Derong Kong. A simple stochastic kinetic transport model. Advances in Applied Probability, 44(3):874– 885, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/41714079.

Dolinsky:2016:RMG

[DK16] Yan Dolinsky and Yuri Kifer. Risk minimization for game options in markets imposing minimal transaction costs. Advances in Applied Probability, 48(3):926–946, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1474296321.

Daley:2010:DST

[DKMJ10] Daryl Daley, Wilfrid S. Kendall, Ilya Molchanov, and Eva Vedel Jensen. Dietrich Stoyan: a tribute on the occasion of his seventieth birthday. Advances in Applied Probability, 42(3):??, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25799059.

Disney:1984:SQL

[DKS84] Ralph L. Disney, Dieter König, and Volker Schmidt. Stationary queue-length and waiting-time distributions in single-server feedback queues. Advances in Applied Probability, 16(2):437– 446, June 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427078.

Dehon:1982:GIR

[DL82] Michel Dehon and Guy Latouche. A geometric interpretation of the relations between the exponential and generalized Erlang distributions. Advances in Applied Probability, 14(4):885–897, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427029.

Daley:2005:DCL

[DL05] Daryl J. Daley and Günter Last. Descending chains, the lilypond model, and mutual-nearest-neighbour matching. Advances in Applied Probability, 37(3):604–628, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037346.

Dia:2011:CDC

[DL11] E. H. A. Dia and D. Lamberton. Connecting discrete and continuous lookback or hindsight options in exponential Lévy models. Advances in Applied Probability, 43(4):1136–1165, December 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/23208042.

Debicki:2018:ENG

[DL18] Krzysztof Debicki and Peng Liu. Extremes of nonstationary Gaussian fluid queues. Advances in Applied Probability, 50(3):887–917, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/extremes-of-nonstationary-gaussian-
fluid-queues/997F20DC349C2C76E673C0DF18E1E4FB.

Dassios:2020:EAF

[DL20] Angelos Dassios and Luting Li. Explicit asymptotics on first passage times of diffusion processes. Advances in Applied Probability, 52(2):681–704, June 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/explicit-asymptotics-on-
first-passage-times-of-diffusion-processes/0796C7466C3E03364A4D3DD4EAC95A92.

delaIglesia:2022:SAB

[dlI22] Manuel D. de la Iglesia. Spectral analysis of bilateral birthdeath processes: some new explicit examples. Advances in Applied Probability, 54(4):1193–1221, December 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/spectral-analysis-of-bilateral-
birthdeath-processes-some-new-explicit-examples/AE2A4C19811549A9D4E0AA67BCA562B3.

Doukhan:2006:SIV

[DLO06] P. Doukhan, A. Latour, and D. Oraichi. A simple integer-valued bilinear time series model. Advances in Applied Probability, 38 (2):559–578, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443456.

Donnelly:1994:ASB

[DLS94] Peter Donnelly, Peter Lloyd, and Aidan Sudbury. Approach to stationarity of the Bernoulli–Laplace diffusion model. Advances in Applied Probability, 26(3):715–727, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427817.

Denuit:2000:CA

[DLS00] Michel Denuit, Claude Lefèvre, and Moshe Shaked. On s-convex approximations. Advances in Applied Probability, 32(4):994– 1010, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428513.

Delgado:2004:LCS

[DLS04] Rosario Delgado, F. Javier López, and Gerardo Sanz. Local conditions for the stochastic comparison of particle systems. Advances in Applied Probability, 36(4):1252–1277, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/4140397.

Daly:2012:SMS

[DLU12] Fraser Daly, Claude Lefèvre, and Sergey Utev. Stein’s method and stochastic orderings. Advances in Applied Probability, 44 (2):343–372, June 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/41714054.

Donati-Martin:1990:PBS

[DM90] C. Donati-Martin. Le problème de Buffon–Synge pour une corde. (French) [The Buffon–Synge problem for a string]. Advances in Applied Probability, 22(2):375–395, June 1990. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427541.

Dryden:1991:GSD

[DM91] I. L. Dryden and K. V. Mardia. General shape distributions in a plane. Advances in Applied Probability, 23(2):259–276, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427747.

Doney:2000:RWC

[DM00] R. A. Doney and R. A. Maller. Random walks crossing curved boundaries: A functional limit theorem, stability and asymptotic distributions for exit times and positions. Advances in Applied Probability, 32(4):1117–1149, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428520.

Dekking:2003:ASR

[DM03] F. M. Dekking and L. E. Meester. An almost sure result for path lengths in binary search trees. Advances in Applied Probability, 35(2):363–376, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428427.

Dieker:2005:AES

[DM05] A. B. Dieker and M. Mandjes. On asymptotically efficient simulation of large deviation probabilities. Advances in Applied Probability, 37(2):539–552, June 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/30037340.

Deijfen:2006:GSR

[DM06] Maria Deijfen and Ronald Meester. Generating stationary random graphs on Z with prescribed independent, identically distributed degrees. Advances in Applied Probability, 38(2):287– 298, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443443.

Duval:2022:NAC

[DM22] Céline Duval and Ester Mariucci. Non-asymptotic control of the cumulative distribution function of Lévy processes. Advances in Applied Probability, 54(3):913–944, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/nonasymptotic-control-of-
the-cumulative-distribution-function-of-levy-processes/
913D47D423BFE4CA1EDB464CC421DEE4.

Dombry:2011:PDP

[DMB11] Clément Dombry, Christian Mazza, and Vincent Bansaye. Phenotypic diversity and population growth in a fluctuating environment. Advances in Applied Probability, 43(2):375–398, June 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/23024511.

Diggle:2000:CBP

[DMC00] Peter J. Diggle, Jorge Mateu, and Helen E. Clough. A comparison between parametric and non-parametric approaches to the analysis of replicated spatial point patterns. Advances in Applied Probability, 32(2):331–343, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428191.

Dereich:2012:TDU

[DMM12] Steffen Dereich, Christian Mönch, and Peter Mörters. Typical distances in ultrasmall random networks. Advances in Applied Probability, 44(2):583–601, June 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/41714064.

Diaz:2009:PEF

[DMPG09] J. Díaz, D. Mitsche, and X. Pérez-Giménez. On the probability of the existence of fixed-size components in random geometric graphs. Advances in Applied Probability, 41(2):344–357, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/27793881.

Diaz:2016:RBG

[DMPPG16] J. Díaz, D. Mitsche, G. Perarnau, and X. Pérez-Giménez. On the relation between graph distance and Euclidean distance in random geometric graphs. Advances in Applied Probability, 48(3):

848–864, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1474296318.

Davis:1988:ASL

[DMR88] Richard A. Davis, Edward Mulrow, and Sidney I. Resnick. Almost sure limit sets of random samples in R^{d}. Advances in Applied Probability, 20(3):573–599, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427036.

Das:2013:LME

[DMR13] Bikramjit Das, Abhimanyu Mitra, and Sidney Resnick. Living on the multidimensional edge: Seeking hidden risks using regular variation. Advances in Applied Probability, 45(1):139– 163, March 2013. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
41806003.

Donati-Martin:1993:SEQ

[DMY93] C. Donati-Martin and M. Yor. On some examples of quadratic functionals of Brownian motion. Advances in Applied Probability, 25(3):570–584, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427524.

Nunno:2023:SSU

[DMYT23] Giulia Di Nunno, Yuliya Mishura, and Anton Yurchenko-Tytarenko. Sandwiched SDEs with unbounded drift driven by Hölder noises. Advances in Applied Probability, 55 (3):927–964, September 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/sandwiched-sdes-with-unbounded-drift-
driven-by-holder-noises/B72D913C3504E8F76010D88C3695669D.

Dalen:1978:TWT

[DN78] G. Dalen and Bent Natvig. On the transient waiting time for GI/M/1 with a ’head of the line’ priority discipline. Advances in Applied Probability, 10(2):315–316, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426915.

Devroye:2002:DAE

[DN02] Luc Devroye and Ralph Neininger. Density approximation and exact simulation of random variables that are solutions of fixedpoint equations. Advances in Applied Probability, 34(2):441– 468, June 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428296.

Davydov:2010:PPA

[DNP10] Y. Davydov, A. Nagaev, and A. Philippe. On the peeling procedure applied to a Poisson point process. Advances in Applied Probability, 42(3):620–630, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/25799061.

DiNardo:2001:CAF

[DNPR01] E. Di Nardo, A. G. Nobile, E. Pirozzi, and L. M. Ricciardi. A computational approach to first-passage-time problems for Gauss–Markov processes. Advances in Applied Probability, 33 (2):453–482, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428263.

Daley:1974:RWP

[DO74] D. J. Daley and D. Oakes. Random walk point processes. Advances in Applied Probability, 6(2):225–226, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426263.

Dobrushin:1977:PFG

[Dob77] R. L. Dobrushin. Point fields from the Gibbsian point of view. Advances in Applied Probability, 9(3):444–445, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426104.

Dombry:2007:WRW

[Dom07] C. Dombry. A weighted random walk model, with application to a genetic algorithm. Advances in Applied Probability, 39(2):550– 568, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443526.

Donnelly:1981:LCG

[Don81] K. P. Donnelly. Linkage, complete genetic information and the probability of relationships being detectable. Advances in Applied Probability, 13(1):2, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426453.

Doshi:1978:OCS

[Dos78] Bharat T. Doshi. Optimal control of the service rate in an M/G/1 queueing system. Advances in Applied Probability, 10 (3):682–701, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426641.

Doubleday:1975:WCG

[Dou75] W. G. Doubleday. Weak convergence of Galton–Watson processes. Advances in Applied Probability, 7(2):250–251, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426062.

Downham:1974:MC

[Dow74] D. Y. Downham. Models of Clines. Advances in Applied Probability, 6(1):7–10, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426199.

Deistler:1984:BLF

[DP84] M. Deistler and B. M. Pötscher. The behaviour of the likelihood function for ARMA models. Advances in Applied Probability, 16 (4):843–866, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427343.

Doumas:2012:CCP

[DP12] Aristides V. Doumas and Vassilis G. Papanicolaou. The coupon collector’s problem revisited: Asymptotics of the variance. Advances in Applied Probability, 44(1):166–195, March 2012. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/23207860.

Dufour:2013:ETC

[DP13] François Dufour and A. B. Piunovskiy. The expected total cost criterion for Markov decision processes under constraints.

Advances in Applied Probability, 45(3):837–859, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1377868541.

Dufour:2015:ICC

[DP15] François Dufour and Alexei B. Piunovskiy. Impulsive control for continuous-time Markov decision processes. Advances in Applied Probability, 47(1):106–127, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1427814583.

Daw:2019:NPE

[DP19] Andrew Daw and Jamol Pender. New perspectives on the Erlang-A queue. Advances in Applied Probability, 51 (1):268–299, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/new-perspectives-on-the-erlanga-
queue/EF38B952E8804582DF0CDE81C97C289E.

Daw:2022:ESE

[DP22] Andrew Daw and Jamol Pender. An ephemerally self-exciting point process. Advances in Applied Probability, 54(2):340–403, June 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/an-
ephemerally-selfexciting-point-process/2694EDF67196138AEC21B19549AEFCB4.

Daw:2023:MCM

[DP23] Andrew Daw and Jamol Pender. Matrix calculations for moments of Markov processes. Advances in Applied Probability, 55(1):126–150, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/matrix-calculations-for-moments-
of-markov-processes/C256F64BE857BD57EFD70D5C3C32E666.

Dayanik:2008:IPD

[DPY08] Savas Dayanik, Warren Powell, and Kazutoshi Yamazaki. Index policies for discounted bandit problems with availability constraints. Advances in Applied Probability, 40(2):377–400, June 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/20443588.

Dunsmuir:1981:PES

[DR81] W. Dunsmuir and P. M. Robinson. Parametric estimators for stationary time series with missing observations. Advances in Applied Probability, 13(1):129–146, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426471.

Darden:1985:DFM

[DR85a] Thomas A. Darden and Michael A. Resnick. DNA fragment mass distributions following molecular recombination. Advances in Applied Probability, 17(3):496–513, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427117.

Davis:1985:TEM

[DR85b] Richard Davis and Sidney Resnick. Tail estimates motivated by extreme-value theory. Advances in Applied Probability, 17 (2):254–256, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427138.

Davis:1989:BPP

[DR89] Richard A. Davis and Sidney I. Resnick. Basic properties and prediction of Max-ARMA processes. Advances in Applied Probability, 21(4):781–803, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427767.

Daley:1992:LTA

[DR92] D. J. Daley and T. Rolski. Light traffic approximations in many-server queues. Advances in Applied Probability, 24(1): 202–218, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427736.

Duchesne:2003:CLR

[DR03] Thierry Duchesne and Jeffrey S. Rosenthal. On the collapsibility of lifetime regression models. Advances in Applied Probability, 35(3):755–772, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428217.

DAuria:2006:DNM

[DR06] Bernardo D’Auria and Sidney I. Resnick. Data network models of burstiness. Advances in Applied Probability, 38(2):373– 404, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443447.

DAuria:2008:IDD

[DR08] Bernardo D’Auria and Sidney I. Resnick. The influence of dependence on data network models. Advances in Applied Probability, 40(1):60–94, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443571.

Ding:2016:OLN

[DR16] Zi Ding and Ilya O. Ryzhov. Optimal learning with non-Gaussian rewards. Advances in Applied Probability, 48(1): 112–136, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1457466158.

Drosen:1986:PJS

[Dro86] James W. Drosen. Pure jump shock models in reliability. Advances in Applied Probability, 18(2):423–440, June 1986. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427307.

Dror:1989:SPG

[Dro89] Moshe Dror. Simple proof for goofspiel: The game of pure strategy. Advances in Applied Probability, 21(3):711–712, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427646.

Dasgupta:2011:GPB

[DRS11] Amites Dasgupta, Rahul Roy, and Anish Sarkar. Geometry of the Poisson Boolean model on a region of logarithmic width in the plane. Advances in Applied Probability, 43(3):616–635, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
23024567.

Daley:2000:LRD

[DRV00] D. J. Daley, T. Rolski, and Rein Vesilo. Long-range dependent point processes and their Palm–Khinchin distributions. Advances in Applied Probability, 32(4):1051–1063, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428516.

Deb:1973:OCB

[DS73] Rajat K. Deb and Richard F. Serfozo. Optimal control of batch service queues. Advances in Applied Probability, 5(2):340–361, August 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426040.

Downham:1976:IEA

[DS76a] D. Y. Downham and S. M. M. Shah. The integral equation approach for models of Clines. Advances in Applied Probability, 8(4):625–628, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1425919.

Downham:1976:SCS

[DS76b] D. Y. Downham and S. M. M. Shah. A sufficiency condition for the stability of an equilibrium. Advances in Applied Probability, 8(1):4–7, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426010.

Daniels:1985:MRW

[DS85] H. E. Daniels and T. H. R. Skyrme. The maximum of a random walk whose mean path has a maximum. Advances in Applied Probability, 17(1):85–99, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427054. See correction [Tjø78].

Daley:1993:TPM

[DS93] D. J. Daley and L. D. Servi. A two-point Markov chain boundary-value problem. Advances in Applied Probability, 25 (3):607–630, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427526.

DSouza:1994:RGG

[D’S94] J. C. D’Souza. The rates of growth of the Galton–Watson process in varying environments. Advances in Applied Probability, 26(3):698–714, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427816.

Daduna:1995:DMN

[DS95] H. Daduna and R. Szekli. Dependencies in Markovian networks. Advances in Applied Probability, 27(1):226–254, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428105.

Dumas:2000:ABF

[DS00] V. Dumas and A. Simonian. Asymptotic bounds for the fluid queue fed by sub-exponential on/off sources. Advances in Applied Probability, 32(1):244–255, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428236.

DallAglio:2003:ZLD

[DS03] Marco Dall’Aglio and Marco Scarsini. Zonoids, linear dependence, and size-biased distributions on the simplex. Advances in Applied Probability, 35(4):871–884, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428489.

Denisov:2007:LAC

[DS07] Denis Denisov and Vsevolod Shneer. Local asymptotics of the cycle maximum of a heavy-tailed random walk. Advances in Applied Probability, 39(1):221–244, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443507.

Dieker:2015:RLQ

[DS15] A. B. Dieker and T. Suk. Randomized longest-queue-first scheduling for large-scale buffered systems. Advances in Applied Probability, 47(4):1015–1038, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1449859798.

Daduna:2016:CFM

[DS16] Hans Daduna and Ryszard Szekli. Correlation formulas for Markovian network processes in a random environment. Advances in Applied Probability, 48(1):176–198, March 2016. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1457466161.

DeSantis:2023:CAP

[DS23] Emilio De Santis and Fabio Spizzichino. Construction of aggregation paradoxes through load-sharing models. Advances in Applied Probability, 55(1):223–244, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/construction-of-aggregation-
paradoxes-through-loadsharing-models/08A262CA3B238D82CDF5332EAF690279.

Dshalalow:1985:MQF

[Dsh85] Jewgeni Dshalalow. On the multiserver queue with finite waiting room and controlled input. Advances in Applied Probability, 17 (2):408–423, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427148.

Daley:1999:VFP

[DSS99] D. J. Daley, H. Stoyan, and D. Stoyan. The volume fraction of a Poisson germ model with maximally non-overlapping spherical grains. Advances in Applied Probability, 31(3):610–624, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428386.

Deng:2017:SRC

[DSS17] Chang-Song Deng, René L. Schilling, and Yan-Hong Song. Subgeometric rates of convergence for Markov processes under subordination. Advances in Applied Probability, 49 (1):162–181, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/subgeometric-rates-of-convergence-
for-markov-processes-under-subordination/4F8C7D5514F3FD2DE8D64B0E35EA8329.

See correction [DSS18].

Deng:2018:SRC

[DSS18] C.-S. Deng, R. L. Schilling, and Y.-H. Song. Subgeometric rates of convergence for Markov processes under subordination — correction. Advances in Applied Probability, 50(3):1005, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/subgeometric-rates-of-convergence-
for-markov-processes-under-subordination-correction/
A0D8171B84A838869B1157B640FFE997. See [DSS17].

Dayar:2011:ILD

[DSSW11] Tuğrul Dayar, Werner Sandmann, David Spieler, and Verena Wolf. Infinite level-dependent QBD processes and matrixanalytic solutions for stochastic chemical kinetics. Advances in Applied Probability, 43(4):1005–1026, December 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/23208036.

DeMeyer:1980:ABB

[DT80] Anne-Marie De Meyer and J. L. Teugels. On the asymptotic behaviour of the busy-period distribution in M/G/1 queues. Advances in Applied Probability, 12(2):301–302, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426577.

Donnelly:1986:AAC

[DT86] Peter Donnelly and Simon Tavaré. The ages of alleles and a coalescent. Advances in Applied Probability, 18(1):1–19, March 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427237.

See correction [Ano86f].

Dao-Thi:2007:ZAQ

[DTM07] Thu-Ha Dao-Thi and Jean Mairesse. Zero-automatic queues and product form. Advances in Applied Probability, 39(2):429– 461, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443521.

DeKok:1984:ASP

[DTV84] A. G. De Kok, H. C. Tijms, and F. A. Van Der Duyn Schouten. Approximations for the single-product productioninventory problem with compound Poisson demand and servicelevel constraints. Advances in Applied Probability, 16(2):378– 401, June 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427075.

Duchamps:2021:FSS

[Duc21] Jean-Jil Duchamps. Fragmentations with self-similar branching speeds. Advances in Applied Probability, 53(4):1149–1189, December 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/fragmentations-
with-selfsimilar-branching-speeds/FA6634BD79929D39F0F449015F3D1CE8.

Dufkova:1977:COD

[Duf77] Vera Dufková. On controlled one-dimensional diffusion processes with unknown parameter. Advances in Applied Probability, 9(1):105–124, March 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425819.

Dufresne:2001:IGB

[Duf01] Daniel Dufresne. The integral of geometric Brownian motion. Advances in Applied Probability, 33(1):223–241, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428449.

Dufresne:2004:LNA

[Duf04] Daniel Dufresne. The log-normal approximation in financial and other computations. Advances in Applied Probability, 36(3): 747–773, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/4140406.

Durham:1975:OBM

[Dur75] S. D. Durham. An optimal branching migration process. Advances in Applied Probability, 7(2):251, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426063.

Durrett:1979:IPS

[Dur79] Richard Durrett. An infinite particle system with additive interactions. Advances in Applied Probability, 11(2):355–383, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426844.

Durrett:1980:IPS

[Dur80] R. Durrett. Infinite particle systems. Advances in Applied Probability, 12(2):268, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426534.

Daley:2015:BMS

[DVN15] D. J. Daley, Johan S. H. Van Leeuwaarden, and Yoni Nazarathy. BRAVO for many-server QED systems with finite buffers. Advances in Applied Probability, 47(1):231–250, March 2015. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1427814589.

Dvoretzky:1977:GM

[Dvo77] Aryeh Dvoretzky. Generalizations of martingales. Advances in Applied Probability, 9(2):193–194, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426341.

Dumbgen:1996:RCR

[DW96] Lutz Dümbgen and Günther Walther. Rates of convergence for random approximations of convex sets. Advances in Applied Probability, 28(2):384–393, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428063.

Dupuis:2002:OSR

[DW02] Paul Dupuis and Hui Wang. Optimal stopping with random intervention times. Advances in Applied Probability, 34(1):141– 157, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428376.

Dimakis:2006:SCS

[DW06] Antonis Dimakis and Jean Walrand. Sufficient conditions for stability of longest-queue-first scheduling: Second-order properties using fluid limits. Advances in Applied Probability, 38 (2):505–521, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443453.

Dwass:1974:PPD

[Dwa74] Meyer Dwass. Poisson process and distribution-free statistics. Advances in Applied Probability, 6(2):359–375, June 1974. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426298.

Donnelly:1999:DRC

[DWH^{+}99] Peter Donnelly, Carsten Wiuf, Jotun Hein, Montgomery Slatkin, W. J. Ewens, and J. F. C. Kingman. Discussion: Recent common ancestors of all present-day individuals. Advances in Applied Probability, 31(4):1027–1035, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428341.

Dhersin:2015:TLE

[DY15] Jean-Stéphane Dhersin and Linglong Yuan. On the total length of external branches for beta-coalescents. Advances in Applied Probability, 47(3):693–714, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1444308878.

Dynkin:1978:SME

[Dyn78] E. B. Dynkin. A stochastic model of economic equilibrium. Advances in Applied Probability, 10(2):263, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426872.

Dassios:2011:DCP

[DZ11a] Angelos Dassios and Hongbiao Zhao. A dynamic contagion process. Advances in Applied Probability, 43(3):814–846, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/23024576.

DeHaan:2011:ERD

[DZ11b] Laurens De Haan and Chen Zhou. Extreme residual dependence for random vectors and processes. Advances in Applied Probability, 43(1):217–242, March 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/23024533.

Ebanks:1978:BPG

[Eba78] Bruce Ebanks. The branching property in generalized information theory. Advances in Applied Probability, 10(4):788–802, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426659.

Eberl:1980:AAD

[Ebe80] W. Eberl. Asymptotic absorption distributions in birth-and-death processes. Advances in Applied Probability, 12(2):298– 299, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426574.

Ebrahimi:1990:BSF

[Ebr90] Nader Ebrahimi. Binary structure functions with dependent components. Advances in Applied Probability, 22(3):627–640, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427461.

Ebrahimi:1993:MRS

[Ebr93] Nader Ebrahimi. Modeling of repairable systems. Advances in Applied Probability, 25(4):926–938, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427799.

Evans:1975:EEP

[EC75] D. A. Evans and K. R. Clarke. Estimation of embedded particle properties from plane section intercepts. Advances in Applied Probability, 7(3):542–560, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426127.

E:2021:CSM

[EC21] Luis H. R. Alvarez E. and Sören Christensen. A class of solvable multidimensional stopping problems in the presence of Knightian uncertainty. Advances in Applied Probability, 53(2):400–424, June 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/class-of-solvable-multidimensional-

stopping-problems-in-the-presence-of-knightian-uncertainty/
36C5030B5FF5C9800743AA01AEA1B4BB.

Eckstein:2019:ELP

[Eck19] Stephan Eckstein. Extended Laplace principle for empirical measures of a Markov chain. Advances in Applied Probability, 51(1):136–167, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/extended-laplace-principle-for-empirical-
measures-of-a-markov-chain/49FA3CD7F6D05537CF199C41851ED7B3.

Economou:2005:GPF

[Eco05] Antonis Economou. Generalized product-form stationary distributions for Markov chains in random environments with queueing applications. Advances in Applied Probability, 37(1):185– 211, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
30037321.

El-Desouky:1990:GWT

[EDH90] B. S. El-Desouky and S. A. Hussen. A generalization of a waiting time problem. Advances in Applied Probability, 22(3):758–760, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427470.

Edler:1978:SSG

[Edl78] L. Edler. Strict supercritical generation-dependent Crump– Mode–Jagers branching processes. Advances in Applied Probability, 10(4):744–763, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426657.

Edmundson:1976:SEI

[Edm76] H. P. Edmundson. Statistical estimation of the inverse of a matrix. Advances in Applied Probability, 8(2):240–241, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425887.

Edwards:1974:KMP

[Edw74] A. W. F. Edwards. On Kimura’s maximum principle in the genetical theory of natural selection. Advances in Applied Probability, 6(1):10–13, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426200.

Edwards:1979:PG

[Edw79] A. W. F. Edwards. Parsing a genealogy. Advances in Applied Probability, 11(1):2–3, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426757.

El-Fattah:1980:RAA

[EF80] Y. M. El-Fattah. Recursive algorithms for adaptive control of finite Markov chains. Advances in Applied Probability, 12(2): 313, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426591.

El-Fattah:1981:GAR

[EF81] Yousri M. El-Fattah. Gradient approach for recursive estimation and control in finite Markov chains. Advances in Applied Probability, 13(4):778–803, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426973.

Ewens:1992:CP

[EFG^{+}92] W. J. Ewens, B. Fristedt, D. R. Gavelek, A. P. Godbole, J. C. Hansen, L. Holst, Z.-X. Hu, J. Jaworski, P. J. Joyce, V. I. Khokhlov, P. Matthews, E. Schmutz, W. Stadje, J. M. Steele, and P. Tetali. Contributed papers. Advances in Applied Probability, 24(4):773–777, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427711.

Eddy:1981:CHS

[EG81] William F. Eddy and James D. Gale. The convex Hull of a spherically symmetric sample. Advances in Applied Probability, 13(4):751–763, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426971.

Ethier:1990:NTL

[EG90] S. N. Ethier and R. C. Griffiths. The neutral two-locus model as a measure-valued diffusion. Advances in Applied Probability, 22(4):773–786, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427561.

Ernst:2017:ATR

[EG17] Philip A. Ernst and Ilie Grigorescu. Asymptotics for the time of ruin in the war of attrition. Advances in Applied Probability, 49(2):388–410, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/asymptotics-for-the-time-of-ruin-
in-the-war-of-attrition/2810503AA1ABF4F5D99D369D81E7AC89.

Embrechts:1979:SDF

[EGV79] Paul Embrechts, Charles M. Goldie, and N. Veraverbeke. Subexponential distribution functions and some applications. Advances in Applied Probability, 11(2):283, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426813.

Embrechts:1985:ACP

[EJMT85] Paul Embrechts, Jens L. Jensen, Makoto Maejima, and J. L. Teugels. Approximations for compound Poisson and Pólya processes. Advances in Applied Probability, 17(3):623–637, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427123.

Ethier:1981:IMN

[EK81] S. N. Ethier and Thomas G. Kurtz. The infinitely-many-neutral-alleles diffusion model. Advances in Applied Probability, 13(3): 429–452, September 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426779.

Eder:2012:PLM

[EK12] Irmingard Eder and Claudia Klüppelberg. Pareto Lévy measures and multivariate regular variation. Advances in Applied Probability, 44(1):117–138, March 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/23207858.

Ernst:2018:HTD

[EKKZ18] Philip A. Ernst, Marek Kimmel, Monika Kurpas, and Quan Zhou. Heavy-tailed distributions in branching process models of secondary cancerous tumors. Advances in Applied Probability, 50(A):99–114, December 2018. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/heavytailed-
distributions-in-branching-process-models-of-secondary-
cancerous-tumors/2C8D051FB99E174F5738298A73597AF8.

Ethier:2005:FTC

[EL05] S. N. Ethier and David A. Levin. On the fundamental theorem of card counting, with application to the game of Trente et Quarante. Advances in Applied Probability, 37(1):90–107, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/30037317.

Ekstrom:2014:ODP

[EL14] Erik Ekström and Bing Lu. The optimal dividend problem in the dual model. Advances in Applied Probability, 46(3):746– 765, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1409319558.

Elishakoff:1977:RVO

[Eli77] I. Elishakoff. Random vibrations of orthotropic plates clamped or simply supported all round. Advances in Applied Probability, 9(2):218–219, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426361.

Ellis:1986:LTS

[Ell86] Steven P. Ellis. A limit theorem for spatial point processes. Advances in Applied Probability, 18(3):646–659, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427181.

Ellis:1987:SOA

[Ell87] Steven P. Ellis. Second-order approximation to the characteristic function of certain point-process integrals. Advances in Applied Probability, 19(3):546–559, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427407.

Elveback:1971:SSD

[Elv71] Lila R. Elveback. Simulation of stochastic discrete-time epidemic models for two agents. Advances in Applied Probability, 3 (2):226–228, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426168.

Elmasry:2011:ASR

[EM11] Amr Elmasry and Hosam Mahmoud. Analysis of swaps in radix selection. Advances in Applied Probability, 43(2):524–544, June 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/23024518.

Emery:1973:EPS

[Eme73] D. J. Emery. Exit problem for a spectrally positive process. Advances in Applied Probability, 5(3):498–520, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425831.

Engelke:2014:SIM

[EMOS14] Sebastian Engelke, Alexander Malinowski, Marco Oesting, and Martin Schlather. Statistical inference for max-stable processes by conditioning on extreme events. Advances in Applied Probability, 46(2):478–495, June 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1401369703.

Ethier:1980:DAM

[EN80] S. N. Ethier and Thomas Nagylaki. Diffusion approximations of Markov chains with two time scales and applications to population genetics. Advances in Applied Probability, 12(1):14– 49, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426492.

Ethier:1988:DAM

[EN88] S. N. Ethier and Thomas Nagylaki. Diffusion approximations of Markov chains with two time scales and applications to population genetics, II. Advances in Applied Probability, 20(3):525–545, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427034.

Englund:1984:ORT

[Eng84] Gunnar Englund. Obtaining remainder term estimates by an inversion technique. Advances in Applied Probability, 16(1):88– 110, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427226.

El-Neweihi:1989:NCC

[ENM89] Emad El-Neweihi and Fan C. Meng. The number of critical connection vectors of L-superadditive structure functions. Advances in Applied Probability, 21(4):930–932, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427775.

El-Neweihi:1978:SMA

[ENPS78] E. El-Neweihi, F. Proschan, and J. Sethuraman. A simple model with applications in structural reliability, extinction of species, inventory depletion and urn sampling. Advances in Applied Probability, 10(1):232–254, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426727.

Edelsbrunner:2017:ESP

[ENR17] Herbert Edelsbrunner, Anton Nikitenko, and Matthias Reitzner. Expected sizes of Poisson–Delaunay mosaics and their discrete Morse functions. Advances in Applied Probability, 49(3):745–767, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/expected-sizes-of-poissondelaunay-
mosaics-and-their-discrete-morse-functions/0C3724270ED11ED3F40FC10977F33259.

El-Neweihi:1993:OAU

[ENS93] Emad El-Neweihi and Jayaram Sethuraman. Optimal allocation under partial ordering of lifetimes of components. Advances in Applied Probability, 25(4):914–925, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427798.

Eriksson:2015:AOV

[EP15] B. Eriksson and M. R. Pistorius. American option valuation under continuous-time Markov chains. Advances in Applied Probability, 47(2):378–401, June 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1435236980.

Eplett:1986:CTA

[Epl86] W. J. R. Eplett. Continuous-time allocation indices and their discrete-time approximation. Advances in Applied Probability, 18(3):724–746, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427185.

Elker:1979:GCT

[EPS79] J. Elker, D. Pollard, and W. Stute. Glivenko-Cantelli theorems for classes of convex sets. Advances in Applied Probability, 11(4): 820–833, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426861.

Embrechts:1994:REG

[ES94] Paul Embrechts and Hanspeter Schmidli. Ruin estimation for a general insurance risk model. Advances in Applied Probability, 26(2):404–422, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427443.

Eun:2004:NDM

[ES04] Do Young Eun and Ness B. Shroff. Network decomposition in the many-sources regime. Advances in Applied Probability, 36 (3):893–918, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/4140414.

Eshel:1972:EPC

[Esh72] Ilan Eshel. Evolution processes with continuity of types. Advances in Applied Probability, 4(3):475–507, December 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425991.

Eshel:1973:EDP

[Esh73] Ilan Eshel. Evolution in diploid populations with continuity of gametic types. Advances in Applied Probability, 5(1):55– 65, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425964.

Ethier:1990:DFA

[Eth90a] S. N. Ethier. The distribution of the frequencies of age-ordered alleles in a diffusion model. Advances in Applied Probability, 22 (3):519–532, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427455.

Ethier:1990:IMN

[Eth90b] S. N. Ethier. The infinitely-many-neutral-alleles diffusion model with ages. Advances in Applied Probability, 22(1):1–24, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427594.

Etheridge:1993:LBT

[Eth93] Alison M. Etheridge. Limiting behaviour of two-level measurebranching. Advances in Applied Probability, 25(4):773–782, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427791.

Edelmann:2014:NIQ

[EW14] Dominic Edelmann and Cornelia Wichelhaus. Nonparametric inference for queueing networks of GEOM^{X} /G/∞ queues in discrete time. Advances in Applied Probability, 46(3):790– 811, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1409319560.

Ewald:2011:IPI

[EX11] Christian-Oliver Ewald and Yajun Xiao. Information: Price and impact on general welfare and optimal investment. An anticipative stochastic differential game model. Advances in Applied Probability, 43(1):97–120, March 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/23024527.

Egami:2014:CSF

[EY14] Masahiko Egami and Kazutoshi Yamazaki. On the continuous and smooth fit principle for optimal stopping problems in spectrally negative Lévy models. Advances in Applied Probability, 46 (1):139–167, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1396360107.

Fackrell:2009:ACM

[Fac09] Mark Fackrell. An alternative characterization for matrix exponential distributions. Advances in Applied Probability, 41(4): 1005–1022, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/27793916.

Faddy:1974:OCF

[Fad74] M. J. Faddy. Optimal control of finite dams: Continuous output procedure. Advances in Applied Probability, 6(4):689–710, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426187.

Faddy:1977:SCM

[Fad77] M. J. Faddy. Stochastic compartmental models as approximations to more general stochastic systems with the general stochastic epidemic as an example. Advances in Applied Probability, 9(3):448–461, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426108.

Faddy:1979:PSD

[Fad79] M. J. Faddy. Partial Stochastic/Deterministic models in epidemic theory. Advances in Applied Probability, 11(2):295– 297, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426826.

Faddy:1990:VBP

[Fad90] M. J. Faddy. On variation in birth processes. Advances in Applied Probability, 22(2):480–483, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427546.

Falin:1984:SCE

[Fal84a] G. I. Falin. On sufficient conditions for ergodicity of multichannel queueing systems with repeated calls. Advances in Applied Probability, 16(2):447–448, June 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427079.

Falin:1984:QIP

[Fal84b] G. I. Falin. Quasi-input process in the M/G/1/∞ queue. Advances in Applied Probability, 16(3):695–696, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427298.

Falin:1988:MBA

[Fal88] G. I. Falin. On a multiclass batch arrival retrial queue. Advances in Applied Probability, 20(2):483–487, June 1988. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427403.

Falin:1989:PQH

[Fal89] G. I. Falin. Periodic queues in heavy traffic. Advances in Applied Probability, 21(2):485–487, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427175.

Farlay:1976:OSC

[Far76] G. Farlay. Optimal stochastic control in terms of martingales. Advances in Applied Probability, 8(2):241, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425888.

Fasen:2005:ERV

[Fas05] Vicky Fasen. Extremes of regularly varying Lévy-driven mixed moving average processes. Advances in Applied Probability, 37 (4):993–1014, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/30037370.

Fasen:2014:LTH

[Fas14] Vicky Fasen. Limit theory for high frequency sampled MCARMA models. Advances in Applied Probability, 46(3):846– 877, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1409319563.

Foley:1983:QDF

[FD83] Robert D. Foley and Ralph L. Disney. Queues with delayed feedback. Advances in Applied Probability, 15(1):162–182, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426988.

Finke:2020:LTS

[FDJ20] Axel Finke, Arnaud Doucet, and Adam M. Johansen. Limit theorems for sequential MCMC methods. Advances in Applied Probability, 52(2):377–403, June 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/limit-theorems-for-sequential-mcmc-
methods/39570E1116306556C8200431D516661D.

Fakinos:1998:OSB

[FE98] D. Fakinos and A. Economou. Overall station balance and decomposability for non-Markovian queueing networks. Advances in Applied Probability, 30(3):870–887, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428155.

Finkelstein:2006:ABG

[FE06] Maxim Finkelstein and Veronica Esaulova. Asymptotic behavior of a general class of mixture failure rates. Advances in Applied Probability, 38(1):244–262, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443437.

Federgruen:1978:PSG

[Fed78] A. Federgruen. On N-person stochastic games with denumerable state space. Advances in Applied Probability, 10(2):452– 471, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426945.

Feichtinger:1976:GSA

[Fei76a] Gustav Feichtinger. On the generalization of stable age distributions to Gani-type person-flow models. Advances in Applied Probability, 8(3):433–445, September 1976. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426137.

Feigin:1976:MLE

[Fei76b] Paul David Feigin. Maximum likelihood estimation for continuous-time stochastic processes. Advances in Applied Probability, 8(4):712–736, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425931.

Feigin:1978:CPP

[Fei78] Paul D. Feigin. A characterization of point processes with the order statistic property. Advances in Applied Probability, 10 (2):301–304, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426904.

Feron:1977:FRC

[Fér77] R. Féron. Fuzzy random correspondences. Advances in Applied Probability, 9(3):426–427, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426089.

Ferrandiz:1993:BGQ

[Fer93] Josep M. Ferrandiz. The BMAP/GI/1 queue with server set-up times and server vacations. Advances in Applied Probability, 25 (1):235–254, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427504.

Fernandez:2007:PRS

[FFG07] Pedro J. Fernández, Pablo A. Ferrari, and Sebastian P. Grynberg. Perfectly random sampling of truncated multinormal distributions. Advances in Applied Probability, 39(4):973–990, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443554.

Fayolle:1986:FEA

[FFH86] Guy Fayolle, Philippe Flajolet, and Micha Hofri. On a functional equation arising in the analysis of a protocol for a multi-access broadcast channel. Advances in Applied Probability, 18(2):441– 472, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427308.

Fomichov:2022:PTD

[FFI22] Vladimir Fomichov, Sandro Franceschi, and Jevgenijs Ivanovs. Probability of total domination for transient reflecting processes in a quadrant. Advances in Applied Probability, 54 (4):1094–1138, December 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/probability-of-total-domination-
for-transient-reflecting-processes-in-a-quadrant/2460BE9BCEC04D8BD08650A29B82EC70.

Feldman:1975:STL

[FFT75] Marcus W. Feldman, Ian Franklin, and Glenys J. Thomson. Selection with three loci. Advances in Applied Probability, 7(1): 7–8, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425843.

Fellous:1981:SPP

[FG81] A. Fellous and J. Granara. Statistique des processus de Poisson stationnaires de sous-variétés linéaires affines. (French) [Statistical of stationary Poisson process of linear affine submanifolds]. Advances in Applied Probability, 13(1):84–92, March 1981. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426468.

Flajolet:2000:FTB

[FG00] Philippe Flajolet and Fabrice Guillemin. The formal theory of birth-and-death processes, lattice path combinatorics and continued fractions. Advances in Applied Probability, 32(3):750– 778, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428412.

Fontbona:2012:QEL

[FGM12] Joaquin Fontbona, Hélène Guérin, and Florent Malrieu. Quantitative estimates for the long-time behavior of an ergodic variant of the telegraph process. Advances in Applied Probability, 44(4): 977–994, December 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/41805987.

Fricker:2017:ASR

[FGRT17] Christine Fricker, Fabrice Guillemin, Philippe Robert, and Guilherme Thompson. Allocation schemes of resources with downgrading. Advances in Applied Probability, 49(2):629–651, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/allocation-
schemes-of-resources-with-downgrading/2D3A98AF060D0A685ED34E31655F7965.

Feng:2017:LTA

[FGZ17] Shui Feng, Fuqing Gao, and Youzhou Zhou. Limit theorems associated with the Pitman–Yor process. Advances in Applied Probability, 49(2):581–602, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/limit-theorems-associated-with-the-
pitmanyor-process/57ABC2FA72085CBC35749E15B02BF5C0.

Fristedt:1974:SPR

[FH74] Bert Fristedt and David Heath. Searching for a particle on the real line. Advances in Applied Probability, 6(1):79–102, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426208.

Fagerholm:2002:SCR

[FH02] Henrik Fagerholm and Göran Högnäs. Stability classification of a Ricker model with two random parameters. Advances in Applied Probability, 34(1):112–127, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428374.

Finch:2004:RCH

[FH04] Steven Finch and Irene Hueter. Random convex hulls: A variance revisited. Advances in Applied Probability, 36(4):981–986, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
4140385.

Fan:2020:LTM

[FHJK20] Jie Yen Fan, Kais Hamza, Peter Jagers, and Fima C. Klebaner. Limit theorems for multi-type general branching processes with population dependence. Advances in Applied Probability, 52(4):1127–1163, December 2020. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/limit-theorems-for-multitype-general-
branching-processes-with-population-dependence/FB30A4DFC6BC4AAE6B0FE02163A732AF.

Fienberg:1975:SMS

[Fie75] Stephen E. Fienberg. Stochastic models for single neuron firing trains: A survey. Advances in Applied Probability, 7(2):259– 260, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426073.

Fielden:1976:ICD

[Fie76a] Christa M. Fielden. Inbreeding coefficients derived from the plenary pedigree. Advances in Applied Probability, 8(1):7–10, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426011.

Fielden:1976:ITL

[Fie76b] Christa M. Fielden. Inbreeding: The three-locus case. Advances in Applied Probability, 8(4):628–630, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425920.

Fielden:1980:IML

[Fie80] Christa Fielden. Inbreeding with multiple loci. Advances in Applied Probability, 12(1):5, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426481.

Fildes:1972:ADB

[Fil72] Robert Fildes. An age dependent branching process with variable lifetime distribution. Advances in Applied Probability, 4(3): 453–474, December 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425990.

Fildes:1974:ADB

[Fil74] Robert Fildes. An age dependent branching process with variable lifetime distribution: The generation size. Advances in Applied Probability, 6(2):291–308, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426294.

Finch:1969:LLS

[Fin69] P. D. Finch. Linear least squares prediction in non-stochastic time series. Advances in Applied Probability, 1(1):111–122, Spring 1969. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426411.

Findeisen:1980:APC

[Fin80] Peter Findeisen. Asymptotic properties of a certain class of Bush-Mosteller learning models. Advances in Applied Probability, 12(4):922–941, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426749.

Finkelstein:1992:SNT

[Fin92] M. S. Finkelstein. Some notes on two types of minimal repair. Advances in Applied Probability, 24(1):226–228, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427740.

Fu:2009:APR

[FJ09] James C. Fu and Brad C. Johnson. Approximate probabilities for runs and patterns in I.I.D. and Markov-dependent multistate trials. Advances in Applied Probability, 41(1):292–308, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/27793874.

Friesen:2020:EAP

[FJKR20] Martin Friesen, Peng Jin, Jonas Kremer, and Barbara Rüdiger. Ergodicity of affine processes on the cone of symmetric positive semidefinite matrices. Advances in Applied Probability, 52(3):825–854, September 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/ergodicity-of-affine-processes-on-
the-cone-of-symmetric-positive-semidefinite-matrices/
5B2F94EF68B3B1CDB329D24F28E1A5CF.

Foley:1989:PCT

[FK89] Robert D. Foley and Peter C. Kiessler. Positive correlations in a three-node Jackson queueing network. Advances in Applied Probability, 21(1):241–242, March 1989. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427212.

Fayolle:1982:SCT

[FKM82] G. Fayolle, P. J. B. King, and I. Mitrani. The solution of certain two-dimensional Markov models. Advances in Applied Probability, 14(2):295–308, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426522.

Fodor:2014:RDP

[FKV14] F. Fodor, P. Kevei, and V. Vígh. On random disc polygons in smooth convex discs. Advances in Applied Probability, 46(4): 899–918, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1418396236.

Fuh:2001:AEM

[FL01] Cheng-Der Fuh and Tze Leung Lai. Asymptotic expansions in multidimensional Markov renewal theory and first passage times for Markov random walks. Advances in Applied Probability, 33 (3):652–673, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428480.

Figueroa-Lopez:2009:NET

[FL09] José Figueroa-López. Nonparametric estimation of timechanged Lévy models under high-frequency data. Advances in Applied Probability, 41(4):1161–1188, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/27793924.

Ferrari:2022:OCH

[FLR22] Giorgio Ferrari, Hanwu Li, and Frank Riedel. Optimal consumption with Hindy–Huang–Kreps preferences under nonlinear expectations. Advances in Applied Probability, 54 (4):1222–1251, December 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/optimal-consumption-with-hindyhuangkreps-
preferences-under-nonlinear-expectations/8B9389A8B359D034256CE17BB97FE5A5.

Follmer:1980:RSL

[FLS^{+}80] Hans Föllmer, Thomas M. Liggett, Klaus Schürger, Frank Spitzer, Maury Bramson, David Griffeath, William A. Little, A. Robertson, and P. Tautu. Random systems with locally interacting objects. Advances in Applied Probability, 12(3):555–559, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426416.

Flynn:2018:FSA

[Fly18] Thomas Flynn. Forward sensitivity analysis for contracting stochastic systems. Advances in Applied Probability, 50(1):102–130, March 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/forward-sensitivity-analysis-for-
contracting-stochastic-systems/2E8BA0A35E49D6067D55A67AED1EB916.

Franz:2012:PDM

[FLZ12] Uwe Franz, Volkmar Liebscher, and Stefan Zeiser. Piecewisedeterministic Markov processes as limits of Markov jump processes. Advances in Applied Probability, 44(3):729–748, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
41714073.

Faris:1988:PAL

[FM88] William G. Faris and Robert S. Maier. Probabilistic analysis of a learning matrix. Advances in Applied Probability, 20(4): 695–705, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427355.

Ferrandiz:1995:SIR

[FM95] J. Ferrándiz and F. Montes. On statistical inference for random closed sets in compact subsets of R^{d}. Advances in Applied Probability, 27(2):344–349, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427829.

Freund:2009:NAT

[FM09] F. Freund and M. Möhle. On the number of allelic types for samples taken from exchangeable coalescents with mutation. Advances in Applied Probability, 41(4):1082–1101, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/27793920.

Foley:2018:YLC

[FM18] R. D. Foley and D. R. McDonald. Yaglom limits can depend on the starting state. Advances in Applied Probability, 50(1):1–34, March 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/yaglom-limits-can-depend-on-the-
starting-state/DAE6D1D2244ACAC4E3B4BE894600B4A6.

Ferrari:1992:ENT

[FMP92] Pablo A. Ferrari, Servet Martínez, and Pierre Picco. Existence of non-trivial quasi-stationary distributions in the birth– death chain. Advances in Applied Probability, 24(4):795–813, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427713.

Foucart:2022:LTC

[FMY22] Clément Foucart, Chunhua Ma, and Linglong Yuan. Limit theorems for continuous-state branching processes with immigration. Advances in Applied Probability, 54(2):599–624, June 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/limit-theorems-
for-continuousstate-branching-processes-with-immigration/
D63EDB70B7DCD3DBE6C79B267D95541F.

Funnemark:1985:BAF

[FN85] Espen Funnemark and Bent Natvig. Bounds for the availabilities in a fixed time interval for multistate monotone systems. Advances in Applied Probability, 17(3):638–665, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427124.

Fill:2013:DCN

[FN13] James Allen Fill and Takehiko Nakama. Distributional convergence for the number of symbol comparisons used by Quick-Select. Advances in Applied Probability, 45(2):425–450, June 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1370870125.

Fan-Orzechowski:2006:ORT

[FOF06] X. Fan-Orzechowski and E. A. Feinberg. Optimality of randomized trunk reservation for a problem with a single constraint. Advances in Applied Probability, 38(1):199–220, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/20443435.

Foucart:2011:DEC

[Fou11] Clément Foucart. Distinguished exchangeable coalescents and generalized Fleming–Viot processes with immigration. Advances in Applied Probability, 43(2):348–374, June 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/23024510.

Fournier:2018:IIC

[Fou18] Julie Fournier. Identification and isotropy characterization of deformed random fields through excursion sets. Advances in Applied Probability, 50(3):706–725, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/identification-and-isotropy-
characterization-of-deformed-random-fields-through-excursion-
sets/7F6D67190CCF23A9031D24D5EE957E84.

Foxall:2014:CMM

[Fox14] Eric Foxall. Convergence and monotonicity for a model of spontaneous infection and transmission. Advances in Applied Probability, 46(2):560–584, June 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1401369707.

Foxall:2016:DCC

[Fox16] Eric Foxall. Duality and complete convergence for multi-type additive growth models. Advances in Applied Probability, 48 (1):32–51, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1457466154.

Fort:1996:CSA

[FP96] Jean-Claude Fort and Gilles Pagès. Convergence of stochastic algorithms: From the Kushner–Clark theorem to the Lyapounov functional method. Advances in Applied Probability, 28(4):1072– 1094, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/ stable/1428165. See correction [Ano98e].

Fuh:2012:SNC

[FP12] Cheng-Der Fuh and Tian-Xiao Pang. A self-normalized central limit theorem for Markov random walks. Advances in Applied Probability, 44(2):452–478, June 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/41714059.

Ferreira:2019:TPZ

[FPP19] M. Ferreira, D. Pinheiro, and S. Pinheiro. Two-player zero-sum stochastic differential games with random horizon. Advances in Applied Probability, 51(4):1209–1235, December 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/twoplayer-zerosum-stochastic-
differential-games-with-random-horizon/A265A06EC7019EE37D93BF57B1879693.

Foss:2022:LTS

[FPS22] Sergey Foss, Timofei Prasolov, and Seva Shneer. Limit theorems and structural properties of the cat-and-mouse Markov chain and its generalisations. Advances in Applied Probability, 54(1):141–166, March 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/limit-theorems-and-structural-properties-
of-the-catandmouse-markov-chain-and-its-generalisations/
550A89B1C3DD5D0B7FE9C52DEDF5DFCE.

Forde:2013:HTO

[FPZ13] Martin Forde, Andrey Pogudin, and Hongzhong Zhang. Hitting times, occupation times, trivariate laws and the forward Kolmogorov equation for a one-dimensional diffusion with memory. Advances in Applied Probability, 45(3):860–875, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1377868542.

Fahady:1971:HTA

[FQVJ71] K. S. Fahady, M. P. Quine, and D. Vere-Jones. Heavy traffic approximations for the Galton–Watson process. Advances in Applied Probability, 3(2):282–300, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426172.

Foschini:1980:CTP

[FR80] G. J. Foschini and M. I. Reiman. Controlling tandem processors. Advances in Applied Probability, 12(2):313–314, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426592.

Feigin:1997:LPE

[FR97] Paul D. Feigin and Sidney I. Resnick. Linear programming estimators and bootstrapping for heavy tailed phenomena. Advances in Applied Probability, 29(3):759–805, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428085.

Faller:2011:ASO

[FR11] Andreas Faller and Ludger Rüschendorf. On approximative solutions of optimal stopping problems. Advances in Applied Probability, 43(4):1086–1108, December 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/23208040.

Feuillet:2012:TBE

[FR12] Mathieu Feuillet and Philippe Robert. On the transient behavior of Ehrenfest and Engset processes. Advances in Applied Probability, 44(2):562–582, June 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/41714063.

Feuillet:2014:SAT

[FR14] Mathieu Feuillet and Philippe Robert. A scaling analysis of a transient stochastic network. Advances in Applied Probability, 46(2):516–535, June 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1401369705.

Franke:1985:APM

[Fra85] Jürgen Franke. ARMA processes have maximal entropy among time series with prescribed autocovariances and impulse responses. Advances in Applied Probability, 17(4):810–840, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1427089. See correction [Ano86e].

Franchi:1989:PSD

[Fra89] J. Franchi. Produits semi-directs de diffusions reelles et lois asymptotiques. (French) [Semi-direct products of real diffusions and asymptotic laws]. Advances in Applied Probability, 21(4): 756–769, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427765.

Frenk:1982:BRS

[Fre82] J. B. G. Frenk. The behavior of the renewal sequence in case the tail of the waiting-time distribution is regularly varying with index −1. Advances in Applied Probability, 14(4):870–884, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427028.

Fristedt:1977:LBS

[Fri77] Bert Fristedt. Local behavior of stable subordinators. Advances in Applied Probability, 9(2):194–195, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426342.

Friesen:1982:ONN

[Fri82] W. O. Friesen. Organization of neural networks in the invertebrates. Advances in Applied Probability, 14(2):196–197, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426507.

Fricker:1986:EFG

[Fri86] C. Fricker. Étude d’une file GI/G/1 à service autonome (avec vacances du serveur). (French) [Study of a GI/G/1 queue in standalone service (with server outages)]. Advances in Applied Probability, 18(1):283–286, March 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427249. See note [Fri87].

Fricker:1987:NMF

[Fri87] Christine Fricker. Note sur un modèle de file GI/G/1 à service autonome (avec vacances du serveur). (French) [Note on a model of GI/G/1 queue with standalone service (with server outages)].

Advances in Applied Probability, 19(1):289–291, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427385. See

[Fri86].

Falgas-Ravry:2020:SCC

[FRLM20] Victor Falgas-Ravry, Joel Larsson, and Klas Markström. Speed and concentration of the covering time for structured coupon collectors. Advances in Applied Probability, 52(2):433–462, June 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/speed-and-concentration-of-the-covering-
time-for-structured-coupon-collectors/10D13221F2E173C9B2823AAA402BC684.

Frostig:2004:UBE

[Fro04] Esther Frostig. Upper bounds on the expected time to ruin and on the expected recovery time. Advances in Applied Probability, 36(2):377–397, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428458.

Frostig:2013:MAR

[Fro13] Esther Frostig. A Markov additive risk process with a dividend barrier. Advances in Applied Probability, 45(2):451–489, June 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1370870126.

Falgas-Ravry:2012:SNN

[FRW12] Victor Falgas-Ravry and Mark Walters. Sharpness in the knearest-neighbours random geometric graph model. Advances in Applied Probability, 44(3):617–634, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/41714068.

Foutz:1979:SIM

[FS79] Robert V. Foutz and R. C. Srivastava. Statistical inference for Markov processes when the model is incorrect. Advances in Applied Probability, 11(4):737–749, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426857.

Freed:1982:PPW

[FS82] D. S. Freed and L. A. Shepp. A Poisson process whose rate is a hidden Markov process. Advances in Applied Probability, 14 (1):21–36, March 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426731.

Fakinos:1989:PFD

[FS89a] D. Fakinos and K. Sirakoulis. Product-form distributions for an M/G/k group-arrival group-departure loss system. Advances in Applied Probability, 21(3):721–724, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427649.

Federgruen:1989:OTR

[FS89b] Awi Federgruen and Kut C. So. Optimal time to repair a broken server. Advances in Applied Probability, 21(2):376–397, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427165.

Federgruen:1991:OTP

[FS91] Awi Federgruen and Kut C. So. Optimality of threshold policies in single-server queueing systems with server vacations. Advances in Applied Probability, 23(2):388–405, June 1991. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427755.

Ferretti:1993:EDP

[FS93] V. Ferretti and D. Sankoff. The empirical discovery of phylogenetic invariants. Advances in Applied Probability, 25(2):290– 302, June 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427653.

Frei:2008:EUI

[FS08] Christoph Frei and Martin Schweizer. Exponential utility indifference valuation in two Brownian settings with stochastic correlation. Advances in Applied Probability, 40(2):401–423, June 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/20443589.

Fasen:2009:FCP

[FS09a] Vicky Fasen and Gennady Samorodnitsky. A fluid cluster Poisson input process can look like a fractional Brownian motion even in the slow growth aggregation regime. Advances in Applied Probability, 41(2):393–427, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/27793884.

Franke:2009:ERW

[FS09b] Brice Franke and Tatsuhiko Saigo. The extremes of random walks in random sceneries. Advances in Applied Probability, 41 (2):452–468, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/27793886.

Ferrari:2016:IIU

[FS16a] Giorgio Ferrari and Paavo Salminen. Irreversible investment under Lévy uncertainty: an equation for the optimal boundary. Advances in Applied Probability, 48(1):298–314, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1457466167.

Froyland:2016:CIA

[FS16b] Gary Froyland and Robyn M. Stuart. Cheeger inequalities for absorbing Markov chains. Advances in Applied Probability, 48(3):631–647, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1474296307.

Federgruen:1978:OEA

[FT78] A. Federgruen and H. C. Tijms. The optimality equation in average cost denumerable state semi-Markov decision problems, recurrency conditions and algorithms. Advances in Applied Probability, 10(2):308, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426908.

Falk:2012:ACD

[FT12] Michael Falk and Diana Tichy. Asymptotic conditional distribution of exceedance counts. Advances in Applied Probability, 44 (1):270–291, March 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/23207864.

Finkelshtein:2018:KBS

[FT18] Dmitri Finkelshtein and Pasha Tkachov. Kesten’s bound for subexponential densities on the real line and its multidimensional analogues. Advances in Applied Probability, 50(2):373–395, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/kestens-bound-for-subexponential-
densities-on-the-real-line-and-its-multidimensional-
analogues/E8B163F408472791735155BAD0FFBB82.

Fuh:1997:CDA

[Fuh97] C. D. Fuh. Corrected diffusion approximations for ruin probabilities in a Markov random walk. Advances in Applied Probability, 29(3):695–712, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428082.

Fuh:2007:AEM

[Fuh07] Cheng-Der Fuh. Asymptotic expansions on moments of the first ladder height in Markov random walks with small drift. Advances in Applied Probability, 39(3):826–852, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/20443542.

Fujimagari:1979:ETD

[Fuj79] Tetsuo Fujimagari. On the extinction time distribution of a branching process in varying environments. Advances in Applied Probability, 11(2):291–292, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426822.

Fujimagari:1980:ETD

[Fuj80] Tetsuo Fujimagari. On the extinction time distribution of a branching process in varying environments. Advances in Applied Probability, 12(2):350–366, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426601.

Forte:1995:SIP

[FV95] B. Forte and E. R. Vrscay. Solving the inverse problem for measures using iterated function systems: A new approach. Advances in Applied Probability, 27(3):800–820, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428134.

Fleischmann:1999:RSG

[FV99] Klaus Fleischmann and Vladimir A. Vatutin. Reduced subcritical Galton–Watson processes in a random environment. Advances in Applied Probability, 31(1):88–111, March 1999. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427993.

Fey:2008:LDE

[FVK08] Anne Fey, Remco Van Der Hofstad, and Marten J. Klok. Large deviations for eigenvalues of sample covariance matrices, with applications to mobile communication systems. Advances in Applied Probability, 40(4):1048–1071, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443624.

Finkelstein:1981:OSR

[FW81] Mark Finkelstein and Robert Whitley. Optimal strategies for repeated games. Advances in Applied Probability, 13(2):415– 428, June 1981. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426692.

Feng:2000:RMT

[FX00] Y. Feng and B. Xiao. Revenue management with two market segments and reserved capacity for priority customers. Advances in Applied Probability, 32(3):800–823, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428414.

Feng:2013:ABM

[FX13] Shui Feng and Jie Xiong. Asymptotic behavior of the Moran particle system. Advances in Applied Probability, 45(2):379–397, June 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1370870123.

Ferrari:2018:OEP

[FY18] Giorgio Ferrari and Shuzhen Yang. On an optimal extraction problem with regime switching. Advances in Applied Probability, 50(3):671–705, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-an-optimal-extraction-problem-
with-regime-switching/24AB3D62423FB4B6095C4AA3ECE1546B.

Federgruen:1985:COP

[FZ85] A. Federgruen and P. Zipkin. Computing optimal (s,s) policies in inventory models with continuous demands. Advances in Applied Probability, 17(2):424–442, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427149.

Foss:1996:VAP

[FZ96] S. G. Foss and S. A. Zuyev. On a Voronoi aggregative process related to a bivariate Poisson process. Advances in Applied Probability, 28(4):965–981, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428159.

Foss:2013:SSR

[FZ13] Sergey Foss and Stan Zachary. Stochastic sequences with a regenerative structure that may depend both on the future and on the past. Advances in Applied Probability, 45(4):1083–1110, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1386857859.

Gual-Arnau:2000:SSC

[GACO00] Ximo Gual-Arnau and Luis M. Cruz-Orive. Systematic sampling on the circle and on the sphere. Advances in Applied Probability, 32(3):628–647, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428404.

Gual-Arnau:2002:VPP

[GACO02] Ximo Gual-Arnau and Luis M. Cruz-Orive. Variance prediction for pseudosystematic sampling on the sphere. Advances in Applied Probability, 34(3):469–483, September 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428241.

Gjessing:2003:FMB

[GAH03] Håkon K. Gjessing, Odd O. Aalen, and Nils Lid Hjort. Frailty models based on Lévy processes. Advances in Applied Probability, 35(2):532–550, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428434.

Galambos:1979:EVT

[Gal79a] Janos Galambos. Extreme value theory in applied probability. Advances in Applied Probability, 11(2):289–290, June 1979. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426820.

Gallisch:1979:MOP

[Gal79b] Eckhardt Gallisch. On monotone optimal policies in a queueing model of M/G/1 type with controllable service time distribution. Advances in Applied Probability, 11(4):870–887, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426864.

Gallardo:1984:CAM

[Gal84] Leonard Gallardo. Comportement asymptotique des marches aléatoires associées aux polynômes de Gegenbauer et applications. (French) [Asymptotic behavior of random walks associated with Gegenbauer polynomials and applications]. Advances in Applied Probability, 16(2):293–323, June 1984. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427071.

Gallo:2011:CUV

[Gal11] Sandro Gallo. Chains with unbounded variable length memory: Perfect simulation and a visible regeneration scheme. Advances in Applied Probability, 43(3):735–759, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/23024573.

Galerne:2016:RFB

[Gal16] Bruno Galerne. Random fields of bounded variation and computation of their variation intensity. Advances in Applied Probability, 48(4):947–971, December 2016. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1482548424.

Gani:1969:RAS

[Gan69] J. Gani. Recent advances in storage and flooding theory. Advances in Applied Probability, 1(1):90–110, Spring 1969. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426410.

Gani:1971:RWE

[Gan71] J. Gani. Recent work in epidemiology at Sheffield. Advances in Applied Probability, 3(2):204–206, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426157.

Gani:1975:SSM

[Gan75] J. Gani. Some stochastic models in linguistic analysis. Advances in Applied Probability, 7(2):232–234, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426048.

Gani:1978:SVA

[Gan78] J. Gani. Some very applied problems in applied probability. Advances in Applied Probability, 10(2):263–265, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426873.

Gani:1979:SIC

[Gan79] J. Gani. Some interesting current directions in epidemic theory. Advances in Applied Probability, 11(2):262, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426793.

Gani:2008:EJA

[Gan08] J. Gani. Editorship of journal of applied probability and advances in applied probability. Advances in Applied Probability, 40(1):??, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443566.

Ganesan:2015:ISR

[Gan15] Ghurumuruhan Ganesan. Infection spread in random geometric graphs. Advances in Applied Probability, 47(1):164–181, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1427814586.

Gart:1971:MMI

[Gar71] John J. Gart. Mathematical models in the interpretation of the interaction between infective agents. Advances in Applied Probability, 3(2):202–203, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426156.

Gardiner:1980:BSF

[Gar80] D. Gardiner. The behaviour of strategy-frequencies in Parker’s model. Advances in Applied Probability, 12(1):5–7, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426482.

Garcia:1995:BDP

[Gar95] Nancy Lopes Garcia. Birth and death processes as projections of higher-dimensional Poisson processes. Advances in Applied Probability, 27(4):911–930, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427928.

Garmo:1999:APC

[Gar99] Hans Garmo. Asymptotic properties of the connectivity number of random railways. Advances in Applied Probability, 31(3): 720–741, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428393.

Gaasemyr:2012:BAM

[Gås12] J. Gåsemyr. Bounds for the availabilities of multistate monotone systems based on decomposition into stochastically independent modules. Advances in Applied Probability, 44(1):292–308, March 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/23207865.

Gates:1985:ATI

[Gat85a] D. J. Gates. Asymptotics of two integrals from optimization theory and geometric probability. Advances in Applied Probability, 17(4):908–910, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427094.

Gates:1985:OCE

[Gat85b] D. J. Gates. On the optimal composition of electricity grids with unreliable units: Solvable models. Advances in Applied Probability, 17(2):367–385, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427146.

Gates:1994:SDP

[Gat94] John Gates. Shape distributions for planar triangles by dual construction. Advances in Applied Probability, 26(2):324–333, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427438.

Gates:1997:SSD

[Gat97] D. J. Gates. Stationary states of a directed planar growth process. Advances in Applied Probability, 29(4):857–877, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427844.

Gatzouras:2000:LCA

[Gat00] Dimitris Gatzouras. On the lattice case of an almost-sure renewal theorem for branching random walks. Advances in Applied Probability, 32(3):720–737, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428410.

Gauthier:2002:EHL

[Gau02] Laurent Gauthier. Excursions height- and length-related stopping times, and application to finance. Advances in Applied Probability, 34(4):846–868, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428325.

Georgii:2003:SMB

[GB03] Hans-Otto Georgii and Ellen Baake. Supercritical multitype branching processes: The ancestral types of typical individuals. Advances in Applied Probability, 35(4):1090–1110, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428501.

Gupta:2017:AFR

[GB17] Ramesh C. Gupta and David M. Bradley. Asymptotic failure rates for a general class of frailty models. Advances in Applied Probability, 49(4):1230–1259, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/asymptotic-failure-rates-
for-a-general-class-of-frailty-models/B0E9F84AA309BAC5E927D3CCDE6E5FA1.

Gomez-Corral:2002:TNB

[GC02] A. Gómez-Corral. On a tandem G-network with blocking. Advances in Applied Probability, 34(3):626–661, September 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428250.

Gandolfi:2016:SES

[GC16] Alberto Gandolfi and Lorenzo Cecconi. SIR epidemics on a scale-free spatial nested modular network. Advances in Applied Probability, 48(1):137–162, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1457466159.

Gibbs:2020:APR

[GC20] Isaac Gibbs and Linan Chen. Asymptotic properties of random Voronoi cells with arbitrary underlying density. Advances in Applied Probability, 52(2):655–680, June 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/asymptotic-properties-of-random-
voronoi-cells-with-arbitrary-underlying-density/3A07C4AD0C27C09891D9D7D9C388F0A3.

Geciauskas:1977:DFD

[Gec77] E. Geciauskas. The distribution function of the distance between two points in a convex domain. Advances in Applied Probability, 9(3):427–428, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426090.

Geman:1980:CBM

[Gem80] Donald Geman. Confluent Brownian motions. Advances in Applied Probability, 12(2):306, June 1980. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1426583.

George:1976:SRP

[Geo76] Larry George. Simulation of a replacement process and complimentary antithetic variates. Advances in Applied Probability, 8 (2):237–238, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425884.

Georgii:1984:FPR

[Geo84] Hans-Otto Georgii. On the ferromagnetic and the percolative region of random spin systems. Advances in Applied Probability, 16(4):732–765, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427339.

Gersch:1977:AME

[Ger77] Will Gersch. ARMA model eigenvalues and an amount of information measure features for automatic classification of multivariate EEG’s. Advances in Applied Probability, 9(2):219, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426362.

Gersch:1980:DSN

[Ger80] Will Gersch. Discrimination in stationary and non-stationary time series. Advances in Applied Probability, 12(2):314–315, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426594.

Gertsbakh:1984:AMR

[Ger84a] I. B. Gertsbakh. Asymptotic methods in reliability theory: A review. Advances in Applied Probability, 16(1):147–175, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427229.

Gertsbakh:1984:OGP

[Ger84b] I. B. Gertsbakh. Optimal group preventive maintenance of a system with observable state parameter. Advances in Applied Probability, 16(4):923–925, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427348.

Gerontidis:1994:SMR

[Ger94] Ioannis I. Gerontidis. Semi-Markov replacement chains. Advances in Applied Probability, 26(3):728–755, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427818.

Gerontidis:1995:MPR

[Ger95] Ioannis I. Gerontidis. Markov population replacement processes. Advances in Applied Probability, 27(3):711–740, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428131.

Getoor:1974:SRP

[Get74] R. K. Getoor. Some remarks on a paper of Kingman. Advances in Applied Probability, 6(4):757–767, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426191. See [Kin73].

Glazebrook:1987:SAS

[GF87] K. D. Glazebrook and N. A. Fay. On the scheduling of alternative stochastic jobs on a single machine. Advances in Applied Probability, 19(4):955–973, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427110.

Gupta:1990:BRE

[GG90] Pushpa L. Gupta and R. D. Gupta. A bivariate random environmental stress model. Advances in Applied Probability, 22 (2):501–503, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427553.

Goldie:1996:PTT

[GG96] Charles M. Goldie and Rudolf Grübel. Perpetuities with thin tails. Advances in Applied Probability, 28(2):463–480, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428067.

Galerne:2012:TDL

[GG12a] B. Galerne and Y. Gousseau. The transparent dead leaves model. Advances in Applied Probability, 44(1):1–20, March 2012.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/23207853.

Ghosh:2012:SLR

[GG12b] Souvik Ghosh and Soumyadip Ghosh. A strong law for the rate of growth of long latency periods in a cloud computing service. Advances in Applied Probability, 44(4):995–1017, December 2012. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
41805988.

Gast:2017:CAT

[GG17] Nicolas Gast and Bruno Gaujal. Computing absorbing times via fluid approximations. Advances in Applied Probability, 49(3):768–790, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/computing-absorbing-times-via-fluid-
approximations/230896F1AE9B9529106E80F2D89DC4E9.

Gubner:1989:BFM

[GGV89] John A. Gubner, B. Gopinath, and S. R. S. Varadhan. Bounding functions of Markov processes and the shortest queue problem. Advances in Applied Probability, 21(4):842–860, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427770.

Guthrie:1975:RIR

[GH75] Gary Lee Guthrie and Paul T. Holmes. On record and interrecord times for a sequence of random variables defined on a Markov chain. Advances in Applied Probability, 7(1):195– 214, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425860.

Goldstein:1980:MCC

[GH80] Martin I. Goldstein and Fred M. Hoppe. Minimal convergence conditions for critical age-dependent processes. Advances in Applied Probability, 12(2):299–300, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426575.

Ghate:2008:DSS

[Gha08] Archis Ghate. Decentralized search on spheres using small-world Markov chains: Expected hitting times and structural properties. Advances in Applied Probability, 40(4):966–978, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/20443620.

Guo:2015:FHO

[GHH15] Xianping Guo, Xiangxiang Huang, and Yonghui Huang. Finitehorizon optimality for continuous-time Markov decision processes with unbounded transition rates. Advances in Applied Probability, 47(4):1064–1087, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1449859800.

Glazebrook:2013:MPI

[GHK13] K. D. Glazebrook, D. J. Hodge, and C. Kirkbride. Monotone policies and indexability for bidirectional restless bandits. Advances in Applied Probability, 45(1):51–85, March 2013. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/41806000.

Guo:2007:ZSG

[GHL07] Xianping Guo and Onésimo Hernández-Lerma. Zero-sum games for continuous-time jump Markov processes in Polish spaces: Discounted payoffs. Advances in Applied Probability, 39(3):645– 668, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443534.

Goncalves:2023:DSP

[GHL23] Branda Goncalves, Thierry Huillet, and Eva Löcherbach. On decay-surge population models. Advances in Applied Probability, 55(2):444–472, June 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-decaysurge-population-models/
42969680B329F4C77347FB95240AF065.

Gail:1988:ANP

[GHT88] H. R. Gail, S. L. Hantler, and B. A. Taylor. Analysis of a non-preemptive priority multiserver queue. Advances in Applied Probability, 20(4):852–879, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427364.

Gail:1996:SAM

[GHT96] H. R. Gail, S. L. Hantler, and B. A. Taylor. Spectral analysis of M/G/1 and G/M/1 type Markov chains. Advances in Applied Probability, 28(1):114–165, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427915.

Gail:1997:NSF

[GHT97] H. R. Gail, S. L. Hantler, and B. A. Taylor. Non-skip-free M/G/1 and G/M/1 type Markov chains. Advances in Applied Probability, 29(3):733–758, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428084.

Ghurye:1987:SML

[Ghu87] S. G. Ghurye. Some multivariate lifetime distributions. Advances in Applied Probability, 19(1):138–155, March 1987. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427377.

Gupta:2010:CER

[GI10] Bhupender Gupta and Srikanth K. Iyer. Criticality of the exponential rate of decay for the largest nearest-neighbor link in random geometric graphs. Advances in Applied Probability, 42(3): 631–658, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25799062.

Gill:1998:CR

[Gil98] Richard D. Gill. Coarsening at random. Advances in Applied Probability, 30(2):278–279, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427948.

Gille:2002:SCD

[Gil02] Wilfried Gille. The set covariance of a dead leaves model. Advances in Applied Probability, 34(1):11–20, March 2002. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428369.

Gnedin:2014:ABC

[GIMM14] Alexander Gnedin, Alexander Iksanov, Alexander Marynych, and Martin Möhle. On asymptotics of the beta coalescents. Advances in Applied Probability, 46(2):496–515, June 2014. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1401369704.

Gittins:1969:ORA

[Git69] J. C. Gittins. Optimal resource allocation in chemical research. Advances in Applied Probability, 1(2):238–270, Autumn 1969. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426219. See

correction [WG91].

Gaver:1980:CRL

[GJ80] D. P. Gaver and P. A. Jacobs. On combinations of random loads. Advances in Applied Probability, 12(2):311, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426588.

Gaver:1984:FBD

[GJL84] D. P. Gaver, P. A. Jacobs, and G. Latouche. Finite birth-and-death models in randomly changing environments. Advances in Applied Probability, 16(4):715–731, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427338.

Griffin:2017:SQS

[GJRS17] A. Griffin, P. A. Jenkins, G. O. Roberts, and S. E. F. Spencer. Simulation from quasi-stationary distributions on reducible state spaces. Advances in Applied Probability, 49 (3):960–980, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/simulation-from-quasistationary-
distributions-on-reducible-state-spaces/8FD473D762D5C79B8DB6BDD8A1E05A6F.

Griffiths:2008:IST

[GJS08] Robert C. Griffiths, Paul A. Jenkins, and Yun S. Song. Importance sampling and the two-locus model with subdivided population structure. Advances in Applied Probability, 40(2):473– 500, June 2008. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443592.

Goldschmidt:2019:SFR

[GK19] Christina Goldschmidt and Eleonora Kreacić. The spread of fire on a random multigraph. Advances in Applied Probability, 51(1):1–40, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/spread-of-fire-on-a-random-multigraph/
A2FF7B15643345A016256391414B2C21.

Gapeev:2021:DOS

[GKL21] Pavel V. Gapeev, Peter M. Kort, and Maria N. Lavrutich. Discounted optimal stopping problems for maxima of geometric Brownian motions with switching payoffs. Advances in Applied Probability, 53(1):189–219, March 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/discounted-optimal-stopping-problems-
for-maxima-of-geometric-brownian-motions-with-switching-
payoffs/56B55566BFB3CEB2C5688BB966CDE85B.

Griffith:1984:SMP

[GKP84] William S. Griffith, Richard J. Kryscio, and Peter Purdue. A stochastic model of particle shattering. Advances in Applied Probability, 16(1):70–87, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427225.

Guo:2021:GIC

[GKPZ21] Xin Guo, Aiko Kurushima, Alexey Piunovskiy, and Yi Zhang. On gradual-impulse control of continuous-time Markov decision processes with exponential utility. Advances in Applied Probability, 53(2):301–334, June 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-gradualimpulse-control-of-continuoustime-
markov-decision-processes-with-exponential-utility/D92622887B1BB939B8E10DC7868E72F9.

Glazebrook:2006:SPS

[GKRH06] K. D. Glazebrook, C. Kirkbride, and D. Ruiz-Hernandez. Spinning plates and squad systems: Policies for bi-directional restless bandits. Advances in Applied Probability, 38(1):95–115, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/20443430.

Grbac:2021:LTT

[GKT21] Zorana Grbac, David Krief, and Peter Tankov. Long-time trajectorial large deviations and importance sampling for affine stochastic volatility models. Advances in Applied Probability, 53(1):220–250, March 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/longtime-trajectorial-large-deviations-
and-importance-sampling-for-affine-stochastic-volatility-
models/573E0321BDD5F413187FBB2151A396AF.

Gusakova:2022:DTD

[GKT22] Anna Gusakova, Zakhar Kabluchko, and Christoph Thäle. The β-Delaunay tessellation: Description of the model and geometry of typical cells. Advances in Applied Probability, 54 (4):1252–1290, December 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/beta-delaunay-tessellation-description-
of-the-model-and-geometry-of-typical-cells/92CB745692FD1E76BE8570230F451126.

Goodey:1996:SMI

[GKW96] Paul Goodey, Markus Kiderlen, and Wolfgang Weil. Section means, integral transforms, and Boolean models. Advances in Applied Probability, 28(2):332–333, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428041.

Gibberd:1979:PAM

[GL79] R. W. Gibberd and H. Lewis. Properties of aggregated Markov chains. Advances in Applied Probability, 11(2):277–278, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426808.

Gaver:1980:FOA

[GL80] D. P. Gaver and P. A. W. Lewis. First-order autoregressive gamma sequences and point processes. Advances in Applied Probability, 12(3):727–745, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426429.

Glynn:1995:LRG

[GL95] Peter W. Glynn and Pierre L’Ecuyer. Likelihood ratio gradient estimation for stochastic recursions. Advances in Applied Probability, 27(4):1019–1053, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427933.

Gleaton:2004:DBS

[GL04a] James U. Gleaton and James D. Lynch. On the distribution of the breaking strain of a bundle of brittle elastic fibers. Advances in Applied Probability, 36(1):98–115, March 2004. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428355.

Gravereaux:2004:PAS

[GL04b] Jean-Bernard Gravereaux and James Ledoux. Poisson approximation for some point processes in reliability. Advances in Applied Probability, 36(2):455–470, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428462.

Gathy:2009:DMS

[GL09] Maude Gathy and Claude Lefèvre. From damage models to SIR epidemics and cascading failures. Advances in Applied Probability, 41(1):247–269, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/27793872.

Glazebrook:1983:SRP

[Gla83] K. D. Glazebrook. Some reward-penalty rules for the multiarmed bandit problem which are asymptotically optimal. Advances in Applied Probability, 15(1):221–222, March 1983. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426995.

Glazebrook:1988:RPD

[Gla88] K. D. Glazebrook. On a reduction principle in dynamic programming. Advances in Applied Probability, 20(4):836–851, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427363.

Glazebrook:1991:SES

[Gla91] K. D. Glazebrook. Strategy evaluation for stochastic scheduling problems with order constraints. Advances in Applied Probability, 23(1):86–104, March 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427513.

Glasserman:1993:RDR

[Gla93a] Paul Glasserman. Regenerative derivatives of regenerative sequences. Advances in Applied Probability, 25(1):116–139, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427499.

Glasserman:1993:SMC

[Gla93b] Paul Glasserman. Stochastic monotonicity and conditional Monte Carlo for likelihood ratios. Advances in Applied Probability, 25(1):103–115, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427498.

Glazebrook:1993:FPD

[Gla93c] K. D. Glazebrook. On a family of prior distributions for a class of Bayesian search models. Advances in Applied Probability, 25(3): 714–716, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427532.

Glasbey:1996:IBI

[Gla96a] C. A. Glasbey. Inference on binary images from binary data. Advances in Applied Probability, 28(3):627–640, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428173.

Glazebrook:1996:UTP

[Gla96b] K. D. Glazebrook. On the undiscounted tax problem with precedence constraints. Advances in Applied Probability, 28(4):1123– 1144, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428167.

Geiss:2020:MSR

[GLL20] Christel Geiss, Céline Labart, and Antti Luoto. Mean square rate of convergence for random walk approximation of forward-backward SDEs. Advances in Applied Probability, 52(3):735–771, September 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/mean-square-rate-of-convergence-
for-random-walk-approximation-of-forwardbackward-sdes/
06BD3DC57BE736270171EDC10E99C5D4.

Grama:2022:LTC

[GLL22] Ion Grama, Ronan Lauvergnat, and Émile Le Page. Limit theorems for critical branching processes in a finite-state-space Markovian environment. Advances in Applied Probability, 54(1):111–140, March 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/limit-theorems-for-critical-branching-
processes-in-a-finitestatespace-markovian-environment/
CAEC7D5F248B269FF0B6C2220D06AFF6.

Gorostiza:1990:MMB

[GLM90] Luis G. Gorostiza and Jose A. Lopez-Mimbela. The multitype measure branching process. Advances in Applied Probability, 22 (1):49–67, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427596.

Gazdzicki:1993:BPL

[GLM93] Pawel Gazdzicki, Ioannis Lambadaris, and Ravi R. Mazumdar. Blocking probabilities for large multirate Erlang loss systems. Advances in Applied Probability, 25(4):997–1009, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427803.

Gulikers:2017:SMC

[GLM17] Lennart Gulikers, Marc Lelarge, and Laurent Massoulié. A spectral method for community detection in moderately sparse degree-corrected stochastic block models. Advances in Applied Probability, 49(3):686–721, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/spectral-method-for-community-
detection-in-moderately-sparse-degreecorrected-stochastic-
block-models/FAF2E125F2D2889CBEFDFA644BA5F57F.

Gracar:2021:PPT

[GLM21] Peter Gracar, Lukas Lüchtrath, and Peter Mörters. Percolation phase transition in weight-dependent random connection models. Advances in Applied Probability, 53(4): 1090–1114, December 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/percolation-phase-transition-in-
weightdependent-random-connection-models/7B85863F4E9E3FB24BA77F538D1A871A.

Ganassali:2022:SAC

[GLM22] Luca Ganassali, Marc Lelarge, and Laurent Massoulié. Spectral alignment of correlated Gaussian matrices. Advances in Applied Probability, 54(1):279–310, March 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/spectral-alignment-of-correlated-
gaussian-matrices/468F92C33622EC7F9A3FE97EB941602D.

Galerne:2015:RMS

[GLR15] Bruno Galerne and Raphael Lachièze-Rey. Random measurable sets and covariogram realizability problems. Advances in Applied Probability, 47(3):611–639, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1444308874.

Gouet:2001:MAS

[GLS01] Raúl Gouet, F. Javier López, and Miguel San Miguel. A martingale approach to strong convergence of the number of records. Advances in Applied Probability, 33(4):864–873, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428309.

Gouet:2005:CLT

[GLS05] Raul Gouet, F. Javier López, and Gerardo Sanz. Central limit theorems for the number of records in discrete models. Advances in Applied Probability, 37(3):781–800, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037355.

Gouet:2011:ANN

[GLS11] Raul Gouet, F. Javier López, and Gerardo Sanz. Asymptotic normality for the number of records from general distributions. Advances in Applied Probability, 43(2):422–436, June 2011. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/23024513.

Gouet:2015:RSO

[GLS15] Raúl Gouet, F. Javier López, and Gerardo Sanz. Records from stationary observations subject to a random trend. Advances in Applied Probability, 47(4):1175–1189, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1449859805.

Glynn:1982:MPG

[Gly82] Peter W. Glynn. On the Markov property of the GI/G/∞ Gaussian limit. Advances in Applied Probability, 14(1):191– 194, March 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426739.

Glynn:1994:PER

[Gly94] Peter W. Glynn. Poisson’s equation for the recurrent M/G/1 queue. Advances in Applied Probability, 26(4):1044–1062, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427904.

Gani:1971:JDR

[GM71] J. Gani and D. R. McNeil. Joint distributions of random variables and their integrals for certain birth–death and diffusion processes. Advances in Applied Probability, 3(2):339–352, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426175.

Griffiths:1979:CBP

[GM79] R. C. Griffiths and R. K. Milne. A class of bivariate Poisson processes. Advances in Applied Probability, 11(2):272–273, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426801.

Goodall:1991:GDS

[GM91] Colin R. Goodall and Kanti V. Mardia. A geometrical derivation of the shape density. Advances in Applied Probability, 23(3): 496–514, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427619.

Goldie:1996:PPA

[GM96] Charles M. Goldie and Ross A. Maller. A point-process approach to almost-sure behaviour of record values and order statistics. Advances in Applied Probability, 28(2):426–462, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428066.

Griffin:1998:RGO

[GM98] P. S. Griffin and R. A. Maller. On the rate of growth of the overshoot and the maximum partial sum. Advances in Applied Probability, 30(1):181–196, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427883.

Gamarnik:2008:SSA

[GM08] David Gamarnik and Petar Momcilović. Steady-state analysis of a multiserver queue in the Halfin–Whitt regime. Advances in Applied Probability, 40(2):548–577, June 2008. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443595.

Griffin:2011:SET

[GM11] Philip S. Griffin and Ross A. Maller. Stability of the exit time for Lévy processes. Advances in Applied Probability, 43(3):712–734, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
23024572.

Garet:2014:GPB

[GM14] Olivier Garet and Régine Marchand. Growth of a population of bacteria in a dynamical hostile environment. Advances in Applied Probability, 46(3):661–686, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1409319554.

Gut:2018:ASP

[GML18] Allan Gut and Anders Martin-Löf. An asymmetric St. Petersburg game with trimming. Advances in Applied Probability, 50(A):115–129, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/an-asymmetric-st-petersburg-game-
with-trimming/548ACDE1E74479A81457AACCAA737473.

Gutierrez:2023:AFV

[GMS23] Abraham Gutierrez, Sebastian Müller, and Stjepan Sebek. On asymptotic fairness in voting with greedy sampling. Advances in Applied Probability, 55(3):999–1032, September 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/on-
asymptotic-fairness-in-voting-with-greedy-sampling/C5E5D5EB9915DD887ECCBABB297D1F8B.

Gustafson:2023:JW

[GN23] Adam Gustafson and Hariharan Narayanan. John’s walk. Advances in Applied Probability, 55(2):473–491, June 2023. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/johns-walk/F268B1A4A0A952F71CC9333D9282C847.

Gneiting:1999:ICF

[Gne99] Tilmann Gneiting. Isotropic correlation functions on ddimensional balls. Advances in Applied Probability, 31(3):625– 631, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428387.

Gneiting:2000:AIC

[Gne00] Tilmann Gneiting. Addendum to ’isotropic correlation functions on d-dimensional balls’. Advances in Applied Probability, 32(4): 960–961, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428510.

Glazebrook:2002:IPC

[GNMA02] K. D. Glazebrook, J. Niño-Mora, and P. S. Ansell. Index policies for a class of discounted restless bandits. Advances in Applied Probability, 34(4):754–774, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428321.

Giorno:1986:SDA

[GNR86] V. Giorno, A. G. Nobile, and L. M. Ricciardi. On some diffusion approximations to queueing systems. Advances in Applied Probability, 18(4):991–1014, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427259.

Giorno:1987:STN

[GNR87] V. Giorno, A. G. Nobile, and L. M. Ricciardi. On some timenon-homogeneous diffusion approximations to queueing systems. Advances in Applied Probability, 19(4):974–994, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427111.

Giorno:1990:ABF

[GNR90] V. Giorno, A. G. Nobile, and L. M. Ricciardi. On the asymptotic behaviour of first-passage-time densities for one-dimensional diffusion processes and varying boundaries. Advances in Applied Probability, 22(4):883–914, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427567.

Ghosh:2017:RTF

[GNR17] Arka Ghosh, Steven Noren, and Alexander Roitershtein. On the range of the transient frog model on Z. Advances in Applied Probability, 49(2):327–343, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-the-range-of-the-transient-frog-
model-on/E9187C66042A98D73D04167566EF28E5.

Giorno:1989:EFP

[GNRS89] V. Giorno, A. G. Nobile, L. M. Ricciardi, and S. Sato. On the evaluation of first-passage-time probability densities via nonsingular integral equations. Advances in Applied Probability, 21 (1):20–36, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427196.

Ghurye:1973:IDL

[GO73] S. G. Ghurye and I. Olkin. Identically distributed linear forms and the normal distribution. Advances in Applied Probability, 5 (1):138–152, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425968.

Godambe:1985:OEF

[God85] V. P. Godambe. Optimal estimation for finite samples in stochastic processes. Advances in Applied Probability, 17(2): 247–249, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427133.

Godbole:1991:PAR

[God91] Anant P. Godbole. Poisson approximations for runs and patterns of rare events. Advances in Applied Probability, 23(4): 851–865, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427680.

Goffard:2019:FRA

[Gof19] Pierre-O. Goffard. Fraud risk assessment within blockchain transactions. Advances in Applied Probability, 51(2):443–467, June 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/fraud-
risk-assessment-within-blockchain-transactions/DC96574C5098794A8345167F69149A44.

Goldie:1977:CTE

[Gol77] Charles M. Goldie. Convergence theorems for empirical Lorenz curves and their inverses. Advances in Applied Probability, 9(4): 765–791, December 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426700.

Goldenschluger:2016:NES

[Gol16] Alexander Goldenschluger. Nonparametric estimation of the service time distribution in the M/G/∞ queue. Advances in Applied Probability, 48(4):1117–1138, December 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1482548431.

Gorroochurn:2001:PDI

[Gor01] Prakash Gorroochurn. Post-data inference of coalescence times and segregating-site distribution in a two-Island model with symmetric migration. Advances in Applied Probability, 33(3): 600–616, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428477.

Gianini-Pettitt:1979:OSB

[GP79] Jacqueline Gianini-Pettitt. Optimal selection based on relative ranks with a random number of individuals. Advances in Applied Probability, 11(4):720–736, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426856.

Greenwood:1980:FIL

[GP80a] Priscilla Greenwood and Jim Pitman. Fluctuation identities for Lévy processes and splitting at the maximum. Advances in Applied Probability, 12(4):893–902, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426747.

Greenwood:1980:FIR

[GP80b] Priscilla Greenwood and Jim Pitman. Fluctuation identities for random walk by path decomposition at the maximum. Advances in Applied Probability, 12(2):291–293, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426564.

Griffiths:1988:IAV

[GP88] R. C. Griffiths and Anthony G. Pakes. An infinite-alleles version of the simple branching process. Advances in Applied Probability, 20(3):489–524, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427033.

Guillotin-Plantard:2002:LDM

[GP02] Nadine Guillotin-Plantard. Large deviations for a Markov chain in a random landscape. Advances in Applied Probability, 34 (2):375–393, June 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428293.

Gad:2015:VOS

[GP15] Kamille Sofie Tågholt Gad and Jesper Lund Pedersen. Variance optimal stopping for geometric Lévy processes. Advances in Applied Probability, 47(1):128–145, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1427814584.

Geluk:2000:CHT

[GPD00] Jaap L. Geluk, Liang Peng, and Casper G. De Vries. Convolutions of heavy-tailed random variables and applications to portfolio diversification and MA(1) time series. Advances in Applied Probability, 32(4):1011–1026, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428514.

Garcia-Pareja:2021:ESC

[GPHK21] Celia García-Pareja, Henrik Hult, and Timo Koski. Exact simulation of coupled Wright–Fisher diffusions. Advances in Applied Probability, 53(4):923–950, December 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/exact-simulation-of-coupled-wrightfisher-
diffusions/68CB174EA6A2D161D1134598B0076CFF.

Goldie:1988:DBS

[GR88] Charles M. Goldie and Sidney Resnick. Distributions that are both subexponential and in the domain of attraction of an extreme-value distribution. Advances in Applied Probability, 20 (4):706–718, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427356.

Grubel:1996:ADT

[GR96] Rudolf Grübel and Uwe Rösler. Asymptotic distribution theory for Hoare’s selection algorithm. Advances in Applied Probability, 28(1):252–269, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427920.

Gentil:2008:USS

[GR08] Ivan Gentil and Bruno Rémillard. Using systematic sampling selection for Monte Carlo solutions of Feynman–Kac equations.

Advances in Applied Probability, 40(2):454–472, June 2008. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443591.

Guerin:2016:JDS

[GR16] Hélène Guérin and Jean-François Renaud. Joint distribution of a spectrally negative Lévy process and its occupation time, with step option pricing in view. Advances in Applied Probability, 48 (1):274–297, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1457466166.

Grandell:1977:PPR

[Gra77] Jan Grandell. Point processes and random measures. Advances in Applied Probability, 9(3):502–526, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426111. See

correction [Ano77e].

Grandell:1982:MMV

[Gra82] Jan Grandell. Mathematical models for the variation of airpollutant concentrations. Advances in Applied Probability, 14 (2):240–256, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426519.

Greenwood:1975:WHM

[Gre75] Priscilla Greenwood. Wiener–Hopf methods, decompositions, and factorisation identities for maxima and minima of homogeneous random processes. Advances in Applied Probability, 7(4): 767–785, December 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426398.

Green:1978:DAT

[Gre78] H. S. Green. Distribution of arrival times in cosmic ray showers. Advances in Applied Probability, 10(4):730–735, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426654.

Greenberg:1987:NCT

[Gre87] Betsy S. Greenberg. On Niu’s conjecture for tandem queues. Advances in Applied Probability, 19(3):751–753, September 1987.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427419.

Grey:1989:NEM

[Gre89a] D. R. Grey. A note on explosiveness of Markov branching processes. Advances in Applied Probability, 21(1):226–228, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427205.

Grey:1989:PRW

[Gre89b] D. R. Grey. Persistent random walks May have arbitrarily large tails. Advances in Applied Probability, 21(1):229–230, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427206.

Glazebrook:2006:SIF

[GRHK06] K. D. Glazebrook, D. Ruiz-Hernandez, and C. Kirkbride. Some indexable families of restless bandit problems. Advances in Applied Probability, 38(3):643–672, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443465.

Griffeath:1975:ETG

[Gri75a] David Griffeath. Ergodic theorems for graph interactions. Advances in Applied Probability, 7(1):179–194, March 1975. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425859.

Gringorten:1975:USM

[Gri75b] Irving I. Gringorten. On the use of stochastic models of simultaneous critical weather conditions in an area. Advances in Applied Probability, 7(2):260–261, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426074.

Grigelionis:1977:RPM

[Gri77] B. Grigelionis. Random point-measures with non-random dual predictable projections. Advances in Applied Probability, 9(3): 439–440, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426100.

Griffiths:1979:ESD

[Gri79a] R. C. Griffiths. Exact sampling distributions from the infinite neutral alleles model. Advances in Applied Probability, 11 (2):326–354, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426843.

Griffiths:1979:TDE

[Gri79b] R. C. Griffiths. A transition density expansion for a multi-allele diffusion model. Advances in Applied Probability, 11(2):310– 325, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426842.

Griffeath:1980:SRB

[Gri80a] David Griffeath. Some remarks on the basic contact processes. Advances in Applied Probability, 12(2):269, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426535.

Griffiths:1980:SMD

[Gri80b] J. D. Griffiths. A stochastic model for delays at a non-signalized pedestrian crossing. Advances in Applied Probability, 12(2): 307, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426584.

Grimmett:1981:CSD

[Gri81] G. R. Grimmett. Critical sponge dimensions in percolation theory. Advances in Applied Probability, 13(2):314–324, June 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426686.

Griffiths:1982:NAS

[Gri82] R. C. Griffiths. The number of alleles and segregating sites in a sample from the infinite-alleles model. Advances in Applied Probability, 14(2):225–239, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426518.

Grishechkin:1992:RBP

[Gri92] Sergei Grishechkin. On a relationship between processor-sharing queues and Crump–Mode–Jagers branching processes. Advances in Applied Probability, 24(3):653–698, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427484.

Grishechkin:1994:GGP

[Gri94] Sergei Grishechkin. GI/G/1 processor sharing queue in heavy traffic. Advances in Applied Probability, 26(2):539–555, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427450.

Griffiths:2006:CLD

[Gri06] Robert C. Griffiths. Coalescent lineage distributions. Advances in Applied Probability, 38(2):405–429, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443448.

Griffiths:2014:FVP

[Gri14] Robert C. Griffiths. The Λ–Fleming–Viot process and a connection with Wright–Fisher diffusion. Advances in Applied Probability, 46(4):1009–1035, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1418396241.

Groenewegen:1978:COC

[Gro78] Luuk P. J. Groenewegen. Characterization of optimal controls. Advances in Applied Probability, 10(2):326–327, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426925.

Groeneboom:2012:CHU

[Gro12] Piet Groeneboom. Convex hulls of uniform samples from a convex polygon. Advances in Applied Probability, 44(2):330– 342, June 2012. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
41714053.

Gan:2017:DAE

[GRR17] Han L. Gan, Adrian Röllin, and Nathan Ross. Dirichlet approximation of equilibrium distributions in Cannings models with mutation. Advances in Applied Probability, 49 (3):927–959, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/dirichlet-approximation-of-equilibrium-
distributions-in-cannings-models-with-mutation/580720B87566A25E1E9AEE7A32B4D64A.

Ghosh:2010:OCS

[GRW10] Arka P. Ghosh, Alexander Roitershtein, and Ananda Weerasinghe. Optimal control of a stochastic processing system driven by a fractional Brownian motion input. Advances in Applied Probability, 42(1):183–209, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/25683812.

Guillemin:2004:TAP

[GRZ04] Fabrice Guillemin, Philippe Robert, and Bert Zwart. Tail asymptotics for processor-sharing queues. Advances in Applied Probability, 36(2):525–543, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428465.

Greenwood:1977:FTM

[GS77a] Priscilla Greenwood and Moshe Shaked. Fluctuation theory for multidimensional processes and applications to diverse storage systems. Advances in Applied Probability, 9(2):195, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426343.

Greenwood:1977:FRW

[GS77b] Priscilla Greenwood and Moshe Shaked. Fluctuations of random walk in R^{d} and storage systems. Advances in Applied Probability, 9(3):566–587, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426115.

Gaenssler:1978:SCR

[GS78] Peter Gaenssler and Winifried Stute. On some criteria for the realization of stochastic processes. Advances in Applied Probability, 10(2):281, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426885.

Godbole:1993:IPA

[GS93] Anant P. Godbole and Andrew A. Schaffner. Improved Poisson approximations for word patterns. Advances in Applied Probability, 25(2):334–347, June 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427656.

Guillemin:1995:TCM

[GS95] Fabrice Guillemin and Alain Simonian. Transient characteristics of an M/M/∞ system. Advances in Applied Probability, 27(3): 862–888, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428137.

Gapeev:2013:BQD

[GS13] Pavel V. Gapeev and Albert N. Shiryaev. Bayesian quickest detection problems for some diffusion processes. Advances in Applied Probability, 45(1):164–185, March 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/41806004.

Grublyte:2014:PSE

[GS14] Ieva Grublyte and Donatas Surgailis. Projective stochastic equations and nonlinear long memory. Advances in Applied Probability, 46(4):1084–1105, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1418396244.

Garavaglia:2019:SPA

[GS19] Alessandro Garavaglia and Clara Stegehuis. Subgraphs in preferential attachment models. Advances in Applied Probability, 51(3):898–926, September 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/subgraphs-in-preferential-attachment-
models/7368A4DC7190D8504F656B1FF6872FC0.

Gneiting:2001:ACB

[GSS01] Tilmann Gneiting, Zoltán Sasvári, and Martin Schlather. Analogies and correspondences between variograms and covariance functions. Advances in Applied Probability, 33(3):617–630, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428478.

Gilchrist:1975:SPB

[GT75] James H. Gilchrist and John B. Thomas. A shot process with burst properties. Advances in Applied Probability, 7(3):527–541, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426126.

Greenwood:1978:GSR

[GT78] P. Greenwood and J. Teugels. Geometrically stopped random walk. Advances in Applied Probability, 10(2):294–295, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426898.

Gani:1984:ITS

[GT84] J. Gani and Pyke Tin. Immigration in the two-sex population process. Advances in Applied Probability, 16(3):687–689, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427294.

Grassmann:2008:SSS

[GT08] Winfried K. Grassmann and Javad Tavakoli. Stochastic and substochastic solutions for infinite-state Markov chains with applications to matrix-analytic methods. Advances in Applied Probability, 40(4):1157–1173, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443628.

Genadot:2012:AFC

[GT12] Alexandre Genadot and Michèle Thieullen. Averaging for a fully coupled piecewise-deterministic Markov process in infinite dimensions. Advances in Applied Probability, 44(3):749–773, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
41714074.

Gergely:1970:MCG

[GTY70] T. Gergely, I. N. Tsukanow, and I. I. Yezhow. Markov chains governed by complicated renewal processes. Advances in Applied Probability, 2(2):287–322, Autumn 1970. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426321.

Guerry:1993:PAS

[Gue93] M. A. Guerry. The probability of attaining a structure in a partially stochastic model. Advances in Applied Probability, 25 (4):818–824, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427793.

Gupta:1980:SCR

[Gup80] Ramesh C. Gupta. Some characterizations of renewal densities with emphasis in reliability. Advances in Applied Probability, 12 (2):311–312, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426589.

Gut:1974:LTS

[Gut74] A. Gut. Limit theorems for some first passage times. Advances in Applied Probability, 6(2):241–242, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426278.

Gut:1990:CSM

[Gut90] Allan Gut. Cumulative shock models. Advances in Applied Probability, 22(2):504–507, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427554.

Gopalan:1976:PAT

[GV76] M. N. Gopalan and P. Venkatachalam. Probabilistic analysis of a two-server n-unit system subject to preventive maintenance and repair. Advances in Applied Probability, 8(2):249–250, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425899.

Gans:1998:ODS

[GV98] Noah Gans and Garrett Van Ryzin. Optimal dynamic scheduling of a general class of parallel-processing queueing systems. Advances in Applied Probability, 30(4):1130–1156, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428033.

Gefferth:2003:NDS

[GVM^{+}03] A. Gefferth, D. Veitch, I. Maricza, S. Molnár, and I. Ruzsa. The nature of discrete second-order self-similarity. Advances in Applied Probability, 35(2):395–416, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428429.

Grinfeld:2015:CMR

[GVW15] Michael Grinfeld, Stanislav Volkov, and Andrew R. Wade. Convergence in a multidimensional randomized Keynesian beauty contest. Advances in Applied Probability, 47(1):57–82, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1427814581.

Guo:2013:ACT

[GVZ13] Xianping Guo, Mantas Vykertas, and Yi Zhang. Absorbing continuous-time Markov decision processes with total cost criteria. Advances in Applied Probability, 45(2):490–519, June 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1370870127.

Gates:1978:ZIM

[GW78] D. J. Gates and M. Westcott. Zone of influence models for competition in plantations. Advances in Applied Probability, 10 (3):499–537, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426632.

Glynn:1991:NVH

[GW91] Peter W. Glynn and Ward Whitt. A new view of the heavytraffic limit theorem for infinite-server queues. Advances in Applied Probability, 23(1):188–209, March 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427517.

Glazebrook:1992:SMS

[GW92] K. D. Glazebrook and Lyn R. Whitaker. Single-machine stochastic scheduling with dependent processing times. Advances in Applied Probability, 24(3):635–652, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427483.

Glasserman:2000:DDB

[GW00] Paul Glasserman and Hui Wang. Discretization of deflated bond prices. Advances in Applied Probability, 32(2):540–563, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428203.

Guo:2013:FIT

[GW13] Fenglong Guo and Dingcheng Wang. Finite- and infinite-time ruin probabilities with general stochastic investment return processes and bivariate upper tail independent and heavy-tailed claims. Advances in Applied Probability, 45(1):241–273, March 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/41806007.

Gao:2020:FCL

[GW20] Fuqing Gao and Yujing Wang. Functional central limit theorems and moderate deviations for Poisson cluster processes. Advances in Applied Probability, 52(3):916–941, September 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/functional-central-limit-
theorems-and-moderate-deviations-for-poisson-cluster-
processes/3F006180D31D83FD752CF3DB55571532.

Ganesh:2007:CDS

[GX07] Ayalvadi Ganesh and Feng Xue. On the connectivity and diameter of small-world networks. Advances in Applied Probability, 39(4):853–863, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443547.

Gelenbe:1978:QSW

[GY78] E. Gelenbe and R. Yasnogorodsky. A queue with server of walking type (autonomous service). Advances in Applied Probability, 10(2):317, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426917.

Guo:2010:NDA

[GY10] Xianping Guo and Liuer Ye. New discount and average optimality conditions for continuous-time Markov decision processes. Advances in Applied Probability, 42(4):953–985, December 2010.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/25799079.

Grigorescu:2016:MVT

[GY16] Ilie Grigorescu and Yi-Ching Yao. Maximizing the variance of the time to ruin in a multiplayer game with selection. Advances in Applied Probability, 48(2):610–630, June 2016. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1465490765.

Gun:2017:FLP

[GY17] Onur Gün and Atilla Yilmaz. Fluid limit for the Poisson encounter-mating model. Advances in Applied Probability, 49 (4):1201–1229, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/fluid-limit-for-the-poisson-encountermating-
model/29555BDC995E4535656C06AEDD4F1E88.

Gyires:1975:ETL

[Gyi75] Béla Gyires. On an extension of a theorem of Linnik and Zinger. Advances in Applied Probability, 7(4):786–800, December 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426399.

Grey:1993:ABE

[GZ93] D. R. Grey and Lu Zhunwei. The asymptotic behaviour of extinction probability in the Smith–Wilkinson branching process. Advances in Applied Probability, 25(2):263–289, June 1993. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427652.

Goldstein:2011:BEB

[GZ11] Larry Goldstein and Haimeng Zhang. A Berry–Esseen bound for the lightbulb process. Advances in Applied Probability, 43(3): 875–898, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/23024578.

Gzyl:1976:SRA

[Gzy76] Cristina Gzyl. Some results about Markov processes on subsets of the state space. Advances in Applied Probability, 8(3): 517–530, September 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426142.

Gzyl:1980:IGT

[Gzy80] Henryk Gzyl. Infinitesimal generators of time-changed processes. Advances in Applied Probability, 12(2):269–270, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426536.

Hill:1979:VQT

[HA79] W. G. Hill and P. J. Avery. Variation in quantitative traits and heterozygosity as influenced by linkage and mating structure. Advances in Applied Probability, 11(1):4–5, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426759.

Hashemi:2020:CFC

[HA20] M. Hashemi and M. Asadi. On component failure in coherent systems with applications to maintenance strategies. Advances in Applied Probability, 52(4):1197–1223, December 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/on-
component-failure-in-coherent-systems-with-applications-
to-maintenance-strategies/2983940AA516442E5F09DC7ED5400ED9.

Hadidi:1974:BPQ

[Had74] N. Hadidi. A basically Poisson queue with non-Poisson output. Advances in Applied Probability, 6(2):254–255, June 1974. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426289.

Hadidi:1976:ASC

[Had76] Nasser Hadidi. Appointment scheme as compared to random service and arrival times from specific distributions. Advances in Applied Probability, 8(2):250–252, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425900.

Hadidi:1979:SME

[Had79] Nasser Hadidi. A stochastic model for the emergence of poets. Advances in Applied Probability, 11(2):305, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426836.

Hadidi:1980:PCQ

[Had80] Nasser Hadidi. Partially correlated queues. Advances in Applied Probability, 12(2):308–310, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426586.

Haggstrom:2000:MRF

[Häg00] Olle Häggström. Markov random fields and percolation on general graphs. Advances in Applied Probability, 32(1):39–66, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428225.

Hahn:1998:ELD

[Hah98] Ute Hahn. Estimation of local density in gradient surface and fibre processes. Advances in Applied Probability, 30(2):279– 280, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427949.

Haigh:1975:GTE

[Hai75] John Haigh. Game theory and evolution. Advances in Applied Probability, 7(1):8–11, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425844.

Haigh:1980:PTP

[Hai80] John Haigh. Persistence times of populations subject to random disasters. Advances in Applied Probability, 12(1):7, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426483.

Hajnal:1976:ANC

[Haj76] J. Hajnal. The absolute numbers of consanguineous marriages. Advances in Applied Probability, 8(4):630–635, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425921.

Hajek:1982:HTO

[Haj82] Bruce Hajek. Hitting-time and occupation-time bounds implied by drift analysis with applications. Advances in Applied Probability, 14(3):502–525, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1426671.

Haken:1982:SAC

[Hak82] H. Haken. Synergetics: An approach to complex dynamic systems. Advances in Applied Probability, 14(2):197, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426508.

Hall:1980:EPN

[Hal80] Peter Hall. Estimating probabilities for normal extremes. Advances in Applied Probability, 12(2):491–500, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426608.

Hall:1985:MPL

[Hal85a] Peter Hall. Macroscopic properties of a linear mosaic. Advances in Applied Probability, 17(2):330–346, June 1985. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427144.

Hall:1985:IOD

[Hal85b] Peter Hall. On inference in a one-dimensional mosaic and an M/G/∞ queue. Advances in Applied Probability, 17(1):210– 229, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427060.

Hallin:1986:NSD

[Hal86] Marc Hallin. Non-stationary q-dependent processes and timevarying moving-average models: Invertibility properties and the forecasting problem. Advances in Applied Probability, 18(1): 170–210, March 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427242.

Hall:1987:ADC

[Hal87] Peter Hall. On the amount of detail that can be recovered from a degraded signal. Advances in Applied Probability, 19(2):371– 395, June 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427424.

Hannan:1978:RCT

[Han78] E. J. Hannan. Rates of convergence for time series regression. Advances in Applied Probability, 10(4):740–743, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426656.

Hannan:1979:LS

[Han79] E. J. Hannan. Linear systems. Advances in Applied Probability, 11(2):263–264, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426794.

Hansen:1996:BTE

[Han96] Martin Bøgsted Hansen. Bayes type estimators for the empty space function. Advances in Applied Probability, 28(2):333, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428042.

Hansen:2007:ALM

[Han07] Niels Richard Hansen. Asymptotics for local maximal stack scores with general loop penalty function. Advances in Applied Probability, 39(3):776–798, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443540.

Harding:1971:PRT

[Har71] E. F. Harding. The probabilities of rooted tree-shapes generated by random bifurcation. Advances in Applied Probability, 3(1):44–77, Spring 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426329.

Harrison:1977:SDL

[Har77] J. Michael Harrison. The supremum distribution of a Lévy process with no negative jumps. Advances in Applied Probability, 9(2):417–422, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426392.

Harrison:1978:DAT

[Har78] J. Michael Harrison. The diffusion approximation for tandem queues in heavy traffic. Advances in Applied Probability, 10(4):

886–905, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426665.

Harrison:1980:MSI

[Har80] J. Michael Harrison. Martingales and stochastic integrals in the theory of securities markets. Advances in Applied Probability, 12(2):270–271, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426537.

Harnqvist:1981:LTP

[Här81] Martin Härnqvist. Limit theorems for point processes generated in a general branching process. Advances in Applied Probability, 13(4):650–668, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426965.

Harrison:1984:NCT

[Har84] P. G. Harrison. A note on cycle times in tree-like queueing networks. Advances in Applied Probability, 16(1):216–219, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427233.

Haslett:1976:CMR

[Has76] John Haslett. The control of a multipurpose reservoir. Advances in Applied Probability, 8(3):592–609, September 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426146.

Haugen:1978:BPG

[Hau78] R. B. Haugen. On the busy period of GI/G/1 queue with two priority levels and feedbacks. Advances in Applied Probability, 10(2):335–336, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426935.

Hansen:1999:FCD

[HBG99] Martin B. Hansen, Adrian J. Baddeley, and Richard D. Gill. First contact distributions for spatial patterns: Regularity and estimation. Advances in Applied Probability, 31(1):15–33, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427988.

Heine:2008:UAD

[HC08] Kari Heine and Dan Crisan. Uniform approximations of discrete-time filters. Advances in Applied Probability, 40(4): 979–1001, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443621.

Helterbrand:1994:SAI

[HCD94] Jeffrey D. Helterbrand, Noel Cressie, and Jennifer L. Davidson. A statistical approach to identifying closed object boundaries in images. Advances in Applied Probability, 26(4):831–854, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427893.

He:1996:QMC

[He96] Qi-Ming He. Queues with marked customers. Advances in Applied Probability, 28(2):567–587, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428072.

He:2016:MLL

[He16] Hui He. Marcinkiewicz law of large numbers for outer products of heavy-tailed, long-range dependent data. Advances in Applied Probability, 48(3):672–690, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1474296309.

Heathcote:1978:PDE

[Hea78] C. R. Heathcote. On parametric density estimators. Advances in Applied Probability, 10(4):735–740, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426655.

Heinrich:1998:CCL

[Hei98] Lothar Heinrich. Contact and chord length distribution of a stationary Voronoi tessellation. Advances in Applied Probability, 30(3):603–618, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428140.

Helland:1975:CEP

[Hel75] Inge S. Helland. The condition for extinction with probability one in a birth, death and migration process. Advances in Applied Probability, 7(1):61–65, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425853.

Henderson:1983:IRM

[Hen83] W. Henderson. Insensitivity and reversed Markov processes. Advances in Applied Probability, 15(4):752–768, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427322.

Henze:1987:FRP

[Hen87] Norbert Henze. On the fraction of random points with specified nearest-neighbour interrelations and degree of attraction. Advances in Applied Probability, 19(4):873–895, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427106.

Hering:1973:ABI

[Her73] H. Hering. Asymptotic behaviour of immigration-branching processes with general set of types. i: Critical branching part. Advances in Applied Probability, 5(3):391–416, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425827.

Hering:1975:ABI

[Her75] H. Hering. Asymptotic behaviour of immigration-branching processes with general set of types. II: Supercritical branching part. Advances in Applied Probability, 7(3):468–494, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426123.

Heuch:1979:EPP

[Heu79] Ivar Heuch. Equilibrium populations of plants with diplodiplo incompatibility. Advances in Applied Probability, 11(1): 3–4, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426758.

Heuch:1980:FII

[Heu80] Ivar Heuch. Further investigations of incompatibility systems in angiosperm plants. Advances in Applied Probability, 12(1): 7, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426484.

Heyde:1974:LTS

[Hey74a] C. C. Heyde. Limit theory for stationary processes via approximating martingales. Advances in Applied Probability, 6(2):196– 197, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426244.

Heyde:1974:EVO

[Hey74b] C. C. Heyde. On estimating the variance of the offspring distribution in a simple branching process. Advances in Applied Probability, 6(3):421–433, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426225.

Heyde:1975:MME

[Hey75] C. C. Heyde. Martingale methods in estimation theory. Advances in Applied Probability, 7(2):234–236, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426049.

Heyde:1977:RCM

[Hey77] C. C. Heyde. On the rate of convergence in the martingale convergence theorem. Advances in Applied Probability, 9(2): 196, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426344.

Heyde:1978:ECC

[Hey78] C. C. Heyde. On an explanation for the characteristic clutch size of some bird species. Advances in Applied Probability, 10(4): 723–725, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426652.

Heyde:1979:CLI

[Hey79] C. C. Heyde. On central limit and iterated logarithm results for subadditive processes. Advances in Applied Probability, 11 (2):283–284, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426814.

Heyde:1985:LIS

[Hey85] C. C. Heyde. Lectures on inference for stochastic processes. Advances in Applied Probability, 17(2):244–245, June 1985. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427130.

Hodge:2015:AOG

[HG15] D. J. Hodge and K. D. Glazebrook. On the asymptotic optimality of greedy index heuristics for multi-action restless bandits. Advances in Applied Probability, 47(3):652–667, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1444308876.

Heuch:1981:SML

[HH81] I. Heuch and M. Homleid. A simple model for linked loci with recombination values depending on the genotype at one locus. Advances in Applied Probability, 13(1):3, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426454.

Holst:1985:SUS

[HH85] Lars Holst and Jürg Hüsler. Sequential urn schemes and birth processes. Advances in Applied Probability, 17(2):257–279, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427140.

Heidergott:2003:TSE

[HH03] Bernd Heidergott and Arie Hordijk. Taylor series expansions for stationary Markov chains. Advances in Applied Probability, 35(4):1046–1070, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428499. See correction [Ano04g].

Horrmann:2014:VZC

[HH14] Julia Hörrmann and Daniel Hug. On the volume of the zero cell of a class of isotropic Poisson hyperplane tessellations. Advances in Applied Probability, 46(3):622–642, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1409319552.

Haan:2020:TAE

[HHBV20] Roland De Haan, Ahmad Al Hanbali, Richard J. Boucherie, and Jan-Kees Van Ommeren. Transient analysis for exponential time-limited polling models under the preemptive repeat random policy. Advances in Applied Probability, 52(1):32–60, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/transient-analysis-for-exponential-
timelimited-polling-models-under-the-preemptive-repeat-
random-policy/BAE973ED2F0BAB24AE3D82446F1DF7D2.

He:2019:MBP

[HHHT19] Qi-Ming He, Gábor Horváth, Illés Horváth, and Miklós Telek. Moment bounds of PH distributions with infinite or finite support based on the steepest increase property. Advances in Applied Probability, 51(1):168–183, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/moment-bounds-of-ph-distributions-
with-infinite-or-finite-support-based-on-the-steepest-
increase-property/B8DA9989530FB6A2685DF9AF69BFD424.

Heidelberger:1979:CSS

[HI79] Philip Heidelberger and Donald L. Iglehart. Comparing stochastic systems using regenerative simulation with Common Random Numbers. Advances in Applied Probability, 11(4):804–819, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426860.

Hill:1974:LDB

[Hil74] William G. Hill. Linkage disequilibrium between neutral genes in finite populations. Advances in Applied Probability, 6(1):13– 15, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426201.

Hill:1975:ASP

[Hil75] William G. Hill. Artificial selection in populations with overlapping generations. Advances in Applied Probability, 7(1):11– 14, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425845.

Hill:1976:EFS

[Hil76a] William G. Hill. Effect of family structure in small populations on selection differentials and correlation of relatives’ performance for quantitative traits. Advances in Applied Probability, 8(4):636–639, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425922.

Hill:1976:LDA

[Hil76b] William G. Hill. Linkage disequilibrium among neutral mutant alleles in finite population. Advances in Applied Probability, 8 (1):10–12, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426012.

Hill:1981:EEP

[Hil81] W. G. Hill. Estimation of effective population size from data on linked genes. Advances in Applied Probability, 13(1):4, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426455.

Hines:1994:EMDa

[Hin94a] W. G. S. Hines. ESS modelling of diploid populations i: Anatomy of one-locus allelic frequency simplices. Advances in Applied Probability, 26(2):341–360, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427440.

Hines:1994:EMDb

[Hin94b] W. G. S. Hines. ESS modelling of diploid populations II: Stability analysis of possible equilibria. Advances in Applied Probability, 26(2):361–376, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427441.

Hiorns:1979:CIH

[Hio79] Robert W. Hiorns. The consequences of intermixture in human populations. Advances in Applied Probability, 11(1):5–7, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426760.

Hiraba:2000:JTF

[Hir00] Seiji Hiraba. Jump-type Fleming–Viot processes. Advances in Applied Probability, 32(1):140–158, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428231.

Hirscher:2017:ODA

[Hir17] Timo Hirscher. Overly determined agents prevent consensus in a generalized Deffuant model on Z with dispersed opinions. Advances in Applied Probability, 49(3): 722–744, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/overly-determined-agents-prevent-
consensus-in-a-generalized-deffuant-model-on-with-dispersed-
opinions/A5822ED91C011AE5DA734A19849FC322.

Hordijk:1976:DTM

[HIS76] Arie Hordijk, Donald L. Iglehart, and Rolf Schassberger. Discrete time methods for simulating continuous time Markov chains. Advances in Applied Probability, 8(4):772–788, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425933.

Hobolth:2000:MSC

[HJ00] Asger Hobolth and Eva B. Vedel Jensen. Modelling stochastic changes in curve shape, with an application to cancer diagnostics. Advances in Applied Probability, 32(2):344–362, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428192.

Hobolth:2002:NDB

[HJ02] Asger Hobolth and Eva B. Vedel Jensen. A note on design-based versus model-based variance estimation in stereology. Advances in Applied Probability, 34(3):484–490, September 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428242.

Hambly:2003:TTP

[HJ03] B. M. Hambly and O. D. Jones. Thick and thin points for random recursive fractals. Advances in Applied Probability, 35 (1):251–277, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428282.

Hambly:2004:RHL

[HJ04] B. M. Hambly and Jonathan Jordan. A random hierarchical lattice: The series-parallel graph and its properties. Advances in Applied Probability, 36(3):824–838, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/4140410.

Hansen:2008:LPR

[HJ08] Jennie C. Hansen and Jerzy Jaworski. Local properties of random mappings with exchangeable in-degrees. Advances in Applied Probability, 40(1):183–205, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443576.

Hamza:2005:EMM

[HJK05] Kais Hamza, Saul Jacka, and Fima Klebaner. The equivalent martingale measure conditions in a general model for interest rates. Advances in Applied Probability, 37(2):415–434, June 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/30037334.

Hirsch:2016:LDP

[HJKP16] Christian Hirsch, Benedikt Jahnel, Paul Keeler, and Robert I. A. Patterson. Large deviation principles for connectable receivers in wireless networks. Advances in Applied Probability, 48(4):1061–1094, December 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1482548429.

Hirsch:2017:TFD

[HJKP17] Christian Hirsch, Benedikt Jahnel, Paul Keeler, and Robert I. A. Patterson. Traffic flow densities in large transport networks. Advances in Applied Probability, 49(4):

1091–1115, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/traffic-flow-densities-in-large-
transport-networks/EB513FE70BEA3D3E2F11746D254B212A.

Huang:2016:MMO

[HJM^{+}16] G. Huang, H. M. Jansen, M. Mandjes, P. Spreij, and K. De Turck. Markov-modulated Ornstein–Uhlenbeck processes. Advances in Applied Probability, 48(1):235–254, March 2016. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1457466164.

Hahn:2003:ISP

[HJVN03] Ute Hahn, Eva B. Vedel Jensen, Marie-Colette Van Lieshout, and Linda Stougaard Nielsen. Inhomogeneous spatial point processes by location-dependent scaling. Advances in Applied Probability, 35(2):319–336, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428423.

Hannan:1984:MLT

[HK84] E. J. Hannan and L. Kavalieris. Multivariate linear time series models. Advances in Applied Probability, 16(3):492–561, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427286.

Heath:1988:LIL

[HK88] David Heath and Robert P. Kertz. Leaving an interval in limited playing time. Advances in Applied Probability, 20(3):635–645, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427039.

Hirayama:1989:EPF

[HK89a] Tetsuji Hirayama and Masaaki Kijima. An extremal property of FIFO discipline in G/IFR/1 queues. Advances in Applied Probability, 21(2):481–484, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427174.

Hunt:1989:CLL

[HK89b] P. J. Hunt and F. P. Kelly. On critically loaded loss networks. Advances in Applied Probability, 21(4):831–841, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427769.

Hall:1990:CIM

[HK90] Peter Hall and Inge Koch. On continuous image models and image analysis in the presence of correlated noise. Advances in Applied Probability, 22(2):332–349, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427539.

Hordijk:1992:ROS

[HK92] Arie Hordijk and Ger Koole. The μc-rule is not optimal in the second node of the tandem queue: A counterexample. Advances in Applied Probability, 24(1):234–237, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427742.

Hordijk:1993:OLR

[HK93] Arie Hordijk and Ger Koole. On the optimality of LEPT and μc rules for parallel processors and dependent arrival processes. Advances in Applied Probability, 25(4):979–996, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427802.

Houdre:2022:NAF

[HK22] Christian Houdré and George Kerchev. Normal approximation for functions of hidden Markov models. Advances in Applied Probability, 54(2):536–569, June 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/normal-approximation-for-functions-
of-hidden-markov-models/9E8FB5AF8BE5A0B9735E13AC9CCB297A.

Hiorns:1976:SMG

[HKH76] R. W. Hiorns, J. Kirby, and G. A. Harrison. Selective outmigration: The genetic effects in human populations. Advances in Applied Probability, 8(1):12–15, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426013.

Heyde:1984:EGM

[HKM^{+}84] C. C. Heyde, Richard J. Kryscio, Douglas R. Miller, Simon Tavaré, Warren W. Esty, Colm Art O’Cinneide, R. R. Kocherlakota, Rakesh K. Shukla, and R. C. Srivastava. Epidemic and genetic models. Advances in Applied Probability, 16(1):26– 29, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427222.

Hall:1991:ENV

[HKT91] Peter Hall, J. W. Kay, and D. M. Titterington. On estimation of noise variance in two-dimensional signal processing. Advances in Applied Probability, 23(3):476–495, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427618.

Haigh:1985:CPM

[HL85] John Haigh and Richard P. Lewis. Conditional probabilities in Moran’s model. Advances in Applied Probability, 17(2):460– 462, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427151.

Hult:2002:MEA

[HL02] Henrik Hult and Filip Lindskog. Multivariate extremes, aggregation and dependence in elliptical distributions. Advances in Applied Probability, 34(3):587–608, September 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428248.

Heyde:2005:SP

[HL05] C. C. Heyde and N. N. Leonenko. Student processes. Advances in Applied Probability, 37(2):342–365, June 2005. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/30037331.

Hult:2006:RVI

[HL06] Henrik Hult and Filip Lindskog. On regular variation for infinitely divisible random vectors and additive processes. Advances in Applied Probability, 38(1):134–148, March 2006. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443432.

Herve:2014:SAM

[HL14] Loïc Hervé and James Ledoux. Spectral analysis of Markov kernels and application to the convergence rate of discrete random walks. Advances in Applied Probability, 46(4):1036–1058, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1418396242.

Hodara:2017:HPV

[HL17] Pierre Hodara and Eva Löcherbach. Hawkes processes with variable length memory and an infinite number of components. Advances in Applied Probability, 49(1):84–107, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/hawkes-processes-
with-variable-length-memory-and-an-infinite-number-of-
components/AD20DE078C062731C5F56D657A139E8B.

Huang:2018:RSS

[HLG18] Yonghui Huang, Zhaotong Lian, and Xianping Guo. Risksensitive semi-Markov decision processes with general utilities and multiple criteria. Advances in Applied Probability, 50(3):783–804, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/risksensitive-semimarkov-decision-
processes-with-general-utilities-and-multiple-criteria/
EC574AF62541158A57864D84F2A1C281.

Heidergott:2016:PAI

[HLLP16] Bernd Heidergott, Haralambie Leahu, Andreas Löpker, and Georg Pflug. Perturbation analysis of inhomogeneous finite Markov chains. Advances in Applied Probability, 48(1):255–273, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1457466165.

Hautphenne:2013:EPB

[HLN13] S. Hautphenne, G. Latouche, and G. Nguyen. Extinction probabilities of branching processes with countably infinitely many types. Advances in Applied Probability, 45(4):1068–1082, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1386857858.

Hug:2014:SPM

[HLPW14] Daniel Hug, Günter Last, Zbyn¯ök Pawlas, and Wolfgang Weil. Statistics for Poisson models of overlapping spheres. Advances in Applied Probability, 46(4):937–962, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1418396238.

Hug:2002:GCD

[HLW02] Daniel Hug, Günter Last, and Wolfgang Weil. Generalized contact distributions of inhomogeneous Boolean models. Advances in Applied Probability, 34(1):21–47, March 2002. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428370.

Hug:2003:DMP

[HLW03] Daniel Hug, Günter Last, and Wolfgang Weil. Distance measurements on processes of flats. Advances in Applied Probability, 35(1):70–95, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428274.

He:2018:SHS

[HLZ18] Hui He, Jingning Liu, and Mei Zhang. On Seneta–Heyde scaling for a stable branching random walk. Advances in Applied Probability, 50(2):565–599, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-senetaheyde-scaling-for-a-stable-
branching-random-walk/53212E3FF7F5C36BE2048F875F717017.

Hines:1981:ASR

[HM81] W. G. S. Hines and W. S. Moore. An analysis of sex in random environments, I. Advances in Applied Probability, 13(3): 453–463, September 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426780.

Heinrich:1996:CLT

[HM96] Lothar Heinrich and Ilya S. Molchanov. Central limit theorem for a class of random measures associated with germ-grain models. Advances in Applied Probability, 28(2):333–334, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428043.

Heinrich:1999:CLT

[HM99] Lothar Heinrich and Ilya S. Molchanov. Central limit theorem for a class of random measures associated with germ-grain models. Advances in Applied Probability, 31(2):283–314, June 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428111.

Hryniv:2010:LTB

[HM10] O. Hryniv and M. Menshikov. Long-time behaviour in a model of microtubule growth. Advances in Applied Probability, 42(1): 268–291, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25683816.

Huillet:2011:DID

[HM11] Thierry Huillet and Servet Martinez. Duality and intertwining for discrete Markov kernels: Relations and examples. Advances in Applied Probability, 43(2):437–460, June 2011. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/23024514.

Hazra:2012:TBR

[HM12] Rajat Subhra Hazra and Krishanu Maulik. Tail behavior of randomly weighted sums. Advances in Applied Probability, 44 (3):794–814, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/41714076.

Hazra:2020:RAC

[HM20] Nil Kamal Hazra and Neeraj Misra. On relative ageing of coherent systems with dependent identically distributed components. Advances in Applied Probability, 52(1):348–376, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/on-relative-
ageing-of-coherent-systems-with-dependent-identically-
distributed-components/74F3052700319C1FBF22243221930E5A.

Hauser:2006:LDB

[HMM06] Raphael Hauser, Servet Martínez, and Heinrich Matzinger. Large deviations-based upper bounds on the expected relative length of longest common subsequences. Advances in Applied Probability, 38(3):827–852, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443473.

Hamadene:2019:PVT

[HMM19] S. Hamadène, R. Martyr, and J. Moriarty. A probabilistic verification theorem for the finite horizon two-player zero-sum optimal switching game in continuous time. Advances in Applied Probability, 51(2):425–442, June 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/probabilistic-verification-theorem-
for-the-finite-horizon-twoplayer-zerosum-optimal-switching-
game-in-continuous-time/F8AA8650216580262E55A58FB926FFE3.

Hart:2003:CCT

[HMS03] Andrew G. Hart, Servet Martínez, and Jaime San Martín. The λ-classification of continuous-time birth-and-death processes. Advances in Applied Probability, 35(4):1111–1130, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428502.

Hart:2006:SMM

[HMV06] Andrew G. Hart, Servet Martínez, and Leonardo Videla. A simple maximization model inspired by algorithms for the organization of genetic candidates in bacterial DNA. Advances in Applied Probability, 38(4):1071–1097, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443488.

Hirsch:2015:FPP

[HNGS15] C. Hirsch, D. Neuhäuser, C. Gloaguen, and V. Schmidt. First passage percolation on random geometric graphs and an application to shortest-path trees. Advances in Applied Probability, 47(2):328–354, June 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1435236978.

Hirsch:2013:CRG

[HNS13] C. Hirsch, D. Neuhäuser, and V. Schmidt. Connectivity of random geometric graphs related to minimal spanning forests. Advances in Applied Probability, 45(1):20–36, March 2013. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/41805998.

Henderson:1994:SDS

[HNT94] W. Henderson, B. S. Northcote, and P. G. Taylor. Statedependent signalling in queueing networks. Advances in Applied Probability, 26(2):436–455, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427445.

Hummel:1998:FTR

[HO98] Hans-Karl Hummel and Joachim Ohser. Flow through a random tessellation. Advances in Applied Probability, 30(2):280– 281, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427950.

Harutyunyan:2009:CLD

[HO09] H. S. Harutyunyan and V. K. Ohanyan. The chord length distribution function for regular polygons. Advances in Applied Probability, 41(2):358–366, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/27793882.

Hoffmann:2007:IDN

[Hof07a] Lars Michael Hoffmann. Intersection densities of nonstationary Poisson processes of hypersurfaces. Advances in Applied Probability, 39(2):307–317, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443514.

Hoffmann:2007:WSW

[Hof07b] Lars Michael Hoffmann. On weak stationarity and weak isotropy of processes of convex bodies and cylinders. Advances in Applied Probability, 39(4):864–882, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443548.

Hognas:1976:RWC

[Hög76] Göran Högnäs. Random walks on compact semigroups. Advances in Applied Probability, 8(2):231–232, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425879.

Hogan:1986:CBD

[Hog86] Michael L. Hogan. Comment on “Corrected Diffusion Approximations in Certain Random Walk Problems”. Journal of Applied Probability, 23(1):89–96, March 1986. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://

www.jstor.org/stable/3214118. See [Sie79].

Hokstad:1979:SSS

[Hok79] Per Hokstad. On the steady-state solution of the M/G/2 queue. Advances in Applied Probability, 11(1):240–255, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426776.

Hokstad:1980:SSS

[Hok80] Per Hokstad. The steady-state solution of the M/K_{2} /m queue. Advances in Applied Probability, 12(3):799–823, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426432.

Hokstad:1985:RWG

[Hok85] Per Hokstad. Relations for the workload of the GI/G/s queue. Advances in Applied Probability, 17(4):887–904, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427092.

Holgate:1976:TNH

[Hol76a] P. Holgate. The total number of heterozygotes before fixation. Advances in Applied Probability, 8(1):15–16, March 1976. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426014.

Holte:1976:GGC

[Hol76b] John M. Holte. A generalization of Goldstein’s comparison lemma and the exponential limit law in critical Crump–Mode– Jagers branching processes. Advances in Applied Probability, 8 (1):88–104, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426023.

Holgate:1978:CMG

[Hol78] P. Holgate. Canonical multiplication in the genetic algebra for linked loci. Advances in Applied Probability, 10(1):10–11, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426711.

Holgate:1980:IGA

[Hol80] Philip Holgate. Inbreeding in genetic algebras. Advances in Applied Probability, 12(1):8, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426485.

Holgate:1981:RFF

[Hol81] P. Holgate. Rates of first fixation in finite multilocus populations. Advances in Applied Probability, 13(1):4–5, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426456.

Holley:1983:TTM

[Hol83] Richard Holley. Two types of mutually annihilating particles. Advances in Applied Probability, 15(1):133–148, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426986.

Holgate:1989:SPS

[Hol89] Philip Holgate. Some power series with random gaps. Advances in Applied Probability, 21(3):708–710, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427645.

Holgate:1990:LPR

[Hol90] Philip Holgate. Lattice points in a random parallelogram. Advances in Applied Probability, 22(2):484–485, June 1990. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427547.

Holmgren:2011:WSD

[Hol11] Cecilia Holmgren. A weakly 1-stable distribution for the number of random records and cuttings in split trees. Advances in Applied Probability, 43(1):151–177, March 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/23024530.

Hoogwerf:1978:MST

[Hoo78] J. Hoogwerf. Mean sojourn time of packets in a computer communications network. Advances in Applied Probability, 10 (2):336–337, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426936.

Hooghiemstra:1989:AVT

[Hoo89] G. Hooghiemstra. Another visit to two halflines. Advances in Applied Probability, 21(4):935–937, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427777.

Hoppe:1985:NMB

[Hop85] Fred M. Hoppe. A note on Markov branching processes. Advances in Applied Probability, 17(2):463–464, June 1985. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427152.

Horowitz:1977:OTO

[Hor77] J. Horowitz. Occupation times and other accumulation processes. Advances in Applied Probability, 9(2):196–197, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426345.

Horalek:1988:NTN

[Hor88] Vratislav Horálek. A note on the time-non-homogeneous Johnson–Mehl tessellation. Advances in Applied Probability, 20 (3):684–685, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427043.

Horvath:1991:WCD

[Hor91] Lajos Horváth. Weak convergence of discrete scattering processes. Advances in Applied Probability, 23(4):733–750, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427673.

Horst:2002:ALI

[Hor02] Ulrich Horst. Asymptotics of locally-interacting Markov chains with global signals. Advances in Applied Probability, 34(2):416– 440, June 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428295.

Horst:2003:SLS

[Hor03] Ulrich Horst. Stability of linear stochastic difference equations in strategically controlled random environments. Advances in Applied Probability, 35(4):961–981, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428495.

Hossjer:2011:CTG

[Hös11] O. Hössjer. Coalescence theory for a general class of structured populations with fast migration. Advances in Applied Probability, 43(4):1027–1047, December 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/23208037.

Houdebert:2018:PRC

[Hou18] Pierre Houdebert. Percolation results for the continuum random cluster model. Advances in Applied Probability, 50(1):231–244, March 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/percolation-results-for-the-continuum-
random-cluster-model/B1C275EAA2C76064BBED49020D8BD9C5.

Hasofer:1978:EOP

[HP78] A. M. Hasofer and P. Petocz. The envelope of an oscillatory process and its upcrossings. Advances in Applied Probability, 10 (4):711–716, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426649.

Harlow:1982:PDS

[HP82] D. Gary Harlow and S. Leigh Phoenix. Probability distributions for the strength of fibrous materials under local load sharing i: Two-level failure and edge effects. Advances in Applied Probability, 14(1):68–94, March 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426734.

Hovey:1983:ISB

[HP83] Peter W. Hovey and Peter Purdue. An inventory system with both delayed and immediate delivery. Advances in Applied Probability, 15(3):691–694, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1426628.

Huang:1991:QPP

[HP91] Wen-Jang Huang and Prem S. Puri. A queueing process with the possibility of customers becoming servers. Advances in Applied Probability, 23(4):957–971, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427686.

Harrison:1996:MGQ

[HP96] Peter G. Harrison and Edwige Pitel. The M/G/1 queue with negative customers. Advances in Applied Probability, 28(2):540– 566, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428071.

Heidergott:2023:GES

[HP23] Bernd Heidergott and Yijie Peng. Gradient estimation for smooth stopping criteria. Advances in Applied Probability, 55(1):29–55, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/gradient-estimation-for-smooth-stopping-
criteria/77A11AA614BD1B9AB8593B411E606C70.

Hellmund:2008:LBC

[HPJ08] Gunnar Hellmund, Michaela Prokesová, and Eva B. Vedel Jensen. Lévy-based Cox point processes. Advances in Applied Probability, 40(3):603–629, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443600.

Henderson:1992:CIB

[HPPT92] W. Henderson, C. E. M. Pearce, P. K. Pollett, and P. G. Taylor. Connecting internally balanced quasi-reversible Markov processes. Advances in Applied Probability, 24(4):934–959, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427720.

Hayen:2000:PTP

[HQ00] Andrew Hayen and Malcolm Quine. The proportion of triangles in a Poisson–Voronoi tessellation of the plane. Advances in Applied Probability, 32(1):67–74, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428226.

Hayen:2002:ACV

[HQ02] A. Hayen and M. P. Quine. Areas of components of a Voronoi polygon in a homogeneous Poisson process in the plane. Advances in Applied Probability, 34(2):281–291, June 2002. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428288.

Harrison:1978:SDR

[HR78] J. Michael Harrison and Sidney I. Resnick. The stationary distribution and recurrence classification of a storage process with general release rule. Advances in Applied Probability, 10(2): 321, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426921.

Hanqin:1989:HTL

[HR89] Zhang Hanqin and Wang Rongxin. Heavy traffic limit theorems for a queueing system in which customers join the shortest line. Advances in Applied Probability, 21(2):451–469, June 1989. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427169.

Hutchinson:2000:RFP

[HR00] John E. Hutchinson and Ludger Rüschendorf. Random fractals and probability metrics. Advances in Applied Probability, 32(4): 925–947, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428508.

Hansen:2004:EWP

[HR04] Jan Hansen and Matthias Reitzner. Electromagnetic wave propagation and inequalities for moments of chord lengths. Advances in Applied Probability, 36(4):987–995, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/4140386.

Heffernan:2005:HRV

[HR05a] Janet Heffernan and Sidney Resnick. Hidden regular variation and the rank transform. Advances in Applied Probability, 37 (2):393–414, June 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/30037333.

Hug:2005:GPV

[HR05b] Daniel Hug and Matthias Reitzner. Gaussian polytopes: Variances and limit theorems. Advances in Applied Probability, 37 (2):297–320, June 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/30037329.

Hurvich:2016:SEA

[HR16] Clifford Hurvich and Josh Reed. Series expansions for the all-time maximum of α-stable random walks. Advances in Applied Probability, 48(3):744–767, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1474296313.

Hug:2004:LPV

[HRS04] Daniel Hug, Matthias Reitzner, and Rolf Schneider. Large Poisson–Voronoi cells and Crofton cells. Advances in Applied Probability, 36(3):667–690, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/4140402.

Hanin:1997:SMC

[HRTY97] L. G. Hanin, S. T. Rachev, A. D. Tsodikov, and A. Yu. Yakovlev. A stochastic model of carcinogenesis and tumor size at detection. Advances in Applied Probability, 29(3):607–628, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428079.

Hallmann:2018:ASS

[HRW18] S. Hallmann, U. Rösler, and M. Wnuk. All solutions of the stochastic fixed point equation of the Quicksort process. Advances in Applied Probability, 50(A):131–140, December 2018.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/all-solutions-
of-the-stochastic-fixed-point-equation-of-the-quicksort-
process/795C0670996532A7E615904BFF214364.

Hanin:1993:OCC

[HRY93] L. G. Hanin, S. T. Rachev, and A. Yu. Yakovlev. On the optimal control of cancer radiotherapy for non-homogeneous cell populations. Advances in Applied Probability, 25(1):1–23, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427493.

Heath:1974:CTG

[HS74] David C. Heath and William D. Sudderth. Continuous-time gambling problems. Advances in Applied Probability, 6(4):651– 665, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426185.

Horn:1977:SRI

[HS77] R. A. Horn and F. W. Steutel. Some results for infinitely divisible distributions in n variables. Advances in Applied Probability, 9(2):227, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426371.

Hochbaum:1982:SGL

[HS82] Dorit Hochbaum and J. Michael Steele. Steinhaus’s geometric location problem for random samples in the plane. Advances in Applied Probability, 14(1):56–67, March 1982. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426733.

Heinrich:1985:NCM

[HS85] Lothar Heinrich and Volker Schmidt. Normal convergence of multidimensional shot noise and rates of this convergence. Advances in Applied Probability, 17(4):709–730, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427084.

Hordijk:1989:CAC

[HS89] Arie Hordijk and Flos Spieksma. Constrained admission control to a queueing system. Advances in Applied Probability, 21 (2):409–431, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427167.

Hoppe:1990:BTI

[HS90] Fred M. Hoppe and Eugene Seneta. Bonferroni-type inequalities and the methods of indicators and polynomials. Advances in Applied Probability, 22(1):241–246, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427607.

Haario:1991:SAP

[HS91] Heikki Haario and Eero Saksman. Simulated annealing process in general state space. Advances in Applied Probability, 23(4): 866–893, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427681.

Hordijk:1992:ERP

[HS92] Arie Hordijk and Flora Spieksma. On ergodicity and recurrence properties of a Markov chain with an application to an open Jackson network. Advances in Applied Probability, 24(2):343– 376, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427696.

Hartfiel:1994:TMS

[HS94] D. J. Hartfiel and E. Seneta. On the theory of Markov set-chains. Advances in Applied Probability, 26(4):947–964, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427899.

Hahn:1998:USE

[HS98] Ute Hahn and Dietrich Stoyan. Unbiased stereological estimation of the surface area of gradient surface processes. Advances in Applied Probability, 30(4):904–920, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428021.

Helmes:2010:CVF

[HS10] Kurt Helmes and Richard H. Stockbridge. Construction of the value function and optimal rules in optimal stopping of onedimensional diffusions. Advances in Applied Probability, 42(1): 158–182, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25683811.

Hug:2011:FGD

[HS11] Daniel Hug and Rolf Schneider. Faces with given directions in anisotropic Poisson hyperplane mosaics. Advances in Applied Probability, 43(2):308–321, June 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/23024507.

Heinrich:2013:CLT

[HS13] Lothar Heinrich and Malte Spiess. Central limit theorems for volume and surface content of stationary Poisson cylinder processes in expanding domains. Advances in Applied Probability, 45(2):312–331, June 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1370870120.

Hug:2014:APR

[HS14] Daniel Hug and Rolf Schneider. Approximation properties of random polytopes associated with Poisson hyperplane processes. Advances in Applied Probability, 46(4):919–936, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1418396237.

Hata:2018:OCI

[HS18] Hiroaki Hata and Shuenn-Jyi Sheu. An optimal consumption and investment problem with partial information. Advances in Applied Probability, 50(1):131–153, March 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/an-
optimal-consumption-and-investment-problem-with-partial-
information/12762BB17077F2124F8B9D9A3E2763EA.

Hsing:1987:ICS

[Hsi87] Tailen Hsing. On the intensity of crossings by a shot noise process. Advances in Applied Probability, 19(3):743–745, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427416.

See comment [Kon89b].

Hsiau:2010:SLC

[Hsi10] Shoou-Ren Hsiau. Selecting the last consecutive record in a record process. Advances in Applied Probability, 42(3):739–760, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25799066.

Heinrich:2005:LTS

[HSS05] Lothar Heinrich, Hendrik Schmidt, and Volker Schmidt. Limit theorems for stationary tessellations with random inner cell structures. Advances in Applied Probability, 37(1):25–47, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/30037314.

Hansen:2008:CRP

[HSS08] Jennie C. Hansen, Eric Schmutz, and Li Sheng. Covering random points in a unit disk. Advances in Applied Probability, 40 (1):22–30, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443568.

Helmes:2018:WCA

[HSZ18] K. L. Helmes, R. H. Stockbridge, and C. Zhu. A weak convergence approach to inventory control using a long-term average criterion. Advances in Applied Probability, 50(4): 1032–1074, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/weak-convergence-approach-to-inventory-
control-using-a-longterm-average-criterion/DDB4F4D8BF2EF6A9F241EE185F744A69.

Hsing:1989:EPS

[HT89] Tailen Hsing and J. L. Teugels. Extremal properties of shot noise processes. Advances in Applied Probability, 21(3):513–525, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427633.

Henderson:1992:DTQ

[HT92] W. Henderson and P. G. Taylor. Discrete-time queueing networks with geometric release probabilities. Advances in Applied Probability, 24(1):229–233, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427741.

Hutzenthaler:2010:TRS

[HT10] Martin Hutzenthaler and Jesse Earl Taylor. Time reversal of some stationary jump diffusion processes from population genetics. Advances in Applied Probability, 42(4):1147–1171, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25799088.

Hofer-Temmel:2017:SPP

[HT17] Christoph Hofer-Temmel. Shearer’s point process, the hardsphere model, and a continuum Lovász local lemma. Advances in Applied Probability, 49(1):1–23, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/shearers-point-process-the-
hardsphere-model-and-a-continuum-lovasz-local-lemma/
11CE8EB91F2EF2090FC7337F5DED0FA5.

Hansen:2015:GRP

[HTO^{+}15] Linda V. Hansen, Thordis L. Thorarinsdottir, Evgeni Ovcharov, Tilmann Gneiting, and Donald Richards. Gaussian random particles with flexible Hausdorff dimension. Advances in Applied Probability, 47(2):307–327, June 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1435236977.

Huang:1990:MEP

[Hua90] Dawei Huang. On the maximal entropy property for ARMA processes and ARMA approximation. Advances in Applied Probability, 22(3):612–626, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427460.

Huang:1997:SFM

[Hua97] D. Huang. Stochastic FM models and non-linear time series analysis. Advances in Applied Probability, 29(4):986–1003, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427850.

Huang:2002:EBM

[Hua02] Haiyan Huang. Error bounds on multivariate normal approximations for word count statistics. Advances in Applied Probability, 34(3):559–586, September 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428247.

Hubner:1978:STN

[Hüb78] Gerhard Hübner. Some tests for non-optimal actions in Markov decision problems with absorption. Advances in Applied Probability, 10(2):309–310, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426910.

Hubner:1980:BGP

[Hüb80] Gerhard Hübner. Bounds and good policies in stationary finitestage Markovian decision problems. Advances in Applied Probability, 12(1):154–173, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426499.

Hueter:1994:CHN

[Hue94] Irene Hueter. The convex Hull of a normal sample. Advances in Applied Probability, 26(4):855–875, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427894.

Hueter:1999:CHS

[Hue99] Irene Hueter. The convex Hull of samples from self-similar distributions. Advances in Applied Probability, 31(1):34–47, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427989.

Hueter:2016:BPG

[Hue16] Irene Hueter. Branching processes in generalized autoregressive conditional environments. Advances in Applied Probability, 48(4):1211–1234, December 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1482548435.

Huillet:2003:ECB

[Hui03a] Thierry Huillet. Energy cascades as branching processes with emphasis on Neveu’s approach to Derrida’s random energy model. Advances in Applied Probability, 35(2):477–503, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428432.

Huillet:2003:RCC

[Hui03b] Thierry Huillet. Random covering of the circle: The size of the connected components. Advances in Applied Probability, 35(3): 563–582, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428208.

Huillet:2004:DPC

[Hui04] Thierry Huillet. On a deposition process on the circle with disorder. Advances in Applied Probability, 36(4):996–1020, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/4140387.

Hunter:1969:MMR

[Hun69] Jeffrey J. Hunter. On the moments of Markov renewal processes. Advances in Applied Probability, 1(2):188–210, Autumn 1969. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426217.

Hunter:1974:RTTa

[Hun74a] J. J. Hunter. Renewal theory in two dimensions. Advances in Applied Probability, 6(2):220–221, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426259.

Hunter:1974:RTTc

[Hun74b] Jeffrey J. Hunter. Renewal theory in two dimensions: Asymptotic results. Advances in Applied Probability, 6(3):546–562, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426233.

Hunter:1974:RTTb

[Hun74c] Jeffrey J. Hunter. Renewal theory in two dimensions: Basic results. Advances in Applied Probability, 6(2):376–391, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426299.

Hunter:1977:RTT

[Hun77] Jeffrey J. Hunter. Renewal theory in two dimensions: Bounds on the renewal function. Advances in Applied Probability, 9(3): 527–541, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426112.

Hunter:1979:QSS

[Hun79] Jeffrey J. Hunter. Queueing and storage systems in parallel. Advances in Applied Probability, 11(2):305–306, June 1979. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426837.

Hunter:1983:FMRa

[Hun83a] Jeffrey J. Hunter. Filtering of Markov renewal queues, i: Feedback queues. Advances in Applied Probability, 15(2):349– 375, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426440.

Hunter:1983:FMRb

[Hun83b] Jeffrey J. Hunter. Filtering of Markov renewal queues, II: Birth–death queues. Advances in Applied Probability, 15(2):376– 391, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426441.

Hunter:1984:FMR

[Hun84] Jeffrey J. Hunter. Filtering of Markov renewal queues, III: Semi-Markov processes embedded in feedback queues. Advances in Applied Probability, 16(2):422–436, June 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427077.

Hunter:1985:FMR

[Hun85] Jeffrey J. Hunter. Filtering of Markov renewal queues, IV: Flow processes in feedback queues. Advances in Applied Probability, 17(2):386–407, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427147.

Hunt:1989:ICL

[Hun89] P. J. Hunt. Implied costs in loss networks. Advances in Applied Probability, 21(3):661–680, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427641.

Hunt:1995:LNU

[Hun95] P. J. Hunt. Loss networks under diverse routing: The symmetric star network. Advances in Applied Probability, 27(1):255– 272, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428106.

Husler:1989:NIT

[Hüs89] J. Hüsler. A note on the independence and total dependence of max I.D. distributions. Advances in Applied Probability, 21 (1):231–232, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427207.

Hutton:1980:RTT

[Hut80] J. Hutton. The recurrence and transience of two-dimensional linear birth and death processes. Advances in Applied Probability, 12(3):615–639, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426423.

Hordijk:1983:AOP

[HV83] Arie Hordijk and Frank A. Van Der Duyn Schouten. Average optimal policies in Markov decision drift processes with applications to a queueing and a replacement model. Advances in Applied Probability, 15(2):274–303, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426437.

Haviv:1984:PBS

[HV84] Moshe Haviv and Ludo Van Der Heyden. Perturbation bounds for the stationary probabilities of a finite Markov chain. Advances in Applied Probability, 16(4):804–818, December 1984.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427341.

Hansen:2001:NEC

[HV01] Martin B. Hansen and Erik W. Van Zwet. Nonparametric estimation of the chord length distribution. Advances in Applied Probability, 33(3):541–558, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428473.

Heubach:1995:SMM

[HW95] Silvia Heubach and Joseph Watkins. A stochastic model for the movement of a white blood cell. Advances in Applied Probability, 27(2):443–475, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427835.

Heyde:2009:FTR

[HW09a] C. C. Heyde and Dingcheng Wang. Finite-time ruin probability with an exponential Lévy process investment return and heavy-tailed claims. Advances in Applied Probability, 41(1):206– 224, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
27793870.

Huber:2009:LBI

[HW09b] Mark L. Huber and Robert L. Wolpert. Likelihood-based inference for Matérn Type-III repulsive point processes. Advances in Applied Probability, 41(4):958–977, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/27793913.

Hug:2019:DBM

[HW19] Daniel Hug and Wolfgang Weil. Determination of Boolean models by densities of mixed volumes. Advances in Applied Probability, 51(1):116–135, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/determination-of-boolean-models-
by-densities-of-mixed-volumes/EB8CAD1D14CBE0CB1B0C39999BE65DAB.

Hwang:1999:APA

[Hwa99] Hsien-Kuei Hwang. Asymptotics of Poisson approximation to random discrete distributions: An analytic approach. Advances in Applied Probability, 31(2):448–491, June 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428118.

Huckemann:2006:PCA

[HZ06] Stephan Huckemann and Herbert Ziezold. Principal component analysis for Riemannian manifolds, with an application to triangular shape spaces. Advances in Applied Probability, 38 (2):299–319, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443444.

He:2007:MED

[HZ07] Qi-Ming He and Hanqin Zhang. On matrix exponential distributions. Advances in Applied Probability, 39(1):271–292, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/20443509.

Iglehart:1971:FLT

[Igl71] Donald L. Iglehart. Functional limit theorems for the queue GI/G/1 in light traffic. Advances in Applied Probability, 3(2): 269–281, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426171.

Iglehart:1973:WCA

[Igl73a] Donald L. Iglehart. Weak convergence in applied probability. Advances in Applied Probability, 5(1):12–13, April 1973. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425948.

Iglehart:1973:WCQ

[Igl73b] Donald L. Iglehart. Weak convergence in queueing theory. Advances in Applied Probability, 5(3):570–594, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425835.

Iglehart:1975:CLT

[Igl75] Donald L. Iglehart. Conditioned limit theorems for random walks. Advances in Applied Probability, 7(2):237, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426050.

Iglehart:1980:ESS

[Igl80] Donald L. Iglehart. Exploiting stochastic structure in the simulation of networks of queues. Advances in Applied Probability, 12 (2):271, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426538.

Iksanov:2002:FPP

[IJ02] Aleksander M. Iksanov and Zbigniew J. Jurek. On fixed points of Poisson shot noise transforms. Advances in Applied Probability, 34(4):798–825, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428323.

Iglehart:1973:ARP

[IL73] Donald L. Iglehart and Austin J. Lemoine. Approximations for the repairman problem with two repair facilities, i: No spares. Advances in Applied Probability, 5(3):595–613, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425836.

Iglehart:1974:ARP

[IL74] Donald L. Iglehart and Austin J. Lemoine. Approximations for the repairman problem with two repair facilities, II: Spares. Advances in Applied Probability, 6(1):147–158, March 1974. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426211.

Istas:1997:EFS

[IL97] Jacques Istas and Catherine Laredo. Estimating functionals of a stochastic process. Advances in Applied Probability, 29(1):249– 270, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427869.

Illsley:1998:MNE

[Ill98] Robert Illsley. The moments of the number of exits from a simply connected region. Advances in Applied Probability, 30 (1):167–180, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427882.

Illsley:2001:ESG

[Ill01] Robert Illsley. The excursions of a stationary Gaussian process outside a large two-dimensional region. Advances in Applied Probability, 33(1):141–159, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428445.

Iyer:2018:CUS

[ILP18] Gautam Iyer, Nicholas Leger, and Robert L. Pego. Coagulation and universal scaling limits for critical Galton–Watson processes. Advances in Applied Probability, 50(2):504–542, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/coagulation-
and-universal-scaling-limits-for-critical-galtonwatson-
processes/AA83743F3EE751AAE422A963B37E088F.

Ishii:2002:TDR

[IM02] H. Ishii and T. Mikami. A two-dimensional random crystalline algorithm for Gauss curvature flow. Advances in Applied Probability, 34(3):491–504, September 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1428243.

Iksanov:2008:NJR

[IM08] Alex Iksanov and Martin Möhle. On the number of jumps of random walks with a barrier. Advances in Applied Probability, 40(1):206–228, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/20443577.

Imhof:1986:TSA

[Imh86] J.-P. Imhof. On the time spent above a level by Brownian motion with negative drift. Advances in Applied Probability, 18(4):1017– 1018, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427261.

Imhof:1992:SPP

[Imh92] J. P. Imhof. A simple proof of Pitman’s 2M-X theorem. Advances in Applied Probability, 24(2):499–501, June 1992. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427703.

Imhof:2008:MTC

[Imh08] Lorens A. Imhof. Multiple-trial conflicts and stochastic evolutionary game dynamics. Advances in Applied Probability, 40(4): 1174–1197, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/20443629.

Ipsen:2021:GDD

[IMS21] Yuguang Ipsen, Ross A. Maller, and Soudabeh Shemehsavar. A generalised Dickman distribution and the number of species in a negative binomial process model. Advances in Applied Probability, 53(2):370–399, June 2021. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/generalised-
dickman-distribution-and-the-number-of-species-in-a-
negative-binomial-process-model/BC4847DD814F3507D6636FB6EEE89610.

Ispany:2005:FLB

[IPV05] M. Ispány, G. Pap, and M. C. A. Van Zuijlen. Fluctuation limit of branching processes with immigration and estimation of the means. Advances in Applied Probability, 37(2):523–538, June 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/30037339.

Iglehart:1983:SER

[IS83] Donald L. Iglehart and Gerald S. Shedler. Statistical efficiency of regenerative simulation methods for networks of queues. Advances in Applied Probability, 15(1):183–197, March 1983. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426989.

Irle:2006:OPP

[IS06] Albrecht Irle and Jörn Sass. Optimal portfolio policies under fixed and proportional transaction costs. Advances in Applied Probability, 38(4):916–942, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443482.

Isokawa:2000:PVT

[Iso00] Yukinao Isokawa. Poisson–Voronoi tessellations in threedimensional hyperbolic spaces. Advances in Applied Probability, 32(3):648–662, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428405.

Ito:1979:RPD

[Ito79] Yoshifusa Ito. Renewal processes decomposable into i.i.d. components. Advances in Applied Probability, 11(2):273, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426802.

Ito:1980:RPD

[Ito80] Yoshifusa Ito. Renewal processes decomposable into i.i.d. components. Advances in Applied Probability, 12(3):672–688, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426426.

Ivanoff:1980:BRF

[Iva80] Gail Ivanoff. The branching random field. Advances in Applied Probability, 12(4):825–847, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426743.

Ivanoff:1983:CBD

[Iva83] B. Gail Ivanoff. The critical branching diffusion in a varying environment. Advances in Applied Probability, 15(3):488–506, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426616.

Ivanoff:1989:MBR

[Iva89] B. Gail Ivanoff. The multitype branching random walk: Temporal and spatial limit theorems. Advances in Applied Probability, 21(3):491–512, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427632.

Iglehart:1970:MCQa

[IW70a] Donald L. Iglehart and Ward Whitt. Multiple channel queues in heavy traffic. I. Advances in Applied Probability, 2(1):150– 177, Spring 1970. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
3518347.

Iglehart:1970:MCQb

[IW70b] Donald L. Iglehart and Ward Whitt. Multiple channel queues in heavy traffic. II: Sequences, networks, and batches. Advances in Applied Probability, 2(2):355–369, Autumn 1970. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426324.

Iyer:2012:PCR

[IY12] Srikanth K. Iyer and D. Yogeshwaran. Percolation and connectivity in “Ab” random geometric graphs. Advances in Applied Probability, 44(1):21–41, March 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/23207854.

Iyer:1988:APM

[Iye88] Srinivas Iyer. Asymptotic performance of a multistate coherent system. Advances in Applied Probability, 20(1):241–243, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427281.

Iyer:2023:DDR

[Iye23] Tejas Iyer. Degree distributions in recursive trees with fitnesses. Advances in Applied Probability, 55(2):407–443, June 2023. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/degree-distributions-
in-recursive-trees-with-fitnesses/E03B3D2F7109958ABC8FFFDCCFC94657.

Jacobs:1974:FIP

[Jac74a] P. A. Jacobs. Functionals of an infinite particle system with independent motions, creation, and Annihilation. Advances in Applied Probability, 6(4):636–650, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426184.

Jacobsen:1974:DRD

[Jac74b] M. Jacobsen. A decomposition result for diffusions. Advances in Applied Probability, 6(2):222–223, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426260.

Jacobs:1980:HTR

[Jac80a] Patricia A. Jacobs. Heavy traffic results for single-server queues with dependent (EARMA) service and interarrival times. Advances in Applied Probability, 12(2):517–529, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426610.

Jacod:1980:CMP

[Jac80b] Jean Jacod. On the construction of a Markov process up to a change of time. Advances in Applied Probability, 12(2):271, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426539.

Jacobsen:2005:TRC

[Jac05] Martin Jacobsen. The time to ruin for a class of Markov additive risk process with two-sided jumps. Advances in Applied Probability, 37(4):963–992, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/30037369.

Jagers:1977:BPC

[Jag77] Peter Jagers. Branching processes and cell proliferation: A survey. Advances in Applied Probability, 9(2):197, June 1977. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426346.

Jagers:2018:APT

[Jag18] Peter Jagers. Applied probability as theoretical science: 50 years in the applied probability community. Advances in Applied Probability, 50(A):1–2, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/applied-probability-as-theoretical-
science-50-years-in-the-applied-probability-community/
69D095889EFAAB09235A2CB5E5F22F32.

Jansen:1984:CES

[Jan84] Uwe Jansen. Conditional expected sojourn times in insensitive queueing systems and networks. Advances in Applied Probability, 16(4):906–919, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427346.

Janson:1986:MFP

[Jan86] Svante Janson. Moments for first-passage and last-exit times, the minimum, and related quantities for random walks with positive drift. Advances in Applied Probability, 18(4):865–879, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427253.

Janson:2018:AFC

[Jan18] Svante Janson. Asymptotics of fluctuations in Crump-Mode–Jagers processes: the lattice case. Advances in Applied Probability, 50(A):141–171, December 2018. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/asymptotics-
of-fluctuations-in-crumpmodejagers-processes-the-lattice-
case/EAE4347F4D6BD0FEB12FC95E67AB3414.

Jansen:2019:CEG

[Jan19] S. Jansen. Cluster expansions for Gibbs point processes. Advances in Applied Probability, 51(4):1129–1178, December 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/cluster-expansions-
for-gibbs-point-processes/896D92275D41FDE356016D50529C69CB.

Janson:2020:MVB

[Jan20] Svante Janson. Mean and variance of balanced Pólya urns. Advances in Applied Probability, 52(4):1224–1248, December 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/mean-and-variance-
of-balanced-polya-urns/D102FBDD1A35B2BC27F2BB41207652D8.

Janson:2023:AND

[Jan23] Svante Janson. Asymptotic normality for m -dependent and constrained U -statistics, with applications to pattern matching in random strings and permutations. Advances in Applied Probability, 55(3):841–894, September 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/asymptotic-normality-for-boldsymbolmdependent-
and-constrained-boldsymbolustatistics-with-applications-

to-pattern-matching-in-random-strings-and-permutations/
1CF7F29EE91684C5AD9391696C90B973.

Joshi:1989:MAG

[JD89] B. H. Joshi and A. D. Dharmadhikari. Maximization of availability of 1-out-of-2:G repairable dependent system. Advances in Applied Probability, 21(3):717–720, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427648.

Jensen:1984:SRR

[Jen84] U. Jensen. Some remarks on the renewal function of the uniform distribution. Advances in Applied Probability, 16(1):214– 215, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427232.

Jensen:1988:USA

[Jen88] J. L. Jensen. Uniform saddlepoint approximations. Advances in Applied Probability, 20(3):622–634, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427038.

Jensen:1995:RVC

[Jen95] Eva B. Vedel Jensen. Rotational versions of the Crofton formula. Advances in Applied Probability, 27(1):87–96, March 1995. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428097.

Jensen:1996:IPG

[Jen96a] E. B. Vedel Jensen. An induction proof of the generalized Blaschke-petkantschin formula. Advances in Applied Probability, 28(2):334, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428044.

Jensen:1996:I

[Jen96b] E. B. Vedel Jensen. Introduction. Advances in Applied Probability, 28(2):328, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428036.

Jensen:2003:JMA

[Jen03] Eva B. Vedel Jensen. Joseph mecke: An appreciation. Advances in Applied Probability, 35(1):1–3, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428269.

Jeulin:2000:OGM

[Jeu00] Dominique Jeulin. Obituary: Georges Matheron. Advances in Applied Probability, 32(4):1190–1192, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428523.

Jalali:1992:DAO

[JH92a] Assad Jalali and Alan G. Hawkes. The distribution of apparent occupancy times in a two-state Markov process in which brief events cannot be detected. Advances in Applied Probability, 24 (2):288–301, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427693.

Jalali:1992:GEA

[JH92b] Assad Jalali and Alan G. Hawkes. Generalised eigenproblems arising in aggregated Markov processes allowing for time interval omission. Advances in Applied Probability, 24(2):302–321, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427694.

Jirina:1974:BMB

[Jir74] Miloslav Jirina. Breakage mechanisms branching processes. Advances in Applied Probability, 6(2):230, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426268.

Jirina:1976:ENH

[Jir76] Miloslav Jirina. Extinction of non-homogeneous Galton–Watson processes. Advances in Applied Probability, 8(2):233, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425880.

Jirak:2013:LTA

[Jir13] M. Jirak. Limit theorems for aggregated linear processes. Advances in Applied Probability, 45(2):520–544, June 2013. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1370870128.

Jia:2021:DBL

[JJL21] Chen Jia, Da-Quan Jiang, and Youming Li. Detailed balance, local detailed balance, and global potential for stochastic chemical reaction networks. Advances in Applied Probability, 53(3):886–922, September 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/detailed-balance-local-detailed-
balance-and-global-potential-for-stochastic-chemical-
reaction-networks/488E848D9717ADA1B1839E6FDBF78820.

Johri:1985:FIS

[JK85] Pravin K. Johri and Michael N. Katehakis. Further insight into the structure of bold and timid policies. Advances in Applied Probability, 17(2):298–307, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427142.

Jensen:1994:USE

[JK94] E. B. Vedel Jensen and K. Kiêu. Unbiased stereological estimation of d-dimensional volume in R^{n} from an isotropic random Slice through a fixed point. Advances in Applied Probability, 26(1):1–12, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427575.

Jara:2011:LTS

[JK11] M. Jara and T. Komorowski. Limit theorems for some continuous-time random walks. Advances in Applied Probability, 43(3):782–813, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/23024575.

Jansen:2023:DCF

[JKM23] Sabine Jansen, Leonid Kolesnikov, and Kilian Matzke. The direct-connectedness function in the random connection model. Advances in Applied Probability, 55(1):179–222, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/directconnectedness-
function-in-the-random-connection-model/445C583DD281BA84DE5AF3ED245536A1.

Jin:2017:EEA

[JKR17] Peng Jin, Jonas Kremer, and Barbara Rüdiger. Exponential ergodicity of an affine two-factor model based on the α-root process. Advances in Applied Probability, 49 (4):1144–1169, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/exponential-ergodicity-of-an-affine-
twofactor-model-based-on-the-root-process/BE0217DD0683D9332302C67FEAE744CD.

Jagers:2007:MPE

[JKS07] Peter Jagers, Fima C. Klebaner, and Serik Sagitov. Markovian paths to extinction. Advances in Applied Probability, 39(2):569– 587, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443527.

Jagers:1973:TTT

[JL73] Peter Jagers and Torgny Lindvall. Three theorems on the thinning of point processes. Advances in Applied Probability, 5(1): 14–15, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425949.

Jacobs:1977:MAM

[JL77] P. A. Jacobs and P. A. W. Lewis. A mixed autoregressivemoving average exponential sequence and point process (EARMA 1, 1). Advances in Applied Probability, 9(1):87–104, March 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425818.

Jacobs:1979:TSG

[JL79] P. A. Jacobs and P. A. W. Lewis. Time series generated by mixtures: The DARMA(p,q) process. Advances in Applied Probability, 11(2):266–267, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426797.

Jelenkovic:1999:ARM

[JL99] Predrag R. Jelenković and Aurel A. Lazar. Asymptotic results for multiplexing subexponential on-off processes. Advances in Applied Probability, 31(2):394–421, June 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428116.

Jacquier:2015:CFI

[JL15] Antoine Jacquier and Matthew Lorig. From characteristic functions to implied volatility expansions. Advances in Applied Probability, 47(3):837–857, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1444308884.

Jacka:2005:CAFa

[JLW05a] Saul D. Jacka, Zorana Lazic, and Jon Warren. Conditioning an additive functional of a Markov chain to stay nonnegative. I. Survival for a long time. Advances in Applied Probability, 37(4):1015–1034, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/30037371.

Jacka:2005:CAFb

[JLW05b] Saul D. Jacka, Zorana Lazic, and Jon Warren. Conditioning an additive functional of a Markov chain to stay nonnegative. II. Hitting a high level. Advances in Applied Probability, 37(4): 1035–1055, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/30037372.

Jasra:2022:UFC

[JLY22] Ajay Jasra, Kody J. H. Law, and Fangyuan Yu. Unbiased filtering of a class of partially observed diffusions. Advances in Applied Probability, 54(3):661–687, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/unbiased-filtering-of-a-
class-of-partially-observed-diffusions/D2E8BC350674D8658E4096C2C081BEC9.

Joffe:1986:WCS

[JM86] A. Joffe and M. Metivier. Weak convergence of sequences of semimartingales with applications to multitype branching processes. Advances in Applied Probability, 18(1):20–65, March 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427238.

Juneja:2013:OPC

[JM13] S. Juneja and M. Mandjes. Overlap problems on the circle. Advances in Applied Probability, 45(3):773–790, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1377868538.

Junca:2019:ORS

[JMFPY19] Mauricio Junca, Harold A. Moreno-Franco, José Luis Pérez, and Kazutoshi Yamazaki. Optimality of refraction strategies for a constrained dividend problem. Advances in Applied Probability, 51(3):633–666, September 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/optimality-of-refraction-strategies-
for-a-constrained-dividend-problem/FEB87FF207567E92536D542909E89556.

Jiang:2001:ORR

[JMJ01] Xiaoyue Jiang, Viliam Makis, and Andrew K. S. Jardine. Optimal Repair/Replacement policy for a general repair model. Advances in Applied Probability, 33(1):206–222, March 2001. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428448.

Jean-Marie:1992:SCQ

[JML92] Alain Jean-Marie and Zhen Liu. Stochastic comparisons for queueing models via random sums and intervals. Advances in Applied Probability, 24(4):960–985, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427721.

Jagers:1984:GCB

[JN84] Peter Jagers and Olle Nerman. The growth and composition of branching populations. Advances in Applied Probability, 16 (2):221–259, June 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427068.

Jensen:1998:TMP

[JN98] Eva B. Vedel Jensen and Linda Stougaard Nielsen. Transformations of Markov point processes. Advances in Applied Probability, 30(2):281, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427951.

Jelenkovic:2010:IRP

[JOC10] Predrag R. Jelenković and Mariana Olvera-Cravioto. Information ranking and power laws on trees. Advances in Applied Probability, 42(4):1057–1093, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/25799084.

Jelenkovic:2012:IRT

[JOC12] Predrag R. Jelenković and Mariana Olvera-Cravioto. Implicit renewal theory and power tails on trees. Advances in Applied Probability, 44(2):528–561, June 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/41714062.

Joffe:1975:RVI

[Jof75] Antole Joffe. Random variables indexed by a tree. Advances in Applied Probability, 7(2):237–238, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426051.

Joffe:1977:RPL

[Jof77] Anatole Joffe. Remark on a paper by S. Litvin. Advances in Applied Probability, 9(2):219–220, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426363. See [Lit69b].

Jordan:2010:DSG

[Jor10] Jonathan Jordan. Degree sequences of geometric preferential attachment graphs. Advances in Applied Probability, 42(2):319– 330, June 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25683821.

Joyce:1991:EFO

[Joy91] Paul Joyce. Estimating the frequency of the oldest allele: A Bayesian approach. Advances in Applied Probability, 23(3):456– 475, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427617.

Jacob:1998:EPB

[JP98] C. Jacob and J. Peccoud. Estimation of the parameters of a branching process from migrating binomial observations. Advances in Applied Probability, 30(4):948–967, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428024.

Jammalamadaka:2000:PLP

[JP00a] S. Rao Jammalamadaka and Mathew D. Penrose. Poisson limits for pairwise and area interaction point processes. Advances in Applied Probability, 32(1):75–85, March 2000. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428227.

Jensen:2000:PMD

[JP00b] Jens Ledet Jensen and Anne-Mette Krabbe Pedersen. Probabilistic models of DNA sequence evolution with context dependent rates of substitution. Advances in Applied Probability, 32 (2):499–517, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428201.

Jones:2014:CCM

[JRR14] Galin L. Jones, Gareth O. Roberts, and Jeffrey S. Rosenthal. Convergence of conditional Metropolis–Hastings samplers. Advances in Applied Probability, 46(2):422–445, June 2014. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1401369701.

Johansen:1980:CAS

[JS80] Søren Glud Johansen and Shaler Stidham, Jr. Control of arrivals to a stochastic input-output system. Advances in Applied Probability, 12(4):972–999, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426752.

Jo:1983:OSR

[JS83] Kyung Y. Jo and Shaler Stidham, Jr. Optimal service-rate control of M/G/1 queueing systems using phase methods. Advances in Applied Probability, 15(3):616–637, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426622.

Jonckheere:2014:SMD

[JS14] Matthieu Jonckheere and Seva Shneer. Stability of multidimensional birth-and-death processes with state-dependent 0homogeneous jumps. Advances in Applied Probability, 46(1): 59–75, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1396360103.

Jelenkovic:2015:DNR

[JS15] Predrag R. Jelenković and Evangelia D. Skiani. Distribution of the number of retransmissions of bounded documents. Advances in Applied Probability, 47(2):425–449, June 2015. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1435236982.

Janson:2019:PAW

[JSS19] Svante Janson, Subhabrata Sen, and Joel Spencer. Preferential attachment when stable. Advances in Applied Probability, 51(4):1067–1108, December 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/preferential-attachment-
when-stable/9F7BDA50951512BDCF3B464154884A77.

Jelenkovic:2013:CHT

[JT13] Predrag R. Jelenković and Jian Tan. Characterizing heavytailed distributions induced by retransmissions. Advances in Applied Probability, 45(1):106–138, March 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/41806002.

Jahnel:2022:SPC

[JT22] Benedikt Jahnel and András Tóbiás. SINR percolation for Cox point processes with random powers. Advances in Applied Probability, 54(1):227–253, March 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/sinr-percolation-for-cox-point-processes-
with-random-powers/37869585A9728948A18CEF2BF5DF919A.

Jedidi:2018:WPP

[JV18] Wissem Jedidi and Stavros Vakeroudis. Windings of planar processes, exponential functionals and Asian options. Advances in Applied Probability, 50(3):726–742, September 2018. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/windings-of-
planar-processes-exponential-functionals-and-asian-options/
1CF769E70E465DE4FBB7D9F75F1655D8.

Janssen:2015:SMS

[JvL15] A. J. E. M. Janssen and J. S. H. van Leeuwaarden. Staffing many-server systems with admission control and retrials. Advances in Applied Probability, 47(2):450–475, June 2015. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1435236983.

Janssen:2008:GEB

[JVZ08] A. J. E. M. Janssen, J. S. H. Van Leeuwaarden, and B. Zwart. Gaussian expansions and bounds for the Poisson distribution applied to the Erlang B formula. Advances in Applied Probability, 40(1):122–143, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443573.

Jordan:2015:PTR

[JW15] Jonathan Jordan and Andrew R. Wade. Phase transitions for random geometric preferential attachment graphs. Advances in Applied Probability, 47(2):565–588, June 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1435236988.

Jarner:2004:CBC

[JY04] Søren F. Jarner and Wai Kong Yuen. Conductance bounds on the L^{2} convergence rate of Metropolis algorithms on unbounded state spaces. Advances in Applied Probability, 36(1): 243–266, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428361.

Jasra:2020:CLT

[JY20] Ajay Jasra and Fangyuan Yu. Central limit theorems for coupled particle filters. Advances in Applied Probability, 52 (3):942–1001, September 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-

probability/article/central-limit-theorems-for-coupled-
particle-filters/4054BBD0481DB7AAFD6E864B8CE9ECAC.

Kalashnikov:1981:CRS

[KA81] V. V. Kalashnikov and S. A. Anichkin. Continuity of random sequences and approximation of Markov chains. Advances in Applied Probability, 13(2):402–414, June 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426691.

Kaijser:1978:MCA

[Kai78] Thomas Kaijser. On Markov chains associated to random systems with complete connection. Advances in Applied Probability, 10(2):295–297, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426899.

Kallenberg:1980:PDB

[Kal80] Olav Kallenberg. Path decomposition at backward times in regenerative sets. Advances in Applied Probability, 12(2):272– 273, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426540.

Kallenberg:2018:GBP

[Kal18] Olav Kallenberg. On the genealogy of branching populations and their diffusion limits. Advances in Applied Probability, 50(A):173–176, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-the-genealogy-of-branching-populations-
and-their-diffusion-limits/28FCA0624DD4D5E6D156801986786EEF.

Kampke:1987:NOC

[Käm87] Thomas Kämpke. Necessary optimality conditions for priority policies in stochastic weighted flowtime scheduling problems. Advances in Applied Probability, 19(3):749–750, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427418.

Kaplan:1973:CTM

[Kap73] Norman Kaplan. A continuous time Markov branching model with random environments. Advances in Applied Probability, 5(1):37–54, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425963.

Karr:1976:LRM

[Kar76] Alan F. Karr. Lévy random measures. Advances in Applied Probability, 8(2):225–226, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425873.

Karr:1980:POP

[Kar80] Alan F. Karr. A partially observed Poisson process. Advances in Applied Probability, 12(2):273, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426541.

Karatzas:1983:CSS

[Kar83] Ioannis Karatzas. A class of singular stochastic control problems. Advances in Applied Probability, 15(2):225–254, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426435.

Karatzas:1985:OSF

[Kar85a] Ioannis Karatzas. An overview of stochastic filtering theory. Advances in Applied Probability, 17(2):249–251, June 1985. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427134.

Karr:1985:ISP

[Kar85b] Alan F. Karr. Inference for stationary point processes. Advances in Applied Probability, 17(2):251–252, June 1985. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427135.

Karr:1986:ISR

[Kar86] Alan F. Karr. Inference for stationary random fields given Poisson samples. Advances in Applied Probability, 18(2):406– 422, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427306.

Karlsen:1990:EMS

[Kar90] Hans Arnfinn Karlsen. Existence of moments in a stationary stochastic difference equation. Advances in Applied Probability, 22(1):129–146, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427601.

Karatapanis:2021:ODS

[Kar21] Konstantinos Karatapanis. One-dimensional system arising in stochastic gradient descent. Advances in Applied Probability, 53(2):575–607, June 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/onedimensional-system-arising-in-
stochastic-gradient-descent/C2552C9770B52C426783AFF1CADAEFF6.

Kasper:1978:SLF

[Kas78] E. F. Kasper. Stochastic lifecycle forecasting processes. Advances in Applied Probability, 10(2):321–322, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426922.

Kassab:1980:GSD

[Kas80] J. Y. Kassab. Generalisation of statistical differentials with applications to estimation of autocorrelation in general linear processes. Advances in Applied Probability, 12(2):315–316, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426595.

Katzenbeisser:1990:JDN

[Kat90] W. Katzenbeisser. On the joint distribution of the number of upper and lower records and the number of inversions in a random sequence. Advances in Applied Probability, 22(4):957–960, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427570.

Kaplan:1985:BRS

[KD85] Norman Kaplan and Thomas Darden. The behavior of the ratio of a small-noise Markov chain to its deterministic approximation. Advances in Applied Probability, 17(4):731–747, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427085.

Karlin:1992:LDM

[KD92] Samuel Karlin and Amir Dembo. Limit distributions of maximal segmental score among Markov-dependent partial sums. Advances in Applied Probability, 24(1):113–140, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427732.

Khalil:1994:STP

[KD94] Z. Khalil and B. Dimitrov. The service time properties of an unreliable server characterize the exponential distribution. Advances in Applied Probability, 26(1):172–182, March 1994. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427585.

Knessl:2008:AAF

[KD08] Charles Knessl and Diego Ernesto Dominici. Asymptotic analysis of a fluid model modulated by an M/M/1 queue. Advances in Applied Probability, 40(3):856–881, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443611.

Krenzler:2016:JNN

[KDO16] Ruslan Krenzler, Hans Daduna, and Sonja Otten. Jackson networks in nonautonomous random environments. Advances in Applied Probability, 48(2):315–331, June 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1465490750.

Keiding:1974:MLE

[Kei74a] N. Keiding. Maximum likelihood estimation in the birthand-death process. Advances in Applied Probability, 6(2):242– 244, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426279.

Keilson:1974:STE

[Kei74b] J. Keilson. Sojourn times, exit times and jitter in multivariate Markov processes. Advances in Applied Probability, 6(4): 747–756, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426190.

Keilson:1974:STF

[Kei74c] J. Keilson. Sojourn times, fluctuation and rare events; the role of jitter. Advances in Applied Probability, 6(2):197–200, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426245.

Keilson:1976:STD

[Kei76] J. Keilson. The structure of transient distribution functions for time-shared computer systems. Advances in Applied Probability, 8(2):219, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425868.

Keiding:1978:SIB

[Kei78] Niels Keiding. Statistical inference in branching processes. Advances in Applied Probability, 10(2):266, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426874.

Kelly:1976:NQ

[Kel76] F. P. Kelly. Networks of queues. Advances in Applied Probability, 8(2):416–432, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425912.

Kelly:1977:ERM

[Kel77] F. P. Kelly. Exact results for the Moran neutral allele model. Advances in Applied Probability, 9(2):197–201, June 1977. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426347.

Kelly:1982:TSB

[Kel82] F. P. Kelly. The throughput of a series of buffers. Advances in Applied Probability, 14(3):633–653, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426678.

Kelly:1986:BPL

[Kel86] F. P. Kelly. Blocking probabilities in large circuit-switched networks. Advances in Applied Probability, 18(2):473–505, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427309.

Kelly:1988:RCS

[Kel88] F. P. Kelly. Routing in circuit-switched networks: Optimization, shadow prices and decentralization. Advances in Applied Probability, 20(1):112–144, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427273.

Kelly:2012:HPM

[Kel12] Michael Kelly. A hierarchical probability model of colon cancer. Advances in Applied Probability, 44(4):1052–1083, December 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/41805990.

Kemperman:1973:PTS

[Kem73] J. H. B. Kemperman. On Poisson type stationary mass distributions in R^{n}. Advances in Applied Probability, 5(1):16, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425950.

Kemperman:1975:EBC

[Kem75] J. H. B. Kemperman. The ergodic behavior of a class of transformations including the a(tan x)+b function. Advances in Applied Probability, 7(2):254, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426068.

Kemp:1987:MRV

[Kem87] Adrienne W. Kemp. The moments of the random variable for the number of returns of a simple random walk. Advances in Applied Probability, 19(2):505–507, June 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427430.

Kennedy:1972:RCQa

[Ken72a] Douglas P. Kennedy. Rates of convergence for queues in heavy traffic. I. Advances in Applied Probability, 4(2):357–381, August 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426004.

Kennedy:1972:RCQb

[Ken72b] Douglas P. Kennedy. Rates of convergence for queues in heavy traffic. II: Sequences of queueing systems. Advances in Applied Probability, 4(2):382–391, August 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1426005.

Kennedy:1974:RWC

[Ken74] D. P. Kennedy. Random walks conditioned on the proportion of time spent on the positive half-line. Advances in Applied Probability, 6(2):250, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426284.

Kendall:1977:DS

[Ken77a] D. G. Kendall. The diffusion of shape. Advances in Applied Probability, 9(3):428–430, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426091.

Kendall:1977:I

[Ken77b] D. G. Kendall. Introduction. Advances in Applied Probability, 9(3):424–425, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426087.

Kent:1977:CDF

[Ken77c] J. T. Kent. Convolutions and diffusions on fractionaldimensional spheres, with applications to the von Mises–Fisher distribution. Advances in Applied Probability, 9(3):436–437, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426098.

Kennedy:1978:SCN

[Ken78a] Douglas P. Kennedy. On sets of countable non-negative matrices and Markov decision processes. Advances in Applied Probability, 10(3):633–646, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426638.

Kent:1978:TRD

[Ken78b] John Kent. Time-reversible diffusions. Advances in Applied Probability, 10(4):819–835, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426661.

Kent:1979:AP

[Ken79] John Kent. Acknowledgement of priority. Advances in Applied Probability, 11(4):888, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1426865.

Kendall:1985:EDS

[Ken85] D. G. Kendall. Exact distributions for shapes of random triangles in convex sets. Advances in Applied Probability, 17(2):308– 329, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427143.

Kendall:1988:SCD

[Ken88a] Wilfrid S. Kendall. Symbolic computation and the diffusion of shapes of triads. Advances in Applied Probability, 20(4):775–797, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427360.

Kennedy:1988:NSS

[Ken88b] D. P. Kennedy. A note on stochastic search methods for global optimization. Advances in Applied Probability, 20(2):476– 478, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427401.

Kendall:1995:HLO

[Ken95] David G. Kendall. How to look at objects in a five-dimensional shape space: Looking at geodesics. Advances in Applied Probability, 27(1):35–43, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428093.

Kendall:1996:DMB

[Ken96] Wilfrid S. Kendall. A diffusion model for Bookstein triangle shape. Advances in Applied Probability, 28(2):334–335, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428045.

Kendall:1998:DMB

[Ken98] Wilfrid S. Kendall. A diffusion model for Bookstein triangle shape. Advances in Applied Probability, 30(2):317–334, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427970.

Kendall:2000:SCD

[Ken00] Wilfrid S. Kendall. Stationary countable dense random sets. Advances in Applied Probability, 32(1):86–100, March 2000. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428228.

Kertz:1978:OPN

[Ker78] R. P. Kertz. On ϵ-optimal plans for a non-stationary dynamic programming model with general total reward function. Advances in Applied Probability, 10(2):310–311, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426911.

Kertz:1986:COV

[Ker86] Robert P. Kertz. Comparison of optimal value and constrained maxima expectations for independent random variables. Advances in Applied Probability, 18(2):311–340, June 1986. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427302.

Kesten:1970:QTMa

[Kes70a] Harry Kesten. Quadratic transformations: A model for population growth. I. Advances in Applied Probability, 2(1):1–82, Spring 1970. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
3518344.

Kesten:1970:QTMb

[Kes70b] Harry Kesten. Quadratic transformations: A model for population growth. II. Advances in Applied Probability, 2(2):179–228, Autumn 1970. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426318.

Kesten:1972:LTSa

[Kes72a] Harry Kesten. Limit theorems for stochastic growth models. I. Advances in Applied Probability, 4(2):193–232, August 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425996.

Kesten:1972:LTSb

[Kes72b] Harry Kesten. Limit theorems for stochastic growth models. II. Advances in Applied Probability, 4(3):393–428, December 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425988.

Kesten:1976:RCM

[Kes76] Harry Kesten. Recurrence criteria for multi-dimensional Markov chains and multi-dimensional linear birth and death processes. Advances in Applied Probability, 8(1):58–87, March 1976. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426022.

Kesten:1978:BDA

[Kes78] Harry Kesten. Branching diffusion with absorption. Advances in Applied Probability, 10(2):267, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426875.

Kesten:1980:TCP

[Kes80] Harry Kesten. On the time constant and path length of firstpassage percolation. Advances in Applied Probability, 12(4): 848–863, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426744.

Kesten:1995:RLT

[Kes95] Harry Kesten. A ratio limit theorem for (sub) Markov chains on 1, 2, …; with bounded jumps. Advances in Applied Probability, 27(3):652–691, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428129.

Khalil:1990:SLC

[KF90] Z. Khalil and G. Falin. A star-like circuit-switching network with multistage service. Advances in Applied Probability, 22(4): 965–967, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427572.

Khan:1980:CLT

[Kha80] M. S. Ali Khan. Central limit theorem for stochastic processes and its application in the theory of dams. Advances in Applied Probability, 12(2):284–285, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426555.

Khan:1986:IDR

[Kha86] M. S. Ali Khan. An infinite dam with random withdrawal policy. Advances in Applied Probability, 18(4):933–951, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427257.

Khorunzhy:2001:SRM

[Kho01] A. Khorunzhy. Sparse random matrices: Spectral edge and statistics of rooted trees. Advances in Applied Probability, 33 (1):124–140, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428444.

Komaki:1992:UMK

[KI92] Fumiyasu Komaki and Yoshiaki Itoh. A unified model for Kakutani’s interval splitting and Rényi’s random packing. Advances in Applied Probability, 24(2):502–505, June 1992. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427704.

Kiderlen:1996:PTS

[Kid96] Markus Kiderlen. On projected thick sections of stationary point processes and Boolean models. Advances in Applied Probability, 28(2):335, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428046.

Kiderlen:1998:BSM

[Kid98] Markus Kiderlen. Blaschke section means of stationary random sets and particle processes. Advances in Applied Probability, 30 (2):281–282, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427952.

Kiderlen:2001:NPE

[Kid01] Markus Kiderlen. Non-parametric estimation of the directional distribution of stationary line and fibre processes. Advances in Applied Probability, 33(1):6–24, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428438.

Kiderlen:2002:DMN

[Kid02] Markus Kiderlen. Determination of the mean normal measure from isotropic means of flat sections. Advances in Applied Probability, 34(3):505–519, September 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1428244.

Kiderlen:2008:EMN

[Kid08] Markus Kiderlen. Estimation of the mean normal measure from flat sections. Advances in Applied Probability, 40(1):31– 48, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443569.

Kieu:1998:PSS

[Kiê98] Kiên Kiêu. Precision of systematic sampling of spatial structures. Advances in Applied Probability, 30(2):282–283, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427953.

Kijima:1992:FMP

[Kij92] Masaaki Kijima. Further monotonicity properties of renewal processes. Advances in Applied Probability, 24(3):575–588, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427480.

Kimmel:1983:PPA

[Kim83] Marek Kimmel. The point-process approach to age- and timedependent branching processes. Advances in Applied Probability, 15(1):1–20, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426979.

Kingman:1973:HMP

[Kin73] J. F. C. Kingman. Homecomings of Markov processes. Advances in Applied Probability, 5(1):66–102, April 1973. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425965. See remarks

[Get74].

Kingman:1977:TSU

[Kin77] J. F. C. Kingman. The thrown string: An unsolved problem. Advances in Applied Probability, 9(3):431, September 1977. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426092.

Kingman:1979:DSM

[Kin79] J. F. C. Kingman. Deterministic and stochastic models in population genetics. Advances in Applied Probability, 11(2):264, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426795.

King:2021:FLR

[Kin21] Christopher King. The fluid limit of a random graph model for a shared ledger. Advances in Applied Probability, 53(1):81–106, March 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/fluid-limit-of-a-random-graph-model-
for-a-shared-ledger/E0BF7F39C81F6E2441CA6C202F870A70.

Kirchner:2023:CCC

[Kir23] Matthias Kirchner. Critical cluster cascades. Advances in Applied Probability, 55(2):357–381, June 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/critical-cluster-cascades/DFCF06FF6F83F88D4329FF7FD29A0682.

Kitagawa:1971:ISA

[Kit71] Tosio Kitagawa. Information science approach to biomathematics and its implications to ecology and quantitative epidemiology. Advances in Applied Probability, 3(2):198–200, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426154.

Kiderlen:2003:EDM

[KJ03] Markus Kiderlen and Eva B. Vedel Jensen. Estimation of the directional measure of planar random sets by digitization. Advances in Applied Probability, 35(3):583–602, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428209.

Koski:2018:ETA

[KJH18] Timo Koski, Brita Jung, and Göran Högnäs. Exit times for ARMA processes. Advances in Applied Probability, 50 (A):191–195, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/exit-times-for-arma-processes/05B360FFC24FDD9A0FEC5CAD3EE2211D.

Karlin:1973:CEC

[KK73] Samuel Karlin and Norman Kaplan. Criteria for extinction of certain population growth processes with interacting types. Advances in Applied Probability, 5(2):183–199, August 1973. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426032.

Kendall:1980:ATD

[KK80] David G. Kendall and Wilfrid S. Kendall. Alignments in twodimensional random sets of points. Advances in Applied Probability, 12(2):380–424, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426603.

Kennedy:1990:LTT

[KK90] Douglas P. Kennedy and Robert P. Kertz. Limit theorems for threshold-stopped random variables with applications to optimal stopping. Advances in Applied Probability, 22(2):396– 411, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427542.

Kennedy:1992:LTS

[KK92] Douglas P. Kennedy and Robert P. Kertz. Limit theorems for suprema, threshold-stopped random variables and last exits of i.i.d. random variables with costs and discounting, with applications to optimal stopping. Advances in Applied Probability, 24(2):241–266, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427691.

Kou:2003:MGS

[KK03] S. C. Kou and S. G. Kou. Modeling growth Stocks via birth– death processes. Advances in Applied Probability, 35(3):641–664, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428212.

Kolossvary:2015:FPP

[KK15] István Kolossváry and Júlia Komjáthy. First passage percolation on inhomogeneous random graphs. Advances in Applied Probability, 47(2):589–610, June 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1435236989.

Kelbert:1990:PTA

[KKSS90] M. Ya. Kelbert, R. P. Kopeika, R. N. Shamsiev, and Yu. M. Sukhov. Perturbation theory approach for a class of hybrid switching networks with small transit flows. Advances in Applied Probability, 22(1):211–229, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427605.

Konstantopoulos:2008:ASF

[KKSS08] Takis Konstantopoulos, Andreas E. Kyprianou, Paavo Salminen, and Marina Sirviö. Analysis of stochastic fluid queues driven by local-time processes. Advances in Applied Probability, 40(4):1072–1103, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443625.

Kobayashi:2006:PMR

[KKT06] Kei Kobayashi, Hideki Kawasaki, and Akimichi Takemura. Parallel matching for ranking all teams in a tournament. Advances in Applied Probability, 38(3):804–826, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443472.

Kolossvary:2016:DDR

[KKV16] István Kolossváry, Júlia Komjáthy, and Lajos Vágó. Degrees and distances in random and evolving Apollonian networks. Advances in Applied Probability, 48(3):865–902, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1474296319.

Karlin:1979:NCM

[KL79] Samuel Karlin and Uri Liberman. A natural class of multilocus recombination processes and related measures of crossover interference. Advances in Applied Probability, 11(3):479–501, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426950.

Karlin:1983:MIC

[KL83] Samuel Karlin and Uri Liberman. Measuring interference in the chiasma renewal formation process. Advances in Applied Probability, 15(3):471–487, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1426615.

Kendall:1987:SED

[KL87] David G. Kendall and Hui-Lin Le. The structure and explicit determination of convex-polygonally generated shape-densities. Advances in Applied Probability, 19(4):896–916, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427107.

Kume:2000:EFM

[KL00] Alfred Kume and Huiling Le. Estimating Fréchet means in Bookstein’s shape space. Advances in Applied Probability, 32(3): 663–674, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428406.

Kume:2003:FMS

[KL03] Alfred Kume and Huiling Le. On Fréchet means in simplex shape spaces. Advances in Applied Probability, 35(4):885–897, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428490.

Kuronen:2013:HCT

[KL13] Mikko Kuronen and Lasse Leskelä. Hard-core thinnings of germ–grain models with power-law grain sizes. Advances in Applied Probability, 45(3):595–625, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1377868531.

Klaassen:1985:SU

[Kla85] Chris A. J. Klaassen. Strong unimodality. Advances in Applied Probability, 17(4):905–907, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427093.

Klebaner:1982:BRW

[Kle82] C. F. Klebaner. Branching random walk in varying environments. Advances in Applied Probability, 14(2):359–367, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426526.

Klebaner:1984:PSD

[Kle84] F. C. Klebaner. On population-size-dependent branching processes. Advances in Applied Probability, 16(1):30–55, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427223.

Kliemann:1980:ETS

[Kli80] W. Kliemann. Ergodic theory of stochastic systems and nonlinear control. Advances in Applied Probability, 12(2):300– 301, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426576.

Kelbert:2005:FRF

[KLRM05] M. Ya. Kelbert, N. N. Leonenko, and M. D. Ruiz-Medina. Fractional random fields associated with stochastic fractional heat equations. Advances in Applied Probability, 37(1):108– 133, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
30037318.

Kochar:2002:TTT

[KLS02] Subhash C. Kochar, Xiaohu Li, and Moshe Shaked. The total time on test transform and the excess wealth stochastic orders of distributions. Advances in Applied Probability, 34(4):826–845, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428324.

Klein:1980:MCT

[KM80] B. Klein and P. D. M. Macdonald. The multitype continuoustime Markov branching process in a periodic environment. Advances in Applied Probability, 12(1):81–93, March 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426495.

Kontovasilis:1995:MMT

[KM95] Kimon P. Kontovasilis and Nikolas M. Mitrou. Markovmodulated traffic with nearly complete decomposability characteristics and associated fluid queueing models. Advances in Applied Probability, 27(4):1144–1185, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427937.

Kieu:1996:ERM

[KM96] Kiên Kiêu and Marianne Mora. Estimating the reduced moments of a random measure. Advances in Applied Probability, 28(2):335–336, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428047.

Kieu:1999:ERM

[KM99] Kiên Kiêu and Marianne Mora. Estimating the reduced moments of a random measure. Advances in Applied Probability, 31(1):48–62, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427990.

Kendall:2000:PSU

[KM00] Wilfrid S. Kendall and Jesper Møller. Perfect simulation using dominating processes on ordered spaces, with application to locally stable point processes. Advances in Applied Probability, 32(3):844–865, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428416.

Kinnison:2010:SMA

[KM10a] Adam L. Kinnison and Peter Mörters. Simultaneous multifractal analysis of the branching and visibility measure on a Galton–Watson tree. Advances in Applied Probability, 42(1): 226–245, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25683814.

Klenke:2010:SOC

[KM10b] Achim Klenke and Lutz Mattner. Stochastic ordering of classical discrete distributions. Advances in Applied Probability, 42 (2):392–410, June 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25683826.

Kholfi:2012:CTZ

[KM12] Sanaa Kholfi and Hosam M. Mahmoud. The class of tenable zero-balanced Pólya urn schemes: Characterization and Gaussian phases. Advances in Applied Probability, 44(3):702–728, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
41714072.

Kalpathy:2014:PFL

[KM14a] Ravi Kalpathy and Hosam Mahmoud. Perpetuities in fair leader election algorithms. Advances in Applied Probability, 46(1): 203–216, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1396360110.

Kobayashi:2014:TAS

[KM14b] Masahiro Kobayashi and Masakiyo Miyazawa. Tail asymptotics of the stationary distribution of a two-dimensional reflecting random walk with unbounded upward jumps. Advances in Applied Probability, 46(2):365–399, June 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1401369699.

Kuchta:2014:MCE

[KM14c] Małgorzata Kuchta and Michal Morayne. Monotone case for an extended process. Advances in Applied Probability, 46(4):1106– 1125, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1418396245.

Kluppelberg:2015:GFL

[KM15] Claudia Klüppelberg and Muneya Matsui. Generalized fractional Lévy processes with fractional Brownian motion limit.

Advances in Applied Probability, 47(4):1108–1131, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1449859802.

Kelbert:2022:MSS

[KMF22] Mark Kelbert and Harold A. Moreno-Franco. On a mixed singular/switching control problem with multiple regimes. Advances in Applied Probability, 54(3):743–782, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/on-
a-mixed-singularswitching-control-problem-with-multiple-
regimes/9C367BEB7920C900B8A1540F6EEBC69C.

Kozubowski:2006:FLM

[KMP06] T. J. Kozubowski, M. M. Meerschaert, and K. Podgórski. Fractional Laplace motion. Advances in Applied Probability, 38 (2):451–464, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443450.

Kent:1996:CCM

[KMW96] John T. Kent, Kanti V. Mardia, and Alistair N. Walder. Conditional cyclic Markov random fields. Advances in Applied Probability, 28(1):1–12, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427910.

Kratz:2016:CFE

[KN16] Marie Kratz and Werner Nagel. On the capacity functional of excursion sets of Gaussian random fields on R^{2}. Advances in Applied Probability, 48(3):712–725, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1474296311.

Knessl:1991:TBR

[Kne91] Charles Knessl. On the transient behavior of the repairman problem. Advances in Applied Probability, 23(2):327–354, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427752.

Knessl:2000:EAS

[Kne00] Charles Knessl. Exact and asymptotic solutions to a PDE that arises in time-dependent queues. Advances in Applied Probability, 32(1):256–283, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428237.

Kijima:1997:LCD

[KNPV97] M. Kijima, M. G. Nair, P. K. Pollett, and E. A. Van Doorn. Limiting conditional distributions for birth–death processes. Advances in Applied Probability, 29(1):185–204, March 1997. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427866.

Kulkarni:1987:CTJ

[KNT87] V. G. Kulkarni, V. F. Nicola, and K. S. Trivedi. The completion time of a job on multimode systems. Advances in Applied Probability, 19(4):932–954, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427109.

Karlin:1987:CLA

[KO87] Samuel Karlin and Friedemann Ost. Counts of long aligned word matches among random letter sequences. Advances in Applied Probability, 19(2):293–351, June 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427422.

Kochar:1990:PSP

[Koc90] Subhash C. Kochar. On preservation of some partial orderings under shock models. Advances in Applied Probability, 22(2):508– 509, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1427555. See correction [NP93].

Kogan:1989:EAC

[Kog89] Yaakov Kogan. Exact analysis for a class of simple, circuitswitched networks with blocking. Advances in Applied Probability, 21(4):952–955, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427782.

Kohn:1978:APT

[Koh78] R. Kohn. Asymptotic properties of time domain Gaussian estimators. Advances in Applied Probability, 10(2):339–359, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426939.

Kohn:1980:SDSa

[Koh80a] R. Kohn. On the spectral decomposition of stationary time series using Walsh functions. I. Advances in Applied Probability, 12(1):183–199, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426501.

Kohn:1980:SDSb

[Koh80b] R. Kohn. On the spectral decomposition of stationary time series using Walsh functions. II. Advances in Applied Probability, 12(2):462–474, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426606.

Kollerstrom:1981:SOH

[Köl81] Julian Köllerström. A second-order heavy traffic approximation for the queue GI/G/1. Advances in Applied Probability, 13(1): 167–185, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426473.

Kolesnik:2014:PDF

[Kol14] Alexander D. Kolesnik. Probability distribution function for the Euclidean distance between two telegraph processes. Advances in Applied Probability, 46(4):1172–1193, December 2014. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1418396248.

Kolodziejek:2018:PLT

[Kol18] Bartosz Kolodziejek. On perpetuities with light tails. Advances in Applied Probability, 50(4):1119–1154, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/on-
perpetuities-with-light-tails/D2500EC4A87DC69D7D14081F6F58E911.

Komaki:1996:HGM

[Kom96] Fumiyasu Komaki. Homogeneous Gaussian Markov processes on general lattices. Advances in Applied Probability, 28(1):189– 206, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1427917. See correction [Ano98f].

Kombrink:2018:RTP

[Kom18] Sabrina Kombrink. Renewal theorems for processes with dependent interarrival times. Advances in Applied Probability, 50(4):1193–1216, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/renewal-theorems-for-processes-with-
dependent-interarrival-times/4D966B53796273588967F566857895DB.

Konig:1977:SPB

[Kön77] D. König. Stochastic processes with basic stationary marked point processes. Advances in Applied Probability, 9(3):440–442, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426101.

Koning:1989:SLR

[Kon89a] A. J. Koning. On single-Lane roads. Advances in Applied Probability, 21(1):142–158, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427201.

Konstantopoulos:1989:CPB

[Kon89b] Panagiotis Konstantopoulos. A comment on the paper ‘On the Intensity of Crossings by a Shot Noise Process”. Advances in Applied Probability, 21(2):473–474, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427171. See [Hsi87].

Kosten:1976:DBM

[Kos76] L. Kosten. Dynamic behaviour of a multi-entry buffer. Advances in Applied Probability, 8(2):243–244, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425892.

Kosten:1978:POT

[Kos78] L. Kosten. Parametrisation of overflow traffic in telecommunication and data handling. Advances in Applied Probability, 10 (2):337–338, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426937.

Koch:2003:STR

[KOS03] Karsten Koch, Joachim Ohser, and Katja Schladitz. Spectral theory for random closed sets and estimating the covariance via frequency space. Advances in Applied Probability, 35(3): 603–613, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428210.

Kotlarski:1980:CPD

[Kot80] I. I. Kotlarski. On a characterization of probability distributions in some stochastic processes. Advances in Applied Probability, 12 (2):283, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426553.

Kotelenez:1987:FNH

[Kot87] Peter Kotelenez. Fluctuations near homogeneous states of chemical reactions with diffusion. Advances in Applied Probability, 19(2):352–370, June 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427423.

Kaspi:1983:ISP

[KP83a] H. Kaspi and D. Perry. Inventory systems of perishable commodities. Advances in Applied Probability, 15(3):674–685, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426625.

Kelly:1983:STC

[KP83b] F. P. Kelly and P. K. Pollett. Sojourn times in closed queueing networks. Advances in Applied Probability, 15(3):638–656, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426623.

Kaspi:1984:ISP

[KP84] H. Kaspi and D. Perry. Inventory systems for perishable commodities with renewal input and Poisson output. Advances in Applied Probability, 16(2):402–421, June 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427076.

Katzenbeisser:1986:NHM

[KP86] Walter Katzenbeisser and Wolfgang Panny. A note on the higher moments of the random variable T associated with the number of returns of a simple random walk. Advances in Applied Probability, 18(1):279–282, March 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427248.

Kirschenhofer:1994:HMN

[KP94] Peter Kirschenhofer and Helmut Prodinger. The higher moments of the number of returns of a simple random walk. Advances in Applied Probability, 26(2):561–563, June 1994. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427452.

Kallsen:2010:VOH

[KP10] Jan Kallsen and Arnd Pauwels. Variance-optimal hedging in general affine stochastic volatility models. Advances in Applied Probability, 42(1):83–105, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/25683808.

Kvitkovicova:2011:AIP

[KP11] Andrea Kvitkovicová and Victor M. Panaretos. Asymptotic inference for partially observed branching processes. Advances in Applied Probability, 43(4):1166–1190, December 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/23208043.

Kuba:2012:LDC

[KP12] Markus Kuba and Alois Panholzer. Limiting distributions for a class of diminishing urn models. Advances in Applied Probability, 44(1):87–116, March 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/23207857.

Knopova:2022:SPA

[KP22] Victoria Knopova and Zbigniew Palmowski. Subexponential potential asymptotics with applications. Advances in Applied Probability, 54(3):783–807, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/subexponential-potential-
asymptotics-with-applications/513B34DAB88C4C02461E57966FCB3360.

Koether:2009:LOG

[KPO09] Robb Koether, Marcus Pendergrass, and John Osoinach. The lying oracle game with a biased coin. Advances in Applied Probability, 41(4):1023–1040, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/27793917.

Karkkainen:2001:EOC

[KPUU01] Salme Kärkkäinen, Antti Penttinen, Nikolai G. Ushakov, and Alexandra P. Ushakova. Estimation of orientation characteristic of fibrous material. Advances in Applied Probability, 33(3): 559–575, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428474.

Keilson:1971:PCD

[KR71a] Julian Keilson and S. Subba Rao. A process with chain dependent growth rate. Part II: The ruin and ergodic problems. Advances in Applied Probability, 3(2):315–338, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426174.

Krus:1971:MTE

[KR71b] V. P. Krus and L. A. Rvachev. The mathematical theory of epidemics: A study of the evolution of resistance in microorganisms. Advances in Applied Probability, 3(2):206–208, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426158.

Karlin:1981:EICa

[KR81a] Samuel Karlin and Yosef Rinott. Entropy inequalities for classes of probability distributions I. The univariate case. Advances in Applied Probability, 13(1):93–112, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426469.

Karlin:1981:EICb

[KR81b] Samuel Karlin and Yosef Rinott. Entropy inequalities for classes of probability distributions II. The multivariate case. Advances in Applied Probability, 13(2):325–351, June 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426687.

Kalashnikov:1985:CPQ

[KR85] V. V. Kalashnikov and S. T. Rachev. Characterization problems in queueing and their stability. Advances in Applied Probability, 17(4):868–886, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427091.

Klebanov:1997:CTQ

[KR97] Lev B. Klebanov and Svetlozar T. Rachev. Computer tomography and quantum mechanics. Advances in Applied Probability, 29(3):595–606, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428078.

Konstantopoulos:2002:CLT

[KR02] Takis Konstantopoulos and Gregory S. Richardson. Conditional limit theorems for spectrally positive Lévy processes. Advances in Applied Probability, 34(1):158–178, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428377.

Kiderlen:2018:DVS

[KR18] Markus Kiderlen and Jan Rataj. Dilation volumes of sets of finite perimeter. Advances in Applied Probability, 50 (4):1095–1118, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/dilation-volumes-of-sets-of-finite-
perimeter/DD02FBCCEDC413C339DFAE5F9A6EFFD2.

Krickeberg:1977:SRF

[Kri77] K. Krickeberg. Statistics of random flats. Advances in Applied Probability, 9(3):432, September 1977. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426093.

Krickeberg:1978:SPP

[Kri78] K. Krickeberg. Statistics of point processes. Advances in Applied Probability, 10(2):267–268, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426876.

Krickeberg:1979:FSP

[Kri79] K. Krickeberg. Filtering spatial processes. Advances in Applied Probability, 11(2):265–266, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426796.

Klebaner:2007:TGW

[KRS07] F. C. Klebaner, U. Rösler, and S. Sagitov. Transformations of Galton–Watson processes and linear fractional reproduction. Advances in Applied Probability, 39(4):1036–1053, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/20443557.

Kim:2009:JRM

[KRW09] Yusik Kim, Rhonda Righter, and Ronald Wolff. Job replication on multiserver systems. Advances in Applied Probability, 41 (2):546–575, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/27793890.

Kryukov:1976:WIR

[Kry76] V. I. Kryukov. Wald’s identity and random walk models for neuron firing. Advances in Applied Probability, 8(2):257–277, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425904.

Kanter:1974:RAI

[KS74a] Marek Kanter and W. L. Steiger. Regression and autoregression with infinite variance. Advances in Applied Probability, 6(4): 768–783, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426192.

Kemperman:1974:SUL

[KS74b] J. H. B. Kemperman and J. C. Smit. Sharp upper and lower bounds for the moments of a martingale. Advances in Applied Probability, 6(2):244–245, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426280.

Kryscio:1975:CEP

[KS75] Richard J. Kryscio and Norman C. Severo. Computational and estimation procedures in multidimensional right-shift processes and some applications. Advances in Applied Probability, 7(2):349–382, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426081.

Kramli:1978:OLF

[KS78] A. Kramli and J. Soltesz. On the optimality of the ’least frequently used’ page replacement algorithm. Advances in Applied Probability, 10(2):338, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426938.

Kotz:1980:SNA

[KS80] S. Kotz and D. N. Shanbhag. Some new approaches to probability distributions. Advances in Applied Probability, 12(4): 903–921, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426748.

Keilson:1983:ETT

[KS83a] J. Keilson and U. Sumita. Evaluation of the total time in system in a Preempt/Resume priority queue via a modified Lindley process. Advances in Applied Probability, 15(4):840–856, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427327.

Keilson:1983:DTM

[KS83b] Julian Keilson and Ushio Sumita. The depletion time for M/G/1 systems and a related limit theorem. Advances in Applied Probability, 15(2):420–443, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426443.

Konig:1990:ECV

[KS90] Dieter König and Volker Schmidt. Extended and conditional versions of the PASTA property. Advances in Applied Probability, 22(2):510–512, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427556.

Keilson:1993:MMM

[KS93] J. Keilson and L. D. Servi. The matrix M/M/∞ system: Retrial models and Markov modulated sources. Advances in Applied Probability, 25(2):453–471, June 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427662.

Khmaladze:2001:CNP

[KS01a] Estate Khmaladze and Eka Shinjikashvili. Calculation of noncrossing probabilities for Poisson processes and its corollaries. Advances in Applied Probability, 33(3):702–716, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428483.

Kim:2001:ARB

[KS01b] Han S. Kim and Ness B. Shroff. On the asymptotic relationship between the overflow probability and the loss ratio. Advances in Applied Probability, 33(4):836–863, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428308.

Knight:2004:IPA

[KS04a] F. B. Knight and J. L. Steichen. An interference problem with application to crystal growth. Advances in Applied Probability, 36(3):725–746, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/4140405.

Ko:2004:RTM

[KS04b] Sung-Seok Ko and Richard F. Serfozo. Response times in M/M/s fork-join networks. Advances in Applied Probability, 36 (3):854–871, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/4140412.

Kim:2006:TBQ

[KS06] Bara Kim and Khosrow Sohraby. Tail behavior of the queue size and waiting time in a queue with discrete autoregressive arrivals. Advances in Applied Probability, 38(4):1116–1131, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443490.

Kabluchko:2009:SSL

[KS09] Zakhar Kabluchko and Evgeny Spodarev. Scan statistics of Lévy noises and marked empirical processes. Advances in Applied Probability, 41(1):13–37, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/27793863.

Kulik:2012:LTL

[KS12] Rafal Kulik and Philippe Soulier. Limit theorems for longmemory stochastic volatility models with infinite variance: Partial sums and sample covariances. Advances in Applied Probability, 44(4):1113–1141, December 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/41805992.

Kouritzin:2016:MLL

[KS16] Michael A. Kouritzin and Samira Sadeghi. Marcinkiewicz law of large numbers for outer products of heavy-tailed, long-range dependent data. Advances in Applied Probability, 48(2):349–368, June 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1465490752.

Krieger:2009:SPM

[KSC09] Abba M. Krieger and Ester Samuel-Cahn. The secretary problem of minimizing the expected rank: a simple suboptimal approach with generalizations. Advances in Applied Probability, 41(4):1041–1058, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/27793918.

Kuroda:1992:LTL

[KT92] Koji Kuroda and Hideki Tanemura. Limit theorem and large deviation principle for the Voronoi tessellation generated by a Gibbs point process. Advances in Applied Probability, 24(1):45– 70, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427729.

Khmaladze:2001:ASC

[KT01] Estate Khmaladze and N. Toronjadze. On the almost sure coverage property of Voronoi tessellation: The R^{1} case. Advances in Applied Probability, 33(4):756–764, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428303.

Kimura:2020:CCS

[KT20]  Masatoshi Kimura and Tetsuya Takine.  Characteriza-

tion of the conditional stationary distribution in Markov chains via systems of linear inequalities. Advances in Applied Probability, 52(4):1249–1283, December 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/characterization-of-the-
conditional-stationary-distribution-in-markov-chains-
via-systems-of-linear-inequalities/C6DCC021D7F7757D68A25021D01E364D.

Kazi-Tani:2019:CMD

[KTR19] Nabil Kazi-Tani and Didier Rullière. On a construction of multivariate distributions given some multidimensional marginals. Advances in Applied Probability, 51(2):487–513, June 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/on-
a-construction-of-multivariate-distributions-given-some-
multidimensional-marginals/50B4F041E28FC52902C273D98229E46A.

Kubota:2019:DBF

[Kub19] Naoki Kubota. Deviation bounds for the first passage time in the frog model. Advances in Applied Probability, 51(1):184–208, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/deviation-bounds-for-the-first-passage-
time-in-the-frog-model/70BAAA233C87279994B4464661709C67.

Kuczek:1983:GGQ

[Kuc83] Thomas Kuczek. On the GA/G/∞ queue. Advances in Applied Probability, 15(2):444–459, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426444.

Kuelbs:1975:LIL

[Kue75] J. Kuelbs. The law of the iterated logarithm for Banach space valued random variables. Advances in Applied Probability, 7 (2):246–247, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426058.

Kufer:1994:ABR

[Küf94] K.-H. Küfer. On the approximation of a ball by random polytopes. Advances in Applied Probability, 26(4):876–892, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427895.

Kunsch:1983:AUI

[Kün83] H. Künsch. Asymptotically unbiased inference for Ising models. Advances in Applied Probability, 15(4):887–888, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427331.

Kurtz:1978:AMS

[Kur78] Thomas G. Kurtz. Applications of multiparameter stopping times. Advances in Applied Probability, 10(2):282–283, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426886.

Kurtz:2018:GCA

[Kur18] Thomas G. Kurtz. Genealogical constructions and asymptotics for continuous-time Markov and continuous-state branching processes. Advances in Applied Probability, 50(A): 197–209, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/genealogical-constructions-and-asymptotics-
for-continuoustime-markov-and-continuousstate-branching-
processes/6D109DC1C0E7E5CF77BCBEDB193F4D02.

Kushner:1994:ALT

[Kus94] Harold J. Kushner. Approximations of large trunk line systems under heavy traffic. Advances in Applied Probability, 26 (4):1063–1094, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427905.

Kennerberg:2018:JLP

[KV18] Philip Kennerberg and Stanislav Volkov. Jante’s law process. Advances in Applied Probability, 50(2):414–439, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/jantes-law-process/
0CE018653A90453DD1DBA45A4C489AE7.

Kendall:1999:QIP

[KVB99] W. S. Kendall, M. N. M. Van Lieshout, and A. J. Baddeley. Quermass-interaction processes: Conditions for stability. Advances in Applied Probability, 31(2):315–342, June 1999. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428112.

Kendall:1987:ODC

[KW87] Wilfrid S. Kendall and Mark Westcott. One-dimensional classical scattering processes and the diffusion limit. Advances in Applied Probability, 19(1):81–105, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427374.

Konstantopoulos:1988:WCS

[KW88] Panagiotis Konstantopoulos and Jean Walrand. On the weak convergence of stochastic processes with embedded point processes. Advances in Applied Probability, 20(2):473–475, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427400.

Kella:1990:DAQ

[KW90a] Offer Kella and Ward Whitt. Diffusion approximations for queues with server vacations. Advances in Applied Probability, 22(3):706–729, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427465.

Konstantopoulos:1990:ENG

[KW90b] Panagiotis Konstantopoulos and Jean Walrand. On the ergodicity of networks of · /GI/1/N queues. Advances in Applied Probability, 22(1):263–267, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427613.

Kendall:2003:IMM

[KW03a] W. S. Kendall and R. G. Wilson. Ising models and multiresolution quad-trees. Advances in Applied Probability, 35(1):96– 122, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428275.

Kou:2003:FPT

[KW03b] S. G. Kou and Hui Wang. First passage times of a jump diffusion process. Advances in Applied Probability, 35(2):504–531, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428433.

Kersting:2018:TAC

[KW18] Götz Kersting and Anton Wakolbinger. The time to absorption in Λ-coalescents. Advances in Applied Probability, 50(A):177–190, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/time-to-absorption-in-coalescents/
E58F417E486D890558BDDDB50EADC4A1.

Klatt:2020:GFF

[KW20] Michael A. Klatt and Steffen Winter. Geometric functionals of fractal percolation. Advances in Applied Probability, 52(4):1085–1126, December 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/geometric-functionals-of-fractal-
percolation/0D537B0C1F39498776269DDD2E6E59F9.

Kwerel:1975:BPU

[Kwe75] Seymour M. Kwerel. Bounds on the probability of the union and intersection of m events. Advances in Applied Probability, 7(2):431–448, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426084.

Khudabukhsh:2022:FCL

[KWRK22] Wasiur R. Khudabukhsh, Casper Woroszylo, Grzegorz A. Rempała, and Heinz Koeppl. A functional central limit theorem for SI processes on configuration model graphs. Advances in Applied Probability, 54(3):880–912, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/functional-central-limit-
theorem-for-si-processes-on-configuration-model-graphs/
8ECE68C7706243E769C7B4BF565C8F3E.

Kallianpur:1994:SME

[KX94] G. Kallianpur and J. Xiong. Stochastic models of environmental Pollution. Advances in Applied Probability, 26(2):377–403, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427442.

Kovchegov:2023:IGW

[KXZ23] Yevgeniy Kovchegov, Guochen Xu, and Ilya Zaliapin. Invariant Galton–Watson trees: metric properties and attraction with respect to generalized dynamical pruning. Advances in Applied Probability, 55(2):643–671, June 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/invariant-galtonwatson-trees-metric-
properties-and-attraction-with-respect-to-generalized-
dynamical-pruning/7DB1713F662D99C51A96EFA18CFC4684.

Kyprianou:1971:VWT

[Kyp71] E. Kyprianou. The virtual waiting time of the GI/G/1 queue in heavy traffic. Advances in Applied Probability, 3(2):249–268, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426170.

Konstantopoulos:1992:SAS

[KZ92] P. Konstantopoulos and Michael A. Zazanis. Sensitivity analysis for stationary and ergodic queues. Advances in Applied Probability, 24(3):738–750, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427487.

Konstantopoulos:1994:SAS

[KZ94] P. Konstantopoulos and M. Zazanis. Sensitivity analysis for stationary and ergodic queues: Additional results. Advances in Applied Probability, 26(2):556–560, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427451.

Kou:2016:FPT

[KZ16] Steven Kou and Haowen Zhong. First-passage times of twodimensional Brownian motion. Advances in Applied Probability, 48(4):1045–1060, December 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1482548428.

Kabluchko:2020:APG

[KZ20] Zakhar Kabluchko and Dmitry Zaporozhets. Absorption probabilities for Gaussian polytopes and regular spherical simplices. Advances in Applied Probability, 52(2):588–616, June 2020. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/absorption-probabilities-
for-gaussian-polytopes-and-regular-spherical-simplices/
3B09449F8DFF00AECC983239BBC6C70F.

Liu:2002:FMD

[LA02] Bin Liu and Attahiru Sule Alfa. A fluid model with data message discarding. Advances in Applied Probability, 34(2):329– 348, June 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428291.

Lacaux:2004:SRS

[Lac04] Céline Lacaux. Series representation and simulation of multifractional Lévy motions. Advances in Applied Probability, 36 (1):171–197, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428358.

Lacker:2020:NEE

[Lac20] Daniel Lacker. A non-exponential extension of Sanov’s theorem via convex duality. Advances in Applied Probability, 52(1):61–101, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/nonexponential-extension-of-sanovs-
theorem-via-convex-duality/F6D597DE19288E1801DBCDF47A5742BA.

Lalley:1984:LTF

[Lal84] S. P. Lalley. Limit theorems for first-passage times in linear and non-linear renewal theory. Advances in Applied Probability, 16 (4):766–803, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427340.

Lam:1993:SMR

[Lam93] C. Teresa Lam. Superposition of Markov renewal processes and applications. Advances in Applied Probability, 25(3):585–606, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427525.

Lambert:2003:CTB

[Lam03] Amaury Lambert. Coalescence times for the branching process. Advances in Applied Probability, 35(4):1071–1089, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428500.

Langen:1982:AMP

[Lan82] Hans-Joachim Langen. Applying the method of phases in the optimization of queuing systems. Advances in Applied Probability, 14(1):122–142, March 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426736.

Lane:1987:BEB

[Lan87] John A. Lane. The Berry–Esseen bound for the Poisson shotnoise. Advances in Applied Probability, 19(2):512–514, June 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427432.

Lantuejoul:1996:CSB

[Lan96] Ch. Lantuéjoul. Conditional simulation of a Boolean model. Advances in Applied Probability, 28(2):336, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428048.

Lanchier:2005:PTD

[Lan05] Nicolas Lanchier. Phase transitions and duality properties of a successional model. Advances in Applied Probability, 37(1):265– 278, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
30037324.

Lanchier:2013:RDI

[Lan13a] N. Lanchier. The role of dispersal in interacting patches subject to an Allee effect. Advances in Applied Probability, 45(4):1182– 1197, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1386857863.

Lanchier:2013:SSM

[Lan13b] N. Lanchier. Stochastic spatial model of producer-consumer systems on the lattice. Advances in Applied Probability, 45(4):1157– 1181, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1386857862.

Last:2004:EPS

[Las04] Günter Last. Ergodicity properties of stress release, repairable system and workload models. Advances in Applied Probability, 36(2):471–498, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428463.

Last:2006:SPP

[Las06] Günter Last. Stationary partitions and Palm probabilities. Advances in Applied Probability, 38(3):602–620, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/20443463.

Latouche:1977:TBQ

[Lat77] Guy Latouche. On the trade-off between queue occupancy and Server’s reward. Advances in Applied Probability, 9(2):217– 218, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426360.

Latouche:1988:PAP

[Lat88] Guy Latouche. Perturbation analysis of a phase-type queue with weakly correlated arrivals. Advances in Applied Probability, 20 (4):896–912, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427366.

Latour:1997:MGP

[Lat97] Alain Latour. The multivariate GINAR(p) process. Advances in Applied Probability, 29(1):228–248, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427868.

Laws:1992:RPQ

[Law92] C. N. Laws. Resource pooling in queueing networks with dynamic routing. Advances in Applied Probability, 24(3):699–726, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427485.

Le:2010:IDS

[LB10] H. Le and D. Barden. On the induced distribution of the shape of the projection of a randomly rotated configuration. Advances in Applied Probability, 42(2):331–346, June 2010. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/25683822.

Gall:2021:CLS

[LBCEN21] Quentin Le Gall, Bartłomiej Błaszczyszyn, Élie Cali, and Taoufik En-Najjary. Continuum line-of-sight percolation on Poisson–Voronoi tessellations. Advances in Applied Probability, 53(2):510–536, June 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/continuum-lineofsight-percolation-
on-poissonvoronoi-tessellations/C48C2562895033BFF3D598470CD9FAB6.

Li:1998:SSR

[LBZ98] Wei Li, W. John Braun, and Yiqiang Q. Zhao. Stochastic scheduling on a repairable machine with Erlang uptime distribution. Advances in Applied Probability, 30(4):1073–1088, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428030.

Leon:1976:EVF

[LC76] Jesús Alberto León and B. Charlesworth. Ecological versions of Fisher’s fundamental theorem of natural selection. Advances in Applied Probability, 8(4):639–641, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425923.

Luczak:1992:GCT

[LC92] Tomasz Luczak and Joel E. Cohen. Giant components in threeparameter random directed graphs. Advances in Applied Probability, 24(4):845–857, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427715.

Li:2008:DPS

[LC08] Junping Li and Anyue Chen. Decay property of stopped Markovian bulk-arriving queues. Advances in Applied Probability, 40 (1):95–121, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443572.

LeGall:1974:SIE

[Le 74] P. Le Gall. Stochastic integral equations applied to telecommunications traffic without delay. Advances in Applied Probability, 6(2):255–257, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426290.

LeBreton:1978:PED

[Le 78] Alain Le Breton. Parameter estimation and design of inputs in a linear dynamical system. Advances in Applied Probability, 10 (2):328–329, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426927.

Le:1989:RSTa

[Le89a] Huiling Le. Random spherical triangles i: Geometrical background. Advances in Applied Probability, 21(3):570–580, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427636.

Le:1989:RSTb

[Le89b] Huiling Le. Random spherical triangles II: Shape densities. Advances in Applied Probability, 21(3):581–594, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427637.

Le:1995:MSS

[Le95] Huiling Le. Mean size-and-shapes and mean shapes: A geometric point of view. Advances in Applied Probability, 27(1):44– 55, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428094.

Le:1996:PMS

[Le96] Huiling Le. On procrustean mean shapes and the shape of the means. Advances in Applied Probability, 28(2):336–337, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428049.

Le:1998:CPM

[Le98] Huiling Le. On the consistency of procrustean mean shapes. Advances in Applied Probability, 30(1):53–63, March 1998. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427875.

Le:2001:LFM

[Le01] Huiling Le. Locating Fréchet means with application to shape spaces. Advances in Applied Probability, 33(2):324–338, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428255.

Leadbetter:1976:PPN

[Lea76] M. R. Leadbetter. Point processes: A non-topological view. Advances in Applied Probability, 8(2):219–221, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425869.

Ledwich:1979:ARC

[Led79] G. F. Ledwich. Adaptive regulator convergence for variable set point. Advances in Applied Probability, 11(2):309, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426841.

Ledoux:2004:LDS

[Led04] James Ledoux. Linear dynamics for the state vector of Markov chain functions. Advances in Applied Probability, 36(4):1198– 1211, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/4140394.

Lee:1979:STN

[Lee79] Alan J. Lee. Sampling theorems for non-stationary random processes. Advances in Applied Probability, 11(2):299–300, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426830.

Lee:1999:OAO

[Lee99a] Duan-Shin Lee. On the optimality and the asymptotic optimality of the smallest weighted available buffer policy. Advances in Applied Probability, 31(4):1118–1150, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428347.

Lee:1999:CLT

[Lee99b] Sungchul Lee. The central limit theorem for Euclidean minimal spanning trees II. Advances in Applied Probability, 31(4): 969–984, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428337.

Lefebvre:1988:UDP

[Lef88] Mario Lefebvre. Utilisation de densités des premier passage en commande optimale stochastique. (French) [Using first-passage densities in stochastic optimal control]. Advances in Applied Probability, 20(1):231–234, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427278.

Lefevre:2007:FCB

[Lef07] Claude Lefèvre. First-crossing and ballot-type results for some nonstationary sequences. Advances in Applied Probability, 39 (2):492–509, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443523.

Lehoczky:1985:IPR

[Leh85] John P. Lehoczky. Inference problems for random-parameter stochastic processes. Advances in Applied Probability, 17(2): 252–253, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427136.

Lehmann:2002:SFP

[Leh02] Axel Lehmann. Smoothness of first passage time distributions and a new integral equation for the first passage time density of continuous Markov processes. Advances in Applied Probability, 34(4):869–887, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428326.

Lemoine:1974:LTG

[Lem74] Austin J. Lemoine. Limit theorems for generalized single server queues. Advances in Applied Probability, 6(1):159–174, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426212.

Lempa:2014:BVC

[Lem14] Jukka Lempa. Bounded variation control of Itô diffusions with exogenously restricted intervention times. Advances in Applied Probability, 46(1):102–120, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1396360105.

Leon:1976:SSV

[Leó76] Jesús Alberto León. Selection and senescence: A variational approach. Advances in Applied Probability, 8(1):17–19, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426015.

Leocard:2009:SSS

[Leo09] Stephanie Leocard. Selective sweep and the size of the hitchhiking set. Advances in Applied Probability, 41(3):731–764, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
27793901.

Lepson:1976:BSI

[Lep76a] Benjamin Lepson. Bounds for skewness and improved moment inequalities for unimodal distributions. Advances in Applied Probability, 8(2):238–239, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425885.

Lepson:1976:SFE

[Lep76b] Benjamin Lepson. Statistical fluctuations of energy losses of charged particles passing through matter. Advances in Applied Probability, 8(2):233–234, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425881.

Letac:1982:FSC

[Let82] Gérard Letac. Les fonctions sphériques d’un couple de Gelfand symétrique et les chaînes de Markov. (French) [Spherical functions of a symmetric Gelfand pair and Markov chains]. Advances in Applied Probability, 14(2):272–294, June 1982. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426521.

Letac:1985:CGD

[Let85] Gérard Letac. A characterization of the gamma distribution. Advances in Applied Probability, 17(4):911–912, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427095.

Leviatan:1977:FMP

[Lev77] T. Leviatan. Functionals of Markov processes. Advances in Applied Probability, 9(2):224–226, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426369.

Leviatan:1978:CMA

[Lev78] Talma Leviatan. Continuation of multiplicative (additive) functionals. Advances in Applied Probability, 10(2):297–299, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426900.

Levy:1979:TRS

[Lev79] Joshua B. Levy. Transience and recurrence of state-dependent branching processes with an immigration component. Advances in Applied Probability, 11(1):73–92, March 1979. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426769. See correction

[Ano79e].

Lewis:1975:MPD

[Lew75] Trevor Lewis. A model for the parasitic disease bilharziasis. Advances in Applied Probability, 7(4):673–704, December 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426396.

Leite:2013:RRT

[LF13] Saul C. Leite and Marcelo D. Fragoso. Reducing response time in fork-join systems under heavy traffic via imbalance control. Advances in Applied Probability, 45(4):1137–1156, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1386857861.

Li:2005:GCR

[LG05] Shuanming Li and José Garrido. On a general class of renewal risk process: Analysis of the Gerber–Shiu function. Advances in Applied Probability, 37(3):836–856, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037358.

Lefevre:2016:SES

[LG16] Claude Lefèvre and Massimiliano Giorgio. SIR epidemics with stages of infection. Advances in Applied Probability, 48(3):768– 791, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1474296314.

Labbe:2007:CDC

[LH07] Chantal Labbé and Andrew J. Heunis. Convex duality in constrained mean-variance portfolio optimization. Advances in Applied Probability, 39(1):77–104, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443501.

Li:1992:NMM

[Li92] Zeng-Hu Li. A note on the multitype measure branching process. Advances in Applied Probability, 24(2):496–498, June 1992.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427702.

Li:1998:IPA

[Li98] Zeng-Hu Li. Immigration processes associated with branching particle systems. Advances in Applied Probability, 30(3):657– 675, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428144.

Li:2009:SES

[Li09] Yuqiang Li. The speed of extinction for some generalized Jiřina processes. Advances in Applied Probability, 41(2):576–599, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/27793891.

Li:2023:OMS

[Li23] Hanwu Li. Optimal multiple stopping problem under nonlinear expectation. Advances in Applied Probability, 55 (1):151–178, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/optimal-multiple-stopping-problem-
under-nonlinear-expectation/AAA744CF0025B8A5EE829706D9E54878.

Liebetrau:1976:TRD

[Lie76] Albert M. Liebetrau. Tests of randomness derived from an estimator of the variance-time curve of a weakly stationary point process. Advances in Applied Probability, 8(2):242, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425889.

Lim:1997:REG

[Lim97] Wei Shi Lim. A rendezvous-evasion game on discrete locations with joint randomization. Advances in Applied Probability, 29 (4):1004–1017, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427851.

Lindgren:1972:WLA

[Lin72] Georg Lindgren. Wave-length and amplitude in Gaussian noise. Advances in Applied Probability, 4(1):81–108, April 1972. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425807.

Lindvall:1974:ASC

[Lin74a] T. Lindvall. Almost sure convergence of Galton–Watson branching processes in varying and random environments. Advances in Applied Probability, 6(2):231, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426269.

Lindvall:1974:LTS

[Lin74b] Torgny Lindvall. Limit theorems for some functionals of certain Galton–Watson branching processes. Advances in Applied Probability, 6(2):309–321, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426295.

Lindvall:1977:PPB

[Lin77] Torgny Lindvall. A probabilistic proof of Blackwell’s renewal theorem. Advances in Applied Probability, 9(2):236, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426381.

Lindgren:1978:PLC

[Lin78] Georg Lindgren. Prediction of level crossings for processes containing non-ergodic components. Advances in Applied Probability, 10(2):290–291, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426893.

Lindgren:1979:PLC

[Lin79] Georg Lindgren. Prediction of level crossings for normal processes containing deterministic components. Advances in Applied Probability, 11(1):93–117, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426770.

Lindgren:1980:EVC

[Lin80] Georg Lindgren. Extreme values and crossings for the χ^{2}-process and other functions of multidimensional Gaussian processes, with reliability applications. Advances in Applied Probability, 12 (3):746–774, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426430.

Lindvall:1992:SCR

[Lin92] T. Lindvall. A simple coupling of renewal processes. Advances in Applied Probability, 24(4):1010–1011, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427724.

Linetsky:2005:TDR

[Lin05] Vadim Linetsky. On the transition densities for reflected diffusions. Advances in Applied Probability, 37(2):435–460, June 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/30037335.

Lindgren:2006:SMS

[Lin06] Georg Lindgren. Slepian models for the stochastic shape of individual Lagrange sea waves. Advances in Applied Probability, 38(2):430–450, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443449.

Lindgren:2010:SDF

[Lin10] G. Lindgren. Slope distribution in front–back asymmetric stochastic Lagrange time waves. Advances in Applied Probability, 42(2):489–508, June 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/25683830.

Lin:2015:NEP

[Lin15] Qian Lin. Nash equilibrium payoffs for stochastic differential games with two reflecting barriers. Advances in Applied Probability, 47(2):355–377, June 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1435236979.

Lipow:1975:BMP

[Lip75] Carla Lipow. A branching model with population size dependence. Advances in Applied Probability, 7(3):495–510, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426124.

Littlewood:1969:PTB

[Lit69a] J. E. Littlewood. On the probability in the tail of a binomial distribution. Advances in Applied Probability, 1(1):43–72, Spring 1969. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426408.

Litwin:1969:DRR

[Lit69b] Samuel Litwin. The distribution of radioactive recovery in randomly cut and sedimented DNA. Journal of Applied Probability, 6(2):275–284, August 1969. CODEN JPRBAM. ISSN 0021-9002 (print), 1475-6072 (electronic). URL http://www.jstor.org/ stable/3211998. See correction [Jof77].

Little:1974:TNR

[Lit74] D. V. Little. A third note on recent research in geometrical probability. Advances in Applied Probability, 6(1):103–130, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426209.

Liu:1990:NCI

[Liu90] Jian Liu. A note on causality and invertibility of a general bilinear time series model. Advances in Applied Probability, 22 (1):247–250, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427608.

Liu:1997:EFS

[Liu97] Quansheng Liu. Sur une équation fonctionnelle et ses applications: une extension du théorème de Kesten–Stigum concernant des processus de branchement. (French) [On a functional equation and its applications: an extension of the theorem of Kesten–Stigum concerning branching processes]. Advances in Applied Probability, 29(2):353–373, June 1997. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428007.

Liu:1998:FPG

[Liu98] Quansheng Liu. Fixed points of a generalized smoothing transformation and applications to the branching random walk. Advances in Applied Probability, 30(1):85–112, March 1998. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427877.

Lalam:2004:EOM

[LJ04] N. Lalam and C. Jacob. Estimation of the offspring mean in a supercritical or near-critical size-dependent branching process. Advances in Applied Probability, 36(2):582–601, June 2004. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428467.

Lalam:2004:MPA

[LJJ04] N. Lalam, C. Jacob, and P. Jagers. Modelling the PCR amplification process by a size-dependent branching process and estimation of the efficiency. Advances in Applied Probability, 36 (2):602–615, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428468.

Liang:1993:SCS

[LK93] Huei-Mei Liang and V. G. Kulkarni. Stability condition for a single-server retrial queue. Advances in Applied Probability, 25 (3):690–701, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427530.

Le:2000:FMS

[LK00] Huiling Le and Alfred Kume. The Fréchet mean shape and the shape of the means. Advances in Applied Probability, 32(1):101– 113, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428229.

Lawrance:1981:NAT

[LL81] A. J. Lawrance and P. A. W. Lewis. A new autoregressive time series model in exponential variables (NEAR(1)). Advances in Applied Probability, 13(4):826–845, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426975.

Lefebvre:1989:FHP

[LL89] Mario Lefebvre and Éric Léonard. On the first hitting place of the integrated Wiener process. Advances in Applied Probability, 21(4):945–948, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427780.

Lam:1991:SRP

[LL91] C. Y. Teresa Lam and John P. Lehoczky. Superposition of renewal processes. Advances in Applied Probability, 23(1):64– 85, March 1991. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427512.

Lee:1997:TPM

[LL97] Shiowjen Lee and J. Lynch. Total positivity of Markov chains and the failure rate character of some first passage times. Advances in Applied Probability, 29(3):713–732, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428083.

Lopes:1998:PES

[LL98] Artur Lopes and Sílvia Lopes. Parametric estimation and spectral analysis of piecewise linear maps of the interval. Advances in Applied Probability, 30(3):757–776, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428150.

Li:2015:RES

[LL15] Xiaoou Li and Jingchen Liu. Rare-event simulation and efficient discretization for the supremum of Gaussian random fields. Advances in Applied Probability, 47(3):787–816, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1444308882.

Liu:2018:EBA

[LL18] Yuanyuan Liu and Wendi Li. Error bounds for augmented truncation approximations of Markov chains via the perturbation method. Advances in Applied Probability, 50(2):645–669, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/error-bounds-for-augmented-truncation-
approximations-of-markov-chains-via-the-perturbation-
method/7FEE4DF31A52C653F02587A5046E6F75.

Lloyd:1973:IGR

[Llo73] Stuart P. Lloyd. An index of genealogical relatedness. Advances in Applied Probability, 5(3):417–438, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425828.

Li:2018:MAT

[LLX18] Xiaoou Li, Jingchen Liu, and Shun Xu. A multilevel approach towards unbiased sampling of random elliptic partial differential equations. Advances in Applied Probability, 50 (4):1007–1031, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/multilevel-approach-towards-unbiased-
sampling-of-random-elliptic-partial-differential-equations/
5ACFB32157F5887F00874AF2F2E7A239.

Lefevre:1990:ALC

[LM90] Claude Lefevre and Xavier Milhaud. On the association of the lifelengths of components subjected to a stochastic environment. Advances in Applied Probability, 22(4):961–964, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427571.

Likhanov:2000:LAL

[LM00] Nikolay Likhanov and Ravi R. Mazumdar. Loss asymptotics in large buffers fed by heterogeneous long-tailed sources. Advances in Applied Probability, 32(4):1168–1189, December 2000. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428522.

Lin:2003:RFP

[LM03] Daming Lin and Viliam Makis. Recursive filters for a partially observable system subject to random failure. Advances in Applied Probability, 35(1):207–227, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428280.

Lin:2004:FPE

[LM04] Daming Lin and Viliam Makis. Filters and parameter estimation for a partially observable system subject to random failure with continuous-range observations. Advances in Applied Probability, 36(4):1212–1230, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/4140395.

Li:2017:TRG

[LM17] Anshui Li and Tobias Müller. On the treewidth of random geometric graphs and percolated grids. Advances in Applied Probability, 49(1):49–60, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-the-treewidth-of-random-geometric-
graphs-and-percolated-grids/E1E962AC607DE38275F963C3CC3A8956.

Lucantoni:1990:SSQ

[LMHN90] David M. Lucantoni, Kathleen S. Meier-Hellstern, and Marcel F. Neuts. A single-server queue with server vacations and a class of non-renewal arrival processes. Advances in Applied Probability, 22(3):676–705, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427464.

Lynch:1983:DOR

[LMP83] James Lynch, Gillian Mimmack, and Frank Proschan. Dispersive ordering results. Advances in Applied Probability, 15(4): 889–891, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427332.

Lopez:2000:SDM

[LMS00] F. Javier López, Servet Martínez, and Gerardo Sanz. Stochastic domination and Markovian couplings. Advances in Applied Probability, 32(4):1064–1076, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428517.

Lopez-Mimbela:1996:CMB

[LMW96] J. Alfredo López-Mimbela and Anton Wakolbinger. Clumping in multitype-branching trees. Advances in Applied Probability, 28(4):1034–1050, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428163.

Lepeltier:2005:RBS

[LMX05] J.-P. Lepeltier, A. Matoussi, and M. Xu. Reflected backward stochastic differential equations under monotonicity and general increasing growth conditions. Advances in Applied Probability, 37(1):134–159, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/30037319.

Lehtonen:1992:SLC

[LN92] T. Lehtonen and H. Nyrhinen. Simulating level-crossing probabilities by importance sampling. Advances in Applied Probability, 24(4):858–874, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427716.

Larsson:2011:EBA

[LN11] Martin Larsson and Johanna Neslehová. Extremal behavior of Archimedean copulas. Advances in Applied Probability, 43(1): 195–216, March 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/23024532.

Leorato:2004:BET

[LO04] S. Leorato and E. Orsingher. Bose–Einstein-type statistics, order statistics and planar random motions with three directions. Advances in Applied Probability, 36(3):937–970, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/4140416.

Loader:1991:LDA

[Loa91] Clive R. Loader. Large-deviation approximations to the distribution of scan statistics. Advances in Applied Probability, 23(4): 751–771, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427674.

Lomnicki:1972:TTS

[Lom72] Z. A. Lomnicki. Two-terminal series-parallel networks. Advances in Applied Probability, 4(1):109–150, April 1972. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425808.

Lomnicki:1978:GSS

[Lom78] A. Lomnicki. Group selection for the self-regulation of population density. Advances in Applied Probability, 10(1):11–14, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426712.

Lorenzen:1979:GSP

[Lor79] Thomas J. Lorenzen. Generalizing the secretary problem. Advances in Applied Probability, 11(2):384–396, June 1979. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426845.

Leorato:2003:AMS

[LOS03] S. Leorato, E. Orsingher, and M. Scavino. An alternating motion with stops and the related planar, cyclic motion with four directions. Advances in Applied Probability, 35(4):1153–1168, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428504.

Lefevre:1990:NSF

[LP90] Claude Lefevre and Philippe Picard. A non-standard family of polynomials and the final size distribution of Reed–Frost epidemic processes. Advances in Applied Probability, 22(1):25– 48, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427595.

Lefevre:1993:USO

[LP93] Claude Lefevre and Philippe Picard. An unusual stochastic order relation with some applications in sampling and epidemic theory. Advances in Applied Probability, 25(1):63–81, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427496.

Lefevre:1996:ATE

[LP96] Claude Lefèvre and Philippe Picard. Abelian-type expansions and non-linear death processes (II). Advances in Applied Probability, 28(3):877–894, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1428185.

Lefevre:1999:AGP

[LP99] Claude Lefèvre and Philippe Picard. Abel-gontcharoff pseudopolynomials and the exact final outcome of SIR epidemic models (III). Advances in Applied Probability, 31(2):532–550, June 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428121.

Li:2021:SNP

[LP21] Bo Li and Guodong Pang. Shot noise processes with randomly delayed cluster arrivals and dependent noises in the large-intensity regime. Advances in Applied Probability, 53 (4):1190–1221, December 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/shot-noise-processes-with-randomly-
delayed-cluster-arrivals-and-dependent-noises-in-the-
largeintensity-regime/13EA17DB664F2B92E8231492E80C87E4.

Leguesdron:1993:TAM

[LPRS93] P. Leguesdron, J. Pellaumail, G. Rubino, and B. Sericola. Transient analysis of the M/M/1 queue. Advances in Applied Probability, 25(3):702–713, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1427531.

Last:2014:MCL

[LPST14] Günter Last, Mathew D. Penrose, Matthias Schulte, and Christoph Thäle. Moments and central limit theorems for some multivariate Poisson functionals. Advances in Applied Probability, 46(2):348–364, June 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1401369698.

Lugannani:1980:SPA

[LR80] Robert Lugannani and Stephen Rice. Saddle point approximation for the distribution of the sum of independent random variables. Advances in Applied Probability, 12(2):475–490, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426607.

Ledoux:1997:SFC

[LR97] James Ledoux and Gerardo Rubino. Simple formulae for counting processes in reliability models. Advances in Applied Probability, 29(4):1018–1038, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427852.

Lachieze-Rey:2011:MPS

[LR11] R. Lachièze-Rey. Mixing properties for STIT tessellations. Advances in Applied Probability, 43(1):40–48, March 2011. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/23024524.

Lin:2012:JVD

[LR12] Kaisheng Lin and Gesine Reinert. Joint vertex degrees in the inhomogeneous random graph model G(\, {√_{⟩|}}). Advances

in Applied Probability, 44(1):139–165, March 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/23207859.

Lanchier:2018:RCS

[LR18] N. Lanchier and S. Reed. The role of cooperation in spatially explicit economical systems. Advances in Applied Probability, 50(3):743–758, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/role-of-cooperation-in-spatially-
explicit-economical-systems/049502661DCF3B3E94897B4E43427C9F.

Leskela:2010:IBP

[LRS10] Lasse Leskelä, Philippe Robert, and Florian Simatos. Interacting branching processes and linear file-sharing networks. Advances in Applied Probability, 42(3):834–854, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/25799070.

Lerche:1989:AEP

[LS89] H. R. Lerche and D. Siegmund. Approximate exit probabilities for a Brownian bridge on a short time interval, and applications. Advances in Applied Probability, 21(1):1–19, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427195.

Li:1993:SMS

[LS93] Haijun Li and Moshe Shaked. Stochastic majorization of stochastically monotone families of random variables. Advances in Applied Probability, 25(4):895–913, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427797.

Last:1996:FCL

[LS96] G. Last and R. Schassberger. A flow conservation law for surface processes. Advances in Applied Probability, 28(1):13–28, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427911.

Last:1998:DSC

[LS98a] G. Last and R. Schassberger. On the distribution of the spherical contact vector of stationary germ-grain models. Advances in Applied Probability, 30(1):36–52, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427874.

Last:1998:AMP

[LS98b] Günter Last and Ryszard Szekli. Asymptotic and monotonicity properties of some repairable systems. Advances in Applied Probability, 30(4):1089–1110, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428031.

Letac:1998:RNT

[LS98c] Gérard Letac and Marco Scarsini. Random nested tetrahedra. Advances in Applied Probability, 30(3):619–627, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428141.

Leipus:2003:RCA

[LS03] Remigijus Leipus and Donatas Surgailis. Random coefficient autoregression, regime switching and long memory. Advances in Applied Probability, 35(3):737–754, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428216.

Last:2011:CAE

[LS11] Günter Last and Ryszard Szekli. Comparisons and asymptotics for empty space hazard functions of germ–grain models. Advances in Applied Probability, 43(4):943–962, December 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/23208033.

Limnios:2013:DTS

[LS13] Nikolaos Limnios and Anatoliy Swishchuk. Discrete-time semi-Markov random evolutions and their applications. Advances in Applied Probability, 45(1):214–240, March 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/41806006.

Lorek:2015:CBS

[LS15] Pawel Lorek and Ryszard Szekli. Computable bounds on the spectral gap for unreliable Jackson networks. Advances in Applied Probability, 47(2):402–424, June 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1435236981.

Lefevre:2021:RPE

[LS21] Claude Lefèvre and Matthieu Simon. Ruin problems for epidemic insurance. Advances in Applied Probability, 53(2):484– 509, June 2021. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL https://www.cambridge.org/
core/journals/advances-in-applied-probability/article/
ruin-problems-for-epidemic-insurance/AC5A6452A66995AD20025904EDCE1A41.

Lindqvist:2016:ESD

[LSH16] Bo H. Lindqvist, Francisco J. Samaniego, and Arne B. Huseby. On the equivalence of systems of different sizes, with applications to system comparisons. Advances in Applied Probability, 48(2):332–348, June 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1465490751.

Lemonnier:2018:SBR

[LSV18] Rémi Lemonnier, Kevin Scaman, and Nicolas Vayatis. Spectral bounds in random graphs applied to spreading phenomena and percolation. Advances in Applied Probability, 50(2):480–503, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/spectral-bounds-in-random-graphs-
applied-to-spreading-phenomena-and-percolation/C8A72B14C2B5081577783AE125D3BF60.

Loris-Teghem:1973:AAS

[LT73] Jacqueline Loris-Teghem. An algebraic approach to some waiting time problems. Advances in Applied Probability, 5(1):16– 17, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425951.

Liu:1994:SSF

[LT94] Zhen Liu and Don Towsley. Stochastic scheduling in inforest networks. Advances in Applied Probability, 26(1):222– 241, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427588.

Ledford:1998:CTB

[LT98] Anthony W. Ledford and Jonathan A. Tawn. Concomitant tail behaviour for extremes. Advances in Applied Probability, 30 (1):197–215, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427884.

Laws:2000:ARF

[LT00] C. N. Laws and Y. C. Teh. Alternative routeing in fully connected queueing networks. Advances in Applied Probability, 32 (4):962–982, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428511.

Ledoux:2001:MBF

[LT01] James Ledoux and Laurent Truffet. Markovian bounds on functions of finite Markov chains. Advances in Applied Probability, 33(2):505–519, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428265.

Lanchier:2015:GBH

[LT15] N. Lanchier and N. Taylor. Galam’s bottom-up hierarchical system and public debate model revisited. Advances in Applied Probability, 47(3):668–692, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1444308877.

Lemaire:2020:TMM

[LTT20] Vincent Lemaire, MichÉle Thieullen, and Nicolas Thomas. Thinning and multilevel Monte Carlo methods for piecewise deterministic (Markov) processes with an application to a stochastic Morris–Lecar model. Advances in Applied Probability, 52(1):138–172, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/thinning-and-multilevel-
monte-carlo-methods-for-piecewise-deterministic-markov-
processes-with-an-application-to-a-stochastic-morrislecar-
model/47F0F47B909929E09727CAB91E10C8FC.

Li:2010:STD

[LTW10] Jinzhu Li, Qihe Tang, and Rong Wu. Subexponential tails of discounted aggregate claims in a time-dependent renewal risk model. Advances in Applied Probability, 42(4):1126–1146, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25799087.

Lukacs:1972:STC

[Luk72] Eugene Lukacs. A survey of the theory of characteristic functions. Advances in Applied Probability, 4(1):1–38, April 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425805.

Lukacs:1977:ST

[Luk77] Eugene Lukacs. Stability theorems. Advances in Applied Probability, 9(2):336–361, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426390.

Luo:2016:SDG

[Luo16] Shangzhen Luo. A stochastic differential game for quadratic– linear diffusion processes. Advances in Applied Probability, 48(4):1161–1182, December 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1482548433.

Ljung:1992:APL

[LW92] Lennart Ljung and Bo Wahlberg. Asymptotic properties of the least-squares method for estimating transfer functions and disturbance spectra. Advances in Applied Probability, 24(2):412– 440, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427698.

Lee:2011:SCT

[LW11] Chihoon Lee and Ananda Weerasinghe. Stationarity and control of a tandem fluid network with fractional Brownian motion input. Advances in Applied Probability, 43(3):847–874, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/23024577.

Liu:2013:DFI

[LX13] Jingchen Liu and Gongjun Xu. On the density functions of integrals of Gaussian random fields. Advances in Applied Probability, 45(2):398–424, June 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1370870124.

Liberman:1977:HOP

[LY77] Varda Liberman and Uri Yechiali. On the hotel overbooking problem: An inventory system with stochastic cancellations. Advances in Applied Probability, 9(2):220, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426364.

Lai:1988:OBP

[LY88] Tze Leung Lai and Zhiliang Ying. Open bandit processes and optimal scheduling of queueing networks. Advances in Applied Probability, 20(2):447–472, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427399.

Lam:1994:CSC

[LY94] C. Teresa Lam and R. H. Yeh. Comparison of sequential and continuous inspection strategies for deteriorating systems. Advances in Applied Probability, 26(2):423–435, June 1994. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427444.

Liu:2012:CRA

[LY12] Jingchen Liu and Xuan Yang. The convergence rate and asymptotic distribution of the bootstrap quantile variance estimator for importance sampling. Advances in Applied Probability, 44 (3):815–841, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/41714077.

Li:2015:PMF

[LY15a] Xiaohu Li and Yinping You. Permutation monotone functions of random vectors with applications in financial and actuarial risk management. Advances in Applied Probability, 47(1):270–291, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1427814591.

Louhichi:2015:EGB

[LY15b] Sana Louhichi and Bernard Ycart. Exponential growth of bifurcating processes with ancestral dependence. Advances in Applied Probability, 47(2):545–564, June 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1435236987.

Lin:2019:OSS

[LY19] Yi-Shen Lin and Yi-Ching Yao. One-sided solutions for optimal stopping problems with logconcave reward functions. Advances in Applied Probability, 51(1):87–115, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/onesided-solutions-for-optimal-
stopping-problems-with-logconcave-reward-functions/9EFABB8657E13FFFD9CB80FDF4702AF0.

Lai:2007:CRW

[LYA07] Tze Leung Lai, Yi-Ching Yao, and Farid Aitsahlia. Corrected random walk approximations to free boundary problems in optimal stopping. Advances in Applied Probability, 39(3):753–775, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443539.

Li:1995:MGQ

[LZ95] Huan Li and Yixin Zhu. On M/G/1 queues with exhaustive service and generalized vacations. Advances in Applied Probability, 27(2):510–531, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427837.

Li:2005:HTA

[LZ05a] Quan-Lin Li and Yiqiang Q. Zhao. Heavy-tailed asymptotics of stationary probability vectors of Markov chains of GI/G/1 type.

Advances in Applied Probability, 37(2):482–509, June 2005. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/30037337.

Li:2005:LTA

[LZ05b] Quan-Lin Li and Yiqiang Q. Zhao. Light-tailed asymptotics of stationary probability vectors of Markov chains of GI/G/1 type. Advances in Applied Probability, 37(4):1075–1093, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
30037374.

Lautensack:2008:RLT

[LZ08] Claudia Lautensack and Sergei Zuyev. Random Laguerre tessellations. Advances in Applied Probability, 40(3):630–650, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443601.

Li:2013:JLT

[LZ13] Bin Li and Xiaowen Zhou. The joint Laplace transforms for diffusion occupation times. Advances in Applied Probability, 45(4):1049–1067, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1386857857.

Last:2017:OZE

[LZ17] Günter Last and Sebastian Ziesche. On the Ornstein– Zernike equation for stationary cluster processes and the random connection model. Advances in Applied Probability, 49 (4):1260–1287, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-the-ornsteinzernike-equation-
for-stationary-cluster-processes-and-the-random-connection-
model/09317DFA3FE1B07BF345D6135927CE02.

Martin:1995:AMG

[MA95a] M. Martin and J. R. Artalejo. Analysis of an M/G/1 queue with two types of impatient units. Advances in Applied Probability, 27 (3):840–861, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428136.

Mytnik:1995:BBD

[MA95b] Leonid Mytnik and Robert J. Adler. Bisexual branching diffusions. Advances in Applied Probability, 27(4):980–1018, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427932.

Ma:2005:STV

[Ma05] Chunsheng Ma. Spatio-temporal variograms and covariance models. Advances in Applied Probability, 37(3):706–725, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
30037351.

Macchi:1975:CAS

[Mac75a] Odile Macchi. The coincidence approach to stochastic point processes. Advances in Applied Probability, 7(1):83–122, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425855.

Macdonald:1975:SAR

[Mac75b] P. D. M. Macdonald. Some asymptotic results for multiphase branching processes. Advances in Applied Probability, 7(2): 251, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426064.

Maejima:1979:NUE

[Mae79] Makoto Maejima. Non-uniform estimates in the local limit theorem for densities. Advances in Applied Probability, 11(2):284– 285, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426815.

Maoui:2007:CDP

[MAF07] Idriss Maoui, Hayriye Ayhan, and Robert D. Foley. Congestiondependent pricing in a stochastic service system. Advances in Applied Probability, 39(4):898–921, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443551.

Mahmoud:2010:GPG

[Mah10] Hosam M. Mahmoud. Gaussian phases in generalized coupon collection. Advances in Applied Probability, 42(4):994–1012, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25799081.

Maisonneuve:1980:TRS

[Mai80] B. Maisonneuve. Topics on regenerative sets and subordinators. Advances in Applied Probability, 12(2):273–274, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426542.

Mairesse:1997:PIR

[Mai97] Jean Mairesse. Products of irreducible random matrices in the (max, +) algebra. Advances in Applied Probability, 29(2):444– 477, June 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428012.

Majewski:2007:LDP

[Maj07] Kurt Majewski. Large deviation properties of constant rate data streams sharing a buffer with long-range dependent traffic in critical loading. Advances in Applied Probability, 39(2):407– 428, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443520.

Malhame:1990:JDM

[Mal90] Roland Malhamé. A jump-driven Markovian electric load model. Advances in Applied Probability, 22(3):564–586, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427458.

Malyshev:1993:NDS

[Mal93] V. A. Malyshev. Networks and dynamical systems. Advances in Applied Probability, 25(1):140–175, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427500.

Malinovskii:2000:PRW

[Mal00] Vsevolod K. Malinovskii. Probabilities of ruin when the safety loading tends to zero. Advances in Applied Probability, 32(3): 885–923, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428418.

Mandl:1974:ECM

[Man74] P. Mandl. Estimation and control in Markov chains. Advances in Applied Probability, 6(1):40–60, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426206.

Mannion:1976:RPI

[Man76] David Mannion. Random packing of an interval. Advances in Applied Probability, 8(3):477–501, September 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426140.

Mannion:1979:RPI

[Man79] David Mannion. Random packing of an interval. II. Advances in Applied Probability, 11(3):591–602, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426956.

Mannion:1988:MCT

[Man88] David Mannion. A Markov chain of triangle shapes. Advances in Applied Probability, 20(2):348–370, June 1988. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427394.

Mannion:1990:CCS

[Man90a] David Mannion. Convergence to collinearity of a sequence of random triangle shapes. Advances in Applied Probability, 22(4): 831–844, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427564.

Mannion:1990:IDS

[Man90b] David Mannion. The invariant distribution of a sequence of random collinear triangle shapes. Advances in Applied Probability, 22(4):845–865, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427565.

Mannion:1994:VTW

[Man94] David Mannion. The volume of a tetrahedron whose vertices are chosen at random in the interior of a parent tetrahedron. Advances in Applied Probability, 26(3):577–596, September 1994.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427809.

Maansson:2007:CBV

[Mån07] Marianne Månsson. A connection between the volume fractions of the stienen model and the dead leaves model. Advances in Applied Probability, 39(1):41–52, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443499.

Manjrekar:2021:SMF

[Man21] Mayank Manjrekar. On spatial matchings: The first-in-first-match case. Advances in Applied Probability, 53(3): 757–800, September 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-spatial-matchings-the-firstinfirstmatch-
case/4B8D918EE1B8E9724947F1BFC47E45F2.

Maruyama:1970:SSM

[Mar70] Takeo Maruyama. Stepping stone models of finite length. Advances in Applied Probability, 2(2):229–258, Autumn 1970. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426319.

Marcus:1975:SED

[Mar75] Allan H. Marcus. Some exact distributions in traffic noise theory. Advances in Applied Probability, 7(3):593–606, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426130.

Mardia:1977:DSG

[Mar77] K. V. Mardia. Distributions on Stiefel and Grassmann manifolds, and their applications. Advances in Applied Probability, 9 (3):435–436, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426097.

Maruyama:1980:OMP

[Mar80] Takeo Maruyama. On an overdominant model of population genetics. Advances in Applied Probability, 12(2):274–275, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426543.

Martyr:2016:DPD

[Mar16] R. Martyr. Dynamic programming for discrete-time finitehorizon optimal switching problems with negative switching costs. Advances in Applied Probability, 48(3):832–847, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1474296317.

Mase:1979:RCC

[Mas79] Shigeru Mase. Random compact convex sets which are infinitely divisible with respect to Minkowski addition. Advances in Applied Probability, 11(4):834–850, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426862.

Masry:1983:NPC

[Mas83] Elias Masry. Non-parametric covariance estimation from irregularly-spaced data. Advances in Applied Probability, 15(1): 113–132, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426985.

Mase:1984:LAN

[Mas84a] Shigeru Mase. Locally asymptotic normality of Gibbs models on a lattice. Advances in Applied Probability, 16(3):585–602, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427288.

Massey:1984:ONQ

[Mas84b] William A. Massey. Open networks of queues: Their algebraic structure and estimating their transient behavior. Advances in Applied Probability, 16(1):176–201, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427230.

Mase:1995:SEP

[Mas95] Shigeru Mase. Stereological estimation of particle size distributions. Advances in Applied Probability, 27(2):350–366, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427830.

Masse:2000:LNV

[Mas00] Bruno Massé. On the LLN for the number of vertices of a random convex Hull. Advances in Applied Probability, 32(3): 675–681, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428407.

Masuyama:2015:EBA

[Mas15] Hiroyuki Masuyama. Error bounds for augmented truncations of discrete-time block-monotone Markov chains under geometric drift conditions. Advances in Applied Probability, 47(1):83–105, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1427814582.

Masuyama:2023:NMI

[Mas23] Hiroyuki Masuyama. A new matrix-infinite-product-form solution for upper block-Hessenberg Markov chains and its quasialgorithmic constructibility. Advances in Applied Probability, 55(1):1–28, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/new-matrixinfiniteproductform-solution-
for-upper-blockhessenberg-markov-chains-and-its-quasialgorithmic-
constructibility/952293123AA6E2F0FA56C8B0530B89EC.

Matheron:1972:EFA

[Mat72] G. Matheron. Ensembles fermés aléatoires, ensembles semimarkoviens et polyèdres poissoniens. (French) [Random closed sets, semi-Markov sets and Poissonian polyhedra]. Advances in Applied Probability, 4(3):508–541, December 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425992.

Matheron:1973:IRF

[Mat73] G. Matheron. The intrinsic random functions and their applications. Advances in Applied Probability, 5(3):439–468, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425829.

Matheron:1974:HPC

[Mat74] G. Matheron. Hyperplans poissoniens et compacts de Steiner. (French) [Poissonian hyperplanes and Steiner compacts]. Advances in Applied Probability, 6(3):563–579, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426234.

Matthes:1977:SHB

[Mat77] K. Matthes. Spatially homogeneous branching processes. Advances in Applied Probability, 9(3):445–446, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426105.

Mattfeldt:1998:SFO

[Mat98] T. Mattfeldt. Scale and first-order effects in the deterministic analysis of stationary isotropic volume processes. Advances in Applied Probability, 30(2):284, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427955.

Mathai:1999:RCP

[Mat99] A. M. Mathai. Random p-content of a p-parallelotope in Euclidean n-space. Advances in Applied Probability, 31(2):343– 354, June 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428113.

Maximov:1977:ETR

[Max77] V. M. Maximov. An ergodic theorem for random products of Euclidean motions. Advances in Applied Probability, 9(3):433, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426094.

Mayo:1978:EST

[May78] O. Mayo. The existence and stability of a three-locus gametophytically-determined self-incompatibility system. Advances in Applied Probability, 10(1):14–15, March 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426713.

Mazars:1989:PMD

[Maz89] N. Mazars. Pseudo-modular decompositions and ’refined bounds’ for the interval reliability and the availability for binary coherent systems. Advances in Applied Probability, 21(1):91– 122, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427199.

Mandjes:2000:OBQ

[MB00] Michel Mandjes and Sem Borst. Overflow behavior in queues with many long-tailed inputs. Advances in Applied Probability, 32(4):1150–1167, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428521.

Moller:2012:TSP

[MB12] Jesper Møller and Kasper K. Berthelsen. Transforming spatial point processes into Poisson processes using random superposition. Advances in Applied Probability, 44(1):42–62, March 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/23207855.

Marquez-Carreras:2009:GFK

[MC09] David Márquez-Carreras. Generalized fractional kinetic equations: Another point of view. Advances in Applied Probability, 41(3):893–910, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/27793907.

McDonald:1976:ERT

[McD76] D. McDonald. Extensions to the renewal theorem. Advances in Applied Probability, 8(2):245, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1425894.

McDonald:1978:LLT

[McD78] D. McDonald. Local limit theorems for sums of integer-valued independent random variables. Advances in Applied Probability, 10(2):288, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426890.

McDiarmid:1981:GPR

[McD81] Colin McDiarmid. General percolation and random graphs. Advances in Applied Probability, 13(1):40–60, March 1981. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426466.

McDiarmid:1983:GFP

[McD83] Colin McDiarmid. General first-passage percolation. Advances in Applied Probability, 15(1):149–161, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426987.

McDonald:1985:IPS

[McD85] D. McDonald. An invariance principle for semi-Markov processes. Advances in Applied Probability, 17(1):100–126, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427055.

McGill:1982:MPD

[McG82] Paul McGill. Markov properties of diffusion local time: A martingale approach. Advances in Applied Probability, 14(4): 789–810, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427024.

McKenzie:1986:AMA

[McK86] Ed McKenzie. Autoregressive moving-average processes with negative-binomial and geometric marginal distributions. Advances in Applied Probability, 18(3):679–705, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427183.

McKenzie:1988:SAM

[McK88] Ed McKenzie. Some ARMA models for dependent sequences of Poisson counts. Advances in Applied Probability, 20(4):822–835, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427362.

McKay:1989:LEB

[McK89] Brendan D. McKay. On Littlewood’s estimate for the binomial distribution. Advances in Applied Probability, 21(2):475– 478, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427172.

McLeish:1975:GMM

[McL75] D. L. McLeish. A generalization of martingales and mixing sequences. Advances in Applied Probability, 7(2):247–249, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426059.

McNeil:1978:I

[McN78] D. R. McNeil. [Introduction]. Advances in Applied Probability, 10(4):703, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426646.

McNamara:1990:PWM

[McN90] J. M. McNamara. The policy which maximises long-term survival of an animal faced with the risks of starvation and predation. Advances in Applied Probability, 22(2):295–308, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427537.

Manfield:1979:AMP

[MD79] D. R. Manfield and T. Downs. An application of multivariate point processes in teletraffic systems. Advances in Applied Probability, 11(2):274, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426803.

Mardia:1989:SDL

[MD89] K. V. Mardia and I. L. Dryden. Shape distributions for Landmark data. Advances in Applied Probability, 21(4):742–755, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427764.

Mardia:1994:SAT

[MD94] K. V. Mardia and I. L. Dryden. Shape averages and their bias. Advances in Applied Probability, 26(2):334–340, June 1994. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427439.

Mead:1975:CSP

[MDB^{+}75] Roger Mead, Peter Diggle, Susannah Brown, Philip Holgate, Peter Clifford, Keith Ord, Howard Ross-Parker, Julian Besag, David Ford, Ronald Martin, Neil Arnason, Denis Mollison, and Eric Renshaw. Conference on spatial processes. Advances in Applied Probability, 7(3):449–460, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426121.

Mecke:1977:CPG

[Mec77] J. Mecke. A characterization of the probability-generating functional for point processes on measurable spaces. Advances in Applied Probability, 9(3):442, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426102.

Mecke:1989:NCP

[Mec89] J. Mecke. A note on Cowan’s process of space division. Advances in Applied Probability, 21(1):235–236, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427209. See [Cow89a].

Mecke:1995:IAP

[Mec95] J. Mecke. Inequalities for the anisotropic Poisson polytope. Advances in Applied Probability, 27(1):56–62, March 1995. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428095.

Mecke:1996:ICM

[Mec96] Joseph Mecke. Inequalities for the cells of mixed stationary Poisson hyperplane tessellations. Advances in Applied Probability, 28(2):338, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428051.

Mecke:1998:IMS

[Mec98a] J. Mecke. Inequalities for mixed stationary Poisson hyperplane tessellations. Advances in Applied Probability, 30(4):921–928, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428022.

Mecke:1998:MAP

[Mec98b] Joseph Mecke. On moments of anisotropic Poisson polygons. Advances in Applied Probability, 30(2):285, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427956.

Mees:1982:CBC

[Mee82] A. I. Mees. Communication in biological cells. Advances in Applied Probability, 14(2):197–198, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426509.

Meidan:1977:GKL

[Mei77] Reuven Meidan. A generalized Karhunen–Loève expansion. Advances in Applied Probability, 9(2):231–232, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426377.

Mejzler:1977:LDE

[Mej77] David Mejzler. Limit distributions for the extreme order statistics. Advances in Applied Probability, 9(2):212–213, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426356.

Melamed:1979:CPT

[Mel79a] Benjamin Melamed. Characterizations of Poisson traffic streams in Jackson queueing networks. Advances in Applied Probability, 11(2):422–438, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426847.

Melamed:1979:PTP

[Mel79b] Benjamin Melamed. On Poisson traffic processes in discretestate Markovian systems with applications to queueing theory. Advances in Applied Probability, 11(1):218–239, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426775. See

correction [Ano79f].

Mellein:1987:NGP

[Mel87] Bernhard Mellein. A nucleation-growth process on the integers. Advances in Applied Probability, 19(1):26–43, March 1987. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427372.

Mergenthaler:1979:SCP

[Mer79] Wolfgang Mergenthaler. A sufficient criterion for particles performing a diffusive motion with state-dependent death rate to die with probability one. Advances in Applied Probability, 11(3): 510–526, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426952.

Mercadier:2006:NBD

[Mer06] Cécile Mercadier. Numerical bounds for the distributions of the maxima of some one- and two-parameter Gaussian processes. Advances in Applied Probability, 38(1):149–170, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/20443433.

Meshalkin:1971:MMC

[Mes71] L. D. Meshalkin. Mathematical methods of chronic diseases study in 1970. Advances in Applied Probability, 3(2):194–196, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426152.

Metivier:1980:PSE

[Met80] Michel Metivier. π^{∗}-processes and stochastic equations. Advances in Applied Probability, 12(2):275–276, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426544.

Meyfroyt:2018:DDT

[Mey18] Thomas M. M. Meyfroyt. Degree-dependent threshold-based random sequential adsorption on random trees. Advances in Applied Probability, 50(1):302–326, March 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/degreedependent-thresholdbased-
random-sequential-adsorption-on-random-trees/7F51AA6B8B1659CD2B604B73BF9B6A9D.

Mardia:1996:DPI

[MGW96] K. V. Mardia, Colin Goodall, and Alistair Walder. Distributions of projective invariants and model-based machine vision. Advances in Applied Probability, 28(3):641–661, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428174.

Miyoshi:1997:SPA

[MH97] Naoto Miyoshi and Toshiharu Hasegawa. Smoothed perturbation analysis for single-server queues with some general service disciplines. Advances in Applied Probability, 29(2):545– 566, June 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428016.

Moller:2008:PDI

[MH08] Jesper Møller and Katerina Helisová. Power diagrams and interaction processes for unions of discs. Advances in Applied Probability, 40(2):321–347, June 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443585.

Mi:1994:BMP

[Mi94] Jie Mi. Burn-in and maintenance policies. Advances in Applied Probability, 26(1):207–221, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427587.

Mihram:1980:SSS

[Mih80] G. Arthur Mihram. Stochastic simulation: Its seed virtually precludes the application of regenerative processes. Advances in Applied Probability, 12(2):317–318, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426599.

Mijnheer:1978:LP

[Mij78] Joop Mijnheer. Limit points of n^{−1/α}S_{n}. Advances in Applied Probability, 10(2):289, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426891.

Mijnheer:1980:APS

[Mij80] Joop Mijnheer. An approximation of partial sums of i.i.d. random variables with infinite variance. Advances in Applied Probability, 12(2):285–286, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426556.

Miles:1969:PFE

[Mil69] R. E. Miles. Poisson flats in Euclidean spaces. Part I: A finite number of random uniform flats. Advances in Applied Probability, 1(2):211–237, Autumn 1969. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426218.

Miles:1971:IRS

[Mil71a] R. E. Miles. Isotropic random simplices. Advances in Applied Probability, 3(2):353–382, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426176.

Miles:1971:PFE

[Mil71b] R. E. Miles. Poisson flats in Euclidean spaces. Part II: Homogeneous Poisson flats and the complementary theorem. Advances in Applied Probability, 3(1):1–43, Spring 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426328.

Milne:1974:IDB

[Mil74] R. K. Milne. Infinitely divisible bivariate Poisson processes. Advances in Applied Probability, 6(2):226–227, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426264.

Miles:1975:DDC

[Mil75] R. E. Miles. Direct derivations of certain surface integral formulae for the mean projections of a convex set. Advances in Applied Probability, 7(4):818–829, December 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426401.

Miles:1980:RTP

[Mil80] R. E. Miles. The random tangential projection of a surface. Advances in Applied Probability, 12(2):425–446, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426604.

Miles:1995:HPL

[Mil95] R. E. Miles. A heuristic proof of a long-standing conjecture of D. G. Kendall concerning the shapes of certain large random polygons. Advances in Applied Probability, 27(2):397–417, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427833.

Miles:1998:LCR

[Mil98] Roger E. Miles. A large class of random tessellations with the classic Poisson polygon distributions. Advances in Applied Probability, 30(2):285, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427957.

Minh:1980:GGQ

[Min80] Do Le Minh. The GI/G/1 queue with uniformly limited virtual waiting times; the finite dam. Advances in Applied Probability, 12(2):501–516, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426609.

Minami:2005:NVG

[Min05] Nariyuki Minami. On the number of vertices with a given degree in a Galton–Watson tree. Advances in Applied Probability, 37 (1):229–264, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/30037323.

Mitra:1987:AAC

[Mit87] Debasis Mitra. Asymptotic analysis and computational methods for a class of simple, circuit-switched networks with blocking. Advances in Applied Probability, 19(1):219–239, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427381.

Mitra:1988:STF

[Mit88] Debasis Mitra. Stochastic theory of a fluid model of producers and consumers coupled by a buffer. Advances in Applied Probability, 20(3):646–676, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1427040.

Miyazawa:1983:DIR

[Miy83] Masakiyo Miyazawa. The derivation of invariance relations in complex queueing systems with stationary inputs. Advances in Applied Probability, 15(4):874–885, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427329.

Miyazawa:1993:IPF

[Miy93] Masakiyo Miyazawa. Insensitivity and product-form decomposability of reallocatable GSMP. Advances in Applied Probability, 25(2):415–437, June 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427660.

Miyoshi:2004:SPS

[Miy04] Naoto Miyoshi. On the subexponential properties in stationary single-server queues: A Palm-martingale approach. Advances in Applied Probability, 36(3):872–892, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/4140413.

Miyazawa:2017:UAL

[Miy17] Masakiyo Miyazawa. A unified approach for large queue asymptotics in a heterogeneous multiserver queue. Advances in Applied Probability, 49(1):182–220, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/unified-approach-for-large-queue-
asymptotics-in-a-heterogeneous-multiserver-queue/5A67D2261345D8698E28FC1A5120A7E1.

Moka:2021:RIS

[MJM21] Sarat Moka, Sandeep Juneja, and Michel Mandjes. Rejectionand importance-sampling-based perfect simulation for Gibbs hard-sphere models. Advances in Applied Probability, 53 (3):839–885, September 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/rejection-and-importancesamplingbased-
perfect-simulation-for-gibbs-hardsphere-models/190AD2BFB1FDAB45B27BE0015293265D.

Mandjes:2001:APT

[MK01] Michel Mandjes and Jeong-Han Kim. Analysis of a phase transition phenomenon in packet networks. Advances in Applied Probability, 33(1):260–280, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428451.

Mollison:1978:CAS

[MKS^{+}78] Denis Mollison, F. P. Kelly, B. W. Silverman, P. J. Green, A. D. Cliff, P. Haggett, B. D. Ripley, C. R. Orton, E. D. Ford, P. A.

Mason, Isobel Clark, D. A. Jones, J. C. Gower, W. G. McGinley, Frank Critchley, Robin Sibson, and D. G. Kendall. Conference on the analysis of spatial data. Advances in Applied Probability, 10(3):491–498, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426631.

Meier:2021:MCA

[MLZ21] Christian Meier, Lingfei Li, and Gongqiu Zhang. Markov chain approximation of one-dimensional sticky diffusions. Advances in Applied Probability, 53(2):335–369, June 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/markov-chain-approximation-
of-onedimensional-sticky-diffusions/56449E6AE5D5E72F7B2F6A290124956B.

Mitra:1983:AEM

[MM83] Debasis Mitra and J. A. Morrison. Asymptotic expansions of moments of the waiting time in closed and open processorsharing systems with multiple job classes. Advances in Applied Probability, 15(4):813–839, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427326.

Morrison:1985:HUA

[MM85] J. A. Morrison and D. Mitra. Heavy-usage asymptotic expansions for the waiting time in closed processor-sharing systems with multiple classes. Advances in Applied Probability, 17(1): 163–185, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427058.

Mackisack:1996:HRT

[MM96a] Margaret S. Mackisack and Roger E. Miles. Homogeneous rectangular tessellations. Advances in Applied Probability, 28(4): 993–1013, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428161.

Miles:1996:FRT

[MM96b] Roger E. Miles and Margaret S. Mackisack. Further random tessellations with the classic Poisson polygon distributions. Advances in Applied Probability, 28(2):338–339, June 1996. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428052.

Mateu:1998:CSS

[MM98] Jorge Mateu and Francisco Montes. A comparative study of simulation methods for marked Gibbs processes. Advances in Applied Probability, 30(2):283–284, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427954.

McCullagh:2006:PP

[MM06] Peter McCullagh and Jesper Møller. The permanental process. Advances in Applied Probability, 38(4):873–888, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/20443479.

McNamara:1991:HDG

[MMC91] J. M. McNamara, S. Merad, and E. J. Collins. The hawk– dove game as an average-cost problem. Advances in Applied Probability, 23(4):667–682, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427669.

Martyr:2022:DTR

[MMP22] Randall Martyr, John Moriarty, and Magnus Perninge. Discretetime risk-aware optimal switching with non-adapted costs. Advances in Applied Probability, 54(2):625–655, June 2022. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/discretetime-
riskaware-optimal-switching-with-nonadapted-costs/87E00698A0D43DCF0365C050F5FD5FC4.

Martinucci:2022:SRT

[MMZ22] Barbara Martinucci, Alessandra Meoli, and Shelemyahu Zacks. Some results on the telegraph process driven by gamma components. Advances in Applied Probability, 54 (3):808–848, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/some-results-on-the-telegraph-process-
driven-by-gamma-components/EBD1A21E96B138EDF54FDE0275B7B069.

Martinez:2014:STT

[MN14] Servet Martínez and Werner Nagel. STIT tessellations have trivial tail σ-algebra. Advances in Applied Probability, 46(3): 643–660, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1409319553.

Martinez:2018:RPP

[MN18] Servet Martínez and Werner Nagel. Regenerative processes for Poisson zero polytopes. Advances in Applied Probability, 50(4):1217–1226, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/regenerative-processes-for-poisson-
zero-polytopes/DE5934750CED54246A4111D1880EE7D1.

Mannersalo:2002:MPS

[MNR02] Petteri Mannersalo, Ilkka Norros, and Rudolf H. Riedi. Multifractal products of stochastic processes: Construction and some basic properties. Advances in Applied Probability, 34(4):888– 903, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428327.

Mecke:2008:IRT

[MNW08] Joseph Mecke, Werner Nagel, and Viola Weiss. The iteration of random tessellations and a construction of a homogeneous process of cell divisions. Advances in Applied Probability, 40 (1):49–59, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443570.

Mohle:1998:CRT

[Möh98a] M. Möhle. Coalescent results for two-sex population models. Advances in Applied Probability, 30(2):513–520, June 1998. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427980.

Mohle:1998:CTM

[Möh98b] M. Möhle. A convergence theorem for Markov chains arising in population genetics and the coalescent with selfing. Advances in Applied Probability, 30(2):493–512, June 1998. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427979.

Mohle:2000:TVD

[Möh00] M. Möhle. Total variation distances and rates of convergence for ancestral coalescent processes in exchangeable population models. Advances in Applied Probability, 32(4):983–993, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428512.

Mohle:2004:TBM

[Möh04] M. Möhle. The time back to the most recent common ancestor in exchangeable population models. Advances in Applied Probability, 36(1):78–97, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428354.

Mohle:2006:NSS

[Möh06] M. Möhle. On the number of segregating sites for populations with large family sizes. Advances in Applied Probability, 38(3): 750–767, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443469.

Molchanov:1995:ERT

[MOK95] Ilya S. Molchanov, Edward Omey, and Eugene Kozarovitzky. An elementary renewal theorem for random compact convex sets. Advances in Applied Probability, 27(4):931–942, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427929.

Mollison:1972:PVS

[Mol72] Denis Mollison. Possible velocities for a simple epidemic. Advances in Applied Probability, 4(2):233–257, August 1972. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425997.

Mollison:1974:PPT

[Mol74] D. Mollison. Percolation processes and tumour growth. Advances in Applied Probability, 6(2):233–235, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426272.

Mollison:1978:MCP

[Mol78] Denis Mollison. Markovian contact processes. Advances in Applied Probability, 10(1):85–108, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426720.

Moller:1989:RTR

[Møl89] J. Møller. Random tessellations in R^{d}. Advances in Applied Probability, 21(1):37–73, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427197.

Moller:1992:RJM

[Møl92] Jesper Møller. Random Johnson-mehl tessellations. Advances in Applied Probability, 24(4):814–844, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427714.

Molchanov:1993:LTC

[Mol93] Ilya S. Molchanov. Limit theorems for convex hulls of random sets. Advances in Applied Probability, 25(2):395–414, June 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427659.

Molchanov:1995:SBM

[Mol95a] Ilya S. Molchanov. Statistics of the Boolean model: From the estimation of means to the estimation of distributions. Advances in Applied Probability, 27(1):63–86, March 1995. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428096.

Moller:1995:GJM

[Møl95b] Jesper Møller. Generation of Johnson-mehl crystals and comparative analysis of models for random nucleation. Advances in Applied Probability, 27(2):367–383, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427831.

Moller:2003:SNC

[Møl03] Jesper Møller. Shot noise Cox processes. Advances in Applied Probability, 35(3):614–640, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428211.

Montes:1995:SIA

[Mon95] Paco Montes. Stereology and image analysis: In honour of Professor L. A. Santaló. Advances in Applied Probability, 27(1): 1, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428090.

Mondal:2017:ZST

[Mon17] Prasenjit Mondal. On zero-sum two-person undiscounted semi-Markov games with a multichain structure. Advances in Applied Probability, 49(3):826–849, September 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/on-
zerosum-twoperson-undiscounted-semimarkov-games-with-
a-multichain-structure/F81F4653FCBF1C21363ADD7597399A11.

Moore:1979:SCL

[Moo79] John B. Moore. On strong consistency of least squares identification algorithms. Advances in Applied Probability, 11(2): 302, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426832.

Moran:1969:SNR

[Mor69] P. A. P. Moran. A second note on recent research in geometrical probability. Advances in Applied Probability, 1(1):73–89, Spring 1969. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426409.

Morton:1973:OCS

[Mor73] Richard Morton. Optimal control of stationary Markov processes. Advances in Applied Probability, 5(1):18–19, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425952.

Morgan:1976:SMG

[Mor76] Byron J. T. Morgan. Stochastic models of grouping changes. Advances in Applied Probability, 8(1):30–57, March 1976. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426021.

Morrison:1994:LPS

[Mor94] J. A. Morrison. Loss probabilities in a simple circuit-switched network. Advances in Applied Probability, 26(2):456–473, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427446.

Moyal:1972:PPN

[Moy72] J. E. Moyal. Particle populations and number operators in quantum theory. Advances in Applied Probability, 4(1):39–80, April 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425806.

Michel:2003:LCB

[MP03] Julien Michel and Katy Paroux. Local convergence of the Boolean Shell model towards the thick Poisson hyperplane process in the Euclidean space. Advances in Applied Probability, 35 (2):354–361, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428425.

McSweeney:2008:ECT

[MP08] John K. McSweeney and Boris G. Pittel. Expected coalescence time for a nonuniform allocation process. Advances in Applied Probability, 40(4):1002–1032, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443622.

McVinish:2010:LLM

[MP10] R. McVinish and P. K. Pollett. Limits of large metapopulations with patch-dependent extinction probabilities. Advances in Applied Probability, 42(4):1172–1186, December 2010. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/25799089.

Medvedev:2018:SNM

[MP18] Alexey Medvedev and Gábor Pete. Speeding up non-Markovian first-passage percolation with a few extra edges. Advances in Applied Probability, 50(3):858–886, September 2018. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/speeding-up-
nonmarkovian-firstpassage-percolation-with-a-few-extra-
edges/8DCAFD32B1521747D3BCE57ADF0A1D7A.

Martinez:1996:SSS

[MPB96] Vicent J. Martínez and María Jesús Pons-Bordería. Statistical studies of the spatial distribution of galaxies. Advances in Applied Probability, 28(2):337–338, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428050.

Makri:2007:SRS

[MPP07] Frosso S. Makri, Andreas N. Philippou, and Zaharias M. Psillakis. Success run statistics defined on an urn model. Advances in Applied Probability, 39(4):991–1019, December 2007. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443555.

Martinez:1998:DAQ

[MPS98] Servet Martinez, Pierre Picco, and Jaime San Martín. Domain of attraction of quasi-stationary distributions for the Brownian motion with drift. Advances in Applied Probability, 30(2):385– 408, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427974.

Maansson:2002:RPN

[MR02] Marianne Månsson and Mats Rudemo. Random patterns of nonoverlapping convex grains. Advances in Applied Probability, 34(4):718–738, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428319.

Moller:2005:PSH

[MR05] Jesper Møller and Jakob G. Rasmussen. Perfect simulation of Hawkes processes. Advances in Applied Probability, 37(3): 629–646, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/30037347.

Mitra:2009:ARV

[MR09] Abhimanyu Mitra and Sidney I. Resnick. Aggregation of rapidly varying risks and asymptotic independence. Advances in Applied Probability, 41(3):797–828, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/27793903.

Martirosyan:2020:ESL

[MR20]  Davit Martirosyan and Philippe Robert.  The equilib-

rium states of large networks of Erlang queues. Advances in Applied Probability, 52(2):617–654, June 2020. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/equilibrium-
states-of-large-networks-of-erlang-queues/95FB887084169030EDD4DBA09E8F3DC2.

Mrkvicka:2001:EVP

[Mrk01] Tomás Mrkvicka. Estimation variances for Poisson processes of compact sets. Advances in Applied Probability, 33(4):765–772, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428304.

Mitra:1986:CFT

[MRSV86] Debasis Mitra, Fabio Romeo, and Alberto Sangiovanni-Vincentelli. Convergence and finite-time behavior of simulated annealing. Advances in Applied Probability, 18(3):747–771, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427186.

Menshikov:2002:LTI

[MRV02] M. V. Menshikov, K. A. Rybnikov, and S. E. Volkov. The loss of tension in an infinite membrane with holes distributed according to a Poisson law. Advances in Applied Probability, 34(2):292– 312, June 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428289.

Morissette:1974:HWL

[MS74] Jean Morissette and Cedric A. B. Smith. The Hardy–Weinberg law and inbreeding. Advances in Applied Probability, 6(1):15– 17, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426202.

Miller:1976:NEW

[MS76] Douglas R. Miller and F. Dennis Sentilles. A note on the extension and weak convergence of some probability measures. Advances in Applied Probability, 8(2):226, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425874.

Marshall:1980:UNP

[MS80a] Albert W. Marshall and Moshe Shaked. Univariate NBU processes. Advances in Applied Probability, 12(2):312–313, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426590.

Mazziotto:1980:FTP

[MS80b] G. Mazziotto and J. Szpirglas. Filtering for two-parameter doubly stochastic Poisson measure. Advances in Applied Probability, 12(2):294, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426566.

Marshall:1983:NBT

[MS83] Albert W. Marshall and Moshe Shaked. New better than used processes. Advances in Applied Probability, 15(3):601–615, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426621.

Masuda:1987:ACP

[MS87] Yasushi Masuda and Ushio Sumita. Analysis of a counting process associated with a semi-Markov process: Number of entries into a subset of state space. Advances in Applied Probability, 19 (4):767–783, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427101.

Meester:1990:CTT

[MS90] Ludolf E. Meester and J. George Shanthikumar. Concavity of the throughput of tandem queueing systems with finite buffer storage space. Advances in Applied Probability, 22(3):764–767, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427472.

Molchanov:1994:APE

[MS94] Ilya Molchanov and Dietrich Stoyan. Asymptotic properties of estimators for parameters of the Boolean model. Advances in Applied Probability, 26(2):301–323, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427437.

Mecke:2001:SCN

[MS01a] Joseph Mecke and Dietrich Stoyan. The specific connectivity number of random networks. Advances in Applied Probability, 33(3):576–583, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428475.

Morato:2001:SAC

[MS01b] Laura M. Morato and Paola Siri. A stochastic algorithm to compute optimal probabilities in the chaos game. Advances in Applied Probability, 33(2):423–436, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428261.

Maier:2003:SIT

[MS03a] Roland Maier and Volker Schmidt. Stationary iterated tessellations. Advances in Applied Probability, 35(2):337–353, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428424.

Molchanov:2003:CWS

[MS03b] Ilya Molchanov and Vadim Scherbakov. Coverage of the whole space. Advances in Applied Probability, 35(4):898–912, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428491.

Morato:2004:ASA

[MS04] Laura M. Morato and Paola Siri. Addendum to ’a stochastic algorithm to compute optimal probabilities in the chaos game’: A new convergence criterion. Advances in Applied Probability, 36 (1):318–323, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428364.

Moller:2010:TSP

[MS10] Jesper Møller and Frederic Paik Schoenberg. Thinning spatial point processes into Poisson processes. Advances in Applied Probability, 42(2):347–358, June 2010. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/25683823.

Molchanov:2011:ETP

[MS11a] Ilya Molchanov and Michael Schmutz. Exchangeability-type properties of asset prices. Advances in Applied Probability, 43 (3):666–687, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/23024570.

Moser:2011:TBM

[MS11b] Martin Moser and Robert Stelzer. Tail behavior of multivariate Lévy–driven mixed moving average processes and SupOU stochastic volatility models. Advances in Applied Probability, 43(4):1109–1135, December 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/23208041.

Mahmoud:2012:JBT

[MS12] Hosam M. Mahmoud and Robert T. Smythe. On the joint behavior of types of coupons in generalized coupon collection. Advances in Applied Probability, 44(2):429–451, June 2012. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/41714058.

Miyoshi:2014:CNM

[MS14] Naoto Miyoshi and Tomoyuki Shirai. A cellular network model with Ginibre configured base stations. Advances in Applied Probability, 46(3):832–845, September 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1409319562.

Milstein:2015:UAC

[MS15] Grigori N. Milstein and John Schoenmakers. Uniform approximation of the Cox–Ingersoll–Ross process. Advances in Applied Probability, 47(4):1132–1156, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1449859803.

Milstein:2016:UAC

[MS16] Grigori N. Milstein and John Schoenmakers. Uniform approximation of the Cox–Ingersoll–Ross process via exact simulation at random times. Advances in Applied Probability, 48(4):1095– 1116, December 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1482548430.

Maller:2017:STA

[MS17] Ross A. Maller and Peter C. Schmidli. Small-time almostsure behaviour of extremal processes. Advances in Applied Probability, 49(2):411–429, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/smalltime-almostsure-behaviour-of-
extremal-processes/215A7F4FCD3A9D04F02D1240C9A8E108.

Muller:2019:DKR

[MS19] Tobias Müller and Merlijn Staps. The diameter of KP-KVB random graphs. Advances in Applied Probability, 51(2):358–377, June 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/diameter-of-kpkvb-random-graphs/
DCC6764E1CB87317BF1110DE2773C924.

Manoharan:1992:PPT

[MSM92] M. Manoharan, Harshinder Singh, and Neeraj Misra. Preservation of phase-type distributions under Poisson shock models. Advances in Applied Probability, 24(1):223–225, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427739.

Miyazawa:1995:SIG

[MSS95] Masakiyo Miyazawa, Rolf Schassberger, and Volker Schmidt. On the structure of an insensitive generalized semi-Markov process with reallocation and point-process input. Advances in Applied Probability, 27(1):203–225, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428104.

Mendivil:2001:RSA

[MSS01] F. Mendivil, R. Shonkwiler, and M. C. Spruill. Restarting search algorithms with applications to simulated annealing. Advances in Applied Probability, 33(1):242–259, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428450.

Menshikov:2001:NTD

[MSV01] M. Menshikov, V. Sidoravicius, and M. Vachkovskaia. A note on two-dimensional truncated long-range percolation. Advances in Applied Probability, 33(4):912–929, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428312.

Meyn:1992:SMP

[MT92] Sean P. Meyn and R. L. Tweedie. Stability of Markovian processes i: Criteria for discrete-time chains. Advances in Applied Probability, 24(3):542–574, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427479.

Meyn:1993:SMPa

[MT93a] Sean P. Meyn and R. L. Tweedie. Stability of Markovian processes II: Continuous-time processes and sampled chains. Advances in Applied Probability, 25(3):487–517, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427521.

Meyn:1993:SMPb

[MT93b] Sean P. Meyn and R. L. Tweedie. Stability of Markovian processes III: Foster–Lyapunov criteria for continuous-time processes. Advances in Applied Probability, 25(3):518–548, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427522.

Miyazawa:1997:GPF

[MT97] Masakiyo Miyazawa and Peter G. Taylor. A geometric productform distribution for a queueing network with non-standard batch arrivals and batch transfers. Advances in Applied Probability, 29(2):523–544, June 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428015.

Moller:2005:GSN

[MT05] Jesper Møller and Giovanni Luca Torrisi. Generalised shot noise Cox processes. Advances in Applied Probability, 37(1):48– 74, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
30037315.

Meester:2006:EBP

[MT06a] Ronald Meester and Pieter Trapman. Estimation in branching processes with restricted observations. Advances in Applied Probability, 38(4):1098–1115, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443489.

Motyer:2006:DRQ

[MT06b] Allan J. Motyer and Peter G. Taylor. Decay rates for quasibirth-and-death processes with countably many phases and tridiagonal block generators. Advances in Applied Probability, 38(2):522–544, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443454.

Milstein:2009:MCM

[MT09] G. N. Milstein and M. V. Tretyakov. Monte Carlo methods for backward equations in nonlinear filtering. Advances in Applied Probability, 41(1):63–100, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/27793865.

Meester:2011:BBC

[MT11] Ronald Meester and Pieter Trapman. Bounding basic characteristics of spatial epidemics with a new percolation model. Advances in Applied Probability, 43(2):335–347, June 2011. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/23024509.

Milstein:2013:PMI

[MT13] G. N. Milstein and M. V. Tretyakov. Probabilistic methods for the incompressible Navier–Stokes equations with space periodic conditions. Advances in Applied Probability, 45(3):742– 772, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1377868537.

Maillardet:2016:QAR

[MT16] R. J. Maillardet and P. G. Taylor. Queues with advanced reservations: an infinite-server proxy for the bookings diary. Advances in Applied Probability, 48(1):13–31, March 2016. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1457466153.

Muche:2005:PVT

[Muc05] L. Muche. The Poisson–Voronoi tessellation: Relationships for edges. Advances in Applied Probability, 37(2):279–296, June 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/30037328.

Muche:2010:CCL

[Muc10] L. Muche. Contact and chord length distribution functions of the Poisson–Voronoi tessellation in high dimensions. Advances in Applied Probability, 42(1):48–68, March 2010. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/25683806.

Mukherjea:1975:RWS

[Muk75] A. Mukherjea. Random walks on semigroups. Advances in Applied Probability, 7(2):255–256, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426070.

Muller:1997:IPM

[Mül97a] Alfred Müller. Integral probability metrics and their generating classes of functions. Advances in Applied Probability, 29(2):429– 443, June 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428011.

Muller:1997:SOG

[Mül97b] Alfred Müller. Stochastic orders generated by integrals: A unified study. Advances in Applied Probability, 29(2):414–428, June 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428010.

Mijatovic:2019:AVR

[MV19] Aleksandar Mijatović and Jure Vogrinc. Asymptotic variance for random walk Metropolis chains in high dimensions: logarithmic growth via the Poisson equation. Advances in Applied Probability, 51(4):994–1026, December 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/asymptotic-variance-for-
random-walk-metropolis-chains-in-high-dimensions-logarithmic-
growth-via-the-poisson-equation/DC80892CE3ADA7B8537E5F463FD899C1.

Mitov:1984:CTB

[MVY84] K. V. Mitov, V. A. Vatutin, and N. M. Yanev. Continuous-time branching processes with decreasing state-dependent immigration. Advances in Applied Probability, 16(4):697–714, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427337.

Matoglu:2019:EAC

[MVY19] Melda Ormeci Matoglu, John H. Vande Vate, and Haiyue Yu. The economic average cost Brownian control problem. Advances in Applied Probability, 51(1):300–337, March 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/economic-average-cost-brownian-
control-problem/B49015F5449D26684AA35011D21DDCBC.

Milne:1972:FRG

[MW72] R. K. Milne and M. Westcott. Further results for Gauss–Poisson processes. Advances in Applied Probability, 4(1):151–176, April 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425809.

Morton:1974:OCD

[MW74] R. Morton and K. H. Wickwire. On the optimal control of a deterministic epidemic. Advances in Applied Probability, 6(4): 622–635, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426183.

McDonald:1978:IPS

[MW78] J. N. McDonald and N. A. Weiss. An infinite particle system with continual input and random death. Advances in Applied Probability, 10(4):764–787, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426658.

Markus:1993:SNL

[MW93] Lawrence Markus and Ananda Weerasinghe. Stochastic nonlinear oscillators. Advances in Applied Probability, 25(3):649– 666, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427528.

Mardia:1994:SSD

[MW94] K. V. Mardia and A. N. Walder. Size-and-shape distributions for paired Landmark data. Advances in Applied Probability, 26(4): 893–905, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427896.

Menshikov:1996:PTM

[MW96a] M. Menshikov and R. J. Williams. Passage-time moments for continuous non-negative stochastic processes and applications. Advances in Applied Probability, 28(3):747–762, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428179.

Miyazawa:1996:SQB

[MW96b] Masakiyo Miyazawa and Ronald W. Wolff. Symmetric queues with batch departures and their networks. Advances in Applied Probability, 28(1):308–326, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427923.

Moller:1998:MCC

[MW98] Jesper Møller and Rasmus Plenge Waagepetersen. Markov connected component fields. Advances in Applied Probability, 30(1): 1–35, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427873.

Meyer:2021:SRV

[MW21]  Nicolas Meyer and Olivier Wintenberger.  Sparse reg-

ular variation. Advances in Applied Probability, 53(4): 1115–1148, December 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/sparse-regular-variation/24B700AD0F14B49051A2C60CC6A6CEBA.

Mendelson:1980:PMO

[MY80] Haim Mendelson and Uri Yechiali. Performance measures for ordered lists in random-access files. Advances in Applied Probability, 12(2):308, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426585.

Mitov:1989:BHB

[MY89] K. V. Mitov and N. M. Yanev. Bellman–Harris branching processes with a special type of state-dependent immigration. Advances in Applied Probability, 21(2):270–283, June 1989. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427160.

McGivney:1999:AGL

[MY99] K. McGivney and J. E. Yukich. Asymptotics for geometric location problems over random samples. Advances in Applied Probability, 31(3):632–642, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1428388.

Martinez:2001:DRC

[MY01a] Servet Martínez and Bernard Ycart. Decay rates and cutoff for convergence and hitting times of Markov chains with countably infinite state space. Advances in Applied Probability, 33(1):188– 205, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428447.

Mitov:2001:LTA

[MY01b] K. V. Mitov and N. M. Yanev. Limit theorems for alternating renewal processes in the infinite mean case. Advances in Applied Probability, 33(4):896–911, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428311.

Matsumoto:2003:DRB

[MY03] Hiroyuki Matsumoto and Marc Yor. On Dufresne’s relation between the probability laws of exponential functionals of Brownian motions with different drifts. Advances in Applied Probability, 35(1):184–206, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428279.

Moller:2011:PRM

[MY11] Jesper Møller and Man Lung Yiu. Probabilistic results for a mobile service scenario. Advances in Applied Probability, 43 (2):322–334, June 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/23024508.

Mikosch:2020:GFC

[MY20] Thomas Mikosch and Jorge Yslas. Gumbel and Fréchet convergence of the maxima of independent random walks. Advances in Applied Probability, 52(1):213–236, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/gumbel-and-frechet-convergence-
of-the-maxima-of-independent-random-walks/11F786EB1A494461E6D32146EA5F4BFD.

Mitov:2018:MBP

[MYH18] Kosto V. Mitov, Nikolay M. Yanev, and Ollivier Hyrien. Multitype branching processes with inhomogeneous Poisson immigration. Advances in Applied Probability, 50(A): 211–228, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/multitype-branching-processes-with-
inhomogeneous-poisson-immigration/AD80F97EF91FA00358B06D300A04512D.

Myllari:2002:LDN

[Myl02] T. Mylläri. Limit distributions for the number of leaves in a random forest. Advances in Applied Probability, 34(4):904–922, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428328.

Mertens:1977:MVB

[MZ77] Jean-François Mertens and Shmuel Zamir. The maximal variation of a bounded martingale. Advances in Applied Probability, 9 (2):228, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426372.

Moller:1996:GTR

[MZ96] Jesper Møller and Sergei Zuyev. Gamma-type results and other related properties of Poisson processes. Advances in Applied Probability, 28(3):662–673, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428175.

Molchanov:1998:VAF

[MZ98] Ilya Molchanov and Sergei Zuyev. Variational analysis of functionals of a Poisson process. Advances in Applied Probability, 30 (2):286, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427958.

Miyazawa:2004:STA

[MZ04] Masakiyo Miyazawa and Yiqiang Q. Zhao. The stationary tail asymptotics in the GI/G/1-Type queue with countably many background states. Advances in Applied Probability, 36(4):1231– 1251, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/4140396.

Mariani:2016:LDE

[MZ16] Mauro Mariani and Lorenzo Zambotti. Large deviations for the empirical measure of heavy-tailed Markov renewal processes. Advances in Applied Probability, 48(3):648–671, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1474296308.

Muravlev:2020:BST

[MZ20] Alexey Muravlev and Mikhail Zhitlukhin. A Bayesian sequential test for the drift of a fractional Brownian motion. Advances in Applied Probability, 52(4):1308–1324, December 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/bayesian-sequential-test-
for-the-drift-of-a-fractional-brownian-motion/8E7BD0333258FDFA38C739C532963D35.

Nagaev:1971:AMP

[Nag71] A. V. Nagaev. Asymptotic methods for problems in the mathematical theory of epidemics. Advances in Applied Probability, 3 (2):222–223, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426165.

Nain:1986:GPR

[Nai86] Philippe Nain. On a generalization of the preemptive resume priority. Advances in Applied Probability, 18(1):255–273, March 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427245.

Nakata:1979:ERT

[Nak79] Tomoichi Nakata. On an estimate of the remainder term in the central limit theorem. Advances in Applied Probability, 11 (2):285–286, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426816.

Nakayama:1996:GCB

[Nak96] Marvin K. Nakayama. General conditions for bounded relative error in simulations of highly reliable Markovian systems. Advances in Applied Probability, 28(3):687–727, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428177.

Nakata:2014:NCO

[Nak14] Toshio Nakata. The number of collisions for the occupancy problem with unequal probabilities. Advances in Applied Probability, 46(1):168–185, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1396360108.

Narita:1987:AAI

[Nar87] Kiyomasa Narita. Asymptotic analysis for interactive oscillators of the van der Pol type. Advances in Applied Probability, 19(1): 44–80, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427373.

Narita:1991:ABV

[Nar91a] Kiyomasa Narita. Asymptotic behavior of velocity process in the Smoluchowski–Kramers approximation for stochastic differential equations. Advances in Applied Probability, 23(2):317– 326, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427751.

Narita:1991:SKA

[Nar91b] Kiyomasa Narita. The Smoluchowski–Kramers approximation for the stochastic Liénard equation with mean-field. Advances in Applied Probability, 23(2):303–316, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427750.

Naasell:1996:QSD

[Nås96] Ingemar Nåsell. The quasi-stationary distribution of the closed endemic SIS model. Advances in Applied Probability, 28(3): 895–932, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428186.

Natvig:1975:IOP

[Nat75] Bent Natvig. On the input and output processes for a general birth-and-death queueing model. Advances in Applied Probability, 7(3):576–592, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426129.

Natvig:1980:IBA

[Nat80] B. Natvig. Improved bounds for the availability and unavailability in a fixed time interval for systems of maintained, interdependent components. Advances in Applied Probability, 12 (1):200–221, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426502.

Natvig:1982:TSH

[Nat82] B. Natvig. Two suggestions of how to define a multistate coherent system. Advances in Applied Probability, 14(2):434– 455, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1426529. See correction [Ano83f].

Natvig:1986:IUB

[Nat86] Bent Natvig. Improved upper bounds for the availabilities in a fixed time interval for multistate monotone systems. Advances in Applied Probability, 18(2):577–579, June 1986. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427314.

Natvig:1989:CSS

[Nat89] Bent Natvig. A comment on the steady-state performance of multistate monotone systems. Advances in Applied Probability, 21(1):237, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427210.

Naveau:2003:ASR

[Nav03] Philippe Naveau. Almost sure relative stability of the maximum of a stationary sequence. Advances in Applied Probability, 35(3): 721–736, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428215.

Nawijn:1991:RIG

[Naw91] Wim M. Nawijn. On a random interval graph and the maximum throughput rate in the system GI/G/1/0. Advances in Applied Probability, 23(4):945–956, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427685.

Nagel:2015:CCC

[NB15] Werner Nagel and Eike Biehler. Consistency of constructions for cell division processes. Advances in Applied Probability, 47(3): 640–651, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1444308875.

Neal:2006:SDA

[Nea06] Peter Neal. Stochastic and deterministic analysis of SIS household epidemics. Advances in Applied Probability, 38(4):943–968, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

20443483. See correction [TF08].

Neal:2014:EBS

[Nea14] Peter Neal. Endemic behaviour of SIS epidemics with general infectious period distributions. Advances in Applied Probability, 46(1):241–255, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1396360112.

Neck:1978:SEA

[Nec78] R. Neck. On some economic applications of stochastic control theory. Advances in Applied Probability, 10(2):327, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426926.

Nerman:2018:PJC

[Ner18] Olle Nerman. Peter Jagers: a collegial perspective. Advances in Applied Probability, 50(A):3–9, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/peter-jagers-a-collegial-
perspective/E6EC166CFF4DD9723CF80AB1FB5DE38C.

Nestoridi:2017:NLR

[Nes17] Evita Nestoridi. A non-local random walk on the hypercube. Advances in Applied Probability, 49(4):1288–1299, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/nonlocal-random-
walk-on-the-hypercube/F21DD5B2A81AB565B32EC66BFB2396D2.

Neuts:1970:TSS

[Neu70] Marcel F. Neuts. Two servers in series, studied in terms of a Markov renewal branching process. Advances in Applied Probability, 2(1):110–149, Spring 1970. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/3518346.

Neuts:1971:QSE

[Neu71] Marcel F. Neuts. A queue subject to extraneous phase changes. Advances in Applied Probability, 3(1):78–119, Spring 1971. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426330.

Neuts:1975:ASQ

[Neu75] Marcel F. Neuts. Algorithmic solutions of queueing problems. Advances in Applied Probability, 7(2):238–239, June 1975. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426052.

Neuts:1976:AWT

[Neu76a] Marcel F. Neuts. Algorithms for waiting-time distributions under various queue disciplines in the M/G/1 queue with a servicetime distribution of phase type. Advances in Applied Probability, 8(2):252, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425901.

Neuts:1976:MFM

[Neu76b] Marcel F. Neuts. Moment formulas for the Markov renewal branching process. Advances in Applied Probability, 8(4):690– 711, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425930.

Neuts:1977:ASS

[Neu77a] Marcel F. Neuts. Algorithms for the steady-state behavior of a queueing model involving several types of customers and changeover times. Advances in Applied Probability, 9(2):201–202, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426348.

Neuts:1977:MGQ

[Neu77b] Marcel F. Neuts. The M/G/1 queue with several types of customers and change-over times. Advances in Applied Probability, 9(3):604–644, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426117.

Neuts:1977:SEF

[Neu77c] Marcel F. Neuts. Some explicit formulas for the steady-state behavior of the queue with semi-Markovian service times. Advances in Applied Probability, 9(1):141–157, March 1977. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425821.

Neuts:1978:MCA

[Neu78] Marcel F. Neuts. Markov chains with applications in queueing theory, which have a matrix-geometric invariant probability vector. Advances in Applied Probability, 10(1):185–212, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426725.

Neuts:1986:CCC

[Neu86a] Marcel F. Neuts. The caudal characteristic curve of queues. Advances in Applied Probability, 18(1):221–254, March 1986. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427244.

Neuts:1986:GPK

[Neu86b] Marcel F. Neuts. Generalizations of the Pollaczek–Khinchin integral equation in the theory of queues. Advances in Applied Probability, 18(4):952–990, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427258.

Neuts:1986:NIE

[Neu86c] Marcel F. Neuts. A new informative embedded Markov renewal process for the PH/G/1 queue. Advances in Applied Probability, 18(2):533–557, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427311.

Neuts:1987:NRP

[Neu87] Marcel F. Neuts. A non-renewal process with renewal counting distributions. Advances in Applied Probability, 19(1):287– 288, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427384.

Ney:1974:ADB

[Ney74a] P. Ney. Age-dependent branching processes. Advances in Applied Probability, 6(2):200–202, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426246.

Ney:1974:CBP

[Ney74b] Peter Ney. Critical branching processes. Advances in Applied Probability, 6(3):434–445, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426226.

Nash:1977:HAO

[NG77] P. Nash and J. C. Gittins. A Hamiltonian approach to optimal stochastic resource allocation. Advances in Applied Probability, 9(1):55–68, March 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425816.

Nicolis:1982:DAL

[Nic82] J. S. Nicolis. A dynamic approach to language. Advances in Applied Probability, 14(2):198–199, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426510.

Nielsen:1998:DVR

[Nie98] Jørgen Nielsen. The distribution of volume reduction induced by isotropic projections. Advances in Applied Probability, 30 (2):287–288, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427960.

Nielsen:1999:DVR

[Nie99] Jørgen Nielsen. The distribution of volume reductions induced by isotropic random projections. Advances in Applied Probability, 31(4):985–994, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428338.

Nijst:1974:PSM

[Nij74] A. G. P. M. Nijst. Propagation of sufficiency and minimal sufficiency in stochastic processes. Advances in Applied Probability, 6(2):245–247, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426281.

Nguyen:2018:TFA

[NLM18] Thi Ngoc Minh Nguyen, Sylvain Le Corff, and Eric Moulines. On the two-filter approximations of marginal smoothing distributions in general state-space models. Advances in Applied Probability, 50(1):154–177, March 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-the-twofilter-approximations-
of-marginal-smoothing-distributions-in-general-statespace-
models/712EEC6894348BED3FBE21B3DABF95A0.

Nagasawa:1979:ATR

[NM79] Masao Nagasawa and Takeo Maruyama. An application of time reversal of Markov processes to a problem of population genetics. Advances in Applied Probability, 11(3):457–478, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426949.

Nino-Mora:2001:RBP

[NM01] José Niño-Mora. Restless bandits, partial conservation laws and indexability. Advances in Applied Probability, 33(1):76– 98, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428442.

Norman:1975:LTS

[Nor75] M. Frank Norman. Limit theorems for stationary distributions. Advances in Applied Probability, 7(3):561–575, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426128.

Norden:1982:DTE

[Nor82] R. H. Norden. On the distribution of the time to extinction in the stochastic logistic population model. Advances in Applied Probability, 14(4):687–708, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427019.

Nosko:1994:HRC

[Nos94] Vladimir P. Nosko. The horizon of the random cone field under a trend. Advances in Applied Probability, 26(3):597–615, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427810.

Nowicki:1988:APD

[Now88] Krzysztof Nowicki. Asymptotic Poisson distributions with applications to statistical analysis of graphs. Advances in Applied Probability, 20(2):315–330, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427392.

Nair:1993:RBI

[NP93] M. G. Nair and P. K. Pollett. On the relationship between μ-invariant measures and quasi-stationary distributions for continuous-time Markov chains. Advances in Applied Probability, 25(1):82–102, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427497. See correction [Ano93g].

Nassar:2017:LTB

[NP17] Elma Nassar and Etienne Pardoux. On the large-time behaviour of the solution of a stochastic differential equation driven by a Poisson point process. Advances in Applied Probability, 49(2):344–367, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-the-largetime-behaviour-of-the-
solution-of-a-stochastic-differential-equation-driven-
by-a-poisson-point-process/C25EC159C4ED798EF64FD17C6EAF414B.

Neuts:2000:RMO

[NPOTC00] Marcel F. Neuts, Rafael Pérez-Ocón, and Inmaculada Torres-Castro. Repairable models with operating and repair times governed by phase type distributions. Advances in Applied Probability, 32(2):468–479, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428199.

Newman:1985:NNV

[NR85] Charles M. Newman and Yosef Rinott. Nearest neighbors and Voronoi volumes in high-dimensional point processes with various distance functions. Advances in Applied Probability, 17(4): 794–809, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427088.

Norros:2006:CPG

[NR06] Ilkka Norros and Hannu Reittu. On a conditionally Poissonian graph process. Advances in Applied Probability, 38(1):59– 75, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443428.

Negrea:2021:AGE

[NR21] Jeffrey Negrea and Jeffrey S. Rosenthal. Approximations of geometrically ergodic reversible Markov chains. Advances in Applied Probability, 53(4):981–1022, December 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/approximations-of-geometrically-
ergodic-reversible-markov-chains/1E014FFA33D724FEC7388AF408A79F99.

Norros:2022:MSD

[NRB22] Ilkka Norros, Hannu Reittu, and Fülöp Bazsó. On model selection for dense stochastic block models. Advances in Applied Probability, 54(1):202–226, March 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/on-model-selection-for-dense-stochastic-
block-models/19D2041A7368B8CD28D24030474D449F.

Newman:1983:NNV

[NRT83] C. M. Newman, Y. Rinott, and A. Tversky. Nearest neighbors and Voronoi regions in certain point processes. Advances in Applied Probability, 15(4):726–751, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427321.

Newman:1977:DBT

[NS77] Charles M. Newman and B. W. Stuck. Discriminating between two Markov processes. Advances in Applied Probability, 9(2):229–230, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426374.

Neuhauser:1993:BAB

[NS93] Claudia Neuhauser and Aidan Sudbury. The biased annihilating branching process. Advances in Applied Probability, 25(1):24– 38, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427494.

Navarro:2013:MRJ

[NSB13] Jorge Navarro, Francisco J. Samaniego, and N. Balakrishnan. Mixture representations for the joint distribution of lifetimes of two coherent systems with shared components. Advances in Applied Probability, 45(4):1011–1027, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1386857855.

Natvig:1986:MRT

[NSHH86] Bent Natvig, Skule Sørmo, Arne T. Holen, and Gutorm Høgåsen. Multistate reliability theory: A case study. Advances in Applied Probability, 18(4):921–932, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427256.

Nummelin:1976:LTR

[Num76] Esa Nummelin. Limit theorems for α-recurrent semi-Markov processes. Advances in Applied Probability, 8(3):531–547, September 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426143.

Nummelin:1979:CPG

[Num79] E. Nummelin. A conservation property for general GI/G/1 queues with an application to tandem queues. Advances in Applied Probability, 11(3):660–672, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426960.

Nummelin:1981:RTQ

[Num81] E. Nummelin. Regeneration in tandem queues. Advances in Applied Probability, 13(1):221–230, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426476.

Nagel:1998:IDQ

[NW98] Werner Nagel and Viola Weiss. Interdependence of directional quantities of planar tessellations. Advances in Applied Probability, 30(2):286–287, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427959.

Nagel:2003:LSS

[NW03] Werner Nagel and Viola Weiss. Limits of sequences of stationary planar tessellations. Advances in Applied Probability, 35(1):123– 138, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428276.

Nagel:2005:CST

[NW05] Werner Nagel and Viola Weiss. Crack STIT tessellations: Characterization of stationary random tessellations stable with respect to iteration. Advances in Applied Probability, 37(4):859– 883, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/30037364.

Nolde:2022:LRM

[NW22] Natalia Nolde and Jennifer L. Wadsworth. Linking representations for multivariate extremes via a limit set. Advances in Applied Probability, 54(3):688–717, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/linking-representations-
for-multivariate-extremes-via-a-limit-set/9A77F8E10602DC13EA26E429CFB5FD21.

Nyrhinen:1998:RDR

[Nyr98] Harri Nyrhinen. Rough descriptions of ruin for a general class of surplus processes. Advances in Applied Probability, 30(4):1008– 1026, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428027.

Nyrhinen:2005:PER

[Nyr05] Harri Nyrhinen. Power estimates for ruin probabilities. Advances in Applied Probability, 37(3):726–742, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/30037352.

Nyrhinen:2016:RGR

[Nyr16] Harri Nyrhinen. On real growth and run-off companies in insurance ruin theory. Advances in Applied Probability, 48(3): 903–925, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1474296320.

Nguyen:1977:IDC

[NZ77] X. X. Nguyen and H. Zessin. Integral and differential characterizations of the Gibbs process. Advances in Applied Probability, 9(3):446–447, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426106.

OBrien:1980:STV

[O’B80] G. L. O’Brien. Scaling transformations for 0,1-valued sequences. Advances in Applied Probability, 12(2):286, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426558.

Orsingher:1989:CDS

[OB89] Enzo Orsingher and Bruno Bassan. On the comparison of the distribution of the supremum of random fields represented by stochastic integrals. Advances in Applied Probability, 21(4): 770–780, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427766.

Ormeci:2005:DAC

[ÖB05] E. L. Örmeci and A. Burnetas. Dynamic admission control for loss systems with batch arrivals. Advances in Applied Probability, 37(4):915–937, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/30037367.

OCinneide:1991:PTD

[O’C91] Colm Art O’Cinneide. Phase-type distributions and invariant polytopes. Advances in Applied Probability, 23(3):515–535, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427620.

OConnell:1995:GBP

[O’C95] Neil O’Connell. The genealogy of branching processes and the age of our most recent common ancestor. Advances in Applied Probability, 27(2):418–442, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427834.

Olvera-Cravioto:2010:DNU

[OC10] Mariana Olvera-Cravioto. On the distribution of the nearly unstable AR(1) process with heavy tails. Advances in Applied Probability, 42(1):106–136, March 2010. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/25683809.

Olvera-Cravioto:2012:AWR

[OC12] Mariana Olvera-Cravioto. Asymptotics for weighted random sums. Advances in Applied Probability, 44(4):1142–1172, December 2012. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
41805993.

Ocafrain:2023:ETA

[Oça23] William Oçafrain. An ergodic theorem for asymptotically periodic time-inhomogeneous Markov processes, with application to quasi-stationarity with moving boundaries. Advances in Applied Probability, 55(2):672–700, June 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/an-ergodic-theorem-for-asymptotically-
periodic-timeinhomogeneous-markov-processes-with-application-
to-quasistationarity-with-moving-boundaries/A8B0E2C3DCF15E359C25FCA6F23DB8FE.

Orsingher:2007:RMF

[OD07] E. Orsingher and A. De Gregorio. Random motions at finite velocity in a non-Euclidean space. Advances in Applied Probability, 39(2):588–611, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443528.

Odoni:1980:AAB

[Odo80] R. W. K. Odoni. The algebraic approach to Bernstein’s problem on Mendelian models of inheritance. Advances in Applied Probability, 12(1):9, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426487.

Offord:1974:ASP

[Off74] A. C. Offord. The application of stochastic processes in function theory. Advances in Applied Probability, 6(2):202–205, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426247.

Okada:1979:CDP

[Oka79] Norio Okada. On convergence to diffusion processes of Markov chains related to population genetics. Advances in Applied Probability, 11(4):673–700, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1426854.

Olsbo:2007:CBV

[Ols07] Viktor Olsbo. On the correlation between the volumes of the typical Poisson Voronoi cell and the typical stienen sphere. Advances in Applied Probability, 39(4):883–892, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/20443549.

Ohser:1995:SSC

[OM95] J. Ohser and F. Mucklich. Stereology for some classes of polyhedrons. Advances in Applied Probability, 27(2):384–396, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427832.

Omey:2006:MSD

[OMS06] E. Omey, F. Mallor, and J. Santos. Multivariate subexponential distributions and random sums of random vectors. Advances in Applied Probability, 38(4):1028–1046, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443486.

ONeill:1995:EMF

[O’N95] Philip O’Neill. Epidemic models featuring behaviour change. Advances in Applied Probability, 27(4):960–979, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427931.

Ohser:1996:SES

[ON96] Joachim Ohser and Werner Nagel. Stereological estimation of the size distribution for systems of random convex polyhedrons. Advances in Applied Probability, 28(2):340–341, June 1996. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428053.

Ohser:1998:SEP

[ON98] Joachim Ohser and Martin Nippe. On stereological estimation of particle size distributions. Advances in Applied Probability, 30 (2):288, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427961.

Ornstein:1977:MSR

[Orn77] Donald Ornstein. Mechanical systems and random processes. Advances in Applied Probability, 9(2):223–224, June 1977. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426367.

Orsingher:1984:DVE

[Ors84] Enzo Orsingher. Damped vibrations excited by white noise. Advances in Applied Probability, 16(3):562–584, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427287.

Orsingher:1990:RMG

[Ors90] Enzo Orsingher. Random motions governed by third-order equations. Advances in Applied Probability, 22(4):915–928, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427568.

Oliveira:2018:DCC

[ORS18] R. I. Oliveira, R. Ribeiro, and R. Sanchis. Disparity of clustering coefficients in the Holme–Kim network model. Advances in Applied Probability, 50(3):918–943, September 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/disparity-of-clustering-
coefficients-in-the-holmekim-network-model/C5B08A9669C875318E5198567021A2BA.

Orsingher:2016:PMS

[ORT16] Enzo Orsingher, Costantino Ricciuti, and Bruno Toaldo. Population models at stochastic times. Advances in Applied Probability, 48(2):481–498, June 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1465490759.

Oesting:2018:ESB

[OS18] Marco Oesting and Kirstin Strokorb. Efficient simulation of Brown–Resnick processes based on variance reduction of Gaussian processes. Advances in Applied Probability, 50 (4):1155–1175, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/efficient-simulation-of-brownresnick-
processes-based-on-variance-reduction-of-gaussian-processes/
70A17B7BA5CBA0B45FE450895004E777.

Osogami:2010:FLC

[Oso10] Takayuki Osogami. A fluid limit for a cache algorithm with general request processes. Advances in Applied Probability, 42 (3):816–833, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/25799069.

Oksendal:2011:OCS

[ØSZ11] Bernt Øksendal, Agnès Sulem, and Tusheng Zhang. Optimal control of stochastic delay equations and time-advanced backward stochastic differential equations. Advances in Applied Probability, 43(2):572–596, June 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/23024520.

Omey:2002:WRF

[OT02] Edward Omey and Jef L. Teugels. Weighted renewal functions: A hierarchical approach. Advances in Applied Probability, 34 (2):394–415, June 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428294.

Owada:2020:LTP

[OT20] Takashi Owada and Andrew M. Thomas. Limit theorems for process-level Betti numbers for sparse and critical regimes. Advances in Applied Probability, 52(1):1–31, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/limit-theorems-for-processlevel-
betti-numbers-for-sparse-and-critical-regimes/8CF5AE51557041B89D278A4801F3C64B.

Ott:1974:IDS

[Ott74] T. J. Ott. Infinite divisibility and stability of finite semi-Markov matrices. Advances in Applied Probability, 6(2):217–218, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426255.

Ott:1975:IDI

[Ott75] Teunis J. Ott. Infinite divisibility and inbeddability of semi-Markov matrices on certain topological groups. Advances in Applied Probability, 7(2):256–257, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426071.

Ott:1976:CFB

[Ott76] Teunis J. Ott. The covariance function of Brownian motion with drift, made ergodic by a reflecting boundary. Advances in Applied Probability, 8(2):234–235, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1425882.

Ott:1977:CFV

[Ott77a] Teunis J. Ott. The covariance function of the virtual waitingtime process in an M/G/1 queue. Advances in Applied Probability, 9(1):158–168, March 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425822.

Ott:1977:SMG

[Ott77b] Teunis J. Ott. The stable M/G/1 queue in heavy traffic and its covariance function. Advances in Applied Probability, 9(1):169– 186, March 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425823.

Ott:1987:SWT

[Ott87a] Teunis J. Ott. On the stationary waiting-time distribution in the GI/G/1 queue, i: Transform methods and almost-phase-type distributions. Advances in Applied Probability, 19(1):240– 265, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427382.

Ott:1987:SSQ

[Ott87b] Teunis J. Ott. The single-server queue with independent GI/G and M/G input streams. Advances in Applied Probability, 19 (1):266–286, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427383.

Ott:2006:RCS

[Ott06] Teunis J. Ott. Rate of convergence for the ’square root formula’ in the Internet Transmission Control Protocol. Advances in Applied Probability, 38(4):1132–1154, December 2006. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443491.

Owada:2022:STC

[OY22] Takashi Owada and D. Yogeshwaran. Sub-tree counts on hyperbolic random geometric graphs. Advances in Applied Probability, 54(4):1032–1069, December 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/subtree-counts-on-hyperbolic-random-
geometric-graphs/C0B310693219D6445EA1B361F34738A3.

Paige:2001:CLS

[Pai01] R. L. Paige. Classification and lumpability in the stochastic Hopfield model. Advances in Applied Probability, 33(4):930–943, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428313.

Pakes:1971:BPS

[Pak71a] A. G. Pakes. A branching process with a state dependent immigration component. Advances in Applied Probability, 3(2): 301–314, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426173.

Pakes:1971:SLT

[Pak71b] A. G. Pakes. Some limit theorems for the total progeny of a branching process. Advances in Applied Probability, 3(1):176– 192, Spring 1971. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426333.

Pakes:1979:LTS

[Pak79] Anthony G. Pakes. Limit theorems for the simple branching process allowing immigration, I. The case of finite offspring mean. Advances in Applied Probability, 11(1):31–62, March 1979. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426767.

Pakes:1986:SPB

[Pak86] Anthony G. Pakes. Some properties of a branching process with group immigration and emigration. Advances in Applied Probability, 18(3):628–645, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1427180.

Pakes:1989:ARE

[Pak89a] Anthony G. Pakes. Asymptotic results for the extinction time of Markov branching processes allowing emigration, I. Random walk decrements. Advances in Applied Probability, 21(2):243– 269, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427159.

Pakes:1989:ABE

[Pak89b] Anthony G. Pakes. On the asymptotic behaviour of the extinction time of the simple branching process. Advances in Applied Probability, 21(2):470–472, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427170.

Pakes:1992:LTN

[Pak92] Anthony G. Pakes. Limit theorems for the numbers of rare mutants: A branching process model. Advances in Applied Probability, 24(4):778–794, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427712.

Pakes:1993:EMB

[Pak93] Anthony G. Pakes. Explosive Markov branching processes: Entrance laws and limiting behaviour. Advances in Applied Probability, 25(4):737–756, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427789.

Pakes:1995:QSL

[Pak95] Anthony G. Pakes. Quasi-stationary laws for Markov processes: Examples of an always proximate absorbing state. Advances in Applied Probability, 27(1):120–145, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428100.

Pakes:1998:LTM

[Pak98] Anthony G. Pakes. A limit theorem for the maxima of the Paracritical simple branching process. Advances in Applied Probability, 30(3):740–756, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428149.

Pakes:2000:NSN

[Pak00] Anthony G. Pakes. The number and sum of near-maxima for thin-tailed populations. Advances in Applied Probability, 32(4): 1100–1116, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428519.

Pakes:2010:CMB

[Pak10] Anthony G. Pakes. Critical Markov branching process limit theorems allowing infinite variance. Advances in Applied Probability, 42(2):460–488, June 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/25683829.

Palacios:1993:FTE

[Pal93] José Luis Palacios. Fluctuation theory for the Ehrenfest urn via electric networks. Advances in Applied Probability, 25(2):472– 476, June 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427663.

Palacios:1994:ALE

[Pal94] José Luis Palacios. Another look at the Ehrenfest urn via electric networks. Advances in Applied Probability, 26(3):820–824, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427822.

Panaretos:2006:DRS

[Pan06] Victor M. Panaretos. The diffusion of Radon shape. Advances in Applied Probability, 38(2):320–335, June 2006. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443445.

Papangelou:1977:CIP

[Pap77] F. Papangelou. Conditional intensities of point processes: Their application to Davidson’s problem and to entropy. Advances in Applied Probability, 9(3):434, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426095.

Papastavridis:1987:LTR

[Pap87] S. Papastavridis. A limit theorem for the reliability of a consecutive-k-out-of-n system. Advances in Applied Probability, 19(3):746–748, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427417.

Papastavridis:1988:WLR

[Pap88] Stavros G. Papastavridis. A Weibull limit for the reliability of a consecutive k-within-m-out-of-n system. Advances in Applied Probability, 20(3):690–692, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427045.

Parzen:1977:IBT

[Par77] Emanuel Parzen. Isomorphism between time series analysis and non-parametric statistical inference. Advances in Applied Probability, 9(2):232–233, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426378.

Parthasarathy:1987:TSM

[Par87] P. R. Parthasarathy. A transient solution to an M/M/1 queue: A simple approach. Advances in Applied Probability, 19(4): 997–998, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427113.

Paroux:1998:QTC

[Par98] Katy Paroux. Quelques théorèmes centraux limites pour les processus poissoniens de droites dans le plan. (French) [Some central limit theorems for Poissonian processes of lines in the plane]. Advances in Applied Probability, 30(3):640–656, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428143.

Paulsen:2007:ODP

[Pau07] Jostein Paulsen. Optimal dividend payments until ruin of diffusion processes when payments are subject to both fixed and proportional costs. Advances in Applied Probability, 39(3):669–689, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443535.

Pan:2019:EAL

[PB19] Yuqing Pan and Konstantin A. Borovkov. The exact asymptotics of the large deviation probabilities in the multivariate boundary crossing problem. Advances in Applied Probability, 51(3):835–864, September 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/exact-asymptotics-of-the-large-deviation-
probabilities-in-the-multivariate-boundary-crossing-
problem/41ECCBAC85C36B90FFEA73BD7F373E20.

Prabhakar:2000:SPP

[PBM00] Balaji Prabhakar, Nicholas Bambos, and T. S. Mountford. The synchronization of Poisson processes and queueing networks with service and synchronization nodes. Advances in Applied Probability, 32(3):824–843, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428415.

Poskitt:1996:MCM

[PC96] D. S. Poskitt and Shin-Ho Chung. Markov chain models, time series analysis and extreme value theory. Advances in Applied Probability, 28(2):405–425, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428065.

Perez:2021:VIM

[PC21] Iker Perez and Giuliano Casale. Variational inference for Markovian queueing networks. Advances in Applied Probability, 53(3):687–715, September 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/variational-inference-for-markovian-
queueing-networks/D35E7DB62BE78D883730A04E617C3DB3.

Penda:2022:CLT

[PD22] S. Valère Bitseki Penda and Jean-François Delmas. Central limit theorem for bifurcating Markov chains under L^{2}- ergodic conditions. Advances in Applied Probability, 54 (4):999–1031, December 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/central-limit-theorem-for-bifurcating-
markov-chains-under-l2ergodic-conditions/4446AF23ADD9CB626DFFB29C4DA2CBD3.

Pearce:1989:ECF

[Pea89] C. E. M. Pearce. Extended continued fractions, recurrence relations and two-dimensional Markov processes. Advances in Applied Probability, 21(2):357–375, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427164.

Peligrad:1983:NTM

[Pel83] Magda Peligrad. A note on two measures of dependence and mixing sequences. Advances in Applied Probability, 15(2):461– 464, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426446.

Pellerey:1993:POU

[Pel93] Franco Pellerey. Partial orderings under cumulative damage shock models. Advances in Applied Probability, 25(4):939–946, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1427800. See correction [Ano94e].

Penrose:1991:CPM

[Pen91] Mathew D. Penrose. On a continuum percolation model. Advances in Applied Probability, 23(3):536–556, September 1991.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427621.

Penrose:1998:EMS

[Pen98] Mathew D. Penrose. Extremes for the minimal spanning tree on normally distributed points. Advances in Applied Probability, 30 (3):628–639, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428142.

Penrose:2002:FSS

[Pen02] Mathew D. Penrose. Focusing of the scan statistic and geometric clique number. Advances in Applied Probability, 34(4):739–753, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428320.

Percus:1985:PTO

[Per85a] Ora E. Percus. Phase transition in one-dimensional random walk with partially reflecting boundaries. Advances in Applied Probability, 17(3):594–606, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1427121.

Perry:1985:ISP

[Per85b] David Perry. An inventory system for perishable commodities with random lifetime. Advances in Applied Probability, 17(1): 234–236, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427063.

Perfekt:1997:EVT

[Per97] Roland Perfekt. Extreme value theory for a class of Markov chains with values in R^{d}. Advances in Applied Probability, 29 (1):138–164, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427864.

Peskun:1975:GCT

[Pes75] P. H. Peskun. Guidelines for choosing the transition matrix in Monte Carlo sampling methods using Markov chains. Advances in Applied Probability, 7(2):261, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426075.

Peschel:1982:MEP

[Pes82a] M. Peschel. Modelling of evolution processes with powderproduct driving forces. Advances in Applied Probability, 14 (2):199–200, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426511.

Pestien:1982:EFE

[Pes82b] Victor C. Pestien. An extended Fatou equation and continuoustime gambling. Advances in Applied Probability, 14(2):309– 323, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426523.

Petruccelli:1982:FIB

[Pet82] Joseph D. Petruccelli. Full-information best-choice problems with recall of observations and uncertainty of selection depending on the observation. Advances in Applied Probability, 14 (2):340–358, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426525.

Petruccell:1984:BCP

[Pet84] Joseph D. Petruccell. Best-choice problems involving recall and uncertainty of selection when the number of observations is random. Advances in Applied Probability, 16(1):111–130, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427227.

Pfeifer:1983:NOT

[Pfe83] Dietmar Pfeifer. A note on the occurrence times of a Pólya– Lundberg process. Advances in Applied Probability, 15(4):886, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427330.

Pfeifer:1985:ACA

[Pfe85a] Dietmar Pfeifer. An average-case analysis for a continuous random search algorithm. Advances in Applied Probability, 17(1): 231–233, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427062.

Pfeifer:1985:RBR

[Pfe85b] Dietmar Pfeifer. On a relationship between record values and Ross’s model of algorithm efficiency. Advances in Applied Probability, 17(2):470–471, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427154.

Pfeifer:1991:SRN

[Pfe91] Dietmar Pfeifer. Some remarks on Nevzorov’s record model. Advances in Applied Probability, 23(4):823–834, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427678.

Purdue:1985:I

[PG85] Peter Purdue and J. Gani. Introduction. Advances in Applied Probability, 17(2):243–244, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427129.

Paulsen:1997:RTS

[PG97] Jostein Paulsen and Håkon K. Gjessing. Ruin theory with stochastic return on investments. Advances in Applied Probability, 29(4):965–985, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427849.

Pitts:1996:CBA

[PGE96] Susan M. Pitts, Rudolf Grübel, and Paul Embrechts. Confidence bounds for the adjustment coefficient. Advances in Applied Probability, 28(3):802–827, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428182.

Paulsen:1999:MCM

[PH99] Jostein Paulsen and Arne Hove. Markov chain Monte Carlo simulation of the distribution of some perpetuities. Advances in Applied Probability, 31(1):112–134, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427994.

Phillips:1978:SRV

[Phi78] M. J. Phillips. Sums of random variables having the modified geometric distribution with application to two-person games. Advances in Applied Probability, 10(3):647–665, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426639.

Perez:2016:MDP

[PHL16] Iker Perez, David Hodge, and Huiling Le. Markov decision process algorithms for wealth allocation problems with defaultable bonds. Advances in Applied Probability, 48(2):392–405, June 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1465490754.

Phoenix:1979:ADT

[Pho79] S. Leigh Phoenix. The asymptotic distribution for the time to failure of a fiber bundle. Advances in Applied Probability, 11 (1):153–187, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426773.

Piau:2001:PBC

[Pia01] Didier Piau. Processus de branchement en champ moyen et réaction PCR. (French) [Branching processes in mean field and the PCR reaction]. Advances in Applied Probability, 33(2):391– 403, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428259.

Piau:2006:HMI

[Pia06] Didier Piau. Harmonic moments of inhomogeneous branching processes. Advances in Applied Probability, 38(2):465–486, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/20443451.

Picinbono:1974:PASa

[Pic74a] B. Picinbono. Properties and applications of stochastic processes with stationary increments. Advances in Applied Probability, 6(2):205–207, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426248.

Picinbono:1974:PASb

[Pic74b] B. Picinbono. Properties and applications of stochastic processes with stationary n th-order increments. Advances in Applied Probability, 6(3):512–523, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426231.

Pickard:1977:AII

[Pic77] D. K. Pickard. Asymptotic inference for an Ising lattice. II. Advances in Applied Probability, 9(3):476–501, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426110.

Pickard:1980:UMF

[Pic80] David K. Pickard. Unilateral Markov fields. Advances in Applied Probability, 12(3):655–671, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426425.

Pickands:1983:EOS

[Pic83] James Pickands III. Extreme order statistics with cost of sampling. Advances in Applied Probability, 15(4):783–797, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427324.

Picard:1985:TEC

[Pic85a] Dominique Picard. Testing and estimating change-points in time series. Advances in Applied Probability, 17(4):841–867, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427090.

Pickard:1985:SID

[Pic85b] David K. Pickard. Statistical inference on discrete Markov fields. Advances in Applied Probability, 17(2):253–254, June 1985. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427137.

Picard:1998:SHM

[Pic98] Philippe Picard. A semi-hybrid model for malaria with limited superinfection. Advances in Applied Probability, 30(4):1027– 1057, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428028.

Picka:2000:VRM

[Pic00] Jeffrey D. Picka. Variance reducing modifications for estimators of standardized moments of random sets. Advances in Applied Probability, 32(3):682–700, September 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428408.

Pierson:1978:CRM

[Pie78] Harry M. Pierson. On a class of random motions of point processes involving interaction. Advances in Applied Probability, 10 (3):613–632, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426637.

Pimentel:2007:MST

[Pim07] Leandro P. R. Pimentel. Multitype shape theorems for first passage percolation models. Advances in Applied Probability, 39(1):53–76, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443500.

Pinsky:2016:SRW

[Pin16] Ross G. Pinsky. The speed of a random walk excited by its recent history. Advances in Applied Probability, 48(1):215–234, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1457466163.

Pitman:1974:ODB

[Pit74] J. W. Pitman. One-dimensional Brownian motion and the threedimensional Bessel process. Advances in Applied Probability, 6 (2):223–224, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426261.

Pitman:1975:ODB

[Pit75] James W. Pitman. One-dimensional Brownian motion and the three-dimensional Bessel process. Advances in Applied Probability, 7(3):511–526, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426125.

Piterbarg:1977:RIV

[Pit77a] V. I. Piterbarg. On the regular increase of the variance of the number of random points in Euclidean space. Advances in Applied Probability, 9(3):443–444, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426103.

Pitman:1977:OMM

[Pit77b] J. W. Pitman. Occupation measures for Markov chains. Advances in Applied Probability, 9(1):69–86, March 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425817.

Pitman:1978:EFT

[Pit78] J. W. Pitman. An extension of de Finetti’s theorem. Advances in Applied Probability, 10(2):268–270, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426877.

Pittel:1986:PRD

[Pit86] Boris Pittel. Paths in a random digital tree: Limiting distributions. Advances in Applied Probability, 18(1):139–155, March 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427240.

See correction [Ano92e].

Pitman:1996:RDD

[Pit96] Jim Pitman. Random discrete distributions invariant under sizebiased permutation. Advances in Applied Probability, 28(2):525– 539, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428070.

Paulin:2018:CPP

[PJCB18] Daniel Paulin, Ajay Jasra, Dan Crisan, and Alexandros Beskos. On concentration properties of partially observed chaotic systems. Advances in Applied Probability, 50(2):440–479, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/on-
concentration-properties-of-partially-observed-chaotic-
systems/1CB3DD888464BFBCC3848EA7785B2758.

Pikovsky:1996:APO

[PK96] Igor Pikovsky and Ioannis Karatzas. Anticipative portfolio optimization. Advances in Applied Probability, 28(4):1095–1122, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428166.

Perry:1989:CPI

[PL89] David Perry and Benny Levikson. Continuous Production/Inventory model with analogy to certain queueing and dam models. Advances in Applied Probability, 21(1):123–141, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427200.

Picard:1990:UAF

[PL90] Philippe Picard and Claude Lefevre. A unified analysis of the final size and severity distribution in collective Reed–Frost epidemic processes. Advances in Applied Probability, 22(2):269– 294, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427536.

Picard:1996:FCB

[PL96] Philippe Picard and Claude Lefèvre. First crossing of basic counting processes with lower non-linear boundaries: A unified approach through pseudopolynomials (i). Advances in Applied Probability, 28(3):853–876, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428184.

Picard:1999:ASM

[PL99] Philippe Picard and Claude Lefèvre. On the algebraic structure in Markovian processes of death and epidemic types. Advances in Applied Probability, 31(3):742–757, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428394.

Platen:2002:ACC

[Pla02] Eckhard Platen. Arbitrage in continuous complete markets. Advances in Applied Probability, 34(3):540–558, September 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428246.

Puente:1996:GDR

[PLPA96] Carlos E. Puente, Miguel M. López, Jorge E. Pinzón, and José M. Angulo. The Gaussian distribution revisited. Advances in Applied Probability, 28(2):500–524, June 1996. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428069.

Pateiro-Lopez:2008:LSA

[PLRC08] Beatriz Pateiro-López and Alberto Rodríguez-Casal. Length and surface area estimation under smoothness restrictions. Advances in Applied Probability, 40(2):348–358, June 2008. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443586.

Park:2008:MMC

[PM08] Hyun Suk Park and Ross Maller. Moment and MGF convergence of overshoots and undershoots for Lévy insurance risk processes. Advances in Applied Probability, 40(3):716–733, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443605.

Piera:2005:PFS

[PMG05] Francisco J. Piera, Ravi R. Mazumdar, and Fabrice M. Guillemin. On product-form stationary distributions for reflected diffusions with jumps in the positive orthant. Advances in Applied Probability, 37(1):212–228, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037322.

Piera:2008:BBP

[PMG08] Francisco J. Piera, Ravi R. Mazumdar, and Fabrice M. Guillemin. Boundary behavior and product-form stationary distributions of jump diffusions in the orthant with state-dependent reflections. Advances in Applied Probability, 40(2):529–547, June 2008. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443594.

Perez-Ocon:1995:HPC

[POGP95] Rafael Pérez-Ocón and M. Luz Gámiz-Pérez. On the HNBUE property in a class of correlated cumulative shock models. Advances in Applied Probability, 27(4):1186–1188, December 1995.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427938.

Pollak:1974:SPE

[Pol74] Edward Pollak. Survival probabilities and extinction times for some multitype branching processes. Advances in Applied Probability, 6(3):446–462, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426227.

Pollard:1975:CPQ

[Pol75] J. M. Pollard. Combinatorial properties of a queueing system with limited availability. Advances in Applied Probability, 7(4): 864–877, December 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426404.

Pollard:1978:FIR

[Pol78] J. H. Pollard. Fluctuating interest rates. Advances in Applied Probability, 10(4):707–711, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426648.

Pollard:1979:WCN

[Pol79] D. Pollard. Weak convergence on non-separable metric spaces. Advances in Applied Probability, 11(2):286–288, June 1979. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426817.

Pollard:1980:NPP

[Pol80] David Pollard. A new proof of Prohorov’s theorem. Advances in Applied Probability, 12(2):295–296, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426570.

Pollett:1986:CRM

[Pol86] P. K. Pollett. Connecting reversible Markov processes. Advances in Applied Probability, 18(4):880–900, December 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427254.

Pollett:1987:PPB

[Pol87] P. K. Pollett. Preserving partial balance in continuous-time Markov chains. Advances in Applied Probability, 19(2):431– 453, June 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427426.

Pollett:1988:RII

[Pol88] P. K. Pollett. Reversibility, invariance and μ-invariance. Advances in Applied Probability, 20(3):600–621, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427037.

Pollett:1991:IMP

[Pol91] P. K. Pollett. Invariant measures for Q-processes when Q is not regular. Advances in Applied Probability, 23(2):277–292, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427748.

Politis:1994:MCM

[Pol94] Dimitris N. Politis. Markov chains in many dimensions. Advances in Applied Probability, 26(3):756–774, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427819.

Pontiggia:2005:TPR

[Pon05] Laura Pontiggia. Two-person red-and-black with bet-dependent win probabilities. Advances in Applied Probability, 37(1):75– 89, March 2005. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
30037316.

Penrose:1996:LDD

[PP96] Mathew D. Penrose and Agoston Pisztora. Large deviations for discrete and continuous percolation. Advances in Applied Probability, 28(1):29–52, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427912.

Papadatos:2002:PAS

[PP02] N. Papadatos and V. Papathanasiou. Poisson approximation for a sum of dependent indicators: An alternative approach. Advances in Applied Probability, 34(3):609–625, September 2002.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428249.

Pitts:2007:AGS

[PP07] Susan M. Pitts and Konstadinos Politis. Approximations for the Gerber–Shiu expected discounted penalty function in the compound Poisson risk model. Advances in Applied Probability, 39(2):385–406, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443519.

Peterson:2015:DBT

[PP15] Nicholas R. Peterson and Boris Pittel. Distance between two random k-out digraphs, with and without preferential attachment. Advances in Applied Probability, 47(3):858–879, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1444308885.

Pinedo:1982:MEM

[PR82] Michael L. Pinedo and Sheldon M. Ross. Minimizing expected makespan in stochastic open shops. Advances in Applied Probability, 14(4):898–911, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427030.

Pollett:1990:DLT

[PR90] P. K. Pollett and A. J. Roberts. A description of the long-term behaviour of absorbing continuous-time Markov chains using a centre manifold. Advances in Applied Probability, 22(1):111– 128, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427600.

Platen:1994:PAS

[PR94] Eckhard Platen and Rolando Rebolledo. Pricing via anticipative stochastic calculus. Advances in Applied Probability, 26(4): 1006–1021, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427902.

Piterbarg:1999:CLT

[PR99] Vladimir Piterbarg and Igor Rychlik. Central limit theorem for wave-functionals of Gaussian processes. Advances in Applied Probability, 31(1):158–177, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427996.

Puhalskii:2000:MGP

[PR00] A. A. Puhalskii and M. I. Reiman. The multiclass GI/PH/N queue in the Halfin–Whitt regime. Advances in Applied Probability, 32(2):564–595, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428204. See correction [Ano04e].

Palmowski:2006:EAB

[PR06] Zbigniew Palmowski and Tomasz Rolski. On the exact asymptotics of the busy period in GI/G/1 queues. Advances in Applied Probability, 38(3):792–803, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443471.

Pekoz:2022:FGR

[PR22] Erol A. Peköz and Sheldon M. Ross. Fair gambler’s ruin stochastically maximizes playing time. Advances in Applied Probability, 54(2):656–659, June 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/fair-gamblers-ruin-stochastically-
maximizes-playing-time/D5B37447C3A8BEFD68D228A2B1B70C7C.

Prabhu:1973:FRW

[Pra73] N. U. Prabhu. Further results on Wiener–Hopf factorization. Advances in Applied Probability, 5(1):19–20, April 1973. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425953.

Preston:1973:GGS

[Pre73] C. J. Preston. Generalized Gibbs states and Markov random fields. Advances in Applied Probability, 5(2):242–261, August 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426035.

Preston:1977:RBG

[Pre77] C. J. Preston. The relation between grand canonical and canonical Gibbs states. Advances in Applied Probability, 9(3):447, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426107.

Priestley:1977:AMT

[Pri77] M. B. Priestley. Applications of multivariate techniques in the study of multivariable stochastic systems. Advances in Applied Probability, 9(2):202–205, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426349.

Prigent:1999:IMC

[Pri99] Jean-Luc Prigent. Incomplete markets: Convergence of options values under the minimal martingale measure. Advances in Applied Probability, 31(4):1058–1077, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428344.

Privault:2015:LTI

[Pri15] Nicolas Privault. Laplace transform identities for the volume of stopping sets based on Poisson point processes. Advances in Applied Probability, 47(4):919–933, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1449859794.

Privault:2022:MMG

[Pri22] Nicolas Privault. Moments of Markovian growth-collapse processes. Advances in Applied Probability, 54(4):1070–1093, December 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/moments-
of-markovian-growthcollapse-processes/CAB250FAE97135B4D14681C830955DA2.

Prorok:1976:TPSa

[Pro76a] Philip C. Prorok. The theory of periodic screening I: Lead time and proportion detected. Advances in Applied Probability, 8 (1):127–143, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/ stable/1426025. See corrections [Ano76e].

Prorok:1976:TPSb

[Pro76b] Philip C. Prorok. The theory of periodic screening II: Doubly bounded recurrence times and mean lead time and detection probability estimation. Advances in Applied Probability, 8(3):

460–476, September 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426139.

Pekoz:2017:JDD

[PRR17] Erol Peköz, Adrian Röllin, and Nathan Ross. Joint degree distributions of preferential attachment random graphs. Advances in Applied Probability, 49(2):368–387, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/joint-degree-distributions-
of-preferential-attachment-random-graphs/45ABC12A41D04E1114950E8649023F33.

Pruscha:1978:SAN

[Pru78] H. Pruscha. Statistical analysis of a non-Markovian model with applications to behavioural processes. Advances in Applied Probability, 10(2):329–330, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426928.

Pontier:1983:AOA

[PS83] Monique Pontier and Jacques Szpirglas. Arrêt optimal avec contrainte. (French) [Optimal stopping with constraints]. Advances in Applied Probability, 15(4):798–812, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427325.

Papiez:1990:DML

[PS90a] Lech S. Papiez and George A. Sandison. A diffusion model with loss of particles. Advances in Applied Probability, 22(3):533–547, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427456.

Pillai:1990:DCM

[PS90b] R. N. Pillai and E. Sandhya. Distributions with complete monotone derivative and geometric infinite divisibility. Advances in Applied Probability, 22(3):751–754, September 1990. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427468. See correction

[Ano91e].

Piccioni:1994:IMC

[PS94a] Mauro Piccioni and Sergio Scarlatti. An iterative Monte Carlo scheme for generating Lie group-valued random variables. Advances in Applied Probability, 26(3):616–628, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427811.

Pollett:1994:EPC

[PS94b] P. K. Pollett and D. E. Stewart. An efficient procedure for computing quasi-stationary distributions of Markov chains with sparse transition structure. Advances in Applied Probability, 26 (1):68–79, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427580.

Perry:2001:DMI

[PS01] David Perry and Wolfgang Stadje. Disasters in a Markovian inventory system for perishable items. Advances in Applied Probability, 33(1):61–75, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428441.

Paganoni:2004:IRU

[PS04] Anna Maria Paganoni and Piercesare Secchi. Interacting reinforced-urn systems. Advances in Applied Probability, 36(3): 791–804, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/4140408.

Penrose:2009:MLE

[PS09] Mathew D. Penrose and Vadim Shcherbakov. Maximum likelihood estimation for cooperative sequential adsorption. Advances in Applied Probability, 41(4):978–1001, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/27793914.

Puplinskaite:2010:ARC

[PS10] Donata Puplinskaite and Donatas Surgailis. Aggregation of a random-coefficient AR(1) process with infinite variance and idiosyncratic innovations. Advances in Applied Probability, 42 (2):509–527, June 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25683831.

Penrose:2011:ANM

[PS11] Mathew D. Penrose and Vadim Shcherbakov. Asymptotic normality of the maximum likelihood estimator for cooperative sequential adsorption. Advances in Applied Probability, 43(3): 636–648, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/23024568.

Pantle:2006:CLT

[PSS06] Ursa Pantle, Volker Schmidt, and Evgueni Spodarev. Central limit theorems for functionals of stationary germ-grain models. Advances in Applied Probability, 38(1):76–94, March 2006. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443429.

Pilipauskaite:2020:JTC

[PSS20] Vytaute Pilipauskaite, Viktor Skorniakov, and Donatas Surgailis. Joint temporal and contemporaneous aggregation of random-coefficient AR(1) processes with infinite variance. Advances in Applied Probability, 52(1):237–265, March 2020. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/joint-temporal-
and-contemporaneous-aggregation-of-randomcoefficient-
ar1-processes-with-infinite-variance/34A5D052D3D282B3BF76C561D60A8E8F.

Papastathopoulos:2017:EEM

[PSTB17] I. Papastathopoulos, K. Strokorb, J. A. Tawn, and A. Butler. Extreme events of Markov chains. Advances in Applied Probability, 49(1):134–161, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/extreme-events-of-markov-chains/
5F5583374C237C8D87FA5394A045ACB1.

Phoenix:1973:ASD

[PT73] S. Leigh Phoenix and Howard M. Taylor. The asymptotic strength distribution of a general fiber bundle. Advances in Applied Probability, 5(2):200–216, August 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426033.

Pruscha:1977:FEV

[PT77] Helmut Pruscha and Radu Theodorescu. Functions of event variables of a random system with complete connections. Advances in Applied Probability, 9(2):228–229, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426373.

Pakes:1981:CAD

[PT81] Anthony G. Pakes and Simon Tavaré. Comments on the age distribution of Markov processes. Advances in Applied Probability, 13(4):681–703, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426967.

Pakes:1985:SPC

[PT85a] Anthony G. Pakes and A. C. Trajstman. Some properties of continuous-state branching processes, with applications to Bartoszyński’s virus model. Advances in Applied Probability, 17 (1):23–41, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427050.

Puri:1985:LBC

[PT85b] Prem S. Puri and Eric S. Tollar. On the limit behavior of certain quantities in a subcritical storage model. Advances in Applied Probability, 17(2):443–459, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427150.

Pandelis:1994:OMS

[PT94] Dimitrios G. Pandelis and Demosthenis Teneketzis. Optimal multiserver stochastic scheduling of two interconnected priority queues. Advances in Applied Probability, 26(1):258–279, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427590.

Platen:2023:NAM

[PT23] Eckhard Platen and Stefan Tappe. No arbitrage and multiplicative special semimartingales. Advances in Applied Probability, 55(3):1033–1074, September 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-

probability/article/no-arbitrage-and-multiplicative-
special-semimartingales/CC32BC3D990D2B7C90563633BA2A0C14.

Pakdaman:2010:FLT

[PTW10] K. Pakdaman, M. Thieullen, and G. Wainrib. Fluid limit theorems for stochastic hybrid systems with application to neuron models. Advances in Applied Probability, 42(3):761–794, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25799067.

Pollett:1992:DAS

[PV92] P. K. Pollett and A. Vassallo. Diffusion approximations for some simple chemical reaction schemes. Advances in Applied Probability, 24(4):875–893, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427717.

Passeggeri:2019:LTM

[PV19] Riccardo Passeggeri and Almut E. D. Veraart. Limit theorems for multivariate Brownian semistationary processes and feasible results. Advances in Applied Probability, 51 (3):667–716, September 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/limit-theorems-for-multivariate-
brownian-semistationary-processes-and-feasible-results/
373E9C7EE56858EDA3593A0FBD776436.

Pinto:2021:OSG

[PvS21] Mónica B. Carvajal Pinto and Kees van Schaik. Optimally stopping at a given distance from the ultimate supremum of a spectrally negative Lévy process. Advances in Applied Probability, 53(1):279–299, March 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/optimally-stopping-at-a-
given-distance-from-the-ultimate-supremum-of-a-spectrally-
negative-levy-process/EAE19AD4619DE7C3BB97C066152E3A42.

Penrose:2004:RMD

[PW04] Mathew D. Penrose and Andrew R. Wade. Random minimal directed spanning trees and Dickman-type distributions. Advances in Applied Probability, 36(3):691–714, September 2004.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/4140403.

Penrose:2006:TLR

[PW06] Mathew D. Penrose and Andrew R. Wade. On the total length of the random minimal directed spanning tree. Advances in Applied Probability, 38(2):336–372, June 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443446.

Penrose:2010:LTR

[PW10] Mathew D. Penrose and Andrew R. Wade. Limit theorems for random spatial drainage networks. Advances in Applied Probability, 42(3):659–688, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/25799063.

Perfilev:2018:LAA

[PW18] Elena Perfilev and Vitali Wachtel. Local asymptotics for the area under the random walk excursion. Advances in Applied Probability, 50(2):600–620, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/local-asymptotics-for-the-area-under-
the-random-walk-excursion/14F47DADAFDAF0530C6760D4C640962B.

Pedarsani:2017:RSF

[PWZ17] Ramtin Pedarsani, Jean Walrand, and Yuan Zhong. Robust scheduling for flexible processing networks. Advances in Applied Probability, 49(2):603–628, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/robust-scheduling-for-flexible-processing-
networks/83106AD90BF0A6F2557D2ED088CCBE38.

Pang:2013:HTL

[PY13] Guodong Pang and David D. Yao. Heavy-traffic limits for a many-server queueing network with switchover. Advances in Applied Probability, 45(3):645–672, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1377868533.

Pyke:1975:BSS

[Pyk75] Ronald Pyke. Brownian sheets: A survey. Advances in Applied Probability, 7(2):239–240, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426053.

Pyke:1979:RDE

[Pyk79] Ronald Pyke. Recent developments in empirical processes. Advances in Applied Probability, 11(2):267–268, June 1979. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426798.

Posner:1986:SMS

[PZ86] M. J. M. Posner and D. Zuckerman. Semi-Markov shock models with additive damage. Advances in Applied Probability, 18(3): 772–790, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427187.

Pang:2016:GGQ

[PZ16] Guodong Pang and Yuhang Zhou. G/G/∞ queues with renewal alternating interruptions. Advances in Applied Probability, 48(3):812–831, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1474296316.

Pirjol:2017:ADT

[PZ17] Dan Pirjol and Lingjiong Zhu. Asymptotics for the discrete-time average of the geometric Brownian motion and Asian options. Advances in Applied Probability, 49(2):446–480, June 2017. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/asymptotics-
for-the-discretetime-average-of-the-geometric-brownian-
motion-and-asian-options/105959C9E64A036DDAC74BAF03D24932.

Qiao:2015:NFS

[QD15] Huijie Qiao and Jinqiao Duan. Nonlinear filtering of stochastic dynamical systems with Lévy noises. Advances in Applied Probability, 47(3):902–918, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1444308887.

Qu:2019:ESL

[QDZ19] Yan Qu, Angelos Dassios, and Hongbiao Zhao. Efficient simulation of Lévy-driven point processes. Advances in Applied Probability, 51(4):927–966, December 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/efficient-simulation-of-
levydriven-point-processes/18AC974C432A3125E5CA6DC8BE1BC33C.

Qu:2011:FIA

[QG11] S. Qu and J. C. Gittins. A forwards induction approach to candidate drug selection. Advances in Applied Probability, 43 (3):649–665, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/23024569.

Qi:2008:FCL

[Qi08] Xin Qi. A functional central limit theorem for spatial birth and death processes. Advances in Applied Probability, 40(3): 759–797, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443607.

Qian:1990:NAI

[Qia90] Wei Qian. A note on asymptotic inference for an asymmetric Ising model around a torus. Advances in Applied Probability, 22 (3):755–757, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427469.

Qiao:2018:UMV

[Qia18] Huijie Qiao. Uniqueness for measure-valued equations of nonlinear filtering for stochastic dynamical systems with Lévy noise. Advances in Applied Probability, 50(2):396–413, June 2018. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/uniqueness-for-
measurevalued-equations-of-nonlinear-filtering-for-stochastic-
dynamical-systems-with-levy-noise/A091518D2FBA4E036B25F972B67B6249.

Quine:2000:GLB

[QS00] M. P. Quine and W. Szczotka. A general linear birth and growth model. Advances in Applied Probability, 32(4):1027–1050, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428515.

Quintanilla:1997:CCP

[QT97] J. Quintanilla and S. Torquato. Clustering in a continuum percolation model. Advances in Applied Probability, 29(2):327– 336, June 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428005.

Quine:1972:MGW

[Qui72] M. P. Quine. The multitype Galton–Watson process with ρ near 1. Advances in Applied Probability, 4(3):429–452, December 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425989.

Rajan:1998:CQS

[RA98] Rajendran Rajan and Rajeev Agrawal. Concavity of queueing systems with NBU service times. Advances in Applied Probability, 30(2):551–567, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427982.

Rahimov:2007:FLT

[Rah07] I. Rahimov. Functional limit theorems for critical processes with immigration. Advances in Applied Probability, 39(4):1054–1069, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443558.

Raic:2007:CRL

[Rai07] Martin Raic. CLT-Related large deviation bounds based on Stein’s method. Advances in Applied Probability, 39(3):731–752, September 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443538.

Ramaswami:1980:GQD

[Ram80] V. Ramaswami. The N/G/1 queue and its detailed analysis. Advances in Applied Probability, 12(1):222–261, March 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426503.

Ramalhoto:1984:BVB

[Ram84] M. F. Ramalhoto. Bounds for the variance of the busy period of the M/G/∞ queue. Advances in Applied Probability, 16(4): 929–932, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427350.

Ranneby:1977:HPS

[Ran77] Bo Ranneby. How to prove strong consistency of maximum likelihood estimates. Advances in Applied Probability, 9(2): 230, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426375.

Ranneby:1980:ETV

[Ran80] Bo Ranneby. The ergodic theorem for D_{q}-valued random elements. Advances in Applied Probability, 12(2):302–303, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426578.

Rao:1983:CSP

[Rao83] B. L. S. Prakasa Rao. Characterization of stochastic processes by stochastic integrals. Advances in Applied Probability, 15(1): 81–98, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1426983. See comment [RSS94].

Rao:2006:SNN

[Rao06] Suhasini Subba Rao. On some nonstationary, nonlinear random processes and their stationary approximations. Advances in Applied Probability, 38(4):1155–1172, December 2006. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/20443492.

Rath:1975:CQH

[Rat75] John H. Rath. Controlled queues in heavy traffic. Advances in Applied Probability, 7(3):656–671, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426134.

Rataj:1996:EOD

[Rat96a] Jan Rataj. Estimation of oriented direction distribution of a planar body. Advances in Applied Probability, 28(2):394–404, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428064.

Rataj:1996:MCM

[Rat96b] Jan Rataj. Mixed curvature measures and a translative integral formula. Advances in Applied Probability, 28(2):341, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428054.

Rataj:1998:DSA

[Rat98] Jan Rataj. Determination of spherical area measures by means of dilation volumes. Advances in Applied Probability, 30(2): 289, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427962.

Rataj:2004:EEN

[Rat04] J. Rataj. On estimation of the Euler number by projections of thin slabs. Advances in Applied Probability, 36(3):715–724, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
4140404.

Ravid:1977:CIA

[Rav77] Itzhak Ravid. Confidence intervals and acceptance tests. Advances in Applied Probability, 9(2):231, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426376.

Ravaska:1983:AMI

[Rav83] Timo Ravaska. On analytical methods for incomplete Markov random fields. Advances in Applied Probability, 15(1):99–112, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426984.

Rathbun:1994:APE

[RC94] Stephen L. Rathbun and Noel Cressie. Asymptotic properties of estimators for the parameters of spatial inhomogeneous Poisson point processes. Advances in Applied Probability, 26(1):122– 154, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427583.

Rehder:1980:ADR

[Reh80] Wulf Rehder. The asymptotic distribution of random molecules. Advances in Applied Probability, 12(3):640–654, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426424.

Reiss:1981:UAD

[Rei81] R.-D. Reiss. Uniform approximation to distributions of extreme order statistics. Advances in Applied Probability, 13(3):533–547, September 1981. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426784.

Reiman:1988:MFQ

[Rei88] Martin I. Reiman. A multiclass feedback queue in heavy traffic. Advances in Applied Probability, 20(1):179–207, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427275.

Rejmanek:1982:DSE

[Rej82] M. Rejmánek. Development of structure at the ecosystem level: Succession and stability. Advances in Applied Probability, 14 (2):200–201, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426512.

Remiche:2000:AIC

[Rem00] Marie-Ange Remiche. Asymptotic independence of counts in isotropic planar point processes of phase-type. Advances in Applied Probability, 32(2):363–375, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428193.

Renshaw:1986:SSS

[Ren86] Eric Renshaw. A survey of stepping-stone models in population dynamics. Advances in Applied Probability, 18(3):581–627, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427179.

Resnick:1972:SMR

[Res72] Sidney I. Resnick. Stability of maxima of random variables defined on a Markov chain. Advances in Applied Probability, 4 (2):285–295, August 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426000.

Resnick:1974:EPR

[Res74a] S. Resnick. Extremal processes and record value times. Advances in Applied Probability, 6(2):247–248, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426282.

Resnick:1974:IEP

[Res74b] Sidney I. Resnick. Inverses of extremal processes. Advances in Applied Probability, 6(2):392–405, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426300.

Resnick:1975:EP

[Res75] Sidney I. Resnick. Extremal processes. Advances in Applied Probability, 7(2):241, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426054.

Resnick:1980:WCP

[Res80] Sidney Resnick. Weak convergence, point processes and regular variation. Advances in Applied Probability, 12(2):296, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426571.

Resnick:1986:PPR

[Res86] Sidney I. Resnick. Point processes, regular variation and weak convergence. Advances in Applied Probability, 18(1):66–138, March 1986. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427239.

Reynolds:1975:CSQ

[Rey75] John F. Reynolds. The covariance structure of queues and related processes: A survey of recent work. Advances in Applied Probability, 7(2):383–415, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426082.

Remerova:2014:RFL

[RFZ14] Maria Remerova, Sergey Foss, and Bert Zwart. Random fluid limit of an overloaded polling model. Advances in Applied Probability, 46(1):76–101, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1396360104.

Ravichandran:1984:SCS

[RG84] N. Ravichandran and A. Gravey. Some comments on the simple queue. Advances in Applied Probability, 16(4):933–935, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427351.

Rangan:1989:RSS

[RG89a] A. Rangan and R. Ester Grace. On the replacement of systems subject to shocks and wear-dependent failure. Advances in Applied Probability, 21(1):238–240, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427211.

Rangan:1989:ORP

[RG89b] A. Rangan and R. Esther Grace. Optimal replacement policies for a deteriorating system with imperfect maintenance. Advances in Applied Probability, 21(4):949–951, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427781.

Robertson:1981:EIC

[RH81] Alan Robertson and W. G. Hill. The estimation of the inbreeding coefficient from deviations from Hardy–Weinberg equilibrium. Advances in Applied Probability, 13(1):5, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426457.

Rumsewicz:1989:IAD

[RH89] M. Rumsewicz and W. Henderson. Insensitivity with agedependent routing. Advances in Applied Probability, 21(2):398– 408, June 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427166.

Rice:1975:EFS

[Ric75] John Rice. Estimated factorisation of the spectral density of a stationary point process. Advances in Applied Probability, 7(4): 801–817, December 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426400.

Rice:1977:GSN

[Ric77] John Rice. On generalized shot noise. Advances in Applied Probability, 9(3):553–565, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426114.

Richard:2011:LTS

[Ric11] M. Richard. Limit theorems for supercritical age-dependent branching processes with neutral immigration. Advances in Applied Probability, 43(1):276–300, March 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/23024536.

Rieder:1975:BDP

[Rie75] Ulrich Rieder. Bayesian dynamic programming. Advances in Applied Probability, 7(2):330–348, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426080.

Riedel:1980:CSP

[Rie80] M. Riedel. Characterization of stable processes by identically distributed stochastic integrals. Advances in Applied Probability, 12(3):689–709, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426427.

Ripley:1977:PPS

[Rip77] B. D. Ripley. The point processes of stochastic geometry. Advances in Applied Probability, 9(3):434–435, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426096.

Rising:1991:AGI

[Ris91] William Rising. Applications of generalized inverses to Markov chains. Advances in Applied Probability, 23(2):293–302, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427749.

Ritter:1994:MST

[Rit94] Gunter Ritter. On the mean sojourn time of jobs in queues with general service disciplines. Advances in Applied Probability, 26 (2):516–538, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427449.

Rosberg:1989:CRD

[RK89] Zvi Rosberg and Parviz Kermani. Customer routing to different servers with complete information. Advances in Applied Probability, 21(4):861–882, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427771.

Risch:1977:LTR

[RL77] Neil Risch and Charles Louy. Laplace transforms for residual waiting times with an application in genetics. Advances in Applied Probability, 9(2):221–222, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426365.

Ramli:2010:SPD

[RL10] Muhamad Azfar Ramli and Gerard Leng. The stationary probability density of a class of bounded Markov processes. Advances in Applied Probability, 42(4):986–993, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/25799080.

Rahme:2021:SMM

[RM21] Youssef Rahme and Pascal Moyal. A stochastic matching model on hypergraphs. Advances in Applied Probability, 53 (4):951–980, December 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/stochastic-matching-model-on-hypergraphs/
4224343660486E48DE9489C729134439.

Resnick:1970:LLM

[RN70] Sidney I. Resnick and Marcel F. Neuts. Limit laws for maxima of a sequence of random variables defined on a Markov chain.

Advances in Applied Probability, 2(2):323–343, Autumn 1970. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426322.

Robertson:1974:VOM

[Rob74a] A. Robertson. The validity of the optimum model. Advances in Applied Probability, 6(1):17–18, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426203.

Robinson:1974:SDE

[Rob74b] P. M. Robinson. Stochastic difference equations with nonintegral differences. Advances in Applied Probability, 6(3):524– 545, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426232.

Robertson:1976:VGF

[Rob76a] A. Robertson. Variation in gene frequency over populations as a test for natural selection. Advances in Applied Probability, 8 (1):19–20, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426016.

Robinson:1976:MDC

[Rob76b] D. R. Robinson. Markov decision chains with unbounded costs and applications to the control of queues. Advances in Applied Probability, 8(1):159–176, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426027.

Robinson:1992:OFC

[Rob92] Derek Robinson. The optimality of fixed channel assignment policies for cellular radio systems. Advances in Applied Probability, 24(2):474–495, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427701.

Robert:2007:SSS

[Rob07] Christian Y. Robert. Stochastic stability of some statedependent growth-collapse processes. Advances in Applied Probability, 39(1):189–220, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443506.

Roberts:2018:PUL

[Rob18] Matthew I. Roberts. The probability of unusually large components in the near-critical Erdős–Rényi graph. Advances in Applied Probability, 50(1):245–271, March 2018. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/probability-
of-unusually-large-components-in-the-nearcritical-erdosrenyi-
graph/661C73E712AEB05AEE03E99A82764337.

Rodland:2006:EDW

[Rød06] Einar Andreas Rødland. Exact distribution of word counts in shuffled sequences. Advances in Applied Probability, 38(1):116– 133, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

20443431. See correction [Ano06f].

Roes:1973:TFQ

[Roe73] P. B. M. Roes. Two finite queues and their duals. Advances in Applied Probability, 5(1):22–24, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1425955.

Roes:1974:VVE

[Roe74] P. B. M. Roes. Variance at variance in eddy diffusion. Advances in Applied Probability, 6(2):224–225, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426262.

Rogers:1989:ITA

[Rog89] L. C. G. Rogers. Ignatov’s theorem: An abbreviation of the proof of Engelen, Tommassen and Vervaat. Advances in Applied Probability, 21(4):933–934, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427776.

Rogers:1990:TSE

[Rog90] L. C. G. Rogers. The two-sided exit problem for spectrally positive Lévy processes. Advances in Applied Probability, 22 (2):486–487, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427548.

Roginsky:1992:CLT

[Rog92] Allen L. Roginsky. A central limit theorem with remainder term for renewal processes. Advances in Applied Probability, 24(2):267–287, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427692.

Roginsky:1994:CLT

[Rog94] Allen L. Roginsky. A central limit theorem for cumulative processes. Advances in Applied Probability, 26(1):104–121, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427582.

Rolski:1978:RCP

[Rol78] Tomasz Rolski. A rate-conservative principle for stationary piecewise Markov processes. Advances in Applied Probability, 10(2):392–410, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426942.

Rolski:1981:QNS

[Rol81] Tomasz Rolski. Queues with non-stationary input stream: Ross’s conjecture. Advances in Applied Probability, 13(3):603– 618, September 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426787.

Rolski:1987:APQ

[Rol87] Tomasz Rolski. Approximation of periodic queues. Advances in Applied Probability, 19(3):691–707, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427413.

Rootzen:1978:LDE

[Roo78] Holger Rootzén. Limit distributions for the error in approximations of stochastic integrals. Advances in Applied Probability, 10 (2):290, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426892.

Rootzen:1983:RCE

[Roo83] Holger Rootzén. The rate of convergence of extremes of stationary normal sequences. Advances in Applied Probability, 15 (1):54–80, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426982.

Rootzen:1988:MES

[Roo88] Holger Rootzén. Maxima and exceedances of stationary Markov chains. Advances in Applied Probability, 20(2):371–390, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427395.

Roos:1993:CPA

[Roo93] Małgorzata Roos. Compound Poisson approximations for the numbers of extreme spacings. Advances in Applied Probability, 25(4):847–874, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427795.

Rosenlund:1973:LNS

[Ros73] Stig I. Rosenlund. On the length and number of served customers of the busy period of a generalised M/G/1 queue with finite waiting room. Advances in Applied Probability, 5(2):379– 389, August 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426042.

Ross:1974:DPG

[Ros74a] Sheldon M. Ross. Dynamic programming and gambling models. Advances in Applied Probability, 6(3):593–606, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426236.

Rossi:1974:SMG

[Ros74b] C. Rossi. Stochastic model of gene decay. Advances in Applied Probability, 6(2):236, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426273.

Rostron:1975:NLF

[Ros75] John Rostron. A non-linear function for maximizing the parentoffspring regression. Advances in Applied Probability, 7(1):14– 16, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425846.

Rosenlund:1976:BPT

[Ros76a] Stig. I. Rosenlund. Busy periods in time-dependent M/G/1 queues. Advances in Applied Probability, 8(1):195–208, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426029.

Rostron:1976:MSG

[Ros76b] J. Rostron. A multivariate study of the genetics of dermal ridges. Advances in Applied Probability, 8(1):20–22, March 1976. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426017.

Rossberg:1978:PLP

[Ros78] H.-J. Rossberg. Phragmén–Lindelöf (PL) theorems in probability theory. Advances in Applied Probability, 10(2):270–272, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426878.

Rostron:1979:GCD

[Ros79] John Rostron. Genetic colonising distance and a problem in algebra. Advances in Applied Probability, 11(1):7–8, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426761.

Rosenkrantz:1980:STV

[Ros80] Walter A. Rosenkrantz. Some theorems for the virtual waiting time process. Advances in Applied Probability, 12(2):310– 311, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426587.

Rostron:1981:GPE

[Ros81] J. Rostron. Genetic parameter estimation. A general-purpose maximum likelihood program. Advances in Applied Probability, 13(1):6, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426458.

Rosen:1989:CSL

[Ros89] Dennis Rosen. Coverage of a square lattice by an inclined rectangle. Advances in Applied Probability, 21(3):705–707, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427644.

Ross:2013:PLP

[Ros13] Nathan Ross. Power laws in preferential attachment graphs and Stein’s method for the negative binomial distribution. Advances in Applied Probability, 45(3):876–893, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1377868543.

Roy:2007:BSM

[Roy07] Emmanuel Roy. Bartlett spectrum and mixing properties of infinitely divisible random measures. Advances in Applied Probability, 39(4):893–897, December 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443550.

Rozanov:1978:SMF

[Roz78] Yu. A. Rozanov. Stochastic Markov fields. Advances in Applied Probability, 10(2):272–273, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426879.

Rozanov:1980:MMP

[Roz80] Yu. A. Rozanov. On multidirectional Markov processes. Advances in Applied Probability, 12(2):276, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426545.

Ramirez-Perez:2001:SNC

[RPS01] Filemon Ramirez-Perez and Robert Serfling. Shot noise on cluster processes with cluster marks, and studies of long range dependence. Advances in Applied Probability, 33(3):631–651, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428479.

Resnick:1973:SEPa

[RR73a] Sidney I. Resnick and Michael Rubinovitch. The structure of extremal processes. Advances in Applied Probability, 5(1):21– 22, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425954.

Resnick:1973:SEPb

[RR73b] Sidney I. Resnick and Michael Rubinovitch. The structure of extremal processes. Advances in Applied Probability, 5(2):287– 307, August 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426037.

Rachev:1990:ASC

[RR90a] S. T. Rachev and L. Rüschendorf. Approximation of sums by compound Poisson distributions with respect to stop-loss distances. Advances in Applied Probability, 22(2):350–374, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427540.

Resnick:1990:MEP

[RR90b] Sidney Resnick and Rishin Roy. Multivariate extremal processes, leader processes and dynamic choice models. Advances in Applied Probability, 22(2):309–331, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427538.

Rachev:1995:PMR

[RR95] S. T. Rachev and L. Rüschendorf. Probability metrics and recursive algorithms. Advances in Applied Probability, 27(3):770–799, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428133.

Rastegar:2021:ASM

[RR21] Reza Rastegar and Alexander Roitershtein. Avalanches in a short-memory excitable network. Advances in Applied Probability, 53(3):609–648, September 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/avalanches-in-a-shortmemory-excitable-
network/9CD370BDB41948E6CCDDEA80EB886CD0.

Reynolds:1971:RWM

[RS71] David S. Reynolds and I. Richard Savage. Random wear models in reliability theory. Advances in Applied Probability, 3(2):229– 248, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426169.

Rao:1986:RRC

[RS86] C. Radhakrishna Rao and D. N. Shanbhag. Recent results on characterization of probability distributions: A unified approach through extensions of Deny’s theorem. Advances in Applied Probability, 18(3):660–678, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427182. See comment

[RSS94].

Recsei:1987:BTI

[RS87] Elaine Recsei and E. Seneta. Bonferroni-type inequalities. Advances in Applied Probability, 19(2):508–511, June 1987. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427431.

Rubinstein:1988:CPA

[RS88] Reuven Y. Rubinstein and Ferenc Szidarovszky. Convergence of perturbation analysis estimates for discontinuous sample functions: A general approach. Advances in Applied Probability, 20 (1):59–78, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427270.

Righter:1992:EBC

[RS92] Rhonda Righter and J. George Shanthikumar. Extension of the bivariate characterization for stochastic orders. Advances in Applied Probability, 24(2):506–508, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427705.

Rachev:1995:LLS

[RS95] Svetlozar T. Rachev and Gennady Samorodnitsky. Limit laws for a stochastic process and random recursion arising in probabilistic modelling. Advances in Applied Probability, 27(1):185– 202, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428103.

Resnick:1997:SHE

[RS97] Sidney Resnick and Cătălin Stărică. Smoothing the Hill estimator. Advances in Applied Probability, 29(1):271–293, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427870.

Rue:1998:BOR

[RS98] Håvard Rue and Anne Randi Syversveen. Bayesian object recognition with Baddeley’s delta loss. Advances in Applied Probability, 30(1):64–84, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427876.

Roquain:2007:ICP

[RS07] Etienne Roquain and Sophie Schbath. Improved compound Poisson approximation for the number of occurrences of any rare word family in a stationary Markov chain. Advances in Applied Probability, 39(1):128–140, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/20443503.

Robert:2009:OSA

[RS09] Philippe Robert and Florian Simatos. Occupancy schemes associated to Yule processes. Advances in Applied Probability, 41 (2):600–622, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/27793892.

Riplinger:2012:APA

[RS12] M. Riplinger and M. Spiess. Asymptotic properties of the approximate inverse estimator for directional distributions. Advances in Applied Probability, 44(4):954–976, December 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/41805986.

Rychlik:2021:LSR

[RS21] Tomasz Rychlik and Fabio Spizzichino. Load-sharing reliability models with different component sensitivities to other components’ working states. Advances in Applied Probability, 53(1):107–132, March 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-

probability/article/loadsharing-reliability-models-with-
different-component-sensitivities-to-other-components-
working-states/73B329ED4CCB97AA2040F0A673CCEA08.

Rao:1994:ICF

[RSS94] C. R. Rao, T. Sapatinas, and D. N. Shanbhag. The integrated Cauchy functional equation: Some comments on recent papers. Advances in Applied Probability, 26(3):825–829, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427823.

See [Rao83, RS86].

Rolski:1999:APS

[RSS99] Tomasz Rolski, Sabine Schlegel, and Volker Schmidt. Asymptotics of Palm-stationary buffer content distributions in fluid flow queues. Advances in Applied Probability, 31(1):235–253, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428000.

Ruschendorf:2016:CTI

[RSW16] Rudger Rüschendorf, Alexander Schnurr, and Victor Wolf. Comparison of time-inhomogeneous Markov processes. Advances in Applied Probability, 48(4):1015–1044, December 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1482548427.

Ren:2023:BBM

[RSY23] Yan-Xia Ren, Renming Song, and Fan Yang. Branching Brownian motion in a periodic environment and uniqueness of pulsating traveling waves. Advances in Applied Probability, 55(2):510–548, June 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/branching-brownian-motion-in-a-periodic-
environment-and-uniqueness-of-pulsating-traveling-waves/
916770E56A68B10BA1C2F7C303C61B8E.

Reitzner:2012:SRV

[RSZ12] M. Reitzner, E. Spodarev, and D. Zaporozhets. Set reconstruction by Voronoi cells. Advances in Applied Probability, 44(4): 938–953, December 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/41805985.

Ramanan:2002:MRF

[RSZM02] Kavita Ramanan, Anirvan Sengupta, Ilze Ziedins, and Partha Mitra. Markov random field models of multicasting in tree networks. Advances in Applied Probability, 34(1):58–84, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428372.

Ross:1990:TEP

[RT90] Keith W. Ross and Danny Tsang. Teletraffic engineering for product-form circuit-switched networks. Advances in Applied Probability, 22(3):657–675, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427463.

Roy:2002:CIB

[RT02] Rahul Roy and Hideki Tanemura. Critical intensities of Boolean models with different underlying convex shapes. Advances in Applied Probability, 34(1):48–57, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428371.

Rigal:2007:NGA

[RT07] L. Rigal and L. Truffet. A new genetic algorithm specifically based on mutation and selection. Advances in Applied Probability, 39(1):141–161, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443504.

Rosenbaum:2021:MPD

[RT21] Mathieu Rosenbaum and Mehdi Tomas. From microscopic price dynamics to multidimensional rough volatility models. Advances in Applied Probability, 53(2):425–462, June 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/from-microscopic-price-dynamics-
to-multidimensional-rough-volatility-models/1F4A19A5F9EF8B04A39BC6CAC5CCC65B.

Rasonyi:2023:CKW

[RT23] Miklós Rásonyi and Kinga Tikosi. Convergence of the Kiefer– Wolfowitz algorithm in the presence of discontinuities. Advances in Applied Probability, 55(2):382–406, June 2023. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/convergence-
of-the-kieferwolfowitz-algorithm-in-the-presence-of-
discontinuities/EB045FE2C64D9B4C60DBD54A598F7E61.

Rubinovitch:1974:QRS

[Rub74] Michael Rubinovitch. Queues with random service output. Advances in Applied Probability, 6(2):207–208, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426249.

Rubinovitch:1978:SRP

[Rub78] Michael Rubinovitch. On subordinates of regenerative phenomena. Advances in Applied Probability, 10(2):273–274, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426880.

Ruben:1981:LCS

[Rub81] Harold Ruben. The lifetimes of configuration states in statistical physics. Advances in Applied Probability, 13(4):704–719, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1426968. See correction [Ano83e].

Rudemo:1972:DSP

[Rud72] Mats Rudemo. Doubly stochastic Poisson processes and process control. Advances in Applied Probability, 4(2):318–338, August 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426002.

Rudemo:1973:PPG

[Rud73] Mats Rudemo. Point processes generated by transitions of Markov chains. Advances in Applied Probability, 5(2):262–286, August 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426036.

Rudemo:1974:CPD

[Rud74] M. Rudemo. Conditional probability distributions for partially observed Markov chains. Advances in Applied Probability, 6 (2):250–251, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426285.

Rudolph:1998:CLT

[Rud98] Andreas Rudolph. A central limit theorem for random coefficient autoregressive models and ARCH/GARCH models. Advances in Applied Probability, 30(1):113–121, March 1998. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427878.

Rukhin:2002:DNW

[Ruk02] Andrew L. Rukhin. Distribution of the number of words with a prescribed frequency and tests of randomness. Advances in Applied Probability, 34(4):775–797, December 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428322.

Ruschendorf:1982:RVM

[Rüs82] Ludger Rüschendorf. Random variables with maximum sums. Advances in Applied Probability, 14(3):623–632, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426677.

Ruschendorf:1991:CSO

[Rüs91] Ludger Rüschendorf. On conditional stochastic ordering of distributions. Advances in Applied Probability, 23(1):46–63, March 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427511.

Rusinek:2010:MRH

[Rus10] Anna Rusinek. Mean reversion for HJMM forward rate models. Advances in Applied Probability, 42(2):371–391, June 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/25683825.

Robbins:1977:OSP

[RW77] Herbert Robbins and John Whitehead. Order statistics and the Pólya urn process. Advances in Applied Probability, 9(2):205– 206, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426350.

Roberts:1984:RR

[RW84] P. H. Roberts and D. E. Winch. On random rotations. Advances in Applied Probability, 16(3):638–655, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427291.

Roehner:1985:AIP

[RW85] Bertrand Roehner and Peter Winiwarter. Aggregation of independent Paretian random variables. Advances in Applied Probability, 17(2):465–469, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427153.

Ritter:1991:MWT

[RW91] Gunter Ritter and Ulrich Wacker. The mean waiting time in a G/G/m/∞ queue with the LCFS-P service discipline. Advances in Applied Probability, 23(2):406–428, June 1991. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427756.

Raoult:2003:RCG

[RW03] J.-P. Raoult and R. Worms. Rate of convergence for the generalized Pareto approximation of the excesses. Advances in Applied Probability, 35(4):1007–1027, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428497.

Rieder:2014:OTW

[RW14] Ulrich Rieder and Marc Wittlinger. On optimal terminal wealth problems with random trading times and drawdown constraints. Advances in Applied Probability, 46(1):121–138, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1396360106.

Righter:1991:SJN

[RX91] Rhonda Righter and Susan H. Xu. Scheduling jobs on nonidentical IFR processors to minimize general cost functions. Advances in Applied Probability, 23(4):909–924, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427683.

Ren:2014:MBB

[RY14] Yan-Xia Ren and Ting Yang. Multitype branching Brownian motion and traveling waves. Advances in Applied Probability, 46 (1):217–240, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1396360111.

Rychlik:1987:RAW

[Ryc87] Igor Rychlik. Regression approximations of wavelength and amplitude distributions. Advances in Applied Probability, 19 (2):396–430, June 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427425.

Rychlik:1992:TBP

[Ryc92] Igor Rychlik. The two-barrier problem for continuously differentiable processes. Advances in Applied Probability, 24(1):71– 94, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427730.

Resnick:2013:AMK

[RZ13] Sidney I. Resnick and David Zeber. Asymptotics of Markov kernels and the tail chain. Advances in Applied Probability, 45 (1):186–213, March 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/41806005.

Sacerdote:1990:SFP

[Sac90] L. Sacerdote. On the solution of the Fokker–Planck equation for a Feller process. Advances in Applied Probability, 22(1):101– 110, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427599.

Sadowsky:1995:PLQb

[Sad95] John S. Sadowsky. The probability of large queue lengths and waiting times in a heterogeneous multiserver queue II: Positive recurrence and logarithmic limits. Advances in Applied Probability, 27(2):567–583, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427839.

Sakurai:2004:NIL

[Sak04] T. Sakurai. Numerical inversion for Laplace transforms of functions with discontinuities. Advances in Applied Probability, 36 (2):616–642, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428469.

Sakagawa:2015:PPC

[Sak15] Hironobu Sakagawa. Persistence probability for a class of Gaussian processes related to random interface models. Advances in Applied Probability, 47(1):146–163, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1427814585.

Salminen:1984:COU

[Sal84] P. Salminen. On conditional Ornstein–Uhlenbeck processes. Advances in Applied Probability, 16(4):920–922, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427347.

Salminen:1988:FHT

[Sal88] Paavo Salminen. On the first hitting time and the last exit time for a Brownian motion to/from a moving boundary. Advances in Applied Probability, 20(2):411–426, June 1988. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427397.

Samorodnitsky:2004:MCT

[Sam04a] Gennady Samorodnitsky. Maxima of continuous-time stationary stable processes. Advances in Applied Probability, 36(3): 805–823, September 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/4140409.

Samuels:2004:WDQ

[Sam04b] Stephen M. Samuels. Why do these quite different best-choice problems have the same solutions? Advances in Applied Probability, 36(2):398–416, June 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428459.

Sandau:1998:EPP

[San98] Konrad Sandau. Estimating parameters of point process models using the fL-function. Advances in Applied Probability, 30 (2):289–290, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427963.

Sarkar:1997:CEO

[Sar97] Anish Sarkar. Co-existence of the occupied and vacant phase in Boolean models in three or more dimensions. Advances in Applied Probability, 29(4):878–889, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427845.

Sarkar:1998:FCH

[Sar98] Anish Sarkar. Finite clusters in high density Boolean models with balls of varying sizes. Advances in Applied Probability, 30 (4):929–947, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428023.

Sarkar:2016:SCT

[Sar16] Amites Sarkar. Secrecy coverage in two dimensions. Advances in Applied Probability, 48(1):1–12, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1457466152.

Sato:1978:CDM

[Sat78] Ken-Iti Sato. Convergence to a diffusion of a multi-allelic model in population genetics. Advances in Applied Probability, 10(3): 538–562, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426633.

Sathe:1984:CSO

[Sat84] Y. S. Sathe. A comment on star-ordering and tail-ordering. Advances in Applied Probability, 16(3):692, September 1984. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427296.

Sathe:1985:IBV

[Sat85] Y. S. Sathe. Improved bounds for the variance of the busy period of the M/G/∞ queue. Advances in Applied Probability, 17(4):

913–914, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427096.

Saunders:1978:LBS

[Sau78] I. W. Saunders. Locating bright spots in a point process. Advances in Applied Probability, 10(3):587–612, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426636.

Saunders:1984:DOD

[Sau84] D. J. Saunders. Dispersive ordering of distributions. Advances in Applied Probability, 16(3):693–694, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427297.

Savits:1975:AST

[Sav75a] T. H. Savits. Applications for space-time harmonic functions to branching processes. Advances in Applied Probability, 7(2):251– 252, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426065.

Savits:1975:STH

[Sav75b] Thomas H. Savits. Space-time harmonic functions and agedependent branching processes. Advances in Applied Probability, 7(2):283–298, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426078.

Sawyer:1976:BDP

[Saw76] Stanley Sawyer. Branching diffusion processes in population genetics. Advances in Applied Probability, 8(4):659–689, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425929.

Sawyer:1977:APE

[Saw77] Stanley Sawyer. Asymptotic properties of the equilibrium probability of identity in a geographically structured population. Advances in Applied Probability, 9(2):268–282, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426386.

Schuh:1977:ABB

[SB77] H.-J. Schuh and A. D. Barbour. On the asymptotic behaviour of branching processes with infinite mean. Advances in Applied Probability, 9(4):681–723, December 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426697.

Schladitz:1998:TOP

[SB98] Katja Schladitz and Adrian J. Baddeley. A third order point process characteristic. Advances in Applied Probability, 30(2): 290, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427964.

Stefanov:2009:RDA

[SB09] Valeri T. Stefanov and Frank Ball. Reward distributions associated with some block tridiagonal transition matrices with applications to identity by descent. Advances in Applied Probability, 41(2):523–545, June 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/27793889.

Stramer:1996:CTT

[SBT96] O. Stramer, P. J. Brockwell, and R. L. Tweedie. Continuoustime threshold AR(1) processes. Advances in Applied Probability, 28(3):728–746, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428178.

Shahani:1977:TMS

[SC77] A. K. Shahani and D. M. Crease. Towards models of screening for early detection of disease. Advances in Applied Probability, 9(4):665–680, December 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426696.

Samuel-Cahn:1983:SAE

[SC83] Ester Samuel-Cahn. Simple approximations to the expected waiting time for a cluster of any given size, for point processes. Advances in Applied Probability, 15(1):21–38, March 1983. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426980.

Senoussi:2000:WHP

[SCA00] R. Senoussi, J. Chadoeuf, and D. Allard. Weak homogenization of point processes by space deformations. Advances in Applied Probability, 32(4):948–959, December 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428509.

Schal:1971:AQS

[Sch71] Manfred Schäl. The analysis of queues with state-dependent parameters by Markov renewal processes. Advances in Applied Probability, 3(1):155–175, Spring 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1426332.

Schal:1973:DPU

[Sch73] Manfred Schäl. Dynamic programming under continuity and compactness assumptions. Advances in Applied Probability, 5 (1):24–25, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425956.

Schassberger:1974:BAS

[Sch74a] R. Schassberger. A broad analysis of single server priority queues with two independent input streams, one of them Poisson. Advances in Applied Probability, 6(4):666–688, December 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426186.

Schmetterer:1974:PGT

[Sch74b] L. K. Schmetterer. Problems of group theory related to probability theory. Advances in Applied Probability, 6(2):209–210, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426250.

Schassberger:1976:LDG

[Sch76a] R. Schassberger. On the limiting distribution of generalized semi-Markov processes. Advances in Applied Probability, 8(2): 245–246, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425895.

Schweitzer:1976:BOP

[Sch76b] Paul Schweitzer. Buffer overflow probabilities using an infinitecapacity model. Advances in Applied Probability, 8(2):244, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425893.

Schwarz:1977:MAE

[Sch77] Gideon Schwarz. Mutual applications of ergodic theory and stochastic processes. Advances in Applied Probability, 9(2):206– 207, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426351.

Schassberger:1978:ISP

[Sch78a] R. Schassberger. The insensitivity of stationary probabilities in networks of queues. Advances in Applied Probability, 10(4): 906–912, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/ stable/1426666. See correction [Ano80e].

Schassberger:1978:ISS

[Sch78b] R. Schassberger. Insensitivity of steady-state distributions of generalized semi-Markov processes with speeds. Advances in Applied Probability, 10(4):836–851, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426662.

Schuh:1978:ABB

[Sch78c] H.-J. Schuh. On the asymptotic behaviour of branching processes with infinite mean. Advances in Applied Probability, 10 (2):305–307, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426906.

Schuh:1979:SRS

[Sch79] H.-J. Schuh. Some results on the structure of irregular branching processes with infinite mean. Advances in Applied Probability, 11(2):292–293, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426823.

Schurger:1981:CBP

[Sch81] Klaus Schürger. A class of branching processes on a lattice with interactions. Advances in Applied Probability, 13(1):14– 39, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426465.

Schuh:1982:SCS

[Sch82a] H.-J. Schuh. Seneta constants for the supercritical Bellman– Harris process. Advances in Applied Probability, 14(4):732–751, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427021.

Schuster:1982:CCB

[Sch82b] P. Schuster. Competition and cooperation of biological macromolecules: An experimentally testable model of Darwinian evolution. Advances in Applied Probability, 14(2):201–202, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426513.

Schassberger:1984:NAM

[Sch84] R. Schassberger. A new approach to the M/G/1 processorsharing queue. Advances in Applied Probability, 16(1):202–213, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427231.

Schassberger:1986:TRI

[Sch86a] R. Schassberger. Two remarks on insensitive stochastic models. Advances in Applied Probability, 18(3):791–814, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427188.

Schilling:1986:MSN

[Sch86b] M. F. Schilling. Mutual and shared neighbor probabilities: Finite- and infinite-dimensional results. Advances in Applied Probability, 18(2):388–405, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1427305.

Schmidt:1987:JQL

[Sch87] Volker Schmidt. On joint queue-length characteristics in infiniteserver tandem queues with heavy traffic. Advances in Applied Probability, 19(2):474–486, June 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427428.

Schassberger:1990:SSA

[Sch90] R. Schassberger. The steady-state appearance of the M/G/1 queue under the discipline of shortest remaining processing time. Advances in Applied Probability, 22(2):456–479, June 1990. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427545.

Scheike:1994:AGV

[Sch94] Thomas H. Scheike. Anisotropic growth of Voronoi cells. Advances in Applied Probability, 26(1):43–53, March 1994. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427577.

Schladitz:1996:EIS

[Sch96] Katja Schladitz. Estimation of the intensity of stationary flat processes. Advances in Applied Probability, 28(2):342, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428056.

Schladitz:1999:SOE

[Sch99a] Katja Schladitz. Surprising optimal estimators for the area fraction. Advances in Applied Probability, 31(4):995–1001, December 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428339.

Schneider:1999:DPF

[Sch99b] Rolf Schneider. A duality for Poisson flats. Advances in Applied Probability, 31(1):63–68, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427991.

Schladitz:2000:EIS

[Sch00] Katja Schladitz. Estimation of the intensity of stationary flat processes. Advances in Applied Probability, 32(1):114–139, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428230.

Schneider:2001:MNM

[Sch01] Rolf Schneider. On the mean normal measures of a particle process. Advances in Applied Probability, 33(1):25–38, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428439.

Schreiber:2002:VAC

[Sch02] Tomasz Schreiber. Variance asymptotics and central limit theorems for volumes of unions of random closed sets. Advances in Applied Probability, 34(3):520–539, September 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428245.

Schneider:2003:NPH

[Sch03a] Rolf Schneider. Nonstationary Poisson hyperplanes and their induced tessellations. Advances in Applied Probability, 35(1): 139–158, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428277.

Schreiber:2003:AGH

[Sch03b] Tomasz Schreiber. Asymptotic geometry of high-density smooth-grained Boolean models in bounded domains. Advances in Applied Probability, 35(4):913–936, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428492.

Schroder:2003:IGB

[Sch03c] Michael Schröder. On the integral of geometric Brownian motion. Advances in Applied Probability, 35(1):159–183, March 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428278.

Schreiber:2005:RDT

[Sch05] Tomasz Schreiber. Random dynamics and thermodynamic limits for polygonal Markov fields in the plane. Advances in Applied Probability, 37(4):884–907, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/30037365.

Schroder:2006:LHDa

[Sch06a] Michael Schröder. On ladder height densities and Laguerre series in the study of stochastic functionals. I. Basic methods and results. Advances in Applied Probability, 38(4):969–994, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/20443484.

Schroder:2006:LHDb

[Sch06b] Michael Schröder. On ladder height densities and Laguerre series in the study of stochastic functionals. II. Exponential functionals of Brownian motion and Asian option values. Advances in Applied Probability, 38(4):995–1027, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443485.

Schroder:2007:CTM

[Sch07] Michael Schröder. Continuous-time methods in the study of discretely sampled functionals of Lévy processes. I. The positive process case. Advances in Applied Probability, 39(1):245– 270, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443508.

Schneider:2009:WFP

[Sch09] Rolf Schneider. Weighted faces of Poisson hyperplane tessellations. Advances in Applied Probability, 41(3):682–694, September 2009. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
27793899.

Schrempp:2016:LDL

[Sch16] Michael Schrempp. The limit distribution of the largest interpoint distance for distributions supported by a d-dimensional ellipsoid and generalizations. Advances in Applied Probability, 48(4):1256–1270, December 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1482548437.

Scott:1973:CLT

[Sco73] D. J. Scott. Central limit theorems for martingales and for processes with stationary increments using a Skorokhod representation approach. Advances in Applied Probability, 5(1):119– 137, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425967.

Simo:1998:MMI

[SD98] A. Simó and E. De Ves. Markov models in image analysis. An ophthalmic application. Advances in Applied Probability, 30 (2):290–291, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427965.

Stern:1991:ASM

[SE91] Thomas E. Stern and Anwar I. Elwalid. Analysis of separable Markov-modulated rate models for information-handling systems. Advances in Applied Probability, 23(1):105–139, March 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427514.

Segers:2003:FCE

[Seg03] Johan Segers. Functionals of clusters of extremes. Advances in Applied Probability, 35(4):1028–1045, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428498.

Segers:2020:OVM

[Seg20] Johan Segers. One- versus multi-component regular variation and extremes of Markov trees. Advances in Applied Probability, 52(3):855–878, September 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/one-versus-multicomponent-regular-
variation-and-extremes-of-markov-trees/F924304944926FA9783FD05B05D4DE8F.

Sekine:2006:NLT

[Sek06] Jun Sekine. A note on long-term optimal portfolios under drawdown constraints. Advances in Applied Probability, 38(3):673– 692, September 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443466.

Seleznjev:1996:LDP

[Sel96] Oleg Seleznjev. Large deviations in the piecewise linear approximation of Gaussian processes with stationary increments.

Advances in Applied Probability, 28(2):481–499, June 1996. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428068.

Seneta:1969:FEG

[Sen69] E. Seneta. Functional equations and the Galton–Watson process. Advances in Applied Probability, 1(1):1–42, Spring 1969. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426407.

Seneta:1974:RVF

[Sen74] E. Seneta. Regularly varying functions in the theory of simple branching processes. Advances in Applied Probability, 6(3): 408–420, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426224.

Seneta:1978:RVG

[Sen78] E. Seneta. A relaxation view of a genetic problem. Advances in Applied Probability, 10(4):716–720, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426650.

Seneta:1979:CESa

[Sen79a] E. Seneta. Coefficients of ergodicity: Structure and applications. Advances in Applied Probability, 11(2):270–271, June 1979. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426800.

Seneta:1979:CESb

[Sen79b] E. Seneta. Coefficients of ergodicity: Structure and applications. Advances in Applied Probability, 11(3):576–590, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426955.

Seneta:1980:CSD

[Sen80] E. Seneta. Computing the stationary distribution for infinite Markov chains. Advances in Applied Probability, 12(2):276– 277, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426546.

Seneta:1986:IG

[Sen86] E. Seneta. An inequality from genetics. Advances in Applied Probability, 18(3):860–861, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427191.

Seneta:1988:PSD

[Sen88] E. Seneta. Perturbation of the stationary distribution measured by ergodicity coefficients. Advances in Applied Probability, 20 (1):228–230, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427277.

Sengupta:1989:IRG

[Sen89a] Bhaskar Sengupta. An invariance relationship for the G/G/1 queue. Advances in Applied Probability, 21(4):956–957, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427783.

Sengupta:1989:MPW

[Sen89b] Bhaskar Sengupta. Markov processes whose steady state distribution is matrix-exponential with an application to the GI/PH/1 queue. Advances in Applied Probability, 21(1):159– 180, March 1989. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427202.

Sennott:1997:CAO

[Sen97] Linn I. Sennott. The computation of average optimal policies in denumerable state Markov decision chains. Advances in Applied Probability, 29(1):114–137, March 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427863.

Seneta:2018:PJS

[Sen18a] Eugene Seneta. Peter Jagers: some reminiscences on his retirement. Advances in Applied Probability, 50(A):11–12, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/peter-
jagers-some-reminiscences-on-his-retirement/3D7F51BDD912300C221151A067585CBD.

Seneta:2018:WLL

[Sen18b] Eugene Seneta. The weak law of large numbers for nonnegative summands. Advances in Applied Probability, 50 (A):241–252, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/weak-law-of-large-numbers-for-nonnegative-
summands/5442E920C4949546A73A37AB6AC05744.

Serfozo:1975:FLT

[Ser75] Richard F. Serfozo. Functional limit theorems for stochastic processes based on embedded processes. Advances in Applied Probability, 7(1):123–139, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425856.

Serfozo:1980:TRC

[Ser80] Richard Serfozo. Thinnings of random cluster fields. Advances in Applied Probability, 12(2):294–295, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426568.

Serfozo:1981:OCR

[Ser81] Richard Serfozo. Optimal control of random walks, birth and death processes, and queues. Advances in Applied Probability, 13(1):61–83, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426467.

Serfozo:1986:HSR

[Ser86] Richard F. Serfozo. Heredity of stationary and reversible stochastic processes. Advances in Applied Probability, 18(2): 574–576, June 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427313.

Serfozo:1989:PFM

[Ser89] Richard F. Serfozo. Poisson functionals of Markov processes and queueing networks. Advances in Applied Probability, 21(3): 595–611, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427638.

Serfozo:2005:RMP

[Ser05] Richard F. Serfozo. Reversible Markov processes on general spaces and spatial migration processes. Advances in Applied Probability, 37(3):801–818, September 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/30037356.

Severo:1971:MRS

[Sev71] Norman C. Severo. Multidimensional right-shift processes. Advances in Applied Probability, 3(2):200–201, Autumn 1971. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426155.

Schweitzer:1979:GCV

[SF79] P. J. Schweitzer and A. Federgruen. Geometric convergence of value-iteration in multichain Markov decision problems. Advances in Applied Probability, 11(1):188–217, March 1979. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426774. See correction

[Ano79f].

Singh:1992:PKC

[SG92] Harshinder Singh and Rameshwar D. Gupta. On the probability that the kth customer finds an M/M/1 queue empty. Advances in Applied Probability, 24(1):238–239, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427743.

Steel:1996:CLT

[SGW96] Mike Steel, Larry Goldstein, and Michael S. Waterman. A central limit theorem for the parsimony length of trees. Advances in Applied Probability, 28(4):1051–1071, December 1996. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428164.

Slud:1989:MLD

[SH89] Eric Slud and Craig Hoesman. Moderate- and large-deviation probabilities in actuarial risk theory. Advances in Applied Probability, 21(4):725–741, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427763.

Sharpe:1978:CPG

[Sha78a] K. Sharpe. The crossings process generated by a stationary χ^{2} process. Advances in Applied Probability, 10(2):291, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426894.

Sharpe:1978:SPC

[Sha78b] Ken Sharpe. Some properties of the crossings process generated by a stationary χ^{2} process. Advances in Applied Probability, 10 (2):373–391, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426941.

Sharpe:1979:UUT

[Sha79] Ken Sharpe. On the use of upcrossings to test for a Gaussian process. Advances in Applied Probability, 11(2):291, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426821.

Sharpe:1980:MPC

[Sha80] M. J. Sharpe. Markov properties of co-optional times. Advances in Applied Probability, 12(2):277–278, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426547.

Shanthikumar:1984:PNB

[Sha84] J. G. Shanthikumar. Processes with new better than used firstpassage times. Advances in Applied Probability, 16(3):667–686, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427293.

Shanthikumar:1987:SMR

[Sha87] J. George Shanthikumar. Stochastic majorization of random variables with proportional equilibrium rates. Advances in Applied Probability, 19(4):854–872, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427105.

Sharma:1997:AGD

[Sha97] Vinod Sharma. Approximations of general discrete time queues by discrete time queues with arrivals modulated by finite chains.

Advances in Applied Probability, 29(4):1039–1059, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427853.

Shah:2003:NSD

[Sha03] Sandeep R. Shah. A note on stochastic domination and conditional thinning. Advances in Applied Probability, 35(4):937–940, December 2003. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428493.

Shank:2009:NNG

[Sha09] Nathan Shank. Nearest-neighbor graphs on the Cantor set. Advances in Applied Probability, 41(1):38–62, March 2009. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/27793864.

Sheng:1985:DBT

[She85] T. K. Sheng. The distance between two random points in plane regions. Advances in Applied Probability, 17(4):748–773, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427086.

Shimizu:1979:SSM

[Shi79] Akinobu Shimizu. The stepping-stone model and a stochastic differential equation. Advances in Applied Probability, 11 (2):293–294, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426824.

Shirayayev:1980:SLT

[Shi80] A. N. Shirayayev. Some limit theorems for simple point processes: Martingale approach. Advances in Applied Probability, 12(2):278–279, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426548.

Shiga:1982:WPI

[Shi82] Tokuzo Shiga. Wandering phenomena in infinite-allelic diffusion models. Advances in Applied Probability, 14(3):457–483, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426669.

Shimkin:1993:ELD

[Shi93] Nahum Shimkin. Extremal large deviations in controlled i.i.d. processes with applications to hypothesis testing. Advances in Applied Probability, 25(4):875–894, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427796.

Shorrock:1974:DTE

[Sho74] R. W. Shorrock. On discrete time extremal processes. Advances in Applied Probability, 6(3):580–592, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426235.

Shorack:1979:EDE

[Sho79] Galen R. Shorack. Extensions of the Darling and Erdős theorem on the maximum of normalized sums. Advances in Applied Probability, 11(2):288–289, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426818.

Siegel:1976:ABD

[Sie76] Martha J. Siegel. The asymptotic behavior of a divergent linear birth and death process. Advances in Applied Probability, 8 (2):315–338, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425907.

Siegmund:1979:CDA

[Sie79] D. Siegmund. Corrected diffusion approximations in certain random walk problems. Advances in Applied Probability, 11(4): 701–719, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/ stable/1426855. See comment [Hog86].

Sigman:1991:NSP

[Sig91] Karl Sigman. A note on a sample-path rate conservation law and its relationship with H = λG. Advances in Applied Probability, 23(3):662–665, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427629.

Silverman:1976:LTD

[Sil76] B. W. Silverman. Limit theorems for dissociated random variables. Advances in Applied Probability, 8(4):806–819, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425935.

Simon:1980:EMR

[Sim80] Burton Simon. Equivalent Markov-renewal processes. Advances in Applied Probability, 12(2):293, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426565.

Simpson:1981:AWT

[Sim81] S. P. Simpson. Ascertainment when the trait has a variable age of onset. Advances in Applied Probability, 13(1):6–7, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426459.

Simatos:2015:SSC

[Sim15] Florian Simatos. State space collapse for critical multistage epidemics. Advances in Applied Probability, 47(3):715–740, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://projecteuclid.org/
euclid.aap/1444308879.

Sin:1998:CSV

[Sin98] Carlos A. Sin. Complications with stochastic volatility models. Advances in Applied Probability, 30(1):256–268, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427887.

Singh:1989:PSP

[SJ89] Harshinder Singh and Kanchan Jain. Preservation of some partial orderings under Poisson shock models. Advances in Applied Probability, 21(3):713–716, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427647.

Sagitov:2018:RDG

[SJ18] Serik Sagitov and Jonas Jagers. Rank-dependent Galton– Watson processes and their pathwise duals. Advances in Applied Probability, 50(A):229–239, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/rankdependent-galtonwatson-
processes-and-their-pathwise-duals/CE6B2EA6102C75C9E56C62278E7BD4E0.

Slater:1973:SSM

[SK73] N. B. Slater and T. C. T. Kotiah. The steady state of a multiserver mixed queue. Advances in Applied Probability, 5(3):614– 631, December 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425837.

Sumita:1985:BLT

[SK85] Ushio Sumita and Masaaki Kijima. The bivariate Laguerre transform and its applications: Numerical exploration of bivariate processes. Advances in Applied Probability, 17(4):683–708, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427083.

Sandau:1996:SPA

[SK96] Konrad Sandau and Haymo Kurz. Some properties of a 2D arterial tree simulated by a stochastic growth model. Advances in Applied Probability, 28(2):341–342, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428055.

Srivastava:2004:BGS

[SK04] Anuj Srivastava and Eric Klassen. Bayesian and geometric subspace tracking. Advances in Applied Probability, 36(1):43– 56, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428352.

Sohn:2008:DWS

[SK08] Eunju Sohn and Charles Knessl. The distribution of wasted spaces in the M/M/∞ queue with ranked servers. Advances in Applied Probability, 40(3):835–855, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443610.

Scheinhardt:2016:CRW

[SK16] Werner R. W. Scheinhardt and Dirk P. Kroese. A comparison of random walks in dependent random environments. Advances in Applied Probability, 48(1):199–214, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1457466162.

Stehr:2020:AVN

[SK20] Mads Stehr and Markus Kiderlen. Asymptotic variance of Newton–Cotes quadratures based on randomized sampling points. Advances in Applied Probability, 52(4):1284–1307, December 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/asymptotic-
variance-of-newtoncotes-quadratures-based-on-randomized-
sampling-points/1F37E1B907A1472811288D996A3B5BF9.

Saunders:1980:PLH

[SKF80] Roy Saunders, Richard J. Kryscio, and Gerald M. Funk. Poisson limits for a hard-core clustering model. Advances in Applied Probability, 12(2):294, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426567.

Saunders:1980:QCR

[SL80] Roy Saunders and Purushottam Laud. Quantile and contour regions for multivariate distributions. Advances in Applied Probability, 12(2):316, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426596.

Slud:1984:MDR

[Slu84] Eric Slud. Multivariate dependent renewal processes. Advances in Applied Probability, 16(2):347–362, June 1984. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427073.

Sivazlian:1978:GRM

[SM78] B. D. Sivazlian and J. F. Mahoney. Group replacement of a multicomponent system which is subject to deterioration only. Advances in Applied Probability, 10(4):867–885, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426664.

Sumita:1987:CPD

[SM87] Ushio Sumita and Yasushi Masuda. Classes of probability density functions having Laplace transforms with negative zeros and poles. Advances in Applied Probability, 19(3):632–651, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427410.

Small:1976:NMP

[Sma76] Robert D. Small. A non-Markovian population growth model with a finite limit distribution and bounded variance. Advances in Applied Probability, 8(2):229–230, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425877.

Sun:2017:MOD

[SMAL17] Yunpeng Sun, Rafael Mendoza-Arriaga, and Vadim Linetsky. Marshall-Olkin distributions, subordinators, efficient simulation, and applications to credit risk. Advances in Applied Probability, 49(2):481–514, June 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/marshallolkin-distributions-subordinators-
efficient-simulation-and-applications-to-credit-risk/
25756CCEE3844AAE5703C85287978C0E.

Smith:1974:SRW

[Smi74] W. L. Smith. Some recent work on branching processes in random environments. Advances in Applied Probability, 6(2):210– 212, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426251.

Smith:1976:MCG

[Smi76] Cedric A. B. Smith. Matrix calculations of genetic probabilities. Advances in Applied Probability, 8(4):641–643, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425924.

Smith:1979:SHQ

[Smi79] Cedric A. B. Smith. Searching for heterozygotes in quantitative characters. Advances in Applied Probability, 11(1):8–9, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426762.

Smith:1980:ECB

[Smi80] Cedric A. B. Smith. Estimation of the correlation between relatives. Advances in Applied Probability, 12(1):8–9, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426486.

Smith:1982:URC

[Smi82] Richard L. Smith. Uniform rates of convergence in extremevalue theory. Advances in Applied Probability, 14(3):600–622, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426676.

Smith:1983:LTA

[Smi83] Richard L. Smith. Limit theorems and approximations for the reliability of load-sharing systems. Advances in Applied Probability, 15(2):304–330, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426438.

Smith:1988:CCE

[Smi88] Richard L. Smith. A counterexample concerning the extremal index. Advances in Applied Probability, 20(3):681–683, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427042.

Smythe:1976:MSP

[Smy76] R. T. Smythe. Multiparameter subadditive processes. Advances in Applied Probability, 8(2):252–254, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425902.

Samaniego:2016:CCS

[SN16] Francisco J. Samaniego and Jorge Navarro. On comparing coherent systems with heterogeneous components. Advances in Applied Probability, 48(1):88–111, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1457466157.

Seaman:1985:GVB

[SO85] John Seaman and Pat Odell. On Goldstein’s variance bound. Advances in Applied Probability, 17(3):679–681, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427126.

Sezer:2011:ABM

[SÖ11] Ali Devin Sezer and Ferruh Özbudak. Approximation of bounds on mixed-level orthogonal arrays. Advances in Applied Probability, 43(2):399–421, June 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/23024512.

Sobczyk:1987:SMF

[Sob87] K. Sobczyk. Stochastic models for fatigue damage of materials. Advances in Applied Probability, 19(3):652–673, September 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427411.

Sonderman:1979:CMSa

[Son79a] David Sonderman. Comparing multi-server queues with finite waiting rooms, i: Same number of servers. Advances in Applied Probability, 11(2):439–447, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426848.

Sonderman:1979:CMSb

[Son79b] David Sonderman. Comparing multi-server queues with finite waiting rooms, II: Different numbers of servers. Advances in Applied Probability, 11(2):448–455, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426849.

Soo:2010:TEM

[Soo10] Terry Soo. Translation-equivariant matchings of coin flips on Z^{d}. Advances in Applied Probability, 42(1):69–82, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/25683807.

Senturia:1973:SMS

[SP73] Jerome Senturia and Prem S. Puri. A semi-Markov storage model. Advances in Applied Probability, 5(2):362–378, August 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426041.

Srinivasan:1980:CFG

[SP80] S. K. Srinivasan and C. Chudalaimuthu Pillai. Correlation functions in the G/M/1 system. Advances in Applied Probability, 12 (2):530–540, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426611.

Spouge:1984:BPS

[Spo84] John L. Spouge. A branching-process solution of the polydisperse coagulation equation. Advances in Applied Probability, 16 (1):56–69, March 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427224.

Spodarev:2001:RIS

[Spo01] Eugene Spodarev. On the Rose of intersections of stationary flat processes. Advances in Applied Probability, 33(3):584–599, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428476.

Sing:1976:CCS

[SR76] C. F. Sing and E. D. Rothman. A consideration of the chi-square test of Hardy–Weinberg equilibrium in a non-multinomial situation. Advances in Applied Probability, 8(1):22–23, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426018.

Srinivasan:1982:PIB

[SR82] S. K. Srinivasan and C. R. Ranganathan. On the parity of individuals in birth and death processes. Advances in Applied Probability, 14(3):484–501, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1426670.

Sigman:2000:CTM

[SR00] Karl Sigman and Reade Ryan. Continuous-time monotone stochastic recursions and duality. Advances in Applied Probability, 32(2):426–445, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428197.

Srinivasan:1985:EGW

[Sri85] C. Srinivasan. Estimation in a Galton–Watson process. Advances in Applied Probability, 17(2):256, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427139.

Simonian:1997:AEB

[SRTM97] A. Simonian, J. W. Roberts, F. Théberge, and R. Mazumdar. Asymptotic estimates for blocking probabilities in a large multi-rate loss network. Advances in Applied Probability, 29(3): 806–829, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428086.

Singer:1977:TIM

[SS77] Burton Singer and Seymour Spilerman. Trace inequalities for mixtures of Markov chains. Advances in Applied Probability, 9 (4):747–764, December 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426699.

Seno:1984:DMT

[SS84a] Shoichiro Seno and Tokuzo Shiga. Diffusion models of temporally varying selection in population genetics. Advances in Applied Probability, 16(2):260–280, June 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427069.

Shanthikumar:1984:DPS

[SS84b] J. G. Shanthikumar and Ushio Sumita. Distribution properties of the system failure time in a general shock model. Advances in Applied Probability, 16(2):363–377, June 1984. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427074.

Sumita:1985:CCC

[SS85] Ushio Sumita and J. George Shanthikumar. A class of correlated cumulative shock models. Advances in Applied Probability, 17 (2):347–366, June 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427145.

Shaked:1987:MHR

[SS87] Moshe Shaked and J. George Shanthikumar. Multivariate hazard rates and stochastic ordering. Advances in Applied Probability, 19(1):123–137, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427376.

Shaked:1988:SCA

[SS88a] Moshe Shaked and J. George Shanthikumar. Stochastic convexity and its applications. Advances in Applied Probability, 20 (2):427–446, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427398.

Sumita:1988:ADC

[SS88b] Ushio Sumita and J. George Shanthikumar. An age-dependent counting process generated from a renewal process. Advances in Applied Probability, 20(4):739–755, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427358.

Shaked:1990:CSS

[SS90] Moshe Shaked and J. George Shanthikumar. Convexity of a set of stochastically ordered random variables. Advances in Applied Probability, 22(1):160–177, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427603.

Shaked:1992:OAR

[SS92] Moshe Shaked and J. George Shanthikumar. Optimal allocation of resources to nodes of parallel and series systems. Advances in Applied Probability, 24(4):894–914, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427718.

Sadowsky:1995:PLQa

[SS95a] John S. Sadowsky and Wojciech Szpankowski. The probability of large queue lengths and waiting times in a heterogeneous multiserver queue i: Tight limits. Advances in Applied Probability, 27(2):532–566, June 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427838.

Shaked:1995:CRP

[SS95b] Moshe Shaked and Ryszard Szekli. Comparison of replacement policies via point processes. Advances in Applied Probability, 27(4):1079–1103, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427935.

Sagitov:2009:MBG

[SS09] Serik Sagitov and Maria Conceição Serra. Multitype Bienaymé–Galton–Watson processes escaping extinction. Advances in Applied Probability, 41(1):225–246, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/27793871.

Stucki:2014:BPG

[SS14] Kaspar Stucki and Dominic Schuhmacher. Bounds for the probability generating functional of a Gibbs point process. Advances in Applied Probability, 46(1):21–34, March 2014. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1396360101.

Sagitov:2015:SNC

[SS15] Serik Sagitov and Maria Conceição Serra. Skeletons of nearcritical Bienaymé–Galton–Watson branching processes. Advances in Applied Probability, 47(2):530–544, June 2015. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1435236986.

Shneer:2020:SMB

[SS20] Seva Shneer and Alexander Stolyar. Stability and moment bounds under utility-maximising service allocations: Finite and infinite networks. Advances in Applied Probability, 52(2):463–490, June 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/stability-and-moment-bounds-under-
utilitymaximising-service-allocations-finite-and-infinite-
networks/A1CB4975F8AFCA80207795E1451B5416.

Sedivy:2013:SIR

[ŠSKB13] Ondřej Šedivý, Jakub Stanek, Blažena Kratochvílová, and Viktor Beneš. Sliced inverse regression and independence in random marked sets with covariates. Advances in Applied Probability, 45(3):626–644, September 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/projecteuclid.org/euclid.aap/1377868532.

Shakkottai:2004:POE

[SSS04] Sanjay Shakkottai, R. Srikant, and Alexander L. Stolyar. Pathwise optimality of the exponential scheduling rule for wireless channels. Advances in Applied Probability, 36(4):1021–1045, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
4140388.

Shepp:1991:PAR

[SSY91] Lawrence A. Shepp, Gordon Simons, and Yi-Ching Yao. On a problem of ammunition rationing. Advances in Applied Probability, 23(3):624–641, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427626.

Scalia-Tomba:1985:AFS

[ST85] Gianpaolo Scalia-Tomba. Asymptotic final-size distribution for some chain-binomial processes. Advances in Applied Probability, 17(3):477–495, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427116.

Stamoulis:1990:STC

[ST90] George D. Stamoulis and John N. Tsitsiklis. On the settling time of the congested GI/G/1 queue. Advances in Applied Probability, 22(4):929–956, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427569.

Szczotka:1994:CLT

[ST94] Władysław Szczotka and Krzysztof Topolski. Conditioned limit theorems for the difference of waiting time and queue length. Advances in Applied Probability, 26(1):242–257, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427589.

Sacerdote:1996:EAA

[ST96] L. Sacerdote and F. Tomassetti. On evaluations and asymptotic approximations of first-passage-time probabilities. Advances in Applied Probability, 28(1):270–284, March 1996. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427921.

Silvestrov:1998:LTE

[ST98] Dmitrii S. Silvestrov and Jozef L. Teugels. Limit theorems for extremes with random sample size. Advances in Applied Probability, 30(3):777–806, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1428151.

Stoev:2004:SPL

[ST04] Stilian Stoev and Murad S. Taqqu. Stochastic properties of the linear multifractional stable motion. Advances in Applied Probability, 36(4):1085–1115, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/4140390.

Schreiber:2010:SOP

[ST10] Tomasz Schreiber and Christoph Thäle. Second-order properties and central limit theory for the vertex process of iteration infinitely divisible and iteration stable random tessellations in the plane. Advances in Applied Probability, 42(4):913–935, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25799077.

Stam:1973:RVT

[Sta73] A. J. Stam. Regular variation of the tail of a subordinated probability distribution. Advances in Applied Probability, 5(2): 308–327, August 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426038.

Stam:1978:HTL

[Sta78a] A. J. Stam. Heavy traffic limit theorems in fluctuation theory. Advances in Applied Probability, 10(4):852–866, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426663.

Stam:1978:LTI

[Sta78b] A. J. Stam. On limit theorems with infinite limiting measure. Advances in Applied Probability, 10(2):274–277, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426881.

Stam:1980:DHF

[Sta80] P. Stam. The distribution of heterogeneity in finite random mating populations. Advances in Applied Probability, 12(1):9– 10, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426488.

Stadje:1990:CPG

[Sta90a] Wolfgang Stadje. The Collector’s problem with group drawings. Advances in Applied Probability, 22(4):866–882, December 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427566.

Stanford:1990:QI

[Sta90b] Robert E. Stanford. On queues with impatience. Advances in Applied Probability, 22(3):768–769, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427473.

Stadje:1992:EVO

[Sta92] Wolfgang Stadje. The expected value of the one-sided Cusum stopping time. Advances in Applied Probability, 24(4):1008– 1009, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427723.

Stadje:1998:APS

[Sta98] Wolfgang Stadje. Asymptotic probabilities in a sequential urn scheme related to the matchbox problem. Advances in Applied Probability, 30(3):831–849, September 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428153.

Sparaggis:1994:SPC

[STC94] Panayotis D. Sparaggis, Don Towsley, and Christos G. Cassandras. Sample path criteria for weak majorization. Advances in Applied Probability, 26(1):155–171, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427584.

Steutel:1973:SRR

[Ste73] F. W. Steutel. Some recent results in infinite divisibility. Advances in Applied Probability, 5(1):25–27, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425957.

Stene:1975:SP

[Ste75a] Jon Stene. Sampling of pedigrees. Advances in Applied Probability, 7(1):18–23, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1425848.

Steutel:1975:STF

[Ste75b] F. W. Steutel. Simple tools in finite and infinite divisibility. Advances in Applied Probability, 7(2):257–259, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426072.

Stene:1976:APB

[Ste76a] J. Stene. Analysis of panmixia based on human family data. Advances in Applied Probability, 8(1):24–27, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426019.

Stene:1976:MTI

[Ste76b] Jon Stene. A model for translocation inheritance. Advances in Applied Probability, 8(4):643–645, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425925.

Stene:1979:AFD

[Ste79] Jon Stene. Ascertainment of family data. Advances in Applied Probability, 11(1):9–10, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426763.

Stene:1980:DRF

[Ste80a] Jon Stene. Detection of regulation of family size after births of affected children. Advances in Applied Probability, 12(1):

10, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426489.

Steutel:1980:SLD

[Ste80b] F. W. Steutel. Stable lattice distributions and continuoustime branching processes. Advances in Applied Probability, 12 (2):286–288, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426559.

Stene:1981:CAD

[Ste81] Jon Stene. Comparisons of assumptions for different ascertainment models. Advances in Applied Probability, 13(1):7–8, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426460.

Stein:1984:SPG

[Ste84] Michael Stein. System parameters governed by jump processes: A model for removal of air pollutants. Advances in Applied Probability, 16(3):603–617, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1427289.

Stegeman:2002:EBH

[Ste02] Alwin Stegeman. Extremal behavior of heavy-tailed ON-Periods in a superposition of ON/OFF processes. Advances in Applied Probability, 34(1):179–204, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428378.

Steinerberger:2015:NBT

[Ste15] Stefan Steinerberger. New bounds for the traveling salesman constant. Advances in Applied Probability, 47(1):27–36, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1427814579.

Stenlund:2018:MUM

[Ste18] David Stenlund. On the Mabinogion urn model. Advances in Applied Probability, 50(2):327–346, June 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/on-the-mabinogion-urn-model/
2F090E9E4BF7001FE252EDEBD813479C.

Stidham:1972:RPT

[Sti72] Shaler Stidham, Jr. Regenerative processes in the theory of queues, with applications to the alternating-priority queue. Advances in Applied Probability, 4(3):542–577, December 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425993.

Stidham:1978:OCS

[Sti78] Shaler Stidham, Jr. Optimal control of stochastic service systems. Advances in Applied Probability, 10(2):277–278, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426882.

Stirzaker:1980:I

[Sti80] David Stirzaker. Inheritance. Advances in Applied Probability, 12(3):574–590, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426421.

Stirzaker:2006:THG

[Sti06] David Stirzaker. Three-handed Gambler’s ruin. Advances in Applied Probability, 38(1):284–286, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443439.

Stoyan:1978:QNI

[Sto78] Dietrich Stoyan. Queueing networks: Insensitivity and a heuristic approximation. Advances in Applied Probability, 10(2): 318, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426918.

Stone:1980:OSM

[Sto80] Lawrence D. Stone. Optimal search for moving targets. Advances in Applied Probability, 12(2):279–280, June 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426550.

Stoka:1998:RVR

[Sto98] Marius Stoka. On a random variable related to a system of convex bodies in the Euclidean space E_{n}. Advances in Applied Probability, 30(2):335–341, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427971.

Stolyar:2015:DST

[Sto15] A. L. Stolyar. Diffusion-scale tightness of invariant distributions of a large-scale flexible service system. Advances in Applied Probability, 47(1):251–269, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1427814590.

Stolyar:2017:LSH

[Sto17] A. L. Stolyar. Large-scale heterogeneous service systems with general packing constraints. Advances in Applied Probability, 49(1):61–83, March 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/largescale-heterogeneous-service-
systems-with-general-packing-constraints/0C4992C0D68C99CB795BAB2E072A9BB0.

Stolyar:2023:LSB

[Sto23] Alexander Stolyar. Large-scale behavior of a particle system with mean-field interaction: Traveling wave solutions. Advances in Applied Probability, 55(1):245–274, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/largescale-behavior-of-a-
particle-system-with-meanfield-interaction-traveling-
wave-solutions/08543CEE0C9F0FA987D8B017F6EDC1F6.

Strobeck:1974:OH

[Str74] Curtis Strobeck. Outcrossing and heterozygosity. Advances in Applied Probability, 6(1):18–20, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426204.

Strobeck:1975:TLM

[Str75] Curtis Strobeck. The two locus model with different recombination values in the two sexes. Advances in Applied Probability, 7(1):23–26, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425849.

Strobeck:1976:AR

[Str76] Curtis Strobeck. The algebra of recombination. Advances in Applied Probability, 8(1):27–29, March 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426020.

Streit:1978:OPS

[Str78] F. Streit. Optimality properties of statistical tests for Poisson point processes. Advances in Applied Probability, 10(2):330– 331, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426929.

Stuart:1978:LMF

[Stu78] Elizabeth Stuart. Linkage and mapping functions in human genetics. Advances in Applied Probability, 10(1):15–17, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426714.

See correction [Ano78e].

Sturgeon:1988:PPC

[Stu88] M. L. Sturgeon. Probabilistic profiles of continuously replenished spherically symmetric clouds. Advances in Applied Probability, 20(2):245–260, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427388.

Saunders:1984:GNS

[STW84] Ian W. Saunders, Simon Tavaré, and G. A. Watterson. On the genealogy of nested subsamples from a haploid population. Advances in Applied Probability, 16(3):471–491, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427285.

Sacerdote:2014:JDF

[STZ14] Laura Sacerdote, Ottavia Telve, and Cristina Zucca. Joint densities of first hitting times of a diffusion process through two time-dependent boundaries. Advances in Applied Probability, 46 (1):186–202, March 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1396360109.

Sazonov:1979:SCC

[SU79] V. V. Sazonov and V. V. Ulyanov. On the speed of convergence in the central limit theorem. Advances in Applied Probability, 11(2):269–270, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426799.

Sudbury:1993:SVI

[Sud93] Aidan Sudbury. The survival of various interacting particle systems. Advances in Applied Probability, 25(4):1010–1012, December 1993. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427804.

Sudbury:1999:HSJ

[Sud99] Aidan Sudbury. Hunting submartingales in the jumping voter model and the biased annihilating branching process. Advances in Applied Probability, 31(3):839–854, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428399.

Surgailis:2008:QAM

[Sur08] Donatas Surgailis. A quadratic ARCH(∞) model with long memory and Lévy stable behavior of squares. Advances in Applied Probability, 40(4):1198–1222, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443630.

Sutton:1979:SMA

[Sut79] G. R. Sutton. A stochastic model of acoustic wave propagation through the ocean. Advances in Applied Probability, 11(2):306– 307, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426838.

Smith:1975:AM

[SV75] Cedric A. B. Smith and Atam Vetta. Assortative mating. Advances in Applied Probability, 7(1):16–18, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425847.

Steutel:1976:IBC

[SV76] F. W. Steutel and K. Van Harn. Interpolating between the compound-geometric and compound-Poisson distributions. Advances in Applied Probability, 8(2):227–228, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425875.

Svane:2014:LDE

[Sva14] Anne Marie Svane. Local digital estimators of intrinsic volumes for Boolean models and in the design-based setting. Advances in Applied Probability, 46(1):35–58, March 2014. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1396360102.

Svensson:1993:DEC

[Sve93] Åke Svensson. Dynamics of an epidemic in a closed population. Advances in Applied Probability, 25(2):303–313, June 1993. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427654.

Steutel:1983:IVB

[SVW83] F. W. Steutel, W. Vervaat, and S. J. Wolfe. Integer-valued branching processes with immigration. Advances in Applied Probability, 15(4):713–725, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427320.

Smythe:1977:FPP

[SW77] R. T. Smythe and John C. Wierman. First-passage percolation on the square lattice. I. Advances in Applied Probability, 9(1): 38–54, March 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425815.

Smythe:1978:FPP

[SW78] R. T. Smythe and John C. Wierman. First-passage percolation on the square lattice. III. Advances in Applied Probability, 10 (1):155–171, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426723.

Stidham:1980:FRA

[SW80] Shaler Stidham, Jr. and Jacob Wijngaard. Foward recursion for average reward Markovian decision processes which are skip-free to the right. Advances in Applied Probability, 12(2):314, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426593.

Sakasegawa:1990:EVD

[SW90] Hirotaka Sakasegawa and Ronald W. Wolff. The equality of the virtual delay and attained waiting time distributions. Advances in Applied Probability, 22(1):257–259, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427611.

Shanthikumar:1991:SGG

[SW91] J. George Shanthikumar and Couchen Wu. On the starshapeness of G/G/c queueing systems. Advances in Applied Probability, 23(2):431–435, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427758.

Sudderth:1992:BBS

[SW92] W. D. Sudderth and A. P. N. Weerasinghe. A Bang–Bang strategy for a finite fuel stochastic control problem. Advances in Applied Probability, 24(3):589–603, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427481.

Shwartz:1993:IRE

[SW93] Adam Shwartz and Alan Weiss. Induced rare events: Analysis via large deviations and time reversal. Advances in Applied Probability, 25(3):667–689, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427529.

Sun:1997:WGA

[SW97] Fengzhu Sun and Michael S. Waterman. Whole genome amplification and branching processes. Advances in Applied Probability, 29(3):629–668, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428080.

Singpurwalla:1998:FMI

[SW98] Nozer D. Singpurwalla and Simon P. Wilson. Failure models indexed by two scales. Advances in Applied Probability, 30(4): 1058–1072, December 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428029.

Setayeshgar:2011:LDF

[SW11] Leila Setayeshgar and Hui Wang. Large deviations for a feedforward network. Advances in Applied Probability, 43(2):545– 571, June 2011. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
23024519.

Shao:2013:EEP

[SW13] Jinghai Shao and Xiuping Wang. Estimates of the exit probability for two correlated Brownian motions. Advances in Applied Probability, 45(1):37–50, March 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/41805999.

Swick:1975:ADD

[Swi75] David A. Swick. Asymptotic distribution of a discrete transform of an arbitrarily sampled homogeneous random field. Advances in Applied Probability, 7(1):140–153, March 1975. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425857.

Song:2014:WCR

[SWL14] Qifan Song, Mingqi Wu, and Faming Liang. Weak convergence rates of population versus single-chain stochastic approximation MCMC algorithms. Advances in Applied Probability, 46(4):1059–1083, December 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1418396243.

Schuhmacher:2008:NMB

[SX08] Dominic Schuhmacher and Aihua Xia. A new metric between distributions of point processes. Advances in Applied Probability, 40(3):651–672, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/20443602.

Smith:1981:GSP

[SY81] N. M. H. Smith and G. F. Yeo. On a general storage problem and its approximating solution. Advances in Applied Probability, 13 (3):567–602, September 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426786.

Simons:1990:SRB

[SY90] Gordon Simons and Yi-Ching Yao. Some results on the bomber problem. Advances in Applied Probability, 22(2):412–432, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427543.

Shanthikumar:1991:BCS

[SY91] J. George Shanthikumar and David D. Yao. Bivariate characterization of some stochastic order relations. Advances in Applied Probability, 23(3):642–659, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427627.

Stefanov:1997:EAR

[SY97] Valeri T. Stefanov and Geoffrey F. Yeo. Explicit asymptotic results for open times in ion channel models. Advances in Applied Probability, 29(4):947–964, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427848.

Syski:1973:EMB

[Sys73] Ryszard Syski. Examples of Martin boundary. Advances in Applied Probability, 5(1):27–29, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1425958.

Syski:1974:KCM

[Sys74] R. Syski. On Keilson’s compensation method. Advances in Applied Probability, 6(2):251–252, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426286.

Syski:1975:EMC

[Sys75] R. Syski. Energy of Markov chains. Advances in Applied Probability, 7(2):254–255, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426069.

Syski:1976:LET

[Sys76] R. Syski. On last-exit time distribution. Advances in Applied Probability, 8(2):246–247, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425896.

Syski:1977:EPM

[Sys77] R. Syski. Ergodic potential of a Markov chain. Advances in Applied Probability, 9(2):207–208, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426352.

Syski:1979:EMC

[Sys79] R. Syski. Energy of Markov chains. Advances in Applied Probability, 11(3):542–575, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1426954.

Syski:1980:EML

[Sys80] R. Syski. Energy of a martingale in the limit. Advances in Applied Probability, 12(2):283–284, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426554.

Syski:1988:FCS

[Sys88] R. Syski. Further comments on the solution of the M/M/1 queue. Advances in Applied Probability, 20(3):693, September 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427046.

Stadje:1990:OSS

[SZ90] Wolfgang Stadje and Dror Zuckerman. Optimal strategies for some repair replacement models. Advances in Applied Probability, 22(3):641–656, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427462.

Song:2011:AIF

[SZ11] Jing-Sheng Song and Paul Zipkin. An approximation for the inverse first passage time problem. Advances in Applied Probability, 43(1):264–275, March 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/23024535.

Szasz:1976:RTM

[Szá76] Domokos Szász. Renewal theory and multicomponent reliability systems. Advances in Applied Probability, 8(2):239–240, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425886.

Szajowski:1993:DST

[Sza93] K. Szajowski. Double stopping by two decision-makers. Advances in Applied Probability, 25(2):438–452, June 1993. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427661.

Szczotka:1986:SRQa

[Szc86a] Władysław Szczotka. Stationary representation of queues. I. Advances in Applied Probability, 18(3):815–848, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427189.

Szczotka:1986:SRQb

[Szc86b] Władysław Szczotka. Stationary representation of queues. II. Advances in Applied Probability, 18(3):849–859, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427190.

Szczotka:1990:EAW

[Szc90] Władysław Szczotka. Exponential approximation of waiting time and queue size for queues in heavy traffic. Advances in Applied Probability, 22(1):230–240, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427606.

Szczotka:1993:ASM

[Szc93] Władysław Szczotka. Asymptotic stationarity of multichannel queues. Advances in Applied Probability, 25(1):203–220, March 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427502.

Szczotka:1999:TSW

[Szc99] Władysław Szczotka. Tightness of the stationary waiting time in heavy traffic. Advances in Applied Probability, 31(3):788–794, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428396.

Sze:1976:MPA

[Sze76] Michael Sze. Markov processes associated with critical Galton– Watson processes with application to extinction probabilities. Advances in Applied Probability, 8(2):278–295, June 1976. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425905.

Szpankowski:1994:SCS

[Szp94] Wojciech Szpankowski. Stability conditions for some distributed systems: Buffered random access systems. Advances in Applied Probability, 26(2):498–515, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427448.

Takacs:1970:DMS

[Tak70] Lajos Takács. On the distribution of the maximum of sums of mutually independent and identically distributed random variables. Advances in Applied Probability, 2(2):344–354, Autumn 1970. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426323.

Takacs:1975:CAMa

[Tak75a] Lajos Takács. Combinatorial and analytic methods in the theory of queues. Advances in Applied Probability, 7(2):241, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426055.

Takacs:1975:CAMb

[Tak75b] Lajos Takács. Combinatorial and analytic methods in the theory of queues. Advances in Applied Probability, 7(3):607–635, September 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426131.

Takacs:1975:SPS

[Tak75c] Lajos Takács. A storage process with semi-Markov input. Advances in Applied Probability, 7(4):830–844, December 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426402.

Takacs:1976:FPTa

[Tak76a] Lajos Takács. On fluctuation problems in the theory of queues. Advances in Applied Probability, 8(2):221–222, June 1976. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1425870.

Takacs:1976:FPTb

[Tak76b] Lajos Takács. On fluctuation problems in the theory of queues. Advances in Applied Probability, 8(3):548–583, September 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426144.

Takahashi:1981:AET

[Tak81] Yukio Takahashi. Asymptotic exponentiality of the tail of the waiting-time distribution in a PH/PH/c queue. Advances in Applied Probability, 13(3):619–630, September 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426788.

Takacs:1984:RFR

[Tak84] Lajos Takács. Random flights on regular graphs. Advances in Applied Probability, 16(3):618–637, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427290.

Takahashi:1988:CME

[Tak88] Rinya Takahashi. Characterizations of a multivariate extreme value distribution. Advances in Applied Probability, 20(1):235– 236, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427279.

Takacs:1991:BEV

[Tak91a] Lajos Takács. A Bernoulli excursion and its various applications. Advances in Applied Probability, 23(3):557–585, September 1991. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427622.

Takagi:1991:AFC

[Tak91b] Hideaki Takagi. Analysis of finite-capacity polling systems. Advances in Applied Probability, 23(2):373–387, June 1991. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427754.

Tambouratzis:1974:PWT

[Tam74] D. G. Tambouratzis. Properties of waiting times of queues under various queue disciplines. Advances in Applied Probability, 6 (2):257–258, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426291.

Tamura:2008:MLT

[Tam08] Takashi Tamura. Maximization of the long-term growth rate for a portfolio with fixed and proportional transaction costs. Advances in Applied Probability, 40(3):673–695, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/20443603.

Tamaki:2011:MPS

[Tam11] Mitsushi Tamaki. Maximizing the probability of stopping on any of the last m successes in independent Bernoulli trials with random horizona. Advances in Applied Probability, 43(3):760–781, September 2011. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
23024574.

Tamaki:2013:OSR

[Tam13] Mitsushi Tamaki. Optimal stopping rule for the no-information duration problem with random horizon. Advances in Applied Probability, 45(4):1028–1048, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://projecteuclid.org/euclid.aap/1386857856.

Tamaki:2016:OSR

[Tam16a] Mitsushi Tamaki. Optimal stopping rule for the full-information duration problem with random horizon. Advances in Applied Probability, 48(1):52–68, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1457466155.

Tamaki:2016:USD

[Tam16b] Mitsushi Tamaki. Urn sampling distributions giving alternate correspondences between two optimal stopping problems. Advances in Applied Probability, 48(3):726–743, September 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1474296312.

Tanny:1981:MBP

[Tan81] David Tanny. On multitype branching processes in a random environment. Advances in Applied Probability, 13(3):464–497, September 1981. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426781.

Tanemura:1996:MSR

[Tan96] Masaharu Tanemura. Methods of simulating random patterns of non-spherical objects and their application. Advances in Applied Probability, 28(2):342–343, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428057.

Taqqu:1975:WCF

[Taq75] Murad Taqqu. Weak convergence to fractional Brownian motion and to the Rosenblatt process. Advances in Applied Probability, 7(2):249, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426060.

Tautu:1971:SME

[Tau71] Petre Tautu. Structural models in epidemiology: An introductory investigation. Advances in Applied Probability, 3(2):196– 198, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426153.

Tautu:1974:RSL

[Tau74] P. Tautu. Random systems of locally interacting cells. Advances in Applied Probability, 6(2):237, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426274.

Tautu:1975:SAF

[Tau75] P. Tautu. A stochastic approach to the French flag problem. Advances in Applied Probability, 7(2):262–263, June 1975. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426076.

Tautu:1980:I

[Tau80] P. Tautu. Introduction. Advances in Applied Probability, 12(3): 545–546, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426414.

Tavare:1976:STC

[Tav76] S. Tavaré. Sojourn times for conditioned Markov chains in genetics. Advances in Applied Probability, 8(4):645–648, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425926.

Tavare:1978:ADM

[Tav78] S. Tavaré. Age distributions for Markov processes in genetics. Advances in Applied Probability, 10(1):17–19, March 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426715.

Tavare:1979:TRA

[Tav79] S. Tavaré. Time reversion and age distributions. Advances in Applied Probability, 11(1):10–12, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426764.

Tavare:1992:I

[Tav92] S. Tavaré. Introduction. Advances in Applied Probability, 24(4): 762–763, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427709.

Taylor:1977:ACP

[Tay77] S. J. Taylor. The asymmetric Cauchy process. Advances in Applied Probability, 9(2):208, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426353.

Taylor:1979:TFF

[Tay79] Howard M. Taylor. The time to failure of fiber bundles subjected to random loads. Advances in Applied Probability, 11(3): 527–541, September 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426953.

Taylor:1989:IPZ

[Tay89] P. Taylor. Insensitivity in processes with zero speeds. Advances in Applied Probability, 21(3):612–628, September 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427639.

Tong:1992:NOD

[TC92] H. Tong and B. Cheng. A note on one-dimensional chaotic maps under time reversal. Advances in Applied Probability, 24(1):219– 220, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427737.

Tykesson:2013:AVH

[TC13] Johan Tykesson and Pierre Calka. Asymptotics of visibility in the hyperbolic plane. Advances in Applied Probability, 45 (2):332–350, June 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1370870121.

Tassiulas:1996:TPQ

[TE96] Leandros Tassiulas and Anthony Ephremides. Throughput properties of a queueing network with distributed dynamic routing and flow control. Advances in Applied Probability, 28(1):285– 307, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427922.

Teodorescu:1982:CSP

[Teo82] N. Teodorescu. Connection, state and phenomena in general systems. Advances in Applied Probability, 14(2):202–203, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426514.

Teugels:1973:RVR

[Teu73] Jozef L. Teugels. Regular variation in renewal and Markov renewal theory. Advances in Applied Probability, 5(1):30–31, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425959.

Teugels:1979:RCM

[Teu79] J. L. Teugels. On the rate of convergence of the maximum of a compound Poisson process. Advances in Applied Probability, 11(2):274–275, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426804.

Touboul:2008:CFH

[TF08] Jonathan Touboul and Olivier Faugeras. A characterization of the first hitting time of double integral processes to curved boundaries. Advances in Applied Probability, 40(2):501–528, June 2008. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

20443593. See correction [Ano09e].

Takine:1990:NMG

[TH90] T. Takine and T. Hasegawa. A note on M/G/1 vacation systems with waiting time limits. Advances in Applied Probability, 22 (2):513–518, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427557.

Thevenin:2022:RST

[Thé22] Paul Thévenin. Random stable-type minimal factorizations of the n -cycle. Advances in Applied Probability, 54(1):1–63, March 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/random-stabletype-minimal-factorizations-
of-the-ncycle/B683BBADF1B57A2346B302E8767E4575.

Thompson:1975:MGI

[Tho75] E. A. Thompson. Multiple gene identities. Advances in Applied Probability, 7(1):26–27, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425850.

Thompson:1976:PGI

[Tho76] E. A. Thompson. A paradox of genealogical inference. Advances in Applied Probability, 8(4):648–650, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1425927.

Thompson:1978:IGI

[Tho78] E. A. Thompson. Impossible gene identity states. Advances in Applied Probability, 10(1):19–22, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426716.

Thompson:1979:EPP

[Tho79] E. A. Thompson. Extinction probabilities and pedigree structure. Advances in Applied Probability, 11(1):12–13, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426765.

Thompson:1980:SSS

[Tho80] E. A. Thompson. Sequential sampling schemes for sibships. Advances in Applied Probability, 12(1):10–11, March 1980. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426490.

Thompson:1981:ICG

[Tho81a] E. A. Thompson. Impossible combinations of gene correlations. Advances in Applied Probability, 13(1):8–9, March 1981. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426461.

Thompson:1981:EPS

[Tho81b] M. E. Thompson. Estimation of the parameters of a semi-Markov process from censored records. Advances in Applied Probability, 13(4):804–825, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426974.

Thompson:1982:OSP

[Tho82] E. A. Thompson. Optimal sampling for pedigree analysis: Tracing a single gene. Advances in Applied Probability, 14(4):752– 762, December 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427022.

Thorisson:1983:CRP

[Tho83] Hermann Thorisson. The coupling of regenerative processes. Advances in Applied Probability, 15(3):531–561, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426618.

Thonnes:1999:PSS

[Thö99] Elke Thönnes. Perfect simulation of some point processes for the impatient user. Advances in Applied Probability, 31(1):69– 87, March 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427992.

Thorisson:2007:PDR

[Tho07] Hermann Thorisson. The Palm-duality for random subsets of d-dimensional grids. Advances in Applied Probability, 39(2):318– 325, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443515.

Tierney:1981:ATF

[Tie81] L. Tierney. The asymptotic time to failure of a bundle of fibers with random slacks. Advances in Applied Probability, 13(3): 548–566, September 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426785.

Tierney:1982:ABT

[Tie82] Luke Tierney. Asymptotic bounds on the time to fatigue failure of bundles of fibers under local load sharing. Advances in Applied Probability, 14(1):95–121, March 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426735.

Tierney:1983:HOP

[Tie83] Luke Tierney. The hazards of optimal proofreading. Advances in Applied Probability, 15(4):892–893, December 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427333.

Tanemura:1998:RTN

[TJM98] M. Tanemura, E. B. Vedel Jensen, and J. Møller. Random tessellation network and some inference problems. Advances in Applied Probability, 30(2):291–292, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427966.

Tjostheim:1976:SGO

[Tjø76a] D. Tjøstheim. Spectral generating operators for non-stationary processes. Advances in Applied Probability, 8(4):831–846, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425937.

Tjostheim:1976:RPA

[Tjø76b] Dag Tjøstheim. On random processes that are almost strict sense stationary. Advances in Applied Probability, 8(4):820–830, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425936.

Tjostheim:1978:SSS

[Tjø78] Dag Tjøstheim. Statistical spatial series modelling. Advances in Applied Probability, 10(1):130–154, March 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426722. See [Ano84e].

Tjostheim:1983:SSS

[Tjø83] Dag Tjøstheim. Statistical spatial series modelling II: Some further results on unilateral lattice processes. Advances in Applied Probability, 15(3):562–584, September 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426619.

Tjostheim:1990:NLT

[Tjø90] Dag Tjøstheim. Non-linear time series and Markov chains. Advances in Applied Probability, 22(3):587–611, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427459.

Thorisdottir:2013:WPL

[TK13] Ó. Thórisdóttir and M. Kiderlen. Wicksell’s problem in local stereology. Advances in Applied Probability, 45(4):925–944, December 2013. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1386857851.

Teng:2015:SMC

[TKF15] Huei-Wen Teng, Ming-Hsuan Kang, and Cheng-Der Fuh. On spherical Monte Carlo simulations for multivariate normal probabilities. Advances in Applied Probability, 47(3):817–836, September 2015. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://projecteuclid.org/
euclid.aap/1444308883.

Torrisi:2020:ASC

[TL20] Giovanni Luca Torrisi and Emilio Leonardi. Almost sure central limit theorems in stochastic geometry. Advances in Applied Probability, 52(3):705–734, September 2020. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/almost-sure-
central-limit-theorems-in-stochastic-geometry/AE55808CE144C1626DBF030D1CB97420.

Terdik:1991:EBB

[TM91] György Terdik and Laurie Meaux. The exact bispectra for bilinear realizable processes with Hermite degree 2. Advances in Applied Probability, 23(4):798–808, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427676.

Thomas:2021:FLT

[TO21] Andrew M. Thomas and Takashi Owada. Functional limit theorems for the Euler characteristic process in the critical regime. Advances in Applied Probability, 53(1):57–80, March 2021. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/functional-limit-
theorems-for-the-euler-characteristic-process-in-the-
critical-regime/61119D2F1378BB75096B06245E69721D.

Todorovic:1976:SRN

[Tod76] P. Todorovic. On the structure of the Radon–Nikodym derivative for a Markov process. Advances in Applied Probability, 8 (2):247–248, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1425897.

Todorovic:1977:SRN

[Tod77] P. Todorovic. On the structure of the Radon–Nikodym derivate for a point process. Advances in Applied Probability, 9(2): 237, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426382.

Tollar:1988:LBM

[Tol88] Eric S. Tollar. On the limit behavior of a multicompartment storage model with an underlying Markov chain. Advances in Applied Probability, 20(1):208–227, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427276.

Tonnelat:1982:OEL

[Ton82] J. Tonnelat. The organization and evolution of living systems as normal consequences of the laws of thermodynamics. Advances in Applied Probability, 14(2):203–205, June 1982. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426515.

Torok:1978:BEQ

[Tör78] T. L. Török. Balance equations in queueing networks. Advances in Applied Probability, 10(2):320, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426920.

Tory:1992:UGF

[TP92] E. M. Tory and D. K. Pickard. Unilateral Gaussian fields. Advances in Applied Probability, 24(1):95–112, March 1992. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427731.

Tawn:2019:APT

[TR19] Nicholas G. Tawn and Gareth O. Roberts. Accelerating parallel tempering: Quantile tempering algorithm (QuanTA). Advances in Applied Probability, 51(3):802–834, September 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/accelerating-parallel-tempering-
quantile-tempering-algorithm-quanta/34F687AEBF9DA5BFD3A6EE43557D42BB.

Trajstman:1974:CGW

[Tra74] A. C. Trajstman. On a conjecture of G. A. Watterson. Advances in Applied Probability, 6(3):489–493, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426229.

Truslove:1975:BPG

[Tru75a] A. L. Truslove. The busy period of the E_{k} /G/1 queue with finite waiting room. Advances in Applied Probability, 7(2):416– 430, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426083.

Truslove:1975:QLG

[Tru75b] A. L. Truslove. Queue length for the E_{k} /G/1 queue with finite waiting room. Advances in Applied Probability, 7(1):215– 226, March 1975. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425861.

Truffet:1997:NCD

[Tru97] Laurent Truffet. Near complete decomposability: Bounding the error by a stochastic comparison method. Advances in Applied Probability, 29(3):830–855, September 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428087.

Thomson:1981:EFO

[TS81] Robert Thomson and Cedric A. B. Smith. Estimating the frequencies of origins of trisomics. Advances in Applied Probability, 13(1):9–11, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426462.

Taylor:1998:CEE

[TS98] Howard M. Taylor and Dennis E. Sweitzer. On the current enhancement at the edge of a crack in a lattice of resistors. Advances in Applied Probability, 30(2):342–364, June 1998. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427972.

Tu:1999:MFM

[TS99] I-Ping Tu and David Siegmund. The maximum of a function of a Markov chain and application to linkage analysis. Advances in Applied Probability, 31(2):510–531, June 1999. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428120.

Trillos:2017:EPU

[TSvB17] Nicolás García Trillos, Dejan Slepcev, and James von Brecht. Estimating perimeter using graph cuts. Advances in Applied Probability, 49(4):1067–1090, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/estimating-perimeter-using-
graph-cuts/D7B6CCD1335F7AB32A6F408EC890083A.

Tuominen:1979:EDE

[TT79a] Pekka Tuominen and Richard L. Tweedie. Exponential decay and ergodicity of general Markov processes and their discrete skeletons. Advances in Applied Probability, 11(4):784–803, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426859.

Tuominen:1979:EEG

[TT79b] Pekka Tuominen and Richard L. Tweedie. Exponential ergodicity of general Markov processes. Advances in Applied Probability, 11(2):279–280, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426810.

Tuominen:1979:RSG

[TT79c] Pekka Tuominen and Richard L. Tweedie. The recurrence structure of general Markov processes. Advances in Applied Probability, 11(2):279, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426809.

Tuominen:1994:SRC

[TT94] Pekka Tuominen and Richard L. Tweedie. Subgeometric rates of convergence of f-ergodic Markov chains. Advances in Applied Probability, 26(3):775–798, September 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427820.

Tang:2004:FIT

[TT04] Qihe Tang and Gurami Tsitsiashvili. Finite- and infinite-time ruin probabilities in the presence of stochastic returns on investments. Advances in Applied Probability, 36(4):1278–1299, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
4140398.

Tucker:1971:AIP

[Tuc71] Howard G. Tucker. On asymptotic independence of the partial sums of positive and negative parts of independent random variables. Advances in Applied Probability, 3(2):404–425, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426178.

Turnbull:1974:ITF

[Tur74] Bruce W. Turnbull. An inequality for the total flow in a dam with non-Markovian inputs. Advances in Applied Probability, 6 (2):218–219, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426256.

Teugels:1975:ALS

[TV75] J. L. Teugels and N. Veraverbeke. On the approach to the limit of successive maxima of partial sums. Advances in Applied Probability, 7(2):249–250, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426061.

Tijms:1978:ICT

[TV78] H. C. Tijms and F. A. Van Der Duyn Schouten. Inventory control with two switch-over levels for a class of finite capacity M/G/1 queueing systems with variable arrival and service rate. Advances in Applied Probability, 10(2):319, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426919.

Taylor:2010:DRB

[TV10] P. G. Taylor and B. Van Houdt. On the dual relationship between Markov chains of Gi/M/1 and M/G/1 type. Advances in Applied Probability, 42(1):210–225, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/25683813.

Tijms:1981:ASS

[TVF81] H. C. Tijms, M. H. Van Hoorn, and A. Federgruen. Approximations for the steady-state probabilities in the M/G/c queue. Advances in Applied Probability, 13(1):186–206, March 1981. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426474.

Tsoucas:1983:NPS

[TW83] Pantelis Tsoucas and Jean Walrand. A note on the processorsharing queue in a quasi-reversible network. Advances in Applied Probability, 15(2):468–469, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426448.

Tsoucas:1984:NSB

[TW84a] Pantelis Tsoucas and Jean Walrand. A note on stochastic bounds for queueing networks. Advances in Applied Probability, 16(4):926–928, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427349.

Turkman:1984:ADM

[TW84b] K. F. Turkman and A. M. Walker. On the asymptotic distributions of maxima of trigonometric polynomials with random coefficients. Advances in Applied Probability, 16(4):819–842, December 1984. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427342.

Taqqu:1987:AFS

[TW87a] Murad S. Taqqu and Walter Willinger. The analysis of finite security markets using martingales. Advances in Applied Probability, 19(1):1–25, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427371.

Tsoucas:1987:ISO

[TW87b] Pantelis Tsoucas and Jean Walrand. On the interchangeability and stochastic ordering of · /M/1 queues in tandem. Advances in Applied Probability, 19(2):515–520, June 1987. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427433.

Turchi:2018:LTR

[TW18] N. Turchi and F. Wespi. Limit theorems for random polytopes with vertices on convex surfaces. Advances in Applied Probability, 50(4):1227–1245, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/limit-theorems-for-random-polytopes-
with-vertices-on-convex-surfaces/87EC9C49CB7C891B5D6A14DDE18B22B0.

Tweedie:1976:CCG

[Twe76] R. L. Tweedie. Criteria for classifying general Markov chains. Advances in Applied Probability, 8(4):737–771, December 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425932.

Tweedie:1977:CCG

[Twe77] R. L. Tweedie. Criteria for classifying general Markov chains. Advances in Applied Probability, 9(2):208–210, June 1977. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426354.

Tweedie:1982:OGS

[Twe82] R. L. Tweedie. Operator-geometric stationary distributions for Markov chains, with application to queueing models. Advances in Applied Probability, 14(2):368–391, June 1982. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426527.

Thale:2012:SST

[TWN12] Christoph Thäle, Viola Weiss, and Werner Nagel. Spatial STIT tessellations: Distributional results for I-segments. Advances in Applied Probability, 44(3):635–654, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/41714069.

Tan:2010:CFJ

[TX10] Sen Tan and Aihua Xia. On customer flows in Jackson queueing networks. Advances in Applied Probability, 42(4):1094–1101, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25799085.

Thoppe:2016:ETD

[TYA16] Gugan C. Thoppe, D. Yogeshwaran, and Robert J. Adler. On the evolution of topology in dynamic clique complexes. Advances in Applied Probability, 48(4):989–1014, December 2016. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1482548426.

Urbanik:1977:GDP

[Urb77] K. Urbanik. Geometric decomposability properties of probability measures. Advances in Applied Probability, 9(3):437–439, September 1977. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426099.

Uquillas:2019:FDT

[US19] Adriana Uquillas and Adilson Simonis. First displacement time of a tagged particle in a stochastic cluster in a simple exclusion process with random slow bonds. Advances in Applied Probability, 51(3):717–744, September 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/first-displacement-time-
of-a-tagged-particle-in-a-stochastic-cluster-in-a-simple-
exclusion-process-with-random-slow-bonds/FD08DCAE45F50196900F6475B7D9FB27.

Vallois:1996:RSR

[Val96] P. Vallois. The range of a simple random walk on Z. Advances in Applied Probability, 28(4):1014–1033, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428162.

VanDerMeulen:1971:TTC

[Van71] Edward C. Van Der Meulen. Three-terminal communication channels. Advances in Applied Probability, 3(1):120–154, Spring 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426331.

VanDerMeulen:1973:MTC

[Van73a] Edward C. Van Der Meulen. Multi-terminal communication channels. Advances in Applied Probability, 5(1):32–33, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425960.

VanMoerbeke:1973:OSF

[Van73b] Pierre Van Moerbeke. On optimal stopping and free boundary problems. Advances in Applied Probability, 5(1):33–35, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425961.

Vanderperre:1974:NHH

[Van74] E. J. Vanderperre. A non-homogeneous Hilbert problem with applications to queueing theory. Advances in Applied Probability, 6(2):258–259, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426292.

VanBeek:1978:I

[Van78a] P. Van Beek. Introduction. Advances in Applied Probability, 10 (2):332–333, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426931.

VanDerDuynSchouten:1978:MGQ

[Van78b] F. A. Van Der Duyn Schouten. An M/G/1 queueing model with vacation times. Advances in Applied Probability, 10(2):316, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426916.

VanHarn:1978:DID

[Van78c] K. Van Harn. Degrees of infinite divisibility and p-functions. Advances in Applied Probability, 10(2):283–285, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426887.

VanNunen:1978:SAM

[Van78d] J. A. E. E. Van Nunen. Successive approximations for Markov decision processes. Advances in Applied Probability, 10(2):311– 312, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426912.

VanDoorn:1980:SMB

[Van80] Erik A. Van Doorn. Stochastic monotonicity of birth–death processes. Advances in Applied Probability, 12(1):59–80, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426494.

VanDenBerg:1983:CCJ

[Van83] J. Van Den Berg. A counterexample to a conjecture of J. M. Hammersley and D. J. A. Welsh concerning first-passage percolation. Advances in Applied Probability, 15(2):465–467, June 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426447.

VanDenBerg:1984:DCS

[Van84] J. Van Den Berg. Disproof of the conjectured subexponentiality of certain functions in percolation theory. Advances in Applied Probability, 16(3):690–691, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427295.

VanDoorn:1985:CEE

[Van85] Erik A. Van Doorn. Conditions for exponential ergodicity and bounds for the decay parameter of a birth–death process. Advances in Applied Probability, 17(3):514–530, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427118.

VanDijk:1988:PTU

[Van88a] Nico M. Van Dijk. Perturbation theory for unbounded Markov reward processes with applications to queueing. Advances in Applied Probability, 20(1):99–111, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427272.

VanOmmeren:1988:EET

[Van88b] J. C. W. Van Ommeren. Exponential expansion for the tail of the waiting-time probability in the single-server queue with batch arrivals. Advances in Applied Probability, 20(4):880–895, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427365.

VanDijk:1990:SPU

[Van90] Nico M. Van Dijk. On a simple proof of uniformization for continuous and discrete-state continuous-time Markov chains. Advances in Applied Probability, 22(3):749–750, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427467.

VanDoorn:1991:QSD

[Van91] Erik A. Van Doorn. Quasi-stationary distributions and convergence to quasi-stationarity of birth–death processes. Advances in Applied Probability, 23(4):683–700, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427670.

VanLieshout:1994:SAN

[Van94] M. N. M. Van Lieshout. Stochastic annealing for nearestneighbour point processes with application to object recognition. Advances in Applied Probability, 26(2):281–300, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427436.

VanLieshout:2006:MLE

[Van06] M. N. M. Van Lieshout. Maximum likelihood estimation for random sequential adsorption. Advances in Applied Probability, 38(4):889–898, December 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/20443480.

Vasata:2016:LRD

[Vaš16] Daniel Vašata. On long-range dependence of random measures. Advances in Applied Probability, 48(4):1235–1255, December 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://projecteuclid.org/euclid.aap/
1482548436.

Vis:2015:HTL

[VBvdM15] P. Vis, R. Bekker, and R. D. van der Mei. Heavy-traffic limits for polling models with exhaustive service and non-FCFS service order policies. Advances in Applied Probability, 47(4): 989–1014, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://projecteuclid.

org/euclid.aap/1449859797.

vanderHofstad:2004:IPT

[vdHK04] Remco van der Hofstad and Marten J. Klok. Improving the performance of third-generation wireless communication systems. Advances in Applied Probability, 36(4):1046–1084, December 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/4140389.

vanderHofstad:2021:RIG

[vdHKV21] Remco van der Hofstad, Júlia Komjáthy, and Viktória Vadon. Random intersection graphs with communities. Advances in Applied Probability, 53(4):1061–1089, December 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/random-intersection-graphs-
with-communities/B575661DEB07B848AEB75251344F4EAB.

vanderHofstad:2020:LTA

[vdHvdHLS20] Remco van der Hofstad, Pim van der Hoorn, Nelly Litvak, and Clara Stegehuis. Limit theorems for assortativity and clustering in null models for scale-free networks. Advances in Applied Probability, 52(4):1035–1084, December 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/limit-theorems-for-assortativity-
and-clustering-in-null-models-for-scalefree-networks/
3A23592361D45F1C837DAA597DDCBBB0.

vanderHofstad:2023:LLS

[vdHvdHM23] Remco van der Hofstad, Pim van der Hoorn, and Neeladri Maitra. Local limits of spatial inhomogeneous random graphs. Advances in Applied Probability, 55(3):793–840, September 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/local-
limits-of-spatial-inhomogeneous-random-graphs/0CF78F79EA76098277B1DA9A09E12975.

Venkataraman:1982:TSA

[Ven82] K. N. Venkataraman. A time series approach to the study of the simple subcritical Galton–Watson process with immigration. Advances in Applied Probability, 14(1):1–20, March 1982. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426730.

Vervaat:1973:SES

[Ver73] Wim Vervaat. Success epochs in a sequence of Bernoulli trials. Advances in Applied Probability, 5(1):35–36, April 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1425962.

Vervaat:1975:SIL

[Ver75] Wim Vervaat. Strassen’s iterated logarithm law for Hilbert space valued random variables. Advances in Applied Probability, 7(2):242–244, June 1975. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426056.

Vervaat:1977:SDE

[Ver77] Wim Vervaat. On a stochastic difference equation. Advances in Applied Probability, 9(2):213–215, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426357.

Vervaat:1978:RBB

[Ver78] W. Vervaat. A relation between the Brownian bridge and the Brownian excursion. Advances in Applied Probability, 10(2): 299, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426901.

Vervaat:1979:SDE

[Ver79] Wim Vervaat. On a stochastic difference equation and a representation of non-negative infinitely divisible random variables. Advances in Applied Probability, 11(4):750–783, December 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426858.

Vetta:1978:NAM

[Vet78] Atam Vetta. Negative assortative mating for a polygenic trait. Advances in Applied Probability, 10(1):22–25, March 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426717.

Vetta:1980:PFF

[Vet80] Atam Vetta. Proof of Fischer’s fundamental theorem of natural selection. Advances in Applied Probability, 12(1):11–13, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426491.

Vetta:1981:FTE

[Vet81] Atam Vetta. Fisher’s theory of evolution of dominance. Advances in Applied Probability, 13(1):11–12, March 1981. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426463.

Voss:2010:SLS

[VGS10] Florian Voss, Catherine Gloaguen, and Volker Schmidt. Scaling limits for shortest path lengths along the edges of stationary tessellations. Advances in Applied Probability, 42(4):936–952, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25799078.

Videla:2020:ERW

[Vid20] Leonardo A. Videla. Out-of-equilibrium random walks. Advances in Applied Probability, 52(3):772–797, September 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/outofequilibrium-
random-walks/84A264F4BF143870B1CCAE3E1CA845B6.

Vitale:1995:VPB

[Vit95] Richard A. Vitale. On the volume of parallel bodies: A probabilistic derivation of the Steiner formula. Advances in Applied Probability, 27(1):97–101, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428098.

Vitale:2001:IVG

[Vit01] Richard A. Vitale. Intrinsic volumes and Gaussian processes. Advances in Applied Probability, 33(2):354–364, June 2001. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428257.

Vere-Jones:1999:FDP

[VJ99] David Vere-Jones. On the fractal dimensions of point patterns. Advances in Applied Probability, 31(3):643–663, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428389.

Vere-Jones:2005:CSS

[VJ05] David Vere-Jones. A class of self-similar random measure. Advances in Applied Probability, 37(4):908–914, December 2005. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/30037366.

Vere-Jones:2008:LTA

[VJ08] David Vere-Jones. A limit theorem with application to Båth’s law in seismology. Advances in Applied Probability, 40(3):882– 896, September 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443612.

VanDerHofstad:2010:CER

[VJV10] Remco Van Der Hofstad, A. J. E. M. Janssen, and Johan S. H. Van Leeuwaarden. Critical epidemics, random graphs, and Brownian motion with a parabolic drift. Advances in Applied Probability, 42(4):1187–1206, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/25799090.

VanBeek:1978:ADA

[VK78] Paul Van Beek and Hubert Keuning. An application of the diffusion approximation technique to store-and-forward computer network models. Advances in Applied Probability, 10(2): 334, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426933.

VanDerMei:1998:EDA

[VL98] R. D. Van Der Mei and H. Levy. Expected delay analysis of polling systems in heavy traffic. Advances in Applied Probability, 30(2):586–602, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427984.

Vera:2007:GCS

[VL07] Francisco Vera and James Lynch. General convex stochastic orderings and related martingale-type structures. Advances in Applied Probability, 39(1):105–127, March 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443502.

Volkovich:2010:AAP

[VL10] Yana Volkovich and Nelly Litvak. Asymptotic analysis for personalized Web search. Advances in Applied Probability, 42(2): 577–604, June 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/25683834.

vanLieshout:2018:NNM

[vL18] M. N. M. van Lieshout. Nearest-neighbour Markov point processes on graphs with Euclidean edges. Advances in Applied Probability, 50(4):1275–1293, December 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/nearestneighbour-markov-
point-processes-on-graphs-with-euclidean-edges/FDE4FD15FE4D79934FC28F73DEBD1B2A.

VonBahr:1980:TLT

[VML80] Bengt Von Bahr and Anders Martin-Löf. Threshold limit theorems for some epidemic processes. Advances in Applied Probability, 12(2):319–349, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426600.

Vasantam:2019:IMF

[VMM19] Thirupathaiah Vasantam, Arpan Mukhopadhyay, and Ravi R. Mazumdar. Insensitivity of the mean field limit of loss systems under SQ(d) routeing. Advances in Applied Probability, 51(4):1027–1066, December 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/insensitivity-of-the-mean-field-
limit-of-loss-systems-under-sqd-routeing/39A940AF56C09E98F868371F350B9ADD.

Voit:1990:CLT

[Voi90] Michael Voit. Central limit theorems for a class of polynomial hypergroups. Advances in Applied Probability, 22(1):68– 87, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427597.

VanDijk:1988:PTM

[VP88] Nico M. Van Dijk and Martin L. Puterman. Perturbation theory for Markov reward processes with applications to queueing systems. Advances in Applied Probability, 20(1):79–98, March 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427271.

Valero:2010:ZTM

[VPCG10] J. Valero, M. Pérez-Casany, and J. Ginebra. On zero-truncating and mixing Poisson distributions. Advances in Applied Probability, 42(4):1013–1027, December 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/25799082.

Vervaat:1980:CFL

[VS80] Wim Vervaat and J. C. Smit. A conditioned functional limit theorem. Advances in Applied Probability, 12(2):296–297, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426572.

VanBeek:1987:VEM

[VS87] K. W. H. Van Beek and A. J. Stam. A variant of the Ehrenfest model. Advances in Applied Probability, 19(4):995–996, December 1987. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427112.

VanOyen:1994:OSS

[VT94] Mark P. Van Oyen and Demosthenis Teneketzis. Optimal stochastic scheduling of forest networks with switching penalties. Advances in Applied Probability, 26(2):474–497, June 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427447.

VanHee:1978:SCD

[VV78] K. M. Van Hee and J. Van Der Wal. Strongly convergent dynamic programming. Advances in Applied Probability, 10 (2):308–309, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426909.

VanLieshout:2001:ESC

[VV01] M. N. M. Van Lieshout and E. W. Van Zwet. Exact sampling from conditional Boolean models with applications to maximum likelihood inference. Advances in Applied Probability, 33(2):339– 353, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1428256. See correction [Ano03e].

Vasileiou:2006:ISM

[VV06] Aglaia Vasileiou and P.-C. G. Vassiliou. An inhomogeneous semi-Markov model for the term structure of credit risk spreads. Advances in Applied Probability, 38(1):171–198, March 2006. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/20443434.

VanHee:1977:SEF

[VW77] K. M. Van Hee and J. Wessels. Strongly excessive functions in Markov decision theory. Advances in Applied Probability, 9 (2):233–235, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426379.

VanCutsem:1994:RTB

[VY94] Bernard Van Cutsem and Bernard Ycart. Renewal-type behavior of absorption times in Markov chains. Advances in Applied Probability, 26(4):988–1005, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427901.

Wilson:1982:SGF

[WA82] Richard J. Wilson and Robert J. Adler. The structure of Gaussian fields near a level crossing. Advances in Applied Probability, 14(3):543–565, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426673.

Wachtel:2009:TPL

[Wac09] Vitali Wachtel. Transition phenomena for ladder epochs of random walks with small negative drift. Advances in Applied Probability, 41(4):1189–1214, December 2009. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/27793925.

Wade:2007:ELL

[Wad07] Andrew R. Wade. Explicit laws of large numbers for random nearest-neighbour-type graphs. Advances in Applied Probability, 39(2):326–342, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443516.

Wahba:1976:ASL

[Wah76] Grace Wahba. The approximate solution of linear operator equations when the data are noisy. Advances in Applied Probability, 8(2):222–223, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425871.

Waksman:1985:PPC

[Wak85] Peter Waksman. Plane polygons and a conjecture of Blaschke’s. Advances in Applied Probability, 17(4):774–793, December 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427087.

Walker:1973:EHC

[Wal73] A. M. Walker. On the estimation of a harmonic component in a time series with stationary dependent residuals. Advances in Applied Probability, 5(2):217–241, August 1973. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426034.

Walker:1974:PAL

[Wal74] A. M. Walker. A property of asymptotic least squares estimates for moving average processes. Advances in Applied Probability, 6(2):239–240, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426277.

Walsh:1981:SMN

[Wal81a] John B. Walsh. A stochastic model of neural response. Advances in Applied Probability, 13(2):231–281, June 1981. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426683.

Walsh:1981:WTD

[Wal81b] John B. Walsh. Well-timed diffusion approximations. Advances in Applied Probability, 13(2):352–368, June 1981. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426688.

Wang:1997:LTA

[Wan97] Chia-Li Wang. Light traffic approximations for regenerative queueing processes. Advances in Applied Probability, 29(4): 1060–1080, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427854.

Wang:2001:SCP

[Wan01] Hui Wang. Some control problems with random intervention times. Advances in Applied Probability, 33(2):404–422, June 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428260.

Wasserman:1978:MBD

[Was78] Stanley S. Wasserman. Models for binary directed graphs and their applications. Advances in Applied Probability, 10(4):803– 818, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426660.

Watterson:1974:STS

[Wat74] G. A. Watterson. The sampling theory of selectively neutral alleles. Advances in Applied Probability, 6(3):463–488, September 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426228.

Watterson:1978:SSG

[Wat78] G. A. Watterson. Simulating samples from a genetic population. Advances in Applied Probability, 10(4):720–723, December 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426651.

Watson:1979:SES

[Wat79a] R. K. Watson. Some extensions of the standard epidemic model. Advances in Applied Probability, 11(2):297, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426827.

Watterson:1979:ETS

[Wat79b] G. A. Watterson. Estimating and testing selection: The twoalleles, genic selection diffusion model. Advances in Applied Probability, 11(1):14–30, March 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426766.

Watterson:1982:MSL

[Wat82] G. A. Watterson. Mutant substitutions at linked nucleotide sites. Advances in Applied Probability, 14(2):206–224, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426517.

Watson:1986:OIE

[Wat86a] G. S. Watson. An ordering inequality for exchangeable random variables. Advances in Applied Probability, 18(1):274–276, March 1986. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427246.

Watson:1986:SRS

[Wat86b] G. S. Watson. The shapes of a random sequence of triangles. Advances in Applied Probability, 18(1):156–169, March 1986. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427241.

Watterson:1992:MNA

[Wat92] G. A. Watterson. The mean number of alleles in multigene families. Advances in Applied Probability, 24(1):1–19, March 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427727.

Waugh:1974:ADS

[Wau74] W. A. O’N. Waugh. Age distribution in some size-dependent populations. Advances in Applied Probability, 6(2):232, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426270.

Waugh:1981:AGW

[Wau81] W. A. O’N. Waugh. Application of the Galton–Watson process to the kin number problem. Advances in Applied Probability, 13 (4):631–649, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426964.

Waymire:1996:MRC

[Way96] Edward C. Waymire. Multiscaling and random cascades. Advances in Applied Probability, 28(2):343–344, June 1996. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1428058.

Williams:1971:HSA

[WB71] Trevor Williams and Rolf Bjerknes. Hyperplasia: The spread of abnormal cells through a plane lattice. Advances in Applied Probability, 3(2):210–211, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426160.

Worz-Busekros:1981:RGA

[WB81] Angelika Wörz-Busekros. Remarks on genetic algebras. Advances in Applied Probability, 13(1):13, March 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426464.

Wang:1990:NKM

[WC90] Yuedong Wang and Jinhua Cao. New kinds of multidimensional IFR distribution. Advances in Applied Probability, 22(1):251– 253, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427609.

Weiss:2011:TRS

[WC11] Viola Weiss and Richard Cowan. Topological relationships in spatial tessellations. Advances in Applied Probability, 43(4): 963–984, December 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/23208034.

Wang:1996:ENS

[WCK96] I-Jeng Wang, Edwin K. P. Chong, and Sanjeev R. Kulkarni. Equivalent necessary and sufficient conditions on noise sequences for stochastic approximation algorithms. Advances in Applied Probability, 28(3):784–801, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428181.

Willmot:2001:LIR

[WCL01] Gordon E. Willmot, Jun Cai, and X. Sheldon Lin. Lundberg inequalities for renewal equations. Advances in Applied Probability, 33(3):674–689, September 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428481.

Weiss:1971:ABG

[WD71] George H. Weiss and Menachem Dishon. Asymptotic behavior of a generalization of Bailey’s simple epidemic. Advances in Applied Probability, 3(2):220–221, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426164.

Wortman:1990:VQM

[WD90] M. A. Wortman and Ralph L. Disney. Vacation queues with Markov schedules. Advances in Applied Probability, 22(3):730– 748, September 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427466.

Wilson:1998:MAR

[WDH98] Helen Wilson, Peter Diggle, and C. Vyvyan Howard. Methods for the analysis of replicated spatial point patterns in clinical neuro-anatomy. Advances in Applied Probability, 30(2):293– 294, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427968.

Weber:1992:ITQ

[Web92] Richard R. Weber. The interchangeability of tandem queues with heterogeneous customers and dependent service times. Advances in Applied Probability, 24(3):727–737, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427486.

Weerasinghe:2002:SSC

[Wee02] Ananda P. N. Weerasinghe. Stationary stochastic control for Itô processes. Advances in Applied Probability, 34(1):128–140, March 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428375.

Weiss:1971:PMT

[Wei71] George H. Weiss. On a perturbation method for the theory of epidemics. Advances in Applied Probability, 3(2):218–220, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426163.

Weiss:1976:SNR

[Wei76] George H. Weiss. The shot noise representation for noise for a free traffic stream. Advances in Applied Probability, 8(2):224– 225, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425872.

Weiss:1977:WRS

[Wei77] Benjamin Weiss. What is a random sequence? Advances in Applied Probability, 9(2):224, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426368.

Weissman:1978:EPL

[Wei78] Ishay Weissman. Estimation of parameters and large quantiles based on the K largest observations. Advances in Applied Probability, 10(2):331–332, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426930.

Weil:1982:ICP

[Wei82] Wolfgang Weil. Inner contact probabilities for convex bodies. Advances in Applied Probability, 14(3):582–599, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426675.

Weiss:1986:NTA

[Wei86] Alan Weiss. A new technique for analyzing large traffic systems. Advances in Applied Probability, 18(2):506–532, June 1986. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427310.

Weil:1995:EMS

[Wei95a] Wolfgang Weil. The estimation of mean shape and mean particle number in overlapping particle systems in the plane. Advances in Applied Probability, 27(1):102–119, March 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428099.

Weiss:1995:AOP

[Wei95b] Gideon Weiss. On almost optimal priority rules for preemptive scheduling of stochastic jobs on parallel machines. Advances in Applied Probability, 27(3):821–839, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428135.

Weil:1997:MSP

[Wei97] Wolfgang Weil. On the mean shape of particle processes. Advances in Applied Probability, 29(4):890–908, December 1997. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427846.

Weil:1998:IFS

[Wei98] Wolfgang Weil. Integral formulas for support functions and their use in the analysis of Boolean models. Advances in Applied Probability, 30(2):292–293, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427967.

Weil:2001:DMV

[Wei01] Wolfgang Weil. Densities of mixed volumes for Boolean models. Advances in Applied Probability, 33(1):39–60, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428440.

Wenocur:1989:RMB

[Wen89] Michael L. Wenocur. A reliability model based on the gamma process and its analytic theory. Advances in Applied Probability, 21(4):899–918, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1427773.

Westcott:1974:ARP

[Wes74] M. Westcott. Asymptotic results for Poisson cluster processes. Advances in Applied Probability, 6(2):227–228, June 1974. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426265.

Wessels:1976:MPS

[Wes76] J. Wessels. Markov programming by successive approximations with respect to weighted supremum norms. Advances in Applied Probability, 8(2):248–249, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1425898.

Westcott:1978:RRR

[Wes78] Mark Westcott. Records and random record processes. Advances in Applied Probability, 10(2):278–280, June 1978. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426883.

Westcott:1979:ZIM

[Wes79] M. Westcott. Zone of influence models for competition in plantations. Advances in Applied Probability, 11(2):307, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426839.

Wesolowski:1990:CRC

[Wes90] Jacek Wesolowski. A constant regression characterization of the gamma law. Advances in Applied Probability, 22(2):488– 490, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427549.

Waymire:1983:APP

[WG83] Ed Waymire and Vijay K. Gupta. An analysis of the Pólya point process. Advances in Applied Probability, 15(1):39–53, March 1983. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426981.

Wang:1991:NGI

[WG91] You-Gan Wang and John Gittins. A note on Gittins indices for pharmaceutical research. Advances in Applied Probability, 23(4): 975–977, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427688. See [Git69].

Whitt:1980:NHT

[WH80] Ward Whitt and Shlomo Halfin. A new heavy-traffic limit theorem for many-server queues. Advances in Applied Probability, 12 (2):295, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426569.

Wilkinson-Herbots:2003:CTV

[WH03] Hilde M. Wilkinson-Herbots. Coalescence times and F_{ST} values in subdivided populations with symmetric structure. Advances in Applied Probability, 35(3):665–690, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428213.

Whittle:1969:SDP

[Whi69] P. Whittle. Sequential decision processes with essential unobservables. Advances in Applied Probability, 1(2):271–287, Autumn 1969. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426220.

Whitt:1970:MCQ

[Whi70] Ward Whitt. Multiple channel queues in heavy traffic. III: Random server selection. Advances in Applied Probability, 2(2):370– 375, Autumn 1970. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426325.

Whitt:1974:CQ

[Whi74] Ward Whitt. The continuity of queues. Advances in Applied Probability, 6(1):175–183, March 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1426213.

White:1980:DSC

[Whi80a] John S. White. The distribution of the serial correlation coefficient for a stationary series. Advances in Applied Probability, 12(2):316–317, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426597.

Whittle:1980:PPIa

[Whi80b] P. Whittle. Polymerisation processes with intrapolymer bonding. I. One type of unit. Advances in Applied Probability, 12 (1):94–115, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426496.

Whittle:1980:PPIb

[Whi80c] P. Whittle. Polymerisation processes with intrapolymer bonding. II. Stratified processes. Advances in Applied Probability, 12 (1):116–134, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426497.

Whittle:1980:PPIc

[Whi80d] P. Whittle. Polymerisation processes with intrapolymer bonding. III. Several types of unit. Advances in Applied Probability, 12(1):135–153, March 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426498.

Whitt:1981:CCP

[Whi81a] Ward Whitt. Comparing counting processes and queues. Advances in Applied Probability, 13(1):207–220, March 1981. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426475.

Whittle:1981:RSL

[Whi81b] P. Whittle. Risk-sensitive Linear/Quadratic/Gaussian control. Advances in Applied Probability, 13(4):764–777, December 1981. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426972.

Whitt:1982:HTL

[Whi82a] Ward Whitt. On the heavy-traffic limit theorem for GI/G/∞ queues. Advances in Applied Probability, 14(1):171–190, March 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426738.

Whittle:1982:ESC

[Whi82b] P. Whittle. The emergence of structure, and critical effects. Advances in Applied Probability, 14(2):205, June 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426516.

Whittle:1983:RMP

[Whi83] P. Whittle. Relaxed Markov processes. Advances in Applied Probability, 15(4):769–782, December 1983. CODEN AAPBBD.

ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427323.

Whittle:1986:PBI

[Whi86] P. Whittle. Partial balance, insensitivity and weak coupling. Advances in Applied Probability, 18(3):706–723, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427184.

Whittle:1991:SMA

[Whi91] P. Whittle. A stochastic model of an artificial neuron. Advances in Applied Probability, 23(4):809–822, December 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427677.

Whitt:1992:PT

[Whi92a] Ward Whitt. H = λG and the Palm transformation. Advances in Applied Probability, 24(3):755–758, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427489.

Whittle:1992:RFR

[Whi92b] P. Whittle. Random fields on random graphs. Advances in Applied Probability, 24(2):455–473, June 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427700.

Whittle:1995:LDE

[Whi95] P. Whittle. Large-deviation expressions for the distribution of first-passage coordinates. Advances in Applied Probability, 27 (3):692–710, September 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1428130.

Whittaker:1996:ARM

[Whi96] J. C. Whittaker. The allocation of resources in a multiple-trial war of attrition conflict. Advances in Applied Probability, 28(3): 933–964, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428187.

Wang:2017:CDS

[WHL17] Zimeng Wang, David J. Hodge, and Huiling Le. Conjugate duality in stochastic controls with delay. Advances in Applied Probability, 49(4):1011–1036, December 2017. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/conjugate-duality-in-stochastic-
controls-with-delay/2FB2F1A9A2D1371B3C0303946D129CE5.

Wu:2010:CEL

[WHZ10] Wei Biao Wu, Yinxiao Huang, and Wei Zheng. Covariances estimation for long-memory processes. Advances in Applied Probability, 42(1):137–157, March 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/25683810.

Wickwire:1977:OIR

[Wic77] K. Wickwire. Optimal immunisation rules for an epidemic with recovery. Advances in Applied Probability, 9(2):222–223, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426366.

Wierman:1977:FPP

[Wie77] John C. Wierman. First-passage percolation on the square lattice, II. Advances in Applied Probability, 9(2):283–295, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426387.

Wierman:1981:BPH

[Wie81] John C. Wierman. Bond percolation on honeycomb and triangular lattices. Advances in Applied Probability, 13(2):298– 313, June 1981. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426685.

Wijngaard:1978:QCE

[Wij78] J. Wijngaard. Quasi-compactness of embedded Markov processes and entry processes. Advances in Applied Probability, 10 (2):300–301, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426902.

Williams:1971:APT

[Wil71] Trevor Williams. An algebraic proof of the threshold theorem for the general stochastic epidemic. Advances in Applied Probability, 3(2):223, Autumn 1971. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426166.

Williams:1974:ESC

[Wil74] Trevor Williams. Evidence for super-critical tumour growth. Advances in Applied Probability, 6(2):237–238, June 1974. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1426275.

Wilson:1979:AMF

[Wil79] S. R. Wilson. Associative mating for familial disorders: Some anomalies. Advances in Applied Probability, 11(2):294–295, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426825.

Willenborg:1985:TSM

[Wil85] Leon Willenborg. The thrown string: A Markov field approach. Advances in Applied Probability, 17(3):607–622, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427122.

Wilson:1988:MFC

[Wil88] Richard J. Wilson. Model fields in crossing theory: A weak convergence perspective. Advances in Applied Probability, 20(4): 756–774, December 1988. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427359.

Willmot:1990:ATB

[Wil90] Gordon E. Willmot. Asymptotic tail behaviour of Poisson mixtures with applications. Advances in Applied Probability, 22 (1):147–159, March 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427602.

Willie:1998:PSS

[Wil98] Helmut Willie. Periodic steady state of loss systems with periodic inputs. Advances in Applied Probability, 30(1):152–166, March 1998. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427881.

Wisniewski:1972:BSP

[Wis72] T. K. M. Wisniewski. Bivariate stationary point processes, fundamental relations and first recurrence times. Advances in Applied Probability, 4(2):296–317, August 1972. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426001.

Wise:1980:NBC

[Wis80] Gary L. Wise. The non-bandlimitedness of a class of random processes. Advances in Applied Probability, 12(2):317, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426598.

Withers:1979:EVS

[Wit79] C. S. Withers. Estimating the variance of the sample mean for time series with applications to robust tests. Advances in Applied Probability, 11(2):302–303, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426833.

Weiss:1994:SOS

[WN94] V. Weiss and W. Nagel. Second-order stereology for planar fibre processes. Advances in Applied Probability, 26(4):906–918, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427897.

Weiss:1999:IDQ

[WN99] Viola Weiss and Werner Nagel. Interdependences of directional quantities of planar tessellations. Advances in Applied Probability, 31(3):664–678, September 1999. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1428390.

Withers:2012:DFL

[WN12] Christopher S. Withers and Saralees Nadarajah. The distribution of Foschini’s lower bound for channel capacity. Advances in Applied Probability, 44(1):260–269, March 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/23207863.

Woess:1985:RWP

[Woe85] Wolfgang Woess. Random walks and periodic continued fractions. Advances in Applied Probability, 17(1):67–84, March 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427053.

Woerner:2007:ILT

[Woe07] Jeannette H. C. Woerner. Inference in Lévy-type stochastic volatility models. Advances in Applied Probability, 39(2):531– 549, June 2007. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
20443525.

Wolfe:1978:UID

[Wol78] Stephen James Wolfe. On the unimodality of infinitely divisible distribution functions. Advances in Applied Probability, 10 (2):285–286, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426888.

Wolf:1980:AIP

[Wol80a] D. Wolf. Approximation of the invariant probability measure of an infinite stochastic matrix. Advances in Applied Probability, 12(3):710–726, September 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.

org/stable/1426428.

Wolfe:1980:CAS

[Wol80b] Stephen James Wolfe. On a continuous analogue of the stochastic difference equation X_{n} = ρX_{n−1} + B_{n}. Advances in Applied Probability, 12(2):286, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1426557.

Wong:2008:GPA

[Won08] Shek-Keung Tony Wong. The generalized perpetual American exchange-option problem. Advances in Applied Probability, 40 (1):163–182, March 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/20443575.

Woods:1974:EMN

[Woo74] A. J. Woods. Epidemic models in non-homogeneous populations. Advances in Applied Probability, 6(2):239, June 1974. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426276.

Woolford:1980:LTA

[Woo80] S. W. Woolford. On limit theorems associated with a storage model. Advances in Applied Probability, 12(2):303–304, June 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1426579.

Worsley:1994:LME

[Wor94] K. J. Worsley. Local maxima and the expected Euler characteristic of excursion sets of χ^{2}, F and t fields. Advances in Applied Probability, 26(1):13–42, March 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427576.

Worsley:1995:BCE

[Wor95] K. J. Worsley. Boundary corrections for the expected Euler characteristic of excursion sets of random fields, with an application to astrophysics. Advances in Applied Probability, 27(4): 943–959, December 1995. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427930.

Worsley:2001:TSU

[Wor01] Keith J. Worsley. Testing for signals with unknown location and scale in a χ^{2} random field, with an application to fMRI. Advances in Applied Probability, 33(4):773–793, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428305.

Weiss:1976:OSR

[WR76a] George H. Weiss and Robert J. Rubin. The ordered spans of a random walk. Advances in Applied Probability, 8(2):236– 237, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1425883.

Whittle:1976:SCC

[WR76b] P. Whittle and J. F. Rudge. Simultaneous communication and control. Advances in Applied Probability, 8(2):365–384, June 1976. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1425909.

Wang:2022:MRD

[WR22] Tiandong Wang and Sidney I. Resnick. Measuring reciprocity in a directed preferential attachment network. Advances in Applied Probability, 54(3):718–742, September 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/measuring-reciprocity-in-
a-directed-preferential-attachment-network/025F69A9B9F1C681026539E4349C9635.

Wright:1992:TPP

[Wri92] Paul E. Wright. Two parallel processors with coupled inputs. Advances in Applied Probability, 24(4):986–1007, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427722.

Weber:1987:OCS

[WS87] Richard R. Weber and Shaler Stidham, Jr. Optimal control of service rates in networks of queues. Advances in Applied Probability, 19(1):202–218, March 1987. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427380.

Wang:2010:SRM

[WS10] Yizao Wang and Stilian A. Stoev. On the structure and representations of max-stable processes. Advances in Applied Probability, 42(3):855–877, September 2010. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/25799071.

Wang:2011:CSS

[WS11] Yizao Wang and Stilian A. Stoev. Conditional sampling for spectrally discrete max-stable random fields. Advances in Applied Probability, 43(2):461–483, June 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/23024515.

Wu:1993:MBD

[Wu93] Yadong Wu. A multilevel birth–death particle system and its continuous diffusion. Advances in Applied Probability, 25(3): 549–569, September 1993. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427523.

Walrand:1980:STO

[WV80] J. Walrand and P. Varaiya. Sojourn times and the overtaking condition in jacksonian networks. Advances in Applied Probability, 12(4):1000–1018, December 1980. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426753.

Weil:1984:DSR

[WW84] Wolfgang Weil and John A. Wieacker. Densities for stationary random sets and point processes. Advances in Applied Probability, 16(2):324–346, June 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.

jstor.org/stable/1427072.

Weber:1991:AIP

[WW91] Richard R. Weber and Gideon Weiss. Addendum to ’on an index policy for restless bandits’. Advances in Applied Probability, 23 (2):429–430, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427757.

Wolff:2003:IPA

[WW03] Ronald W. Wolff and Chia-Li Wang. Idle period approximations and bounds for the GI/G/1 queue. Advances in Applied Probability, 35(3):773–792, September 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http:/

/www.jstor.org/stable/1428218.

Wang:2021:KJN

[WW21] Chia-Li Wang and Ronald W. Wolff. Kelly and Jackson networks with interchangeable, cooperative servers. Advances in Applied Probability, 53(2):463–483, June 2021. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/kelly-and-jackson-networks-
with-interchangeable-cooperative-servers/CB1EA37C244DDAA948883C3DC67DAC8B.

Wei:2012:ODS

[WWY12] Jiaqin Wei, Rongming Wang, and Hailiang Yang. On the optimal dividend strategy in a regime-switching diffusion model. Advances in Applied Probability, 44(3):886–906, September 2012. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/41714080.

Wang:2019:OLC

[WZ19] Wenyuan Wang and Zhimin Zhang. Optimal loss-carry-forward taxation for Lévy risk processes stopped at general draw-down time. Advances in Applied Probability, 51 (3):865–897, September 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/optimal-losscarryforward-taxation-
for-levy-risk-processes-stopped-at-general-drawdown-
time/941185A6C321DC978BF0D2AF30CC2EBE.

Wang:2020:DPR

[WZ20] Wenyuan Wang and Xiaowen Zhou. Draw-down Parisian ruin for spectrally negative Lévy processes. Advances in Applied Probability, 52(4):1164–1196, December 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/drawdown-parisian-ruin-for-
spectrally-negative-levy-processes/05A5764841D1511DE7419AE8B2B15343.

Xu:2018:NSC

[XCWC18] Hui Xu, Fengyang Cheng, Yuebao Wang, and Dongya Cheng. A necessary and sufficient condition for the subexponentiality of the product convolution. Advances in Applied Probability, 50(1):57–73, March 2018. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/necessary-and-sufficient-condition-
for-the-subexponentiality-of-the-product-convolution/
DE667E3A9D50D51DC156152BDF3ED283.

Xu:2023:FCD

[XHW23] Chuang Xu, Mads Christian Hansen, and Carsten Wiuf. Full classification of dynamics for one-dimensional continuous-time Markov chains with polynomial transition rates. Advances in Applied Probability, 55(1):321–355, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/full-classification-of-dynamics-
for-onedimensional-continuoustime-markov-chains-with-
polynomial-transition-rates/E94F590AA003981FDED8AA5CD22CDD1A.

Xu:1992:JRI

[Xu92] Susan H. Xu. On a job resequencing issue in parallel processor stochastic scheduling. Advances in Applied Probability, 24(4): 915–933, December 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427719.

Xu:2011:SFL

[Xu11] Fang Xu. The sampling formula and Laplace transform associated with the two-parameter Poisson–Dirichlet distribution. Advances in Applied Probability, 43(4):1066–1085, December 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/23208039.

Xia:2015:NAS

[XY15] Aihua Xia and J. E. Yukich. Normal approximation for statistics of Gibbsian input in geometric probability. Advances in Applied Probability, 47(4):934–972, December 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1449859795.

Xu:1996:DRJ

[XZ96] Susan H. Xu and Y. Quennel Zhao. Dynamic routing and jockeying controls in a two-station queueing system. Advances in Applied Probability, 28(4):1201–1226, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428170.

Xu:2020:IQS

[XZ20] Jiaming Xu and Yuan Zhong. Improved queue-size scaling for input-queued switches via graph factorization. Advances in Applied Probability, 52(3):798–824, September 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/improved-queuesize-scaling-
for-inputqueued-switches-via-graph-factorization/1E2856862C8B0AFDD7B5428FF6D184A3.

Yao:2000:SNA

[YA00] J.-F. Yao and J.-G. Attali. On stability of nonlinear AR processes with Markov switching. Advances in Applied Probability, 32(2):394–407, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428195.

Yao:2010:VPV

[Yao10] Yi-Ching Yao. On variances of partial volumes of the typical cell of a Poisson–Voronoi tessellation and large-dimensional volume degeneracy. Advances in Applied Probability, 42(2):359– 370, June 2010. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
25683824.

Yashchin:1985:UAA

[Yas85] Emmanuel Yashchin. On a unified approach to the analysis of two-sided cumulative sum control schemes with headstarts. Advances in Applied Probability, 17(3):562–593, September 1985. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1427120.

Yi:2020:MSS

[YBC20] He Yi, Narayanaswamy Balakrishnan, and Lirong Cui. On the multi-state signatures of ordered system lifetimes. Advances in Applied Probability, 52(1):291–318, March 2020. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/on-the-multistate-signatures-
of-ordered-system-lifetimes/ABC4F45331ED61EF7B586C131ED948F7.

Yang:1996:SSQ

[YC96] Tao Yang and M. L. Chaudhry. On steady-state queue size distributions of the discrete-time GI/G/1 queue. Advances in Applied Probability, 28(4):1177–1200, December 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428169.

Ycart:1984:LRA

[Yca84] Bernard Ycart. L’application des representations au calcul explicite sur certaines chaînes de Markov finies. (French) [The application of the representations to explicit calculation on some finite Markov chains]. Advances in Applied Probability, 16(3): 656–666, September 1984. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427292.

Yeh:1988:NOR

[Yeh88] Lam Yeh. A note on the optimal replacement problem. Advances in Applied Probability, 20(2):479–482, June 1988. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://www.jstor.org/stable/1427402.

Yeh:1990:RRM

[Yeh90] Lam Yeh. A repair replacement model. Advances in Applied Probability, 22(2):494–497, June 1990. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427551.

Yeo:1978:SAR

[Yeo78] G. F. Yeo. On a superposition of alternating renewal processes in traffic. Advances in Applied Probability, 10(2):304, June 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1426905.

Yin:2000:CGS

[Yin00] G. Yin. Convergence of a global stochastic optimization algorithm with partial step size restarting. Advances in Applied Probability, 32(2):480–498, June 2000. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428200.

Yin:2023:PFI

[Yin23] Mei Yin. Parking functions: interdisciplinary connections. Advances in Applied Probability, 55(3):768–792, September 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL https://www.cambridge.org/core/journals/
advances-in-applied-probability/article/parking-functions-
interdisciplinary-connections/5AAE9D7437260405B88DFA01B518E962.

Yamashita:1998:GPF

[YM98] Hideaki Yamashita and Masakiyo Miyazawa. Geometric product form queueing networks with concurrent batch movements. Advances in Applied Probability, 30(4):1111–1129, December 1998.

CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (elec-

tronic). URL http://www.jstor.org/stable/1428032.

Yor:1992:SEF

[Yor92] Marc Yor. On some exponential functionals of Brownian motion. Advances in Applied Probability, 24(3):509–531, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1427477.

Yadin:1979:RIM

[YS79] M. Yadin and R. Syski. Randomization of intensities in a Markov chain. Advances in Applied Probability, 11(2):397– 421, June 1979. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1426846.

Yeung:1994:MPF

[YS94] Raymond W. Yeung and Bhaskar Sengupta. Matrix productform solutions for Markov chains with a tree structure. Advances in Applied Probability, 26(4):965–987, December 1994. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427900.

Yasodharan:2023:LTB

[YS23] Sarath Yasodharan and Rajesh Sundaresan. Large-time behaviour and the second eigenvalue problem for finitestate mean-field interacting particle systems. Advances in Applied Probability, 55(1):85–125, March 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

https://www.cambridge.org/core/journals/advances-in-
applied-probability/article/largetime-behaviour-and-
the-second-eigenvalue-problem-for-finitestate-meanfield-
interacting-particle-systems/0130B85AD92C6C2DB51F9C1FED5367B7.

Yu:2008:IKR

[Yu08] Yaming Yu. On an inequality of Karlin and Rinott concerning weighted sums of I.I.D. random variables. Advances in Applied Probability, 40(4):1223–1226, December 2008. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/20443631.

Yuan:2022:KMR

[Yua22] Linglong Yuan. Kingman’s model with random mutation probabilities: convergence and condensation I. Advances in Applied Probability, 54(1):311–335, March 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/kingmans-model-with-random-mutation-
probabilities-convergence-and-condensation-i/4627B811916730C84FEE600F9A93F2C9.

Yuh:1982:SOC

[Yuh82] Yih-Shyh Yuh. Second-order corrections for Brownian motion approximations to first-passage probabilities. Advances in Applied Probability, 14(3):566–581, September 1982. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1426674.

Yunus:1986:SDA

[Yun86] M. Naim Yunus. A simple derivation of an approximation to the transient blocking probability in a loss system. Advances in Applied Probability, 18(3):862–863, September 1986. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427192.

Yang:2022:OEC

[YY22] Yue Yang and Xiang Yu. Optimal entry and consumption under habit formation. Advances in Applied Probability, 54(2):433–459, June 2022. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/optimal-entry-and-consumption-under-
habit-formation/ABE4A4B1C317529013A6DDC8D10A8478.

Yadin:1978:ASC

[YZ78] M. Yadin and S. Zacks. Adaptation of the service capacity in a queueing system which is subjected to a change in the arrival rate at unknown epoch. Advances in Applied Probability, 10(3): 666–681, September 1978. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1426640.

Yang:2001:SNL

[YZ01] Hailiang Yang and Lianzeng Zhang. Spectrally negative Lévy processes with applications in risk theory. Advances in Applied Probability, 33(1):281–291, March 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428452.

Yin:2002:CSS

[YZ02] G. Yin and Hanqin Zhang. Countable-state-space Markov chains with two time scales and applications to queueing systems. Advances in Applied Probability, 34(3):662–688, September 2002. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/

1428251. See correction [Ano04f].

Yin:2003:DTS

[YZB03] G. Yin, Q. Zhang, and G. Badowski. Discrete-time singularly perturbed Markov chains: Aggregation, occupation measures, and switching diffusion limit. Advances in Applied Probability, 35(2):449–476, June 2003. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428431.

Zachary:1991:BLN

[Zac91] S. Zachary. On blocking in loss networks. Advances in Applied Probability, 23(2):355–372, June 1991. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1427753.

Zahle:1996:DMU

[Zäh96] Martina Zähle. Dimensions of measures under orthogonal projections. Advances in Applied Probability, 28(2):344–345, June 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064

(electronic). URL http://www.jstor.org/stable/1428059.

Zakai:1977:SIP

[Zak77] Moshe Zakai. Stochastic integrals in the plane. Advances in Applied Probability, 9(2):210–212, June 1977. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1426355.

Zazanis:1992:APD

[Zaz92] Michael A. Zazanis. Analyticity of Poisson-driven stochastic systems. Advances in Applied Probability, 24(3):532–541, September 1992. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427478.

Zirbel:1996:DPS

[ZÇ96] Craig L. Zirbel and Erhan Çinlar. Dispersion of particle systems in Brownian flows. Advances in Applied Probability, 28(1):53– 74, March 1996. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427913.

Zayas-Caban:2019:AOH

[ZCJW19] Gabriel Zayas-Cabán, Stefanus Jasin, and Guihua Wang. An asymptotically optimal heuristic for general nonstationary finite-horizon restless multi-armed, multi-action bandits. Advances in Applied Probability, 51(3):745–772, September 2019. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL https://www.cambridge.org/core/
journals/advances-in-applied-probability/article/an-
asymptotically-optimal-heuristic-for-general-nonstationary-
finitehorizon-restless-multiarmed-multiaction-bandits/
CD95BDA1544D52BB3CA91ADA1804640B. See corrigendum [ZCLJW23].

Zayas-Caban:2023:AOH

[ZCLJW23] Gabriel Zayas-Cabán, Jiaxin Liang, Stefanus Jasin, and Guihua Wang. An asymptotically optimal heuristic for general nonstationary finite-horizon restless multi-armed, multi-action bandits: Corrigendum. Advances in Applied Probability, 55 (3):1075–1083, September 2023. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL https://

www.cambridge.org/core/journals/advances-in-applied-
probability/article/an-asymptotically-optimal-heuristic-
for-general-nonstationary-finitehorizon-restless-multiarmed-
multiaction-bandits-corrigendum/FEE34691647983AFD164EF5705C2A48E.

See [ZCJW19].

Zhu:1996:STC

[ZCWS96] Li-Xing Zhu, Ping Cheng, Gang Wei, and Pei-De Shi. A stopping time concerning sphere data and its applications. Advances in Applied Probability, 28(3):674–686, September 1996. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1428176.

Zhao-Guo:1988:CEH

[ZG88] Chen Zhao-Guo. Consistent estimates for hidden frequencies in a linear process. Advances in Applied Probability, 20(2):295– 314, June 1988. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1427391.

Zeevi:2004:ETD

[ZG04] Assaf Zeevi and Peter W. Glynn. Estimating tail decay for stationary sequences via extreme values. Advances in Applied Probability, 36(1):198–226, March 2004. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/1428359.

Zhang:1997:SAI

[Zha97] Hanqin Zhang. Strong approximations of irreducible closed queueing networks. Advances in Applied Probability, 29(2):498– 522, June 1997. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428014.

Zhang:2015:OTD

[Zha15] Hongzhong Zhang. Occupation times, drawdowns, and drawups for one-dimensional regular diffusions. Advances in Applied Probability, 47(1):210–230, March 2015. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1427814588.

Zhang:2014:APM

[ZHCC14] Li-Xin Zhang, Feifang Hu, Siu Hung Cheung, and Wai Sum Chan. Asymptotic properties of multicolor randomly reinforced Pólya urns. Advances in Applied Probability, 46(2):585–602, June 2014. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-

6064 (electronic). URL http://projecteuclid.org/euclid.
aap/1401369708.

Zirbel:2001:LOH

[Zir01] Craig L. Zirbel. Lagrangian observations of homogeneous random environments. Advances in Applied Probability, 33(4):810– 835, December 2001. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1428307.

Ziedins:1989:LTL

[ZK89] I. B. Ziedins and F. P. Kelly. Limit theorems for loss networks with diverse routing. Advances in Applied Probability, 21(4): 804–830, December 1989. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://www.jstor.org/

stable/1427768.

Zhou:2009:CMC

[ZL09] Hua Zhou and Kenneth Lange. Composition Markov chains of multinomial type. Advances in Applied Probability, 41(1):270– 291, March 2009. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
27793873.

Zhao:1998:IBS

[ZLB98] Yiqiang Q. Zhao, Wei Li, and W. John Braun. Infinite blockstructured transition matrices and their properties. Advances in Applied Probability, 30(2):365–384, June 1998. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL

http://www.jstor.org/stable/1427973.

Zhang:2016:DPW

[ZM16] Panpan Zhang and Hosam Mahmoud. The degree profile and weight in Apollonian networks and k-trees. Advances in Applied Probability, 48(1):163–175, March 2016. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

projecteuclid.org/euclid.aap/1457466160.

Zuyev:1999:SSG

[Zuy99] Sergei Zuyev. Stopping sets: Gamma-type results and hitting properties. Advances in Applied Probability, 31(2):355– 366, June 1999. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428114.

Zwart:2000:FQF

[Zwa00] A. P. Zwart. A fluid queue with a finite buffer and subexponential input. Advances in Applied Probability, 32(1):221– 243, March 2000. CODEN AAPBBD. ISSN 0001-8678 (print),

1475-6064 (electronic). URL http://www.jstor.org/stable/
1428235.

Zhu:2016:OFD

[ZY16] Jinxia Zhu and Hailiang Yang. Optimal financing and dividend distribution in a general diffusion model with regime switching. Advances in Applied Probability, 48(2):406–422, June 2016. CO-DEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic).

URL http://projecteuclid.org/euclid.aap/1465490755.

Zhang:2011:ARM

[ZYY11] Zhimin Zhang, Hailiang Yang, and Hu Yang. On the absolute ruin in a map risk model with debit interest. Advances in Applied Probability, 43(1):77–96, March 2011. CODEN AAPBBD. ISSN 0001-8678 (print), 1475-6064 (electronic). URL http://

www.jstor.org/stable/23024526.